A Complete Bibliography of Publications in the Journal of Cell Biology: 1980–1984

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

04 June 2019 Version 1.00

Title word cross-reference

1 [KSD^{+}83, Sch81b]. 10 [BMF84, PWSW80, TIK81]. 110 [HM83a]. 12 [WCZG84]. 13 [WCZG84]. 2 [DWP81, FFEL81, HA80, MSB^{+}83, SP84a]. 21 [ME82]. 32 [OKA84]. 34 [GBE84]. 5 [SC84b, WA83]. 56 [WES84]. 58 [NS83a]. 6 [BCF84, GGK82, GK84, GSS^{+}84, GGK83]. 8 [JZ81]. 95 [BM83]. ^{+} [PF83b, ZKF84]. ^{2+} [BPM81, BYS82, CH80a, FJ82c, GWS84, JSDM82, JSMT82, KKM84, MSP84, MA82, NJR80, PG84, PGD^{+}84, POS80, SW84, Sch84, SM83c, Wal82a, Wal82b, YAF81]. α [Bad81, FB80, MLSR83, RH81]. β [KFV83].

-acetate [WCZG84]. -actinin [FB80]. -activated [PG84]. -adrenergic [MSB^{+}83]. -ATPase [ES81a, JSDM82, MS84e, MM81c, POS80, PF83b, ZKF84]. -azacytidine [SC84b]. -bromodeoxyuridine-blocked [WA83]. -bungarotoxin-binding [Bad81]. -calmodulin [KDMS83, MSP84]. -calmodulin-dependent [KM82]. -cell [JZ81]. -containing [Rag81, WH80]. -deficient [GGK83].

-dependent [BYS82, CH80a, GWS84, JSMT82, MB82b, PGD^{+}84, YAF81]. -from [FP81a]. -independent [CF80, PGD^{+}84]. -induced [PS82]. -kdalton [BM83, GBE84, HM83a, NS83a, WES84]. -kilodalton [OKA84]. -lactamase [KFV83]. -like [KSD^{+}83]. -macroglobulin [DWP81]. -mannosidase [RH81]. -mediated [BPM81]. -methyladenine [Sch81b]. -microglobulin [FFEL81, SP84a]. -myristate [WCZG84]. -nm [BMF84, ME82, PWSW80, TIK81]. -phosphate [BCF84, GGK82, GK84, GSS^{+}84, GGK83]. -ribonucleoprotein [BK83a]. -sensitive [MA82]. -sequestering [Wal82a, Wal82b]. -tubulin [MLSR83]. -ylcarbamate [HA80].

/calmodulin [Sch84]. /calmodulin-dependent [Sch84]. /H [BC84b, WCZG84].

000 [FYC82]. 000- [CVC^{+}84]. 000-dalton [Bre83, BFST83, BYS82, CP82b, CVC^{+}84, FJ82b, FJ82c, GCH84, GKL^{+}84, GLP80, JOA84, LYG81, MWH83, NCH83, OT82, PDK83, RCM^{+}83, SSSJ83]. 000-mol [DBG^{+}81]. 000-mol-wt [CG84, HM84c, MBLvdM84, PV84, RBH^{+}84, Vah83, WS84]. 000-Mr [FYC82].

1 [KB81c, MWH83, Vah83]. 1- [MSB^{+}83]. 100 [IMS82, TFGC84, WSB^{+}80]. 100-A [WSB^{+}80]. 105 [OFO83]. 125I [CB82, BLG^{+}80, CVGO81, CGA^{+}82, LSM83, Ott83, SBB81]. 125I-actin [SBB81]. 125I-epidermal [CGA^{+}82]. 125I-human [BLG^{+}80]. 125I-insulin [CVGO81]. 125I-thrombin [CB82]. 125I-vinculin [Ott83]. 125I-wheat [LSM83]. 14.5-d [CCKMH82]. 1A [BLV84].

2 [APZL80, BV83, ED84, EW82, OWMB82, WSM^{+}82]. 2-Br-alpha-ergocryptine [APZL80]. 2-mannosidase [HTGB84]. 20 [MGK83]. 21 [HKSF80]. 240 [GGW83]. 25-Dihydroxyvitamin [MKD^{+}84]. 260/240 [PFM^{+}84, GGW83]. 2A [WNB82].

3 [CQB84, DMD80]. 3’-diaminobenzidine [CQB84]. 31P [GOSW82]. 35 [BF82c, CL81, FB82]. 3H [BO81a, BKO81, ML84a, SL80, TTGA82, UAL84]. 3T3 [AKM^{+}82, FNS81, JB81a, JB81b, NCBP84, NNRR80, PL83b, SGW^{+}82, SM82b, SBVW82, VPMF84, WR83]. 3T3-L1 [NNRR80].

4 [SEE84]. 42s [MLZD83]. 431 [ABG81, CMC81]. 450 [MMO^{+}81]. 45Ca [MHLS81]. 48- [WES84]. 4F2 [ASM^{+}84].

5 [MEKT84, WB83b, DMD80, DSD80b, DSD80a, TD83, EPO^{+}84, SHHR84]. 5’-AMP [DMD80, DSD80b, DSD80a, TD83]. 5’-nucleotidase [MEKT84]. 5’-phosphate [WB83b]. 5-triazido-2 [MF84a]. 50- [NS83a]. 51 [FYC82].

5S [KSG81].

6-trinitrobenzene [MF84a]. 6301 [LYG81].

7SL [WB83a].

9 [BTJ82, GLMW84]. 90 [NBMD81]. 95 [CL80b]. 95-Kdalton [CL80b].

A-431 [ABG81, CMC81]. A-colloidal [SG81]. A-induced [JHCC84]. a-protein [ZP80]. a/b [KTW83, MS80a, SBG^{+}81]. A1 [BR84a]. A431 [Bar82a, HT83, WCZG84]. AB2 [RMAF80]. ABH [Kho82]. abilities [GCL83]. ability [GLMW84, SSCP80]. abluminal [SSP82b]. Abnormal [HDB82, WCJ83, SKN83, WWKW83]. Abnormality [HWK^{+}84]. abnormally [LBM82]. Absence [KHL82, NB81, Sta81, YKJF80, BRH82, HOMR82, IER84, TOK^{+}84]. absent [FYC82]. absorptive [GN84, Mad83]. absorptive-cell [Mad83]. Abundance [BFST83]. Acanthamoeba [BO83, CBP84, Bow80, BOH81, KKP84a, KP84, KKP84b, MFP80, OT80a, Pol82, Pol84b, TP82, TRC^{+}84]. acceptor [LYG81]. acceptors [BRS^{+}82]. accessibility [dPMMF^{+}81]. accessory [MA82]. accompanied [IY84c, KPPC83, Mad83]. accompanies [BL83b, HH81b, LBM82, LH84]. accompany [WG82a]. accompanying [Bro81]. according [EBS84, SHT^{+}83]. accounts [MCPB80]. accumulation [Asc84b, BS82a, BB82c, FMS^{+}84, HMKM82, KSC82, KC82, MS84c, RKK82, Wal82a, WLFD84, WB83b]. acetate [JB81b, MR82, PWWS80, PWSW80, WCZG84]. acetates [SNdC^{+}84]. acetyl [RBO83, Rot84]. acetylated [VVBZ84]. acetylated-low [VVBZ84]. Acetylcholine [AKT84, HMR82, AWB82, AF83b, AH84, ABSC81, Bar82b, BAF84, Blo84, BN83, BNGN84, BU84, BF84b, BMA84, CSR^{+}81, CW80, Con84, GNW^{+}84, Har80b, HE81, HF83, KHS80b, MS84c, NB81, Pen83, PP84, Pen84, PFP^{+}82, PB83b, SSHP82, SH83, SHT^{+}84, SWF84, SB82b, SN80, SA83b, WRH81]. acetylcholine-receptor-rich [CSR^{+}81]. acetylcholinesterase [DK81a]. acetylglucosamine [VC84]. acetylglucosaminidase [RMJ84]. acetylglucosaminyltransferase [CK82]. acetylmannosamine [BO81a, BKO81]. acetyltransferase [JRM^{+}80]. Achlya [KCGH84]. acid [ASC84a, AF84, AC81b, BO81a, BKO81, BBCL82, FS80, GMO84, JVMRK82, KCGH84, MP83, MG81, OHA^{+}84, PA83, RJVS81, SGW^{+}82, Sea82, SG84a, SSC80, UAL84]. acid-glutaraldehyde-saponin [MP83]. acidic [HSME84]. Acidification [MTMT83, YFM83, AF83a, HM84b, MPC84, SNdC^{+}84, TKM83]. acidosomes [AF83a]. acids [BALR83, CPS84, Cre81, KRB^{+}84c]. acinar [BBR83, IJ82a, IJ82b, IHJ82, Oli82, PE83, RMJ83, SDGYJ84]. acquire [GBB^{+}82]. acquired [CFW84]. acquisition [FML83, TH84]. Acroschismus [RN83]. acrosinlike [ZCH80]. Acrosomal [IT82, TI82b, BM83, DTBF82]. acrosome [MM81c, PV84, SSG80]. across [AR81, GBW82, GWB82, MD80a, MD80b, OJJ^{+}84, SJC80, WIB81, WB81a, WB81b]. act [SMT80]. ACTH [OBH84, Sur82]. ACTH/ [ME81]. ACTH/beta [Sur82]. ACTH/beta-endorphin [Sur82]. Actin [ASG83, CSHP84, FBRP84, HTS80, MB84b, OB81, PSSS82, TJ80, TBD81, TS83b, TEDS83, TBCM83, AP84, AFI^{+}81, AUT83, AFSF^{+}81, BYW81, BP83, BRH82, BSH83b, BS82c, BM83, BFST83, BS81a, BYS82, BRS81, BKM82, CBW81, CCJC82, CF80, CRC^{+}82, CT83, CT84, CV82, CVC^{+}84, DTBF82, DGS80, FD82, FP81b, FT80, FLH^{+}81, FA82, FA83, FSVW80, FK80, GWS84, Gla83a, Gla83b, GGW83, GHL84, HLS81, HCP81, HT82b, HTFH82, HM84c, HM84d, HM83a, JTM80, KZS81, Kan80, Kan83, KOST80, KA84, Kir80, KC80b, Leh81, Let83, LTGL80, LFS^{+}81, LGHI84, Mab83, MB82b, MP83, MBR84, MA82, MGW^{+}80, MPW82, ML84d, NSI80, NSI82, NRW82, NAH83, OB82, OT82, OBH84, OKE81, OH84, PWM81, PS82, PFM^{+}84, PG83c, PM81, PSM84, Pol84b, PNYO83, RSL81, RS82, RFUC84]. actin [SS80b, SDL81, SWP81, SNPE84, SSP84a, SKKC84, Sma81, SBB81, SHBS80, Sto84, SLL80, SIM^{+}84, TWH80, TRSS81, TDM80b, TT84b, TP82, TTI^{+}81, TTI^{+}83, TAB80, WWPL84, WS84, Wan84, WMW84, YHS^{+}82, YRS83, YPR^{+}82, YAF81, YMF84, ZS82, ZS84, dCSB84, Pol84a, SBB81, TTI84]. actin-based [Kan80]. Actin-binding [HTS80, BFST83, CVC^{+}84, LFS^{+}81, NAH83, OT82, RSL81, RS82, SBB81, SHBS80, SIM^{+}84, TAB80]. actin-bundling [MB82b]. actin-bundling/depolymerizing [MB82b]. actin-calmodulin- [HM83a]. Actin-dependent [OB81, OB82]. actin-depolymerizing [Mab83]. Actin-independent [ASG83]. Actin-membrane [MB84b, BKM82]. actin-myosin [OT82]. acting [EH81, RACGB81]. actinin [BH83, BF82b, CP82b, FFRF83, FB80, GDTS81, HRBJG80, KPPC83, LdMM^{+}84, LMYM84, MS80b, NDC80]. actinins [CL80b, EM84]. actinogelin [MA82]. actins [LS80]. Action [MBB83, Sch82, BS81a, CBP84, Fis82c, FJ82a, NBSW81, PWM81, RS83, TP82]. actions [GLBS81]. activatable [EB83, GCS81a, GCS81b]. activated [AFK^{+}84, MF84a, RKH82, vDB80, PG84]. activates [OB82]. Activation [GDH82, BHSC83, CHI^{+}81, Gil83, Gol82, GN81, IER84, LZR82, LCA84, LA84, MHCG80, MNW81, MR82, PGD^{+}84, RF80b, SEE84, TD83, WB82]. activator [DVB^{+}81, EB83, GMJR82, JB81a, JB81b, KS84, LSN^{+}84, LTBS80, SLK^{+}84, YKJF80]. activators [GPR83, LL82b]. Active [HF80, BE82, BDP^{+}84, EPO^{+}84, FNNFS84, HAS^{+}81, Ko84, NB81, RCJS82, YK84]. actively [GBM84]. activities [FQBN81, GIKR84, NDS84, SH82b]. activity [BM80, BFST83, BL82b, CK82, FR81, GDM^{+}80, GCF^{+}83, GCS81a, GCS81b, HCP81, HTGB84, JB81a, KP84, KKBG83, KB81b, LFGR82, LBM82, LFS^{+}81, ML84b, MTMT83, MCPB80, Mil81a, MM81c, MHW83, MWH83, NRMB84, NPG^{+}80, PBG80, Pie82, RVM83, SKN83, SGH^{+}83, SDM84, SBM84, SBVW82, Sur82, Vah83, WMP^{+}84b]. acto [Vah83]. actomyosin [SM83c, KP84]. acts [HWS82]. acute [BJG83]. Acyl [RFML84]. Acyl-Coa [RFML84]. adaptation [DSD80b, TD83, Van83, WB83b, DSD80a]. adapted [EDK80]. adaption [VV83]. Adaptive [Sin80]. added [Moo81]. Addition [SGHS^{+}80, SJM81, HZM80, HFW82, ZB82]. adenocarcinoma [JHCC84]. Adenosine [EH81, BL82b, MSP84, MW80, PB84b, VSD84]. adenosinetriphosphatase [AFK^{+}84]. adenyl [CVGM83]. adenylate [HS81, TD83, VKHH82]. ADH [KMF81, MKD80]. ADH-induced [KMF81]. ADH-stimulated [MKD80]. adherence [CKB^{+}80a]. adherent [GSB81, MMLT84]. adhering [DH81]. adherons [SLKB83]. Adhesion [CHRT83, CFML83, IER84, MS80a, SB82a, BGE84, BC83, CRH81, CL80a, CG84, CEC84, DWS84, EF84, GGVR82, GSMR83, Gri80, GE84, HJ80, HJ81, JKHT81, JH84, KKM81, LNB84, LLB84, MBZB81, MCN^{+}84, Mar80, OFO83, OYNT83, RCH81, RH81, Rau84, RGE83, SL82a, SLKB83, SB84, SGPB84, SR81a, SM80, SCM83, SB80c, TGPC82]. adhesion-deficient [SL82a]. adhesion-mediating [SL82a]. adhesions [MHCG80]. Adhesive [DGD^{+}84, GGL80, HEA^{+}82, LSC83, SR81b, STR84, YK84]. adhesiveness [HLM^{+}84b]. Adhesivity [EL84]. adipocytes [SJ80]. administration [PAGG81, PYC^{+}82]. ADP [KGAB81, LN82, LMN83, Pol84a]. ADP- [LMN83]. ADP-actin [Pol84a]. ADP-induced [LN82]. ADP-ribosylation [KGAB81]. adrenal [AT83, OBH84, PBWK83]. adrenergic [HS83, MSB^{+}83]. adrenocortical [MR80b, MB82c]. adrenocorticotropin [BK84, GLP80, MGK83]. adsorbed [DWP81]. adsorptive [GSOG80, GSH^{+}84, MS83a, PvD83]. adult [ASM^{+}84, BS82b, BB82a, CEC84, CFN^{+}83, JWG84, JSDM82, NCM^{+}83, RMS^{+}82, RMF83, SFS81, VB81, WWCD84, WRM83, WRH81, AMM84a]. Aequorin [TBR80]. affect [AMCG83, WSH84]. affected [KOLV81, SHO^{+}84]. affecting [BB82c, CAZF84, FK80, SLLP82]. affects [BYS82, HM84c]. affinity [HSS82, JNS^{+}84, LGHI84, SF84b, SPEL84, SK83]. African [GIKR84]. after [BHJ83, BO81a, BKO81, BBLY83, BSS83b, BK84, DSD80b, ERB83, GS84e, HS83, HR81b, HHS^{+}84, HS84, KLF80, KML83, KOS83, LR81, LSM^{+}84, LN82, LA84, MWI^{+}84, MTH84, MB81, Nac80, NKH82, PAS84, PBS84, PHM81, RF84, SGHS^{+}80, SSL^{+}84a, SLS^{+}84]. against [AB80, BMF84, CLJ^{+}81, ERN83, FGS^{+}82, HAS^{+}81, LFS^{+}83, MS84a, MWI^{+}84, OAOB81, OMG82, PBPD83, PG83b, PKB84, RBH^{+}84, SBC81, SSP84a, WDE83]. agarose [DHHG83, RFLR81]. agarose-embedded [DHHG83]. age [JRB80, Per81b]. agent [LA80b]. agents [AS81a, SJ80]. agglutinin [EL84, LSM83, LA80b, MMKL81, MRPK82, OMG82, PV84, SH82c, SSC80, VS82, dSTK81]. agglutinin-resistant [SSC80]. aggregated [Cre81]. Aggregates [AF83b, BAF84, CBW81, GOA83, GAA83, HBS^{+}82, MKD80, NB81, OT82]. Aggregation [SM80, AKT84, GNW^{+}84, HFW82, LFF^{+}81]. Aggregation-dependent [SM80]. Agrobacterium [WPKB83]. AI [SSJS84]. AII [SSJS84]. Ala [SGM^{+}84]. albicans [SM83b]. albumin [BJS81, MNW81, SSZ83, SWvK^{+}83, SHP83]. algae [NW82]. algal [HMM82]. alkaline [BL82b, CSR^{+}81, RPB81]. Allogromia [TA81, TKA83]. allow [BHP83, KZS81]. almost [NRW82]. along [HA84a, ML84a]. alpha [AMCG83, APZL80, AL84, BH83, BNM82, BF82b, CMF82, CP82b, CL80b, CBL84, DWP81, EM84, FFRF83, FBS^{+}81, GDTS81, GWWP83, HAEC81, HTGB84, HRBJG80, HBB82, KSSR84, KPPC83, LR83b, LdMM^{+}84, LMYM84, LL81, MSB^{+}83, MA83b, MS80b, NDC80, PDK83, SGH^{+}83, SSLB84, TAC84, VMMC84, WWS83, WW83]. alpha- [AMCG83, KSSR84, PDK83]. alpha-actinin [BH83, CP82b, FFRF83, GDTS81, HRBJG80, KPPC83, LdMM^{+}84, MS80b, NDC80]. alpha-actinin-enriched [LMYM84]. alpha-actinins [CL80b, EM84]. alpha-amylase [MA83b]. alpha-ethinylestradiol [HAEC81]. alpha-fetoprotein [CMF82]. alpha-glucose [AL84]. alpha-granule-deficient [GWWP83]. alpha-granules [SSLB84]. alpha-latrotoxin [VMMC84]. alpha-mannose [AL84]. alpha-toxin [FBS^{+}81]. alpha-tubulin [BNM82, CBL84, LR83b, WWS83, WW83]. alphavirus [KHS^{+}80a]. alter [RSP^{+}83a]. Alteration [GBWR83, HWBC82, MTH84, VS82, HEG81, Sin80, WLCS84, YBG81]. Alterations [BVGS82, PS81, BS82b, LFAG84, Mad83, MS84a, RS83, RKK82, TQ82]. altered [DBBN81, IY84b, LK83, MJM83, VTSS80, MG83b]. Alternate [TT84a]. alveolar [BVGS82, KK82, TSS82, VB81, VKE^{+}81]. alveolar-macrophage [KK82]. amacrine [SSJS84]. ameba [HT82a]. amebae [RFUC84]. amebas [TWH80]. amebocytes [OR81]. ameboid [TWH80, TBR80, YMF84]. amiloride [BC84b]. amiloride-sensitive [BC84b]. amines [SBF84]. Amino [SG84a, BALR83, CPS84, GMO84, MG81, OHA^{+}84, PA83, RJVS81, KCGH84]. aminoacyl [Deu84]. aminoacyl-tRNA [Deu84]. aminopeptidase [RH83, RH84a]. aminotransferase [CL81, O’B80]. amnion [CLFC83, LVL^{+}82, RLT81]. amoeba [GK81a]. amoeboid [NRW82, RW82b]. among [TAC84]. amorphin [CP82b]. amounts [MACK84, SM83a]. AMP [AM81, CGG84, DK81b, DMD80, DSD80b, DSD80a, FTM82, GSMHF84, GLBS81, LLS81, MR80a, MMC83, MKB81, MF84e, MSD82b, NKT80, OB82, OBH84, Ste83a, Ste84b, TD83, WR83]. AMP-dependent [LLS81, MF84e, MSD82b, Ste83a, Ste84b]. amphibian [BDP^{+}84, FCO83, Har80b, KSD^{+}83, LM84, MBC80a, SHT^{+}84, SY80, WS81a]. amphiphilic [BTJ82]. Amplification [AFK^{+}84, LBM82]. Amplified [KHS82, FDSWW84, FKS^{+}81]. amylase [MA83b]. analog [OHA^{+}84, SP84b]. analog-modified [OHA^{+}84]. analogue [RJVS81, SBSL84, YSS82]. analogues [SSG80, YRS83]. analyses [BS82d, OGS82]. Analysis [BP83, BLH82, BG83, KLC83, KJ82, Kuo82, RZSW83, RV80, RMJ84, SMS84a, TH84, AHPL81, BMF82, BB80a, BP80, BCC^{+}83, BOH81, BPB^{+}81, BBS81, CQB84, DF84, FO80, FGHC80, Fis82a, GDM^{+}82, GR80, GBG82, GDH82, GAS83b, HAF82, HA84b, HAFG82, HPRL81, IJ82a, JL84, Jef82, KL82, KHL82, KOB82, LL82a, MFP80, MSB80, ME84, MIRS84, MNW81, MYR81, MC80, NRMB84, OBC^{+}84, PG84, PvD83, Ric80, SHHP81, SSL^{+}84a, SLS^{+}84, SF81, SB80b, SSJS84, SGKC82, SEPB81, SBP81, SDL82, SSF81, SR81b, SCS82, SHWS83, SLL80, TTGA82, TJB83, WES84, WLFD84, ZSL82, ZFP84, dLTvdSB83, MSC80a, SBP81]. Analytical [AC81b, MB80a, WTSAC^{+}81]. analyzing [TOG83]. anaphase [CMM81, JL84, NKH82, Nic83, SB80a, WHJ83]. Anchorage [LTBS80, CFW84, CRGS82, JHCC84]. anchorage-independent [CFW84]. anchorin [MBLvdM84]. angiotensin [GAAG82]. Angiotensinogen [BCC^{+}83]. angle [LP83a, PL83a, UE83]. animal [MBZB81]. anionic [BF82a, RKF80, SSP81, SSSP81, SSP82b]. anionic-lipid [BF82a]. ankyrin [ML84c]. annulate [SS84b]. antagonist [BSP83]. antagonists [GGS83]. antennae [WOBR80]. anterior [APZL80, SF81]. anterograde [MMKL81]. anterogradely [TI80]. Anti [PATR82, BF82b]. anti-alpha-actinin [BF82b]. Anti-immunoglobulin [PATR82]. antiactin [BF82b, LS80]. antiactivator [EB83]. antiactivator/antiplasmin [EB83]. Antibodies [BBLY83, FHG^{+}81, HCK80, LRW82, RSP^{+}83a, RBH^{+}84, YF81, ABBG82, AB80, BBWS83, BMF84, CLJ^{+}81, DF84, DSRF84, ED84, Fro84, GRR^{+}84, GV82, GKL^{+}84, HP81, HO84b, HAS^{+}81, HM84e, KKP84a, KP84, KKP84b, KWM82, KDL84, Lan82, LMYM84, LFS^{+}83, MB84c, MS84a, MWI^{+}84, MDA83, NKH81, OAOB81, OMG82, PBPD83, Pip84, RMS^{+}82, SHT^{+}84, SSG^{+}83b, SWF84, SSP84a, SDGYJ84, SF84b, TOG83, TLPTV83, WDE83, WB83a, WWCD84, WCY^{+}83, WMENS82, ERN83]. Antibody [WS81b, ASM^{+}84, CVT84, CF84a, CD84, EB84, FGS^{+}82, KKP84a, KML83, LQ82, LGF^{+}84, LS80, MPU84, MLD82, MSD82b, NLD^{+}82, PKB84, PHM81, SHHP81, SHPH84, SGPB84, SLS81, SB80c, TFAP81, TAC84, WWS83, WW83, WES84]. antibody-binding [KKP84a]. antibody-dependent [PHM81]. antibody-induced [KML83]. anticoordinate [SKS84b]. antidiuretic [HBS^{+}82]. antifibronectin [MK84b]. antifolate [LBM82]. antifolate-resistant [LBM82]. antigen [ASM^{+}84, Bur82a, CF84a, CF84b, HEO82, HPG83, IWM83, MHNT81, SP84a, TFAP81]. antigenic [ERH82, LFAG84, dPMMF^{+}81, Pol84b]. antigenically [SB84, SF84b]. antigens [BBWS83, CBS^{+}84, CHH^{+}83, DTT81, DBL82, EWG82, EW82, EZ84, HS80b, Kho82, MNP^{+}84, MP84c, PBFC^{+}84, Pip84, RB82a, SBC81, SP82a, SSE84]. antimonate [WH80]. antimyosin [BF82b]. antiparallel [WM81]. antipeptide [SSG^{+}83b]. antiplasmin [EB83]. antireceptor [ABBG82, MPU84]. Antisera [HS80b, DKDB81, MNP^{+}84, PG83b]. antiserum [CK80, DBL82, SBC81, TOG83]. antitubulin [KWM82]. anucleate [MG81]. aortic [CCCR81, MM81a, SBCS^{+}82]. apical [FHG^{+}81, MBP^{+}83, MS84b, PS83b, PAS84, PBS84, RIP^{+}84, RKMF84, RBPSB84]. apolipoprotein [SGA80, ZRW80]. apoplast [SGKC82]. apoprotein [BR84a, WC83]. apparatus [ACTK84, BKL^{+}82, BO81b, BLS^{+}80, DR83, FP81a, Fle81, HSB^{+}82, Hep80, LOL81, MACK84, PRK^{+}82, RIP^{+}84, RS84, SBK^{+}84, SP84b, TLPTV83, VC84, WHJ83]. Apparent [CH80d, CH83b]. Appearance [HBR81, MGMS84, WR81, BM82b, PMP81b, SC81, WA84, WD83]. Application [BBS81, SEPB81, SBP81, FHFL82, SF84b, Wol80b]. applications [CKP84]. applied [SR83b]. approach [BTHL81, KZS81, SEE84, dHLC84]. aqueous [MK84a]. arachidonic [ASC84a]. arachnoidal [KSMF84]. Architecture [EL83, CP83, FSKJ81, GBM84, TOG83]. area [GRB82, OCT81]. areas [ASG83, BOH81, BN83, GAR^{+}84]. arm [GH82a, TTUI83]. arms [MW80, ONO82]. Arnold [TA81]. Arrangement [Lan80, LOL81, SB83b, WW83, WFR84]. arrangements [MBB83]. array [HPG83, KKMK84, SA82]. arrays [AFSF^{+}81, HCMK81, KNHK84, KNK84, Man82, WSO^{+}81, WD83]. Arrest [CH80c, CKJ83, IY84c, KS82, LR81, SHW^{+}82b, WB81b]. arrested [MLNK83]. Arsenate [KSSP83]. arterial [VSS83, WH83]. arteries [WH83]. arthropod [Bro81]. articular [PPRR82]. artifact [FP81a, MHC80]. artificial [SBP81]. ascending [EHK80]. ascites [CCJC82, GOSW82, GKZ81, JHCC84]. ascorbate [DPK84, SBCS^{+}82]. asialoglycoprotein [GSS^{+}82a, MNST82, SFU84, SBF84, TKM83, Wei80]. asialoglycoproteins [HWAK83, WKAH84, ZH82]. Asparagine [DCS84, DWS84, DR83]. Asparagine-linked [DCS84, DWS84, DR83]. aspartate [NLPH82]. aspect [ASG83, SSP82b]. aspects [SPHE80]. aspergillus [OM83, GBM84]. assay [BM83, RACGB81, SCS84, SCS82]. assays [Fis82b]. assemble [AFSF^{+}81]. assembled [BL83a, LL82a, SHT^{+}83, SD82, WIB81, WB81a, WC84]. assembles [TBCM83]. assembling [CVC^{+}84]. Assembly [CHS82, CP81, EDP82, Ful84, LSM^{+}84, RKF80, RBPS83, ZK83, BSG81, BB80a, BP80, BM83, BCP^{+}81, BYS82, CWBB83, CTKJ83, CT84, CCWR84, DB83, FKK83, GWG82, HWBC82, Hat81, Her80, HM84c, HM84d, HPRL81, JB81d, JB82, Kan83, KWSB84, LR83b, LL82a, LL81, LM84, MWP84, MWM^{+}81, ML83, ML84c, MM81b, OB81, OLH80, PS82, PBW80, RA81, RPB81, RBS84, RBMM82, SMS84a, SKCJW84, SBGD84, TdS83a, TRSS81, Wal84, Wei82, WMP83]. assembly-disassembly [CTKJ83, DB83]. assembly-with [RA81]. associated [ATW83, AF83b, BM82b, BG82, BK83b, Blo84, BV83, BLV84, BSV84, BL82a, BW80, BN83, BB80c, BB80d, CCJC82, COBC84, CMC83, DBG^{+}81, DPS80, DCK84, FDSWW84, FHG^{+}81, FWG^{+}81, Ful84, GBS80, GLS83, HTKC82, HAS^{+}81, HPM82, HM83b, HM84e, IWM83, Law83, LV83, LLS81, LLS82, LW81, LNA81, LV81, LOL81, MS84a, MPW82, MP82, MB80b, Mur80, MHW83, MWH83, OOP^{+}81, PWO84, PKB84, PL84, POYF81, PPR^{+}82a, PL83c, RMT84, RQHR81, SDL81, SKS^{+}84a, Sch84, SW81b, SD80, SDK^{+}84, SD82, TFGC84, TBD81, TQ82, TQS84, TQG84, UM82, VDT81, Val82, VBT84, WCY^{+}83, WRM83, WSM^{+}82, WD83, WWES83, WOBR80, Woo80, dCSB84, LQ82, SKS82]. Association [BV83, BNB^{+}82, CPLC80, CN80, DYP83, EKSZ80, GSMHF84, KC80b, NM84, PPR^{+}84, Sin82, SSSJ83, SD82, WSB^{+}80, ASG83, BLG^{+}80, BGNT81, CF80, DHHG83, DMB84, FLH^{+}81, HSGD83, JNS^{+}84, KOST80, LBA^{+}84, MB84b, PWM81, RCM^{+}83, RCJS82, SDG^{+}84]. Associations [RS84, BTHL81, Fri83, NKH82]. Asterias [ER84]. asters [SSBB82]. Astrocyte [LR81]. astrocytes [LDV83, SFS81, TIK81]. Astroglial [HL81, MdV80]. Asymmetric [SDZ84, SM81, BCHG83]. Asynchronous [SC81]. ATP [BDPS80, BSSR84, Bur82b, HSC84, MS83b, OK82, SCS84, TT84a, TWG83, TTUI83, WC84, YFM83, YT82]. ATP-dependent [Bur82b, SCS84, TTUI83, YFM83]. ATP-disintegrated [TT84a]. ATP-induced [WC84]. ATP-sensitive [HSC84]. ATPase [BHJV80, BSSR84, DPS80, ES81a, HSC84, JW83, JSDM82, JSMT82, KP84, MS84e, MM81c, MHW83, MWH83, POS80, PF83b, RBH^{+}84, SSBR84, SHZN83, SMJ80, Ste84b, TOT^{+}84, Vah83, YT82, ZKF84]. Atrial [GSS82b]. AtT [MGK83]. AtT-20 [MGK83]. attached [Kuo82, RBPS83]. Attachment [KSMF84, ACF84, BKM82, CEC84, FBST84, GAR^{+}84, HSW82, HEA^{+}82, JKHT81, JH84, LC81, MTT80, PSC83, RK80, SPBB84, SK83, VS82, VC84, WYK^{+}84, vEvV81]. Attractant [BWM84]. Attractant-induced [BWM84]. aureus [FBS^{+}81, MFP82]. autoantibodies [BPB^{+}81, DTT81]. autoinhibit [YK84]. autointernalization [BYW81]. autonomy [Baj82, DK84b]. autophagic [MAG82, SG84a]. autoradiographic [MHLS81, PYC^{+}82, RMJ83, TAGG80]. autoradiography [Fis82a, GDL80, HLSWH82, SL80]. autoregulatory [CH83b]. autosomal [JL84]. auxotrophic [ZHTH80]. availability [CNH80]. average [SHT^{+}83]. Avian [NGL83, AMK83, AMJ^{+}81, BMF82, BL83a, BDP^{+}84, BL82a, COS83, FML83, GLBH82, GL83a, GRL82, HE81, LWW^{+}83, MCN^{+}84, ML84c, RCM^{+}83, SS84c]. axenic [KC83]. axial [ME82, SSJS84]. axolemma [GPDM81, WE80, YDL^{+}84]. axolemma- [YDL^{+}84]. axon [ML84d, SW81b, SW81a, TQG84]. Axonal [BGM81, BTHL81, BTL84, FG80b, LGB84, TAGG80, BL80b, CBL84, HSB^{+}82, HLH81, HHT^{+}84, HGP84, MMKL81, ML82, PYC^{+}82, ST83, SFB^{+}84, WE80]. Axonally [SW81b, Ber80b, GL81, GPDM81, LSM83, LW81, SW81a, WSB^{+}80]. axonemal [TT81]. axonemes [BFW80, ONO82, OAOB81]. axons [BNM82, Bur84, Can82, Hir82, NBM82, SR82b, SM83a, SP84c, SBGP84, SSP84b, TI80]. axoplasm [GS84b, ML84d]. axoplasmic [BNM82, EP80, ML82, NBM82, PAGG81, Ste82a]. axostyle [WL80, YT82]. azacytidine [SC84b]. azurophilic [BSS83a, BSS83b].

b [BPM80b, BHSC83, GL81, KTW83, MS80a, MW80, RCJS82, SBG^{+}81, TQG84, GA84, LFC^{+}80, LL81, ML84b, MR80b, NBSW81, ORA^{+}84, PWM81, POYF81, PATR82, SCS80, TQG84, WG82a, ZRW80]. B-cell [ORA^{+}84]. b-cytochrome [BHSC83]. B’-iminodipropionitrile [TQG84]. B-mediated [WG82a]. B16 [MF84b]. b5 [BM80, MCPB80, RVB80, Sta81]. b559 [ZP80]. b6 [MS83b]. baby [GQW83, Gri80, MG83a, OMG82, WWDYW83]. backwards [SHRL84]. bacteria [EHS80, Tam82]. bacterial [KFV83]. bacteriorhodopsin [SMG81]. bacterium [VF82]. balance [KW84, KMF84]. BALB [FNS81]. BALB/c [FNS81]. BALB/c-3T3 [FNS81]. band [CP82b, CN80, FT80, JSC84, MS84d, SMAS81, SD80, SS84c, WD83]. Banding [SS80b, LBM82]. bands [NJSC83]. barbed [BM83]. barium [LHHH82]. barley [Hep80]. Barr [NPG^{+}80]. barrier [GBWR83]. barriers [SGKC82]. Basal [GW82, KBDK84, AF83b, AKT84, BWW^{+}80, DB81a, LLM82, MS84c, MFVPK82, OCT81, OOP^{+}81, Oli82, PB84a, SH81b, Wat84, WWWC80, WCJ83]. basal-body [WCJ83]. based [Ino81b, Kan80, VVBZ84]. Basement [FFGC84, ACMH83, BVGS82, Ekb81, FGS^{+}82, FML83, GLMW84, GRZ83, HBO84, HTKC82, ILL83, KLF80, KHF81, KR82a, LLM82, LGF^{+}84, MRFF80, MPF83a, OWMB82, RMAF80, San82, SSP82b, VB81, WWCD84]. bases [Max82, PO81]. basic [DM83, OP81, OWPC82, RPHA81, SS84a]. basis [FBP82, MM81b, SHT^{+}83, TWH80, TBR80, WL80]. basolateral [CVGM83, MSRB84, RIP^{+}84]. basophil [TTGA82]. basophilic [MLP84, MHB84]. BC3H1 [MCG82]. be [EN84, WKH80, YBG81]. Bead [Bro81, BF84a]. beads [Bec84, BF84a, Gri80, LGHI84, MG83a, TEL84]. bearing [AC83, BGS81, SC82b]. beat [HW83]. beating [KW84]. become [TT84b]. becomes [SG82a]. been [VSBL83]. beetroot [MB80a]. before [BM82b, BSS83b, CLHA80, CKS82, LSM^{+}84, MS84b, MTMT83, Ric80]. Beginning [OR81]. beginnings [PB81]. Behavior [Hea80, HC83, CLFC83, DHS^{+}84, FTM82, KML83, Kub82, MKW^{+}82, MB81, TBCM83]. behavioral [HHT80, ZLK81]. beneath [HH81a, LSM83]. bent [TEDS83]. benzimidazole [HA80]. benzimidazole- [HA80]. benzo [SCSC81]. benzyl [SNdC^{+}84]. berghei [WWES83]. berghei-infected [WWES83]. Bergmann [SLS81]. Beroë [TT84a]. Beta [SP84a, AMCG83, BBB82, DBL82, EPO^{+}84, FFEL81, GK84, GLP80, HC84, HS83, KSSR84, LL81, MSB^{+}83, NLN^{+}83, PAGG81, PYC^{+}82, PD80, PDK83, RFML84, RMJ84, SHHR84, Ste84b, SC84c, KHJG84, SGKD83, TS83a]. beta-adrenergic [HS83]. beta-cell [DBL82]. beta-endorphin [Sur82]. beta-galactosidase [NLN^{+}83]. beta-globin [PD80]. beta-glucuronidase [GK84]. beta’-iminodipropionitrile [PAGG81, PYC^{+}82]. beta-lipotropin [GLP80]. beta-oxidation [RFML84]. beta-tubulin [AMCG83, HC84, KSSR84, PDK83]. between [BC84a, BPDR83, BO83, BTHL81, CKB^{+}80a, CKB^{+}80b, CH81, DANH84, FJ82a, Fri83, GK81a, GS84d, GE84, HCP81, Her80, Hir82, HT82b, JB81a, KW84, KC83, KKP84b, LLS82, LS81b, LS80, Mas84, MS80a, MSC80b, NW82, ONO82, PG82, RMU84, RUB84, RGE83, SSJ83, SHH82, SSE84, SG82a, SH81a, SCP^{+}83, SK83, STR84, SSSJ83, SD82, Sur82, TDM80a, VN82, WP80, WM81, WD84, Wri81, WA83, WLCS84, WNB82, YHS^{+}82]. BHK [WWDYW83, HKSF80]. BHK-21 [WWDYW83, HKSF80]. Bidirectional [GS84b, TTI84, MSC80b]. bifunctional [RKM^{+}81]. bilayer [CAZF84]. bile [CHH^{+}83, RH83, RH84a, RH84b]. biliary [SFU84]. biliprotein [LYG81]. bind [GHL84, HCK80, KKP84a, Wei80, YF81]. bindin [GGVR82]. Binding [BLG^{+}80, GWWP83, Gol82, GB82, HG80a, MLM83, MLHM84, MTM83, TTI^{+}83, Bad81, BCSC82, BGNT81, BVAR83, BTJ82, BBWS83, BFST83, BMS83, BWV84, CBS83, CVC^{+}84, EMA83, EL84, GGK82, GS80a, GWS84, GFLG^{+}84, GBL81, GE84, HAEC81, HWK^{+}84, HTS80, HSS82, HM83a, HHD83, ISGS81, KKG84, KKP84a, KL81, KNH80, LFC^{+}80, LSC83, LOC81, LH84, LFS^{+}81, LGHI84, MG83a, MLP84, MKB82, MK84b, Moo81, MRB84, NAH83, OT82, Ott83, PS83a, PWM81, PV84, RZSW83, ROW^{+}83, RSL81, RS82, RMJ83, RMJ84, SHHP81, SHW^{+}82a, SSF81, SH82a, SBB81, SSC80, SHBS80, SR83a, SSP84b, SNS83, SIM^{+}84, TTGA82, TV83, TRN^{+}82, TAB80, VS82, VSS83, VKE^{+}81, WB81a, WGRB80, WMP^{+}84b, WWWC80, dSTK81, CGS81, MLZD83, VVP82]. binds [BS81a, CB82, MPH82, PH81, SGPB84, WIB81]. Biochemical [ARL^{+}80, ERN83, FCL^{+}82, IJ82a, LN81, NRMB84, OWMB82, SLL80, SPHE80, BLG^{+}80, BPM^{+}80a, BTL84, BS82d, DK84a, FKB83, JRM^{+}80, JRB80, LSBPG84, Luc84, MAG82, RMJ83, SG84b, WOBR80, ZH82, KHJG84, NCM^{+}83, QGW84, SDC80, WWS83, YDL^{+}84]. Biogenesis [ML84c, JB81d, JB82, PJ83, SA82]. biological [DK84a]. Biologically [BDP^{+}84, YK84]. biology [Wei83, PB81]. biomatrix [RGM^{+}80]. biophysical [WOBR80]. Biosynthesis [BR83, BR84a, CEC^{+}83, GPHV83, HO84a, HBO84, Ioz84, LNB84, LLB84, MA83b, RKP^{+}82, RMH^{+}84, WM84, BZ83, BM80, CH80d, CSW80, GGK83, GB81, GFLG^{+}84, GKZ81, MCPB80, PLMS81, SMAS81, Sin80, VC84, BPM80b]. Biosynthetic [SBG^{+}81, SHZN83, ACMH83]. biotin [MLSW82]. biotin-labeled [MLSW82]. bipolar [HTS80]. bipolarity [TH84]. bird [TS83b, TEDS83]. birefringence [HHSS81, HHS^{+}81]. bisphosphate [RBMM82]. bivalents [JL84]. bladder [CFN^{+}83, HBS^{+}82, KMF81, MKD80, NCM^{+}83, RV80, SCSC81]. blastoderm [WMW84]. blastomeres [JZ81, Sob84]. bleaching [LSM^{+}84]. blebs [TWW82]. block [ERB83, FNNFS84, JGSH82]. blockage [KI83]. blocked [BCLT^{+}80, EPO^{+}84, NS83b, WA83]. blocking [LTGL80, SMH84]. blocks [EMC83, HH81b, IY84c, MSD82a, OM83, OT82, PV84]. blood [BL80a, BBLY83, Kho82, MD82, SCN^{+}81, vDB80]. blots [SF84b]. boar [AdS83, AdS84]. bodies [BS82c, CHS82, CHCS82, DHS^{+}84, GW80, OOP^{+}81, ORA^{+}84, VC84, WOM82]. body [Bro81, WCJ83]. bonding [Car84]. Bone [NvIvGdH^{+}82, AHMS82, BHA82, BvdMN84, LSBPG84, MTG^{+}82, PPR^{+}84, RNH^{+}83, SR84a, STRP83, SHJA82, SB83a, SGKD83, WR81, WC83, WHD^{+}84]. bone-forming [BvdMN84]. bone-specific [WHD^{+}84]. border [BP82, Bur82b, CL80b, FWG^{+}81, GDTS81, HTFH82, HKCM83, HMG80, KM82, KNBSF84, MS80b, MPW82, MBC^{+}84, PFM^{+}84, SWB84]. borders [GGW83]. both [AA83, BTF^{+}83, FJ82c, HA84a, LL82b, MKB82, ROB^{+}84, SPS^{+}84, SMG81, YKJF80]. botulinum [HH81b]. bound [AC81b, ABG81, Bar82b, CDLJ81, GN84, MR81b, Mec81a, Mec81b, MP84b, MP81, MP82, RKP^{+}82, VDT81]. boundary [Bro81, WLFD84]. Bovine [GCL83, Kla80, LRT^{+}82, AT83, CKJ83, GDH^{+}83, GZR^{+}80, GHGT81, GLMW84, ISGS81, KJBR82, LTBS80, LL82b, MGMS84, MMS84b, MM81a, Mil81b, MWM^{+}81, PG83b, PG83a, PPR82b, RLP84, SSG^{+}83a, WHD^{+}84, ZSR^{+}82, ZGN^{+}80]. BP80 [WWWC80]. Br [APZL80]. Brain [MPAM80, SSHP82, AMM84a, AMM84b, BV83, CGCS80, DCK84, GCL83, KJBR82, LRT^{+}82, MNP^{+}84, NGL83, SHWS83, SS84c, TZW83, TTI^{+}83, TM84, WWWC80, YSJ81]. brain-derived [TM84]. brains [MWM^{+}81]. branching [HTS80, TS83a]. breakdown [MLNK83, PBW80, SDT81]. bridge [DPS80, GK81d, GVK81]. bridging [LLS82, WM81]. Brij [DB83]. Brij-lysed [DB83]. broad [DKDB81]. bromide [GB82]. bromide-derived [GB82]. bromodeoxyuridine [O’B80, WA83]. brucei [OBC^{+}84]. Brush [HMG80, MS80b, MBC^{+}84, BP82, Bur82b, CL80b, FWG^{+}81, GDTS81, GGW83, HTFH82, HKCM83, KM82, KNBSF84, MPW82, PFM^{+}84, SWB84]. Brush-border [HMG80, MS80b, CL80b]. Bud [SBJ84, AP84, SHPH84]. budding [RBPS83]. buffering [LWWP84]. Bufo [SLIF80]. bull [LFR80]. bundle [BRH82, BT84, CL82, FK80, HF80, MB82b]. bundles [GGW83, OKE81, OM82, Wan84]. bundling [GN81, TM84]. bundling/depolymerizing [MB82b]. bungarotoxin [Bad81]. burst [NPG^{+}80]. butyrate [MGSC84, VSLS80]. butyrylcholinesterase [DK81a]. BW5147 [EK84].

C [YBG81, Can83, CD84, DSRF84, Gal81, HMKM82, LDF^{+}84, Moo81, ORA^{+}84, PGD^{+}84, RMS^{+}82, RMF83, SGA80, YBG81]. c-3T3 [FNS81]. C-fibers [Can83]. C-peptide [ORA^{+}84]. C-protein [DSRF84, RMS^{+}82, RMF83]. C-value [Gal81]. C3b [HSFM82]. C5b [BTJ82]. C5b-9 [BTJ82]. C6 [ASA83, HS83, HS81, WSM80]. C6-cells [HS83]. Ca [KKM84, BPM81, BYS82, BWM84, CH80a, FJ82c, GWS84, JSDM82, JSMT82, KDMS83, KM82, KJBR82, MB82b, MSP84, MA82, NJR80, PG84, PGD^{+}84, SW84, Sch84, SM83c, Wal82a, Wal82b, WH80, YAF81]. Caenorhabditis [CT82, CKE81, NRW82, RW82b, WAN81]. caffeine [YS84]. caffeine-treated [YS84]. cage [ZHM80]. calcification [GBF84, NvIvGdH^{+}82, PPR^{+}84, SKA^{+}83]. Calcitonin [DVB^{+}81, WGRB80]. Calcium [BFW80, BLHC81, ER84, GG80, LA80a, MR82, OYNT83, SS80a, SGSS^{+}81, TMW82, TPR82, UE83, WMG84, YS84, BHSA82, BL80a, BL83b, BHJV80, BP82, BMA84, CV82, EWG82, FT80, GG83, GS84a, Gil83, HH81b, KW84, Kie81, LWWP84, Mar80, MSB80, MGW^{+}80, NCBP84, Pen84, PTH83, PB84b, PATR82, SBSM84, SW84, SF84a, SBC82, SY83, TRN^{+}82, TRG84, UPLH84, WS84, WB82, WHJ80]. calcium- [CV82]. calcium-buffering [LWWP84]. Calcium-dependent [OYNT83, BL80a, MGW^{+}80, WS84]. Calcium-induced [GG80, WMG84, SY83]. Calcium-labile [SS80a]. Calcium-mediated [UE83]. calcium-modulated [SBSM84]. calcium-regulated [SF84a]. calcium-sequestering [Kie81]. Calcofluor [Her80]. calf [SMJB83, SS84c]. calf-brain [SS84c]. caliber [HGP84]. californica [KHS80b]. Calmodulin [BHJV80, DMB84, GGS83, GCF^{+}83, RN81, WS81a, BSP83, BCSC82, BTHL81, BWV84, CGS81, CBS83, GW80, GWP82, GCS81a, GCS81b, HOT81, HMG80, HM83a, KM82, KJBR82, LDM80, MDM^{+}81, NAK82, NAK83, PBPD83, Pen84, RZSW83, SCS80, SDG^{+}84, SSMH82, TRG84, TTI^{+}83, WWWC80, CGS81, KDMS83, MSP84]. calmodulin- [GCS81b]. calmodulin-activatable [GCS81a]. calmodulin-binding [BCSC82, BWV84, CBS83, RZSW83, WWWC80, CGS81]. calmodulin-dependent [Sch84]. Calmodulin-microtubule [DMB84]. Calmodulin-stimulated [GCF^{+}83]. calsequestrin [JSCM83, JC84]. calspectin [TTI^{+}83]. calvaria [ECGvdRP83, RUM^{+}82, SKA^{+}83]. CaM [WWWC80, BGE84, OFO83]. cAMP [BF82c, CL81, DMD80, DSD80b, DSD80a, FB82, FJ82c, FS84, IMS82, VV83, VS84]. cAMP-dependent [BF82c, FB82, FS84, IMS82]. cAMP-mediated [VV83]. can [CRGS82, EN84, GH82b, HA84a, KGAB81, LTE84, NCBP84, WKH80, YK84]. canalicular [MSRB84, RH83, RH84a, RH84b]. canaliculus [CHH^{+}83]. Canavanine [ML83]. cancer [GSPS82, NLN^{+}81, Wei83]. Candida [SM83b]. canine [MRHS81, MBP^{+}83, MSB^{+}83, PS83b, PAS84, PBS84, RIP^{+}84, RBPS83, RBPSB84, TOT^{+}84]. cannot [VSD84]. capacitation [SH82b]. capacity [LWWP84, LH84]. capillaries [RKF80, SSP82b]. Capillary [MW83a, GMJR82, MOV83, SSP81, SSSP81, SSP82a, SLS83]. capillary-like [MOV83]. capped [RSB84]. Capping [KPPC83, BM83, BNH81, CBP84, HRBJG80, HDRK83, NAK82, PATR82, SKWF84, TBR80]. capsule [FML83]. captured [CL81, OR81]. Carbohydrate [GGVR82, BHJ83, CRH81, CEC84, Hat81, RCH81, SR83a]. carbohydrate-binding [SR83a]. carbohydrate-derivatized [Hat81]. carbohydrate-reactive [CRH81, RCH81]. carbonate [FHFL82]. carboxyl [BGPB80]. carboxyl-methylase [BGPB80]. carboxylase [RBMM82]. carboxymethylcellulose [HWBC82]. carcinoma [ABG81, Bar82a, CMC81, GOA83, HBR81, HT83, Ioz84, JVMRK82, MGSC84, SLK^{+}84, WCZG84]. cardiac [CHI^{+}81, DANH84, Gla83a, GSS82b, PSC83, SGB^{+}81, TDS83b, Tok83, TS81, YRS83]. carolinensis [HT82a]. carrier [HBS^{+}81]. carteri [GK82]. Cartilage [HHS^{+}84, HS84, BL82b, DPK84, KPG^{+}82, KMP^{+}82, PST^{+}83, PPRR82, PPR^{+}82a, PPR^{+}84, STRP83, WR81]. cartilage-specific [STRP83]. casein [GRDR82]. castellanii [CBP84]. cat [DK81a, GBLH83]. catalytic [BF82c, FB82, MSD82b]. catalyze [KGAB81]. catalyzed [FC82]. catecholamine [SLRP82]. cathepsin [GPHV83]. cationized [VVM81]. cations [AW81, FNS81, GS83, GSMR83, PP80a, Ste82a]. caudatum [AF84, HHT80]. Caulobacter [SA82]. cause [GNW^{+}84, MHB84, NCBP84]. caused [KSR^{+}80]. causes [KDMS83, SSHP82, VC83, WB81b]. cavitation [KMCAS80]. CDC24 [SAP81]. cDNA [MIRS84]. Cell [AT83, Axe81, BZ83, BGS^{+}83, CG84, FW80, FTM82, GWK84, Hat81, HEA^{+}82, HLM^{+}84b, HS81, Ino81a, LSC83, LSJT83, MCN^{+}84, MTT80, MGLF83, MF84d, OFO83, Rao81, RF80a, RRD83, SBGD84, SSF81, SB83a, SA82, SB80c, SHFK83, TGPC82, TPHL80, WMH81, ZH82, ZLK81, AW81, AW82, AFW84a, APZL80, ASA83, AA83, AHMS82, ASG83, ASG84, BSP83, Bar82a, BZ84, BO81a, BKO81, BBLY83, BC84a, BC84b, BBWS83, BDP^{+}84, BCC^{+}83, BBB82, BC80, BW80, BN83, BB80c, BWW^{+}80, Bur82b, BK84, BZN82, BS81c, BKM82, Cab83, Can80, CMM81, CP83, CB82, CRH81, CARK81, CBS^{+}84, CGM81, CKJ83, CS82, COBC84, CL80a, CP82a, CH80d, CEC84, DK84a, DKDB81, DBK80, DB83, DWS84, DSKS84, DS80, DHS^{+}84, DBL82, EWG82, EW83, EKS^{+}82, EHESF82, FNNFS84, FFEL81, FS80]. cell [FSBR83, FSVW80, FSW^{+}83, FNS81, Fri82a, Fri82b, FCNS84, GDRS80, GDM^{+}82, GKKPB83, GB83, GJRO^{+}83, GAS82, GAR^{+}84, GF84, GBS80, GW80, GGVR82, GGK83, GDH82, GSH80, GHGT81, GSK83, GSMR83, GK81d, GK82, Gri80, GHH82, GE84, GGL80, HSB^{+}83, HWK^{+}84, HJ80, Har80a, HER81, HCP81, HEO82, HO84b, HBR^{+}84, HF80, HRBJG80, HCH^{+}81, HS80b, HK80b, Ioz84, ISGS81, IHJ82, JOA84, JOH^{+}84, JB81c, JB81d, JDTW80, JZ81, JHCC84, KSR^{+}80, KUM^{+}84, KSF84, KBD82, KWM82, KA84, KvRK^{+}84, KI83, LTE84, LA80a, LAAS84, LA83, LM84, LKJ84, LNB84, Mad83, MW83a, MW83b, MKD^{+}84, MBZB81, MSD82a, MMS^{+}84a, Mar80, MRHS81, MS84b, MNST82, MEKT84, MP83, MPJ80, MdV80, MPF83a, MLM83, MBC80a, MSB^{+}83, MKB81, MB82c, MBC80b, MFVPK82, MMAO83, MHNT81, MDA83, NRMB84]. cell [NS83a, NPG^{+}80, NW83, NCP84, NS83b, OB81, OYNT83, OCT81, Oli82, OWMB82, OMG82, PF83a, Pen83, Per81a, PWT80, POYF81, PWT82, RGB84, RCH81, RH81, Rau84, Reb82, Ric80, RHK82, RCJS82, RVH84, RKMF84, RBPF80, RUB84, RDY^{+}83, RGE83, SB80b, SWGB80, SBG80, SMS84b, SDL82, SSG^{+}83a, SMT80, SC82b, SF84a, SLKB83, SHW^{+}82b, Sin80, SCP^{+}83, SAP81, SB83b, SDGYJ84, SCM83, Sol81, SHB80, SS80c, SDM84, SBM84, SK83, SSC80, SMTT83, SR83a, SC84b, SWvK^{+}83, SR83b, SKA^{+}83, SH81b, TEL84, TOT^{+}84, Tam82, TT83, TFGC84, TS83b, TQG84, VKHH82, VS84, VY82, WH81, WWKW83, WYK^{+}84, WOM^{+}81, WFGB83, WHSK80, WC80, WG81, WSM^{+}82, WW83, WSLSJ83, WLCS84, YF81, ZF84, ZB82, ZSM^{+}83, dSK84, dHLC84, dLTvdSB83, BSS83b, JB82, PB81, LLB84, ORA^{+}84]. cell-axon [TQG84]. cell-contact [ASG83]. Cell-cycle [SSF81, BZN82, KBD82, Ric80, JB82]. cell-derived [GDRS80]. cell-free [LM84, POYF81, RUB84]. cell-mediated [BBLY83]. cell-protein [PF83a]. Cell-specific [MF84d]. Cell-substrate [Axe81, GAS82]. Cell-substratum [CG84, MCN^{+}84, MTT80, BN83, CS82, DKDB81, Gri80, WYK^{+}84]. cell-surface [LA83, SR83a]. cell-to-cell [MBZB81]. Cell-type-related [SB83a]. cells [AMCG83, AT83, AMS82, ASWR83, ALE84, AdS84, AKM^{+}82, AUT83, ABG81, AH84, APZL80, AA83, AFK^{+}84, AMJ^{+}81, Atk81, ABSC81, BJS81, BSH83a, BGD80, BE82, BHJ83, BBR83, Bec84, BHA82, BZ83, BZ84, BKB80, BS83, BGNT81, BAWL81, BB82b, BPM81, BG82, BC82, BL83a, BV83, BMF84, BMH^{+}83, BM80, BPM80b, BCC^{+}83, BNH81, BBB82, BGE84, BC80, BFL^{+}83, BS80, BPB^{+}81, BB80b, Bre83, BLHC81, BN83, BNGN84, BR84b, BCP^{+}81, BLS^{+}80, BBS81, BB80d, BWM84, BvdMN84, BZN82, BF82b, BGS81, BF82c, CWBB83, Can80, CMM81, CTKJ83, CSS80, CB81, CEC^{+}83, CBW81, CCJC82, CARK81, CKB^{+}80a, CLHA80, CLFC83, CGM81, CVGM83, CPS84, CH80c, CMC83, CNSW80, CMC81, CD84, CMF82, CWS^{+}80, CFCS82, CFW84, CF83, CGG84, Con84, CH80d, CK82, CL81, CFML83]. cells [DKDB81, DF84, DK84b, DB81a, DVB^{+}81, DBG^{+}81, DMB84, DH81, DTT81, DCS84, DWM83, DBBN81, EKSZ80, EZ84, EPM84, EK84, EF84, EMC83, ELS^{+}83, EPRS81, EPE^{+}80, EM81, ERB83, EHK80, FSB81, FN80, FFRF83, FHDL82, FC82, FDKF82, FP81b, FBP82, Fis82c, FB82, FDSWW84, FSF^{+}80, FHG^{+}81, FSW^{+}83, FJ82a, Fri83, Fuk80, FPFP81, GGK82, GK84, GSPS82, GDTS81, GRB82, GSS^{+}82a, GSG82, GOSW82, Gla83b, GGW83, GSB81, GNW^{+}84, GN84, GGV81, Goo83, GWP82, GMO84, GHGT81, GCL83, GLMW84, GMSK81, GSK83, GV82, GL82, GKZ81, GMRB81, GH82b, GLS83, GDRS83, GQW83, Gri80, GMJR82, GOA83, GE84, GS80b, GLBS81, GAAG82, GS84f, GE80, HSB^{+}83, HH83a, HR80, HL81, Hat81, HLM84a, HA84a, HER81, HOMR82, HBS^{+}82, HBR81, HO84a, HT80, HKSF80, HH83b, Hep80]. cells [Her83, HHSS81, HHS^{+}81, HH81a, HT82b, HTFH82, HLM^{+}84b, HT83, HA80, HS81, HM83a, HSC80, HC83, HEG81, HPG83, HF83, IRMM84, IY84a, Ioz84, IDF83, ISGS81, IJ82a, IJ82b, IWM83, JD82, JB81a, JB81b, JWG84, JWBC81, JVMRK82, JSC84, JHCC84, KUM^{+}84, KCM83, KSMF84, KWSB84, KRB84a, KOLV81, Kho82, KBD82, Kla80, KKM81, KML83, KBDK84, KSK^{+}82, KKM84, KS84, KB81a, KB81b, LSM83, LdMM^{+}84, LC81, LK83, LFAG84, LTBS80, LVL^{+}82, Law83, LPB84, LBJ84, LL82b, LW81, LNA81, LBM82, LCK84, LQ82, LS81b, LA83, LNB84, MHSW81, MKW^{+}82, ME81, MB84a, MW83b, MB84c, MKD^{+}84, MBWH83, MTKH84, MBP^{+}83, MS84b, MPJ80, MSGP81, MG83a, MF84b, MLP84, MMC83, MGSC84, MG81, MHRR83, MA83a, MMLT84, MB81, MR80b, MR81b, Mec81a, Mec81b, MKK^{+}82, MCPB80]. cells [MHB84, MFP82, MB82c, MMJ^{+}81, MA82, MWA^{+}80, MOV83, MMR^{+}84, ML83, MGK83, MGW^{+}80, MC80, MDJ82, MDSH80, MKD80, MCG82, MPC84, MSD82b, NSI80, NSI82, NLN^{+}81, NCL84, NCBP84, NvIvGdH^{+}82, NSW82, O’B80, OLHH81, OYNT83, OGS82, OLH^{+}81, Oli82, Olm81, OBH84, OM82, PF83a, PLP81, PBPD83, PBFC^{+}84, PE83, PS83b, PAS84, PBS84, PPSP83, PB84a, PR83, PBWK83, Pie82, PTH83, PT83, PWT82, PGD^{+}84, PFP^{+}82, QGW83, RS81, RRRS82, RSP83b, RGB84, RMABB82, Rau84, RUM^{+}82, RIP^{+}84, RPM83, ROB^{+}84, RBPS83, RBPSB84, RRD83, RMJ83, RB82b, Rot84, RUB84, RRS82, RGE83, SCS80, SSL^{+}84a, SO80, SS80b, SSJ83, SP82a, SHPH84, SKN83, SSL^{+}84b, SHHR84, SKS^{+}84a, SW84, SNPE84, SG80, SGW^{+}82, Sch83a, Sch83b, SFS81, SSE84, SG82a, SBCS^{+}82, SBF84, SMTP80, SM82b, SSF81]. cells [SS84a, SR81a, SR81b, SW81a, SB83a, Sma81, SLS81, SB82c, SN80, SBSL84, SDK^{+}84, SSC80, SD84, Ste83a, SLL80, SVW^{+}83, SLRP82, SCM84, SCSC81, Sur82, TTGA82, TI82a, TDM80b, TS83b, TEDS83, TT84b, TWWG84, TLPTV83, TIS^{+}81, TQS84, TKM83, UAL84, VS82, VPMF84, Van83, VSBL83, VSLS80, VEL80, VBPL84, VRR80, VH82, VY82, VVBZ84, WS83, WOM82, WM84, WB84, WR83, WWKW83, WYK^{+}84, WHSK80, WC80, Wat84, WSM^{+}82, WP82a, WGF83, WCZG84, WHD^{+}84, WSO^{+}81, WD83, WH83, WWDYW83, Wri84b, YKJF80, YAF81, YDL^{+}84, ZGJ^{+}84, ZSE80, dSTK81, dLTvdSB83, vMS83, HS83]. Cellular [OCA^{+}84, SKE84, BVAR83, BLV84, DSKS84, GDM^{+}82, HMDP80, KUM^{+}84, KKG84, LOC81, MPH82, MR80a, MBC^{+}84, PWT80, Rot84, SW84, SHJA82, SAP81, TWWG84, TSGS83, YHS^{+}82, YF83]. cellulose [GBS80, HWBC82, MB80b]. center [EDP82, FP81a, VF82, VA83]. centers [GRB82, GAR^{+}84, GMSK81, GSK83, MJ83, NJSC83, SKN83, Tuc84, WOBR80]. Central [AHPL81, CG84, DJ81, DCG83, DHL84, LOL81, MER^{+}83, MS84a, OWPC82, OK80, SW81b]. Central-pair [AHPL81, LOL81, OK80]. Centrifugation [SL82b, KMCAS80, LL82a]. centriolar [KC80b]. Centriole [KRB84a, KB81a, NJSC83, KCB84, KB81b, WWKW83]. Centrioles [VY82, CN80, KC80b, OOP^{+}81, RHK82]. Centripetal [RW82b]. centromere [MHNT81]. centromeres [BAC^{+}84]. centrosomal [SM82b, SM82a]. centrosome [KKMK84]. centrosomes [BKB80, KNHK84, KKMK84, KNK84, KCB84, KB81b]. cerebellar [CBL84, HL81, Hat81, HLM84a, LR83a]. cerebellum [FYC82, HM84e, LDM80]. cerebral [CKJ83, DWM83, MdV80]. Cerebroside [TFAP81, TZW83]. cerevisiae [FS80, AP84, Har80a, IY84b, IY84c, KHL82, MJM83, MB84d, NLPH82, RF80a, RF80b, SBJ84, SAP81, WH82]. certain [GCH84, LK84]. cervical [DK81a, GSPS82]. CF0 [MS83b]. CF0-CF1 [MS83b]. CF1 [MS83b]. CFU [NW83]. cGMP [SSG80, VV83, VV84]. cGMP-dependent [VV84]. cGMP-specific [VV84]. chabaudi [TMW82]. chabaudi-infected [TMW82]. chain [BMF82, BNH81, BNB^{+}82, COS83, EH81, KWSB84, PKS81, SDG^{+}84, WB81b, WA83, ZHTH80]. chains [BTF^{+}83, BHP83, KWSB84, KH84, Nac81, SP84a, Wri81]. Change [FSW^{+}83, DTBF82, HSB^{+}83, HFW82, PF81, Sch81b]. Changes [AFW84a, APZL80, BG82, BvGJ83, CBS^{+}84, GBE84, GRZ83, NSI80, NSI82, SSZ83, SBCS^{+}82, SM83a, SW81a, BSR84, BSH83b, BGE84, Bro81, BWM84, CGG81, CGG84, FDSWW84, GHY81, GDMG81, GGV81, GK81c, HH83a, HT82a, HR81b, HS81, JRM^{+}80, JTM80, KKM84, KRB^{+}84c, LR81, LSJT83, LP83b, LN82, LWWP84, LLB84, MR82, SJ80, SGSS^{+}81, SA83a, TTUI83, UE83, WWPL84, WMP^{+}84a, WN81, WP82a, WD84, WHJ83, IT82]. channel [EMA83]. channels [SSP84b]. Chaos [HT82a]. Chara [NBSW81, NSW82, PL84]. characean [NW82]. character [PR83]. characteristic [HSS82]. characteristics [FSVW80, GH82b, KPG^{+}82, KR82b, KJBR82]. Characterization [CVT84, CAL^{+}82, FQBN81, GHWR80, GB80, HE81, HM83a, LFGR82, MKG81, MLD82, MA82, RMF83, RB80, SDL82, Sow84, WOM^{+}81, ZRW80, AC82, ASG84, BO80, BSS83a, CGCS80, CMC83, CNSW80, CD84, CKE81, Den80, EF84, FKB83, FBB82, FWG^{+}81, GWB82, GPR83, HJ80, IRMM84, IHJ82, KHF81, KSF84, KDL80, KMV82, LFS^{+}81, ML84a, MB80a, MAG82, MB82b, MA83a, MD80a, MB82c, Mil81b, MM82, NSI82, NKH81, NW83, OBC^{+}84, OM82, PST^{+}83, PG83b, PFM^{+}84, PKS81, QCB84, RSL81, SBSM84, SBC81, SP82b, SG80, SCS84, SHJA82, SSSP81, SSP84a, SDK^{+}84, SBM84, SG84b, SR84b, SIM^{+}84, TTGA82, VKHH82, WWS83, WTK83, ZGJ^{+}84, ZHM80, CHH^{+}83, KC80a, MP82, SWGB80, WH83, ZGN^{+}80]. charge [BF84a, GF84]. chase [LNB84]. Chediak [IBB^{+}81, NFG82]. chelator [LTE84]. Chemical [SR82a, FCKP84, MBC80b]. chemically [FHDL82, RRRS82]. chemoattractant [NFS81, SPEL84]. chemoattractants [PS81]. Chemoattraction [VSD84]. chemokinesis [BG83]. Chemokinetic [WLFD84]. Chemotactic [BSKFW83, HM84d, SGH^{+}83, SZ80, WWPL84, HFW82, JNS^{+}84, MGMS84, SGM^{+}84, SGS^{+}82, SDZ84, Van83, Zig81, ZSL82]. chemotaxis [BHSA82, BG83, GSMHF84, VSD84, ZLK81]. Chick [CF84a, CF84b, OT80b, ALE84, BS82a, BL82b, BF84b, CG84, CDvdM80, CH82, CRGS82, GRR^{+}84, GBF84, GT84, HT82b, LSM83, LSBPG84, Let83, LZR82, MMKL81, MU82, NDRL80, PST^{+}83, Pie82, PF83b, SLR^{+}83, SG82a, SMTP80, TRN^{+}82, TMS84, WG82a, WBG83, Wri81, WA83]. chick-rat [SLR^{+}83]. chicken [BR83, BR84a, BHR81, BB82a, BB82b, BGE84, BWV84, DSRF84, EM84, FGS^{+}82, GDTS81, GAS82, GBE84, HC84, KLC83, LMYM84, LFS^{+}83, MWI^{+}84, ML83, MF84d, OMM82, PL83c, RCM^{+}83, RMS^{+}82, RMF83, SHJA82, SDG^{+}84, SPHE80, TDS83c, TDS83b, Tok83, TS81, Woo80]. chimera [PB84a]. Chinese [AMCG83, AMS82, BHA82, BKB80, BS83, BS80, Cab83, CWBB83, CPS84, CK82, ERB83, HJ80, KRB84a, KvRK^{+}84, KB81a, KB81b, LBM82, MMS^{+}84a, RPM83, ROB^{+}84, SSC80, WC80, ZHTH80]. Chironomus [KB81c]. chitin [Her80, MBC80b]. Chlamydomonas [LR83b, BFW80, HG80c, MLSR83, OO82, SM80, SCM83, AHPL81, BSR84, BM82c, BL83b, BLH82, DGC83, DHL84, GO80, GW80, GDH82, Goo83, HMKM82, HW83, HWJW84, HPRL81, JB81d, JB82, JR80, KW84, KSSR84, KJ82, MHCG80, PHRL81, RW82a, SRGB83, SMGB84, SHRL84, SBC82, SS80c, WOBR80, WD84, WCJ83]. chloride [EDK80, GCKAA83]. chloride-secreting [EDK80]. chlorophyll [DRF84, KTW83, MS80a, SBG^{+}81, ZP80]. chlorophyll-protein [DRF84]. Chloroplast [GO80, SMGB84, MS80a, OKA84, OO82, SRGB83, SA83a]. Chloroplast-cytoplasmic [GO80]. Chloroplasts [Bog81, RZSW83, RBMM82, ZP80]. Chloroquine [GNGTS80, BCF84]. chlorotetracycline [WHJ80]. cholecystokinin [RMJ83, RMJ84]. cholecystokinin-binding [RMJ83, RMJ84]. Cholera [KGAB81, Fis82c, KPPC83]. Cholesterol [CNH80, AC83, HDRK83, HSCS80, MFA83, SHH82, Sin80, SBBS84]. cholesteryl [FMS^{+}84, MA83a]. choline [JRM^{+}80]. cholinergic [CK80, JRM^{+}80, JRB80, MMIGK82, MTM83, PJM^{+}81]. chondroblasts [PST^{+}83]. chondrocalcin [PPR^{+}84]. chondrocyte [GSG82, HHS^{+}84]. chondrocytes [DPK84, DVvdMvdM81, GBF84, GT84, KPG^{+}82, KMP^{+}82, MTT80, MBLvdM84, RUM^{+}82]. chondrogenesis [ZS84]. chondrogenic [SHPH84]. Chondroitin [HCJV83, AMM84a, AMM84b, LH83]. choriogonadotropin [Asc84b, LA83]. chorionic [HBB82, HHPB83]. choroid [VVM81]. chromaffin [AT83, Cre81, HH83a, PBFC^{+}84, PBWK83, SLRP82]. Chromatin [BAC^{+}84, AW81, ACH^{+}81, AHRG82, ARHG84, AAW^{+}84, Bal82, BBWS83, DBBN81, HK80b, KNK84, LP83a, dPMMF^{+}81, RSDH82, SC81, WFR84, ZF84]. chromatins [GC80]. chromatography [PK81, Sta81]. chromatography-mass [Sta81]. Chromatophores [MP84a, Ste84a]. chromosomal [EH81, EPO^{+}84, FDSWW84, GBL81, HAS^{+}81, KB81c, RPHA81, WRK81]. Chromosome [Gal81, MB81, PHST84, AB81, CMM81, EL83, EHCR84, GRDR82, KHS82, KMI82, KLD^{+}82, KW81a, Kub82, Kuo82, PL83a, WWKW83, WH82]. chromosomes [BKB80, CPLC80, CLJ^{+}81, EL83, HR80, HMB^{+}82, KB81c, LP83a, LBM82, Mil81a, PL83a, Sch81a]. chronic [TQG84]. chymotryptic [PKS81]. CII [MBLvdM84]. Cilia [Gib81, HR81a, ARL^{+}80, Den80, DPS80, LN81, MKG81, MW80, OK80, RR80b, SSA80, SSMH82, TTUI83, WZ80]. Ciliary [TT81, ARL^{+}80, BHJV80, Den80, ERN83, LN81, MKG81, RSP^{+}83a, Ste83b]. ciliate [Yao82]. ciliated [GWP82]. cilium [CSHP84]. circulation [LHHH82]. circulatory [LR81]. circumferential [OKE81, OM82]. cis [Cre81]. cis-Unsaturated [Cre81]. cisternae [CC82, GBB^{+}82, GQW83, RB82b, SSCF84, YS84]. Claims [BRS81]. clam [Mur80, SR84b]. clamp [JGSH82]. class [Sch81a, SP84a, SNdC^{+}84]. classes [NS83a]. Classification [EBS84]. Clathrin [GL81, AW82, ATW83, BSSR84, BHP83, CP81, KH84, MTKH84, MP83, OHA^{+}84, PSR82, RKMF84, RBPF80, SSBR84, SBBS84, ZK83]. clathrin-coated [MTKH84, MP83, OHA^{+}84, RBPF80]. clear [SCD82]. Cleavage [GK81d, CMM81, HOW80, KOB82, KFV83, LS81a, NDC80, WMW84, WN81]. cleaving [KRN83, SS80b]. clofenapate [RRRS82]. clofibrate [RFML84]. clofibrate-treated [RFML84]. Clogging [KR82a]. Clonal [WG82b, GSH80, KBD82, SW84, SMT80, SKA^{+}83]. clone [SL82a]. cloned [CSK^{+}80, GKKPB83, KSG81, KKBG83, MIRS84]. clones [AHMS82, GDM^{+}82]. clots [CGW82]. cluster [Pen83, PP84]. clustered [CB82]. clustering [MHB84]. clusters [Blo84, BN83, BMA84, Con84, Pen84, PB83b, RBPS83]. CNS [BNM82, CKFS82, NBM82]. Co [AL84, CGA^{+}82, ELS^{+}83, HRBJG80, KPPC83, RKP^{+}82, SS84a, SKKC84]. co- [RKP^{+}82]. co-capping [HRBJG80, KPPC83]. co-cultured [SS84a]. co-distribution [SKKC84]. Co-localization [CGA^{+}82]. Co-migration [ELS^{+}83]. Co-translational [AL84]. CO2 [SG82a]. Coa [RFML84, RQHR81]. Coalescence [PBG80]. coat [GT84, PL84, RKMF84, SBBS84]. Coated [MFP82, PSBF83, BF84b, CGA^{+}82, CRH81, CBWW80, CGG84, Gri80, GML^{+}80, HSFM82, HJ80, Heu80, KC83, KPW^{+}83, MTKH84, MP83, MFA83, MB84d, MDA83, OHA^{+}84, PSR82, PL84, PvD83, PK81, PDK83, RBPF80, RFLR81, SCS80, SBBS84, SHWS83, VKE^{+}81, WLFD84, WP82b, ZK83]. coats [BSSR84, CP81, SSBR84, ZK83]. cobalamin [LAAS84]. cochlea [TS83b, TEDS83]. cochlear [TDM80b]. Codistribution [HOMR82, RMAF80, RB82b]. coefficient [SSL^{+}84a]. colcemid [ERB83]. Colchicine [WMP^{+}84b, BRR^{+}83, BG82, BBB82, BMCGO^{+}82, MF84c, PE83, PVPB82, SMH84]. Colchicine-binding [WMP^{+}84b]. cold [BHA82, BTL84, Gri80, JL84, SB80a, SEM80]. cold-fracture [SEM80]. cold-recovering [JL84]. cold-stable [SB80a]. cold-treated [JL84]. coli [SR82a]. Collagen [MMAO83, BT84, BF84a, BWW^{+}80, CEC^{+}83, CGM81, CCKMH82, CDvdM80, DB81a, DWS84, ELST81, FMS^{+}84, FGS^{+}82, FML83, GSG82, GBF84, Gol82, GB82, GH82b, HJ80, KNH80, KMP^{+}82, LLM82, LBJ84, LFS^{+}83, LGF^{+}84, ML84a, MHGM82, MKB82, MBLvdM84, MOV83, MRB84, PPRR82, RMAF80, SSZ83, SMJB83, San82, SAW83, SB82a, SB84, SGPB84, SK83, SSCP80, TH84, TSGS83]. collagen-binding [KNH80]. Collagenase [AFW84b, AFW84a, GMJR82, JWG84]. collagenous [GJRO^{+}83, KKM81, LPB84]. collagens [MWI^{+}84]. collapse [BMF84]. collar [TQS84]. colloidal [DWP81, GSvdLS81, HLSWH82, Law84, dHLC84, SG81]. colon [Ioz84]. colony [BGS81, GS80b, VC83]. colony-stimulating [BGS81, GS80b, VC83]. colostrum [Kla80]. columnar [UAL84]. comb [TT81]. combination [TOK^{+}84]. combined [KOS83, Sta81]. comes [Per81b]. Commitment [SMTP80, BLHC81]. common [GLP80, OYNT83]. communicating [BDH82]. Communication [WLCS84, MKK^{+}82, MYR81]. Compaction [HCMK81]. Comparative [HAFG82, WRK81, AUT83, WOBR80]. compared [Let83, NGL83, NDC80]. Comparison [BWW^{+}80, CGG84, CL80b, GLMW84, IJ82b, NDS84, RLP84, BLH82, BB82c, CMM81, DCS84, FFS^{+}82, GCL83, JKHT81, LRT^{+}82, PST^{+}83, PG84, TSGS83, TAB80, MS80b]. compartment [MK84a, MPU84, OHA^{+}84, VN82]. Compartmentalization [TdS84, HSRL81, VEL80]. Compartmentation [DR83, BAWL81, SJC80]. compartments [BT84, BPM80b, BOH81, GN84, ML84a, Rot84, SSO84]. competence [SCP^{+}83]. competitive [BDP^{+}84]. complement [BTJ82, SHW^{+}82a, TJB83, WLCS84]. complementary [CSK^{+}80, DCK84, GKKPB83, KKBG83, WB83a]. complementation [HFSK82, HSTK83]. complete [CKB^{+}80b]. completion [FNHSS84]. complex [AHPL81, Asc84b, BW82, BTJ82, BM80, BNB^{+}82, Bro81, CVGM83, CF84b, DRF84, Deu84, ERH82, FP81a, GL81, GOKK82, GWS84, GIKR84, GQW83, GR81, HM84c, KTW83, LOC81, LK84, LRW82, MS80a, MSP84, MB80b, Mur84a, Mur84b, PE83, PBS84, PPSP83, QGW83, RMU84, SBG^{+}81, SSE84, SDGYJ84, SG84b, TAC84, TIS^{+}81, UM82, WTSAC^{+}81, WLFD84, WKAH84]. complex-coated [WLFD84]. complex-enriched [SG84b]. complexed [Pfe84]. complexes [BW81, BN83, CD84, FMS^{+}84, FBB82, GBS80, IMB83, JNS^{+}84, JOM80, KPPC83, KC80b, MK84b, MFA83, MP84b, MS83b, NB81, PM83, SBGP84, TT83, TD82]. complexity [MR81b, MP81]. component [BSV84, BHSC83, Bre83, GL81, HMG80, KPW^{+}83, LLS81, MD80a, MD80b, MB80b, ONO82, PSC83, RN81, SHP83]. Components [GNW^{+}84, SP84c, SBGP84, SSP84b, BL80b, BH83, BFL^{+}83, DYP83, DB80, ES82, FFGC84, GRR^{+}84, GRL82, GRZ83, IMB83, KK82, LC81, LOC81, MW83a, Mas84, MIS^{+}80, OLHH81, PK81, PDK83, PHRL81, PT83, RW82b, RH82, SHHR84, SvB81, SBCS^{+}82, SH81a, SB83a, SB83b, SDK^{+}84, SA83a, SS84b, VRR80, WMG84]. compose [CB84a]. composed [FBB82]. composition [AHPL81, FWP84, FFS^{+}82, IBB^{+}81, KNBSF84, LD81, MSUC80, MGG83, MSC83, NGL83, PHRL81, Sin80, SD80, CBJ^{+}82]. compounds [SNdC^{+}84]. Concanavalin [PG82, Fis82a, JHCC84, SCD82, dSTK81]. concentration [BWM84, Gol82, NDC80, WS84]. Concentrations [KST84, BYW81, GL80b, TP82]. concept [KKHK82]. Concomitant [HER81, WD83]. condensed [AW81, HR80]. conditional [MMS^{+}84a]. conditional-lethal [MMS^{+}84a]. conditions [CT83, MLMS81, PSSS82]. conductance [GCKAA83, SNdC^{+}84]. cone [ALE84, BM82a, BBF84, BS82d, PB84b, TJ80]. cone-like [ALE84]. cones [BSNP82, Let83, PCB83]. confers [BHJV80]. configuration [BZ83, GS84c]. configuration-related [BZ83]. conformation [BAC^{+}84, LGF^{+}84, PNYO83, WWPL84]. conformation-dependent [LGF^{+}84]. Congo [Her80]. conjugate [HSS82]. conjugates [HAEC81, OGS82]. conjugation [BR84a, HHT80]. conjunctival [THT^{+}84]. Connectin [MSK^{+}84]. connecting [CSHP84]. Connective [RGM^{+}80, JWG84]. connexon [GS84c]. consequence [HB84]. Consequences [CSR^{+}81, Zig81, LL81, ZLK81]. conservation [WNB82]. consistency [GK81c]. consistent [FA83]. consisting [CF84b]. constants [PM81]. constituent [CC82]. constituents [GAS82, MTG^{+}82]. constitutive [IY84b]. constraints [TWW82]. constriction [CMM81]. Construction [RC83]. contact [AT83, ASG83, BZ84, BN83, COBC84, CP82a, GRB82, GAR^{+}84, MSGP81, WYK^{+}84]. contact-mediated [AT83]. contacts [Axe81, BC84a, CS82, CW80, GAS82, Sin82]. contain [BRS81, CH82, GCKAA83, MACK84, PPR^{+}82a]. containing [ACTK84, AC82, CQB84, HOW80, IDF83, KKMK84, KHS82, KS82, KRB^{+}84c, LSBPG84, LBM82, MACK84, MP81, MP82, PD80, SB83a, Ste83b, TKM83, WPKB83, ZF84, Rag81, WH80]. content [BSS83a, BSS83b, BBS81, FBRP84, GS84a, HP82a, LKR^{+}83, RLP84]. contents [AR81, AMS82, CLHA80, SSLB84, YS84]. continually [TT83]. continuity [CKB^{+}80b]. Continuous [CFCS82, WSM80, Cab83]. contract [SSCP80]. Contractile [TWH80, TBR80, Bur82b, CHI^{+}81, FAP81, GK81c, Kan80, Kan83, NSI80, NSI82, Pol81, SBSM84, SMS84a, SMS84b]. contractility [HKCM83, OKE81]. contraction [Bur82b, BSNP82, CTKJ83, CGW82, CS83, KM82, LA80a, OB82, PCB83, PB84b, PNYO83, WC84, vDB80]. contractions [BS82a]. contrast [Ino81b]. contrasting [Sea82]. contribute [KMI82]. Contribution [Sto84, FBST84]. Control [AS81a, ASA83, BCLT^{+}80, HGP84, LTAP^{+}81, MKB81, NAK83, SB83b, Wri84a, AA83, BZ83, BZ84, BFW80, BO81b, CH83b, FAP81, Har80a, KW84, KSI83, KJ82, MLZD83, OT80a, SSJS84, SHW^{+}82b, SG84a, SBVW82, SC84b]. controlling [GHGT81, LA80b]. controls [MCG82]. Convergent [PR83]. conversion [KNK84]. converting [FQBN81, NDS84]. convoluta [MTI80, MTM83]. cooperation [GDL80]. Coordinate [ME81, HOMR82, SKS84b]. coordinated [SNPE84]. coordination [CH80d]. coprophila [Kub82]. corallina [PL84]. cord [RGE83, SFB^{+}84]. core [AHRG82, BP82, DGC83, ERH82, GB81, HMG80, MB82b, MS80b, OLHH81, PB83a, PM83, ST83]. cores [PM81]. cornea [CDvdM80, LFS^{+}83, NNA^{+}83]. corneal [BT84, FFGC84, GLMW84, THT^{+}84]. correlate [AFW84a]. correlated [HOR84, PF83b, WG81]. correlates [TQ82]. Correlation [JB81a, PA83, SK83, Sur82, WD84, YHS^{+}82, BAF84, GMO84, LWWP84, SA83b]. Correlative [JRM^{+}80]. correspondence [RUB84]. cortex [BRH82, BSH83b, CKJ83, CBL84]. Cortical [VKHH82, KMV82, RF84, SL82b]. cortices [BC84a]. cortisol [NDRL80]. cortisol-induced [NDRL80]. cotransfer [WHSK80]. cotyledons [CHS82, CHCS82]. could [RW82b]. coupled [BSSR84, KB81b]. Coupling [KS82, SHW^{+}82b, NW82, PCOK82, SS81]. course [Gla83a, MNW81]. covalent [FG80b]. covalently [SC82a]. crane [JL84, SLL80, SL82b]. crane-fly [SLL80, SL82b]. cranial [DJ81]. crassa [Man82, SHZN83]. crawling [DH81]. crayfish [Fri83, dCSB84]. creatine [EKSZ80, FM80, WDE83]. crenating [AS81a]. crest [BDP^{+}84, BF84a, RDY^{+}83, SR83b]. crinophagic [ORA^{+}84]. critical [ATW83]. critical-point-dried [ATW83]. Cross [Hir82, CV82, CBP84, DPS80, EP80, EPRS81, FT80, GBLH83, GGW83, GV82, HSC84, JGSH82, MB82a, MGW^{+}80, OCT81, RBH^{+}84, SHT^{+}84, SPHE80, WM81]. cross- [GV82]. cross-bridging [WM81]. cross-filaments [MB82a]. cross-link [FT80]. Cross-linker [Hir82]. cross-linking [DPS80, EP80, MGW^{+}80]. cross-links [CV82, GGW83, HSC84]. cross-reacting [SHT^{+}84]. cross-reactivity [CBP84]. cross-reacts [RBH^{+}84]. cross-reinnervated [GBLH83]. cross-sectional [OCT81]. cross-species [JGSH82]. cross-striated [EPRS81, SPHE80]. cruzi [Wil84]. cryomicrodissection [HM84a]. Crystalline [AFI^{+}81, CH83a]. crystallization [MU82]. crystals [BU84, CRC^{+}82, GR80, KTW83, MJ83]. CsCl [LL82a]. CSF [BGS81, VC83]. CSF-1 [BGS81, VC83]. ctenophore [TT81, TT84a]. culture [APZL80, Bar82a, BGD80, Bec84, BCC^{+}83, Can80, CVGM83, CW80, DBK80, DVvdMvdM81, ECGvdRP83, EPM84, EM84, FP81b, FM80, GW82, GBF84, GHGT81, HSB^{+}83, HPM^{+}83, Hat81, Kho82, Kla80, KPG^{+}82, LDV83, MLMS81, MKK^{+}82, MWA^{+}80, MDSH80, NS83a, NW80, OLHH81, OLH^{+}81, PLP81, PTH83, PT83, PFP^{+}82, RMABB82, RUM^{+}82, RE80, RGE83, RNH^{+}83, SMJB83, SB80b, ST83, TLPTV83, VRR80, WB84, WP82a, WH83, WSM80, YKJF80]. Cultured [CARK81, FCNS84, LN82, LL82b, AT83, ABSC81, BJS81, BS82a, BSG81, BW81, BLG^{+}80, BHA82, BZ83, BZ84, BAWL81, Blo84, BV83, BFL^{+}83, BN83, BR84b, BCP^{+}81, BLS^{+}80, BMA84, CSS80, CB81, CEC^{+}83, CVGO81, CCCR81, CS82, CMF82, Con84, CH80d, CS83, DBG^{+}81, DMB84, DANH84, EKSZ80, Fis82c, GL80a, GDMG81, GM84, GSG82, GPHV83, GNW^{+}84, GDRS83, GMJR82, GAAG82, HG80a, HH83a, HCJV83, HCH^{+}81, IYVV83, IDF83, JRM^{+}80, JRB80, KSSP83, KML83, KBDK84, LdMM^{+}84, LFAG84, LBJ84, Let83, LCK84, LQ82, LA83, LNB84, MER^{+}83, MCN^{+}84, MBWH83, MPMBC80, MS84b, Max82, MMC83, MA83a, MLHM84, MIS^{+}80, MMR^{+}84, MDJ82, MSC83, MSD82b, NHF^{+}83, NNA^{+}83, NNRR80, OLH80, PST^{+}83, PLMS81, PvD83, PR83, PBWK83, PF83b, RRS82, SCS80, SSL^{+}84b, SDT81, SNPE84]. cultured [SG80, SSG^{+}83a, SBCS^{+}82, SS84a, SR81b, Sma81, SG82b, SD84, SBVW82, SA83b, SLRP82, SCM84, TI82a, VEL80, WBPO80, WC80, WMG84, Wat84, WG82b, WHK^{+}84, WLCS84, YDL^{+}84, ZGJ^{+}84, ZFP84]. cultures [ALE84, ASA83, BKT84, BJG83, BWW^{+}80, CAL^{+}82, IRMM84, LS82, MW83a, MdV80, Pen83, PL83b, RF80a, RGM^{+}80, SGKD83, SY80, SHFK83, WG81, YSJ81, ZS84, ZH82]. current [Cle84, FCNS84, GL80a, KCGH84]. current-voltage [GL80a]. currents [KRN83, KCGH84, SCM84]. Cuticle [CKE81]. cuticular [TEDS83]. cutting [BM83]. Cyanobacterial [WGG80, GBG82, GWG82]. cyanobacterium [NBBS83]. cyanogen [GB82]. cyclase [CVGM83, HS81, TD83, VKHH82]. cycle [BO80, BBWS83, BC80, BZN82, CBS^{+}84, CKJ83, CRGS82, FSW^{+}83, FNS81, GWK84, HS81, JB81c, JB81d, KBD82, KA84, KB81a, KB81b, MA83a, PBWK83, Ric80, RF80a, SDL82, SBGD84, SHW^{+}82b, SSF81, SAP81, SB83b, SS80c, VY82, ZSM^{+}83, dLTvdSB83, JB82]. cycled [WSM^{+}82]. Cyclic [DMD80, DSD80b, DSD80a, Ste84b, TD83, AM81, BLS^{+}80, CGG84, DK81b, FTM82, GSMHF84, GCS81a, GLBS81, LLS81, MR80a, MMC83, MSP84, MKB81, MF84e, MSD82b, NKT80, OB82, OT80a, OBH84, PB84b, Ste83a, VVP82, VSD84, WR83]. cycling [CH80d, RS81]. cycloheximide [HAF82, MLNK83]. cycloheximide-arrested [MLNK83]. Cytidine [Sch83a, Sch83b]. cytidine-dependent [Sch83b]. Cytochalasin [SM81, BS81a, ML84b, MR80b, NBSW81, PWM81, POYF81, Sch82, WG82a]. Cytochalasins [LTGL80, PWM81, YHS^{+}82]. Cytochemical [Rot84, BF82c, FB82]. cytochemistry [CVAK80, CQB84, JRM^{+}80, dSTK81, dST82]. cytochrome [BM80, BHSC83, MMO^{+}81, MS83b, RVB80, RCJS82, Sta81, BPM80b, ZP80, MCPB80]. cytodifferentiation [BBR83]. cytoimmunochemical [RPHA81]. cytokeratin [FSW^{+}83, MFVPK82, SKCJW84]. cytokeratins [BZ84]. cytokineplast [MD82]. cytokinesis [Can80, Fuk80]. Cytological [HBB82, HHPB83, KZS81, LTAP^{+}81, SEE84]. Cytomatrix [Ste84a, Cle84, Ful84, Mas84, MP84a, Por84, VBT84]. cytometry [MPC84, Ric80]. cytoplasm [BR84b, CP83, DS80, GK81a, HFSK82, HOR84, HHT80, KST84, MSD82a, SSL^{+}84a, SWA83, SMGB84, SH81a]. Cytoplasmic [BS80, DB83, FD82, HLS81, HSC84, HA80, LR83a, OP81, Pol81, RTP82, Sat84, SR82b, Sea82, ASG83, BK83a, BP80, BR84b, BCP^{+}81, CVC^{+}84, DHS^{+}84, FN80, FLH^{+}81, GKKPB83, GDH^{+}83, GWK84, GO80, GKZ81, GK81d, GVK81, HP81, HCP81, HA84b, Kan80, LGB84, LL81, LM84, LS80, MP83, MKD80, NCH83, NBSW81, NSW82, PP84, Pip84, PATR82, RUB84, SHW^{+}82a, SDC80, SNdC^{+}84, Sto84, SR84b, TMRT84, TAC84, TQS84, TPR82, WSLSJ83, WL80]. cytoplasmically [CVT84]. cytoplast [WG82a]. cytoplasts [KKMK84, LWW^{+}83]. Cytoskeletal [BH83, GR81, Pen83, TTI80, BYW81, BCSC82, BL80b, CKP84, CBJ^{+}82, CCWR84, FWP84, GAR^{+}84, Gol82, HM83b, Jef82, JSC84, KGAB81, LC81, LCA84, MB84c, MR81a, MIS^{+}80, PFP^{+}82, SvB81, Sch82, SR81a, SM83a, TOG83, VS82, VBPL84, WMP^{+}84a]. Cytoskeleton [Nac80, PWT82, TWWG84, ATW83, AA83, BNB^{+}82, Bre83, CML^{+}81, Con84, EKSZ80, GSMHF84, GGV81, GHL84, HH81a, HGW84, JNS^{+}84, JHCC84, KOS83, LV83, LN82, LFS^{+}81, LGHI84, MBC^{+}84, ML82, MM81b, MG83b, PG82, PFM^{+}84, RMT84, SMS84b, SD80, SWB84, TdS83a, TKA83, Wal84, WP82a, WHK^{+}84, ZS84, SH81b]. cytoskeleton-associated [LV83]. cytoskeleton-related [GGV81]. Cytoskeleton-secretory [PWT82]. cytoskeletons [EDP82, HK80a, Law84]. cytosol [LM84, MOS84, PSR82, SSO84]. cytosolic [LWWP84, SW84]. cytosolic-free [LWWP84]. Cytostructural [SB82c]. cytotoxic [GRB82].

D [EPO^{+}84, KHJG84, RBO83, Rot84, SHHR84, SGKD83, TS83a, CCKMH82, GPHV83, Sch82, TRN^{+}82]. D-dependent [TRN^{+}82]. D2 [LNB84, LLB84]. D2-cell [LLB84]. D3 [MKD^{+}84]. D3-induced [MKD^{+}84]. dalton [Bre83, BFST83, BYS82, CP82b, CVC^{+}84, FJ82b, FJ82c, GCH84, GKL^{+}84, GLP80, JOA84, LYG81, MWH83, NCH83, OT82, PDK83, RCM^{+}83, SSSJ83]. daltons [LV83]. Damage [Leh81, FBS^{+}81, OKA84]. danica [TPPH80]. Darby [MBP^{+}83, MSB^{+}83, PS83b, PAS84, PBS84, RIP^{+}84, RBPS83, RBPSB84]. dark [SS80c]. data [ME84]. deadaptation [VV83]. death [KSR^{+}80]. decondensation [BPM81]. decorated [SMG81]. Decoration [BC84a, TH81, WS81b]. decrease [LH84]. Decreased [SLSK84, SC84b]. dedifferentiated [FDSWW84]. deep [CH80b, HH81a, Hir82, HKCM83, Law84]. deep-etch [HH81a, HKCM83]. deep-etched [Law84]. deep-etching [CH80b, Hir82]. deeply [HT82b]. defect [ASC84a, MMS^{+}84a]. Defective [GSPS82, BS81c, DK81b, HJ80, KvRK^{+}84, RPM83, Sin80, WAN81, WCJ83, Wri84a]. deferens [MHLS81]. deficiency [THT^{+}84]. deficient [GK84, GWWP83, GGK83, SL82a, TEL84, PG84]. defined [CAL^{+}82, FHDL82, vMS83]. Definitive [LC80]. deflagellation [BSR84]. deformability [SLSK84]. degeneration [Can83, GRZ83, PBW80, SB80b]. Degradation [GAS^{+}83a, HF83, NBMD81, BHJ83, BALR83, DB81a, Fis82c, MHRR83, MLHM84, MPS^{+}83, MP84b, SMJB83, SG84a, SC84a, TWG83, ZCH80]. degrade [COBC84]. Degranulation [HFW82, BSS83b, DHS^{+}84]. degree [MP81]. degrees [HMKM82]. dehydration [Sma81]. dehydrogenase [PLP81, RFML84, Wie81]. delta [AWB82]. demembranated [KW84]. demonstrate [KKP84b, SSG^{+}83b]. demonstrated [DCG83, DHHG83, FKS^{+}81, Sta81]. Demonstration [GDM^{+}80, OKE81, CBL84, LNA81, SFS81, SC84a, TSS82, TP82, WGRB80]. denatured [FMS^{+}84]. dendrite [BM82b]. dendritic [LR83a]. denervated [DK81a]. Denervation [MIRS84]. Dense [BS82c, MGK83, OWPC82, ST83]. dense-core [ST83]. densities [CGCS80, CGS81, EK84, FYC82, GCS81a, GCS81b, MPAM80, SBC81]. Density [Bad81, GWQ^{+}84, QGW84, AdS84, ABG81, ABBG82, BR83, BR84a, BW81, BW82, CGS81, CBS83, CGA^{+}82, CQB84, GSPS82, GHGT81, HAEC81, ISGS81, MFA83, PS83a, QCB84, RR80a, RDY^{+}83, SBC81, SLLP82, SHH82, SSG^{+}83b, VSS83, VRR80, VVBZ84, ZRW80, QGW83]. density-dependent [VRR80]. denticulata [GBS80]. depactin [Mab83]. dependence [BDPS80, CW80, LS81b, OJJ^{+}84, Rau84, FW80, JRB80]. dependent [BL80a, BYS82, BLS^{+}80, Bur82b, BGS^{+}83, BMA84, BF82c, CH80a, CCCR81, CF80, FB82, FS84, GWS84, IMS82, JPS84, JSMT82, KM82, KBD82, LLS81, LGF^{+}84, MB82b, MGW^{+}80, MF84e, MSD82b, OB81, OB82, OYNT83, PSR82, PHM81, PGD^{+}84, ROW^{+}83, RJVS81, Sch83b, Sch84, SCS84, SBJ84, SM80, SBBS84, Ste83a, Ste84b, SD82, TRN^{+}82, TTUI83, Val82, VV84, VRR80, WB81b, WS84, WH82, YFM83, YAF81, vdWSCH81]. dephosphorylation [SS81]. depleted [DRF84]. depletion [GO80, MMC83, PTH83]. depolarization [HH81b]. depolymerization [KDMS83, KR82b]. depolymerizing [Mab83, MB82b]. Deposition [ACMH83, AKT84, Car84, MKB82]. derivatives [HWBC82]. derivatized [Hat81]. derived [BPDR83, DK84a, Fle81, GDRS80, GB82, GLMW84, HHPB83, HHD83, KWM82, KORBS82, ME81, MFA83, RGB84, RUM^{+}82, SDL82, SGS^{+}82, SCS82, SCP^{+}83, SC84b, SKA^{+}83, TM84, WOM^{+}81, WC83, OHA^{+}84, VRR80]. desensitized [Nac81]. desmin [BL82a, DF84, IDF83, TDS83c, TDS83b, TMS84]. desmin-containing [IDF83]. desmosomal [KSMF84, WMG84]. Desmosome [DHS^{+}80, FWG^{+}81]. desmosome-associated [FWG^{+}81]. desmosomes [GS81]. destined [HSB^{+}82, RIP^{+}84]. detached [Can82, Can83]. detaches [NLD^{+}82]. detachment [CKB^{+}80b]. detect [EW83]. detectable [RS83]. detected [GKL^{+}84, MTMT83, PB83b, SS84a, TAC84]. Detection [HA84a, IDF83, NKH81, Ott83, SGA80, SBB81, SBBS84, TRSS81, WHJ80, DSRF84, HHSS81, HLSWH82, RMS^{+}82, SP82a, SB80c, WSM^{+}82, GBW82]. detergent [BSNP82, DYP83, GSMHF84, Goo83, SA83b]. detergent-extracted [Goo83]. detergent-insoluble [DYP83, GSMHF84]. detergent-lysed [BSNP82]. detergents [GSB81, MGLF83]. determinant [FSW^{+}83, GB83, VRR80]. determinants [MS84d, Pol84b, Sol81, SS84c]. determination [LGF^{+}84]. determine [GMRB81]. determined [CSK^{+}80, HA80, KB81a, SSZ83, WB83a]. determining [RKP^{+}82]. developed [BNGN84]. developing [AMM84b, BCHR83, BAF84, CHS82, CHCS82, FN80, GLBH82, HL81, HPG83, IJ82b, MGMS84, NCP84, SSHP82, SFS81, SBS80, TMS84, TIS^{+}81, WWCD84, WRH81, WA84]. Development [MPJ80, RCJS82, RK81, AKT84, BHR81, BDP^{+}84, COS83, CFN^{+}83, DHS^{+}80, DCK84, ERH82, ER84, FKK83, FK80, GO80, GFLG^{+}84, GK81d, HC84, HO84b, HBB82, JRM^{+}80, JRB80, MR80a, NHMT84, NCM^{+}83, PJM^{+}81, PL83c, Rag81, SB80b, SG82a, SA82, SB82b, VMRR80, WD83, FB82]. Developmental [DH83, FML83, FPFP81, MS84a, LLB84, SB83a, SF84b]. developmentally [AAW^{+}84]. Dexamethasone [BGD80, BE82, LK83]. diameter [FGHC80]. diaminobenzidine [CQB84]. diatoms [LPH83, TPHL80]. dibutyryl [CGG84]. dichloro [EPO^{+}84, SHHR84]. Dictyostelium [BCHR83, BC84a, BFST83, BYS82, BWM84, CV82, CVC^{+}84, DH83, DMD80, DSD80b, DSD80a, EW83, Fuk80, FTM82, GSMHF84, GHL84, JFS84, KRC82, KC83, LFS^{+}81, LGHI84, MR80a, PKS81, RJVS81, RFUC84, SSP84a, SHB80, TD83, VV83, VVP82, Van83, VV84, VS84, VSD84, VTSS80, WTK83, WB83b, YPR^{+}82, YF83, YMF84]. Differences [DHS^{+}84, Let83, SSJ83, ZF84, AHMS82, CH81, PR83, UAL84]. Different [MM81a, MFVPK82, CGCS80, EK84, GL83a, GC80, JC84, RLP84, SSO84, SHHR84, SCSC81, TSGS83, VTSS80, WLCS84, YHS^{+}82, YF81, ZF84, ZSM^{+}83]. Differential [AC81a, BZ84, EM84, GLBS81, HG80b, RS81, SRCG84, SSC80, WNB82, YDL^{+}84, BC80, KMCAS80, SK81]. differentiate [JVMRK82]. Differentiated [SSP81, SSSP81, SSP82a, YSJ81, RUM^{+}82, WA83]. differentiating [BWM84, GOA83, Olm81, Wat84]. Differentiation [BSH83a, CF83, NCL84, PJM^{+}81, TS81, AS81b, BS82b, BLHC81, CSH81, CWS^{+}80, EHESF82, FW80, GJRO^{+}83, GLMW84, GA84, GBE84, HSB^{+}83, HS80a, HC84, HBR81, HPG83, JFS84, KMF84, KS82, LTE84, LH84, MKD^{+}84, MR80a, MTKH84, MGSC84, MMS84b, NSI80, NSI82, PLP81, PR83, SR84a, SHW^{+}82b, SY83, SKA^{+}83, VH82, WR81, WC83, Wil84, Wri84a, ZB82, RKF80]. differentiation-defective [Wri84a]. differentiations [HH81a]. differently [BCHR83, HLM84a]. Differing [dPMMF^{+}81]. diffraction [KL82, LP83a, PL83a]. diffuse [AW81]. diffusibility [TWW82]. Diffusion [BW82, EWG82, MK84a, SSL^{+}84a, WV84, EW82, EZ84, JOA84, KOB82, MLP84, MPK84, SGKC82, WZ83, SBGP84]. Diffusive [Pai84]. digestible [BO83]. digestion [KLF80, TGPC82]. Digestive [AS81b]. digital [TMRT84]. digitonin [SO82]. digitonin-permeabilized [SO82]. Dihydrofolate [LCK84, FDSWW84, KHS82, LBM82]. Dihydroxyvitamin [MKD^{+}84]. dimensinal [PC80]. dimensional [ATW83, CKP84, FWP84, GR80, HSW82, Heu80, HPM82, MJ83, MS80b, Mur84b, NAH83, NBBS83, SAW83, FO80]. dimeric [WE80]. dimeric-particles [WE80]. dimers [KRB84b]. dimethyl [FK80, Fuk80, YF83]. dimethyloxazolidinedione [JE81]. diminish [SSCP80]. dimorphic [SM83b]. diphosphatidylglycerol [BGM81]. Diphtheria [SO80, DS80, KUM^{+}84, LFC^{+}80, MSD82a, SSO84]. diploid [BZN82, NBMD81, Rao81]. Direct [BM83, BF82c, FB82, KFS81, KOST80, KKP84a, MKK^{+}82, PM81, SKWF84, WGRB80, WN81, ML84d, SB80b, SW84, SC84a, VN82]. directed [BG83, CLJ^{+}81, RFUC84]. direction [HW83]. directions [HA84a]. directly [BR84b, SBJ84, SSA80]. Disappearance [PHM81]. disassembles [CT83]. Disassembly [Mur84a, CTKJ83, DB83, MKB81, RPB81, SM81]. disc [BP80]. discharge [CLHA80, DTBF82, WB82]. disclosed [Bar82b, VTSS80]. discoideum [BCHR83, BFST83, BYS82, CV82, CVC^{+}84, DH83, DMD80, DSD80b, DSD80a, EW83, FTM82, GSMHF84, GHL84, KRC82, KC83, LFS^{+}81, LGHI84, MR80a, RJVS81, TD83, VV83, VVP82, Van83, VV84, VSD84, VTSS80, WTK83, WB83b]. Discoidin [BCHR83]. discovery [Bai81, HHD83]. Discrete [CH82]. discriminates [BO83, GS84d]. diseases [WES84]. disintegrated [TT84a]. disintegration [WZ80]. disk [BBF84, CSHP84, RH82, RKH82, OMM82]. disks [Fri82b, GS84c]. dispersal [SHHR84]. dispersive [NLN^{+}81]. disrupted [KMF84]. disruption [CML^{+}81, GSK83, KML83, ZS84]. disruptors [TdS83a]. disrupts [BC84a]. Dissection [GQW83, QGW83, DHL84, FCKP84, GGL80, HG80c, Luc84]. dissimilarity [ONO82]. dissociable [MMC83]. dissociated [BGNT81, BF82c, CLHA80, FB82, SMTT83, YSJ81]. dissociates [MF84e]. Dissociation [BSSR84, KSR^{+}80, NPG^{+}80, CT84, HWAK83, SMH84, WKAH84]. dissociative [RMU84]. distal [HZM80, HEO82, TV83]. Distinct [VBPL84, AC81a, BPM81, GRR^{+}84, LDF^{+}84, NDS84, SB84, TWG83, TSS82]. distinctive [KNBSF84]. distinguish [SFB^{+}84]. distinguishes [LS80, RSB84, WMP^{+}84b]. Distribution [AUT83, BBP80, BN83, BGS81, CLJ^{+}81, FFRF83, FB80, GLBH82, GK81a, GMSK81, HOW80, Har80b, LSN^{+}84, Law84, MMO^{+}81, MNST82, MBC80b, MHNT81, Rag81, Rod80, SH83, SHBS80, TT83, WMW84, AW81, AP84, ABBG82, AMK83, AC83, Bad81, BLV84, BSV84, BPM^{+}80a, BPB^{+}81, BFST83, BB80d, Bur84, CBS^{+}84, DCG83, DTT81, ED84, EKWR84, EP80, EM84, FLM84, Fis82a, GG83, GF84, Gla83b, GBE84, GE80, HSFM82, HCP81, HHPB83, JD82, JB81b, KRB84a, LVL^{+}82, LAAS84, MTR81, MHRR83, Mec81a, MP81, OFO83, PWT80, PMP81a, RACGB81, RRD83, RKP^{+}82, RBO83, SHH82, SSP81, SSP82a, SSP82b, SKKC84, SP84c, SPBB84, SDZ84, TTGA82, TWH80, TS83b, WH81, WB83a, WD84, WHJ80, YAF81, ZH82, dLTvdSB83, AC81b, BM80]. Distributions [TMS84, GDTS81, GL83a, SHWS83]. disulfide [Car84, CF84b]. disulfide-linked [CF84b]. divalent [AW81, GS83, GSMR83, PP80a, SHB80, Ste82a]. divergent [SC84c]. diversity [CT82, EDK80, Fri82a]. dividing [JD82, WS83]. division [Cab83, CMM81, CRGS82, Fuk80, Har80a, Ino81a, Kub82, MBC80a, RGB84, TPHL80, WWKW83]. DMO [JE81]. DNA [BR84c, CSK^{+}80, CWS^{+}80, DK84b, DBBN81, DCK84, GKKPB83, GS84f, HR80, HAF82, HBS^{+}81, KHS82, KKBG83, Kuo82, MLSW82, MR80b, PL83b, Rao81, RS81, RSP83b, SPBB84, VPMF84, WB83a, WR83, WHSK80, Yao82, ZHTH80]. DNA-mediated [HBS^{+}81, WHSK80]. DNase [OBH84]. do [HSFM82]. docking [PWT82]. dodecyl [ZRW80]. does [GMRB81, KMI82, SCP^{+}83]. dog [LAAS84]. dog-ileum [LAAS84]. dolichol [GBWR83]. dolichyl [RACGB81]. Domain [FGS^{+}82, GKKPB83, MKB82, MPK84, RH83, RH84a, RH84b]. domain-specific [RH83, RH84a, RH84b]. domains [AC81a, BF82a, CHH^{+}83, HWM83, KKHK82, KH84, KRN83, LFC^{+}80, PB83b, RIP^{+}84, Ste83a, VBPL84]. domes [SCM84]. donors [FR81]. dorsal [GB80, SFB^{+}84]. dorsi [RMF83]. double [FN80, GSvdLS81]. double-label [FN80]. double-labeling [GSvdLS81]. doublet [HW83, TT84a, WMP^{+}84b]. doubly [RIP^{+}84]. Down [SC84a, VC83]. DRB [EPO^{+}84]. dried [ATW83, HK80a, MHLS81, SGKC82, SMG81]. driven [ACF84, DB80]. droebachiensis [POS80]. Drosophila [BWM80, CPLC80, CL82, DK81b, FSB81, FP81b, FBB82, Fri82b, HMB^{+}82, HAS^{+}81, KSI83, KI83, LV81, San81, SLLP82, SF84b, SS84b, SEE84, VSBL83, WB84, WMW84]. Drug [SHHR84]. Drug-induced [SHHR84]. drying [FSFA84]. Dual [GMVF84, CRGS82, MPC84]. Dualistic [YK84]. Duchenne [GDMG81]. ducts [GHWR80]. Dunce [DK81b]. duodenum [TRN^{+}82]. duplication [BPB^{+}81]. Durable [IY84a]. during [AR81, ASWR83, AW82, AMRE81, AB81, AKT84, Asc84b, BMF82, BS82b, BO80, BBWS83, BPM81, BHSC83, BNH81, Bow80, BOH81, BVGS82, BNM82, BvGJ83, CLFC83, CBS^{+}84, CH80b, Che81, CWS^{+}80, CGW82, COS83, CBL84, DMD80, DSD80a, DCK84, DHS^{+}84, ER84, EM84, FGHC80, FSW^{+}83, Fri82b, GF83, GHB82, GAR^{+}84, GHY81, GK81b, GL83b, GA84, GVK81, GBE84, GPDM81, GBB^{+}82, GHH82, GRZ83, HC84, Hea80, HBR81, HA84b, HBR^{+}84, HHT80, HFW82, HBB82, JOH^{+}84, JGSH82, JB81c, JFS84, Kan83, KHS82, KRB84a, KA84, KMF84, KOB82, Kub82, LR83b, LWW^{+}83, LS81a, LSJT83, LAAS84, MTKH84, MFA83, MPS^{+}83, NSI80, NSI82, NHMT84, NBM82, O’B80, PBG80, PADH83, PBW80, PMP81b, PWT82, PNYO83, Rag81, RW82a, RWD84, RUM^{+}82, RIP^{+}84, RF80b, RW82c, RKK82, RFUC84]. during [SB80a, SB80b, SSBB82, SDT81, SMTP80, SH82b, SW81a, SCN^{+}81, SAP81, SA82, SJ80, SY83, SHBS80, TBR80, TAGG80, TPHL80, THT^{+}84, VMRR80, VKE^{+}81, WMW84, WFGB83, WR81, WC83, WH80, Wie81, WG82a, WHJ80, ZB82, dLTvdSB83, vDB80]. Dye [SG82a, GK81b]. dying [HPG83]. Dynamic [SKS^{+}84a, MM81b, PG83c, TKA83]. Dynamics [FK80, Jen82, BW81, BC82, DH84, GWK84, PP80a, PP80b, SLS^{+}84, SSL^{+}84b, SB82c]. Dynein [POS80, BBPR81, BHJV80, DPS80, GH82a, HSC84, JW83, MW80, ONO82, OAOB81, Pip84, TTUI83, WM81, YT82, TH81]. Dynein-like [POS80, DPS80, HSC84]. dynein-microtubule [WM81]. dystrophic [PB84a]. dystrophy [GDMG81, GFLG^{+}84].

E2 [GKKPB83]. each [TS83b]. ear [HT82b]. Early [HMDP80, HOR84, PP84, AO84, ERH82, ER84, EKWR84, ELST81, FKK83, LDV83, MR82, SSHP82, SHPH84, SB83a, YSS82, IY84c]. ECHO [BG83]. ectopic [WRH81]. ectosymbiotic [Tam82]. edge [Che81, Sma81]. EDTA [MGG83]. Effect [BRR^{+}83, HSB^{+}83, ISGS81, KHJG84, LHHH82, LFF^{+}81, MTI80, MBC80a, PE83, PO81, PVPB82, RA81, RBS84, RS82, SCN^{+}81, SH82c, AS81a, AC81a, BBB82, DGC83, GO80, GSK83, HG80b, HF83, JB81b, LFR80, OFH^{+}82, ROW^{+}83, SBF84, TGPC82]. Effects [AB80, GBM84, GSG82, HH83a, HSCS80, MGG83, NJR80, NSW82, OAOB81, PWM81, POYF81, PTH83, RE80, VS84, VSLS80, ASA83, BF84a, BMCGO^{+}82, CH81, FM80, Gol82, HMR82, HA80, KJBR82, KWSB84, KHS80b, LUT80, MLP84, MMC83, MLMS81, MW80, PL83a, PP80a, PP80b, Per81a, PWT80, SRCG84, SJM81, TS83a, TMA^{+}83, WWS83, YHS^{+}82, YT82, DPK84, MR80a, WW83]. efficient [LK84, vMS83]. efflux [BL83b]. egg [BRH82, BSH83b, CH80b, GGVR82, HOW80, HHS^{+}81, HM84c, Kan80, KDL80, KL81, KRN83, KMV82, Pip84, SH82a, WMP^{+}84b]. eggs [EKWR84, GG83, GWK84, Gil83, JE81, KNHK84, KJBR82, LS81a, LM84, MBC80a, RF84, RB83, RPHA81, SHHP81, SS80a, SSBB82, TJ80, WS84, WN81, WZ83]. Ehrlich [GOSW82, GKZ81]. eIF3 [OLHH81]. either [Kan80]. ejaculated [KNH80]. Elastase [CVAK80]. Elastase-like [CVAK80]. elastic [CHRB84]. Elastin [MLMS81, MGMS84, SGM^{+}84]. elastogenic [MMS84b]. electric [CK80, CSR^{+}81, EN84, Fro84, GNW^{+}84, HCK80, Sea82, Sow84]. Electrical [SCM84, FCNS84, JGSH82, SR83b, Sur82]. electricus [EMA83]. electrodes [MR82]. electrogenic [GCKAA83]. electrolyte [SBP81]. electromigration [MLP84]. Electron [DK81a, GC80, HP81, OFH^{+}82, PP81, TTI^{+}81, AO84, BKL^{+}82, BM83, BBS81, CKP84, CGA^{+}82, DGS80, ECM81, FCKP84, GDL80, GL82, GAS83b, HAF82, HLSWH82, KL82, KB81a, LCA84, MWI^{+}84, MSB80, MS84e, OOZF80, PPSP83, PM81, SSJS84, SEPB81, SBP81, SHH82, SKKC84, SGSS^{+}81, Ste82a, SHWS83, TTI80, WSB^{+}80, Wol80b, ZRW80, dHLC84, BB80a, TI80]. electron-probe [SGSS^{+}81]. electrophoresis [HPM82, PC80, RFLR81, VKHH82]. Electrophoretic [SHHP81]. electrophoretically [Sta81]. Electrophorus [EMA83]. electrophysiological [BNGN84]. elegans [CT82, CKE81, NRW82, RW82b, WAN81]. Elemental [SGKC82, BBS81]. elements [CBJ^{+}82, RS84, WW83]. Elevation [OB82]. elicitation [SHB80]. elicited [PKB84]. Elimination [Yao82]. elongate [TBD81]. Elongation [DGC83, BRH82, FW80, KHL82, LPH83, LTGL80, OB81, TI82b, TH84, WB81b, ZHTH80]. elsewhere [YBG81]. elution [SHB80]. embedded [DHHG83]. embedding [HHS^{+}84, HS84]. embeddment [CKP84]. embeddment-free [CKP84]. embryo [ALE84, BWV84, CB82, CCKMH82, GT84, LMYM84, ML83, NDRL80, OT80b, RCM^{+}83, SBS80, SS84b, WKW82, WG82a, WBG83]. embryogenesis [Rag81]. embryonal [GOA83, HBR81, JVMRK82, MGSC84]. Embryonic [EN84, ASM^{+}84, BS82b, BB82a, BDP^{+}84, BGE84, BL82a, BNGN84, BL82b, CG84, CDvdM80, CEC84, DF84, DGD^{+}84, Ekb81, GBF84, GGL80, Hat81, LSBPG84, LFS^{+}83, MPJ80, NLD^{+}82, OLH^{+}81, PFP^{+}82, RRS82, SSL^{+}84a, SG82a, SHJA82, SB82b, TS83a, YSJ81]. embryos [BDP^{+}84, FBB82, GK81d, GK82, Let83, MLNK83, MU82, POS80, RB83, SLS^{+}84, SF84b, TS81]. Emergence [GA84, RUM^{+}82]. emphasis [BWW^{+}80, RA81]. emtB [WC80]. Encapsulation [OLHH81]. encoding [RC83]. end [BM83, CT83, HZM80, TBCM83, VH82]. end- [VH82]. endo [RMJ84]. endocervical [CNSW80]. endochondral [PPR82b, PPR^{+}84, SR84a]. endocrine [MKK^{+}82, MMAO83, SD82]. Endocytic [Oli82, AR81, Max82, MSC83, SPS^{+}84, TKM83, YFM83]. endocytose [CKS82]. Endocytosis [MH84, SMMC83, ZF81, Asc84b, Bow80, BOH81, DWP81, DH84, GSOG80, GSH^{+}84, HHS83, KvRK^{+}84, KI83, MFA83, MS83a, PBWK83, PSBF83, QCB84, RPM83, SCS80, SDT81, SG82b, TBR80, TSS82, VSS83, WBPO80]. Endodermal [GLMW84]. Endogenous [CHRB84, CB84b, SDGYJ84, SHB80, BHR81, BRS^{+}82, BB82a, BB82b, BZN82, DB83, DR82, EW83, FJ82a, Sol81, SHO^{+}84, LN81]. endomembranes [SVW^{+}83]. endometrial [RS83]. endometrium [GHH82]. endoplasm [NW82, SWA83]. endoplasmic [ACTK84, AWB82, BPM^{+}80a, BVC83, Bro81, CHS82, CHCS82, DYP83, FNHSS84, FHFL82, FFS^{+}82, GS80a, GBW82, GWB82, GIKR84, HTGB84, LK84, LRW82, MACK84, ML84a, MD80a, MD80b, RMH^{+}84, TLPTV83, UPLH84, WIB81, WB81a, WB81b, Wal82a, Wal82b]. endorphin [Sur82, ME81]. endosomal [ROB^{+}84]. Endosome [MPC84, HA84b]. Endosomes [GSH^{+}84, PAS84]. endosperm [EJM82, JD82, Jen82, MBB83, WHJ80]. Endothelial [DK84a, CARK81, DWM83, FDKF82, FCNS84, GDRS80, GF84, GCL83, GLMW84, GMSK81, GSK83, GMJR82, ISGS81, KBDK84, LS82, LTBS80, LL82b, MHSW81, MKW^{+}82, MW83a, MOV83, MMR^{+}84, MDJ82, STR84, VVBZ84, WOM82, WM84, WGF83]. endothelium [DBK80, HSS82, MM81a, OJJ^{+}84, SSP81, SSSP81, SSP82a, SSP82b, SLS83, VSS83]. endplates [WRH81]. ends [BS81a, MPW82, TBD81]. energy [LYG81, NLN^{+}81, SJM81, TRSS81, WD84]. enhance [FMS^{+}84]. Enhanced [TWW82]. enhancement [AH84, GMO84, PBG80]. enhances [CT84, GNGTS80, Ino81b]. enoyl [RQHR81]. enriched [LMYM84, RNH^{+}83, SBVW82, SG84b, WTSAC^{+}81, YDL^{+}84, TZW83]. enrichment [CGCS80]. enter [GN84, SSO84]. enterocytes [SGA80]. entire [MDA83]. entities [SB84]. Entry [KUM^{+}84, CRGS82, DS80, HKSF80, IYVV83, KCGH84, MOS84, MRHS81, SO80]. enucleated [RVM83]. enucleation [WG82a]. env [WSH84]. envelope [EB84, FSKJ81, GMRB81, LM84, MMO^{+}81, MLNK83, SF84b, WD83]. envelope-associated [WD83]. enveloped [AC81a, CFW84, RBPS83]. Enzymatic [FBP82, GGL80]. enzymatically [DB80]. enzyme [BM80, BPM80b, DK81b, GGK82, GNGTS80, GPR83, KLF80, LSJT83, MTMT83, MCPB80, SSBR84, SCN^{+}81]. enzymes [BPM^{+}80a, CVAK80, EK84, FNNFS84, FNSB80, Fle81, Mas84, NDS84, RFML84, RACGB81, RKH82]. enzymic [QCB84]. eosinophils [CKB^{+}80a, CKB^{+}80b, KCB84]. ependymal [SFS81, SLS81]. Epidermal [Bar82a, GGV81, GSH80, SMT80, SM82b, SM82a, SHT^{+}83, AKM^{+}82, BS82b, BSH83a, BMH^{+}83, BPDR83, BBG^{+}84, CMC81, CGG84, CH81, DH84, EBS84, HEG81, JWG84, KOLV81, LH84, LA83, SSE84, SRCG84, SY83, SMTT83, SC84a, WG81, WES84, WP82b, YSS82, ZB82, HDB82, CGA^{+}82]. epidermis [MFVPK82, RWD84, Wat84, WMENS82]. epidermoid [Bar82a, CMC81, GSPS82, WCZG84]. epididymis [BvGJ83]. epilation [HDB82]. epinectin [EF84]. Epinemin [Law83, Law84]. epiphyseal [PPR82b]. Epithelia [GH82b, BSG81, GW82, MFVPK82, NS83a, PR83, THT^{+}84]. Epithelial [CFN^{+}83, FWP84, FCKP84, GOA83, MBC83, NCM^{+}83, ASG84, BZ84, BFL^{+}83, Bur82b, CLFC83, CGM81, CFCS82, DKDB81, DB81a, EPM84, EF84, FHG^{+}81, FFGC84, GDTS81, GGW83, GLMW84, GMRB81, GDRS83, HBS^{+}82, HO84a, HEO82, HO84b, HH81a, HTFH82, HM83a, KSF84, KOLV81, Kla80, KML83, Lan82, LVL^{+}82, LA80a, LPB84, MS84b, MSB^{+}83, MMJ^{+}81, MGW^{+}80, MKD80, NNA^{+}83, OM82, PTH83, SSJ83, SSG^{+}83a, SSMH82, SCM84, SH81b, SCSC81, YKJF80, ZSE80, VY82]. Epithelial-mesenchymal [CFN^{+}83, NCM^{+}83]. epitheliomas [MFVPK82]. epithelium [CFN^{+}83, DGS80, EDK80, FKB83, GWP82, GAA83, HSME84, HO84a, HBO84, MPMBC80, MIS^{+}80, MM81a, Mil81b, NCM^{+}83, NW80, OKE81, VVM81]. epitopes [MWI^{+}84]. epsilon [MLP84]. Epstein [NPG^{+}80]. equilibria [ML84d]. equilibrium [CQB84, LL82a, PB84a]. Equimolar [Ste82b, ME81, MACK84]. ergocryptine [APZL80]. errors [WG82b]. Erythrocyte [SMAS81, AMRE81, AS81a, CN80, CF80, EL84, Fis82b, FLH^{+}81, HSCS80, IBB^{+}81, JTM80, LWW^{+}83, LHS82, LBA^{+}84, LZR82, LA80b, MTKH84, NGL83, NAK83, OK82, PF81, Per81a, PG83c, SP82a, SP82b, SLR^{+}83, SJS84, SH82c, Sow84, TTI80, TTI^{+}81, TTI84, dST82]. erythrocyte-specific [MTKH84]. erythrocytes [BCSC82, CBJ^{+}82, FH84, FBST84, GRL82, GAS83b, JSC84, KRB^{+}84c, LBA^{+}84, MTW82, MMVO83, NJSC83, SD80, SS84c, TMW82, WWES83, Woo80]. erythroid [BL83a, BLHC81, ML83, ML84c, NW83, SP82a, SBSL84]. erythroleukemia [BLHC81, CWS^{+}80, EMC83, NCL84, VH82]. erythroleukemic [MTKH84]. erythrophore [BP83]. erythrophores [SO82]. erythropoiesis [EMC83]. Erythropoietin [VC83]. Escherichia [SR82a]. esophageal [MM81a, Mil81b, THT^{+}84]. especially [EH81]. essential [MD80a, MD80b]. Establishment [CHH^{+}83, Reb82, SSCP80, CKB^{+}80b]. ester [EB83, FMS^{+}84, MA83a]. esters [BC84b]. Estimation [IYVV83]. estradiol [APZL80]. Estrogen [RS83, SDG^{+}84]. etch [HH81a, HKCM83, TTUI83]. etched [HT82b, Law84]. etching [CH80b, Hir82]. ethinylestradiol [HAEC81]. Ethylmaleimide [PCB83, MBC80a]. eucaryotes [BTF^{+}83, IY84a, IY84c]. eucaryotic [Deu84, Luc84, Tam82, TT83]. Euglena [CB84b, GHB82, NJR80, RB80, RB82a]. eukaryotes [Gib81]. evaluation [ME84, PBFC^{+}84]. event [ACF84, AAW^{+}84, BPM81, WKAH84]. events [AW82, CH80c, EKWR84, EPE^{+}80, HMDP80, KSD^{+}83, RKM^{+}81, RPHA81, SSHP82, Sch81c, SB83b, SEE84, YHS^{+}82]. Evidence [AR81, ASWR83, BVAR83, Ber80b, BZN82, BF84b, FCO83, HSB^{+}82, HZM80, HH83b, IMS82, JC84, KMI82, MB80b, MKD80, NKT80, RR80a, SSO84, SPS^{+}84, YRS83, AB81, AS81b, Baj82, BM82b, BAWL81, BVC83, BTL84, BS81a, BBG^{+}84, CKB^{+}80a, CKB^{+}80b, CVGM83, DS80, EDP82, FCL^{+}82, FS84, GOSW82, GC80, HG80a, Her80, HTGB84, HH81b, Kie81, KCB84, KRB84b, LFGR82, MBP^{+}83, MLSR83, MS84d, MMIGK82, NDS84, RB82a, SKCJW84, SW84, SB84, SLL80, TH81, UE83, VN82, WFR84, ZSM^{+}83, vDB80, BPM80b, CLHA80, LC80, MCPB80, ML80, OT80a, Ric80, SBG80, SDT81]. evoked [NFS81]. examination [LCA84, ZLK81]. exchange [BC84b, PSSS82, WCZG84]. excision [AL84]. excitability [HHT80, RSP^{+}83a]. excitable [ARL^{+}80, ERN83, LN81]. Excitation [VV83]. exclude [AFSF^{+}81]. Exclusion [MTKH84, MLZD83]. exhibit [Bec84]. exhibits [MPK84]. existence [KUM^{+}84, LC80, Pip84]. exists [SPEL84]. exit [BTF^{+}83, HOR84]. exocrine [FJ82a, FJ82b, Oli82, SJC80, SKS84b, TDM80a]. Exocytosis [AMS82, BAWL81, ACF84, BCLT^{+}80, BPM81, HH83a, KMF81, LWWP84, OR81, PBWK83, RLP84]. exocytosis-specific [RLP84]. exocytosis/endocytosis [PBWK83]. Exocytotic [PBFC^{+}84, LTAP^{+}81]. exogenous [BALR83, NSW82, RDY^{+}83]. expansion [GS84a, PMP81b, RF80a]. Experimental [HG80c, MPMBC80]. experiments [AB81]. explanted [SR81b]. Export [NS83b, ES82]. exposed [PS81]. exposes [SL82b]. exposing [QCB84]. exposure [PBFC^{+}84]. express [CAL^{+}82, FSBR83, GH82b]. expressed [GCH84]. expresses [HEO82]. Expression [ALE84, COBC84, GKKPB83, KMF84, KKBG83, LVL^{+}82, PL83c, SG82b, SCSC81, THT^{+}84, AFW84a, BB80b, BGS^{+}83, COS83, DCK84, EBS84, EM84, FSVW80, GF83, GSG82, HC84, HK80b, KZS81, KLC83, LFAG84, LCK84, MJM83, MSB^{+}83, MF84d, NHMT84, NCL84, RBS84, SSG^{+}83a, SHFK83, WES84, WA84, WGF83, ZLK81]. extension [DTBF82, PWSW80]. extensive [PSC83, RKMF84]. extent [OT80a, RCH81]. external [RB82a, SSG^{+}83b]. Externalization [BHR81]. externally [CHH^{+}83]. extinction [MNW81]. Extracellular [AKM^{+}82, BAF84, BT84, BR84c, DVvdMvdM81, Hay81, MMS84b, AHMS82, BWM84, Car84, CF84b, DGD^{+}84, GAS^{+}83a, GFLG^{+}84, HJ81, HOMR82, HTKC82, LKJ84, MKB82, MLMS81, MRB84, NLD^{+}82, PTH83, PPR^{+}84, SR84a, SL82a, SBCS^{+}82, SH81b, TSGS83, WHK^{+}84]. extracellularly [CHRB84]. extrachromosomal [CDLJ81]. extract [Kan80, SSHP82, STRP83]. extractability [SA83b]. extracted [Goo83, TI82a, LCA84]. extraction [Bar82b, CB84a, MD80b]. extracts [MFP80, WSM^{+}82]. extrasynaptic [San82]. extruded [GS84b, RN81]. extrusion [BCLT^{+}80, Wei82]. exudate [VC83]. eyes [GFLG^{+}84, PPR^{+}82a].

f [MS83b, BRS81, CBW81, CF80, FLH^{+}81, LGHI84, MWH83, PNYO83, RMJ84, WMW84]. F-1 [MWH83]. F-actin [BRS81, CBW81, CF80, FLH^{+}81, LGHI84, PNYO83, WMW84]. f/b6 [MS83b]. F1 [Ste84b]. F1-ATPase [Ste84b]. F9 [GOA83, GAA83]. Fab’ [MKB82]. facile [RMU84]. facilitated [SO80]. factor [ABSC81, Bar82a, BKT84, BRAC83, BG82, BMH^{+}83, BPDR83, BBG^{+}84, BGS81, CGA^{+}82, CMC81, CGG81, CGG84, CH81, DH84, GDRS80, GWK84, GGV81, GSH80, GLS83, GL80b, GS80b, GB80, GLBS81, HFW82, HHD83, HEG81, IMS82, KUM^{+}84, KOLV81, LFGR82, LRT^{+}82, LNA81, LAAS84, LH84, LA83, MLZD83, MLNK83, OFH^{+}82, RNH^{+}83, SMT80, SSE84, SRCG84, SGS^{+}82, SM82b, SM82a, SCS82, SCP^{+}83, SC84a, SC84b, TSGS83, VS82, VPMF84, VC83, VRR80, WP82b, YSS82]. factor-binding [LH84]. factor-cobalamin [LAAS84]. factor-induced [HFW82, HEG81, SM82a, VPMF84]. factor-responsive [BG82]. Factors [CB81, GHGT81, WGF83, ASWR83, FK80, GCL83, MFP80]. Failure [EW83, KvRK^{+}84]. falciparum [GAS83b, LBA^{+}84, OK82, Per81a]. falciparum-infected [GAS83b]. false [DTBF82]. false-discharge [DTBF82]. family [GRDR82]. fasciculation [RE80]. fast [GL83a, HSB^{+}82, IYVV83, PYC^{+}82]. fat [Bro81, FHG^{+}81]. Fate [CML^{+}81, SDT81, SSF84, ZH82]. fates [BO81a, BKO81]. fatty [Cre81, KRB^{+}84c]. Fc [MLP84, MPS^{+}83, MP84b, MPU84, Pfe84]. FCCP [KHS^{+}80a]. features [BCHG83, VB81]. feeder [AKM^{+}82]. feeding [SBCS^{+}82]. feet [CC82, FSFA84, KZS81]. feline [IRMM84]. female [BRR^{+}83]. fenestrae [SLS83]. fenestrated [SSP82b]. ferax [Hea80]. Ferritin [MEKT84, CGA^{+}82, HSS82, HP81, KLF80, VVM81]. ferritin-labeled [HP81]. ferritin-low [CGA^{+}82]. fertilization [BL83b, CH80b, DGC83, EKWR84, GS84e, HOW80, JGSH82, KL81, MP84c, RF84, SSBB82, SBC82, TJ80, WAN81, WN81]. fertilization-defective [WAN81]. fertilized [RPHA81, WZ83]. Fetal [SMJB83, WHD^{+}84, JWG84, LNB84, MGMS84, PPR82b, RUM^{+}82]. fetoprotein [CMF82]. few [SFB^{+}84]. fiber [DANH84, GSS82b, GRZ83, HWS82, HAFG82, Jen82, PG83b, PG83a, RK81, San82, SGB^{+}81, WCJ83, ZSR^{+}82]. fiber-defective [WCJ83]. fiber-like [DANH84]. fibers [AF83b, BF82b, CHRB84, CS83, GLBH82, GBLH83, GL83a, HWJW84, MSK^{+}84, MGG83, PSC83, PNYO83, SB80a, SSJ83, Sin82, SKKC84, WGF83, Can83]. fibril [BT84]. fibrillar [AFRS83]. fibrillary [HSME84]. fibrils [Ber80a, GBS80]. fibrinogen [LA84]. Fibroblast [FNSB80, Gri80, Che81, EB83, EN84, GCL83, LRT^{+}82, LQ82, MKB82, MNW81, RH81, SDL82, WG82a, YF81]. fibroblastic [Gla83b, ZGJ^{+}84]. fibroblasts [AFW84a, AFW84b, ASM^{+}84, BW81, BW82, BHJ83, BT84, BRAC83, BWV84, CGA^{+}82, CS82, CDvdM80, CH82, CRGS82, CHRT83, DGD^{+}84, EW82, FB80, GAS82, GDMG81, GPHV83, Gol82, GB82, GK81b, HG80a, HG80b, HJ81, HCJV83, Heu80, JOA84, KFS81, Kla80, LVL^{+}82, LUT80, LMYM84, MB84b, ML84a, Max82, MLM83, MLHM84, MLMS81, MGMS84, MMS84b, MS83a, NHF^{+}83, NBMD81, OT80b, OLH80, PLMS81, PvD83, PWT80, RCM^{+}83, Rao81, SMJB83, SSJ83, SDT81, SSG^{+}83b, SGH^{+}83, SGM^{+}84, SGS^{+}82, SS84a, Sin82, SSF84, SBVW82, SK81, SSCP80, SC84a, TH84, Wan84, WKW82, WG82a, WBG83, WG82b, WHK^{+}84, VN82]. fibroin [Liz80, TOK^{+}84]. Fibronectin [CRGS82, CKHF80, ECM81, EKS^{+}82, GF83, GHH82, GML^{+}80, HJ81, MG83a, NNA^{+}83, SBS80, ACMH83, BE82, BRAC83, BDP^{+}84, CRH81, COBC84, FMS^{+}84, GWWP83, Gri80, HJ80, HER81, HOMR82, HSC80, HC83, JKHT81, JH84, KOST80, LSC83, LLM82, LUT80, MRFF80, MF84b, MKB82, MK84b, MLM83, OMG82, PF83a, PWT80, RCH81, RDY^{+}83, San82, SKE84, Sin82, SKKC84, SSF84, SK83, SK81, TGPC82, TMA^{+}83, VKE^{+}81, WR81, WLCS84, YK84, YKJF80, vdWSCH81]. fibronectin-coated [HJ80]. Fibronectin-independent [HJ81]. Fibronectin-mediated [GML^{+}80, MG83a, TGPC82]. Fibronectins [HY82]. fibronexus [SKKC84]. fibrosarcoma [MW83b]. field [Sow84, SR83b]. field-induced [Sow84]. fields [EN84]. filaggrin [RWD84]. Filament [DJ81, HK80a, SM83b, BRH82, BM83, BS81a, BP82, CTKJ83, CP83, CT83, CS83, DSKS84, EDP82, FSB81, FWP84, FA83, FBRP84, FSW^{+}83, GMVF84, GV82, KKP84a, KML83, LTGL80, OB81, PWSW80, Pol84b, PL83c, SKCJW84, SSG^{+}83a, SDK^{+}84, TBCM83, TDS83c, TDS83b, WCY^{+}83, Wan84, YF81]. filament-associated [PL83c, WCY^{+}83]. filament-microtubule [DSKS84]. Filamentous [GSB81, PG84, SL82b, BR84b]. filaments [AFRS83, AFSF^{+}81, ASG84, BSS81, BL83a, BV83, BMF84, BFL^{+}83, BL82a, BRS81, CLFC83, CV82, DF84, DSKS84, FSF^{+}80, FKS^{+}81, GHL84, GRL82, HTS80, HT82b, HTFH82, IDF83, KSMF84, KL82, KS83, KOLV81, LVL^{+}82, Law83, Leh81, LKR^{+}83, MSK^{+}84, MB82a, MPAM80, MP83, ME82, MM81a, Mil81b, MGW^{+}80, MPW82, MG83b, NAH83, OT82, PSSS82, PP84, PM81, Pol82, PSM84, SS80b, SDL81, SKCJW84, SJG84, SHT^{+}83, SCSC81, TI82a, TDM80b, TBD81, TS83b, TEDS83, TT84b, Tok83, TG83, TIK81, WB84, WRM83, WS84, WW83, WSB^{+}80, Woo80, YPR^{+}82, ZS82, ZHM80]. Filamin [NDC80, BH83, GL83a, GL83b, LdMM^{+}84]. Filipin [MFA83, BN83, NB81, PB83b, SBBS84]. Filipin-cholesterol [MFA83]. filipin-sterol [BN83, NB81]. Filopodelike [YF83]. filopodial [ML84b, RK80]. filtration [RKF80]. Fimbrin [BW80]. final [BCLT^{+}80, BPM81, GHGT81]. Fine [PPSP83]. Firm [Fri83]. First [Sch81c, HOW80, KHS82, Kub82, RN83]. fish [Gil83]. Five [FG81]. Five-hour [FG81]. fixation [MP83, MF84a]. fixed [AUT83, FB80]. flagella [AHPL81, AB80, BLH82, CB84b, DHL84, GM84, Gib81, GW80, HR81a, HW83, HPRL81, KW84, KJ82, LR83b, LOL81, LFR80, Oku80, PCOK82, PHRL81, BLH82]. Flagellar [HWJW84, MHCG80, RB82a, SCM83, AdS84, BFW80, BM82c, FW80, GHB82, HG80c, JR80, KJ82, LR83b, LD81, LL81, MR81a, MLSR83, NJR80, OAOB81, RW82a, RB80, SBSM84, SM80]. flagellar-specific [RB80]. Flagellated [Tam82]. flagellates [Wal84, YT82]. flagellum [Luc84, RB82a]. flexible [ECM81]. flourescently [FB80]. flow [AS81b, Can82, KOB82, MSC80b, MPC84, Ric80, RW82b, RW82c, SCM84, VKHH82]. fluid [BAWL81, GSOG80, GSH^{+}84, GN84]. fluid-phase [GN84]. Fluidity [IBB^{+}81, KC82]. fluorescein [KFS81, LSM^{+}84, SSL^{+}84a, SLS^{+}84, TOG83]. fluorescein-labeled [KFS81, LSM^{+}84, SSL^{+}84a, SLS^{+}84, TOG83]. fluoresceinated [HSC80]. Fluorescence [Moo81, NBSW81, Axe81, BO80, CF83, GHY81, JWBC81, LSM^{+}84, MPC84, SSL^{+}84a, SLS^{+}84, TRSS81]. Fluorescent [BW81, OGS82, Bec84, PNYO83, SP84b, SKWF84, SSF84, TPR82, WS83, WSLSJ83]. fluorescently [GS84b, PLMS81, SMS84a, SMS84b, SKS^{+}84a, TWH80]. fluorescently-labeled [SMS84b]. fluorescinated [FB82]. flux [BHSA82, Pfe84, SS81]. fly [JL84, SLL80, SL82b]. foam [MA83a]. focal [GAS82, Sin82]. focusing [BFL^{+}83]. Fodrin [LW81, CBS83]. folds [PHM81, SH83]. follicle [CGM81, Her83]. follicles [FHDL82, GW82, NW80]. follicular [BGNT81]. following [ABD84, MB84c, PAGG81, PYC^{+}82, WWDYW83]. foot [ML84a]. foraminifera [TA81, TKA83]. forbesi [ER84]. force [DTBF82, HWS82, JPS84, KMI82, Nic83]. forces [HH83a, KMF81]. forest [HKSF80, WKH80, GKKPB83, GQW83, KKKH84, KKBG83, MBWH83, WMH81]. fork [RF80b]. form [AC81a, HPG83, JB81a, MFA83, MM81a, Mil81b, SS84c, WWS83, WW83]. Formation [Ekb81, Fuk80, ZSM^{+}83, AMCG83, BRH82, BT84, BO81b, BWW^{+}80, BvdMN84, CSHP84, CGM81, CGG81, CGG84, Con84, DVvdMvdM81, FNNFS84, GW82, GHY81, GBS80, GGSB83, GMO84, GVK81, GDRS83, GOA83, GAA83, HMDP80, HBO84, Her80, Heu80, JNS^{+}84, KKP84a, KHL82, KBDK84, LM84, MP80, NvIvGdH^{+}82, Pen84, PPR^{+}84, RN83, SBJ84, SM83b, TRG84]. formed [GWS84, MJ83, SCM84, WPKB83, Wri81]. forming [BvdMN84, TQS84]. forms [Bar82b, CEC84, FSBR83, HLM84a, MR81a, MLSR83, OT80b, PA83, SF84b]. formyl [NFS81, JNS^{+}84]. formyl-methionyl-leucyl-phenylalanine [NFS81]. formylmethionylleucylphenylalanine [BG83]. found [CCKMH82]. fraction [AC82, Ber80a, FBB82, FJ82b, KL81, MMIGK82, PSR82, SWGB80, SP82b, SBVW82]. Fractionation [Ber80a, DB81b]. fractions [BRS^{+}82, BPM^{+}80a, BBP80, GIKR84, HP82a, HP82b, Man82, PVPB82, RA81, RZSW83, SG84b, YDL^{+}84]. fracture [AdS83, AMRE81, AS81b, AC83, BNGN84, CH80b, FGHC80, Fis82a, PF83b, SLLP82, SEM80, TdS84, TJB83, ZF81, ZFP84, dSTK81, dST82, dSK84, vDB80, BP80]. fracture-label [AdS83, TdS84]. fractured [GS80a, dSTK81, dST82]. fragment [FNSB80, GLP80, LFC^{+}80]. fragmented [BBCL82]. fragments [EB84, GHL84, MHRR83, PKS81, SJS84, YK84]. Fragmin [SM83c]. framework [CKP84, FWP84, FPFP81, Jef82, PFP^{+}82]. Free [BHA82, ALE84, Bar82a, CKP84, CFCS82, FHDL82, Kla80, LWWP84, LM84, MR81b, Mec81b, MP81, MPK84, POYF81, PATR82, RVB80, RFML84, RUM^{+}82, RTP82, RUB84, SW84, SK81, TD82, TPR82, VKHH82, WSM80]. free-flow [VKHH82]. Freeze [AMRE81, AC83, BNGN84, FGHC80, SLLP82, TJB83, dSTK81, dST82, AS81b, CH80b, FSFA84, Fis82a, GS80a, GDL80, HK80a, HH81a, Hir82, HKCM83, HA80, HHS^{+}84, HS84, MHLS81, PF83b, SGKC82, SMG81, TTUI83, ZF81, ZFP84, vDB80, BP80]. freeze-dried [HK80a, MHLS81, SGKC82, SMG81]. freeze-drying [FSFA84]. freeze-etch [TTUI83]. Freeze-fracture [AMRE81, AC83, BNGN84, FGHC80, SLLP82, TJB83, dSTK81, dST82, AS81b, CH80b, Fis82a, PF83b, ZF81, ZFP84, vDB80, BP80]. freeze-fractured [GS80a, dSTK81, dST82]. freeze-substitution [GDL80, HA80]. freezing [CH80c, CH80b, HHS^{+}84, HS84, OR81]. frequency [MP81, NDC80]. fresh [ZFP84]. freshly [SR81b]. friend [CWS^{+}80, EMC83]. Frog [KS83, Bur82a, Bur84, CH80a, FGHC80, HR81b, Hir82, Ko84, LHHH82, MSK^{+}84, MH84, MG83b, NB81, PPSP83, SM83a, SL80, UPLH84]. frozen [BR84b, BBS81, DCG83, FBH83, GSvdLS81, GS84c, HT82b, JSMT82, JSCM83, Law84, RGR80, RMAF80, SEPB81, SBP81, SR82b, Tok83]. frozen-hydrated [BBS81, SEPB81, SBP81]. fucosyl [SG82b, ZB82]. Fucosylated [BS82d]. fucosyltransferase [FBP82]. Function [CGS81, GCS81a, GCS81b, BDP^{+}84, BBB82, BLH82, CBJ^{+}82, Deu84, ES82, GVK81, HSB^{+}83, HH83a, HWS82, HPRL81, JOA84, JSC84, KP84, MMS^{+}84a, MBB83, MP81, MP82, OFH^{+}82, OMG82, SA83a, SLIF80, TPHL80, TWWG84, WL80, WLCS84, YK84]. Functional [Baj82, BM82c, EB84, GAAG82, RVM83, SB80a, CT82, FHDL82, HBR81, MD82, MSRB84, SO82, TRC^{+}84, ACTK84]. functionally [TSS82]. functions [BO80, CAL^{+}82, CVC^{+}84, KOB82, MJM83, MBC^{+}84, QGW83, ROB^{+}84, WH82]. Fundulus [EDK80]. fungal [HA80]. Fungi [AB81]. fungus [Hea80]. furrow [CMM81, KOB82, NDC80]. further [SKCJW84, MSC83]. Fusarium [AB81]. fused [CKB^{+}80b, PB83a, RUB84]. fusing [FPFP81]. Fusion [WKH80, ACF84, BSP83, BF82a, BCLT^{+}80, CKB^{+}80a, CKS82, CH80c, CAZF84, CNH80, Cre81, GDH82, HMM82, KMF81, KC80a, KKKH84, KKBG83, OT80a, OLH^{+}81, PRK^{+}82, RRD83, SDC80, Sow84, WMH81, WSLSJ83, KSC82].

G [BS83, AL84, BKL^{+}82, MHRR83, ML80, Mil80, PAS84, PBS84, Rod80, TTI84, WFGB83]. G-actin [TTI84]. G0 [IY84a, IY84c, Ric80]. G1 [ASWR83, HR80, IY84c]. G2 [CKJ83, GSH80]. galactolipid [MS83b]. galactosamine [RBO83, Rot84]. Galactose [BRS^{+}82, WH81, GBB^{+}82, QCB84]. galactose-exposing [QCB84]. Galactose-specific [WH81]. galactosidase [NLN^{+}83]. galactoside [Wei80]. galactosyl [SB83a]. galactosyl-containing [SB83a]. galactosyltransferase [Fle81, RB82b, SH82a, SH82b, SVW^{+}83]. gametes [GDH82, MHCG80, SM80]. ganglia [RE80, WWWC80]. ganglion [DK81a, LSM83, SW81a, SFB^{+}84]. ganglioside [KPPC83, Pie82]. gangliosides [SKWF84, SSF84]. Gap [BCHG83, GS84c, PP80a, PP80b, FG81, Her80, HAFG82, KG80, RGR80, SNdC^{+}84, TdS83a, UE83, ZCR80]. gap-junction [FG81]. garfish [CE80]. garland [KI83]. gas [Sta81]. gastric [RBH^{+}84]. gastrulation [BDP^{+}84]. gel [CGM81, FO80, HPM82, NAH83, Ott83, PC80, RFLR81, SSCP80, YAF81]. gel-sol [YAF81]. gelatin [GML^{+}80, HSC80, MKB82, VKE^{+}81, vdWSCH81]. gelatin-binding [MKB82]. gelatin-coated [GML^{+}80, VKE^{+}81]. gelatin-latex [vdWSCH81]. gelation [BFST83, MFP80, Pol84b, WC84]. gels [GSG82, GBF84, GH82b, HEA^{+}82, Kan80, Kan83, SBB81, TH84]. gelsolin [LK83, WYK^{+}84, YAF81]. Gene [LBM82, AFW84a, AFW84b, ABD84, BB80b, GSG82, GRDR82, HK80b, HBS^{+}81, KZS81, KSI83, LCK84, LZR82, MLZD83, NHMT84, NCL84, SAP81, SEE84, SHFK83, SC84c, WA84, WSH84, WH82]. general [DCG83]. Generation [GM84, LWW^{+}83, Sch81c]. Genes [FNHSS84, EPO^{+}84, FG80a, FP81b, FDSWW84, HC84, HAS^{+}81, KSG81, KHS82, KSSR84, KLD^{+}82, Sch81a, SHHR84, WC80, Wri84b]. Genetic [DHL84, Luc84, HPRL81, Mil81a]. Genetics [Raf84, BCC^{+}83]. genome [YBG81]. genomic [BB82c]. geometry [JZ81, SR84a]. GERL [BO81b]. germ [EL84, LSM83, LA80b, MMKL81, MRPK82, OMG82, PV84, SH82c, SSC80, VS82, dSTK81]. germinal [LS81b]. GH3 [SW84]. ghost [ML82, Sow84]. ghosts [JTM80, OK82, PF81, SP82b]. giant [Liz80]. gill [SSMH82]. gizzard [GDTS81]. gland [AMK83, AC82, DSKS84, GPR83, GLP80, HS80a, HE81, LP83b, MTH84, SSG^{+}83a, SMTP80, SHO^{+}84, SDK^{+}84, WOM^{+}81]. glands [BMCGO^{+}82, TDM80a, TS83a]. glia [WE80]. Glial [HSME84, ALE84, GE84, HLM84a, JVMRK82, MPAM80, SLS81, WE80, YF81]. glial-axonal [WE80]. glial-free [ALE84]. glioma [ASA83, AA83, HS81, WSM80]. globin [PD80]. globular [GC80, ZF84]. globule [FHG^{+}81]. globulin [Gri80]. glomerular [KLF80, KHF81, KR82a, KHJG84, KSF84, RKF80, RBO83, SC82b]. glomerulus [CKHF80]. Glucocorticoid [AA83, LKJ84, MG81, GB83]. Glucocorticoids [LBJ84, CMF82]. Glucose [LK84, AL84, GKZ81]. glucosylate [CB84b]. glucuronidase [GK84]. Glutamine [NDRL80]. glutaraldehyde [Leh81, MP83]. Gly [SGM^{+}84, SGM^{+}84]. Glycera [MTI80, MTM83]. glycerinated [CTKJ83]. glycerolipid [JB81d]. glycine [AH84]. glycocalyx [RBO83]. glycocomponents [TdS84]. glycoconjugate [CMC83]. glycoconjugates [PMP81a, PMP81b, VSLS80]. Glycogen [RKK82, GKZ81]. glycol [KLC83, RRD83, WSLSJ83, Wol80b]. glycol-induced [KLC83]. glycol-mediated [WSLSJ83]. glycol-resistant [RRD83]. glycolipids [MWA^{+}80, SC82a]. glycolytic [Mas84]. glycopeptide [KL81]. glycopeptides [CKJ83]. glycophorin [dST82]. glycoprotein [AWB82, BGD80, BKL^{+}82, BS83, CF84b, ED84, Fle81, FSBR83, FR81, GKKPB83, GFLG^{+}84, HJ81, JOA84, JOH^{+}84, KNBSF84, KKBG83, MBWH83, MS84a, MBP^{+}83, MS84b, MPF83a, ML80, OFO83, OFH^{+}82, PS83b, RMT84, RB80, RB82a, RBPF80, RF81, SGPB84, SWvK^{+}83, TFGC84, TQ82, TQS84, TQG84, WSH84, ZB82]. Glycoproteins [MDA83, AdS83, AdS84, ACC83, BHJ83, BO81a, BKO81, BvGJ83, GHB82, GZR^{+}80, GS81, GPDM81, HTKC82, HY82, JHCC84, KHS^{+}80a, LK84, PG82, RIP^{+}84, SC82a, SH81a, ZGN^{+}80]. glycosaminoglycan [LSC83, MMR^{+}84, TS83a, SGKD83]. glycosaminoglycan-binding [LSC83]. glycosaminoglycan-rich [MMR^{+}84]. glycosaminoglycans [GZR^{+}80, KLF80, MB84a, ZGN^{+}80]. glycosidases [CRH81, RCH81]. glycosomes [OBC^{+}84]. Glycosylation [GMRB81, ACC83, BVC83, GBWR83, GB81, KWSB84, MA83b, PBG80, PRK^{+}82, PB83a, PA83, Rot84]. Glycosyltransferase [GIKR84]. glycosyltransferases [BBP80, CB84b]. glyoxysomes [TE81]. GM1 [KPPC83, Pie82]. GMP [AM81, VVP82]. GMP-binding [VVP82]. goblet [BB82b, Rot84, SN80]. goblin [ML84c]. gold [DWP81, GSvdLS81, HAEC81, HLSWH82, KC83, Law84, SG81, dHLC84]. gold-labeled [dHLC84]. gold-labelled [Law84]. golgi [Bro81, BRS^{+}82, BKL^{+}82, BM80, BBP80, BO81b, CVT84, DR83, FP81a, Fle81, FR81, GCKAA83, GSH^{+}84, GIKR84, GBB^{+}82, GQW83, HSB^{+}82, HTGB84, HP82a, HP82b, LQ82, LK84, LRW82, ML84a, OHA^{+}84, PRK^{+}82, PE83, PBS84, PVPB82, QGW83, RIP^{+}84, ROB^{+}84, RS84, RB82b, RF81, RMU84, SBK^{+}84, SP84b, SDGYJ84, SWvK^{+}83, TV83, TZW83, TLPTV83, TIS^{+}81, VC84, WW83, WTSAC^{+}81, WP82b]. Golgi-associated [LQ82]. Golgi-derived [OHA^{+}84]. Golgi-enriched [TZW83]. Golgi-like [FR81]. gonadotropin [FS84, HBB82, HHPB83, Reb82]. gonadotropin-induced [FS84]. gonadotropin-responsive [Reb82]. gonial [SLL80]. GP140 [Car84]. gp85 [TFGC84]. gradients [LL82a, SBP81, VSD84]. granulation [SKKC84]. granule [BVAR83, BO81b, BSS83b, BK84, CLHA80, GS84a, GDH^{+}83, GZR^{+}80, GWWP83, RLP84, SGH^{+}83]. granule-specific [BK84]. granules [BSS83a, BSS83b, Cre81, DHS^{+}84, FQBN81, Fri83, GZR^{+}80, KMV82, MGK83, NDS84, PBFC^{+}84, SDC80, SSLB84, SD82, TDM80a, ZGN^{+}80, ZSM^{+}83]. granulocytes [BG83, CF83, JNS^{+}84]. greatly [Ino81b]. grids [WSB^{+}80]. ground [BR84b, HM84a]. group [GBL81, Kho82, KSD^{+}83, KST84, PF83b, WRK81]. group- [KSD^{+}83]. groups [DM83]. grow [EPM84]. growing [GBM84, PJ83, SP84c, SBGP84, SSP84b]. grown [DRF84, HPM^{+}83]. Growth [DBK80, ML84b, MMC83, SDM84, SBM84, SBVW82, SC84b, AP84, AA83, AHMS82, BLG^{+}80, Bar82a, BKT84, BM82b, BRAC83, BG82, BMH^{+}83, BPDR83, BVGS82, BBG^{+}84, CGA^{+}82, CARK81, CMC81, CFCS82, CGG81, CGG84, CH81, CHRT83, DH84, EKS^{+}82, FS80, GDRS80, GJRO^{+}83, GK81b, GGV81, GSH80, GCL83, GLS83, GL80b, GB80, GLBS81, HG80b, HHD83, HEG81, IY84b, IY84c, JDTW80, KC83, KOLV81, KKM81, KS82, LTBS80, LRT^{+}82, Let83, LR84, LH84, LA83, MBZB81, NRMB84, NCBP84, PPR82b, SMT80, SSE84, SBCS^{+}82, SHW^{+}82b, SRCG84, SGS^{+}82, SM82b, SM82a, SCS82, SCP^{+}83, SW81b, SK81, SC84a, TAGG80, TSGS83, VS82, VPMF84, VRR80, WP82b, YSS82]. Growth-inhibitory [SDM84, SBM84, NRMB84]. gruberi [FW80, LD81, Wal84]. Gryllotalpa [KW81a]. GTP [GBWR83, PBG80, PB83a]. GTP-stimulated [PB83a]. guanine [SPEL84]. guanosine [WB83b]. Guinea [MB82c, BF82a, Mad83, MHLS81, MP84c, MPK84, SSG80, SLSK84].

H [BC84b, WCZG84, WSB^{+}80, BDPS80, BF82c, CL81, EW82, FB82, GCKAA83, RBH^{+}84]. H-2 [EW82]. H-35 [BF82c, CL81, FB82]. h1 [AAW^{+}84, ACH^{+}81, SC81]. h1-like [AAW^{+}84]. H2B [dPMMF^{+}81, San81]. H5 [ACH^{+}81]. hadacidin [RJVS81]. Haemanthus [EJM82, JD82, MBB83, WHJ80]. haemopoiesis [SGKD83]. Hagfish [SDK^{+}84, DSKS84]. hair [HT82b, TDM80b, TS83b, TEDS83]. half [FG81, JL84, KHS82]. half-bivalents [JL84]. half-life [FG81]. Halobacteria [SS81]. halves [Fis82b]. hamster [AMCG83, AMS82, BHA82, BKB80, BS83, BS80, Cab83, CWBB83, CPS84, CK82, ERB83, GWP82, GQW83, Gri80, HJ80, KRB84a, KvRK^{+}84, KB81a, KB81b, LBM82, MMS^{+}84a, MG83a, MM81c, OMG82, RRRS82, RPM83, ROB^{+}84, SDL82, Sin82, SSC80, WC80, WWDYW83, ZHTH80]. handling [SEPB81]. Hantzschia [TPHL80]. happens [TEDS83]. haptoglobin [BJ83]. haptotactic [MF84b]. haptotaxis [MPF83a]. harvesting [DRF84, KTW83, MS80a, SBG^{+}81]. having [SC84b]. HD33 [GKZ81]. Head [BB80a, AdS83, IT82]. Head-to-tail [BB80a]. healing [FFGC84]. heart [DF84, Har80b]. Heat [LV81, Atk81, FP81b, IY84a, IY84b, MMS^{+}84a, MF84d, San81, SF84b, SSSJ83, VSBL83, WSM^{+}82, WWWC80]. heat-labile [WWWC80]. heat-sensitive [MMS^{+}84a]. Heat-shock [LV81, Atk81, FP81b, San81, SSSJ83]. heavy [BMF82, KWSB84, KH84, MBC80a, PKS81, PCB83, SP84a, SMG81]. heavy-chain [PKS81]. HeLa [AFK^{+}84, BC82, BC80, BB80c, DBBN81, GE80, MP83, MSGP81, MHRR83, PL83a, RB82b, SM82b, SC81, SVW^{+}83, WSM^{+}82]. helical [GB82, LGF^{+}84, WFR84]. hemagglutinin [RBPSB84]. Hemidesmosome [GGSB83, TRG84]. hemoflagellate [OBC^{+}84]. hemopoietic [SB83a]. Henle [EHK80]. heparan [AF83b, HOMR82, Ioz84, KLF80, LFGR82, LSC83, LLM82, RVH84, SK83, WHK^{+}84]. Heparin [MB84a, FMS^{+}84, vdWSCH81]. heparinized [KS82]. heparinlike [CARK81]. Hepatic [HAEC81, HP82a, BBP80, HP82b, PRK^{+}82, RA81]. hepatocyte [EWB83, HWM83, HM83b, MNST82, OFO83, RR80a, RH83, RH84a, RH84b, TOT^{+}84, WH81]. hepatocytelike [RRRS82]. hepatocytes [BJS81, BJG83, BO81a, CHH^{+}83, DH84, FCL^{+}82, ITT80, JKHT81, JH84, OYNT83, RRQ^{+}84, RGM^{+}80, SG84a, SY80, SHP83, Wei80, ZH82]. hepatocytic [BPM^{+}80a]. hepatoma [BJS81, BGD80, BE82, BHJ83, BCC^{+}83, BF82c, CMF82, CL81, FB82, FDSWW84, GJRO^{+}83, MG81, MNW81, O’B80, RVH84, SBF84, TKM83]. hepatoma-mouse [MNW81]. heteroclitus [EDK80]. heterodimers [Ste82b]. Heterogeneity [HP82b, TAC84, FML83, HW83, Rao81, RMF83, SC84c, WWCD84]. Heterogeneous [GBL81, CD84, MSGP81, PM83, BPM^{+}80a]. heterokaryons [EWG82, KLC83, LVL^{+}82, LZR82, MNW81, SLR^{+}83, Wri81, WA83, Wri84a]. heterologous [MKK^{+}82]. Heterotypic [GE84]. hexadactyla [KW81a]. hexanucleotide [YBG81]. Higashi [IBB^{+}81, NFG82]. High [ATW83, AdS84, KSD^{+}83, MLSW82, SF81, Ste82a, ASC84a, BR83, BR84a, BCSC82, BM82b, BL82a, BHP83, GBL81, GCH84, GLS83, HJ81, HSS82, HHS^{+}84, HS84, IY84a, IT82, JNS^{+}84, KST84, LV81, LCA84, MSRB84, NHF^{+}83, PS83a, Pol84b, RDY^{+}83, SCM83, STR84, SIM^{+}84, WRK81, dSK84, KMCAS80]. high-affinity [HSS82, JNS^{+}84]. high-mobility-group [KST84, WRK81]. High-resolution [ATW83, MLSW82]. high-salt-resistant [LV81]. high-yield [BHP83, KMCAS80]. Higher [ARHG84, ACH^{+}81, ABD84, IY84c, MB80b, RSP83b, TWWG84, WFR84]. higher-order [ACH^{+}81, WFR84]. Highly [GL83a, Fri82a, HSGD83, MMR^{+}84, SC84c]. hindlimb [RK81]. Histamine [HSS82, HBR^{+}84]. histamine-ferritin [HSS82]. histidine [LBA^{+}84]. histidine-rich [LBA^{+}84]. histocompatibility [SSE84]. Histone [HK80b, AHRG82, DBBN81, EB84, HAP82, dPMMF^{+}81, MGSC84, PL83a, SHHP81, San81, SC81, WA84, WNB82]. Histones [ERH82, ACH^{+}81, AAW^{+}84]. historical [ES81b]. history [DB81b, Por84]. HL [CF83, JB81c, MKD^{+}84, RCJS82]. HL-60 [CF83, JB81c, MKD^{+}84, RCJS82]. HMG [KB81c]. HMG-1 [KB81c]. HMG1 [WRK81]. HMG2 [WRK81]. HMWP [BM82b]. hnRNA [EH81, PC80, vEvV81]. hnRNP [PM83, CLJ^{+}81]. homeostasis [DTP84, TPR82]. homogenates [OT80a]. homologous [GL83a, LKJ84, MKK^{+}82, SHT^{+}83]. homology [KKP84b]. hoof [MM81a]. hooks [ME84]. horizontal [AH84]. Hormonal [SKS84b, BS81b, GPR83, SMTP80]. Hormonally [WP82a, CAL^{+}82]. Hormone [FJ82a, FJ82b, FJ82c, MF84e, APZL80, ASA83, AA83, Asc84b, AHMS82, BLG^{+}80, CVGM83, CFCS82, HMDP80, Har80a, HBS^{+}82, KKM84, SW84, SMT80, SHJA82, VKHH82]. Hormone-induced [FJ82a, FJ82b, FJ82c, MF84e, HBS^{+}82, KKM84]. hormone-receptor [Asc84b]. hormone-sensitive [VKHH82]. hormone-supplemented [CFCS82]. hormones [BHA82, GJRO^{+}83, Sch81b]. host [FBST84, Per81a]. hour [FG81]. HR [GLMW84]. HTC [MG81]. Human [ASG84, BBPR81, CKS82, FFEL81, HHD83, LNA81, VKE^{+}81, AKM^{+}82, ABG81, Asc84b, ASM^{+}84, BSG81, BSH83a, BW81, BW82, BLG^{+}80, Bar82a, BC84b, BTJ82, BBWS83, BHSC83, BFL^{+}83, BPB^{+}81, BG83, BZN82, CVGO81, CGA^{+}82, CKB^{+}80a, CLHA80, CLFC83, CMC81, CVAK80, DTP84, DHS^{+}84, EB83, EWG82, FH84, FDKF82, FLH^{+}81, FBST84, GDMG81, GPHV83, GK81b, GCL83, GV82, GLP80, HPM^{+}83, HG80a, HG80b, HFW82, HT83, HM84b, HBB82, HHPB83, HSCS80, Ioz84, JB81c, JNS^{+}84, KSMF84, Kho82, KMCAS80, KS82, KRB^{+}84c, LHS82, LVL^{+}82, LUT80, LWWP84, LA83, LA84, LA80b, MHSW81, MKW^{+}82, MD82, MKD^{+}84, MTKH84, MS84a, MLM83, MS83a, MBC83, MFVPK82, MDJ82, Nac80, NLN^{+}81, NFG82, NHF^{+}83, NBMD81, OFH^{+}82, OK82, PLP81, PF81, Per81a, PWT80, PJE80, Rao81, RCJS82, RVM83]. human [RSL81, RLT81, SGA80, SSE84, SHW^{+}82a, SDC80, SP84a, SB82a, SJS84, SCN^{+}81, SBB81, SSF84, SC84a, TJB83, TTI80, TTI^{+}81, TTI84, TKM83, WOM82, WM84, WG81, WMG84, Wat84, WP82a, WCZG84, WLFD84, WG82b, WMENS82, WLCS84, YKJF80, ZRW80, ZF81, ZFP84, dST82, vDB80, BLG^{+}80]. hv1 [WA84]. hyaluronate [GT84, MTT80]. hyaluronidase [OT80b]. hybrid [HLSWH82, KUM^{+}84, SKE84]. hybridization [HLSWH82, Jef82, SSZ83]. Hybrids [Nac81, HK80b, KUM^{+}84, MNW81, SG82b, Wri84a]. Hydra [HMM82]. hydratase [RFML84, RQHR81]. hydratase-dehydrogenase [RFML84]. hydrated [BBS81, SEPB81, SBP81, TH84]. hydrocortisone [ASA83]. Hydrodynamic [NW82]. hydrogen [PP80b]. Hydrolase [HB84]. hydrolases [RKP^{+}82]. hydrolysis [BSSR84]. hydrophilic [BTJ82]. hydrostatic [BSH83b]. hydroxylase [AT83, BJS81, GSH80]. hydroxystilbamidine [Liz80]. Hyoscyamus [Rag81]. hyperacetylation [MGSC84]. hyperosmotically [BO81b]. hyperproliferative [WES84]. Hypertrophy [BSS81]. Hypertrophy-induced [BSS81]. hyphal [HA80]. Hypomethylation [CWS^{+}80]. hypothalamo [BO81b]. hypothalamo-neurohypophysial [BO81b]. hypothalamus [DK84a]. hypothesis [EHS80]. hypoxia [LSJT83].

I.V [BPM^{+}80a]. Ia [SC82b]. Ia-bearing [SC82b]. Identification [Bur82a, BWV84, CC82, CBS83, CB84a, DWS84, DR82, GOKK82, HPM82, IRMM84, KSF84, LDF^{+}84, LC81, MTR81, MD80a, MB84d, Mur80, PK81, PDK83, PJE80, RWD84, RMJ83, San81, Sch81a, SF84a, SF84b, VVBZ84, WTK83, YAF81, BJS81, BTHL81, CCKMH82, DPS80, LSBPG84, MNP^{+}84, MSGP81, PS83a, SHHP81, TI80, HO84a, HM83b, LFS^{+}81, MWH83]. identified [Fro84, YSJ81]. identifying [Sch81b]. Identity [MHW83, MWH83, RQHR81]. IEF [BFL^{+}83]. IgE [MHB84, MLP84, MHB84]. IgG [AR81, RK84]. IGS [LdMM^{+}84]. II [AMM84b, BCHR83, BCSC82, BKO81, BO80, BSS83b, BF82c, CRH81, CF84b, CL82, DK81a, DHHG83, DPS80, FO80, FJ82b, GKKPB83, GWB82, GSS82b, GCS81b, GAAG82, HBO84, HHS^{+}81, HM83b, HS84, JB82, JRB80, KJBR82, KKP84a, KP84, KC82, KMF84, KMP^{+}82, LGHI84, MACK84, MCPB80, MBLvdM84, MP82, MSC80b, MWH83, NSI82, NCM^{+}83, PST^{+}83, PS83b, PBS84, PMP81b, Pol82, QGW83, QGW84, RH84a, SWGB80, SBP81, SHW^{+}82b, Sie81, SSSP81, SBGP84, SDK^{+}84, SBM84, SBVW82, TI82b, TEDS83, TKA83, WB81a, Wal82b, WW83, WB82, WOBR80, ZGN^{+}80]. IIb [SGPB84]. III [ARL^{+}80, BPM80b, BBS81, CGS81, CDvdM80, DPK84, DMD80, EWG82, FGHC80, FJ82c, GZR^{+}80, GHL84, HSGD83, JSC84, KOB82, MSC83, OT80a, RH81, RH84b, SBG80, SDT81, SSP84b, TPHL80, WB81b, Wei82, WH83]. IIIA [MLZD83]. ileum [GN84, LAAS84]. illuminated [Axe81]. image [HA84b, Ino81b, MS84e, SSL^{+}84a, TMRT84]. images [SEM80]. imaginal [Fri82b]. imaging [AO84]. imbalanced [CML^{+}81]. imidoesters [RKM^{+}81]. iminodipropionitrile [PAGG81, PYC^{+}82, TQG84]. immature [BL82b]. immobilization [ABSC81]. immobilized [CL80a, SB82a, SF84b]. Immune [GWP82, Mar80, MK84b, MP84b, WLFD84]. immunization [PBPD83]. immunoadsorption [MFP82, MBK82, RH84b]. immunoaffinity [WSB^{+}80]. Immunochemical [BMF82, RQHR81, LFAG84, TRC^{+}84]. Immunocytochemical [AMM84a, AMM84b, BJS81, BDH82, BMS83, CBWW80, CBL84, DF84, FBH83, GSS^{+}82a, HM84e, KKG84, LUT80, LAAS84, MRPK82, NCP84, NLN^{+}83, OWPC82, RB82b, SFS81, SLK^{+}84, TRN^{+}82, TIS^{+}81, TQS84, WWWC80, CSHP84, LNA81, LS81b, LFS^{+}83]. immunocytochemically [EW83]. immunocytochemistry [DHHG83, FLM84, ZB83]. Immunoelectron [BS83, CS82, DBG^{+}81, GDTS81, TDS83c, TDS83b, GSvdLS81, IDF83, LDM80, MEKT84, MDA83, PPRR82, SG81]. immunoelectronmicroscopic [SHW^{+}82a]. immunoferritin [JSMT82, JSCM83, RMAF80, TOT^{+}84]. Immunofluorescence [BB80c, MNW81, WSO^{+}81, WD83, BSVS80, BF82b, DCG83, FN80, FFGC84, LFAG84, MF84a, PBFC^{+}84, RSB84, SKKC84, TS81, YMF84, DTT81, TMS84]. Immunofluorescent [BLS^{+}80, JOM80, NCH83, SP82a, KC80b, MDM^{+}81, PPR82b]. immunoglobulin [GA84, HTGB84, HDRK83, KWSB84, LS81b, MHRR83, Mec81a, Mec81b, MHB84, NAK82, PATR82, Rod80, SFU84, SHB80]. immunoglobulins [SDT81]. immunogold [LdMM^{+}84]. Immunohistochemical [HPG83, ACMH83, DGS80, GRR^{+}84]. immunohistological [Ekb81]. Immunolocalization [LV83, WOM82, WMENS82, SHHR84, WKA^{+}83]. immunologic [CBP84, FSW^{+}83]. Immunological [KSK^{+}82, ONO82, PST^{+}83, RB82a, PS83b]. immunologically [PPR^{+}82a]. Immunoprecipitation [MB84c]. Immunoreactive [SS84a]. immunoreactivity [EBS84, LSN^{+}84]. immunostaining [MK84b]. impaired [CWBB83]. impairing [GNGTS80, HH81b]. impairs [ROB^{+}84]. imperfecti [AB81]. implantation [GAS^{+}83a, PAS84, PBS84]. implanted [MBP^{+}83, PS83b]. Implications [Kir80, HHS^{+}84, NBM82, PWM81, Rot84, SB80a, Sch81b, WL80]. import [FNHSS84]. Importance [SR84a]. Improved [MP83, MF84a, SF84b, YMF84]. IMR [NBMD81]. IMR-90 [NBMD81]. in-vitro-assembled [SD82, WIB81, WB81a]. inactivated [OKA84]. including [FSVW80, LSC83]. incomplete [SDT81]. incorporated [WE80]. incorporating [MMLT84]. Incorporation [GPDM81, SSF84, AH84, BO81a, BKO81, CKFS82, Gla83a, SKMA82, SMAS81, SHH82, UAL84]. increase [BSS81, BL83b, FBP82, JFS84]. Increased [BS82a, KLF80, CK82, CS83, GL80b, SC84b, VVBZ84]. Increases [Mad83, BE82, LBM82, MBZB81, SW84, WS84]. incubation [GBWR83, PBG80, WDE83]. independent [ASG83, BPM80b, CFW84, CF80, HJ81, KB81b, LTBS80, Pfe84, PGD^{+}84, TWG83, TD83]. indicator [TPR82]. indigestible [BO83]. indirect [MDM^{+}81]. individual [BW81, Ber80b, KKM84]. induce [Cre81, KK82, MMR^{+}84]. induced [AFW84b, APZL80, BSR84, BSS81, BMF84, BMH^{+}83, BCF84, BWM84, CKJ83, CMC81, CP82a, CQB84, FA82, FJ82a, FJ82b, FJ82c, FS84, Fuk80, GG80, GMO84, HBS^{+}82, HFW82, HEG81, KML83, KLC83, KKM84, KRB^{+}84c, LP83b, LMN83, LTGL80, LFF^{+}81, Mad83, MF84c, MF84e, NDRL80, Rag81, RRRS82, RKM^{+}81, SHHR84, SM82a, Sow84, SY83, SCSC81, TdS83a, TM84, VPMF84, WWPL84, WMG84, WC84, WP82a, WKH80, YF83, ZCH80, BG83, JHCC84, KMF81, LN82, MKD^{+}84, PS82]. Inducers [RSP83b, PLP81, RS81]. induces [AFSF^{+}81, ABSC81, BP82, GSH80, JVMRK82, KSSP83, LK83, MGSC84, MLNK83, PG82, PT83, SSBB82, SNPE84, SP84a, SM82b, SCM83, SM83c, WLCS84, YSS82]. inducibility [Rao81]. inducing [Sch81b]. Induction [CL81, JZ81, Kan80, LH83, RRQ^{+}84, SSG80, Wri84b, ZS84, BHSA82, BL80a, CFN^{+}83, ELST81, FK80, MMO^{+}81, MMS84b, NCM^{+}83, STRP83, SY80, KS82]. Inefficient [ABG81]. inexcitable [HSTK83]. infected [AMJ^{+}81, BS83, GQW83, GAS83b, LBA^{+}84, NPG^{+}80, QGW83, RIP^{+}84, SSZ83, TMW82, WWES83, MTW82]. infection [ML80, Mil80, OK82, WWDYW83]. Infectious [MRHS81]. inflammation [RKK82]. Influence [CGM81, SB82b, BHA82, BBF84, BG83, GBF84, KC82, LKJ84, MS84c, SGW^{+}82, Sma81]. influenced [EN84]. influencing [CB81, WGF83]. influenza [MRHS81, RBPSB84, WMH81]. influx [HH81b, MM81c]. inheritance [BS80]. inhibit [GGS83, Her80, KHS^{+}80a, LTGL80, PCB83, SNS83, VPMF84]. inhibitable [BB82a]. inhibited [HBR^{+}84, HMM82]. inhibiting [HS80b]. Inhibition [CMF82, CS83, EHESF82, KP84, LR84, MR80b, ML80, Oku80, OLH^{+}81, OMG82, PL83b, RJVS81, WKW82, ZHTH80, AH84, BZN82, Fuk80, GS83, GSK83, GL82, KKP84a, LKJ84, MF84c, OBH84, SMT80, SRCG84, SGPB84, VRR80, WWDYW83, Wil84, SDM84]. inhibitor [BDP^{+}84, CARK81, EPO^{+}84, FB82, FS84, Liz80, MBZB81, SBM84, TS83a]. inhibitors [ASWR83, KOS83, YT82]. Inhibitory [Per81a, ASA83, NRMB84, SDM84, SBM84, SBVW82]. inhibits [BS82a, BSP83, Bar82a, BDP^{+}84, EH81, GNGTS80, GGV81, GQW83, HWAK83, HCH^{+}81, HM84b, KL81, MP80, MKB82, ML83, SSBB82, SBC82, Vah83, WCZG84]. inhition [Mil80]. inihibition [Bro81]. Initial [AW82, BM82b, GR80, MM82, RGE83]. initiate [ECGvdRP83]. initiated [MTMT83]. Initiation [HR80, RF84, BR84c, EH81, ES82, LR84, LFR80, WR83, Mec81b]. injected [HSRL81, MHRR83, ML84a]. injection [BO81a, BKO81, KNHK84]. injured [GL82]. injury [SB80b]. inner [ONO82, SHH82]. Inositol [AH84]. insect [SB80a]. insertion [BPM80b, Kan83, PG83a]. insoluble [DYP83, FHG^{+}81, GSMHF84, Gri80, TFGC84]. insufficient [Mar80]. Insulin [GMO84, KCM83, ORA^{+}84, BBB82, FCL^{+}82, HWK^{+}84, HBR81, MPC84, PSBF83, PVPB82, SSF81, CVGO81]. Insulin-induced [GMO84]. insulin-resistant [HWK^{+}84]. insulinoma [SF84a]. intact [BBCL82, FC82, FJ82a, HS83, LCA84, MMJ^{+}81, MPF83b, Pie82, SHJA82, TPR82]. integral [BM80, BPM80b, GR80, IMB83, MCPB80, PG83b, PG83a, RBS84, RH82, TFGC84, LGHI84]. Integrated [BPM^{+}80a]. integration [AWB82, GB81, MBC^{+}84, TWWG84]. Integrity [BS81c, DGC83, EWB83]. intense [FGHC80]. intensification [HA84b]. interact [HLM84a]. Interaction [LOC81, MBR84, MMJ^{+}81, MG83b, PSM84, PFP^{+}82, SMS84b, SSE84, WP80, DGD^{+}84, Gri80, GS80b, KML83, Mab83, MB84b, MCN^{+}84, MBLvdM84, MMVO83, OT82, PFM^{+}84, PJE80, SvB81, SCP^{+}83, SSF84, STR84, TQG84, YPR^{+}82]. Interactions [BPDR83, HT82b, LLS82, MBWH83, Mas84, MW80, MRB84, AW82, CFN^{+}83, DKDB81, Den80, DPS80, DSKS84, GJRO^{+}83, GDTS81, GAR^{+}84, NCM^{+}83, PG82, PF83a, Ped83, PWT82, Rau84, RGE83, SKS^{+}84a, SSP84a, TDM80a, WM81]. intercalated [GS84c]. Intercellular [LS82, MYR81, GS81, MF84e, SR81b, HS84]. Intercompartmental [RMU84]. Interconversion [Kan83, KNHK84]. interdigitating [AFSF^{+}81]. interferon [NRMB84, PWT80, WMP^{+}84a]. interleukins [MMR^{+}84]. Intermediate [FSF^{+}80, SJG84, WB84, ASG84, BSS81, BV83, BFL^{+}83, BRS81, CS83, DF84, DSKS84, EDP82, FSB81, FWP84, GMVF84, GV82, GRL82, HTFH82, IDF83, MS83a, MM81a, Mil81b, MG83b, NW80, PLMS81, PL83c, SSG^{+}83a, SDK^{+}84, SHT^{+}83, SCSC81, TDS83c, TDS83b, Tok83, WCY^{+}83, WRM83, WW83, Woo80, YF81]. intermediate-filament [CS83]. Intermediate-sized [FSF^{+}80, WB84, BFL^{+}83]. intermediates [MS84d, RWD84, SPS^{+}84]. Intermittent [Gib80, DRF84]. intermittent-light-grown [DRF84]. intermolecular [Car84]. internal [Axe81, LWW^{+}83, LW81]. Internalization [Fis82c, HT83, MPS^{+}83, MP84b, MPU84, Rud80, ABG81, BLG^{+}80, BRR^{+}83, ELS^{+}83, GSPS82, HAEC81, KC82, SSF81]. internalized [FCL^{+}82]. internally [BO83]. interphase [BLV84, FSW^{+}83, KNHK84, KKMK84, KNK84, OM83, SS80b]. Interplay [WR83]. interrelations [GO80]. interrelationship [PL84]. interspecies [SKE84]. interspecific [HSTK83]. intervals [O’B80]. intestinal [BB82b, BM83, Bur82b, CL80b, DGS80, FWG^{+}81, GDTS81, GGW83, HH81a, HTFH82, HM83a, Mad83, MB82a, MGW^{+}80, MPW82, PFM^{+}84, Rot84, ZSE80]. intestine [AR81, Rod80, UAL84]. Intra [HSTK83]. Intra- [HSTK83]. Intracellular [BBR83, BKL^{+}82, Cle84, FNS81, GE80, HT80, JE81, MWP84, MHRR83, NLN^{+}81, RBPSB84, SSJS84, SHP83, WKAH84, AMK83, BR83, BP83, BS83, BALR83, Bro81, Bur82a, BR84c, DGC83, FHFL82, GOSW82, GS84d, GBB^{+}82, GWQ^{+}84, HWAK83, HM84a, HF83, JFS84, KHS^{+}80a, KSR^{+}80, KC80a, KKM84, LS81a, LAAS84, LWWP84, LLB84, ML80, NM84, PST^{+}83, PLMS81, PR84, QGW84, RIP^{+}84, RKK82, RBS84, ROW^{+}83, SP84a, SO82, SLRP82, TTGA82, TSS82, TdS84, WM84, WN81, YAF81, ZH82, GZR^{+}80, Wal82a]. intracellularly [TPR82]. Intralysosomal [KSC82, KC82, MSC80a, PO81]. Intramembrane [KW81b, FGHC80, MB80b, MKD80, PT83, PF83b, RV80, RRD83, SHH82, SP84c, dLTvdSB83]. intramembranous [AO84]. intramolecular [Lan82]. Intranuclear [RN83]. intraorganellar [LLS82]. intrinsic [BDH82, FBH83, LNA81, LAAS84, PR83]. introduced [GS84b]. introducing [vMS83]. Introduction [PB81]. invasion [AMRE81, FBST84, HHS^{+}84, Per81a, WB82, Wei82]. invertebrate [Wal82a, Wal82b]. invertebrates [LC80]. inverted [KORBS82]. investigation [LPH83]. Investigations [DRF84]. Involement [ZCH80]. Involucrin [BSG81, WG81]. involved [AF83a, CL80a, DWS84, EK84, GO80, GSOG80, HS80b, IMS82, LC81, LA83, MB84b, MLD82, PvD83, QCB84, WC80]. Involvement [GK81c, OK82, BM82b, DS80, KSD^{+}83, SKS82, SSC80, YRS83, PAS84, PBS84]. involving [Wri84a]. iodide [CVGM83]. iodinatable [WG82a]. iodination [MSUC80, MSC80a]. Ion [KRN83, KCGH84, MR82, MM81c, Pfe84, SGKC82]. ion-flux-independent [Pfe84]. ion-selective [MR82]. Ionic [BMH^{+}83, LP83b, WHJ83, SGSS^{+}81]. ionophore [ACC83, LUT80, SLRP82]. ionophore-mediated [SLRP82]. ionophores [Max82]. ions [AMK83, DTP84, EWG82, Kie81, MR80a, PP80b]. iris [BKT84]. iron [KvRK^{+}84, LTE84]. irradiated [SKN83]. irradiation [BCHG83, KCB84]. irradiations [LPH83]. islet [FQBN81, MKK^{+}82, MMAO83, NM84, NDS84, SF84a]. islet-like [MMAO83]. islets [KCM83]. isoactin [RRS82, WBG83]. isoelectric [BFL^{+}83, EBS84]. Isoenzymes [PLP81]. Isoforms [RMS^{+}82, DSRF84]. isolated [AC81b, BJS81, BC84a, BFW80, Bre83, BBCL82, CGM81, CKJ83, DRF84, DCK84, ECGvdRP83, EHESF82, FFRF83, FQBN81, GHWR80, GIKR84, GS84c, HF80, HMG80, HC83, IMB83, ITT80, KHF81, KCM83, KJBR82, KSSR84, KHL82, LHS82, LP83a, LSJT83, LOC81, LMYM84, MS80a, MSRB84, MSPF83, NvIvGdH^{+}82, OMM82, PL83b, PM81, POS80, RH84a, RBMM82, SS80a, SG84a, SBBS84, SHWS83, TEL84, VKHH82, WZ80, WTSAC^{+}81, ZSR^{+}82, ZH82, ZP80, ZSE80]. Isolation [AHMS82, Blo84, BL82a, Cab83, CGCS80, CMC83, CNSW80, EF84, FBB82, FSBR83, FWG^{+}81, FHFL82, GS81, HJ80, HWM83, HM83b, KG80, KDL80, KL81, KMV82, LD81, Liz80, MW83b, MAG82, Mil81b, MMIGK82, MM82, NW83, OM82, RSL81, SP82b, SIM^{+}84, TIK81, ZGJ^{+}84, ZHM80, ASG84, CKE81, EHK80, KC80b, MFP82, PADH83, PL83a, PD80, RGM^{+}80, SBSM84, SG80, Val82, VVBZ84, WC80, BSS83a, Den80, GWB82, KPG^{+}82, RH84b, SDK^{+}84]. isometric [CGW82]. isoproterenol [HS83, MTH84]. isozymes [GLBH82, GBLH83, KKP84b]. IV [BCHG83, CEC^{+}83, DWS84, DSD80b, FGS^{+}82, FML83, LLM82, LGF^{+}84, MR80a, MWI^{+}84, MR81b, RMAF80, SGKD83]. IX [ERN83].

J774 [MKB81, TEL84]. Journal [PB81]. junction [AKT84, BCHG83, BH83, Bur82a, CB84a, CH80a, FG81, GS84c, GDRS83, HLS81, HR81b, Ko84, Mad83, MH84, PP80a, PP80b, PT83, SH83, UE83, WE80, ZCR80]. junctional [AMJ^{+}81, CC82, FSFA84, MTH84, SNdC^{+}84, SG84b]. junctions [BDH82, BBCL82, EWB83, EDK80, FGHC80, Fri82b, GG83, HAFG82, KG80, LC80, LS82, MPMBC80, MIS^{+}80, MDA83, NB81, PTH83, RGR80, SSHP82, SR81a, SB82b, TdS83a, TT83, ZSR^{+}82]. juxtaposed [WE80].

K-transport [SMJ80]. K562 [ELS^{+}83]. karyophilic [FKS84]. karyoskeletal [KSK^{+}82]. KB [GMO84]. kdalton [BM83, GBE84, HM83a, NS83a, WES84, CL80b]. Keratin [BS82b, KOLV81, AFRS83, BFL^{+}83, CLFC83, EDP82, KML83, KOS83, LVL^{+}82, ME82, MFVPK82, NS83a, SF84a, SHT^{+}83, SCSC81, THT^{+}84, WES84, WMENS82, ZGJ^{+}84]. keratin-like [ZGJ^{+}84]. keratin-type [BFL^{+}83]. Keratinization [HDB82]. keratinocyte [PR83]. keratinocytes [AKM^{+}82, BSG81, WMG84, WES84]. keratins [BFL^{+}83, EBS84, KMF84]. kidney [CFCS82, CAL^{+}82, Ekb81, ES81a, GLMW84, GQW83, Gri80, HER81, HOMR82, HEO82, KKG84, KNBSF84, MRFF80, MHGM82, MRHS81, MBP^{+}83, MG83a, MSB^{+}83, OMG82, PS83b, PAS84, PBS84, RKF80, RIP^{+}84, RKMF84, RBPS83, RBPSB84, RMAF80, TFAP81, VKHH82, WWDYW83]. kidneys [SKE84]. killifish [EDK80]. killing [MGLF83]. kilodalton [OKA84]. kinase [BNB^{+}82, BF82c, CH82, FB82, GCF^{+}83, GCH84, GCS81b, GBE84, IMS82, LLS81, MF84e, MSD82b, PGD^{+}84, Sch84, SDG^{+}84, Ste83a, VDT81, WDE83]. kinases [BLS^{+}80, CFW84, NCH83]. kinds [ZF84]. Kinetic [ZSL82, BO80, BAWL81]. Kinetics [MS84d, SDC80, BF82c, CRH81, CHCS82, GGK82, TI82b, VVP82, VV84]. Kinetochore [BPB^{+}81, CL82, EHCR84, EM81, ERB83, HWS82, Jen82, MB81, MHNT81, SB80a, TPHL80]. kinetochores [Hea80]. knobs [GAS83b, LBA^{+}84].

L [CL80a, EZ84, VN82, WHSK80]. L-fibroblasts [VN82]. L1 [NNRR80]. L2 [PG84]. L2-deficient [PG84]. L929 [LC81, LK83]. Label [dSK84, AdS83, CL80a, FN80, TdS84]. Label-fracture [dSK84]. labeled [FB80, GS84b, HP81, KFS81, LSM^{+}84, MLSW82, PLMS81, SSL^{+}84a, SLS^{+}84, SC82a, SMS84a, SMS84b, SKS^{+}84a, TOG83, TWH80, dHLC84]. labeling [EH81, FG80b, GSvdLS81, JSMT82, JSCM83, PMP81a, PMP81b, dSK84]. labelled [Law84]. labile [DM83, SS80a, WWWC80]. lacking [HWJW84, KKMK84, WOBR80]. lacrimal [BMCGO^{+}82]. lactamase [KFV83]. lactate [PLP81, Wie81]. lactating [FHG^{+}81]. lactogen [HBB82, HHPB83]. lactoperoxidase [CL80a]. lactose [BB82a]. lactose-inhibitable [BB82a]. laevis [AH84, CML^{+}81, ERH82, FKS^{+}81, GWK84, KSK^{+}82, LS81a, MLNK83, RB83, SP84a]. laggards [JL84]. lamella [DH81]. lamellae [SS84b]. lamellar [BT84]. lamellipodial [ML84b]. lamin [LDF^{+}84]. lamina [AF83b, AKT84, BWW^{+}80, DB81a, FBB82, FPFP81, GW82, KBDK84, LLM82, MS84c]. laminas [PB84a]. Laminin [LDV83, MER^{+}83, MF84b, San82, ACMH83, BF84a, CHRT83, DWS84, GAA83, HER81, HOMR82, IER84, JKHT81, LLM82, MRFF80, MW83b, MPF83a, PF83a, WWCD84]. lampbrush [Sch81a]. lapse [HA84b]. large [CMC83, CF84b, KMF84, KSD^{+}83, MHB84, RBMM82, ST83, TPHL80, UM82]. large-scale [MHB84]. Laser [KCB84, AB81, GK81b, MB81, SLS^{+}84]. late [HH83b]. Lateral [EW82, EZ84, JOA84, MLP84, EWG82, HSFM82, Oli82, PT83, TWW82, WZ83, dLTvdSB83]. Latex [BF84a, Gri80, GML^{+}80, MG83a, TEL84, VKE^{+}81, vdWSCH81]. laticollaris [TA81, TKA83]. latipes [Gil83]. latissimus [RMF83]. latrotoxin [VMMC84]. layer [CH80b, GOA83, Wat84]. layers [HPG83, MLM83]. Ld [EZ84]. lead [IY84c]. leading [DH81, Sma81, WH82]. leaf [NHMT84]. Lectin [PMP81a, PMP81b, TV83, BHR81, BB82b, CHRB84, DR82, EW83, FBP82, GGVR82, GSMR83, MP80, RB83, SSP82a, SHB80]. Lectin-binding [TV83]. lectin-resistant [FBP82]. lectins [BB82a, CRH81, RCH81]. leg [OMM82]. Legionella [HM84b]. Length [WD80, ARHG84, HWS82, SHT^{+}83, TG83, TS83b]. Length-tension [WD80]. Lens [GDL80, BDH82, FML83, FBH83, HSME84, HO84a, HBO84, HAFG82, NGL83, PG83b, PG83a, SG82a, ZSR^{+}82]. lesion [TJB83]. less [SG82a]. lethal [MMS^{+}84a]. Lettré [GKZ81]. leucine [RH83, RH84a]. leucyl [NFS81]. leukemia [MKD^{+}84, MLP84, MHB84, NSI80, NSI82, TTGA82]. leukemic [BC84b, PLP81]. Leukocyte [MMR^{+}84, RLT81, Zig81, ZLK81]. leukocytes [BSS83a, BSS83b, DLZ83, MD82, NFS81, NFG82, RKK82, RVM83, SDC80, SH81a, SCN^{+}81, SZ80, SDZ84]. leupeptin [HF83, OLH^{+}81]. leuS [WC80]. level [BK84, HLSWH82, LBM82, SBGD84, TOK^{+}84]. levels [BB82c, DGS80, GCH84, KW84, RSP83b, SM83c]. Lewis [SLK^{+}84]. Leydig [LA83, MC80, Reb82]. Lidocaine [SBC82]. life [FG81, GHGT81]. ligament [MGMS84, MMS84b, SMJB83]. ligamentum [MLMS81]. Ligand [ROW^{+}83, DH84, Gol82, JNS^{+}84, KOB82, MS83a, WKAH84]. Ligand-dependent [ROW^{+}83]. ligand-receptor [JNS^{+}84, KOB82, WKAH84]. ligands [CKS82, SFU84, WLCS84]. ligation [WLCS84]. Light [NHMT84, Sch83b, BBF84, BNH81, BNB^{+}82, BHP83, COS83, DRF84, DGS80, KTW83, MF84a, MS80a, Nac81, RKH82, SBG^{+}81, SDG^{+}84, SS80c, WP80, WD84, Wri81, WA83, WB83b]. Light- [Sch83b]. light-activated [MF84a, RKH82]. light-chain [COS83, WA83]. light-harvesting [DRF84, KTW83, MS80a, SBG^{+}81]. Light-regulated [NHMT84]. light-scattering [WB83b]. like [ALE84, AAW^{+}84, BHP83, CVAK80, DPS80, DANH84, FR81, HSC84, IRMM84, KSD^{+}83, MKD^{+}84, MMAO83, MOV83, PFM^{+}84, POS80, WPKB83, ZGJ^{+}84]. limb [EHK80, SHPH84, ZS84]. limited [KKP84b, PSSS82, RF80a]. limulus [BM83, DTBF82, KL82, OR81, WD80]. line [BC84b, BK84, DCS84, GJRO^{+}83, GSH80, HWK^{+}84, HEO82, HBR^{+}84, HCH^{+}81, JB81c, KWM82, MKD^{+}84, MRHS81, MPF83a, MSB^{+}83, NRMB84, NPG^{+}80, OWPC82, OWMB82, PJM^{+}81, Reb82, RCJS82, RVH84, SSG^{+}83a, SRCG84, SC84b, SKA^{+}83, TEL84, TKM83, WWKW83, WOM^{+}81, ZSM^{+}83, vdWSCH81, YRS83]. lineages [GK81d, GK82]. linear [HWS82]. lines [FSBR83, GB83, GGK83, WC80, WDE83, YRS83]. link [FT80, PPRR82, PPR^{+}82a, PPR82b]. Linkage [WC80, SH81a]. linked [CF84b, DCS84, DWS84, DR83, PDK83, WC80, EK84, LK84]. linker [Hir82]. linking [CC82, DPS80, EP80, LFGR82, MGW^{+}80]. links [CV82, GGW83, HSC84]. Lipid [HSFM82, PB83b, BF82a, BGM81, DHS^{+}84, HSCS80, KKHK82, LFC^{+}80, MHC80, Mil80, PLMS81, SJM81]. lipid-binding [LFC^{+}80]. lipids [BR84a, CB84b, CH80d, vMS83]. lipophilic [LTE84]. Lipoprotein [FMS^{+}84, ABG81, ABBG82, BR83, BR84a, BW81, BW82, CGA^{+}82, GSPS82, HAEC81, HP82b, MFA83, RR80a, SSG^{+}83b, VSS83, VVBZ84, ZRW80]. lipoprotein-gold [HAEC81]. Lipoprotein-heparin-fibronectin-denatured [FMS^{+}84]. lipoprotein-receptor [BW82]. lipoproteins [PS83a]. liposomes [MS80a, MS83b, OLHH81]. lipotropin [GLP80]. liquid [IBB^{+}81]. live [Gib80, MB84c]. liver [AC81b, ACTK84, BR83, BR84a, BKO81, BRS^{+}82, BRR^{+}83, BM80, BPM80b, BBG^{+}84, CSH81, DH84, EHESF82, FG81, Fle81, FFS^{+}82, GSS^{+}82a, GSS^{+}84, GF84, GBWR83, HAFG82, KKG84, KORBS82, LSJT83, MACK84, MAG82, MMO^{+}81, MEKT84, MBR84, MSRB84, MCPB80, MYR81, MP81, MP82, NLN^{+}83, PBG80, PC80, PVPB82, QCB84, RVB80, RFML84, RACGB81, RQHR81, RH84a, SSZ83, SDG^{+}84, SBBS84, SHWS83, SG84b, WH81, WTSAC^{+}81]. living [AUT83, BMF84, FB80, Gla83b, GK81b, HA84a, HHSS81, HHS^{+}81, HSC80, JWBC81, MSGP81, SWA83, SKS^{+}84a, TWH80, Wan84]. lizard [BBF84].

LLC [DVB^{+}81]. loaded [KKM84]. loads [Mad83]. Localization [BM80, BPM80b, CW80, DSRF84, DGS80, FN80, GRDR82, GLP80, JSDM82, KB81c, LSM83, LLM82, LDM80, LS81b, MSB80, MCPB80, MSD82b, OMM82, PB83a, SWP81, SC82b, WH80, YMF84, ZB83, AMM84a, AMM84b, AM81, BDH82, BHSC83, BGPB80, Bre83, BSS83b, BLS^{+}80, BB80c, BMS83, BMCGO^{+}82, BF82c, CGA^{+}82, CSHP84, CVGM83, CBWW80, CHH^{+}83, CBP84, CKHF80, DK81a, DBG^{+}81, EHS80, ES81a, FBH83, FB82, GSS^{+}82a, GSS^{+}84, GFLG^{+}84, GWP82, HP81, HTKC82, HBB82, HM84e, HPG83, Ioz84, JOM80, JSMT82, JSCM83, KKG84, KKP84a, KC80b, KSK^{+}82, LdMM^{+}84, LUT80, LAAS84, LFS^{+}83, MRFF80, MDM^{+}81, MHGM82, MEKT84, MA82, MRPK82, NLPH82, NCH83, NCP84, NDRL80, NLN^{+}83, OFO83, OWPC82, PST^{+}83, PLMS81, PF83a, PG83b, PKS81, PPR82b, RB80, RH83, RH84a, RMJ83, RB82b, Rot84, SBG80]. localization [SBK^{+}84, SWF84, SHJA82, SLK^{+}84, SSMH82, TOT^{+}84, TFAP81, TRN^{+}82, TDS83c, TDS83b, TIS^{+}81, VMMC84, VKE^{+}81, WM84, WWWC80, dSTK81, FJ82b]. Localized [FS80, MP84c, BCHR83, BBLY83, CHRB84, CVAK80, EK84, FTM82, HSME84, KDMS83, MPK84, RB83, WYK^{+}84]. located [KHS82]. Location [CCKMH82, RFUC84, FHG^{+}81, SHO^{+}84, WSH84, WCJ83]. locations [CH82, MCPB80]. locomotion [CRGS82, GAR^{+}84, HSW82, HM84d, NRW82, RFUC84]. locus [SEE84]. Loligo [GS84b, LOL81]. long [BTHL81, ECM81, Kla80, RGM^{+}80, SCP^{+}83, SGKD83]. long-term [BTHL81, Kla80, RGM^{+}80, SCP^{+}83, SGKD83]. longitudinal [WRM83]. longitudinally [Tok83]. longitudinally-oriented [Tok83]. loop [EHK80, SPBB84]. lose [GH82b]. Loss [SC82a, BGE84, HER81, KK82, PADH83, SM81]. Low [BL80a, LP83a, PL83a, ABG81, ABBG82, BCHG83, BW81, BW82, CGA^{+}82, EDK80, GSPS82, GSG82, GBF84, GBG82, GE80, HAEC81, HHS^{+}84, HS84, LBM82, MOS84, MFA83, QCB84, RR80a, SMJB83, SO80, SSG^{+}83b, UE83, VSS83, VVBZ84, WKH80, ZRW80]. low-density [VSS83]. low-irradiation [BCHG83]. low-resistance [EDK80]. Low-temperature [BL80a]. Luminal [Bur84, HBS^{+}82, MKD80, RS83, SSP81, SSSP81, SSP82a]. luminescence [TBR80]. lung [BVGS82, CHRB84, DHS^{+}84, LBM82, SLK^{+}84, VB81]. lymphatic [OJJ^{+}84]. lymphoblastoid [SCS80]. Lymphocyte [HSW82, HRBJG80, Ric80, SAW83, BNB^{+}82, FC82, GA84, KW81b, Mar80, SDM84, SR83a]. lymphocytes [BLG^{+}80, CVGO81, DTP84, GRB82, HOR84, HDRK83, KPPC83, LH83, MF84d, PATR82, RSB84, RTP82, SKWF84, STR84, TPR82]. lymphoid [GB83, HLM^{+}84b, SDM84, SBM84]. lymphoma [EK84, GGK82, MMC83, Ste83a, BNH81]. lypolysis [SJ80]. lypolytic [SJ80]. lysed [BSNP82, CMM81, DB83]. Lysosomal [Asc84b, GGK82, AF84, BLG^{+}80, DS80, FNSB80, GNGTS80, KSR^{+}80, KC82, MTMT83, MBR84, NLN^{+}83, RBH^{+}84, RKP^{+}82, SCN^{+}81, SPS^{+}84, VN82]. lysosome [GHY81, HMM82, KC80a, PWSW80]. lysosomes [Bai81, GK84, GDMG81, KSC82, MS83a, MSC80b, OP81, PAS84, SDT81]. lysosomotropic [MGLF83, SBF84]. Lytechinus [SS80a].

M [Bar82b, EPRS81, GWK84, SPHE80, WDE83]. M-lines [WDE83]. M-phase-specific [GWK84]. M-protein [EPRS81, SPHE80]. machinery [MD82]. macrocilia [TT84a]. macroglobulin [DWP81]. macromolecules [FCNS84, GSG82, GN84, HO84a, HBO84, JFGC82, MMLT84, MGK83, Rud80, STRP83, TAB80]. macronuclear [WA84]. macronuclear-specific [WA84]. macronuclei [WA84]. macronucleus [Wol80a]. macrophage [ASC84a, GL82, HPG83, KK82, MKD^{+}84, MA83a, MPS^{+}83, MP84b, MPU84, MKB81, MMVO83, NRMB84, PHM81, VKE^{+}81, vdWSCH81]. macrophage-injured [GL82]. macrophage-like [MKD^{+}84]. macrophage-specific [HPG83]. macrophagelike [HCH^{+}81, TEL84, WYF81]. macrophages [AW82, ATW83, AUT83, FMS^{+}84, GL80a, GHY81, GML^{+}80, HPM82, IYVV83, KC80a, LC81, MSC83, OP81, PWM81, PWWS80, PWSW80, PS81, PO81, SPEL84, SG82b, SHBS80, SNS83, TSS82, VC83, WBPO80, WMP^{+}84a, WYF81]. made [NAH83, SMGB84]. Madin [MBP^{+}83, MSB^{+}83, PS83b, PAS84, PBS84, RIP^{+}84, RBPS83, RBPSB84]. magnesium [MW80, YS84, Mar80]. magnetic [GOSW82]. main [BDH82, DRF84, FBH83]. maintained [FSVW80]. maintaining [TQS84]. maintenance [HO84b, MMS84b, KMP^{+}82]. maize [BB81, HSRL81, NHMT84]. major [AFW84b, BSV84, CBS83, CBJ^{+}82, ED84, FHG^{+}81, GOKK82, HO84a, HAFG82, HMKM82, HMG80, HWM83, JOH^{+}84, JFGC82, JHCC84, KSF84, KLD^{+}82, KSK^{+}82, NS83b, OWPC82, PST^{+}83, PF83b, RN81, SSE84, SF84b, VSBL83, VRR80]. make [KWSB84]. malaria [LBA^{+}84]. malarial [AMRE81, FBST84, OK82, Per81a]. males [CL82]. malignant [HLM^{+}84b, VEL80, YKJF80]. Malorientation [JL84]. Mammalian [PPR^{+}82a, ASWR83, Bal82, BZ83, BPB^{+}81, BCP^{+}81, BS81c, Can80, CHRB84, DPK84, DJ81, DMB84, DS80, EKSZ80, FKB83, GLBH82, GDM^{+}80, HGW84, HCK80, HS80b, HK80b, KFV83, KPG^{+}82, KMP^{+}82, MHNT81, MGG83, MM82, PKB84, PM83, RV80, SKN83, SSL^{+}84b, SG80, SSSJ83, TD82, WH81, WWKW83, WHSK80, ZHM80, vMS83]. mammals [WNB82]. mammary [DKDB81, DK84b, DB81a, EPM84, HSB^{+}83, HS80a, JHCC84, LPB84, PTH83, PT83, SSG^{+}83a, SHFK83, WOM^{+}81, YKJF80]. mammotrophs [SF81]. Manipulation [DKDB81]. Mannans [SNS83]. manner [BYS82, HSC84, WS84]. Mannose [TSS82, AL84, BCF84, GGK82, GK84, GSS^{+}84, GGK83, NHF^{+}83]. mannose- [BCF84]. Mannose-specific [TSS82]. mannose-type [NHF^{+}83]. mannosidase [HTGB84, RH81]. mannosyl [SG82b]. mannosyl-fucosyl [SG82b]. mansoni [CKS82, SSZ83, SC82a, SCD82]. many [BKM82, LW81]. MAP [BV83, BLV84, BSV84, PWO84, PAO84, VDT81, WSM^{+}82]. MAP-1A [BLV84]. MAP-2 [WSM^{+}82]. Mapping [CEC84, KDL84, MLSW82, MWI^{+}84, MS80b, Ste83a]. MAPs [SDL81, Val82, WSM^{+}82]. marginal [CN80, DBK80, MS84d, NJSC83, SD80, SS84c, JSC84]. marinus [SLIF80]. marker [BPM^{+}80a, CF84a, FYC82, KCM83, RH83, RH84a, RH84b]. markers [Lan82, NS83a, SY83, TV83, THT^{+}84, WHSK80, WES84]. marmorata [CSR^{+}81]. marrow [IRMM84, RNH^{+}83, SB83a, SGKD83, WR81, WC83]. marrow-derived [WC83]. masking [FSW^{+}83]. Mass [SHWS83, Sta81, SHT^{+}83]. masses [SHT^{+}83]. masses-per-unit-length [SHT^{+}83]. Mast [GDM^{+}82, CLHA80, CH80c, DHS^{+}84, HBR^{+}84, TTGA82, ZSM^{+}83]. mastocytoma [ZSM^{+}83]. material [Blo84, Bur84]. mating [Har80a, MJM83]. mating-type [MJM83]. matrices [DGD^{+}84, HSW82, HSC80, HC83, KKM81, LSC83, MOV83, NLD^{+}82, SAW83, SD84]. matrix [AKM^{+}82, AHMS82, BAF84, Ber80a, BT84, Car84, CBS^{+}84, CVC^{+}84, CHRT83, DPK84, DVvdMvdM81, ECGvdRP83, EP80, FWP84, FBB82, GZR^{+}80, GSG82, GAS^{+}83a, HJ81, Hay81, HOMR82, HTKC82, HS84, IMB83, KPG^{+}82, KMP^{+}82, LWW^{+}83, LGB84, LKJ84, MW83a, MTG^{+}82, MKB82, MLMS81, MMS84b, MMAO83, MMR^{+}84, MRB84, PC80, PHST84, PPRR82, SR84a, Sat84, SR82a, SBCS^{+}82, SDG^{+}84, SPBB84, Sto84, SH81b, TSGS83, WHD^{+}84, WHK^{+}84, ZGN^{+}80, vEvV81]. matrix-intermediate [FWP84]. matrix-specific [MMS84b]. Maturation [MLNK83, BS82b, BvGJ83, CBL84, FS84, GDM^{+}82, GBE84, LAAS84, LK84, MKD^{+}84, RCJS82, Sch81b, TV83]. maturation-inducing [Sch81b]. Maturation-promoting [MLNK83]. mature [CF83, JSC84, LWW^{+}83, LL82a]. maturing [WV84]. may [STR84, YBG81]. mays [BB82c]. MDCK [HEO82]. means [FHFL82]. measured [JE81, KKM84, MPC84, SMH84]. Measurement [OCT81, ML84d, PM81, TMRT84, VKHH82, WN81]. Measurements [Nic83, SLS^{+}84, Wal82b]. measuring [Fis82b]. Mechanical [PT83, MGG83, NKH82, SSA80]. Mechanics [AB81]. Mechanism [BS81a, Che81, FO80, GVK81, HKCM83, JPS84, JHCC84, MOS84, OH84, PS82, SJC80, ST83, SN80, TP82, VTSS80, YPR^{+}82, CBP84, EHS80, FBS^{+}81, JGSH82, JH84, KCGH84, LLS82, MLZD83, PWM81, SM82a, TI82b, WSLSJ83]. Mechanisms [SKMA82, FAP81, GGL80, LP83b, LA83, SB83b, SMTT83]. Mechanochemical [PCOK82, KP84]. medaka [Gil83]. media [ASC84a]. medial [GQW83]. mediate [LGHI84, ZK83]. mediated [AR81, AT83, Asc84b, BHSA82, BBLY83, BZN82, CG84, DWP81, DH84, FCO83, GE84, GML^{+}80, HHS83, HBS^{+}81, MKD^{+}84, MCN^{+}84, MG83a, MS80a, MFA83, MPS^{+}83, ML84c, NHF^{+}83, PSBF83, QCB84, RPM83, RK84, SCS80, SFU84, Sch81b, Sea82, SLRP82, TGPC82, UE83, VTSS80, WHSK80, WSLSJ83, YSS82, BPM81, BGE84, VV83, WG82a]. mediates [FNSB80, LSC83, RGE83, WB81a]. mediating [GGVR82, PJE80, SL82a]. mediation [HJ81]. mediator [FD82]. medium [CFCS82, CAL^{+}82, EHESF82, FHDL82, KS82, Rud80, WSM80]. medulla [ES81a]. medullary [EHK80]. megakaryocyte [LMN83, RNH^{+}83]. megakaryocyte-enriched [RNH^{+}83]. megakaryocytes [LN82, ZFP84]. megaplasmid [WPKB83]. Meiosis [CL82, Kub82, FS84, WS81a]. meiotic [Kub82, Mur80, RN83, SR84b]. Melanesian [SLSK84]. melanocytes [HPM^{+}83]. melanogaster [BWM80, CPLC80, CL82, DK81b, FSB81, FBB82, HMB^{+}82, KSI83, KI83, SEE84]. melanoma [FBP82, HWK^{+}84, MF84b, SSF81]. melanophores [SWP81]. meliloti [WPKB83]. Membrane [BP80, CCJC82, GG83, GN84, JD82, KKKH84, MR81b, Mec81a, Mec81b, MTW82, MSC83, OKA84, PRK^{+}82, PJ83, RCM^{+}83, RW82c, Rob81, TDM80a, WWES83, dCSB84, vDB80, ARL^{+}80, AW82, ATW83, AdS83, AMRE81, ACF84, AS81a, AF84, ACC83, AC81a, ACMH83, AWB82, AC82, ASG83, Bar82b, BHJ83, BF82a, BCLT^{+}80, BC84a, BS83, BTJ82, BM80, BPM80b, BNB^{+}82, BOH81, BvGJ83, BSS83a, BSS83b, BKM82, BF84b, CKB^{+}80a, CKB^{+}80b, CKS82, CH80c, CFW84, CF80, CHH^{+}83, CH80d, DH83, DB80, Den80, DPS80, ERN83, EL84, Fis82b, FGS^{+}82, FML83, FLH^{+}81, FHG^{+}81, Fri82a, FFGC84, FFS^{+}82, FBS^{+}81, GR80, GKKPB83, GDTS81, GAS82, GBS80, GBW82, GBWR83, GB81, GHL84, GMRB81, GS84d, GDRS83, GBB^{+}82, GQW83, GWQ^{+}84, GE84, GRZ83, HSTK83, HJ81, HBO84, HHT80, HRBJG80]. membrane [HFW82, HB84, HCMK81, HM83a, HP82a, HM83b, HSCS80, IYVV83, ILL83, JOA84, JB81d, JB82, JR80, JWBC81, KHS^{+}80a, KLF80, KHF81, KR82a, KK82, KC80a, KC82, KDL80, KMCAS80, KRN83, KI83, KW81b, KTW83, LSM83, LR81, LHS82, LNA81, LN81, LHHH82, LFS^{+}81, LGHI84, MB84b, Man82, MTKH84, MB80a, MBP^{+}83, MS84b, MTR81, MPF83a, MFA83, MSRB84, MCPB80, MKG81, MHCG80, MD80a, MD80b, MBK82, MH84, MBLvdM84, MMVO83, ML84c, MPW82, MMIGK82, MR82, MP81, MP82, MSC80a, MSC80b, MHW83, Mur84a, Mur84b, NKH81, NCH83, NCP84, OWMB82, OHA^{+}84, OO82, PBFC^{+}84, PG83b, PG83a, Per81a, PL84, PS83b, PAS84, PBS84, PWWS80, PJE80, POYF81, PTH83, PWT82, PV84, QGW84, RSP^{+}83a, RKM^{+}81, RA81, RIP^{+}84, RB80, RBS84, RH83, RH84a, RH84b, RH82, RBPF80, RMU84]. membrane [RC83, RLP84, RLT81, SMAS81, SWGB80, San82, SLSK84, SJC80, SDT81, SLLP82, SHH82, SSG^{+}83b, SJS84, SJM81, SSP82b, SLS83, SP84b, SP84c, SBGP84, SH82c, SBSL84, SA83a, SBBS84, SMMC83, SSP84b, SLIF80, TT83, TWW82, TGPC82, TFGC84, TBD81, TdS84, TJB83, TTI80, TTI^{+}81, TTI84, VB81, VF82, VMMC84, VN82, VSLS80, WKH80, WG82a, WSLSJ83, WZ83, WD84, WHJ80, ZP80, ZSE80, ZFP84, vMS83]. Membrane-associated [CCJC82, WWES83, dCSB84, ATW83, DPS80, FHG^{+}81, HM83b, LNA81, MPW82, POYF81, TBD81]. membrane-binding [HM83a]. Membrane-bound [GN84, MR81b, Mec81a, Mec81b, Bar82b, MP81, MP82]. membrane-calcium [WHJ80]. membrane-cytoskeleton [BNB^{+}82]. membrane-enveloped [CFW84]. membrane-inexcitable [HSTK83]. membrane-microfilament [GDTS81]. membrane-microtubule [Mur84a, Mur84b]. membrane-particle [AW82]. membrane-producing [OWMB82]. membrane-spanning [GKKPB83]. Membranes [Hep80, ASC84a, ARL^{+}80, AS81b, AC81b, AC83, BHJ83, BVAR83, BPM^{+}80a, Bow80, BVGS82, BB81, BC83, CCJC82, CSR^{+}81, CAZF84, DRF84, Den80, Ekb81, EK84, FR81, Fro84, FHFL82, FFS^{+}82, GS80a, GCKAA83, GMO84, GLMW84, GWQ^{+}84, HBS^{+}82, HCJV83, HT82b, HTKC82, HWM83, HM83b, IBB^{+}81, JFGC82, KKHK82, KMCAS80, KHS80b, LLM82, LN81, LGF^{+}84, LGHI84, MRFF80, MP83, MS80a, MBR84, MTM83, MKD80, NB81, NM84, OKA84, PL84, PL83b, PS81, QGW84, RS83, RBH^{+}84, RN83, RK80, RMAF80, RH84a, RH82, RKH82, RMJ83, RF81, RUB84, SKMA82, SMAS81, Sea82, SWF84, Sin80, Sow84, SDM84, SBM84, SEM80, Ste83b, SMG81, TIS^{+}81, VPMF84, WB81a, WB81b, WWCD84, WTSAC^{+}81, WV84, WOBR80, ZKF84, dST82]. membranous [Hir82, TI80]. meningiomal [KSMF84]. meristematic [WSO^{+}81]. meromyosin [MBC80a, PCB83]. meromyosins [WP80]. mesangium [RMAF80, SC82b]. mesenchymal [CFN^{+}83, GH82b, NCM^{+}83, ZS84]. mesenchyme [CFN^{+}83, ELST81, NCM^{+}83, SHPH84]. meshwork [PP84]. meshworks [BR84b, Sma81]. Messenger [Jef82, PD80, CSK^{+}80, HHPB83, HSRL81, LDF^{+}84, Ped83, RC83, SKS82]. metabolic [GDL80, JB81a, SCN^{+}81, ZF81]. metabolically [DM83]. Metabolism [PLMS81, BHSA82, FH84, HMR82, Sch83a]. metabolites [SGW^{+}82]. metal [SMG81]. metal-decorated [SMG81]. metanephric [ELST81]. metaphase [EL83, EHCR84, HWS82, KNHK84, KNK84, LP83a, PL83a, WHJ83]. metaphase/anaphase [WHJ83]. metastatic [DCS84]. methionyl [NFS81, DYP83]. methionyl-tRNA [DYP83]. method [CKP84, Hir82, HKCM83, JE81, LdMM^{+}84, ME84, MMLT84, MSC80a, OCT81, SF84b, Tok83, dSK84, vMS83]. methods [Cle84, SFU84]. methyl [DM83, HA80, RRRS82]. methyladenine [Sch81b]. methylase [BGPB80]. methylated [PS81]. methylation [DK84b]. mevalonic [SGW^{+}82]. Mg [POS80]. Mg2 [JSDM82, JSMT82, NJR80, RTP82]. MgCl2 [TP82]. mice [BO81b, SSZ83, TQS84]. Micrasterias [GBS80]. microanalysis [NLN^{+}81]. microbeam [AB81, LPH83]. microbicidal [BHSC83]. Microbodies [TE81, OBC^{+}84]. Microdomains [KNBSF84, SSP81, SSSP81, SSP82a]. microelectrode [IYVV83]. microfibril [Her80, MB80b]. microfibril-synthesizing [MB80b]. Microfilament [GAR^{+}84, Sch81b, BC82, BW80, GL81, GDTS81, GN81, KOS83, OM82, PWT80]. microfilament-associated [BW80]. Microfilament-mediated [Sch81b]. Microfilament-organizing [GAR^{+}84]. Microfilaments [SG80, BC84a, DGC83, HF80, LMYM84, MG83b, OT82, POYF81, SCS80, TH84]. microglial [HPG83]. microglobulin [FFEL81, SP84a]. Microinjected [Bec84, BMF84, Gla83b, KFS81, MBC80a, NBMD81, WRK81]. Microinjection [HFSK82, KDMS83, WS83, BBLY83, HHT80, MB84c]. microirradiation [MB81]. micromanipulation [NKH82]. micromolar [FT80, SM83c, TP82]. micronuclei [AAW^{+}84]. microphotometric [Wal82b, HHSS81]. microprobe [MSB80]. microscope [BB80a, BKL^{+}82, DK81a, DGS80, GDTS81, GAS83b, GC80, HLSWH82, Moo81, PPRR82, SKKC84, TTI^{+}81, WSB^{+}80, ZRW80]. microscope-analysis [GAS83b]. microscopic [AO84, BS83, BM83, CGA^{+}82, CS82, DBG^{+}81, HAF82, HP81, KL82, LCA84, MEKT84, MWI^{+}84, SSJS84, TDS83c, TDS83b, TI80]. microscopy [BCHG83, BM82a, BBS81, CKP84, ECM81, FCKP84, GSvdLS81, GDL80, Ino81b, IT82, IDF83, JWBC81, KB81a, LDM80, MF84a, MS84e, MDA83, OFH^{+}82, OOZF80, PP81, PPSP83, PM81, SEPB81, SBP81, SG81, Ste82a, SHWS83, TS81, TTI80, WSO^{+}81, WD83, Wol80b, dHLC84]. microsomal [ACTK84, GBW82, MACK84, PBG80, RVB80, SJC80, WB81a, WB81b]. microsomes [AC81b, BM80, GBWR83, KORBS82, KFV83, MA83b, NDS84, PBG80, PB83a, WTSAC^{+}81]. Microspectrofluorometry [TMRT84, KKM84]. microsporidian [WB82, Wei82]. Microtrabecular [EP80]. microtubular [AHPL81, RFUC84, TKA83, WL80]. Microtubule [BK83b, Den80, DPS80, GSK83, KB81b, RPB81, SKN83, SDL81, VBT84, WSM^{+}82, AFSF^{+}81, BM82b, BV83, BLV84, BSV84, BB80c, BB80d, CL82, DBG^{+}81, DB83, DMB84, DSKS84, DCK84, GBM84, GRB82, GMSK81, GLS83, HA84a, HZM80, HM84e, IWM83, Jen82, JPS84, KNHK84, KKMK84, KNK84, KHL82, KOS83, LLS81, LLS82, LOL81, LL81, ME84, MKB81, MS84d, MBB83, MF84c, Mur80, Mur84a, Mur84b, NJSC83, OCA^{+}84, OLH80, PWO84, PKB84, PWSW80, Raf84, RA81, RBS84, SLS^{+}84, SKS^{+}84a, Sch84, SD80, SM81, SD82, SS84c, TT84a, TAC84, Tuc84, VDT81, Val82, WZ80, WM81, WWS83, WSO^{+}81, WD83, WL80]. Microtubule-associated [BK83b, SDL81, VBT84, WSM^{+}82, BM82b, BV83, BLV84, BSV84, BB80c, BB80d, DBG^{+}81, DCK84, GLS83, HM84e, IWM83, LLS81, LLS82, LOL81, Mur80, PWO84, PKB84, SKS^{+}84a, Sch84, SD80, SD82, VDT81, Val82]. microtubule-dependent [JPS84]. Microtubule-membrane [Den80, DPS80]. microtubule-myosin [AFSF^{+}81]. Microtubule-nucleating [KB81b]. Microtubule-organizing [GSK83, SKN83, GRB82, NJSC83, Tuc84]. Microtubules [BBB82, BMCGO^{+}82, BP83, BB80a, BKB80, BG82, BV83, BLV84, BTL84, BCP^{+}81, BB80c, Bur84, CT82, DHL84, EM80, EM81, EJM82, ERB83, Fri83, HR81a, HA84a, HLH81, HHT^{+}84, HHS^{+}81, Hir82, HSC84, HA80, IWM83, Jen82, KFS81, KDMS83, Kie81, Lan80, LL81, MPH82, MR81a, MHW83, MWH83, MG83b, NKH82, OK80, PWO84, PSM84, RR80a, RS84, SB80a, SMH84, SBK^{+}84, SSJS84, SBGD84, Ste82b, SD82, TH81, TAC84, TPPH80, TH84, Tuc84, TM84, Val82, WR83, WM81, WSM^{+}82, WW83, WC84]. microvascular [FDKF82, HSS82, KBDK84]. microvillar [BRH82, BM83, BP82, CCJC82, MPW82]. microvilli [BW80, HM83a, MB82a, MGW^{+}80, TJ80, dCSB84]. microvillus [Bre83, GGW83, HMG80, MB82b, MS80b, MGW^{+}80, PM81]. midbody [EM80, MM82, NDC80]. migrate [MP84c]. migrating [GMSK81, GH82b, UAL84]. Migration [MPF83a, SLR^{+}83, BDP^{+}84, ELS^{+}83, GSK83, MF84b, MBC83, NNA^{+}83, RDY^{+}83, RLT81, SAW83, SR83b, Wat84]. migratory [Fri83]. milk [FHG^{+}81, SK81, SHFK83]. mimics [SSHP82]. mineralization [ECGvdRP83, HHS^{+}84]. mineralizing [LSBPG84]. minor [MNP^{+}84, PK81]. miotic [Hep80]. MIP [FBH83]. MIP26 [BDH82]. missorting [SFU84]. mistake [Ste84b]. Mitochondria [ES81b, BGM81, BM80, BS81c, GK81b, JB81c, Man82, MWH83]. mitochondria-specific [BGM81]. Mitochondrial [LL82a, BVAR83, BS81c, CPS84, CF83, FFS^{+}82, GL82, JWBC81, Man82, SHZN83, SHH82, Ste84b, WTSAC^{+}81]. mitochondrial-specific [CF83]. mitogen [DTP84, RGB84, SBG80, WCZG84]. mitogen-stimulated [WCZG84]. mitogenesis [SM82a]. Mitogenic [FKB83, SCP^{+}83]. mitogens [CCCR81, DK84a, EB83, LKJ84]. Mitosis [RHK82, AB81, Baj82, BO80, CLFC83, Hea80, HBR^{+}84, KRB84a, KOB82, MBB83, RS81, WFGB83, WH80, WHJ80]. Mitotic [CP83, ASWR83, BO80, BLV84, BLS^{+}80, CWBB83, CLFC83, CBS^{+}84, GBM84, HZM80, HH83b, Ino81a, IWM83, KHL82, KB81b, LPH83, ME84, MB81, MBB83, Nic83, PKB84, POS80, SS80a, SMH84, SB83b, WHJ83, ZHM80]. mixtures [MHC80]. MM [WDE83]. MM-creatine [WDE83]. Mobility [SA83b, WYF81, GAS82, GBL81, KSD^{+}83, KST84, SSF84, WRK81]. mobilization [GHB82, SW84]. mode [EBS84]. modeccin [SSO84]. model [ACF84, Bal82, Can80, CMM81, DB83, DHS^{+}80, Fri82b, GDM^{+}82, GAS^{+}83a, Pfe84, SHW^{+}82b, SO82, WE80]. models [IMS82, KW84, MP84a, NW82, PCB83, PB84b, SMS84b]. modes [NBSW81]. Modification [ZB82, BM82c, CH80b, CBL84, MBB83, NBM82, SR82a]. Modifications [BF82a, LLB84]. modified [LR83b, OHA^{+}84, PCB83, MBC80a]. modifiers [Gol82]. modifying [DM83]. modular [HY82]. modulate [AW81, MB84a]. modulated [SBSM84, SY83]. modulates [CNH80]. modulating [WR83]. Modulation [ACC83, LPB84, LWWP84, TSGS83, AA83, BALR83, DTP84, GPR83, HG80a, HM84d, JNS^{+}84, MW83a, MPMBC80, SZ80, Zig81, ZSL82, SMT80]. modulations [HS80a, dLTvdSB83]. moieties [VEL80]. moiety [BHJ83]. mol [BB80d, CG84, DBG^{+}81, HM84c, MBLvdM84, PV84, RQHR81, RBH^{+}84, TFGC84, Vah83, WS84, WSM^{+}82]. mold [KCGH84]. Molecular [EMA83, ES82, EPE^{+}80, GZR^{+}80, ZGN^{+}80, BCSC82, BM82b, BL82a, GSG82, GBF84, GLS83, GE80, HJ81, IY84a, JB81a, JB81b, JW83, KZS81, KPW^{+}83, LLS81, MS84a, MFVPK82, NS83a, Pol84b, SMJB83, SCM83, SIM^{+}84, TWW82, Wei83, WES84, SBVW82]. molecule [CEC84, ECM81, FA83, GGW83, GE84, LNB84, LLB84, RGE83]. molecules [BO81a, BKO81, KOST80, LS82, Mec81a, MLD82, OYNT83, OGS82, PPR^{+}82a, SM80, SB80c, SR83a, SMG81, SH81b, TGPC82]. molluscan [CN80, NJSC83]. Monensin [HWAK83, KHS^{+}80a, MSD82a, AC81a, LUT80, QGW83, WKW82, GQW83]. monitored [TPR82]. Monitoring [JWBC81]. Monoclonal [BBWS83, CD84, ED84, GV82, HAS^{+}81, KKP84b, Lan82, LFS^{+}83, MWI^{+}84, Pip84, SSP84a, WWCD84, WES84, ASM^{+}84, BMF84, CVT84, CF84a, DF84, DSRF84, FGS^{+}82, Fro84, GKL^{+}84, HO84b, HM84e, KKP84a, KP84, KWM82, KDL84, LQ82, LMYM84, LGF^{+}84, MDA83, NLD^{+}82, NKH81, PBPD83, PKB84, RMS^{+}82, SHT^{+}84, SHPH84, SWF84, SGPB84, SDGYJ84, SLS81, TAC84, WCY^{+}83, WWS83, WW83, WMENS82]. monocytes [BSKFW83, HM84b, LH83, SGM^{+}84, WLCS84]. monoether [ZRW80]. Monolayer [Fis82a, ALE84, ASA83, DVvdMvdM81, HSFM82, SY80, ZH82]. monolayer-coated [HSFM82]. monolayers [CGM81, FCNS84, MMAO83]. Monomer [ML84d, CT84, PPRR82, PPR82b]. Monomer-polymer [ML84d]. mononuclear [BvdMN84, GS80b, IRMM84, WC83]. monophosphate [MSP84, PB84b, VSD84]. monopolar [Baj82]. monovalent [LUT80, MPU84]. MOPC [GK84, HTGB84]. mordanting [BBCL82]. morphogenesis [BT84, CF84a, GVK81, HA80, SAP81, SA82, TS83a, WAN81]. morphogenetic [AP84, OKE81]. morphogenetic-mutant [AP84]. Morphological [JRM^{+}80, JRB80, MA83a, BLG^{+}80, Bow80, Bro81, BS82d, CFN^{+}83, FCL^{+}82, LP83b, MAG82, SDC80, SG84b, ZH82, MBP^{+}83]. morphologically [MPF83b]. Morphology [KML83, MSPF83, WPKB83, APZL80, CGG81, CGG84, DBK80, EMA83, FH84, KC83, KPG^{+}82, LK83, LA80b, SSCF84, SGW^{+}82, SM81, Sol81, ZGJ^{+}84]. Morphometric [BOH81, MC80, FGHC80, OCT81, OBC^{+}84]. Most [WG82a]. Motile [Goo83, SSA80, HT80, MD82, NW82, PCB83]. Motility [All81, KRC82, BP83, BP82, Bur82b, EN84, EKS^{+}82, HG80c, KCB84, PWT80, SO82, TA81, TKA83, VS84, WL80]. Motion [DH81, Bec84, HSFM82, PHST84]. motoneurons [SM83a]. motor [SGKC82]. mount [KB81a]. mounts [LCA84]. Mouse [KWSB84, BHJ83, BJ83, BJG83, BO81a, BKO81, BNH81, BBB82, CB82, DKDB81, DK84b, DCK84, ECGvdRP83, EWG82, EW82, EPM84, EK84, FG81, FBP82, FSBR83, GGK82, GM84, GAS^{+}83a, GDL80, GLMW84, GRDR82, Hat81, HDB82, HPG83, IMB83, JZ81, KPPC83, KG80, KOLV81, KC80a, KLD^{+}82, LSN^{+}84, LPB84, LCK84, LH84, LA83, ME81, Max82, MF84b, MMC83, MA83a, MNW81, MGK83, OP81, PWM81, PAO84, PWWS80, PWSW80, PT83, PO81, RWD84, SFS81, SSF81, SH82a, Sob84, SLS81, Ste83a, SG84b, SKA^{+}83, SNS83, Sur82, TS83a, TIK81, VN82, VC83, WWCD84, WMP^{+}84a, WHSK80, WC83, Wie81, WZ83, WWWC80, ZSE80, ZFP84]. mouse-derived [GLMW84]. mouse-human [EWG82]. move [SHRL84]. Movement [FKS84, AB80, Baj82, BLH82, BG83, CMM81, Che81, HF80, JOH^{+}84, KMI82, KOB82, MR82, NRW82, PMP81b, PT83, RW82b, SCS84, TWH80, TBR80, TI80, WWKW83, WW83, YMF84]. movements [HA84b, JPS84, SSBB82]. MP26 [PG83b, PG83a]. Mr [EPRS81, FYC82, LV83, SWF84, SPHE80]. mRNA [BB82c, MR81b, Mec81a, Mec81b, MDSH80, MP82, OLHH81, SSZ83, SMTP80]. mRNAs [HBB82]. mRNP [Liz80]. mucosa [RBH^{+}84]. mucus [SN80]. multifunctional [HY82]. multilayered [ILL83]. multinuclear [Fuk80]. Multiple [MR81a, MLSR83, MCG82, SMTT83, Ber80b, RHK82]. multiplicity [DGD^{+}84]. multivalent [CKS82, FNSB80]. murine [BLHC81, BGS81, DCS84, ED84, FS84, GS80b, HDRK83, JOA84, MRFF80, MW83b, MGSC84, NCL84, Reb82, SP82a, SP82b, TEL84, VSLS80, VH82, WWES83, vdWSCH81]. muscarinic [Har80b, HE81]. muscle [AF83b, ABSC81, Bad81, BSP83, BHR81, BAF84, BSS81, BS82c, BSVS80, BL82a, BN83, BNGN84, BBCL82, CC82, CTKJ83, CSS80, CARK81, CB84a, CCCR81, CF84a, CW80, Con84, CS83, COS83, DSRF84, EM84, EPRS81, FN80, FFRF83, FSFA84, FSVW80, FAP81, FM80, FPFP81, GLBH82, GBLH83, GDTS81, GNW^{+}84, GL83a, GSS82b, GHGT81, GKL^{+}84, GRZ83, GAAG82, HF83, IDF83, JSDM82, JSMT82, JSCM83, KL82, KS83, KLC83, LFAG84, LBJ84, LKJ84, LS80, MP80, MB84a, MSK^{+}84, MHLS81, MS84c, MIRS84, MPF83b, Moo81, MGG83, MCG82, NLD^{+}82, OMM82, OLH^{+}81, PSC83, PFP^{+}82, PNYO83, RMS^{+}82, RMF83, RRS82, RK81, RGE83, SSCF84, San82, SHT^{+}84, SGB^{+}81, SBCS^{+}82, SGSS^{+}81, SPHE80, TI82a, TWW82, TDS83c, TDS83b, Tok83, TG83, WP80, WD80, WDE83, WRM83, WRH81]. muscle [WH83, Wri84b, YS84]. muscle-contractile [FAP81]. muscle-specific [MCG82]. muscles [TS81]. muscular [GDMG81]. Mutant [RPM83, SHRL84, AP84, AMJ^{+}81, BCLT^{+}80, BLH82, BS81c, CWBB83, CGA^{+}82, DCS84, EZ84, GFLG^{+}84, HHT80, HDB82, HWJW84, IY84b, KvRK^{+}84, KSI83, KI83, KJ82, MMS^{+}84a, MKG81, SDL82, SS84a, Sin80, TOK^{+}84, WCJ83, ZHTH80]. Mutants [Har80a, AHPL81, Cab83, DK81b, FNNFS84, GBG82, GDH82, HSTK83, JR80, KRC82, KKKH84, LTAP^{+}81, MJM83, NRMB84, NS83b, VTSS80, WAN81, WGG80, WOBR80, ZSM^{+}83]. mutation [OM83, ROB^{+}84, Ste83a]. Mutations [AMCG83, WSH84, BB82c, GBM84, SLLP82, WC80]. Mycoplasma [GSB81]. Myelin [MWA^{+}80, TQG84, CKFS82, DM83, HCMK81, MS84a, OWPC82, PBW80, SS84a, TIS^{+}81, TQ82, TQS84, YDL^{+}84]. Myelin-associated [TQG84, MS84a, TQ82, TQS84]. myelin-enriched [YDL^{+}84]. Myelin-specific [MWA^{+}80]. myelinating [TQS84, TQG84]. myeloid [NSI80, NSI82, NPG^{+}80]. myeloma [KWSB84, MR81b, Mec81a, Mec81b]. myoblast [CNH80, EHESF82, KLC83, SL82a, Wri84a]. myoblasts [AFSF^{+}81, DGD^{+}84, KSSP83, LH84, Wri81, WA83]. myocardial [JC84, KG80]. myocardium [GSS82b, GS84c, SGB^{+}81]. myocytes [CHI^{+}81, DANH84, Gla83a, Wri81, WA83]. myoepithelial [FSF^{+}80, WOM^{+}81]. myofibril [PSC83, TI82a]. myofibril-sarcolemma [PSC83]. myofibrillar [CS83, SMS84a]. myofibrils [DANH84, Moo81, Nac81]. myofibroblast [SKKC84]. myogenesis [Atk81, BMF82, EM84, GF83, GL83b]. myogenic [MDSH80, OLHH81, SHPH84]. myolemma [CHI^{+}81]. myomesin [EPRS81]. Myosin [CSS80, COS83, GBLH83, GSS82b, LFAG84, SGB^{+}81, Sob84, AFSF^{+}81, BMF82, BS82a, BC84a, BNH81, BNB^{+}82, DGS80, FC82, GLBH82, HP81, HCP81, HTFH82, Kan83, KM82, KMI82, KKP84b, Nac81, NW82, NDC80, OT82, PWM81, PG84, RFUC84, SCS84, SDG^{+}84, SHBS80, TI82a, Wri81, WA83, YMF84, KKP84a, PKS81, Pol82]. myosin-extracted [TI82a]. myosins [BSVS80, FN80, WP80]. myotendinous [CF84a, CF84b]. myotube [MP80]. myotubes [BS82a, Blo84, BF84b, BMA84, PF83b, SA83b, TMS84]. myristate [JB81b, MR82, PWWS80, PWSW80, WCZG84]. myxamoeba [VSD84].

N [PCB83, RMJ84, ZRW80, BO81a, BKO81, BGE84, BG83, CK82, JNS^{+}84, LK84, MBC80a, RBO83, Rot84, VC84]. N-acetylglucosamine [VC84]. N-acetylglucosaminyltransferase [CK82]. N-acetylmannosamine [BO81a, BKO81]. N-CAM [BGE84]. n-dodecyl [ZRW80]. N-Ethylmaleimide-modified [PCB83, MBC80a]. N-formyl [JNS^{+}84]. N-formylmethionylleucylphenylalanine-induced [BG83]. N-linked [LK84]. NADH [BM80, BPM80b, MCPB80]. NADH-cytochrome [MCPB80, BM80]. Naegleria [FW80, LD81, Wal84]. Narcine [HCK80]. Nascent [BTF^{+}83, AL84, BR83, BR84a, DANH84, MD80a, MD80b]. native [GB82, HMB^{+}82, HS84, KRB^{+}84c, PG84, SSCP80]. natural [BSG81]. naturally [RSB84]. nature [BR83, SHT^{+}84, YK84]. Nd [TOK^{+}84]. near [CDLJ81, EH81, PR84]. negative [MHC80, Sea82]. nematode [RW82c]. Neocurtilla [KW81a]. neonatal [RWD84, WOM^{+}81]. neoplasms [HHPB83]. neoplastic [BBR83, GL82, SCSC81]. neostigmine [SB82b]. Nerve [HEG81, AMM84a, BKT84, BG82, BMH^{+}83, BWW^{+}80, CE80, CW80, CGG81, CGG84, DCG83, DHHG83, GLS83, GL80b, GPDM81, GB80, GLBS81, Har80b, HH81b, HCK80, HSGD83, MSB80, MTM83, PF83a, PBW80, Pen83, SRCG84, TIK81, VS82, VMMC84]. nerve-muscle [CW80]. nerves [DJ81, PB84a]. nervous [AMM84a, AMM84b, BSV84, CBWW80, CG84, DCG83, MS84a, SFS81, SW81b, SLS81, YF81]. network [PSC83, TA81, TKA83, TTI80, WRM83]. networks [BL80b, CP83, CKP84, JTM80, MOV83, Sch82, SL82b]. Neural [RDY^{+}83, RGE83, ALE84, BDP^{+}84, BF84a, CG84, CEC84, Law83, Pie82, SLKB83, SR83b, Wri84b]. neural-specific [CG84]. Neurite [Rau84, BG82, CMC83, LFGR82, LR84, PJ83, RGB84, SWGB80, SBG80]. neurite-associated [CMC83]. neurite-derived [RGB84]. neurites [GL80b, GB80, Let83, ML84b, RE80, SSJS84]. neuritic [MER^{+}83]. neuroblastoma [CMC83, GSOG80, Olm81, Rau84, dLTvdSB83]. neurofilament [CSW80, HGW84, HGP84, MWM^{+}81, NBM82, YF81]. neurofilaments [Hir82, LLS81, LLS82, ME82, WS81b]. neurohypophysial [BO81b]. neuromuscular [AKT84, BH83, Bur82a, CH80a, FGHC80, HLS81, HR81b, HH81b, Ko84, MH84, NB81, SSHP82, SH83, SB82b]. neuron [PJ83]. neuronal [CT82, FD82, GE84, HGW84, LNB84, MTR81, MPAM80, MHW83, MWH83, RMT84]. neurons [ABSC81, BK83b, BO81b, CGG81, GL80b, HL81, HLM84a, HPG83, JRM^{+}80, JRB80, JVMRK82, KS84, Let83, MER^{+}83, PJM^{+}81, PMP81a, PMP81b, RGE83, ST83, SS84a, SFB^{+}84, Ste82a, YSJ81]. neuropathy [TQG84]. neurosecretory [MTR81]. neurospora [Man82, LL82a, SHZN83]. neurotoxin [MTM83]. neurotransmitter [FGHC80]. neurotransmitters [PGD^{+}84]. neutral [PG84]. Neutrophil [KMCAS80, BHSC83, MR82, SH81a, WWPL84]. neutrophils [BHSA82, CKB^{+}80a, CKB^{+}80b, CKS82, CVAK80, GDH^{+}83, HSFM82, HFW82, HM84d, LWWP84, MBC83, WLFD84]. newborn [AR81, RKF80, SKA^{+}83]. Newly [WFGB83, BNGN84, DCK84, GWQ^{+}84, HSB^{+}82, KHF81, NM84, PJ83, QGW84, RF81, SC82a, SC81]. newt [GK81c, KSG81, KCB84]. nexin [EB83]. NH2 [SSG^{+}83b]. NH2-terminus [SSG^{+}83b]. Ni2 [LFR80]. Nicotinic [JFGC82]. nidulans [GBM84, OM83]. niger [Rag81]. Nitella [HF80]. nitrobenzoxadiazole [BYW81]. nitrobenzoxadiazole-phallacidin [BYW81]. nitrocellulose [SF84b]. nitrogen [KMCAS80, RF80a]. nitrogen-limited [RF80a]. Nitzschia [TPHL80]. nm [BMF84, ME82, PWSW80, TIK81]. no [NRW82]. node [WE80]. nodule [WPKB83]. nodule-like [WPKB83]. nodules [WPKB83]. non [Law83, SSJS84]. non-neural [Law83]. non-uniform [SSJS84]. nonactivated [vDB80]. nonclustered [ABSC81]. noncoated [CB82]. noncollagen [LBJ84]. Nonconverted [OHA^{+}84]. noncycling [RS81]. nondiffusive [Pai84]. nondirected [BG83]. Nonequilibrium [VVP82]. Nonerythrocyte [BKM82, MB84c]. nonerythroid [SBSL84]. nonfilamentous [WS84]. nonfusing [GDH82]. nonhelical [SKCJW84]. Nonhistone [DTT81, EPO^{+}84, HAS^{+}81]. nonimmunological [LS81b]. nonionic [GSB81]. nonkeratinized [PR83]. Nonlinear [GL80a]. Nonlysosomal [AF83a, BBG^{+}84]. nonmetastatic [DCS84]. nonmitogenic [YSS82]. nonmuscle [Bre83, IDF83, MA82]. nonpreferred [CT83]. Nonrandom [KW81a]. nonreceptor [Bar82b]. nonsecreting [KWM82]. nontoxic [BYW81]. nontranscriptionally [CH83b]. noradrenaline [ST83]. Normal [PB84a, CGA^{+}82, CKB^{+}80b, CLFC83, GKKPB83, GBLH83, HS80a, HOMR82, HWJW84, IJ82a, MYR81, MC80, NFG82, Rag81, RSP83b, RGM^{+}80, SHW^{+}82a, SS84a, SK81, TLPTV83, VEL80, WYF81, YKJF80]. normally [HH81b]. Notophthalmus [KSG81]. Novel [WDE83, CF84b, EF84, Fuk80, GDH^{+}83, LV83, Liz80, MSC80a, OT82, Sch81a, dHLC84]. nuchae [MLMS81]. Nuclear [MLZD83, NLPH82, Ped83, Ber80a, CBS^{+}84, CD84, DTT81, EB84, FLM84, FO80, FKS84, FWP84, FBB82, FKK83, FSKJ81, FK80, Fuk80, GOKK82, GOSW82, GKZ81, IMB83, JOM80, LDF^{+}84, LV81, LOC81, LM84, MWP84, MWKP84, MMO^{+}81, MSGP81, MLNK83, PADH83, PD80, PC80, PM83, SGHS^{+}80, SSBB82, SDG^{+}84, SF84b, SPBB84, SS84b, SD84, UM82, WD83, WMP83, ZCH80, vEvV81]. nuclease [LV81]. nuclease-resistant [LV81]. Nucleated [MPW82, BCSC82, BKB80, CBJ^{+}82, JSC84, PM81, SD80]. nucleating [KB81b, SKN83]. nuclei [FCO83, GBL81, KNHK84, KSSR84, LWW^{+}83, LP83a, LOC81, LTGL80, MHNT81, SLR^{+}83, SD84, TD82]. nuclei-induced [LTGL80]. nucleolar [BBLY83, FKS^{+}81, KDL84, SHHR84]. nucleoli [FKS^{+}81, KSK^{+}82, SSSJ83]. nucleolus [BBLY83, Mil81a]. nucleoplasm [GK81c]. nucleoprotein [PM83]. nucleoside [OOP^{+}81, PSR82]. Nucleosome [HMB^{+}82]. nucleosomes [ERH82, OOZF80, ZF84]. nucleotidase [MEKT84]. nucleotide [BLS^{+}80, GBWR83, GCS81a]. nucleotide-dependent [BLS^{+}80]. nucleotides [OT80a, SPEL84]. Nucleus [Woo80, AW81, BBWS83, CP83, GK81a, KST84, LCK84, Yao82]. Nucleus-associated [Woo80]. number [FGHC80, KJ82, OCT81, SKN83, TS83b, WCJ83]. numbers [ZF84]. numerical [dLTvdSB83]. nutrient [RJVS81]. nutrient-dependent [RJVS81].

O [EK84, LSBPG84, Rot84]. O-glycosylation [Rot84]. O-linked [EK84]. O-phosphoserine- [LSBPG84]. O-phosphothreonine-containing [LSBPG84]. O2 [LWWP84]. O2-production [LWWP84]. observation [BW81]. Observations [LS81a, TT84b, AC83, BBCL82, CFN^{+}83, DGC83, NCM^{+}83, BO81a, BKO81]. occidentalis [BBF84]. occludentes [SG84b]. Occluding [MIS^{+}80, EDK80, MPMBC80, PTH83]. occur [HA84a, RKK82]. Occurrence [WKA^{+}83, Fuk80, NDC80, PAO84, RH81]. occurring [BKM82, PBG80]. occurs [GW80, MS84b, Mec81b, SGHS^{+}80]. Ochromonas [TPPH80]. octaethyleneglycol [ZRW80]. octamer [ERH82]. old [Sat84, WRH81]. olfactory [Bur84, CE80]. oligodendrocytes [MWA^{+}80]. oligodendroglial [MdV80]. Oligomeric [Bar82b]. oligomers [CHS82, KRB84b, MHB84, MM81b]. oligomycin [BS80]. oligosaccharide [DR83, GGK82]. oligosaccharides [DCS84, DWS84, EK84, HTGB84, LK84, NHF^{+}83, TV83]. Oncogenes [Wei83]. one [GWS84, TWG83, TBCM83]. one-to-one [GWS84]. only [GCH84, SHRL84, TBCM83]. Onset [WC83, MGMS84]. onto [EDP82]. oocyte [FCO83, FS84, GK81c, SR84b]. oocytes [CML^{+}81, ER84, FO80, FKS84, GWK84, HSRL81, KSD^{+}83, KSK^{+}82, Mab83, MLZD83, Mur80, RN83, RB83, Sch81b, SP84a, WS81a]. openings [SLS83]. opercular [EDK80]. opsin [GB81, NCP84]. opsonized [MMVO83, NPG^{+}80]. optic [MMKL81, TIK81]. optical [HHSS81, HHS^{+}81, KL82]. order [ACH^{+}81, ARHG84, ABD84, PBW80, WFR84]. ordered [Man82]. ordering [SEE84]. organ [BKT84, CK80, CSR^{+}81, Fro84, GNW^{+}84, HCK80]. Organelle [NNRR80, HA84b, SBSM84]. organelles [CQB84, Hir82, PAGG81, QCB84, SKMA82, SSJS84, TI80]. Organization [AW81, BMA84, FG80a, HTFH82, HGW84, MB82a, SR81a, Sma81, TPPH80, WOBR80, BAF84, BRH82, BSH83b, BCP^{+}81, BB82c, FWP84, FBRP84, FSW^{+}83, GZR^{+}80, GR81, HHT^{+}84, HK80a, HSCS80, Jen82, KML83, KW81b, LR83a, Lan82, Let83, MKB82, MS84d, MOV83, MG83b, OCA^{+}84, Pen83, PWT80, PPRR82, RSDH82, SBK^{+}84, SDL81, SWP81, SA83a, TDM80b, TEDS83, TDS83c, TDS83b, Tuc84, WMP^{+}84a, WWS83, WW83, ZSR^{+}82, ZKF84, ZGN^{+}80, ZF84]. Organizational [MKW^{+}82]. organized [MMR^{+}84, TT84b, WSO^{+}81]. organizing [EDP82, GRB82, GAR^{+}84, GMSK81, GSK83, KKMK84, NJSC83, SKN83, Tuc84]. organoids [MMAO83]. organs [LFS^{+}83, SGKC82]. Orientation [Fle81, CGM81, CL82, EN84, FA83, HLH81, HWJW84, SSG^{+}83b, WCJ83, Zig81]. oriented [CVT84, CHH^{+}83, Tok83]. origin [ASM^{+}84, MWH83, RRQ^{+}84, RF80b, SKE84, MHW83]. orotic [UAL84]. Oryzias [Gil83]. oscillations [BWM84]. oscillatory [Baj82]. OsFeCN [Wal82a]. osmium [Sma81]. osmolality [HH83a]. osmotic [HH83a, KMF81, Mad83]. osmotically [ACF84]. ossification [AO84, PPR82b]. osteoblastlike [NvIvGdH^{+}82]. Osteoblasts [ECGvdRP83]. Osteoclast [BvdMN84, IRMM84, MTG^{+}82]. osteoclast-like [IRMM84]. osteoclasts [MCN^{+}84]. osteogenic [SKA^{+}83]. other [BKO81, BW80, BB81, CKP84, HWBC82, IY84a, LFS^{+}83, MP83, MFVPK82, RKK82]. ouabain [AFK^{+}84, MMJ^{+}81, SL80]. Outer [HW83, BS82d, CSHP84, GH82a, Man82, ONO82, TTUI83, WTSAC^{+}81, WMP^{+}84b]. outer-membrane [Man82]. outgrowth [BG82, LFGR82, Rau84, RE80]. outgrowths [GAS^{+}83a]. ovalbumin [SMTP80]. ovalocytes [SLSK84]. ovarian [JDTW80]. ovaries [PS83a]. ovary [AMCG83, AMS82, BHA82, BKB80, BS83, BS80, Cab83, CWBB83, CPS84, CK82, ERB83, HJ80, KRB84a, KvRK^{+}84, KB81a, KB81b, MMS^{+}84a, RPM83, ROB^{+}84, SSC80]. overlay [Ott83, SBB81]. overproduction [SS84b]. Oversized [JR80]. oviduct [GRR^{+}84, SMTP80]. ovomucoid [CSK^{+}80]. oxidase [BHSC83, RFML84, RCJS82]. oxidation [RFML84]. oxygen [SCN^{+}81].

P [MMO^{+}81]. P-450 [MMO^{+}81]. P0 [TIS^{+}81]. p230 [LV83]. p50 [WCY^{+}83]. P700 [ZP80]. P700-chlorophyll [ZP80]. Packing [ABD84, CP81, SHT^{+}83, TEDS83]. pad [ML84a]. pair [AHPL81, DHL84, LOL81, OK80]. Palade [WOM82]. palindrome [CDLJ81]. pallidin [DR82]. pallidum [DR82]. pancreas [FJ82a, FJ82b, GHWR80, GCF^{+}83, IJ82a, RRRS82, RRQ^{+}84, SKS84b]. pancreatic [BBR83, DBL82, IJ82a, IJ82b, KCM83, MPJ80, MMAO83, ORA^{+}84, PE83, RMJ83, SJC80, SDGYJ84, SD82, VMRR80, dSTK81]. papillary [BBS81, JSDM82]. Papua [SLSK84]. paracrystals [FA82]. paradox [Gal81]. parafollicular [BGNT81]. parallel [GCKAA83, PBG80, WM81]. parallels [AKT84, MGMS84]. paralyzed [JR80]. Paramecium [GS84a, ERN83, ARL^{+}80, AS81b, AF83a, AF84, BCLT^{+}80, BPM81, FG80a, GGS83, GS83, HFSK82, HSTK83, HHT80, LTAP^{+}81, LN81, MDM^{+}81, MKG81, OK80, PWT82, RSP^{+}83a, RN81]. Parameters [CAZF84]. paramyosin [LKR^{+}83]. paranemin [PL83c]. paranodal [WE80]. parasite [AMRE81, CKB^{+}80b, OK82, Per81a]. parasites [CKB^{+}80a, FBST84]. Parathyroid [SHJA82]. parental [SSC80]. parfollicular [BGNT81]. parietal [ACMH83, GGV81, ILL83, LNA81]. parotid [AC82, FJ82a, FJ82b, LP83b, MTH84, SHO^{+}84]. part [EHCR84, GL81]. Partial [CKB^{+}80b, KHF81, MB82b, MBK82, MWM^{+}81, AC82, ABSC81, CKE81, FKB83, KDL80, LFS^{+}81, SBSM84, SP82b, ZHM80, Ber80a, SSSP81, SBM84]. partially [BE82, LL82a]. participate [PSBF83, STR84]. Participation [AHRG82, GDRS83, NAK82, Pen84, SB84]. particle [AW82, GBS80, GBW82, GWB82, HA84a, HBS^{+}82, HCMK81, MLZD83, MKD80, NPG^{+}80, PWM81, RRD83, SLLP82, UM82, VKE^{+}81, WB83a, vDB80]. particles [BO83, DH81, FGHC80, GSvdLS81, GML^{+}80, HP82b, KC83, KW81b, Liz80, LM84, MP82, PD80, PF83b, RV80, SHH82, SL82a, SP84c, WE80, dLTvdSB83, vdWSCH81, WGG80]. particular [BWW^{+}80]. partition [dST82]. pass [HSB^{+}82]. passage [FCNS84]. passaged [WBG83]. patch [Wei80]. pathology [BBPR81]. pathway [BF84a, BBG^{+}84, GBWR83, ML84a, MRHS81, SF81, SHZN83, Sch81b, WH82]. Pathways [GSOG80, GS84d, Oli82, SBG^{+}81, SPS^{+}84]. patients [BPB^{+}81, NFG82]. Pattern [Gla83a, LZR82, PBW80, FS80, UAL84, UE83, WDE83]. patterning [SA82]. patterns [GK81d, Hat81, HAF82, TT84a]. pawn [MKG81]. PC [GSH80]. PC12 [BMH^{+}83, BG82, CGG84, GLS83, GLBS81, HEG81, PGD^{+}84, RGB84, SRCG84, VS82]. pea [CHS82, CHCS82, RZSW83, RBMM82]. pealei [GS84b, LOL81]. peckii [RN83]. PEG [Wol80b]. pellucida [ED84, SH82a]. pelvic [BL82b]. pennate [TPHL80]. peptidase [BWM80]. peptide [Bar82b, BDP^{+}84, HM84d, JNS^{+}84, ME81, MS80b, SGM^{+}84, SZ80, SDZ84, WWPL84, YK84, Zig81, ZSL82, ORA^{+}84]. peptide-induced [WWPL84]. peptides [BDP^{+}84, BWM80, GB82, MGMS84, SKCJW84]. perfused [DH84, LSJT83, NSW82, PS83a]. periaxonal [TQS84]. Pericellular [GT84, MMR^{+}84]. periciliary [PPSP83]. perikarya [Can82, Can83, SM83a]. perimeter [UM82]. period [ASWR83]. periodic [SA82]. periodicity [ME82, TQ82]. Peripheral [Fro84, KS84, AMM84a, DJ81, DCG83, FBB82, MER^{+}83, MS84d, OWPC82, PF83a, SW81b, SCN^{+}81, SS84c, TQ82]. periphery [FSKJ81, LW81]. peritoneal [GML^{+}80, OP81, PWWS80, PO81, SNS83, VC83, WMP^{+}84a]. Permanently [FSVW80]. permeability [AMJ^{+}81, GDL80, JFGC82, KLF80, MTH84, MBC83]. permeabilized [Can80, CMM81, SO82]. permeation [PK81]. permissive [RR80a]. peroxidase [CQB84]. peroxidase-containing [CQB84]. peroxisomal [RFML84, RQHR81]. peroxisome [RRRS82, RQHR81]. peroxisome-proliferation-associated [RQHR81]. peroxisomes [FFS^{+}82, TE81]. perpendicular [HTS80]. persist [GS84e]. Persistence [SCP^{+}83]. perspectives [ZS82]. perturb [HSFM82]. perturbation [BK84]. PF [GLMW84]. pH [BR84c, CV82, DTP84, GS84a, GHY81, Gil83, GOSW82, HT80, HT82a, JFS84, JE81, KC82, LS81a, LS81b, MOS84, Max82, MPC84, PO81, RPB81, RTP82, SO80, TMRT84, WN81, WKH80]. pH-regulated [CV82]. phagocytes [BvdMN84, WC83]. phagocytic [GS80b, KSC82, SDC80]. Phagocytosis [vdWSCH81, AW82, HCH^{+}81, HPM82, HPG83, MG83a, MPS^{+}83, NPG^{+}80, PHM81, SCN^{+}81, SHBS80, SNS83, TEL84, VTSS80, WMP^{+}84a]. phagocytosis-associated [HPM82]. phagocytotic [VTSS80]. phagolysosome [GHY81]. phagolysosomes [AF84, OT80a]. Phagosome [HMM82, KC80a, AF83a, Bow80, GHY81, HM84b]. Phagosome-lysosome [HMM82, KC80a]. phagosomes [HT82a, MTMT83]. phallacidin [BYW81]. Phalloidin [CT84, PNYO83]. phallotoxin [NBSW81]. phase [BJG83, BM82b, GWK84, GSOG80, GSH^{+}84, GN84, HM84a, MR82, O’B80, Ric80, RF80a, RF80b, KHS82]. phasic [CHI^{+}81]. phenomenon [PBG80]. Phenotype [CCCR81, FBP82, KMP^{+}82, MB84a]. Phenotype-dependent [CCCR81]. Phenotypic [BGE84, MW83a, MMS84b]. phenylalanine [BJS81, NFS81]. pheochromocytoma [BG82, BMH^{+}83, GSH80, HEG81, SRCG84]. Phorbol [BC84b, EB83, PWWS80, PWSW80, JB81b, MR82, WCZG84]. phosphatase [AF84, BL82b, FS80]. Phosphate [DM83, BCF84, GGK82, GK84, GSS^{+}84, GGK83, RACGB81, WB83b]. phosphatidylcholine [HH83b, KRB^{+}84c, SP84b]. phosphatidylinositol [Sch83b]. phosphodiesterase [DK81b, GCS81a, VV84]. Phosphoinositide [FH84]. phosphoinositides [AH84]. phosphokinase [EKSZ80]. Phospholipid [BHSA82, CAZF84, CH80d, CL81, FFS^{+}82, MS83b, PJ83, PS81, Sin80]. phospholipids [RA81]. phosphomannan [FNSB80]. Phosphomannosyl [STR84]. phosphoprotein [BS81b, DCG83, DHHG83, HSGD83, SF84a]. phosphoproteins [LSBPG84, SHO^{+}84, SS81, WWES83]. phosphorylase [OOP^{+}81]. phosphorylated [FJ82b, NHF^{+}83, PDK83]. Phosphorylation [BNH81, CLFC83, EPO^{+}84, FC82, OO82, Sch84, BZ83, BL80a, CCWR84, DBBN81, FJ82a, FJ82b, FJ82c, GGK82, GMVF84, HWK^{+}84, KM82, LLS81, LN81, PF81, PL83a, PKS81, PHRL81, RCM^{+}83, SA83a, Ste83a, Ste84b, WD84]. phosphoserine [LSBPG84]. phosphothreonine [LSBPG84]. phosphotyrosine [CH82]. photobleaching [SSL^{+}84a, SLS^{+}84]. Photochemical [DPS80]. photoinduced [NJR80]. photoreceptor [AC83, BS82d, CSHP84, RH82, RKH82, SLLP82, Sch83a, Sch83b, UPLH84, Wal82a, Wal82b]. photoreceptors [BM82a, BP80, LHHH82, NCP84, dCSB84]. Photosensitive [SS81]. photosynthetic [MJ83, SJM81, VF82]. photosystem [VA83, WOBR80]. phragmoplast [EM80]. phycobiliprotein [OGS82]. phycobilisome [LYG81]. phycobilisomes [GBG82, GWG82, LYG81, WGG80]. phycoerythrin [GBG82, GWG82]. Phylogenetic [SCS82]. Physarum [BGS^{+}83, CDLJ81, HAF82, Nac81, OT82, OH84, SWA83, SBGD84, SIM^{+}84]. Physical [TRC^{+}84, VH82]. physiologic [BTHL81]. physiological [CT83, EN84, LTAP^{+}81, PSSS82, SHW^{+}82b, Ste84b, BCLT^{+}80]. physiologically [AAW^{+}84]. phytohemagglutinin [VC84]. phytohormone [ZB83]. picture [FA83]. pig [BF82a, Mad83, MHLS81, MB82c, MP84c, MPK84, SSG80]. pigment [Fri83, JB82, WGG80]. pigmented [OKE81, OM82]. pigmentosum [GS84f]. pilosebaceous [MFVPK82]. pinocytic [AMS82, KSC82, MSUC80]. pinocytosed [BAWL81]. Pinocytosis [BALR83, VN82, BO80, DLZ83, FNSB80, Gol82, GNGTS80, PvD83, PWWS80, RJVS81, WMP^{+}84a, WKW82, WWDYW83]. pinosomes [HT82a]. Pit [CGG81, CGG84]. pitfall [MK84b]. pits [CGA^{+}82, MTKH84, MDA83, PvD83, WP82b]. pituitary [APZL80, BB80b, BK84, GCL83, GPR83, GLP80, KBD82, LRT^{+}82, ME81, MGK83, SF81, SW84, SMT80, Sur82]. PK1 [DVB^{+}81]. placenta [HBB82, HHPB83]. placental [HBB82, HHPB83]. planar [ACF84, CAZF84]. plant [Jen82, WSO^{+}81, WD83]. plants [DRF84, MB80b, TWWG84]. plaques [AF83b, BC83, KSMF84, SR81a]. Plasma [AC82, Fri82a, PL84, AdS83, AC81a, ASG83, BJG83, BVAR83, Bow80, BvGJ83, CHH^{+}83, DH83, GBS80, GMRB81, GDRS83, GWQ^{+}84, Gri80, HEA^{+}82, HCJV83, HHD83, HWM83, HM83b, JOA84, KK82, KKG84, KDL80, KMCAS80, KS82, LSM83, LGHI84, MBP^{+}83, MS84b, MLM83, MSRB84, MBK82, MMVO83, MMIGK82, MTM83, MSC80b, NCH83, PG83b, PG83a, PS83b, PAS84, PL83b, PS81, QGW84, RIP^{+}84, RBS84, RH83, RH84a, RH84b, RMJ83, SDT81, SSG^{+}83b, Sin80, SP84b, SP84c, SBGP84, SDM84, SBM84, SEM80, SMMC83, SSP84b, TT83, TMA^{+}83, VPMF84, VMMC84, VN82, WKH80, WTSAC^{+}81, vMS83, GHL84]. plasma-binding [HHD83]. Plasma-membrane [Fri82a]. plasmalemma [RKMF84, TDM80a]. plasmalemmal [PMP81b]. plasmin [EB83]. plasmin-activatable [EB83]. Plasminogen [GPR83, GMJR82, DVB^{+}81, EB83, JB81a, JB81b, KS84, LSN^{+}84, LTBS80, LL82b, SLK^{+}84, YKJF80]. plasmodia [OT82, SIM^{+}84]. Plasmodium [LBA^{+}84, SWA83, GAS83b, MTW82, OK82, Per81a, TMW82, WWES83]. Plasmodium-infected [MTW82]. plastic [DCG83]. plasticity [FD82]. plate [PPR82b, TEDS83]. Platelet [GN81, LCA84, SGS^{+}82, SB84, SGPB84, BPDR83, CCWR84, HHD83, MLHM84, NKH81, PG82, RNH^{+}83, SGH^{+}83, SCS82, SCP^{+}83, SC84b, vDB80, VRR80]. Platelet-collagen [SB84, SGPB84]. Platelet-derived [SGS^{+}82, BPDR83, HHD83, SCS82, SCP^{+}83, SC84b, VRR80]. platelet-specific [RNH^{+}83]. platelets [BL80a, FBRP84, GWWP83, IER84, LA84, Nac80, PJE80, RSL81, SB82a, SBB81, SSLB84, ZF81, vDB80]. plates [TT81]. platinum [HK80a]. play [RR80a]. plectin [WKA^{+}83]. pleiotropic [KWSB84]. pleiotropically [KvRK^{+}84, RPM83]. Pleurodeles [GK81c, Sch81a]. plexiform [HPG83]. plexus [VVM81]. plus [BWW^{+}80, FT80]. pneumoniae [GSB81]. pneumophila [HM84b]. podocalyxin [KSF84]. podocytes [RBO83]. point [ATW83, EBS84]. points [RK80]. polar [JB81d]. Polarity [BKB80, EM80, EJM82, ERB83, HLH81, WM81, BS82c, CGM81, EM81, GH82b, HO84b, HW83, JZ81, Kir80, ME84, MSRB84, NW80, SS80b, TH81, YF83, ZLK81]. Polarization [CVGM83, WBPO80, HHSS81, HHS^{+}81, Ino81b]. polarization-based [Ino81b]. polarized [Fri82a, RBPS83]. poliovirus [MOS84]. pollen [Rag81]. poly [BK83a, MSGP81, Rag81, SGHS^{+}80]. polyacrylamide [HEA^{+}82, SBB81]. polyadenylate [MP81, MP82]. polyadenylate-containing [MP81, MP82]. polyamine [MMC83]. Polyanion [KC82]. polyanions [KSC82, KC82]. polycephalum [HAF82, SIM^{+}84]. polyclonal [BMF84, MNP^{+}84]. polyethylene [KLC83, RRD83, WSLSJ83, Wol80b]. polylysine [FA82, Hat81]. polymer [ML84d]. polymerase [SLR^{+}83]. Polymerization [Pol84a, BB80a, FFEL81, Her80, KR82b, LTGL80, MPW82, MF84c, OH84, PM81, Pol82, SMH84, TTI84, WWS83, NSI80]. polymerized [BHA82, PH81]. polymers [Kir80, ML82]. Polymorphism [FA82, AFI^{+}81]. Polymorphonuclear [RLT81, BSS83a, BSS83b, DLZ83, MD82, NFG82, RKK82, RVM83, SDC80, SZ80, SDZ84, ZLK81]. Polypeptide [FFS^{+}82, AHPL81, ACTK84, BTF^{+}83, BSV84, BDH82, BC80, BZN82, CG84, CL80b, FKB83, FBH83, GOKK82, Har80a, LV83, LYG81, MACK84, MSUC80, MWH83, PDK83, PHRL81, RQHR81, SBG^{+}81, SC84c, VPMF84, WB83a, WSB^{+}80, WD84]. polypeptide-mediated [BZN82]. polypeptide-specific [WB83a]. polypeptides [BB82c, CLFC83, CHS82, CKFS82, FO80, GL83a, GL83b, HAFG82, HGW84, HMKM82, JFGC82, LW81, MM81a, Mil81b, MA83b, MFVPK82, OKA84, SF84b, WMENS82, ZK83]. polysomal [MP81]. polysome [MP81, MP82]. polysomes [Liz80, MR81b, RFML84, RKP^{+}82, WIB81, WB81a]. polyspermy [JGSH82]. Polysphondylium [DR82]. polystyrene [Bec84, CL80a, CFML83, LC81, MG83a]. polytene [CPLC80, CLJ^{+}81, KB81c]. polytomella [MR81a, MLSR83]. polyvalent [MP84b]. pool [HF83, KSD^{+}83, RB82a, SHHP81]. pools [Ber80b, Olm81, TTGA82, TSS82]. population [Rao81, RSB84]. populations [MP81, MP82, RMU84, VKHH82]. porcine [DVB^{+}81, FHDL82]. pore [FBB82, GOKK82, SR82a, SS84b, UM82]. porelike [Man82]. pores [FKS84]. porin [SR82a]. portion [VDT81]. portions [San82]. position [BTF^{+}83, PA83, TS83b]. positioning [HL81]. positive [GGK83]. Possible [SKS82, SR83a, CVC^{+}84, FD82, HH83a, HS80b, KUM^{+}84, LLS82, MLZD83, MPU84, PL84, RH81, RR80b, SBC82, Ste82a, TI82b, Wol80a]. post [LLB84, RKP^{+}82]. post-translational [LLB84, RKP^{+}82]. posterior [RMF83]. postmitotic [AFSF^{+}81, Ric80]. postnatal [BVGS82, JRB80]. Postreceptor [MSP84]. Postsynaptic [SBC81, Bad81, Bur82a, CGCS80, CGS81, CBS83, FYC82, Fro84, GCS81a, GCS81b, GR81, HLS81, JFGC82, MPAM80, Sea82, SWF84]. Posttranslational [BNM82, NBM82]. posttranslationally [LR83b]. Potassium [MM81c, ASC84a, AFK^{+}84]. potassium-activated [AFK^{+}84]. potential [IYVV83, JWBC81, MTW82, MR82]. Poterioochromonas [Her80]. pp60src [SR81a]. pre [KFV83]. pre- [KFV83]. preadipocytes [NNRR80]. preaggregation [FTM82]. prealbumin [KKG84]. preceding [BF82a]. precipitates [WH80]. precursor [BL83a, GLP80, LDF^{+}84, MLSR83, MA83b, SHPH84, Ste84b]. Precursors [MWKP84, MTG^{+}82, ML84a, PD80, SMGB84]. predominantly [GHL84]. Preface [McI81, Sie81]. Preferential [BHJ83, LLS81, SSP82b, PJM^{+}81, SSP81]. preformed [CGM81]. Prefracture [SEM80]. Preganglionically [DK81a]. pregnancy [GHH82]. prekeratin [FSF^{+}80]. preliminary [SG84b]. prelysosomal [GN84, MPU84, SPS^{+}84, WKAH84]. prematurely [HR80]. Preparation [MdV80, PG83b, SSCF84, SD84, AC82, HSGD83, SBC81, TFAP81]. preparations [CBS83, EMA83, LRT^{+}82, LM84, WTSAC^{+}81]. prepared [SSP84a]. preparing [CKP84, TOG83]. Preprophase [WD83]. preproteins [BB81]. presage [BS82b]. Presence [MDM^{+}81, RKMF84, TQ82, ASWR83, BCSC82, Cab83, GC80, HAF82, JC84, ML84b, SD80, SM83c, WSM^{+}82, YKJF80, GCS81a, GCS81b, HWM83]. present [BW80, GAS83b, HMB^{+}82, ORA^{+}84, VSBL83, YBG81]. preservation [HS84, MP83]. pressure [BSH83b, HHS^{+}84, HS84]. presumptive [GBF84, PP84]. presynaptic [Har80b, MSB80, MBK82, MMIGK82, Pen83]. prevalence [KRB84b]. prevented [OLH^{+}81]. preventing [CT84]. prezein [BB81]. Primary [CSK^{+}80, CHI^{+}81, LFC^{+}80, SY80, BJG83, CAL^{+}82, CH82, CRGS82, GBS80, LS82, LDV83, SG82b]. primate [WH83]. primates [BB80d, DBG^{+}81]. prior [MP84c]. Pro [SGM^{+}84, ME81]. pro-ACTH [ME81]. pro-ACTH/ [ME81]. proadipocyte [KS82, SHW^{+}82b]. probably [EH81]. probe [DWP81, MIRS84, SGSS^{+}81, WB83a]. probes [BDP^{+}84, BF84a, DCK84, ED84, MLSW82, ME84, SG81, WSLSJ83]. probing [SR82a]. procaryotes [BTF^{+}83]. procaryotic [TT83]. procaryotic-eucaryotic [TT83]. procedure [CQB84, PL83a, TOG83, Val82]. process [BCLT^{+}80, BM83, BMCGO^{+}82, DTBF82, GA84, RMU84, TI82b]. processes [FD82, LTAP^{+}81, RUB84]. Processing [HTGB84, AAW^{+}84, BVC83, BNM82, BB81, BB82c, CHS82, DR83, HT83, Ino81b, MWP84, MA83b, MS84e, NDS84, PD80, Ped83, Per81b, RWD84, RBS84, RKP^{+}82, RMH^{+}84, TMRT84, WM84]. procollagen [LUT80, MK84b, PST^{+}83]. produce [CARK81, LL82b, MMR^{+}84, OT80b, SMTP80]. produced [GLMW84, LDV83, Nic83]. producing [OWMB82]. product [AFW84b, SAP81]. Production [PF83a, PBPD83, AHMS82, BCC^{+}83, CHRT83, GMJR82, JWG84, JPS84, KMI82, LTBS80, LFS^{+}83, MLMS81, MGMS84, MNW81, SMT80, LWWP84]. products [ME81, SHHR84, WSH84]. profilaggrin [RWD84]. profilaments [Lan80]. profile [HH83b]. profiles [IJ82b]. profilin [OH84, TBCM83, TP82, TRC^{+}84]. progenitor [NW83]. progenitors [IRMM84]. Progesterone [GRR^{+}84, HS80a]. proglucagon [FQBN81]. proglucagon-converting [FQBN81]. program [BGD80]. Programmed [HC84]. Progression [DK84b, SCSC81]. prohormone [NDS84]. prohormone-processing [NDS84]. prohormones [NM84]. proinsulin [FQBN81, ORA^{+}84, OHA^{+}84]. proinsulin- [FQBN81]. projection [VDT81]. projections [YF83]. Prolactin [KBD82, APZL80, BRR^{+}83, BB80b, GZR^{+}80, PVPB82, TLPTV83, ZGN^{+}80]. proliferating [FSVW80]. Proliferation [JB81c, ASA83, GCL83, GLMW84, Kla80, LTE84, LKJ84, RQHR81, SB80b, SWGB80, SBG80, SMT80, SHW^{+}82b, SHFK83]. proliferative [GHGT81, YDL^{+}84, ZSM^{+}83]. proliferator [RRRS82]. proline [ML84a]. proline-injected [ML84a]. prometaphase [TPHL80, CL82]. promote [CRGS82, MF84b]. promoted [MOV83]. promoter [SNPE84, WCZG84]. promotes [EKS^{+}82, FLH^{+}81, HTS80, LFGR82, MER^{+}83, MMAO83, NNA^{+}83]. promoting [MLNK83]. promyelocytic [CF83, MKD^{+}84, RCJS82]. pronounced [NKT80]. pronuclear [LM84]. propagation [MHSW81]. proparathyroid [HMDP80]. propel [RW82b, Tam82]. propeptides [GB82]. Properties [CVC^{+}84, SL82a, WZ80, ALE84, BHP83, BYS82, CBP84, ERH82, GLBH82, HHSS81, HHS^{+}81, HM84a, HM84c, HF83, IBB^{+}81, JRM^{+}80, JRB80, KKMK84, LSM^{+}84, Mab83, MGG83, RVH84, RS82, SWA83, SBCS^{+}82, SH82c, SBVW82, TRC^{+}84, VS82, VTSS80, Wal82b, WTK83, YSJ81]. prophase [OM83]. proposal [TPHL80]. propulsion [SLIF80]. prosomatostatins [NDS84]. prostatic [CFN^{+}83, NCM^{+}83]. protamine [ZCH80]. Protease [KHS80b, CPLC80, COBC84, EB83, PG84]. protection [Leh81]. protective [LA80b]. Protein [CDLJ81, DHHG83, GBW82, GWB82, PADH83, AC81b, AWB82, AL84, BK83a, BCSC82, BL80a, BBLY83, Ber80a, BGNT81, BALR83, BTJ82, BPM81, BV83, BLV84, BSV84, BVC83, BM83, BM80, BPM80b, BGPB80, BL82a, BW80, Bre83, BFST83, BYS82, BLS^{+}80, BB80d, BS82d, BS81c, BC83, BMCGO^{+}82, BF82c, CBS83, CBWW80, CHS82, CHCS82, CV82, CH82, CBP84, DRF84, DBG^{+}81, DCG83, DCK84, EMA83, EHS80, ES82, EF84, EPRS81, EHESF82, FG81, FCO83, FKS84, FWP84, FYC82, FB82, FHG^{+}81, FKS^{+}81, FJ82a, FJ82b, FJ82c, FM80, GR80, GBWR83, GCF^{+}83, GCH84, GCS81b, GLS83, GKL^{+}84, HTS80, HSME84, HM84c, HAS^{+}81, HM83a, HHD83, HM84e, HSGD83, IMS82, IJ82b, IWM83, JR80, KSD^{+}83, KNH80, KLC83, KLD^{+}82, KSK^{+}82, KTW83, KB81c, LD81, Law83, LBA^{+}84, LBJ84]. protein [LLS81, LN81, LOC81, LQ82, Mab83, MW83b, MB82b, MLZD83, MTR81, MCPB80, MLD82, ML80, Mil80, MA82, MBLvdM84, Moo81, MMIGK82, MGG83, MF84e, MPK84, NAH83, NCH83, NBM82, OJJ^{+}84, ONO82, OT82, OKA84, OWPC82, PF83a, PWO84, PG83b, PG83a, PFM^{+}84, PKB84, PAS84, PBS84, PV84, Pol84b, PPRR82, PPR^{+}82a, PPR82b, PPR^{+}84, PGD^{+}84, RBH^{+}84, RBS84, ROW^{+}83, RSL81, RS82, RMU84, RUB84, RC83, RLP84, RNH^{+}83, SMAS81, San81, SBGD84, SKS84b, SR82a, SKS82, SBG^{+}81, Sch84, SWF84, SG84a, SZ81, SD80, SG81, SBGP84, SF84b, SBSL84, SS81, SA83a, SY83, Ste83a, SHBS80, SHFK83, SSSJ83, SIM^{+}84, TTGA82, TWG83, TRN^{+}82, TIS^{+}81, Vah83, VDT81, VVP82, VSBL83, VC84, WOM82, WM84, WIB81, WB81a, WB81b, WCY^{+}83, WS84, WFGB83, WC80, WG82b, WWWC80, YK84, YAF81]. protein [ZP80, vEvV81, MSD82b, DSRF84, EPRS81, OMM82, Ped83, RMS^{+}82, RMF83, SPHE80]. protein-actin [HM84c]. protein-bound [AC81b]. protein-lipid [Mil80]. protein-requiring [ML80]. proteinaceous [FBB82]. proteinase [ZCH80]. Proteins [BFL^{+}83, VMRR80, AKM^{+}82, Atk81, BJG83, BBR83, BM82b, Ber80b, BK83b, BTHL81, Bro81, BSS83b, BB80c, BMS83, BWV84, BKM82, CB81, CGS81, CRH81, CB84a, CPS84, CVT84, CL80a, CD84, CHCS82, CML^{+}81, CVC^{+}84, CH82, CS83, CSW80, DH83, DJ81, DM83, EPO^{+}84, ERN83, FSB81, FNHSS84, FP81b, FG80b, Fro84, GKKPB83, GK81a, GBL81, GK81c, GV82, GMRB81, GS84d, GDRS83, GBB^{+}82, GQW83, GWQ^{+}84, GS84e, HOR84, HSB^{+}82, HEA^{+}82, HPM82, HM83b, HSRL81, IY84a, IY84b, IJ82a, JKHT81, JOM80, KGAB81, KKG84, KSSP83, KDL84, KSD^{+}83, KKBG83, KW81b, KST84, LPB84, LLS82, LV81, LOL81, LGHI84, MB84b, MTKH84, MS84a, MNP^{+}84, MSGP81, MKG81, MD80a, MD80b, MWA^{+}80, MGK83, MS80b, MF84d, MSC80a, MSC80b, MSC83, MRB84, MCG82, Mur80, NSI80, NSI82, NBMD81]. proteins [NKH81, NSW82, NS83b, Ott83, OO82, Pai84, PYC^{+}82, Per81a, PC80, PJE80, Pol81, PL83c, PM83, QGW84, QCB84, RCH81, RW82a, RZSW83, RPHA81, RMJ84, RLP84, SKMA82, SMS84a, SMS84b, SDL81, SJC80, SKS^{+}84a, SKS82, SSG^{+}83a, SRGB83, SMGB84, Sch84, SGH^{+}83, SS84a, SR81a, SM83a, SW81b, SW81a, SBB81, SCM83, SFB^{+}84, SWB84, SD82, TGPC82, Val82, VBT84, WIB81, WB81a, WB81b, WSM^{+}82, WHD^{+}84, WG82a, WYF81, WZ83, WRK81, YF81, ZGJ^{+}84, ZP80, ZSE80, ARL^{+}80]. Proteoglycan [TS83a, AF83b, AKT84, AMM84a, AMM84b, BK84, DB81a, HOMR82, Ioz84, KMP^{+}82, LFGR82, LLM82, LH83, PST^{+}83, PPRR82, PPR^{+}82a, PPR82b, RVH84]. Proteoglycans [WH83, HS84, KHF81, KHJG84, PPR82b, WHK^{+}84]. proteoheparan [OWMB82]. proteolysis [HMR82]. Proteolytic [AAW^{+}84, MA83b, SMJB83, TGPC82]. proteus [GK81a]. proton [KCGH84, SS81, YFM83]. protoplasts [GBM84]. protozoan [OBC^{+}84]. protrusions [GAS83b]. provide [HL81, Lan82]. provides [DTBF82]. proximal [CFCS82, CAL^{+}82, KNBSF84, RKMF84, TV83]. proximal-to-distal [TV83]. Proximodistal [Can83]. pseudopodia [OCT81]. pseudopodial [RW82b]. PtK1 [CMM81, EDP82, EM81, IWM83]. PtK2 [FSW^{+}83, MB81]. pulmonary [CLHA80]. pulse [LNB84]. pulse-chase [LNB84]. pulses [FTM82]. pump [CVGM83, EMC83, GCKAA83, MMJ^{+}81, YFM83]. Purification [Bre83, GW80, MPF83b, OBC^{+}84, Pol84b, RFLR81, SR84b, ASG84, BHP83, CQB84, DR82, FKB83, KMCAS80, MFP80, MBK82, MWM^{+}81, QCB84, SSBR84, SF84b, NSI82]. Purified [MP80, VPMF84, HCK80, HSGD83, KRB84b, MD82, MTG^{+}82, OT82, RBH^{+}84, RF81, SMJ80, Sta81, ZKF84].

Purine [OOP^{+}81, ZHTH80]. purine-auxotrophic [ZHTH80]. purple [SMG81]. purpuratus [HOW80, HAP82, KR82b, KJBR82, PV84]. Putative [FYC82, LDF^{+}84]. pyrene [SCSC81]. pyrene-induced [SCSC81]. pyriformis [Den80]. pyrophosphatase [RB82b]. pyrophosphorylase [GKZ81]. PYS [OWMB82]. PYS-2 [OWMB82].

quaking [TQS84]. Qualitative [PG84]. quality [Ino81b]. quantal [MTI80]. Quantification [Pie82, SHHP81]. Quantitation [BB82a, MW80, PS83b, SHB80, BO80]. Quantitative [HHSS81, HHS^{+}81, RACGB81, SR81b, TOT^{+}84, dLTvdSB83, BLG^{+}80, CGA^{+}82, Fis82a, GS84a, MYR81, PYC^{+}82, PBFC^{+}84, SAW83, SCS84, TAGG80, KOB82]. quantities [Fis82b]. questions [Sat84]. Quick [HH81a, CH80c, CH80b, Hir82, HT82b, HKCM83, Law84]. Quick-freeze [HH81a, Hir82, HKCM83]. quick-freezing [CH80c, CH80b]. quick-frozen [HT82b, Law84]. quiescence [GG80, ZSM^{+}83]. quin2 [KKM84]. quin2-loaded [KKM84].

r [Bar82b]. rabbit [AFW84a, AFW84b, BHSA82, BSS83a, BSS83b, CNSW80, CAL^{+}82, DBK80, ES81a, FFGC84, HSB^{+}83, HO84a, HBO84, KNH80, NFS81, NNA^{+}83, RNH^{+}83, SSCF84, SW81b, SB83a, SHFK83, SCSC81, THT^{+}84, Vah83, VKHH82, ZCH80]. Radial [HPRL81, PHRL81, BLH82]. radial-spoke [BLH82]. radio [SCS82]. radio-receptor [SCS82]. Radioautographic [LSBPG84, ML84a, UAL84, BRR^{+}83, TTGA82]. radioautography [BO81a, BKO81, BKL^{+}82, WGRB80]. radioimmunoassay [HHD83]. radioiodinated [Fis82a]. radiolabeling [SFU84]. raise [Max82]. ram [WV84]. Raman [BM82a]. Ranvier [WE80]. Rapid [AMJ^{+}81, BSR84, CMC81, GL80b, JNS^{+}84, SMH84, TKM83, WWDYW83, BAWL81, BHP83, CGG81, CGG84, LH84, Mad83, MPU84, MOV83, MPC84, NCL84, OR81, SNPE84, SR82b, SM82b, SN80, TdS83a, TOG83]. rapid-freezing [OR81]. rapid-frozen [SR82b]. rapidly [BC84b, GS84c, KOLV81, Max82, RGR80]. Rat [KWM82, Wei80, AR81, AC81b, ACTK84, ACMH83, APZL80, AM81, ASA83, AA83, AC82, Bad81, BJS81, BKT84, BE82, BHJ83, BO81a, BKO81, BRS^{+}82, BRR^{+}83, BGM81, Blo84, BMH^{+}83, BM80, BPM80b, BCC^{+}83, BB80b, BvGJ83, BBS81, BMA84, BBG^{+}84, BMCGO^{+}82, CEC^{+}83, CCKMH82, CKHF80, DWM83, DGS80, DH84, ES81a, EKS^{+}82, FBH83, Fle81, FDSWW84, FSVW80, FWG^{+}81, FJ82b, FFS^{+}82, GJRO^{+}83, GSS^{+}82a, GSS^{+}84, GHWR80, GBWR83, GN84, GCF^{+}83, GLP80, HHS83, HER81, HOMR82, HBR^{+}84, HWM83, HM83b, HM84e, ILL83, ITT80, IJ82a, IJ82b, IHJ82, JKHT81, JH84, JDTW80, JSDM82, JSMT82, JSCM83, KCM83, KKG84, KBD82, KSSP83, KLC83, KORBS82, KST84, LSJT83, LP83b, LDV83, LDM80, LS81b, LNB84, MHGM82, MAG82, MMO^{+}81, MNST82, MEKT84, MTH84, MdV80, MLP84]. rat [MG81, MBR84, MSRB84, MCPB80, MHB84, MNW81, MYR81, MWA^{+}80, MC80, MP81, MP82, NCP84, NLN^{+}83, PS83a, PBG80, PRK^{+}82, PE83, PC80, PR83, PVPB82, QCB84, RVB80, RFML84, RACGB81, RQHR81, RRQ^{+}84, RKF80, RUM^{+}82, RVH84, Rod80, RKMF84, RGM^{+}80, RH84a, RMJ83, RK81, SF81, SLR^{+}83, SHHR84, Sch83a, Sch83b, SGA80, SG84a, SR81a, SJ80, SDGYJ84, SHO^{+}84, SBBS84, SA83b, SHP83, TZW83, TG83, TIS^{+}81, UAL84, VB81, VVM81, VMRR80, VSLS80, WYK^{+}84, WOM^{+}81, WGRB80, WWS83, WW83, WRH81, WTSAC^{+}81, WKW82, WYF81, WSM80, Wri81, WA83, YSJ81, ZH82, MACK84]. rate [GHGT81, HOR84, HF83, MR80a, OT80a, PM81, RF80b, SMH84, SH83]. rates [EWG82]. rather [SSA80]. ratio [CH80d]. ratios [NLN^{+}81]. rats [GFLG^{+}84, HAEC81, ML84a]. ray [BBS81, GR80, LP83a, NLN^{+}81, PL83a, SEPB81, SBP81, UE83]. rDNA [CDLJ81, YBG81]. reaches [MS84b]. react [MS84a]. reacting [GV82, SHT^{+}84, WWS83, WW83]. reaction [DTBF82, IT82, LS81b, MJ83, MM81c, PV84, SSG80, TI82b, VF82, VA83, WOBR80]. reactivated [AB80, GG80, PCB83, PB84b]. Reactivation [Bur82b, BSNP82, IMS82, LWW^{+}83, SLR^{+}83, YT82]. reactive [CRH81, DTT81, EW83, RCH81]. reactivity [CBP84]. reacts [RBH^{+}84]. read [VSD84]. reagents [OFH^{+}82]. realized [PB84a]. Reappearance [HS83]. Rearrangement [KOS83, Fri82b]. rearrangements [BBWS83, KA84]. Reassembly [LL81, ZK83]. reassociation [MD80b, TTI^{+}81]. recently [GK82]. Receptor [DH84, HHS83, QCB84, RK84, SFU84, AR81, ABG81, ABBG82, AWB82, AKT84, Asc84b, BW81, BW82, BHSA82, Bar82b, BAF84, Blo84, BN83, BU84, BMA84, CSR^{+}81, CVGM83, Con84, DWP81, DR82, ELS^{+}83, FCL^{+}82, FNSB80, GGK82, GK84, GSMHF84, GRR^{+}84, GSS^{+}82a, GSS^{+}84, GBW82, GWB82, GGK83, GNGTS80, Gri80, HG80a, HWAK83, HWK^{+}84, HS80a, HMR82, HT83, HF83, JNS^{+}84, KHS80b, KOB82, LAAS84, LA84, MW83b, MKD^{+}84, MNST82, MFA83, MSB^{+}83, MPS^{+}83, ML84c, NB81, NHF^{+}83, OMG82, Pen83, PP84, Pen84, PSBF83, PFP^{+}82, PB83b, RPM83, SCS80, SSHP82, SHT^{+}84, SSG^{+}83b, Sch81b, SBF84, SWF84, SHJA82, SCS82, SPEL84, SG82b, SC84a, SZ80, SDZ84, TSS82, VMMC84, WH81, WBPO80, WKAH84, Zig81, ZSL82]. receptor-adenyl [CVGM83]. receptor-bound [ABG81]. receptor-deficient [GK84]. Receptor-mediated [DH84, HHS83, QCB84, RK84, SFU84, AR81, Asc84b, BHSA82, DWP81, MKD^{+}84, MFA83, MPS^{+}83, ML84c, NHF^{+}83, PSBF83, RPM83, SCS80]. receptor-positive [GGK83]. receptors [AF83b, ABSC81, BPDR83, BNGN84, BCF84, BF84b, BGS81, CB82, CP82a, CW80, GNW^{+}84, GAAG82, HSFM82, Har80b, HS80a, HBR81, HSS82, HS83, HRBJG80, HE81, HDRK83, HF83, JNS^{+}84, LA83, MLP84, MS84c, MPS^{+}83, MP84b, MPU84, MHB84, OMG82, Pfe84, Rod80, SH83, SSE84, SSP82a, SB82b, STR84, SC84a, SA83b, VS82, VC83, Wei80, WRH81, WLCS84]. recirculation [SR83a]. Recognition [FBST84, NHF^{+}83, SLS81, AWB82, CRH81, GBW82, GWB82, RCH81, RH81, VTSS80, WH81, WIB81, WB81a, WB81b, WB83a]. recognized [DBL82, WCY^{+}83]. recognizes [ASM^{+}84, LQ82]. recollections [PB81]. Reconstituted [Mil80, EMA83, RUB84, SMJ80]. Reconstitution [MS83b, PM83, SKCJW84, Ste83b, DB80, Mur84a, TI82a, DRF84]. reconstruction [MB82b]. recovering [JL84]. Recovery [KORBS82, AS81a, ERB83, DSD80b]. Recruitment [MTG^{+}82]. Recycling [CCWR84, SBF84, ABBG82, CH80a, GNGTS80, GA84, HHS83, HB84, MPU84, MSC83, SMMC83, TSS82]. red [BBLY83, Her80]. Redistribution [BCF84, CP82a, JOH^{+}84, LA84, PYC^{+}82, SSLB84, ZSE80, CVGO81, Fri82b, PWSW80, SSL^{+}84a, SLS^{+}84, SSHP82, SKWF84, vDB80]. reduce [SNdC^{+}84]. Reduced [TOK^{+}84]. reduces [BS82a, DB81a]. reducing [OFH^{+}82]. reductase [BM80, BPM80b, FDSWW84, KHS82, LBM82, LCK84, MCPB80]. reduction [AMJ^{+}81, FS84, NCBP84, SM83c]. Reevaluation [BP82, KRB84b]. reexamination [FM80]. Reexpression [Kho82, MNW81]. reference [HM84a]. reference-phase [HM84a]. reflect [SCP^{+}83]. reflection [Axe81]. reflects [FS80]. Reformation [SS84c, MS84d]. regard [PWM81]. regenerated [Bad81]. regenerating [MS84c, MYR81, RB82a, SM83a]. Regeneration [Ko84, CE80, GHB82, GPDM81, GRZ83, MER^{+}83, RW82a, RR80b, SW81a, TAGG80]. region [GBF84, NDC80, TT83]. Regional [RBO83, NBM82, SC84c, PMP81a]. Regionalization [AdS83, WZ83, WV84]. regions [ACTK84, CGCS80, CB82, FHG^{+}81, JC84, KHS82, MACK84, dPMMF^{+}81, MP84c, OWPC82, PSC83, SL82b, WYK^{+}84]. regular [HPG83]. Regulated [BB80b, FP81b, AAW^{+}84, CH83b, LCK84, NHMT84, SF84a, SPEL84, TBCM83, CV82]. regulates [BGD80, BLHC81]. Regulation [ACH^{+}81, BKT84, BJ83, CTKJ83, FDKF82, FS84, GJRO^{+}83, GLS83, GS84f, HPM^{+}83, HMKM82, JB81b, JWG84, JDTW80, MR80a, MGW^{+}80, OLH80, PB84b, SGKD83, SS80c, SA83a, VA83, Wie81, WA83, AT83, BJG83, BCC^{+}83, BS81b, CH80d, DH83, FNS81, FJ82c, GG83, GS84a, GOSW82, GMVF84, HOT81, HAP82, LA83, MG81, OCA^{+}84, OH84, RR80b, ROW^{+}83, SBGD84, SKS84b, Sin80, SF84b, SC84a, TP82, VC83, EWG82, FT80]. regulatory [MMC83, OLHH81, SS81, Ste82a, Ste83a, YAF81]. Reichert [ILL83]. reinhardi [GO80, KSSR84, OO82]. reinhardrii [JR80]. reinhardtii [DHL84, HMKM82, JB81d, JB82, KJ82, SHRL84, WD84, WCJ83]. reinnervated [GBLH83]. related [AA83, BZ83, DCS84, FSB81, FA83, GGW83, GGV81, HLM^{+}84b, HW83, MB84a, PPR^{+}82a, SB83a, SF84b, TS83b, WMP^{+}84a, YBG81]. Relation [HCP81, BBR83, CP82b, EL84, HS80a, MB84c, WD80]. Relationship [AP84, KC83, SHH82, WHK^{+}84, BP80, CH80d, DMD80, DANH84, Fis82c, MPAM80, SM82a, FJ82a, HBO84, HRBJG80]. relationships [GL80a, GRB82, JOA84, NNRR80]. relative [BTF^{+}83, BFST83, JWBC81, PMP81b]. relaxation [Oku80, PCOK82]. relaxed [PNYO83]. relay [DMD80, DSD80b, DSD80a, TD83]. Release [PSR82, ASC84a, BKT84, CB81, FGHC80, HBR^{+}84, HR81b, KvRK^{+}84, LSJT83, MTI80, PGD^{+}84, RRS82, SCN^{+}81, SGSS^{+}81, TWW82]. released [WB81b]. releasing [APZL80, KKM84, SW84, SMT80]. relevance [GBWR83, YF83]. removal [Con84, KLF80, LK84, OKA84, SBBS84, VC84]. removes [SSBR84]. renal [BBS81, DVB^{+}81, EHK80, KSF84, SMJ80, WGRB80]. Reorganization [FHDL82, PAGG81, Wan84, FPFP81, LCA84, MMAO83, SNPE84, WMG84]. reoxygenation [LSJT83]. repair [GS84f, OKA84]. repeat [ARHG84, HMB^{+}82]. repeated [HDB82, Sch81a, YBG81]. repeating [BK83a, SGM^{+}84]. replace [LTE84]. replacement [KRB^{+}84c, OKA84]. replica [TTUI83]. replicas [HEA^{+}82, HK80a, PF83b]. replicates [KHS82]. Replication [RF80b, DBBN81, HAF82]. replicational [SPBB84]. reported [GK82]. repression [ABD84, NCL84]. required [AWB82, FNHSS84, LK84, MM81c]. requirement [MOS84, NCBP84, OB81, SC84b]. requirements [MD80b, RDY^{+}83, SK81]. requiring [Cab83, CWBB83, ML80, TWG83]. resealed [OK82]. reserpine [ST83]. reserpine-resistant [ST83]. reserve [BVC83, CHCS82]. residual [Ber80a]. residues [BO81a, BKO81, RBO83, Rot84, SSC80, ZB82]. resistance [BS80, EDK80, Mad83, SMJB83]. resistant [CK82, FBP82, HWK^{+}84, IY84b, LV81, LBM82, NRMB84, RRD83, ST83, SSC80, DCS84]. resolution [ATW83, IT82, MLSW82, PPSP83, SF81, dSK84]. resolved [HP82a]. Resonance [BM82a, GOSW82]. Respective [KNK84]. respectively [OT80a]. respiratory [NPG^{+}80]. Response [RRRS82, SH81b, BSKFW83, BSH83b, CCCR81, CGG81, CGG84, DMD80, DSD80b, DSD80a, ELST81, FJ82b, FTM82, GL80b, GB80, HEG81, IY84c, MR80a, RS81, SWGB80, Sch81b, SCP^{+}83, SF84b, SC84b, TSGS83, VV83]. responses [AHMS82, BMH^{+}83, LSC83, NJR80, OBH84, Sch81c, SCN^{+}81, WB83b, YSS82, YDL^{+}84]. responsible [BK83a]. responsive [BG82, FJ82c, Reb82]. responsiveness [MGMS84]. rest [Nac80]. restimulation [SMTP80]. resting [HOR84, IY84c]. Restoration [HHT80]. restrict [FCNS84]. Restricted [GAS82, O’B80]. result [SS84b]. results [CS83, NCL84, SKS84b]. resumption [WS81a]. Retention [HBS^{+}82]. reticulocytes [HHS83, MF84d, TWG83]. reticulopodial [TA81, TKA83]. reticulum [ACTK84, AWB82, BPM^{+}80a, BVC83, Bro81, CH83a, CHI^{+}81, CHS82, CHCS82, DYP83, FNHSS84, FHFL82, FFS^{+}82, GS80a, GBW82, GWB82, GIKR84, HTGB84, JSDM82, JSMT82, JC84, LHS82, LK84, LRW82, MACK84, ML84a, MD80a, MD80b, PL84, RMH^{+}84, SSCF84, SGSS^{+}81, TLPTV83, UPLH84, WIB81, WB81a, WB81b, Wal82a, Wal82b]. reticulum-golgi [Bro81, ML84a]. retina [ALE84, AMM84a, BMS83, DK84a, GDM^{+}80, HPG83, MMKL81, SLKB83, SL80]. retinal [AH84, BP80, BGE84, BS82d, BSNP82, GFLG^{+}84, LSM83, LHHH82, NCP84, OB81, OB82, OM82, PPSP83, Pie82, PCB83, PB84b, SSJS84, SW81a]. retinas [NDRL80]. Retinoic [JVMRK82]. retinoid [BMS83]. retinoid-binding [BMS83]. retinol [GFLG^{+}84, KKG84, LOC81, ROW^{+}83]. retinol-binding [GFLG^{+}84, KKG84, LOC81, ROW^{+}83]. retinol/cellular [LOC81]. retinula [Fri83]. retraction [Che81, GL80b]. Retrieval [AF84, KI83, PBFC^{+}84]. retrograde [ST83]. retrogradely [TI80]. return [Ric80]. Reuber [BF82c, FB82]. reveal [BBWS83, Pip84, RBH^{+}84, WWCD84]. revealed [AHPL81, AdS83, BCHG83, BLH82, FCKP84, HSS82, HK80a, Hir82, MS84e, MDA83, NLN^{+}81, PPSP83, SHT^{+}84, SFU84, SHH82, SMG81, TdS84, VBPL84, Wal82a, Wal82b, vDB80]. reveals [HEA^{+}82, HW83, SC84c]. reversal [NW80, PBW80, TT81]. Reversible [DLZ83, KSI83, KI83, AMJ^{+}81, PP80a, PP80b, WLCS84, ZSM^{+}83]. reversibly [Max82, SBC82]. reversion [SDL82, SC84b]. review [ES81b]. revision [GK82]. rhabdomeral [dCSB84]. Rheological [SWA83]. Rhizobium [WPKB83]. rhodamine [CF83, Gla83a, Gla83b, GK81b]. rhodamine-actin [Gla83a, Gla83b]. rhodopseudomonas [VF82]. ribbon [HWBC82, WFR84]. ribofuranosylbenzimidazole [EPO^{+}84, SHHR84]. ribonuclease [Liz80]. ribonucleoprotein [Ber80a, CD84, FKK83, JOM80, MWP84, MSGP81, PD80, PM83, TD82, WMP83, BK83a]. ribonucleoproteins [FLM84]. ribophorins [ACTK84, KORBS82, MACK84, RMH^{+}84]. ribosomal [FG80a, FJ82b, LL82a, SRGB83, SMGB84]. Ribosome [GS80a, BTF^{+}83, LL82a]. Ribosomes [SZ81, ACTK84, KORBS82, MACK84, MR81b, Mec81a, Mec81b, MU82, RVB80]. ribosylation [KGAB81]. ribulose [RBMM82]. rice [MA83b]. rich [CSR^{+}81, LBA^{+}84, MMR^{+}84]. ridge [PPSP83]. rigidity [EL84]. rigor [PCOK82, PNYO83]. rigor-wave [PCOK82]. ring [Bur82b, SM83b]. Rings [SLS83, Bro81, Sob84]. rise [SLRP82]. RNA [ABD84, BSR84, CSK^{+}80, EH81, FKK83, GE80, HHPB83, Jef82, KSG81, MWP84, MWKP84, MLZD83, MSGP81, MP81, PD80, Ped83, Per81b, Rag81, RC83, SGHS^{+}80, SBGD84, SLR^{+}83, SKS82, UAL84, WB83a]. RNA-associated [SKS82]. RNA-binding [MLZD83]. RNA-protein [Ped83]. RNAs [HSRL81, LDF^{+}84]. Rod [GBG82, GWG82, BM82a, PPSP83, RH82, RKH82]. rodent [CBWW80, CFN^{+}83, GCH84, NCM^{+}83]. rods [OB81, OB82, Sch81c]. Role [BRAC83, CHCS82, Con84, GKZ81, GAA83, HDRK83, KMF81, KKMK84, KKM81, LHS82, LMN83, MKB82, MBLvdM84, OBH84, PPR82b, SCS80, SBK^{+}84, TRG84, VRR80, BP83, BSSR84, BvdMN84, BR84c, Car84, CHI^{+}81, CRGS82, DK84b, DRF84, GG83, GS83, GT84, GSMR83, HOT81, HH83a, HHS^{+}81, KM82, KCB84, MPU84, MS83a, PGD^{+}84, RH81, RR80a, RA81, RR80b, RKP^{+}82, SKCJW84, SLKB83, SH82a, SB83b, SBC82, Ste82a, SR83a, TH84, TQS84, WG82b, WB82]. Roles [Gil83, LM84, SAP81, KNK84, TWG83, TSS82]. room [KRB84b]. root [GB80, SFB^{+}84, WPKB83]. rootlet [LD81]. rootlets [GGW83]. roots [SBSM84]. rose [MF84c]. rosettes [MBWH83]. rotary [FSFA84, MWI^{+}84, YT82]. rotary-shadowing [FSFA84]. Rotation [OK80, TT81]. rotational [HWJW84]. rough [DYP83, GS80a, GBWR83, KORBS82, LK84, LRW82, ML84a, PBG80, TLPTV83]. rounding [CMC81].

Rous [CH81, DWM83, EPE^{+}80, MTT80, SR81a, WYK^{+}84, WSH84, WBG83]. Rows [WE80]. rRNA [SHHR84]. rudiments [IJ82b]. ruffling [BC83, GMO84]. rupture [KSR^{+}80].

S [Vah83, COBC84, KUM^{+}84, BTJ82, KHS82, KJBR82, O’B80, RF80b]. S-phase [KHS82]. S115 [DK84b]. S3 [DBBN81]. S49 [Ste83a]. sac [ACMH83, CCKMH82, ILL83]. saccharide [VEL80]. saccharides [MPJ80]. saccharomyces [FS80, AP84, Har80a, IY84b, IY84c, KA84, KHL82, MJM83, MB84d, NLPH82, RF80a, RF80b, SBJ84, SAP81, WH82]. salivary [HMB^{+}82, TS83a]. salt [AMK83, HE81, LV81, PSSS82]. saltatory [Bec84]. Samanea [SGKC82]. same [BTF^{+}83, HA84a, RIP^{+}84, SWvK^{+}83]. saponin [Blo84, MP83, PB83b]. Saprolegnia [Hea80]. sarcolemma [PSC83]. sarcoma [AMJ^{+}81, CH81, DWM83, MTT80, RCM^{+}83, SR81a, VSLS80, WYK^{+}84, WSH84, WBG83, EPE^{+}80]. sarcomere [RMF83]. sarcomeres [WRM83]. sarcoplasmic [CH83a, CHI^{+}81, JSDM82, JSMT82, JC84, SSCF84, SGSS^{+}81]. satellite [MLSW82]. saturation [VRR80]. Saxitoxin [SSP84b]. scaffold [EHCR84, FWP84, Kuo82]. scaffolds [EL83]. scale [MHB84, SHT^{+}83]. scallop [CH83a, Nac81, SSMH82]. Scanning [GAS83b, BBS81, PPSP83, SEPB81, SBP81, SHWS83, dHLC84]. scattering [WB83b]. Sceloporus [BBF84]. Schistosoma [CKS82, SSZ83, SC82a, SCD82]. Schistosome [CRC^{+}82]. Schistosomula [SCD82, CKB^{+}80a, CKB^{+}80b, CKS82, SC82a]. Schwann [BWW^{+}80, CB81, CEC^{+}83, EKS^{+}82, KS84, MPF83a, MWA^{+}80, PB84a, RGB84, SB80b, SWGB80, SBG80, SS84a, TIS^{+}81, TQS84, TQG84, YDL^{+}84]. Schwannoma [PF83a]. Sciara [Kub82]. sciatic [PBW80]. scleroderma [BPB^{+}81]. SDS [HEA^{+}82]. Sea [KZS81, AB80, BRH82, BSH83b, CH80b, EKWR84, GG80, Gib80, GGVR82, HOW80, HAP82, HHS^{+}81, HM84c, IMS82, JE81, Kan80, KDL80, KL81, KMV82, ONO82, Oku80, PCOK82, Pip84, POS80, RF84, SHHP81, SS80a, SSL^{+}84a, SLS^{+}84, SSBB82, SBS80, TJ80, WS83, WS84, WMP^{+}84b, WNB82]. search [BC80, RLP84]. seawater [EDK80]. seawater-adapted [EDK80]. second [SB80c]. secondary [GBS80, KSC82, MSC80b]. seconds [RS83]. secretagogue [FJ82a, FJ82c, LP83b]. secretagogue-induced [LP83b]. secretagogue-responsive [FJ82c]. secretagogues [FJ82b]. secrete [KS84, SMJB83]. secreted [FSBR83, LPB84, MGK83, VMRR80]. secreting [EDK80]. Secretion [BB82b, CLHA80, APZL80, BE82, BDPS80, BBB82, BK84, BMCGO^{+}82, DMD80, FS80, GGS83, GS83, GNGTS80, HB84, JDTW80, LNA81, LH83, ME81, MDJ82, NKT80, RMABB82, RR80a, ROW^{+}83, SN80, SLRP82, Sur82, TOK^{+}84, TDM80a, VKE^{+}81, WC83]. secretions [EB83]. Secretory [BPM81, AC82, BGD80, BBR83, BVAR83, BO81b, Bro81, BK84, BMCGO^{+}82, CML^{+}81, FNNFS84, FQBN81, GS84a, GS84d, IJ82a, IJ82b, LK84, MB84a, MGK83, NDS84, NS83b, PBWK83, PWT82, RLP84, SF81, SJC80, SD82, SHP83, WIB81, WB81a]. section [PvD83, TTI80]. sectional [OCT81]. sections [BBS81, CKP84, DCG83, FLM84, FBH83, GSvdLS81, JSMT82, JSCM83, OOZF80, RMAF80, SGKC82, SEPB81, SBP81, Tok83]. sedimentation [RS82]. seed [BVC83, MA83b]. seeded [BB80a, HC83]. segment [CSHP84]. segments [BS82d]. Segregation [ACTK84, MACK84, GMRB81, HCMK81, KW81a, SB83a, WRH81, WKAH84, WH82]. selected [SR81a]. Selection [CK82, AFK^{+}84, FO80, WG82b]. Selective [CPS84, MSUC80, RUM^{+}82, RRS82, Wat84, ASC84a, FSW^{+}83, GO80, KK82, LFR80, MR82, MRB84, WB81a, PJ83]. selectively [EH81]. Self [MM81b]. Self-assembly [MM81b]. seminiferous [FKB83]. semliki [HKSF80, GKKPB83, GQW83, KKKH84, KKBG83, MBWH83, WKH80, WMH81]. senescence [GK81b]. sensing [WB83b]. sensitive [AMJ^{+}81, BC84b, KI83, MJM83, MMS^{+}84a, MA82, SG82a, VKHH82, WWKW83, WBG83, DCS84, HSC84]. sensitivity [BHJV80, GB83, KJBR82, Kie81, O’B80]. Sensory [Van83, GL80b]. separate [GN84, MdV80, NDS84]. separated [GW82, HBS^{+}82, MB82c]. separates [SM81]. Separation [ACF84, BGNT81, EHK80, SHPH84, AB81, Ber80a, QGW83, SM82b, SM82a, WWKW83]. septa [Fri82b]. Septate [Fri82b]. Sequence [SC84c, CSK^{+}80, EKWR84, ES82, PA83, WB81b]. sequence-dependent [WB81b]. sequences [Yao82]. sequestering [Kie81, Wal82a, Wal82b]. sequestration [SG84a]. Serial [MHSW81, PvD83, SSJS84]. Serial-section [PvD83]. Serotonin [TTGA82, BGNT81, RR80b, YSJ81, TTGA82]. Serum [SK81, Bar82a, BTJ82, CCCR81, CFCS82, FHDL82, GS84f, Kla80, LKJ84, RUM^{+}82, SC84b, VRR80, WSM80, ZRW80]. serum- [FHDL82]. Serum-free [SK81, Bar82a, CFCS82, Kla80, RUM^{+}82, WSM80]. serum-stimulated [GS84f]. serves [TWG83]. several [KST84, RB83, TAB80]. severely [CPS84]. severin [GWS84, YPR^{+}82]. severs [WS84]. sexual [SBC82]. Sgs [SEE84]. Sgs-4 [SEE84]. shadowed [ME82]. shadowing [FSFA84, MWI^{+}84]. Shape [JTM80, KRB^{+}84c, AFW84a, AS81a, GHH82, HSB^{+}83, HFW82, LHS82, NAK83, PF81, SSJS84]. shared [RLP84]. shares [Pol84b]. shearing [SL82b, TT83]. shedding [BBF84]. sheets [AFI^{+}81, GMSK81, OMM82]. shibirets1 [KI83]. Shift [CQB84, ELST81]. shifts [KK82]. shiverer [PB84a, SS84a]. shock [Atk81, FP81b, IY84a, IY84b, LV81, MF84d, San81, SF84b, SSSJ83, VSBL83]. shock-resistant [IY84b]. shocked [VSBL83]. short [ARHG84, BMCGO^{+}82, DB81b, KWSB84, Por84]. short-term [BMCGO^{+}82]. shown [BO81a, BKO81]. shows [IY84b]. shunted [CHI^{+}81]. shuttle [VN82]. sialic [AC81b, BO81a, BKO81, SSC80]. sialoglycoprotein [SP82a, SP82b]. sialoglycoproteins [FBST84]. sialoprotein [KSF84]. sialyltransferase [Fle81]. sidedness [JFGC82]. sides [GHL84]. Signal [AWB82, BWM80, EHS80, ES82, GBW82, GWB82, MSP84, MF84e, Pfe84, WB81b, WB83a, WIB81, WB81a, WB81b]. signaling [DMD80, DSD80b, DSD80a, Pfe84]. signalling [SBC82]. signals [DGC83, Van83]. significance [OKE81]. silica [KSR^{+}80]. silk [Liz80]. silkworm [TOK^{+}84]. silver [SHHR84]. simian [ITT80]. similar [ERH82, IY84c, MWH83]. Similarities [CH81, DK84a, LFC^{+}80]. similarly [SMT80]. Single [SMG81, ER84, EKWR84, FFRF83, GRDR82, HA84a, HT80, KSI83, MB81, MCPB80, MGG83, MPC84, RMF83, ROB^{+}84, RBPS83, VS84, WLCS84]. singlet [LL81]. sinus [CFN^{+}83, NCM^{+}83]. sinusoids [GF84]. site [BO81a, BKO81, CSHP84, EH81, ISGS81, MWP84, PKS81, WB81b]. site-specific [WB81b]. Sites [GL82, SRGB83, Ste83a, Bad81, BVAR83, BCP^{+}81, CS82, COBC84, CEC84, GS80a, HSS82, KKP84a, dPMMF^{+}81, MS84c, MMVO83, PS83a, PWM81, RKF80, RB83, RMJ83, Sch81b, SSP81, SSSP81, SSP82b, SPBB84, TV83, TTI^{+}83, VTSS80, WGRB80, dSTK81]. situ [BF82b, FJ82b, HSS82, HLSWH82, Jef82, LGF^{+}84, NNA^{+}83, PS83a, RSB84, SSZ83, SD84, TI82a, VSS83, Wal82a, WGF83]. size [KJ82, WG81, SP84c]. sized [BFL^{+}83, FSF^{+}80, WB84]. sizes [HF83]. Sizing [SG81]. skeletal [AF83b, Bad81, BSVS80, BL82a, BBCL82, CC82, CB84a, DSRF84, EM84, FN80, FSFA84, FPFP81, GLBH82, GBLH83, GKL^{+}84, GRZ83, JSMT82, JSCM83, KS83, LS80, MP80, MSK^{+}84, MIRS84, MPF83b, Moo81, MGG83, NLD^{+}82, PG84, RMS^{+}82, RRS82, SSCF84, SHT^{+}84, SBSL84, TI82a, TDS83c, TS81, TG83, WP80, WDE83, WRM83, WGRB80, Wol80a, Wri81, YS84]. skeletal- [EM84]. skeletin [TDS83c, TDS83b]. skeleton [FKS^{+}81, LBA^{+}84, ML84c, SJS84, TFGC84]. skin [MLM83, MFVPK82, SHW^{+}82a]. skinned [MSK^{+}84]. sliding [KHL82, TT84a, WZ80]. slime [DSKS84, SDK^{+}84]. sloughing [SCD82]. Slow [BL80b, BSVS80, Can82, GL81, GL83a]. slowly [CT83]. Small [FKK83, MHB84, LS82, Mad83, MWP84, MWKP84, Rod80, UAL84, WMP83]. smaller [ME81]. smooth [BSS81, BS82c, CTKJ83, CARK81, CCCR81, EM84, FFRF83, FSVW80, GDTS81, GHGT81, GAAG82, LFAG84, LBJ84, LKJ84, MB84a, MHLS81, SBCS^{+}82, Wal82a, Wal82b, WH83]. smooth-muscle-type [EM84]. Sodium [MGSC84, ACC83, AFK^{+}84, EMA83, FHFL82, LMN83, SSP84b, SLRP82, VSLS80]. sodium- [AFK^{+}84]. sol [YAF81]. solation [BP82, MGW^{+}80]. solenoid [AHRG82]. solid [HSFM82]. solidissima [SR84b]. solubilization [CLHA80]. solubilized [DB80, EMA83, SH81b]. soluble [BL83a, Gol82, GB82, KSD^{+}83, PG84]. Solution [TMA^{+}83, FFEL81, KRB84b]. solutions [SBP81]. Solvent [HM84a]. somatic [Ber80b, BCC^{+}83, KUM^{+}84, KSK^{+}82, Sin80, YBG81, Yao82]. Somatostatin [GS84d]. Some [ZS82, Cle84, EPO^{+}84, LSC83, SSHP82, WTK83]. soon [SGHS^{+}80]. sorted [MGK83]. Sorting [BK84, MS84b, AR81, SFU84]. source [RGB84]. space [TQS84]. span [GHGT81]. spanning [GKKPB83]. Spatial [GRB82, HHT^{+}84, SPBB84, Tuc84, BF82c, EKWR84, SB83b, SA83a, VSD84]. special [RA81]. Specialization [MMVO83, NBM82, PP84, RK81]. Specializations [RK80]. species [GSG82, GE80, JGSH82, KRB^{+}84c, TP82]. Specific [BSS83a, BSS83b, DBL82, GS80b, IY84c, MK84b, SY83, SSMH82, VMMC84, Wil84, ABD84, BBWS83, BGM81, BK84, CK80, CG84, CF83, DCG83, DHHG83, EPRS81, FHG^{+}81, GE84, HOR84, HAS^{+}81, HHD83, HPG83, HSGD83, ISGS81, JB81a, KKP84a, KP84, Lan82, MB84c, MCN^{+}84, MTKH84, MTR81, MMS84b, MLD82, MWA^{+}80, MFVPK82, MMIGK82, MF84d, MCG82, MSD82b, PWO84, QGW83, RW82a, Ric80, RB80, RH83, RH84a, RH84b, RLP84, RNH^{+}83, SSG^{+}83a, STRP83, TOG83, TT83, TSS82, TLPTV83, THT^{+}84, WH81, WB81b, WB83a, WGRB80, WA84, WHD^{+}84, Yao82, ZF84, GWK84, VV84]. specifically [WWS83, WW83]. Specification [Sol81]. specificities [GB82]. specificity [FGS^{+}82, FBST84, GGK82, GGVR82, JZ81, LFS^{+}83, RCH81, SWGB80, SR81b, TP82]. Specimen [SEPB81]. Spectrin [CF80, FT80, FLH^{+}81, PF81, RSB84, GGW83, JTM80, MB84c, ML84c, MM81b, NGL83, OK82, PFM^{+}84, TTI^{+}81, TTI^{+}83]. spectrin-actin [JTM80]. Spectrin-dependent [CF80]. spectrin-like [PFM^{+}84]. spectrin-related [GGW83]. spectrins [BKM82, Pol84b]. spectrometry [Sta81]. Spectroscopic [Fis82b, AO84]. spectroscopy [GHY81]. spectrum [DKDB81]. speed [Ino81b]. Sperm [GS84e, SH82b, WAN81, AdS83, AdS84, AB80, BBPR81, Bal82, BF82a, BGPB80, DTBF82, GG80, Gib80, GGVR82, GC80, IT82, KL81, KC80b, LOL81, LFR80, LM84, MM81c, MP84c, MPK84, ONO82, Oku80, PCOK82, PV84, SH82a, SLL80, TI82b, TBCM83, VBPL84, ZCH80]. sperm-egg [GGVR82, KL81]. spermatids [GM84, SLL80]. spermatocyte [IMB83]. spermatocytes [JL84, SB80a, SLL80, SL82b]. spermatogenesis [Wie81]. spermatozoa [BvGJ83, IMS82, KNH80, RW82b, SSG80, WV84]. spermatozoan [NRW82, SLIF80]. spermiogenesis [RW82c]. sphaeroides [VF82]. spike [KKBG83, MBWH83, ML80]. spinach [RZSW83, ZP80]. spinal [RE80, RGE83, SFB^{+}84]. Spindle [NKH82, SLS^{+}84, AMCG83, Baj82, BLV84, CWBB83, EJM82, HHS^{+}81, Ino81a, IWM83, Jen82, KR82b, KJBR82, Kie81, KHL82, Kub82, LPH83, ME84, MBB83, Mur80, Nic83, PHST84, RN83, SMH84, SB83b, SR84b, TH81, TPPH80, ZHM80]. spindles [KHL82, PKB84, POS80, SS80a]. spines [CRC^{+}82, LR83a]. Spisula [SR84b]. spleen [PBPD83]. spoke [BLH82]. spokes [HPRL81, PHRL81]. spontaneous [BS82a, MTI80]. spore [WB82, Wei82]. spread [ASM^{+}84, TD82]. spreading [AUT83, BZ84, CRH81, HS80b, LN82, LMN83, Nac80, PWWS80, Sob84, SB82c, SHBS80, SMTT83]. sprouting [PMP81a, PMP81b, PJ83]. squamous [NS83a]. squid [GS84b, LOL81]. SRP [WIB81, WB81a, WB81b]. stability [BSR84, BTL84, HH83a, JZ81, Kir80, LHS82, LGF^{+}84, MMS84b, WSM^{+}82]. Stabilization [RKM^{+}81, AHRG82, Con84, LCK84, ML84c, VH82]. stabilized [WSO^{+}81]. Stable [ML82, DM83, FSBR83, MD82, NCBP84, SB80a]. Stacking [MHC80]. stage [FP81a, SB83a, VH82]. stages [NW80, RBPF80, SCSC81]. staining [BYW81, LdMM^{+}84, MP83, MHC80, SHHR84, Tok83, WDE83, Wal82a]. stalk [PHRL81]. standard [SBP81]. Staphylococcus [FBS^{+}81, MFP82]. starfish [ER84, IMS82, Mab83, Sch81b]. start [SBJ84]. starved [CPS84]. state [BTJ82, HOR84, HS84, IY84c, PNYO83, WS84]. states [PG83c, SPEL84, ZSM^{+}83]. static [SWB84]. stationary [Sin82]. statocyst [SSA80]. Status [GK81b, RBS84]. stays [OHA^{+}84]. steady [WS84]. steady-state [WS84]. stem [GSMR83, OYNT83, YSJ81]. step [AFK^{+}84, BCLT^{+}80, ML80]. steps [BPM81, BVC83, WM84]. stereo [OOZF80]. stereocilia [TDM80b, TS83b, TEDS83]. stereological [BPM^{+}80a, MHLS81, GWQ^{+}84]. sterna [GBF84]. steroid [DK84b, JDTW80]. steroidogenesis [HOT81]. sterol [BN83, CH80d, FDKF82, NB81, TT83]. sterol-specific [TT83]. sterols [SLS83]. stiffness [OAOB81]. stimulate [BC84b, EB83, LBJ84]. Stimulated [HBR^{+}84, GCF^{+}83, GS84f, MHCG80, SJ80, SN80, SSLB84, WCZG84, MKD80, PB83a]. stimulates [DVB^{+}81, EMC83, PWWS80, PWSW80, RGB84, SDG^{+}84, WMP^{+}84a]. stimulating [BGS81, CVGM83, GS80b, VC83]. Stimulation [GS83, NFS81, SLRP82, BMH^{+}83, BL82b, DSD80a, GBWR83, MTH84, SWGB80, SKS84b, SGKD83]. stimulator [LTE84]. stimulators [JWG84]. stimuli [SSA80, TSGS83]. stimulus [SHO^{+}84]. stimulus-affected [SHO^{+}84]. Stoichiometry [LA80b]. stomatitis [AL84, BKL^{+}82, BS83, FSBR83, FR81, ML80, PAS84, PBS84, RBPF80, RF81, SWvK^{+}83, WFGB83, WMH81, WWDYW83]. stomatitus [Mil80]. storage [BBR83, CHS82, HSRL81, KBD82, KSD^{+}83, TTGA82]. stored [ERH82, SHHP81]. strain [KBD82, SMT80]. strains [KC83, SHRL84, WPKB83]. strand [ECM81]. stranded [KS83]. stratified [NS83a]. streaming [NBSW81, NSW82]. Stress [BF82b, DANH84, KSSP83, SSJ83, Sin82, WGF83]. stressed [BO81b]. stretched [MSK^{+}84]. Striated [FAP81, SBSM84, EPRS81, HWJW84, SPHE80, WD80, WCJ83]. striatum [AM81]. stringent [IY84c]. stripped [GBWR83, PBG80]. Strongylocentrotus [KR82b, KJBR82, HAP82, HOW80, PV84, POS80]. Structural [BRH82, CT82, EWB83, EDK80, EM81, GDMG81, HR81b, HH81b, IHJ82, KA84, KH84, KLD^{+}82, MSRB84, PG83c, SvB81, TAB80, VB81, WL80, ZKF84, BSH83b, BBWS83, DGC83, Fri83, GT84, HW83, Kan80, Kan83, KKP84b, LFC^{+}80, Ste83a, SHT^{+}83, TTUI83, ZSR^{+}82, LL82a]. structurally [JC84, WSO^{+}81]. Structure [JW83, LKR^{+}83, Man82, MSGP81, MS84e, MP81, MP82, Pol82, RGR80, RVH84, SHT^{+}84, SD80, SLIF80, AFI^{+}81, ACH^{+}81, ARHG84, BK83a, Bal82, BPB^{+}81, BR84b, Bur84, CH83a, CP82b, Deu84, DBBN81, EP80, FSFA84, FCKP84, Gal81, GB82, HMB^{+}82, HK80b, JOA84, KL82, KHS80b, KRB84b, Kub82, KTW83, LFC^{+}80, LR81, LP83a, LL81, LGF^{+}84, Mad83, MGW^{+}80, Mur84b, NAH83, OFH^{+}82, PD80, PL84, PPSP83, RV80, Rob81, SMS84a, SKN83, SR82b, Sea82, SJM81, SJ80, Sto84, TMA^{+}83, UE83, WFR84, WL80, ZCR80, vDB80]. structures [ABD84, Bad81, BCHG83, BW80, DPS80, DANH84, EP80, GSB81, GC80, HLS81, HOW80, HZM80, KZS81, LV81, MB84c, MCN^{+}84, MP83, MPF83b, RFUC84, SHW^{+}82a, TT81, WPKB83, YHS^{+}82]. studied [HKCM83, HM84a, HSGD83, KUM^{+}84, KNHK84, TS81, WMP83]. Studies [CRH81, CSW80, DCK84, GGK83, HO84b, HSC80, JGSH82, Kie81, PFM^{+}84, PKB84, PNYO83, RCH81, RH81, RBMM82, SB80b, SWGB80, SBG80, TA81, TKA83, WWKW83, WSLSJ83, ARL^{+}80, ACMH83, AC83, BCLT^{+}80, BS83, BG82, BPM^{+}80a, BNGN84, CS82, CLJ^{+}81, DTT81, DGS80, ERN83, FGHC80, FFGC84, GS84a, GDTS81, GHY81, GWQ^{+}84, Gri80, GHH82, HHSS81, HHS^{+}81, JRM^{+}80, JRB80, KMF81, KHJG84, KKG84, LP83a, LFAG84, LTAP^{+}81, LN81, LNB84, MWM^{+}81, MSC83, NFG82, NS83a, NBSW81, OWMB82, PL83a, PT83, QGW84, SHW^{+}82a, SDC80, SKKC84, SO82, Ste82a, TDS83c, TDS83b, TQS84, WRK81, YDL^{+}84, ZF81, ZFP84, AMM84b]. Study [CE80, CSH81, AMRE81, AC81b, AUT83, BLG^{+}80, BRR^{+}83, BSVS80, Bow80, CGA^{+}82, CH80b, Cle84, DWP81, DSKS84, Ekb81, FSFA84, FDSWW84, GRR^{+}84, GDL80, KZS81, KW81a, Lan82, MHLS81, MBC80b, Moo81, MR82, PYC^{+}82, PPRR82, Por84, RPHA81, RMAF80, RMJ83, SLLP82, SGW^{+}82, SAW83, SGSS^{+}81, TAGG80, TMS84, TTI^{+}81, TTUI83, WTSAC^{+}81, WOBR80, WB83b, YMF84, ZRW80, ZH82]. studying [Can80, CMM81, MP84a]. Subcellular [BHSC83, Gla83b, HSRL81, SHO^{+}84, VEL80, WB83a, AC81b, BPM80b, JB81b, MCPB80, RACGB81, RKP^{+}82, WTSAC^{+}81, WSH84, Mec81a]. subclass [HCK80, MGK83]. subcompartments [TV83]. subconfluent [VPMF84]. subfibers [MW80]. subfractionation [SD84, WTSAC^{+}81]. subfractions [ACTK84, BSS83a, HP82a, MACK84, SC81]. subfragment [BC84a, PCB83]. subfragment-1 [BC84a, PCB83]. subjects [NFG82]. submembrane [JHCC84]. Submicromolar [KW84]. subnuclear [FBB82]. suborganellar [RZSW83]. subpopulations [MP82, OT80a]. subsequent [BO81a, BKO81, OLHH81, WKAH84]. subset [PWO84]. substance [BR84b, HM84a]. substances [OP81]. Substituted [SNdC^{+}84]. substitution [GDL80, HA80, HHS^{+}84, HS84]. substrata [GJRO^{+}83, Hat81, LPB84, WLFD84]. Substrate [JKHT81, JH84, Axe81, Blo84, FFEL81, GAS82, GCH84, GBE84, NCH83, PG84, RCM^{+}83, RK80, SK83]. substrate-associated [Blo84]. substrates [CHRT83, KC83, LN81]. substratum [BN83, CS82, CG84, DKDB81, Gri80, MCN^{+}84, MTT80, WYK^{+}84]. Substructure [GH82a, AFRS83, GBG82, GWG82, Wol80a]. Subunit [FSFA84, AWB82, CML^{+}81, HZM80, HHPB83, MGG83, MSD82b, NGL83, PSSS82, SHZN83, Ste83a, Ste84b, SHT^{+}83, TOK^{+}84]. subunits [BF82c, CF84b, FB82, Lan80, LL82a, LOL81, Man82, MWM^{+}81, Pip84, RBMM82, SBG^{+}81, SHT^{+}83, VF82, VBT84, ZF84]. successive [ML84a]. suckling [GN84]. Suctoria [Rud80]. sufficient [Mar80]. sugar [SHB80]. sugars [Sta81]. suggestions [Deu84]. suggests [GG83]. sulfate [AF83b, AMM84a, AMM84b, HOMR82, HCJV83, Ioz84, KLF80, LFGR82, LSC83, LLM82, LH83, RVH84, SK83, WHK^{+}84]. sulfated [ED84, GZR^{+}80, MGK83, RVH84]. sulfates [OWMB82]. sulfatoxygalactosylacylalkylglycerol [LS81b]. sulfhydryl [ZCH80]. sulfhydryl-induced [ZCH80]. sulfotransferase [TFAP81, TZW83]. sulfoxide [FK80, Fuk80, YF83]. sulphated [HO84a, HBO84, HTKC82]. superior [DK81a]. supersensitivity [MIRS84]. supplemented [CFCS82]. support [GCL83, GLMW84, TQS84]. supports [Kla80]. suppresses [WMP^{+}84a]. supranucleosomal [ZF84]. surf [Mur80, SR84b]. Surface [ABBG82, BO80, CVGO81, GF84, KOB82, MNP^{+}84, SJ80, AdS84, AF83b, AA83, AC82, ASM^{+}84, BM82c, BOH81, BW80, BF84a, BSS83b, BF84b, CKS82, CP82a, CRC^{+}82, CGG81, CGG84, DBL82, EWG82, EW83, FNNFS84, FFEL81, FS80, FSBR83, FLH^{+}81, FPFP81, GKKPB83, GGVR82, GGV81, GA84, GSMR83, GAS83b, GS84e, HER81, HEO82, HO84b, HH81a, HRBJG80, HG80c, HDRK83, HPM82, HS80b, Ioz84, JOH^{+}84, JNS^{+}84, JHCC84, KCM83, Kho82, KL81, KNH80, LSC83, LBA^{+}84, LSJT83, LS81b, LA83, LA84, MW83b, MS84b, MNST82, MEKT84, MPJ80, MHB84, MTT80, MDA83, MP84c, MPK84, NCH83, NS83b, OFO83, PG82, PBFC^{+}84, PHM81, PMP81a, PFP^{+}82, RS83, RMT84, RGB84, RH81, Rau84, RKF80, RW82b, RBPSB84, RB82a, RBS84, RBPF80, SBG80, SC82a, SCD82, Sch81b]. surface [Sea82, SH81a, SH82a, SH82b, SSP81, SSSP81, SSP82a, SKKC84, SB83a, SA82, SCM83, SHB80, SK83, SSC80, SR83a, SWvK^{+}83, SH81b, TOT^{+}84, TGPC82, TFGC84, TMA^{+}83, WH81, WFGB83, WG82a, WYF81, ZH82, ZB82]. Surfaces [RH82, RKH82, CB82, CRH81, CFML83, HSFM82, HSW82, Oli82, PR84, SMG81, Wei80, dSK84, dHLC84]. surround [SLS83]. surrounds [ZHM80]. susceptibility [BG82]. suspended [GH82b]. suspension [ASA83, GW82, NW80, RBPS83]. suspensions [BWM84]. SV3T3 [PL83b]. SV40 [BSH83a, NCBP84]. swimming [Gib80]. Swiss [NCBP84, SGW^{+}82]. Switching [GL83b]. symbionts [HMM82]. sympathetic [CGG81, JRM^{+}80, JRB80, ST83]. symport [KCGH84]. synapses [DCG83, MNP^{+}84]. Synapsin [DCG83, DHHG83, HSGD83]. synaptic [CK80, CH80a, CW80, DHHG83, FD82, HCK80, HSGD83, JRM^{+}80, KSI83, MTR81, MS84c, MH84, San82]. Synaptonemal [IMB83]. synaptosomes [DHHG83, MMIGK82]. synchronized [DTT81, DBBN81]. syncytial [WMW84]. syndrome [IBB^{+}81, NFG82]. Synechococcus [LYG81]. synechocystis [GWG82, WGG80, GBG82]. Synemin [GRL82, PL83c]. synergistic [GJRO^{+}83, GLBS81]. synexin [Cre81]. synovial [AFW84a, AFW84b]. synthase [GKZ81]. Synthesis [Atk81, BJG83, CKFS82, CDvdM80, DPK84, GHB82, GBF84, HTKC82, KPG^{+}82, KMP^{+}82, MDJ82, RVB80, RW82a, Wal84, Wri81, ZP80, AT83, ABD84, BSR84, BSG81, BKT84, BGD80, BE82, BJ83, BZ84, BC80, BS81c, BR84c, BB82c, CPS84, CMF82, CHCS82, CH83b, CML^{+}81, DMD80, EK84, FM80, GSG82, HOR84, HR80, HH83b, HAP82, IY84a, IY84b, JB81d, JB82, LK83, LBJ84, LH83, MR80b, MLMS81, ML83, MCG82, PG83a, PL83b, PS81, Rao81, RS81, RSP83b, RMABB82, RKP^{+}82, RNH^{+}83, SGHS^{+}80, SBGD84, SKS84b, SKS82, Sch83b, SRGB83, SZ81, SCM83, SGKD83, SY80, TS83a, TSGS83, VPMF84, VA83, VKE^{+}81, WR83, WC80, WG81, Wie81, WA83]. synthesize [WHD^{+}84]. synthesized [ERH82, GWQ^{+}84, HSB^{+}82, HO84a, IJ82a, KHF81, NM84, OWMB82, PJ83, QGW84, RFML84, RVH84, RF81, SMGB84, SC81, TOK^{+}84, VMRR80, WFGB83, WH83]. synthesizing [MB80b, WIB81, WB81a]. synthetase [DYP83, Deu84, MS83b, NDRL80, Pie82]. synthetic [BDP^{+}84, RC83, Wei80, WG82b]. system [ACF84, AS81b, BCSC82, BGM81, BC82, BSV84, BO81b, BLH82, CBWW80, CBJ^{+}82, CG84, DK81b, DCG83, DSKS84, GK81d, GVK81, HHSS81, Hir82, JSC84, Kie81, MS84a, MSC80a, MSC80b, MSC83, POYF81, RFML84, RCJS82, RUB84, RNH^{+}83, SFS81, SLS81, WH81, WP82b, YF81]. systems [Raf84, SW81b].

T [BNH81, GRB82, Mar80]. T-lymphoma [BNH81]. tagged [MPU84]. tail [BB80a, KKP84a]. tails [AHRG82, BGPB80]. Tandemly [YBG81, Sch81a]. tannic [BBCL82, MP83, Sea82]. tannic-acid-mediated [Sea82]. tarantula [LKR^{+}83]. target [GRB82, SCP^{+}83]. Targeting [GK84]. tau [DCK84, LFF^{+}81]. Taxol [AFSF^{+}81, CWBB83, MPH82, MF84c, PH81, SSBB82, TM84, Cab83, LR84, MBB83, Val82]. taxol-dependent [Val82]. Taxol-induced [MF84c, TM84]. Taxol-requiring [CWBB83]. technique [BO80, Ott83]. techniques [BCC^{+}83, HM84a, LS81b, SEPB81]. tectum [MMKL81]. teleost [BSNP82, OB81, OB82, PB84b]. telopeptides [GB82]. telson [KL82]. Temperature [KK82, OJJ^{+}84, AMJ^{+}81, BL80a, BG83, HHS^{+}84, HS84, KI83, KRB84b, MJM83, WWKW83]. temperature-sensitive [AMJ^{+}81, KI83, MJM83, WWKW83]. template [HL81]. Temporal [EKWR84, GHY81, BF82c]. tendon [CF84a]. tension [PT83, SCN^{+}81, SM83c, WD80]. Teratocarcinoma [GSMR83, OYNT83, PJM^{+}81]. term [BTHL81, BMCGO^{+}82, Kla80, RGM^{+}80, SCP^{+}83, SGKD83]. Terminal [BTJ82, VH82, BHJ83, BBP80, BLHC81, Bur82b, CC82, DCG83, DHHG83, HSGD83, KMF84, LH84, LYG81, MTM83, Pen83, Rot84, SSCF84, SY83, VMMC84, YS84]. terminal-specific [DCG83, DHHG83, HSGD83]. terminally [Wat84]. terminals [Har80b, HCK80, MSB80]. termini [CDLJ81]. terminus [SSG^{+}83b]. termite [YT82]. test [HFSK82]. testicular [LS81b]. testis [MC80]. tests [ACTK84]. tetanized [SGSS^{+}81]. tetraethylammonium [LHHH82]. Tetrahymena [AAW^{+}84, Den80, Goo83, MW80, NKT80, TTUI83, WZ80, WA84, Wol80a, YBG81, Yao82]. tetraurelia [ARL^{+}80, ERN83, FG80a, LTAP^{+}81, LN81, MDM^{+}81, MKG81, PWT82, RSP^{+}83a, RN81]. tetrodotoxin [EMA83]. tetrodotoxin-binding [EMA83]. tetroxide [Sma81]. th [PFP^{+}82]. their [AFI^{+}81, Can82, Can83, CP82b, CHCS82, CVC^{+}84, DGC83, EBS84, EP80, FDSWW84, GG83, GCL83, HWAK83, HOMR82, HS84, IRMM84, KC83, Mil80, NKH82, OLHH81, PL84, RIP^{+}84, RKP^{+}82, SCD82, SSSP81, SG82b, SSLB84, TSS82, TBD81, VVBZ84, WM84, WKAH84, WOBR80, YF83]. theophylline [DBBN81]. theories [HHS^{+}84]. Thermal [LGF^{+}84, Gol82]. thermophila [WA84]. these [BO81a, BKO81]. thiamine [RB82b]. thick [CTKJ83, EHK80, KL82, KS83, LKR^{+}83]. Thin [TG83, ECM81, FLM84, OOZF80, PP84, SGKC82, TTI80]. Thin-section [TTI80]. those [ERH82, YHS^{+}82]. thread [DSKS84, SDK^{+}84]. threads [SM83c]. Three [BB82c, GR80, Heu80, Mur84b, NAH83, NBBS83, ATW83, CKP84, CHH^{+}83, CH82, CEC84, FWP84, HSW82, HWM83, KS83, PDK83, RLP84, SAW83, TDM80a, WC80]. Three-dimensional [GR80, Heu80, Mur84b, NAH83, NBBS83, ATW83, CKP84, FWP84, HSW82, SAW83]. three-stranded [KS83]. thrombin [BSKFW83, CB82, HG80a, HG80b, ISGS81, LMN83, SSLB84]. thrombin-induced [LMN83]. thrombin-stimulated [SSLB84]. Thrombocytopoiesis [ZFP84]. Thrombospondin [RMABB82, MLHM84, MDJ82, MRB84]. throughout [BC80, SH83]. thummi [KB81c]. Thylakoid [JB81d, JB82, DRF84, HMKM82, WOBR80, WD84, ZP80]. thyone [IT82, TI82b, TBCM83]. thyroid [BGNT81, CGM81, CVGM83, FHDL82, GW82, Her83, Kho82, NW80, WP82a]. thyroid-stimulating [CVGM83]. Thyrotropin [KKM84, SW84, APZL80, FHDL82, SMT80]. thyrotropin-free [FHDL82]. Thyrotropin-releasing [KKM84, SW84, APZL80, SMT80]. tight [BBWS83, EWB83, GDRS83, LC80, Mad83, MTH84, PT83]. tight-binding [BBWS83]. tight-junction [Mad83]. tightly [CDLJ81]. time [Gla83a, HA84b, MNW81]. time-course [MNW81]. time-lapse [HA84b]. Timing [WA84, SB83b]. tip [BBS81, HA80, MHCG80]. Tissue [SSG^{+}83a, AMM84a, AMM84b, BSG81, BGD80, Bec84, Bro81, Can80, DB81b, FFEL81, FM80, GSvdLS81, JWG84, KSMF84, LL82b, MTR81, MdV80, OFO83, PWO84, RGM^{+}80, SSZ83, SB80b, Sea82, SKKC84, WB84]. tissue-type [LL82b]. tissues [BKO81, BB82a, BB80d, BGS81, GCH84, GBE84, KST84, PAO84, PPR^{+}82a, WWCD84, WGRB80, WKA^{+}83, YAF81]. toad [HBS^{+}82, KMF81, MKD80, SW81a, SLIF80]. tomato [Vah83]. tonofilament [Lan82]. tonofilaments [FWG^{+}81]. Topographical [WRH81]. Topography [TZW83, VF82, JFGC82, PTH83, WBPO80, Wal82a]. topology [Rot84, SR82a, SVW^{+}83]. Torpedo [CSR^{+}81, Fro84, GNW^{+}84, JFGC82, KHS80b, MMIGK82, SWF84]. total [Axe81]. toxin [DS80, Fis82c, FBS^{+}81, HH81b, KUM^{+}84, KGAB81, KPPC83, LFC^{+}80, MSD82a, SO80, SSO84]. toxin-ganglioside [KPPC83]. toxins [KUM^{+}84]. tracer [ZFP84]. tracheal [GWP82]. track [KC83]. tract [MFVPK82]. Traction [HWS82]. traffic [RF84]. trailing [Che81]. trans [GBB^{+}82, GQW83, HTGB84, RB82b]. trans-Golgi [RB82b]. transcabamoylase [NLPH82]. Transcellular [KCGH84, SHP83]. transcribed [Sch81a, SHHR84]. transcript [LCK84]. Transcription [KSSR84, EPO^{+}84, FP81b, FKK83, FW80, KSG81, MLZD83]. transcriptional [KSD^{+}83, SHHR84]. Transcytosis [Her83, PAS84, PBS84, VSS83]. transduce [SSA80]. transduction [MSP84, Pfe84]. transendothelial [FCNS84]. Transepithelial [MBP^{+}83, PS83b, Mad83, MBC83]. transfer [BHJ83, BRS^{+}82, HBS^{+}81, LS82, MLZD83, MF84e, PJ83, RB82a, RMU84, SHH82, SG82a, SJM81, TRSS81]. transferred [MKD80]. transferrin [ELS^{+}83, HHS83, HT83, KvRK^{+}84, LTE84, SWvK^{+}83]. Transformation [DWM83, ITT80, WBG83, COBC84, NCBP84, SSCP80, YAF81]. transformation-associated [COBC84]. Transformation-sensitive [WBG83]. transformed [BSH83a, BGE84, CBW81, CLFC83, CH82, EPE^{+}80, HER81, HOMR82, MTT80, RCM^{+}83, RSP83b, SC82a, SR81a, SK81, VSLS80, WYK^{+}84, WBG83]. transglutaminase [FFEL81]. Transient [BL83b, FR81, JFS84, VV84, VC84, DGC83, IYVV83, JNS^{+}84, SDG^{+}84]. transients [ER84, LWWP84]. Transit [WP82b, BZN82]. transition [BTJ82, OM83, WHJ83]. translates [GO80]. translation [FW80, Mec81b]. Translational [HAP82, AL84, LLB84, RKP^{+}82]. translocating [SB82c]. Translocation [WIB81, WB81a, WB81b, ACTK84, BHSC83, BF84a, GBW82, GWB82, HA84b, MACK84, MD80a, MD80b, MLD82]. translocations [MP84a]. Transmembrane [Pfe84, SH81a, AdS83, AdS84, AWB82, BKL^{+}82, Fis82a, SHT^{+}84, SR82a, SS81]. transmission [CKP84, HH81b, KSI83, SHWS83]. transmit [SSA80]. transmitter [HR81b, MTI80]. transplantation [SP84a]. Transport [BBG^{+}84, KFV83, RBPF80, RF81, RUB84, SP84b, SVW^{+}83, AR81, ACC83, BBR83, BKL^{+}82, BS83, BL80b, BGM81, BTHL81, Bro81, BNM82, BF84b, CHCS82, FDKF82, FG80b, GZR^{+}80, GPHV83, GS84b, GDL80, GMO84, GL82, GBB^{+}82, GQW83, HSB^{+}82, HA84a, HGP84, KHS^{+}80a, LGB84, LNA81, LLB84, MMKL81, MSD82a, MBP^{+}83, MG81, MS83a, NBM82, OJJ^{+}84, PYC^{+}82, PS83b, RW82a, RIP^{+}84, RK84, RBPSB84, ROW^{+}83, RMU84, SJC80, SFU84, ST83, SP84a, Ste82a, SHP83, TMW82, VVM81, WSH84, SMJ80]. transported [Ber80b, GL81, GKKPB83, GPDM81, LSM83, LW81, SW81b, SW81a, SWvK^{+}83, TI80, WFGB83, WSB^{+}80]. transporting [MPMBC80, MIS^{+}80]. transthyretin [KKG84]. transverse [CC82, WRM83]. trapped [CML^{+}81, TPR82]. traverse [EB84, RIP^{+}84]. treadmilling [Kir80]. treated [GSB81, HAEC81, JL84, PLP81, QGW83, RFML84, SHHR84, WKW82, YS84]. treatment [CSR^{+}81, FHFL82, HS83, KOS83, MGG83, SG82a, SB82b]. triad [CB84a, MPF83b]. triads [FSFA84, MSPF83]. triazido [MF84a]. trichocyst [BCLT^{+}80, LTAP^{+}81]. trichocysts [RN81]. Trifluoperazine [BSP83, HCH^{+}81]. triggered [RCH81]. triggers [WS81a]. trimeric [SB82a]. trinitrobenzene [MF84a]. triphosphate [MW80, PSR82]. triple [GB82]. tripolar [KRB84a]. Triton [IMS82, LCA84, TFGC84]. Triton-extracted [LCA84]. tRNA [DYP83, Deu84]. trophoblast [GAS^{+}83a]. Tropomyosin [BHP83, LMYM84, DGS80, GDTS81, Leh81, SG80]. Tropomyosin-enriched [LMYM84]. Tropomyosin-like [BHP83]. troponin [TS81]. Trypanosoma [OBC^{+}84, Wil84]. trypanosomes [GIKR84]. TTB [MF84a]. tube [KZS81, WB82, Wei82, Wei82]. Tubular [BU84, CFCS82, DVB^{+}81, SMTP80]. tubule [CAL^{+}82, DGC83, GLMW84, HEO82, KNBSF84, RKMF84]. tubules [CC82]. Tubulin [BCP^{+}81, KPW^{+}83, ME84, NFG82, Olm81, SSL^{+}84b, AMCG83, AP84, AB80, BSR84, BHA82, BM82b, BVAR83, BMF84, BTL84, BNM82, CH83b, CBL84, DB83, GBM84, HOW80, HC84, KR82b, KJBR82, KSSR84, KA84, Kir80, KRB84b, LR83b, LSM^{+}84, LOL81, LL81, LFF^{+}81, MR81a, MLSR83, MHC80, ML84d, NFS81, OAOB81, PH81, PDK83, SSL^{+}84a, SMH84, SLS^{+}84, SBGD84, Ste83b, SC84c, SR84b, SS84c, TAC84, WS83, WWS83, WW83, WTK83, WD83]. tubulin-binding [BVAR83]. tubulin-containing [HOW80]. tubulins [BGS^{+}83, WMP^{+}84b]. Tumor [WCZG84, BB80b, BK84, CCJC82, DKDB81, DK84b, DCS84, DWS84, GOSW82, GKZ81, IJ82a, IJ82b, IHJ82, KBD82, LA83, ME81, Mar80, MPF83a, MGK83, Reb82, SNPE84, SDM84, Sur82, WYF81]. tumoral [TLPTV83]. tumorigenicity [HLM^{+}84b]. tunicamycin [BCF84, CK82, OLH^{+}81]. tunicamycin-resistant [CK82]. turning [GB80]. Turnover [RPHA81, BAWL81, BPM80b, CS83, HF83, SH83, SM80, SWB84]. TW [GGW83, PFM^{+}84]. TW-260 [GGW83]. TW-260/240 [GGW83]. twist [DTBF82, OK80]. Two [EK84, FSB81, HLM84a, MJM83, MJ83, PC80, BB82a, BL80b, BVC83, BMS83, HSW82, HEA^{+}82, HTKC82, HPM82, IY84b, JC84, KW84, MFP80, Mil81b, MS80b, OT80b, RFML84, RF84, RWD84, RBPF80, RMU84, SFU84, SBG^{+}81, SF84b, SCM83, SPEL84, TTGA82, TWG83, TSS82, TPHL80, TSGS83, WC80, WGG80, ZSM^{+}83, FO80]. Two-dimensinal [PC80]. Two-dimensional [MJ83, HSW82, HPM82, MS80b, FO80]. two-way [RF84]. Type [DB81a, AP84, BFL^{+}83, BLH82, BGS^{+}83, CEC^{+}83, CW80, CDvdM80, DWS84, EZ84, ELST81, EM84, FGS^{+}82, FML83, FSF^{+}80, FSVW80, GDH^{+}83, Gol82, GB82, GDH82, HHT80, KMF84, LSN^{+}84, LLM82, LL82b, LFS^{+}83, LGF^{+}84, MOS84, MJM83, MHGM82, MWI^{+}84, MKG81, MBLvdM84, MRB84, NHF^{+}83, SSG^{+}83a, SLK^{+}84, SB83a, WAN81, WOBR80, ZF84, PST^{+}83]. type-dependent [BGS^{+}83]. type-specific [SSG^{+}83a, ZF84]. types [BO81a, BKO81, BKM82, CSS80, CGCS80, GB82, GSS82b, MB82c, RMAF80, RLP84, SGB^{+}81, TTGA82, YF81]. tyrosinase [HPM^{+}83]. Tyrosine [O’B80, AT83, CH82, CL81, GSH80, GCH84, GBE84, NCH83, RCM^{+}83]. tyrosine-protein [CH82]. tyrosinolation [NFS81, NFG82]. Tyrosyl [CFW84]. tyrosylated [WWS83, WW83].

U1 [MWP84, WMP83]. U4 [MWKP84]. ubiquitin [TWG83]. UDP [GKZ81]. ultrahigh [PPSP83]. ultramicrotomy [PP81]. Ultrastructural [AM81, BBCL82, ES81a, FLM84, GSS^{+}84, JSMT82, JSCM83, LdMM^{+}84, MRFF80, MHGM82, RSDH82, SWF84, AB81, CVAK80, DSRF84, FFGC84, HA80, KW81a, MSB80, MBC80b, OWMB82, SPHE80, VN82, VKE^{+}81, WOBR80, RH84a]. Ultrastructure [CGW82, ILL83, NW80, SJS84, SMJ80, TD82, CSR^{+}81, CW80, GVK81, HHS^{+}84, HS84, LD81, NBBS83, Sma81, TA81]. ultrastructure-preservation [HS84]. ultrathin [FBH83, GSvdLS81, JSMT82, JSCM83, RMAF80]. Ultraviolet [LPH83]. umbilical [MKW^{+}82]. uncoated [MFA83]. uncoating [BSSR84, SSBR84]. uncoupling [ASA83]. under-sulfated [RVH84]. undergo [PSSS82]. undergoing [Atk81, SLR^{+}83]. underlying [TTI80]. undifferentiated [Wri81, ZSM^{+}83]. undulating [SLIF80]. unembedded [FCKP84]. unfertilized [HM84c, SHHP81, SSBB82, WS84, WZ83]. unicellular [NBBS83]. uniform [SSJS84, TH81, TG83]. unilamellar [CAZF84]. unique [DSKS84, GV82, HH83b, IWM83, KZS81, SMJB83, WTK83, ZK83]. uniqueness [MP81]. unit [SJS84, SHT^{+}83]. univalent [FNS81, SB80c]. unlinked [WHSK80]. Unphosphorylated [WYK^{+}84]. unraveled [AFRS83]. unresponsive [Har80a]. Unsaturated [Cre81]. unstimulated [FBRP84]. untreated [JL84]. unusually [KMF84]. unwarranted [BRS81]. upon [AC83, Bar82b, IYVV83, MGG83, PSR82, SBCS^{+}82, SBBS84, SD82, WDE83]. Uptake [MMKL81, MDSH80, PS83a, UPLH84, BB81, FCO83, GML^{+}80, NHF^{+}83, OLHH81, OP81, PVPB82, QCB84, ST83, VVBZ84]. urchin [AB80, BRH82, BSH83b, CH80b, EKWR84, GG80, Gib80, GGVR82, HOW80, HAP82, HHS^{+}81, HM84c, IMS82, JE81, KZS81, Kan80, KDL80, KL81, KMV82, ONO82, Oku80, PCOK82, Pip84, POS80, RF84, SHHP81, SS80a, SSL^{+}84a, SLS^{+}84, SSBB82, SBS80, TJ80, WS83, WS84, WMP^{+}84b]. urchins [WNB82]. uridine [UAL84]. urinary [CFN^{+}83, KMF81, KLD^{+}82, NCM^{+}83, RV80]. urogenital [CFN^{+}83, NCM^{+}83]. urokinase [LSN^{+}84, LL82b, SLK^{+}84]. urokinase-type [LSN^{+}84, SLK^{+}84]. Use [CL80a, GSvdLS81, SSG^{+}83b, BBCL82, ERH82, GK81b, LFS^{+}83, PK81, TOG83, WSB^{+}80]. using [Blo84, BM83, Can80, CH80b, CLJ^{+}81, DKDB81, DSRF84, FG80b, HM84e, JGSH82, KWM82, KJ82, Liz80, MLSW82, MNP^{+}84, MF84a, MLD82, MSD82b, PKB84, PNYO83, RMJ84, SLS^{+}84, SMS84a, SWF84, SF84b, SEE84, SS84c, TLPTV83, WSLSJ83, WMENS82, YMF84, ZB83]. uterine [WP80]. utilization [MDSH80, RW82a, RGM^{+}80].

V [DSD80a, LFS^{+}83, MHGM82, MWI^{+}84, Mec81a, MRB84]. vacuolar [MMS^{+}84a, MSC80a, MSC80b, MSC83]. vacuolation [OP81]. vacuole [MB80a, MSC83, OT80a]. vacuoles [KSC82, MAG82, SDC80]. Val [SGM^{+}84]. Val-Gly-Val-Ala-Pro-Gly [SGM^{+}84]. validation [FB82]. value [Gal81]. vanadate [Oku80]. Vanadium [EMC83]. variable [KJ82]. variant [SL82a, SC84b, WA84]. Variants [TEL84, FDSWW84, HJ80, Wri84a, WNB82]. variations [JB81d, JB82]. variegatus [SS80a]. variety [YAF81]. various [BO81a, BKO81, BPM80b, BFL^{+}83, CGCS80, CPS84, KRB^{+}84c, MCPB80, RA81, SJ80]. vas [MHLS81]. vascular [BSS81, FSVW80, GHGT81, GCL83, GAAG82, HHS^{+}84, ISGS81, LS82, LFAG84, LBJ84, LKJ84, MB84a, WGF83]. vasopressin [RKM^{+}81]. vasopressin-induced [RKM^{+}81]. vasoproliferative [GDM^{+}80]. vectorial [MD80b]. vehicles [ST83]. vein [MKW^{+}82]. venom [MTI80]. ventral [SFB^{+}84]. Ventricular [SGB^{+}81]. venules [HSS82, STR84]. vertebrae [BSVS80]. vertebrate [BH83, BS82c, BMS83, CTKJ83, CD84, FSB81, Pol84b, SH83, Sch81c, WP80, WRM83]. very [RR80a]. vesicle [AR81, ACF84, CBWW80, Heu80, HSGD83, JRM^{+}80, MTR81, MFA83, MHC80, MFP82, MH84, PL84, PDK83, PWT82]. vesicle-associated [PL84]. vesicle-planar [ACF84]. vesicle-specific [MTR81]. vesicle-tubulin [MHC80]. vesicles [AF83a, BP80, BF84b, CK80, CH80a, CAZF84, CL81, DHHG83, Fle81, GS84b, GSH^{+}84, GE84, HCK80, HSGD83, KPW^{+}83, KMCAS80, KORBS82, MBWH83, Max82, MFA83, MSRB84, MSUC80, MS83a, Mil80, MB84d, MHW83, PBG80, PSR82, PvD83, PK81, PSBF83, PWT82, RBPF80, RFLR81, SCS80, ST83, SLS83, SMJ80, SBBS84, SHWS83, SWvK^{+}83, TZW83, TKM83, YFM83, ZK83]. Vesicular [SWvK^{+}83, VVM81, AL84, BKL^{+}82, BS83, FSBR83, FR81, Kie81, LNA81, ML80, Mil80, PAS84, PBS84, RBPF80, RF81, SSO84, WFGB83, WMH81, WWDYW83]. vesiculation [BP82]. vessel [DBK80]. VI [LN81, Mec81b]. via [Bur82b, SWvK^{+}83, Wei80]. Video [Ino81b, HA84b, IT82, SSL^{+}84a]. view [FG80b]. viewed [AW82]. views [ATW83, Cle84]. VII [AC81b, TWH80, TBR80, WWKW83]. VIII [OFH^{+}82, WTSAC^{+}81]. VIII/Von [OFH^{+}82]. villin [BM83, MB82b]. Vimentin [BL83a, BS81b, BZ83, BZ84, BL82a, CLFC83, GRL82, KSMF84, LVL^{+}82, Law83, ML83, NCL84, SFS81, SCSC81, TMS84]. vinblastine [LFF^{+}81]. vinblastine-induced [LFF^{+}81]. Vinculin [PSC83, ASG83, BH83, GDTS81, Ott83, SNPE84, Sin82, Ott83]. vinculin-binding [Ott83]. Viral [GBB^{+}82, RIP^{+}84, ACC83, FR81, GMRB81, GQW83, MBP^{+}83, PS83b, RF81]. virally [CH82, QGW83]. viridis [HMM82]. virtually [FYC82]. virus [AMJ^{+}81, AL84, BKL^{+}82, BS83, BG83, CFW84, CH81, DWM83, EPE^{+}80, FSBR83, GKKPB83, GQW83, HKSF80, ITT80, KKKH84, KKBG83, MBWH83, MRHS81, MTT80, ML80, Mil80, MG83b, NPG^{+}80, PAS84, PBS84, RBPSB84, RBPF80, SR81a, SWvK^{+}83, VSLS80, WYK^{+}84, WFGB83, WKH80, WWDYW83, WSH84, WBG83]. virus-3 [MG83b]. virus-infected [NPG^{+}80]. virus-transformed [MTT80, SR81a, VSLS80, WYK^{+}84]. viruses [AC81a, RCM^{+}83, RBPS83, WMH81]. visceral [CCKMH82]. viscoelastic [SH82c]. Viscometric [MFP80]. viscosity [ZS82]. viscous [NW82]. vision [Sch81c]. visual [BGM81]. Visualization [AO84, GBS80, ME82, OOZF80, PBWK83, Tok83, VSS83, BM83, BF82b, EP80, Heu80, HH81a, KFS81, LSBPG84, Mil81a, SKWF84, UAL84]. visualized [ABBG82, BKL^{+}82, ECM81, FLM84, GS80a, GDL80, IT82, Sea82, TLPTV83]. vitamin [ROW^{+}83, TRN^{+}82, THT^{+}84]. vitelline [CH80b]. vitellogenin [SY80]. vitro [BMF82, BB80a, BKB80, BVC83, BBB82, BB80c, BL82b, BB81, BB82c, CSH81, DPK84, DHS^{+}80, EHESF82, FC82, FJ82c, FR81, GS84a, GGSB83, GB81, GK81b, GL83b, GLMW84, HL81, HLM84a, HO84a, HZM80, HF80, KSG81, KFS81, KM82, KR82b, KSSR84, KOST80, KPG^{+}82, KMP^{+}82, LSM^{+}84, LN81, LH84, MP80, MHSW81, MF84b, MB82c, MU82, MOV83, MF84c, MM81b, OJJ^{+}84, OT80a, OKE81, OO82, PF83a, PBPD83, PH81, PG83a, Per81a, SDL81, SKCJW84, SHW^{+}82a, STRP83, SCS84, SH82b, SK83, SO82, SHT^{+}83, SKA^{+}83, SD82, SS84c, TdS83a, TRG84, WIB81, WB81a, WM81, WWS83, WC84, WMP83, YHS^{+}82]. vivo [BMF82, BYW81, BHR81, BRR^{+}83, BVC83, BB80c, FG80b, GGK82, GHH82, HWBC82, Kie81, Kir80, LP83a, LN81, OKE81, OO82, Pai84, PF83a, RPHA81, SMH84, SKKC84, SWB84, TMS84, WGRB80, WWS83]. voltage [GL80a, JGSH82, LCA84, Ste82a]. volume [JTM80, MKB81, YS84]. volumes [BOH81]. Volvox [GK81d, GVK81, GK82]. vs [KOB82]. VSV [Mil80].

wall [ILL83, MBC80b, VB81]. Wallerian [PBW80, SB80b]. walls [GBS80]. waltlii [GK81c, Sch81a]. was [KUM^{+}84]. washes [PC80]. Water [PR84, Cle84, KCGH84]. wave [LFR80, PCOK82]. waveform [BFW80, HWJW84]. way [RF84]. Weak [Max82, PO81]. weakly [OP81]. web [Bur82b]. Weibel [WOM82]. weight [BCSC82, BM82b, BL82a, GSG82, GBF84, GLS83, GE80, HJ81, IY84a, JB81a, JB81b, JW83, LLS81, MFVPK82, Pol84b, SMJB83, SCM83, SIM^{+}84]. weights [MS84a]. WGA [DCS84]. WGA-resistant [DCS84]. WGA-sensitive [DCS84]. WGAr [DCS84]. WGAs [DCS84]. Wheat [PV84, EL84, LA80b, MMKL81, MRPK82, OMG82, SH82c, SSC80, VS82, dSTK81, LSM83]. wheat-germ [dSTK81]. wheeleri [RN83]. whereas [GCH84]. which [OT82, RGB84, SSHP82]. white [Her80]. whole [KB81a, LCA84, TOG83]. whole-mount [KB81a]. wide [MTR81]. Widespread [BLV84, BSV84, BB80d, GCH84]. width [TS83b]. wild [AP84, BLH82, EZ84, GDH82, HHT80, MKG81, WAN81, WOBR80]. wild-type [AP84, BLH82, EZ84, GDH82, HHT80, MKG81, WAN81]. Willebrand [OFH^{+}82, WOM82, WM84]. withdrawal [SMTP80]. within [BRS^{+}82, BBWS83, BO81b, GSG82, HGW84, LGF^{+}84, MB84c, ONO82, PLMS81, RS83, RMF83, SR81a, SM83a, TI82a, WE80]. without [GKKPB83, HBS^{+}81, ML84b, TT81]. wounds [FFGC84]. wt [BB80d, CG84, DBG^{+}81, HM84c, MBLvdM84, PV84, RQHR81, RBH^{+}84, TFGC84, Vah83, WS84, WSM^{+}82].

x [SKN83, BBS81, GR80, IMS82, LP83a, NLN^{+}81, PL83a, SEPB81, SBP81, TFGC84, UE83, Wie81]. X-100 [IMS82]. X-100-insoluble [TFGC84]. x-irradiated [SKN83]. X-ray [BBS81, GR80, LP83a, NLN^{+}81, PL83a, SEPB81, SBP81, UE83]. xenogenic [DBL82]. Xenopus [SR83b, AH84, BNGN84, CML^{+}81, ERH82, FO80, FKK83, FKS^{+}81, GG83, GWK84, HSRL81, KNHK84, KRN83, KSK^{+}82, LS81a, MLZD83, MLNK83, RB83, SP84a, WN81]. Xenos [RN83]. xeroderma [GS84f]. xyloside [KHJG84, TS83a]. xylosides [SGKD83].

y-1 [GO80, HMKM82]. Yeast [FNNFS84, SNS83, BAC^{+}84, FNHSS84, IY84a, KA84, KHL82, MBC80b, NS83b, RSDH82, SBJ84, SM83b, SEM80]. yield [BHP83, MSRB84, KMCAS80]. ylcarbamate [HA80]. yolk [ACMH83, CCKMH82, ILL83]. young [Rod80].

Z [CP82b, OMM82, YRS83]. Z-band [CP82b]. Z-disk [OMM82]. Z-line [YRS83]. Z-lines [YRS83]. Z-protein [OMM82]. Zea [BB82c]. zein [BB82c, HSRL81]. zona [ED84, SH82a]. zonal [MB82c]. zone [GRZ83, Ko84]. zones [NB81]. Zonulae [SG84b]. ZP [ED84]. ZP-2 [ED84]. zymosan [SNS83].

References

Armelin:1983:GHM

[AA83] M. C. Armelin and H. A. Armelin. Glucocorticoid hormone modulation of both cell surface and cytoskeleton related to growth control of rat glioma cells. Journal of Cell Biology, 97 (2):459–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

              org/content/97/2/459.

Allis:1984:PPH

[AAW^{+}84] C. D. Allis, R. L. Allen, J. C. Wiggins, L. G. Chicoine, and R. Richman. Proteolytic processing of h1-like histones in chromatin: a physiologically and developmentally regulated event in Tetrahymena micronuclei. Journal of Cell Biology, 99(5): 1669–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

              org/content/99/5/1669.

Asai:1980:EAA

[AB80] D. J. Asai and C. J. Brokaw. Effects of antibodies against tubulin on the movement of reactivated sea urchin sperm flagella. Journal of Cell Biology, 87(1):114–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

              http://jcb.rupress.org/content/87/1/114.

Aist:1981:MCS

[AB81] J. R. Aist and M. W. Berns. Mechanics of chromosome separation during mitosis in Fusarium (Fungi imperfecti): new evidence from ultrastructural and laser microbeam experiments. Journal of Cell Biology, 91(2):446–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/91/2/446.

Anderson:1982:SDR

[ABBG82] R. G. Anderson, M. S. Brown, U. Beisiegel, and J. L. Goldstein. Surface distribution and recycling of the low density lipoprotein receptor as visualized with antireceptor antibodies. Journal of Cell Biology, 93(3):523–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/93/3/523.

Andersson:1984:PSG

[ABD84] K. Andersson, B. Björkroth, and B. Daneholt. Packing of a specific gene into higher order structures following repression of RNA synthesis. Journal of Cell Biology, 98(4):1296–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
98/4/1296.

Anderson:1981:IIR

[ABG81] R. G. Anderson, M. S. Brown, and J. L. Goldstein. Inefficient internalization of receptor-bound low density lipoprotein in human carcinoma A-431 cells. Journal of Cell Biology, 88 (2):441–??, February 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/88/2/441.

Axelrod:1981:FNI

[ABSC81] D. Axelrod, H. C. Bauer, M. Stya, and C. N. Christian. A factor from neurons induces partial immobilization of nonclustered acetylcholine receptors on cultured muscle cells. Journal of Cell Biology, 88(2):459–??, February 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/88/2/459.

Alonso:1981:DEM

[AC81a] F. V. Alonso and R. W. Compans. Differential effect of monensin on enveloped viruses that form at distinct plasma membrane domains. Journal of Cell Biology, 89(3):700–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/89/3/
700.

Amar-Costesec:1981:ASM

[AC81b] A. Amar-Costesec. Analytical study of microsomes and isolated subcellular membranes from rat liver. VII. Distribution of protein-bound sialic acid. Journal of Cell Biology, 89 (1):62–??, April 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/89/1/62.

Arvan:1982:PMR

[AC82] P. Arvan and J. D. Castle. Plasma membrane of the rat parotid gland: preparation and partial characterization of a fraction containing the secretory surface. Journal of Cell Biology, 95 (1):8–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/95/1/8.

Andrews:1983:FFS

[AC83] L. D. Andrews and A. I. Cohen. Freeze-fracture studies of photoreceptor membranes: new observations bearing upon the distribution of cholesterol. Journal of Cell Biology, 97(3): 749–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/3/749.

Alonso-Caplen:1983:MGT

[ACC83] F. V. Alonso-Caplen and R. W. Compans. Modulation of glycosylation and transport of viral membrane glycoproteins by a sodium ionophore. Journal of Cell Biology, 97(3):659– ??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/3/659.

Akabas:1984:SOD

[ACF84] M. H. Akabas, F. S. Cohen, and A. Finkelstein. Separation of the osmotically driven fusion event from vesicle-planar membrane attachment in a model system for exocytosis. Journal of Cell Biology, 98(3):1063–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/3/1063.

Allan:1981:RHO

[ACH^{+}81] J. Allan, G. J. Cowling, N. Harborne, P. Cattini, R. Craigie, and H. Gould. Regulation of the higher-order structure of chromatin by histones H1 and H5. Journal of Cell Biology, 90 (2):279–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/2/279.

Amenta:1983:DFL

[ACMH83] P. S. Amenta, C. C. Clark, and A. Martinez-Hernandez. Deposition of fibronectin and laminin in the basement membrane of the rat parietal yolk sac: immunohistochemical and biosynthetic studies. Journal of Cell Biology, 96(1):104–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/1/
104.

Amar-Costesec:1984:SPT

[ACTK84] A. Amar-Costesec, J. A. Todd, and G. Kreibich. Segregation of the polypeptide translocation apparatus to regions of the endoplasmic reticulum containing ribophorins and ribosomes. I. Functional tests on rat liver microsomal subfractions. Journal of Cell Biology, 99(6):2247–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/6/2247.

Aguas:1983:RTG

[AdS83] A. P. Aguas and P. Pinto da Silva. Regionalization of transmembrane glycoproteins in the plasma membrane of boar sperm head is revealed by fracture-label. Journal of Cell Biology, 97(5):1356–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/5/1356.

Aguas:1984:HDT

[AdS84] A. P. Aguas and P. Pinto da Silva. High density of transmembrane glycoproteins on the flagellar surface of boar sperm cells. Journal of Cell Biology, 99(2):655–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/2/655.

Allen:1983:NVA

[AF83a] R. D. Allen and A. K. Fok. Nonlysosomal vesicles (acidosomes) are involved in phagosome acidification in Paramecium. Journal of Cell Biology, 97(2):566–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/2/566.

Anderson:1983:AAR

[AF83b] M. J. Anderson and D. M. Fambrough. Aggregates of acetylcholine receptors are associated with plaques of a basal lamina heparan sulfate proteoglycan on the surface of skeletal muscle fibers. Journal of Cell Biology, 97(5):1396–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/5/
1396.

Allen:1984:RLM

[AF84] R. D. Allen and A. K. Fok. Retrieval of lysosomal membrane and acid phosphatase from phagolysosomes of Paramecium caudatum. Journal of Cell Biology, 99(6):1955–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/6/
1955.

Aebi:1981:CAS

[AFI^{+}81] U. Aebi, W. E. Fowler, G. Isenberg, T. D. Pollard, and P. R. Smith. Crystalline actin sheets: their structure and polymorphism. Journal of Cell Biology, 91(2):340–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/91/2/340.

Ash:1984:ASP

[AFK^{+}84] J. F. Ash, R. M. Fineman, T. Kalka, M. Morgan, and B. Wire. Amplification of sodium- and potassium-activated adenosinetriphosphatase in HeLa cells by ouabain step selection. Journal of Cell Biology, 99(3):971–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/3/971.

Aebi:1983:FSK

[AFRS83] U. Aebi, W. E. Fowler, P. Rew, and T. T. Sun. The fibrillar substructure of keratin filaments unraveled. Journal of Cell Biology, 97(4):1131–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/4/1131.

Antin:1981:TIP

[AFSF^{+}81] P. B. Antin, S. Forry-Schaudies, T. M. Friedman, S. J. Tapscott, and H. Holtzer. Taxol induces postmitotic myoblasts to assemble interdigitating microtubule-myosin arrays that exclude actin filaments. Journal of Cell Biology, 90(2):300–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
90/2/300.

Aggeler:1984:CCS

[AFW84a] J. Aggeler, S. M. Frisch, and Z. Werb. Changes in cell shape correlate with collagenase gene expression in rabbit synovial fibroblasts. Journal of Cell Biology, 98(5):1662–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/5/
1662.

Aggeler:1984:CMG

[AFW84b] J. Aggeler, S. M. Frisch, and Z. Werb. Collagenase is a major gene product of induced rabbit synovial fibroblasts. Journal of Cell Biology, 98(5):1656–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/5/1656.

Anderson:1984:IIP

[AH84] R. E. Anderson and J. G. Hollyfield. Inositol incorporation into phosphoinositides in retinal horizontal cells of Xenopus laevis: enhancement by acetylcholine, inhibition by glycine. Journal of Cell Biology, 99(2):686–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/2/686.

Aubin:1982:IBC

[AHMS82] J. E. Aubin, J. N. Heersche, M. J. Merrilees, and J. Sodek. Isolation of bone cell clones with differences in growth, hormone responses, and extracellular matrix production. Journal of Cell Biology, 92(2):452–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/92/2/452.

Adams:1981:CPM

[AHPL81] G. M. Adams, B. Huang, G. Piperno, and D. J. Luck. Central-pair microtubular complex of Chlamydomonas flagella: polypeptide composition as revealed by analysis of mutants. Journal of Cell Biology, 91(1):69–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/91/1/69.

Allan:1982:PCH

[AHRG82] J. Allan, N. Harborne, D. C. Rau, and H. Gould. Participation of core histone “tails” in the stabilization of the chromatin solenoid. Journal of Cell Biology, 93(2):285–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/93/2/
285.

Alitalo:1982:EMP

[AKM^{+}82] K. Alitalo, E. Kuismanen, R. Myllylä, U. Kiistala, S. Asko-Seljavaara, and A. Vaheri. Extracellular matrix proteins of human epidermal keratinocytes and feeder 3T3 cells. Journal of Cell Biology, 94(3):497–??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/94/3/497.

Anderson:1984:ARA

[AKT84] M. J. Anderson, F. G. Klier, and K. E. Tanguay. Acetylcholine receptor aggregation parallels the deposition of a basal lamina proteoglycan during development of the neuromuscular junction. Journal of Cell Biology, 99(5):1769–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/5/
1769.

Atkinson:1984:CTE

[AL84] P. H. Atkinson and J. T. Lee. Co-translational excision of alpha-glucose and alpha-mannose in nascent vesicular stomatitis virus G protein. Journal of Cell Biology, 98(6):2245–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
98/6/2245.

Adler:1984:ECL

[ALE84] R. Adler, J. D. Lindsey, and C. L. Elsner. Expression of cone-like properties by chick embryo neural retina cells in glial-free monolayer cultures. Journal of Cell Biology, 99(3): 1173–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/3/1173.

Allen:1981:M

[All81] R. D. Allen. Motility. Journal of Cell Biology, 91(3):??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
91/3/148s.

Ariano:1981:ULC

[AM81] M. A. Ariano and A. I. Matus. Ultrastructural localization of cyclic GMP and cyclic AMP in rat striatum. Journal of Cell Biology, 91(1):287–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/91/1/287.

Abraham:1983:MAB

[AMCG83] I. Abraham, M. Marcus, F. Cabral, and M. M. Gottesman. Mutations in alpha- and beta-tubulin affect spindle formation in Chinese hamster ovary cells. Journal of Cell Biology, 97(4): 1055–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/4/1055.

Atkinson:1981:RRR

[AMJ^{+}81] M. M. Atkinson, A. S. Menko, R. G. Johnson, J. R. Sheppard, and J. D. Sheridan. Rapid and reversible reduction of junctional permeability in cells infected with a temperaturesensitive mutant of avian sarcoma virus. Journal of Cell Biology, 91(2):573–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/91/2/573.

Andrews:1983:DII

[AMK83] S. B. Andrews, J. E. Mazurkiewicz, and R. G. Kirk. The distribution of intracellular ions in the avian salt gland. Journal of Cell Biology, 96(5):1389–??, May 1983. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/5/1389.

Aquino:1984:ILCa

[AMM84a] D. A. Aquino, R. U. Margolis, and R. K. Margolis. Immunocytochemical localization of a chondroitin sulfate proteoglycan in nervous tissue. I. Adult brain, retina, and peripheral nerve. Journal of Cell Biology, 99(3):1117–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/3/
1117.

Aquino:1984:ILCb

[AMM84b] D. A. Aquino, R. U. Margolis, and R. K. Margolis. Immunocytochemical localization of a chondroitin sulfate proteoglycan in nervous tissue. II. Studies in developing brain. Journal of Cell Biology, 99(3):1130–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/3/1130.

Aikawa:1981:FFS

[AMRE81] M. Aikawa, L. H. Miller, J. R. Rabbege, and N. Epstein. Freeze-fracture study on the erythrocyte membrane during malarial parasite invasion. Journal of Cell Biology, 91(1): 55–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/91/1/55.

Adams:1982:EPC

[AMS82] C. J. Adams, K. M. Maurey, and B. Storrie. Exocytosis of pinocytic contents by Chinese hamster ovary cells. Journal of Cell Biology, 93(3):632–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/93/3/632.

Arsenault:1984:VEI

[AO84] A. L. Arsenault and F. P. Ottensmeyer. Visualization of early intramembranous ossification by electron microscopic and spectroscopic imaging. Journal of Cell Biology, 98(3): 911–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/3/911.

Adams:1984:RAT

[AP84] A. E. Adams and J. R. Pringle. Relationship of actin and tubulin distribution to bud growth in wild-type and morphogeneticmutant Saccharomyces cerevisiae. Journal of Cell Biology, 98 (3):934–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/3/934.

Antakly:1980:CCM

[APZL80] T. Antakly, G. Pelletier, F. Zeytinoglu, and F. Labrie. Changes of cell morphology and prolactin secretion induced by 2-Br-alpha-ergocryptine, estradiol, and thyrotropin-releasing hormone in rat anterior pituitary cells in culture. Journal of Cell Biology, 86(2):377–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/86/2/377.

Abrahamson:1981:ESE

[AR81] D. R. Abrahamson and R. Rodewald. Evidence for the sorting of endocytic vesicle contents during the receptor-mediated transport of IgG across the newborn rat intestine. Journal of Cell Biology, 91(1):270–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/91/1/270.

Allan:1984:HOS

[ARHG84] J. Allan, D. C. Rau, N. Harborne, and H. Gould. Higher order structure in a short repeat length chromatin. Journal of Cell Biology, 98(4):1320–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/4/1320.

Adoutte:1980:BSE

[ARL^{+}80] A. Adoutte, R. Ramanathan, R. M. Lewis, R. R. Dute, K. Y. Ling, C. Kung, and D. L. Nelson. Biochemical studies of the excitable membrane of Paramecium tetraurelia. III. Proteins of cilia and ciliary membranes. Journal of Cell Biology, 84 (3):717–??, March 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/84/3/717.

Alhanaty:1981:CEM

[AS81a] E. Alhanaty and M. P. Sheetz. Control of the erythrocyte membrane shape: recovery from the effect of crenating agents. Journal of Cell Biology, 91(3):884–??, December 1981. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/91/3/884.

Allen:1981:DSM

[AS81b] R. D. Allen and L. A. Staehelin. Digestive system membranes: freeze-fracture evidence for differentiation and flow in Paramecium. Journal of Cell Biology, 89(1):9–??, April 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/89/1/9.

Armelin:1983:CRC

[ASA83] M. C. Armelin, R. C. Stocco, and H. A. Armelin. Control of rat C6 glioma cell proliferation: uncoupling of the inhibitory effects of hydrocortisone hormone in suspension and monolayer cultures. Journal of Cell Biology, 97(2):455–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/97/2/455.

Aderem:1984:SDA

[ASC84a] A. A. Aderem, W. A. Scott, and Z. A. Cohn. A selective defect in arachidonic acid release from macrophage membranes in high potassium media. Journal of Cell Biology, 99(4): 1235–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/4/1235.

Ascoli:1984:LAH

[Asc84b] M. Ascoli. Lysosomal accumulation of the hormone-receptor complex during receptor-mediated endocytosis of human choriogonadotropin. Journal of Cell Biology, 99(4):1242–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/4/1242.

Avnur:1983:AIA

[ASG83] Z. Avnur, J. V. Small, and B. Geiger. Actin-independent association of vinculin with the cytoplasmic aspect of the plasma membrane in cell-contact areas. Journal of Cell Biology, 96 (6):1622–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/6/1622.

Aynardi:1984:HEC

[ASG84] M. W. Aynardi, P. M. Steinert, and R. D. Goldman. Human epithelial cell intermediate filaments: isolation, purification, and characterization. Journal of Cell Biology, 98(4):1407–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
98/4/1407.

Azzarone:1984:MAR

[ASM^{+}84] B. Azzarone, H. Suarez, M. C. Mingari, L. Moretta, and A. S. Fauci. 4f2 monoclonal antibody recognizes a surface antigen on spread human fibroblasts of embryonic but not of adult origin. Journal of Cell Biology, 98(3):1133–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/3/1133.

Adlakha:1983:EPI

[ASWR83] R. C. Adlakha, C. G. Sahasrabuddhe, D. A. Wright, and P. N. Rao. Evidence for the presence of inhibitors of mitotic factors during G1 period in mammalian cells. Journal of Cell Biology, 97(6):1707–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/6/1707.

Acheson:1983:CCM

[AT83] A. L. Acheson and H. Thoenen. Cell contact-mediated regulation of tyrosine hydroxylase synthesis in cultured bovine adrenal chromaffin cells. Journal of Cell Biology, 97(3):925– ??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/3/925.

Atkinson:1981:SHS

[Atk81] B. G. Atkinson. Synthesis of heat-shock proteins by cells undergoing myogenesis. Journal of Cell Biology, 89(3):666–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
89/3/666.

Aggeler:1983:HRT

[ATW83] J. Aggeler, R. Takemura, and Z. Werb. High-resolution threedimensional views of membrane-associated clathrin and cytoskeleton in critical-point-dried macrophages. Journal of Cell Biology, 97(5):1452–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/97/5/1452.

Amato:1983:DAS

[AUT83] P. A. Amato, E. R. Unanue, and D. L. Taylor. Distribution of actin in spreading macrophages: a comparative study on living and fixed cells. Journal of Cell Biology, 96(3):750–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/3/
750.

Aaronson:1981:OCN

[AW81] R. P. Aaronson and E. Woo. Organization in the cell nucleus: divalent cations modulate the distribution of condensed and diffuse chromatin. Journal of Cell Biology, 90(1):181–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/90/1/
181.

Aggeler:1982:IED

[AW82] J. Aggeler and Z. Werb. Initial events during phagocytosis by macrophages viewed from outside and inside the cell: membrane-particle interactions and clathrin. Journal of Cell Biology, 94(3):613–??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/94/3/613.

Anderson:1982:SRP

[AWB82] D. J. Anderson, P. Walter, and G. Blobel. Signal recognition protein is required for the integration of acetylcholine receptor delta subunit, a transmembrane glycoprotein, into the endoplasmic reticulum membrane. Journal of Cell Biology, 93 (2):501–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/93/2/501.

Axelrod:1981:CSC

[Axe81] D. Axelrod. Cell-substrate contacts illuminated by total internal reflection fluorescence. Journal of Cell Biology, 89 (1):141–??, April 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/89/1/141.

Bloom:1984:CCY

[BAC^{+}84] K. S. Bloom, E. Amaya, J. Carbon, L. Clarke, A. Hill, and E. Yeh. Chromatin conformation of yeast centromeres. Journal of Cell Biology, 99(5):1559–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/5/1559.

Bader:1981:DDB

[Bad81] D. Bader. Density and distribution of α-bungarotoxin-binding sites in postsynaptic structures of regenerated rat skeletal muscle. Journal of Cell Biology, 88(2):338–??, February 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/88/2/338.

Bayne:1984:EMO

[BAF84] E. K. Bayne, M. J. Anderson, and D. M. Fambrough. Extracellular matrix organization in developing muscle: correlation with acetylcholine receptor aggregates. Journal of Cell Biology, 99(4):1486–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/4/1486.

Bainton:1981:DL

[Bai81] D. F. Bainton. The discovery of lysosomes. Journal of Cell Biology, 91(3):??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/91/3/66s.

Bajer:1982:FAM

[Baj82] A. S. Bajer. Functional autonomy of monopolar spindle and evidence for oscillatory movement in mitosis. Journal of Cell Biology, 93(1):33–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/93/1/33.

Balhorn:1982:MSC

[Bal82] R. Balhorn. A model for the structure of chromatin in mammalian sperm. Journal of Cell Biology, 93(2):298–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/93/2/
298.

Besterman:1983:PID

[BALR83] J. M. Besterman, J. A. Airhart, R. B. Low, and D. E. Rannels. Pinocytosis and intracellular degradation of exogenous protein: modulation by amino acids. Journal of Cell Biology, 96(6):1586–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/6/1586.

Barnes:1982:EGF

[Bar82a] D. W. Barnes. Epidermal growth factor inhibits growth of A431 human epidermoid carcinoma in serum-free cell culture. Journal of Cell Biology, 93(1):1–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/93/1/1.

Barrantes:1982:OFM

[Bar82b] F. J. Barrantes. Oligomeric forms of the membrane-bound acetylcholine receptor disclosed upon extraction of the M(r) 43,000 nonreceptor peptide. Journal of Cell Biology, 92(1): 60–??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/1/60.

Besterman:1981:EPF

[BAWL81] J. M. Besterman, J. A. Airhart, R. C. Woodworth, and R. B. Low. Exocytosis of pinocytosed fluid in cultured cells: kinetic evidence for rapid turnover and compartmentation. Journal of Cell Biology, 91(3):716–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/91/3/716.

Bergen:1980:HTP

[BB80a] L. G. Bergen and G. G. Borisy. Head-to-tail polymerization of microtubules in vitro. Electron microscope analysis of seeded assembly. Journal of Cell Biology, 84(1):141–??, January 1980.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/84/1/141.

Brennessel:1980:REP

[BB80b] B. A. Brennessel and D. K. Biswas. Regulated expression of the prolactin gene in rat pituitary tumor cells. Journal of Cell Biology, 87(1):6–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/87/1/6.

Bulinski:1980:ILH

[BB80c] J. C. Bulinski and G. G. Borisy. Immunofluorescence localization of HeLa cell microtubule-associated proteins on microtubules in vitro and in vivo. Journal of Cell Biology, 87(3): 792–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/87/3/792.

Bulinski:1980:WDM

[BB80d] J. C. Bulinski and G. G. Borisy. Widespread distribution of a 210,000 mol wt microtubule-associated protein in cells and tissues of primates. Journal of Cell Biology, 87(3):802–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
87/3/802.

Burr:1981:VUP

[BB81] F. A. Burr and B. Burr. In vitro uptake and processing of prezein and other maize preproteins by maize membranes. Journal of Cell Biology, 90(2):427–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/90/2/427.

Beyer:1982:QTE

[BB82a] E. C. Beyer and S. H. Barondes. Quantitation of two endogenous lactose-inhibitable lectins in embryonic and adult chicken tissues. Journal of Cell Biology, 92(1):23–??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/92/1/23.

Beyer:1982:SEL

[BB82b] E. C. Beyer and S. H. Barondes. Secretion of endogenous lectin by chicken intestinal goblet cells. Journal of Cell Biology, 92 (1):28–??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/1/28.

Burr:1982:TMZ

[BB82c] F. A. Burr and B. Burr. Three mutations in Zea mays affecting zein accumulation: a comparison of zein polypeptides, in vitro synthesis and processing, mRNA levels, and genomic organization. Journal of Cell Biology, 94(1):201–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/94/1/201.

Boyd:1982:MBC

[BBB82] A. E. Boyd, W. E. Bolton, and B. R. Brinkley. Microtubules and beta cell function: effect of colchicine on microtubules and insulin secretion in vitro by mouse beta cells. Journal of Cell Biology, 92(2):425–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/92/2/425.

Brunschwig:1982:UOI

[BBCL82] J. P. Brunschwig, N. Brandt, A. H. Caswell, and D. S. Lukeman. Ultrastructural observations of isolated intact and fragmented junctions of skeletal muscle by use of tannic acid mordanting. Journal of Cell Biology, 93(3):533–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/93/3/533.

Bernstein:1984:ILC

[BBF84] S. A. Bernstein, D. J. Breding, and S. K. Fisher. The influence of light on cone disk shedding in the lizard, Sceloporus occidentalis. Journal of Cell Biology, 99(2):379–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/99/2/379.

Burwen:1984:TEG

[BBG^{+}84] S. J. Burwen, M. E. Barker, I. S. Goldman, G. T. Hradek, S. E. Raper, and A. L. Jones. Transport of epidermal growth factor by rat liver: evidence for a nonlysosomal pathway. Journal of Cell Biology, 99(4):1259–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/4/1259.

Bennett:1983:ANP

[BBLY83] F. C. Bennett, H. Busch, M. A. Lischwe, and L. C. Yeoman. Antibodies to a nucleolar protein are localized in the nucleolus after red blood cell-mediated microinjection. Journal of Cell Biology, 97(5):1566–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/97/5/1566.

Bretz:1980:DTG

[BBP80] R. Bretz, H. Bretz, and G. E. Palade. Distribution of terminal glycosyltransferases in hepatic Golgi fractions. Journal of Cell Biology, 84(1):87–??, January 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/84/1/87.

Baccetti:1981:HDS

[BBPR81] B. Baccetti, A. G. Burrini, V. Pallini, and T. Renieri. Human dynein and sperm pathology. Journal of Cell Biology, 88(1): 102–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/88/1/102.

Becich:1983:ITS

[BBR83] M. J. Becich, M. Bendayan, and J. K. Reddy. Intracellular transport and storage of secretory proteins in relation to cytodifferentiation in neoplastic pancreatic acinar cells. Journal of Cell Biology, 96(4):949–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/4/949.

Bulger:1981:ASE

[BBS81] R. E. Bulger, R. Beeuwkes, and A. J. Saubermann. Application of scanning electron microscopy to X-ray analysis of frozenhydrated sections. III. Elemental content of cells in the rat renal papillary tip. Journal of Cell Biology, 88(2):274–??, February 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
88/2/274.

Bhorjee:1983:MAS

[BBWS83] J. S. Bhorjee, S. L. Barclay, A. Wedrychowski, and A. M. Smith. Monoclonal antibodies specific for tight-binding human chromatin antigens reveal structural rearrangements within the nucleus during the cell cycle. Journal of Cell Biology, 97 (2):389–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/2/389.

Bravo:1980:SDP

[BC80] R. Bravo and J. E. Celis. A search for differential polypeptide synthesis throughout the cell cycle of HeLa cells. Journal of Cell Biology, 84(3):795–??, March 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/84/3/795.

Blikstad:1982:DMS

[BC82] I. Blikstad and L. Carlsson. On the dynamics of the microfilament system in HeLa cells. Journal of Cell Biology, 93 (1):122–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/93/1/122.

Burridge:1983:NPA

[BC83] K. Burridge and L. Connell. A new protein of adhesion plaques and ruffling membranes. Journal of Cell Biology, 97(2):359–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
97/2/359.

Bennett:1984:DMS

[BC84a] H. Bennett and J. Condeelis. Decoration with myosin subfragment-1 disrupts contacts between microfilaments and the cell membrane in isolated Dictyostelium cortices. Journal of Cell Biology, 99(4):1434–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/4/1434.

Besterman:1984:PER

[BC84b] J. M. Besterman and P. Cuatrecasas. Phorbol esters rapidly stimulate amiloride-sensitive Na+/H+ exchange in a human leukemic cell line. Journal of Cell Biology, 99(1):340–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/1/

Bouhnik:1983:APR

[BCC^{+}83] J. Bouhnik, D. Cassio, E. Coezy, P. Corvol, and M. C. Weiss. Angiotensinogen production by rat hepatoma cells in culture and analysis of its regulation by techniques of somatic cell genetics. Journal of Cell Biology, 97(2):549–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/97/2/549.

Brown:1984:RMP

[BCF84] W. J. Brown, E. Constantinescu, and M. G. Farquhar. Redistribution of mannose-6-phosphate receptors induced by tunicamycin and chloroquine. Journal of Cell Biology, 99(1): 320–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/1/320.

Baker:1983:GJS

[BCHG83] T. S. Baker, D. L. Caspar, C. J. Hollingshead, and D. A. Goodenough. Gap junction structures. IV. Asymmetric features revealed by low-irradiation microscopy. Journal of Cell Biology, 96(1):204–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/1/204.

Barondes:1983:DDI

[BCHR83] S. H. Barondes, D. N. Cooper, and P. L. Haywood-Reid. Discoidin I and discoidin II are localized differently in developing Dictyostelium discoideum. Journal of Cell Biology, 96(1): 291–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/1/291.

Beisson:1980:CMF

[BCLT^{+}80] J. Beisson, J. Cohen, M. Lefort-Tran, M. Pouphile, and M. Rossignol. Control of membrane fusion in exocytosis. Physiological studies on a Paramecium mutant blocked in the final step of the trichocyst extrusion process. Journal of Cell Biology, 85(2):213–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/85/2/213.

Brinkley:1981:TAS

[BCP^{+}81] B. R. Brinkley, S. M. Cox, D. A. Pepper, L. Wible, S. L. Brenner, and R. L. Pardue. Tubulin assembly sites and the organization of cytoplasmic microtubules in cultured mammalian cells. Journal of Cell Biology, 90(3):554–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/90/3/554.

Bartelt:1982:CSN

[BCSC82] D. C. Bartelt, R. K. Carlin, G. A. Scheele, and W. D. Cohen. The cytoskeletal system of nucleated erythrocytes. II. presence of a high molecular weight calmodulin-binding protein. Journal of Cell Biology, 95(1):278–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/95/1/278.

Bok:1982:ILL

[BDH82] D. Bok, J. Dockstader, and J. Horwitz. Immunocytochemical localization of the lens main intrinsic polypeptide (MIP26) in communicating junctions. Journal of Cell Biology, 92(1): 213–??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/1/213.

Boucaut:1984:BAS

[BDP^{+}84] J. C. Boucaut, T. Darribère, T. J. Poole, H. Aoyama, K. M. Yamada, and J. P. Thiery. Biologically active synthetic peptides as probes of embryonic development: a competitive peptide inhibitor of fibronectin function inhibits gastrulation in amphibian embryos and neural crest cell migration in avian embryos. Journal of Cell Biology, 99(5):1822–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/5/
1822.

Berglindh:1980:ADH

[BDPS80] T. Berglindh, D. R. Dibona, C. S. Pace, and G. Sachs. ATP dependence of H+ secretion. Journal of Cell Biology, 85 (2):392–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/85/2/392.

Baumann:1982:DIS

[BE82] H. Baumann and D. Eldredge. Dexamethasone increases the synthesis and secretion of a partially active fibronectin in rat hepatoma cells. Journal of Cell Biology, 95(1):29–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/95/1/
29.

Beckerle:1984:MFP

[Bec84] M. C. Beckerle. Microinjected fluorescent polystyrene beads exhibit saltatory motion in tissue culture cells. Journal of Cell Biology, 98(6):2126–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/6/2126.

Berezney:1980:FNM

[Ber80a] R. Berezney. Fractionation of the nuclear matrix. I. Partial separation into matrix protein fibrils and a residual ribonucleoprotein fraction. Journal of Cell Biology, 85(3):641–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/85/3/
641.

Berry:1980:EMS

[Ber80b] R. W. Berry. Evidence for multiple somatic pools of individual axonally transported proteins. Journal of Cell Biology, 87(2): 379–??, November 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/87/2/379.

Bearer:1982:MAL

[BF82a] E. L. Bearer and D. S. Friend. Modifications of anioniclipid domains preceding membrane fusion in guinea pig sperm. Journal of Cell Biology, 92(3):604–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/92/3/604.

Byers:1982:SFC

[BF82b] H. R. Byers and K. Fujiwara. Stress fibers in cells in situ: immunofluorescence visualization with antiactin, antimyosin, and anti-alpha-actinin. Journal of Cell Biology, 93(3):804–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
93/3/804.

Byus:1982:DCL

[BF82c] C. V. Byus and W. H. Fletcher. Direct cytochemical localization of catalytic subunits dissociated from cAMP-dependent protein kinase in Reuber H-35 hepatoma cells. II. Temporal and spatial kinetics. Journal of Cell Biology, 93(3):727–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
93/3/727.

Bronner-Fraser:1984:LBP

[BF84a] M. Bronner-Fraser. Latex beads as probes of a neural crest pathway: effects of laminin, collagen, and surface charge on bead translocation. Journal of Cell Biology, 98(6):1947–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
98/6/1947.

Bursztajn:1984:ECV

[BF84b] S. Bursztajn and G. D. Fischbach. Evidence that coated vesicles transport acetylcholine receptors to the surface membrane of chick myotubes. Journal of Cell Biology, 98(2): 498–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/2/498.

Bravo:1983:PII

[BFL^{+}83] R. Bravo, S. J. Fey, P. M. Larsen, N. Coppard, and J. E. Celis. Proteins IEF (isoelectric focusing) 31 and IEF 46 are keratin-type components of the intermediate-sized filaments: keratins of various human cultured epithelial cells. Journal of Cell Biology, 96(2):416–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/2/416.

Brier:1983:ARG

[BFST83] J. Brier, M. Fechheimer, J. Swanson, and D. L. Taylor. Abundance, relative gelation activity, and distribution of the 95,000dalton actin-binding protein from Dictyostelium discoideum. Journal of Cell Biology, 97(1):178–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/1/178.

Bessen:1980:CCW

[BFW80] M. Bessen, R. B. Fay, and G. B. Witman. Calcium control of waveform in isolated flagellar axonemes of chlamydomonas. Journal of Cell Biology, 86(2):446–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/86/2/446.

Black:1982:CCS

[BG82] M. M. Black and L. A. Greene. Changes in the colchicine susceptibility of microtubules associated with neurite outgrowth: studies with nerve growth factor-responsive PC12 pheochromocytoma cells. Journal of Cell Biology, 95(2):379–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
95/2/379.

Bultmann:1983:ADN

[BG83] B. D. Bültmann and H. Gruler. Analysis of the directed and nondirected movement of human granulocytes: influence of temperature and ECHO 9 virus on N-formylmethionylleucylphenylalanine-induced chemokinesis and chemotaxis. Journal of Cell Biology, 96(6):1708–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/6/1708.

Baumann:1980:DRP

[BGD80] H. Baumann, T. D. Gelehrter, and D. Doyle. Dexamethasone regulates the program of secretory glycoprotein synthesis in hepatoma tissue culture cells. Journal of Cell Biology, 85(1):1–??, April 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/85/1/1.

Brackenbury:1984:PCL

[BGE84] R. Brackenbury, M. E. Greenberg, and G. M. Edelman. Phenotypic changes and loss of N-CAM–mediated adhesion in transformed embryonic chicken retinal cells. Journal of Cell Biology, 99(6):1944–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/6/1944.

Blaker:1981:ATM

[BGM81] W. D. Blaker, J. F. Goodrum, and P. Morell. Axonal transport of the mitochondria-specific lipid, diphosphatidylglycerol, in the rat visual system. Journal of Cell Biology, 89(3):579–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
89/3/579.

Bernd:1981:SDT

[BGNT81] P. Bernd, M. D. Gershon, E. A. Nunez, and H. Tamir. Separation of dissociated thyroid follicular and parfollicular cells: association of serotonin binding protein with parafollicular cells. Journal of Cell Biology, 88(3):499–??, March 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/88/3/499.

Bouchard:1980:LPC

[BGPB80] P. Bouchard, C. Gagnon, D. M. Phillips, and C. W. Bardin. The localization of protein carboxyl-methylase in sperm tails. Journal of Cell Biology, 86(2):417–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/86/2/417.

Byrne:1981:DCB

[BGS81] P. V. Byrne, L. J. Guilbert, and E. R. Stanley. Distribution of cells bearing receptors for a colony-stimulating factor (CSF-1) in murine tissues. Journal of Cell Biology, 91(3):848–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
91/3/848.

Burland:1983:CTD

[BGS^{+}83] T. G. Burland, K. Gull, T. Schedl, R. S. Boston, and W. F. Dove. Cell type-dependent expression of tubulins in Physarum. Journal of Cell Biology, 97(6):1852–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/6/
1852.

Bloch:1983:CCV

[BH83] R. J. Bloch and Z. W. Hall. Cytoskeletal components of the vertebrate neuromuscular junction: vinculin, alpha-actinin, and filamin. Journal of Cell Biology, 97(1):217–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/97/1/217.

Beertsen:1982:FPT

[BHA82] W. Beertsen, J. N. M. Heersche, and J. E. Aubin. Free and polymerized tubulin in cultured bone cells and Chinese hamster ovary cells: the influence of cold and hormones. Journal of Cell Biology, 95(2):387–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/95/2/387.

Baumann:1983:PDT

[BHJ83] H. Baumann, E. Hou, and G. P. Jahreis. Preferential degradation of the terminal carbohydrate moiety of membrane glycoproteins in rat hepatoma cells and after transfer to the membranes of mouse fibroblasts. Journal of Cell Biology, 96(1): 139–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/1/139.

Blum:1980:CCC

[BHJV80] J. J. Blum, A. Hayes, G. A. Jamieson, and T. C. Vanaman. Calmodulin confers calcium sensitivity on ciliary dynein AT-Pase. Journal of Cell Biology, 87(2):386–??, November 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/87/2/386.

Brodsky:1983:TLP

[BHP83] F. M. Brodsky, N. J. Holmes, and P. Parham. Tropomyosinlike properties of clathrin light chains allow a rapid, high-yield purification. Journal of Cell Biology, 96(3):911–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/3/
911.

Barondes:1981:EEC

[BHR81] S. H. Barondes and P. L. Haywood-Reid. Externalization of an endogenous chicken muscle lectin with in vivo development. Journal of Cell Biology, 91(2):568–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/91/2/568.

Bareis:1982:PMC

[BHSA82] D. L. Bareis, F. Hirata, E. Schiffmann, and J. Axelrod. Phospholipid metabolism, calcium flux, and the receptor-mediated induction of chemotaxis in rabbit neutrophils. Journal of Cell Biology, 93(3):690–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/93/3/690.

Borregaard:1983:SLB

[BHSC83] N. Borregaard, J. M. Heiple, E. R. Simons, and R. A. Clark. Subcellular localization of the b-cytochrome component of the human neutrophil microbicidal oxidase: translocation during activation. Journal of Cell Biology, 97(1):52–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/97/1/52.

Baumann:1983:RMH

[BJ83] H. Baumann and G. P. Jahreis. Regulation of mouse haptoglobin synthesis. Journal of Cell Biology, 97(3):728– ??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/3/728.

Baumann:1983:SRA

[BJG83] H. Baumann, G. P. Jahreis, and K. C. Gaines. Synthesis and regulation of acute phase plasma proteins in primary cultures of mouse hepatocytes. Journal of Cell Biology, 97(3): 866–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/3/866.

Baker:1981:IIP

[BJS81] R. E. Baker, L. S. Jefferson, and R. Shiman. Immunocytochemical identification of phenylalanine hydroxylase and albumin in cultured hepatoma cells and isolated rat hepatocytes. Journal of Cell Biology, 90(1):145–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/90/1/145.

Baer:1983:PRR

[BK83a] B. W. Baer and R. D. Kornberg. The protein responsible for the repeating structure of cytoplasmic poly(A)- ribonucleoprotein. Journal of Cell Biology, 96(3):717–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
96/3/717.

Black:1983:MAP

[BK83b] M. M. Black and J. T. Kurdyla. Microtubule-associated proteins of neurons. Journal of Cell Biology, 97(4):1020–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
97/4/1020.

Burgess:1984:SSA

[BK84] T. L. Burgess and R. B. Kelly. Sorting and secretion of adrenocorticotropin in a pituitary tumor cell line after perturbation of the level of a secretory granule-specific proteoglycan. Journal of Cell Biology, 99(6):2223–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/6/2223.

Bergen:1980:PMN

[BKB80] L. G. Bergen, R. Kuriyama, and G. G. Borisy. Polarity of microtubules nucleated by centrosomes and chromosomes of Chinese hamster ovary cells in vitro. Journal of Cell Biology, 84 (1):151–??, January 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/84/1/151.

Bergeron:1982:ITT

[BKL^{+}82] J. J. Bergeron, G. J. Kotwal, G. Levine, P. Bilan, R. Rachubinski, M. Hamilton, G. C. Shore, and H. P. Ghosh. Intracellular transport of the transmembrane glycoprotein G of vesicular stomatitis virus through the Golgi apparatus as visualized by electron microscope radioautography. Journal of Cell Biology, 94(1):36–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/94/1/36.

Burridge:1982:NSA

[BKM82] K. Burridge, T. Kelly, and P. Mangeat. Nonerythrocyte spectrins: actin-membrane attachment proteins occurring in many cell types. Journal of Cell Biology, 95(2):478–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/95/2/478.

Bennett:1981:SISb

[BKO81] G. Bennett, F. W. Kan, and D. O’Shaughnessy. The site of incorporation of sialic acid residues into glycoproteins and the subsequent fates of these molecules in various rat and mouse cell types as shown by radioautography after injection of [3H]N-acetylmannosamine. II. Observations in tissues other than liver. Journal of Cell Biology, 88(1):16–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/88/1/16.

Barth:1984:RNG

[BKT84] E. M. Barth, S. Korsching, and H. Thoenen. Regulation of nerve growth factor synthesis and release in organ cultures of rat iris. Journal of Cell Biology, 99(3):839–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/3/
839.

Bennett:1980:LTI

[BL80a] W. F. Bennett and G. Lynch. Low-temperature induction of calcium-dependent protein phosphorylation in blood platelets. Journal of Cell Biology, 86(1):280–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/86/1/280.

Black:1980:SCA

[BL80b] M. M. Black and R. J. Lasek. Slow components of axonal transport: two cytoskeletal networks. Journal of Cell Biology, 86 (2):616–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/86/2/616.

Breckler:1982:INH

[BL82a] J. Breckler and E. Lazarides. Isolation of a new high molecular weight protein associated with desmin and vimentin filaments from avian embryonic skeletal muscle. Journal of Cell Biology, 92(3):795–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/3/795.

Burch:1982:VSA

[BL82b] W. M. Burch and H. E. Lebovitz. In vitro stimulation of alkaline phosphatase activity in immature embryonic chick pelvic cartilage by adenosine. Journal of Cell Biology, 93(2):338–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
93/2/338.

Blikstad:1983:VFA

[BL83a] I. Blikstad and E. Lazarides. Vimentin filaments are assembled from a soluble precursor in avian erythroid cells. Journal of Cell Biology, 96(6):1803–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/6/1803.

Bloodgood:1983:TIC

[BL83b] R. A. Bloodgood and E. N. Levin. Transient increase in calcium efflux accompanies fertilization in Chlamydomonas. Journal of Cell Biology, 97(2):397–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/2/397.

Barazzone:1980:BIL

[BLG^{+}80] P. Barazzone, M. A. Lesniak, P. Gorden, E. Van Obberghen, J. L. Carpentier, and L. Orci. Binding, internalization, and lysosomal association of 125I-human growth hormone in cultured human lymphocytes: a quantitative morphological and biochemical study. Journal of Cell Biology, 87(2):360– ??, November 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/87/2/360.

Brokaw:1982:AMC

[BLH82] C. J. Brokaw, D. J. Luck, and B. Huang. Analysis of the movement of Chlamydomonas flagella: the function of the radialspoke system is revealed by comparison of wild-type and mutant flagella. Journal of Cell Biology, 92(3):722–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/92/3/
722.

Bridges:1981:CRC

[BLHC81] K. Bridges, R. Levenson, D. Housman, and L. Cantley. Calcium regulates the commitment of murine erythroleukemia cells to terminal erythroid differentiation. Journal of Cell Biology, 90(2):542–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/90/2/542.

Bloch:1984:IAR

[Blo84] R. J. Bloch. Isolation of acetylcholine receptor clusters in substrate-associated material from cultured rat myotubes using saponin. Journal of Cell Biology, 99(3):984–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/3/
984.

Browne:1980:ILC

[BLS^{+}80] C. L. Browne, A. H. Lockwood, J. L. Su, J. A. Beavo, and A. L. Steiner. Immunofluorescent localization of cyclic nucleotidedependent protein kinases on the mitotic apparatus of cultured cells. Journal of Cell Biology, 87(2):336–??, November 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/87/2/336.

Bloom:1984:WCD

[BLV84] G. S. Bloom, F. C. Luca, and R. B. Vallee. Widespread cellular distribution of MAP-1A (microtubule-associated protein 1A) in the mitotic spindle and on interphase microtubules. Journal of Cell Biology, 98(1):331–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/1/331.

Borgese:1980:LBNa

[BM80] N. Borgese and J. Meldolesi. Localization and biosynthesis of NADH-cytochrome b5 reductase, an integral membrane protein, in rat liver cells. I. Distribution of the enzyme activity in microsomes, mitochondria, and Golgi complex. Journal of Cell Biology, 85(3):501–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/85/3/501.

Barry:1982:RRM

[BM82a] B. Barry and R. Mathies. Resonance Raman microscopy of rod and cone photoreceptors. Journal of Cell Biology, 94(2): 479–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/94/2/479.

Bernhardt:1982:IPD

[BM82b] R. Bernhardt and A. Matus. Initial phase of dendrite growth: evidence for the involvement of high molecular weight microtubule-associated proteins (HMWP) before the appearance of tubulin. Journal of Cell Biology, 92(2):589–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
92/2/589.

Bloodgood:1982:FMC

[BM82c] R. A. Bloodgood and G. S. May. Functional modification of the Chlamydomonas flagellar surface. Journal of Cell Biology, 93(1):88–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/93/1/88.

Bonder:1983:DEM

[BM83] E. M. Bonder and M. S. Mooseker. Direct electron microscopic visualization of barbed end capping and filament cutting by intestinal microvillar 95-kdalton protein (villin): a new actin assembly assay using the limulus acrosomal process. Journal of Cell Biology, 96(4):1097–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/4/1097.

Bursztajn:1984:OAR

[BMA84] S. Bursztajn, J. L. McManaman, and S. H. Appel. Organization of acetylcholine receptor clusters in cultured rat myotubes is calcium dependent. Journal of Cell Biology, 98(2): 507–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/2/507.

Busson-Mabillot:1982:MPS

[BMCGO^{+}82] S. Busson-Mabillot, A. M. Chambaut-Guérin, L. Ovtracht, P. Muller, and B. Rossignol. Microtubules and protein secretion in rat lacrimal glands: localization of short-term effects of colchicine on the secretory process. Journal of Cell Biology, 95 (1):105–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/95/1/105.

Bader:1982:IAM

[BMF82] D. Bader, T. Masaki, and D. A. Fischman. Immunochemical analysis of myosin heavy chain during avian myogenesis in vivo and in vitro. Journal of Cell Biology, 95(3):763–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/95/3/
763.

Blose:1984:NFI

[BMF84] S. H. Blose, D. I. Meltzer, and J. R. Feramisco. 10-nm filaments are induced to collapse in living cells microinjected with monoclonal and polyclonal antibodies against tubulin. Journal of Cell Biology, 98(3):847–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/3/847.

Boonstra:1983:IRG

[BMH^{+}83] J. Boonstra, W. H. Moolenaar, P. H. Harrison, P. Moed, P. T. van der Saag, and S. W. de Laat. Ionic responses and growth stimulation induced by nerve growth factor and epidermal growth factor in rat pheochromocytoma (PC12) cells. Journal of Cell Biology, 97(1):92–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/97/1/92.

Bunt-Milam:1983:ILT

[BMS83] A. H. Bunt-Milam and J. C. Saari. Immunocytochemical localization of two retinoid-binding proteins in vertebrate retina. Journal of Cell Biology, 97(3):703–??, September 1983. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/97/3/703.

Bridgman:1983:DFS

[BN83] P. C. Bridgman and Y. Nakajima. Distribution of filipin-sterol complexes on cultured muscle cells: cell-substratum contact areas associated with acetylcholine receptor clusters. Journal of Cell Biology, 96(2):363–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/2/363.

Bourguignon:1982:AML

[BNB^{+}82] L. Y. Bourguignon, M. L. Nagpal, K. Balazovich, V. Guerriero, and A. R. Means. Association of myosin light chain kinase with lymphocyte membrane-cytoskeleton complex. Journal of Cell Biology, 95(3):793–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/95/3/793.

Bridgman:1984:FFE

[BNGN84] P. C. Bridgman, S. Nakajima, A. S. Greenberg, and Y. Nakajima. Freeze-fracture and electrophysiological studies of newly developed acetylcholine receptors in Xenopus embryonic muscle cells. Journal of Cell Biology, 98(6):2160–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/6/
2160.

Bourguignon:1981:PML

[BNH81] L. Y. Bourguignon, M. L. Nagpal, and Y. C. Hsing. Phosphorylation of myosin light chain during capping of mouse Tlymphoma cells. Journal of Cell Biology, 91(3):889–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
91/3/889.

Brown:1982:PPA

[BNM82] B. A. Brown, R. A. Nixon, and C. A. Marotta. Posttranslational processing of alpha-tubulin during axoplasmic transport in CNS axons. Journal of Cell Biology, 94(1):159–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/94/1/
159.

Berlin:1980:SFD

[BO80] R. D. Berlin and J. M. Oliver. Surface functions during mitosis. II. Quantitation of pinocytosis and kinetic characterization of the mitotic cycle with a new fluorescence technique. Journal of Cell Biology, 85(3):660–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/85/3/660.

Bennett:1981:SISa

[BO81a] G. Bennett and D. O’Shaughnessy. The site of incorporation of sialic acid residues into glycoproteins and the subsequent fates of these molecules in various rat and mouse cell types as shown by radioautography after injection of [3H]Nacetylmannosamine. I. Observations in hepatocytes. Journal of Cell Biology, 88(1):1–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/88/1/1.

Broadwell:1981:GAG

[BO81b] R. D. Broadwell and C. Oliver. Golgi apparatus, GERL, and secretory granule formation within neurons of the hypothalamo-neurohypophysial system of control and hyperosmotically stressed mice. Journal of Cell Biology, 90(2):474–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
90/2/474.

Bowers:1983:ADI

[BO83] B. Bowers and T. E. Olszewski. Acanthamoeba discriminates internally between digestible and indigestible particles. Journal of Cell Biology, 97(2):317–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/2/317.

Bogorad:1981:C

[Bog81] L. Bogorad. Chloroplasts. Journal of Cell Biology, 91(3): ??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/91/3/256s.

Bowers:1981:MAV

[BOH81] B. Bowers, T. E. Olszewski, and J. Hyde. Morphometric analysis of volumes and surface areas in membrane compartments during endocytosis in Acanthamoeba. Journal of Cell Biology, 88(3):509–??, March 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/88/3/509.

Bowers:1980:MSP

[Bow80] B. Bowers. A morphological study of plasma and phagosome membranes during endocytosis in Acanthamoeba. Journal of Cell Biology, 84(2):246–??, February 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/84/2/246.

Besharse:1980:MAR

[BP80] J. C. Besharse and K. H. Pfenninger. Membrane assembly in retinal photoreceptors I. Freeze-fracture analysis of cytoplasmic vesicles in relationship to disc assembly. Journal of Cell Biology, 87(2):451–??, November 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/87/2/451.

Burgess:1982:RBB

[BP82] D. R. Burgess and B. E. Prum. Reevaluation of brush border motility: calcium induces core filament solation and microvillar vesiculation. Journal of Cell Biology, 94(1):97–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/94/1/
97.

Beckerle:1983:ARM

[BP83] M. C. Beckerle and K. R. Porter. Analysis of the role of microtubules and actin in erythrophore intracellular motility. Journal of Cell Biology, 96(2):354–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/2/354.

Brenner:1981:KSD

[BPB^{+}81] S. Brenner, D. Pepper, M. W. Berns, E. Tan, and B. R. Brinkley. Kinetochore structure, duplication, and distribution in mammalian cells: analysis by human autoantibodies from scleroderma patients. Journal of Cell Biology, 91(1):95–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
91/1/95.

Bowen-Pope:1983:IBR

[BPDR83] D. F. Bowen-Pope, P. E. Dicorleto, and R. Ross. Interactions between the receptors for platelet-derived growth factor and epidermal growth factor. Journal of Cell Biology, 96(3):679–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
96/3/679.

Bolender:1980:ISB

[BPM^{+}80a] R. P. Bolender, D. Paumgartner, D. Muellener, G. Losa, and E. R. Weibel. Integrated stereological and biochemical studies on hepatocytic membranes. I.V. Heterogeneous distribution of marker enzymes on endoplasmic reticulum membranes in fractions. Journal of Cell Biology, 85(3):577–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/85/3/577.

Borgese:1980:LBNb

[BPM80b] N. Borgese, G. Pietrini, and J. Meldolesi. Localization and Biosynthesis of NADH–Cytochrome b(5) reductase, an integral membrane protein, in rat liver cells. III. Evidence for the independent insertion and turnover of the enzyme in various subcellular compartments. Journal of Cell Biology, 86 (1):38–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/86/1/38.

Bilinski:1981:SPD

[BPM81] M. Bilinski, H. Plattner, and H. Matt. Secretory protein decondensation as a distinct, Ca^{2+}-mediated event during the final steps of exocytosis in Paramecium cells. Journal of Cell Biology, 88(1):179–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/88/1/179.

Banerjee:1983:BHD

[BR83] D. Banerjee and C. M. Redman. Biosynthesis of high density lipoprotein by chicken liver: nature of nascent intracellular high density lipoprotein. Journal of Cell Biology, 96(3):651–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
96/3/651.

Banerjee:1984:BHD

[BR84a] D. Banerjee and C. M. Redman. Biosynthesis of high density lipoprotein by chicken liver: conjugation of nascent lipids with apoprotein A1. Journal of Cell Biology, 99(6):1917–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/6/1917.

Bridgman:1984:SCD

[BR84b] P. C. Bridgman and T. S. Reese. The structure of cytoplasm in directly frozen cultured cells. I. Filamentous meshworks and the cytoplasmic ground substance. Journal of Cell Biology, 99(5):1655–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/5/1655.

Burns:1984:EID

[BR84c] C. P. Burns and E. Rozengurt. Extracellular Na+ and initiation of DNA synthesis: role of intracellular pH and K+. Journal of Cell Biology, 98(3):1082–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/3/1082.

Bitterman:1983:RFG

[BRAC83] P. B. Bitterman, S. I. Rennard, S. Adelberg, and R. G. Crystal. Role of fibronectin as a growth factor for fibroblasts. Journal of Cell Biology, 97(6):1925–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/6/1925.

Bretscher:1983:PDP

[Bre83] A. Bretscher. Purification of an 80,000-dalton protein that is a component of the isolated microvillus cytoskeleton, and its localization in nonmuscle cells. Journal of Cell Biology, 97 (2):425–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/2/425.

Begg:1982:SOA

[BRH82] DA Begg, LI Rebhun, and H. Hyatt. Structural organization of actin in the sea urchin egg cortex: microvillar elongation in the absence of actin filament bundle formation. Journal of Cell Biology, 93(1):24–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/93/1/24.

Brodie:1981:BRE

[Bro81] D. A. Brodie. Bead rings at the endoplasmic reticulumgolgi complex boundary: morphological changes accompanying inihibition of intracellular transport of secretory proteins in arthropod fat body tissue. Journal of Cell Biology, 90 (1):92–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/1/92.

Bergeron:1983:ECI

[BRR^{+}83] J. J. Bergeron, L. Resch, R. Rachubinski, B. A. Patel, and B. I. Posner. Effect of colchicine on internalization of prolactin in female rat liver: an in vivo radioautographic study. Journal of Cell Biology, 96(3):875–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/3/875.

Buckley:1981:CIF

[BRS81] I. K. Buckley, T. R. Raju, and M. Stewart. Claims that intermediate filaments contain F-actin are unwarranted. Journal of Cell Biology, 90(2):309–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/90/2/309.

Bergeron:1982:GTE

[BRS^{+}82] J. J. Bergeron, R. A. Rachubinski, R. A. Sikstrom, B. I. Posner, and J. Paiement. Galactose transfer to endogenous acceptors within Golgi fractions of rat liver. Journal of Cell Biology, 92(1):139–??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/92/1/139.

Breen:1980:CIO

[BS80] G. A. Breen and I. E. Scheffler. Cytoplasmic inheritance of oligomycin resistance in Chinese hamster ovary cells. Journal of Cell Biology, 86(3):723–??, September 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/86/3/723.

Brown:1981:MAC

[BS81a] S. S. Brown and J. A. Spudich. Mechanism of action of cytochalasin: evidence that it binds to actin filament ends. Journal of Cell Biology, 88(3):487–??, March 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/88/3/487.

Browning:1981:VPU

[BS81b] E. T. Browning and M. M. Sanders. Vimentin: a phosphoprotein under hormonal regulation. Journal of Cell Biology, 90(3): 803–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/3/803.

Burnett:1981:IMM

[BS81c] K. G. Burnett and I. E. Scheffler. Integrity of mitochondria in a mammalian cell mutant defective in mitochondrial protein synthesis. Journal of Cell Biology, 90(1):108–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/90/1/108.

Bandman:1982:IKI

[BS82a] E. Bandman and R. C. Strohman. Increased K+ inhibits spontaneous contractions reduces myosin accumulation in cultured chick myotubes. Journal of Cell Biology, 93(3):698–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/93/3/
698.

Banks-Schlegel:1982:KAD

[BS82b] S. P. Banks-Schlegel. Keratin alterations during embryonic epidermal differentiation: a presage of adult epidermal maturation. Journal of Cell Biology, 93(3):551–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/93/3/551.

Bond:1982:DBA

[BS82c] M. Bond and A. V. Somlyo. Dense bodies and actin polarity in vertebrate smooth muscle. Journal of Cell Biology, 95(2): 403–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/95/2/403.

Bunt:1982:FPR

[BS82d] A. H. Bunt and J. C. Saari. Fucosylated protein of retinal cone photoreceptor outer segments: morphological and biochemical analyses. Journal of Cell Biology, 92(2):269–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/92/2/269.

Bergmann:1983:IMS

[BS83] J. E. Bergmann and S. J. Singer. Immunoelectron microscopic studies of the intracellular transport of the membrane glycoprotein (G) of vesicular stomatitis virus in infected Chinese hamster ovary cells. Journal of Cell Biology, 97(6): 1777–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/6/1777.

Banks-Schlegel:1981:IST

[BSG81] S. Banks-Schlegel and H. Green. Involucrin synthesis and tissue assembly by keratinocytes in natural and cultured human epithelia. Journal of Cell Biology, 90(3):732–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/90/3/
732.

Banks-Schlegel:1983:DHE

[BSH83a] S. P. Banks-Schlegel and P. M. Howley. Differentiation of human epidermal cells transformed by SV40. Journal of Cell Biology, 96(2):330–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/2/330.

Begg:1983:CSO

[BSH83b] D. A. Begg, E. D. Salmon, and H. A. Hyatt. The changes in structural organization of actin in the sea urchin egg cortex in response to hydrostatic pressure. Journal of Cell Biology, 97(6):1795–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/6/1795.

Bar-Shavit:1983:CRM

[BSKFW83] R. Bar-Shavit, A. Kahn, J. W. Fenton, and G. D. Wilner. Chemotactic response of monocytes to thrombin. Journal of Cell Biology, 96(1):282–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/1/282.

Burnside:1982:RCD

[BSNP82] B. Burnside, B. Smith, M. Nagata, and K. Porrello. Reactivation of contraction in detergent-lysed teleost retinal cones. Journal of Cell Biology, 92(1):199–??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/92/1/199.

Bar-Sagi:1983:TCA

[BSP83] D. Bar-Sagi and J. Prives. Trifluoperazine, a calmodulin antagonist, inhibits muscle cell fusion. Journal of Cell Biology, 97(5):1375–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/5/1375.

Baker:1984:RCT

[BSR84] E. J. Baker, J. A. Schloss, and J. L. Rosenbaum. Rapid changes in tubulin RNA synthesis and stability induced by deflagellation in Chlamydomonas. Journal of Cell Biology, 99(6): 2074–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/6/2074.

Berner:1981:HII

[BSS81] P. F. Berner, A. V. Somlyo, and A. P. Somlyo. Hypertrophyinduced increase of intermediate filaments in vascular smooth muscle. Journal of Cell Biology, 88(1):96–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/88/1/96.

Brown:1983:SAGa

[BSS83a] W. J. Brown, W. A. Shannon, and W. J. Snell. Specific and azurophilic granules from rabbit polymorphonuclear leukocytes. I. Isolation and characterization of membrane and content subfractions. Journal of Cell Biology, 96(4):1030–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/4/
1030.

Brown:1983:SAGb

[BSS83b] W. J. Brown, W. A. Shannon, and W. J. Snell. Specific and azurophilic granules from rabbit polymorphonuclear leukocytes. II. Cell surface localization of granule membrane and content proteins before and after degranulation. Journal of Cell Biology, 96(4):1040–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/4/1040.

Braell:1984:DCC

[BSSR84] W. A. Braell, D. M. Schlossman, S. L. Schmid, and J. E. Rothman. Dissociation of clathrin coats coupled to the hydrolysis of ATP: role of an uncoating ATPase. Journal of Cell Biology, 99 (2):734–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/2/734.

Bloom:1984:WDM

[BSV84] G. S. Bloom, T. A. Schoenfeld, and R. B. Vallee. Widespread distribution of the major polypeptide component of MAP 1 (microtubule-associated protein 1) in the nervous system. Journal of Cell Biology, 98(1):320–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/1/320.

Bormioli:1980:SMV

[BSVS80] S. P. Bormioli, S. Sartore, M. Vitadello, and S. Schiaffino. ”slow” myosins in vertebrae skeletal muscle. An immunofluorescence study. Journal of Cell Biology, 85(3):672–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/85/3/
672.

Birk:1984:ECM

[BT84] D. E. Birk and R. L. Trelstad. Extracellular compartments in matrix morphogenesis: collagen fibril, bundle, and lamellar formation by corneal fibroblasts. Journal of Cell Biology, 99(6):2024–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/6/2024.

Bernabeu:1983:NPC

[BTF^{+}83] C. Bernabeu, E. M. Tobin, A. Fowler, I. Zabin, and J. A. Lake. Nascent polypeptide chains exit the ribosome in the same relative position in both eucaryotes and procaryotes. Journal of Cell Biology, 96(5):1471–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/5/1471.

Brady:1981:ATC

[BTHL81] S. T. Brady, M. Tytell, K. Heriot, and R. J. Lasek. Axonal transport of calmodulin: a physiologic approach to identification of long-term associations between proteins. Journal of Cell Biology, 89(3):607–??, June 1981. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/89/3/607.

Bhakdi:1982:TMC

[BTJ82] S. Bhakdi and J. Tranum-Jensen. Terminal membrane C5b-9 complex of human complement: transition from an amphiphilic to a hydrophilic state through binding of the S protein from serum. Journal of Cell Biology, 94(3):755– ??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/94/3/755.

Brady:1984:ATA

[BTL84] S. T. Brady, M. Tytell, and R. J. Lasek. Axonal tubulin and axonal microtubules: biochemical evidence for cold stability. Journal of Cell Biology, 99(5):1716–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/5/
1716.

Brisson:1984:TCA

[BU84] A. Brisson and P. N. Unwin. Tubular crystals of acetylcholine receptor. Journal of Cell Biology, 99(4):1202–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/4/
1202.

Burden:1982:IIP

[Bur82a] S. Burden. Identification of an intracellular postsynaptic antigen at the frog neuromuscular junction. Journal of Cell Biology, 94(3):521–??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/94/3/521.

Burgess:1982:RIE

[Bur82b] D. R. Burgess. Reactivation of intestinal epithelial cell brush border motility: ATP-dependent contraction via a terminal web contractile ring. Journal of Cell Biology, 95(3):853–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
95/3/853.

Burton:1984:LMM

[Bur84] P. R. Burton. Luminal material in microtubules of frog olfactory axons: structure and distribution. Journal of Cell Biology, 99(2):520–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/2/520.

Bloom:1983:AMA

[BV83] G. S. Bloom and R. B. Vallee. Association of microtubuleassociated protein 2 (MAP 2) with microtubules and intermediate filaments in cultured brain cells. Journal of Cell Biology, 96(6):1523–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/6/1523.

Bernier-Valentin:1983:ETB

[BVAR83] F. Bernier-Valentin, D. Aunis, and B. Rousset. Evidence for tubulin-binding sites on cellular membranes: plasma membranes, mitochondrial membranes, and secretory granule membranes. Journal of Cell Biology, 97(1):209–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/97/1/209.

Bollini:1983:VVP

[BVC83] R. Bollini, A. Vitale, and M. J. Chrispeels. In vivo and in vitro processing of seed reserve protein in the endoplasmic reticulum: evidence for two glycosylation steps. Journal of Cell Biology, 96(4):999–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/4/999.

Burger:1984:OFM

[BvdMN84] E. H. Burger, J. W. van der Meer, and P. J. Nijweide. Osteoclast formation from mononuclear phagocytes: role of boneforming cells. Journal of Cell Biology, 99(6):1901–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/6/1901.

Brown:1983:CPM

[BvGJ83] C. R. Brown, K. I. von Glos, and R. Jones. Changes in plasma membrane glycoproteins of rat spermatozoa during maturation in the epididymis. Journal of Cell Biology, 96(1):256–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
96/1/256.

Brody:1982:AAB

[BVGS82] J. S. Brody, C. A. Vaccaro, P. J. Gill, and J. E. Silbert. Alterations in alveolar basement membranes during postnatal lung growth. Journal of Cell Biology, 95(2):394–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/95/2/394.

Bretscher:1980:FNM

[BW80] A. Bretscher and K. Weber. Fimbrin, a new microfilamentassociated protein present in microvilli and other cell surface structures. Journal of Cell Biology, 86(1):335–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/86/1/335.

Barak:1981:FLD

[BW81] L. S. Barak and W. W. Webb. Fluorescent low density lipoprotein for observation of dynamics of individual receptor complexes on cultured human fibroblasts. Journal of Cell Biology, 90(3):595–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/90/3/595.

Barak:1982:DLD

[BW82] L. S. Barak and W. W. Webb. Diffusion of low density lipoprotein-receptor complex on human fibroblasts. Journal of Cell Biology, 95(3):846–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/95/3/846.

Brennan:1980:SPS

[BWM80] M. D. Brennan, T. G. Warren, and A. P. Mahowald. Signal peptides and signal peptidase in Drosophila melanogaster. Journal of Cell Biology, 87(2):516–??, November 1980. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/87/2/516.

Bumann:1984:AIC

[BWM84] J. Bumann, B. Wurster, and D. Malchow. Attractant-induced changes and oscillations of the extracellular Ca++ concentration in suspensions of differentiating Dictyostelium cells. Journal of Cell Biology, 98(1):173–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/1/173.

Burgess:1984:ICB

[BWV84] W. H. Burgess, D. M. Watterson, and L. J. Van Eldik. Identification of calmodulin-binding proteins in chicken embryo fibroblasts. Journal of Cell Biology, 99(2):550–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/99/2/550.

Bunge:1980:CNC

[BWW^{+}80] M. B. Bunge, A. K. Williams, P. M. Wood, J. Uitto, and J. J. Jeffrey. Comparison of nerve cell and nerve cell plus Schwann cell cultures, with particular emphasis on basal lamina and collagen formation. Journal of Cell Biology, 84(1): 184–??, January 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/84/1/184.

Brown:1982:DPD

[BYS82] S. S. Brown, K. Yamamoto, and J. A. Spudich. A 40,000dalton protein from Dictyostelium discoideum affects assembly properties of actin in a Ca^{2+}-dependent manner. Journal of Cell Biology, 93(1):205–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/93/1/205.

Barak:1981:VSC

[BYW81] L. S. Barak, R. R. Yocum, and W. W. Webb. In vivo staining of cytoskeletal actin by autointernalization of nontoxic concentrations of nitrobenzoxadiazole-phallacidin. Journal of Cell Biology, 89(2):368–??, May 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/89/2/368.

Ben-Zeev:1983:CCR

[BZ83] A. Ben-Ze’ev. Cell configuration-related control of vimentin biosynthesis and phosphorylation in cultured mammalian cells.

Journal of Cell Biology, 97(3):858–??, September 1983. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/97/3/858.

Ben-Zeev:1984:DCC

[BZ84] A. Ben-Ze’ev. Differential control of cytokeratins and vimentin synthesis by cell–cell contact and cell spreading in cultured epithelial cells. Journal of Cell Biology, 99(4):1424–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/4/
1424.

Burmer:1982:EEP

[BZN82] G. C. Burmer, C. J. Zeigler, and T. H. Norwood. Evidence for endogenous polypeptide-mediated inhibition of cell-cycle transit in human diploid cells. Journal of Cell Biology, 94 (1):187–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/94/1/187.

Cabral:1983:ICH

[Cab83] F. R. Cabral. Isolation of Chinese hamster ovary cell mutants requiring the continuous presence of taxol for cell division. Journal of Cell Biology, 97(1):22–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/1/22.

Chung:1982:CPR

[CAL^{+}82] S. D. Chung, N. Alavi, D. Livingston, S. Hiller, and M. Taub. Characterization of primary rabbit kidney cultures that express proximal tubule functions in a hormonally defined medium. Journal of Cell Biology, 95(1):118–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/95/1/118.

Cande:1980:PCM

[Can80] W. Z. Cande. A permeabilized cell model for studying cytokinesis using mammalian tissue culture cells. Journal of Cell Biology, 87(2):326–??, November 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/87/2/326.

Cancalon:1982:SFA

[Can82] P. Cancalon. Slow flow in axons detached from their perikarya. Journal of Cell Biology, 95(3):989–??, December 1982. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/95/3/989.

Cancalon:1983:PDC

[Can83] P. Cancalon. Proximodistal degeneration of C-fibers detached from their perikarya. Journal of Cell Biology, 97(1):6–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/1/
6.

Carter:1984:RID

[Car84] W. G. Carter. The role of intermolecular disulfide bonding in deposition of GP140 in the extracellular matrix. Journal of Cell Biology, 99(1):105–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/1/105.

Castellot:1981:CEC

[CARK81] J. J. Castellot, M. L. Addonizio, R. Rosenberg, and M. J. Karnovsky. Cultured endothelial cells produce heparinlike inhibitor of smooth muscle cell growth. Journal of Cell Biology, 90(2):372–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/90/2/372.

Cohen:1984:PAF

[CAZF84] F. S. Cohen, M. H. Akabas, J. Zimmerberg, and A. Finkelstein. Parameters affecting the fusion of unilamellar phospholipid vesicles with planar bilayer membranes. Journal of Cell Biology, 98(3):1054–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/3/1054.

Carey:1981:FIR

[CB81] D. J. Carey and R. P. Bunge. Factors influencing the release of proteins by cultured Schwann cells. Journal of Cell Biology, 91(3):666–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/91/3/666.

Carney:1982:TBC

[CB82] D. H. Carney and J. S. Bergmann. 125i-thrombin binds to clustered receptors on noncoated regions of mouse embryo cell surfaces. Journal of Cell Biology, 95(3):697–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/95/3/697.

Caswell:1984:IEP

[CB84a] A. H. Caswell and J. P. Brunschwig. Identification and extraction of proteins that compose the triad junction of skeletal muscle. Journal of Cell Biology, 99(3):929–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/3/
929.

Chen:1984:EGG

[CB84b] S. J. Chen and G. B. Bouck. Endogenous glycosyltransferases glucosylate lipids in flagella of Euglena. Journal of Cell Biology, 98(5):1825–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/5/1825.

Cohen:1982:CSN

[CBJ^{+}82] W. D. Cohen, D. Bartelt, R. Jaeger, G. Langford, and I. Nemhauser. The cytoskeletal system of nucleated erythrocytes. I. Composition and function of major elements. Journal of Cell Biology, 93(3):828–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/93/3/828.

Cumming:1984:IDA

[CBL84] R. Cumming, R. D. Burgoyne, and N. A. Lytton. Immunocytochemical demonstration of alpha-tubulin modification during axonal maturation in the cerebellar cortex. Journal of Cell Biology, 98(1):347–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/1/347.

Cooper:1984:ACC

[CBP84] J. A. Cooper, J. D. Blum, and T. D. Pollard. Acanthamoeba castellanii capping protein: properties, mechanism of action, immunologic cross-reactivity, and localization. Journal of Cell Biology, 99(1):217–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/1/217.

Carlin:1983:IFM

[CBS83] R. K. Carlin, D. C. Bartelt, and P. Siekevitz. Identification of fodrin as a major calmodulin-binding protein in postsynaptic density preparations. Journal of Cell Biology, 96(2): 443–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/2/443.

Chaly:1984:CDN

[CBS^{+}84] N. Chaly, T. Bladon, G. Setterfield, J. E. Little, J. G. Kaplan, and D. L. Brown. Changes in distribution of nuclear matrix antigens during the mitotic cell cycle. Journal of Cell Biology, 99(2):661–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/2/661.

Carley:1981:FAA

[CBW81] W. W. Carley, L. S. Barak, and W. W. Webb. F-actin aggregates in transformed cells. Journal of Cell Biology, 90(3): 797–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/3/797.

Cheng:1980:ILC

[CBWW80] T. P. Cheng, F. I. Byrd, J. N. Whitaker, and J. G. Wood. Immunocytochemical localization of coated vesicle protein in rodent nervous system. Journal of Cell Biology, 86(2):624–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
86/2/624.

Cadwell:1982:ICJ

[CC82] J. J. Cadwell and A. H. Caswell. Identification of a constituent of the junctional feet linking terminal cisternae to transverse tubules in skeletal muscle. Journal of Cell Biology, 93(3):543–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/93/3/543.

Chamley-Campbell:1981:PDR

[CCCR81] J. H. Chamley-Campbell, G. R. Campbell, and R. Ross. Phenotype-dependent response of cultured aortic smooth muscle to serum mitogens. Journal of Cell Biology, 89(2):379–??, May 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
89/2/379.

Carraway:1982:MAA

[CCJC82] K. L. Carraway, R. F. Cerra, G. Jung, and C. A. Carraway. Membrane-associated actin from the microvillar membranes of ascites tumor cells. Journal of Cell Biology, 94(3):624– ??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/94/3/624.

Clark:1982:LIC

[CCKMH82] C. C. Clark, J. Crossland, G. Kaplan, and A. Martinez-Hernandez. Location and identification of the collagen found in the 14.5-d rat embryo visceral yolk sac. Journal of Cell Biology, 93(2):251–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/93/2/251.

Cox:1984:RPP

[CCWR84] A. C. Cox, R. C. Carroll, J. G. White, and G. H. Rao. Recycling of platelet phosphorylation and cytoskeletal assembly. Journal of Cell Biology, 98(1):8–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/1/8.

Choi:1984:MAC

[CD84] Y. D. Choi and G. Dreyfuss. Monoclonal antibody characterization of the C proteins of heterogeneous nuclear ribonucleoprotein complexes in vertebrate cells. Journal of Cell Biology, 99(6):1997–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/6/1997.

Cheung:1981:PTB

[CDLJ81] M. K. Cheung, D. T. Drivas, V. C. Littau, and E. M. Johnson. Protein tightly bound near the termini of the Physarum extrachromosomal rDNA palindrome. Journal of Cell Biology, 91 (1):309–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/91/1/309.

Conrad:1980:STI

[CDvdM80] G. W. Conrad, W. Dessau, and K. von der Mark. Synthesis of type III collagen by fibroblasts from the embryonic chick cornea. Journal of Cell Biology, 84(3):501–??, March 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/84/3/501.

Cancalon:1980:SRG

[CE80] P. Cancalon and J. S. Elam. Study of regeneration in the garfish olfactory nerve. Journal of Cell Biology, 84(3):779–??, March 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
84/3/779.

Carey:1983:BTI

[CEC^{+}83] D. J. Carey, C. F. Eldridge, C. J. Cornbrooks, R. Timpl, and R. P. Bunge. Biosynthesis of type IV collagen by cultured rat Schwann cells. Journal of Cell Biology, 97(2):473–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/2/
473.

Crossin:1984:MTC

[CEC84] K. L. Crossin, G. M. Edelman, and B. A. Cunningham. Mapping of three carbohydrate attachment sites in embryonic and adult forms of the neural cell adhesion molecule. Journal of Cell Biology, 99(5):1848–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/5/1848.

Cohen:1980:SDI

[CF80] C. M. Cohen and S. F. Foley. Spectrin-dependent and - independent association of F-actin with the erythrocyte membrane. Journal of Cell Biology, 86(2):694–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/86/2/694.

Collins:1983:DPH

[CF83] J. M. Collins and K. A. Foster. Differentiation of promyelocytic (HL-60) cells into mature granulocytes: mitochondrialspecific rhodamine 123 fluorescence. Journal of Cell Biology, 96(1):94–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/1/94.

Chiquet:1984:CMAa

[CF84a] M. Chiquet and D. M. Fambrough. Chick myotendinous antigen. I. A monoclonal antibody as a marker for tendon and muscle morphogenesis. Journal of Cell Biology, 98(6):1926–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
98/6/1926.

Chiquet:1984:CMAb

[CF84b] M. Chiquet and D. M. Fambrough. Chick myotendinous antigen. II. A novel extracellular glycoprotein complex consisting of large disulfide-linked subunits. Journal of Cell Biology, 98 (6):1937–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/6/1937.

Chuman:1982:CGP

[CFCS82] L. Chuman, L. G. Fine, A. H. Cohen, and M. H. Saier. Continuous growth of proximal tubular kidney epithelial cells in hormone-supplemented serum-free medium. Journal of Cell Biology, 94(3):506–??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/94/3/506.

Curtis:1983:ACP

[CFML83] A. S. Curtis, J. V. Forrester, C. McInnes, and F. Lawrie. Adhesion of cells to polystyrene surfaces. Journal of Cell Biology, 97(5):1500–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/5/1500.

Cunha:1983:EMI

[CFN^{+}83] G. R. Cunha, H. Fujii, B. L. Neubauer, J. M. Shannon, L. Sawyer, and B. A. Reese. Epithelial-mesenchymal interactions in prostatic development. I. morphological observations of prostatic induction by urogenital sinus mesenchyme in epithelium of the adult rodent urinary bladder. Journal of Cell Biology, 96(6):1662–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/6/1662.

Clinton:1984:TKA

[CFW84] G. M. Clinton and J. Finley-Whelan. Tyrosyl kinases acquired from anchorage-independent cells by a membrane-enveloped virus. Journal of Cell Biology, 99(3):788–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/99/3/788.

Cole:1984:CSA

[CG84] G. J. Cole and L. Glaser. Cell-substratum adhesion in embryonic chick central nervous system is mediated by a 170,000mol-wt neural-specific polypeptide. Journal of Cell Biology, 99(5):1605–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/5/1605.

Carpentier:1982:CLE

[CGA^{+}82] J. L. Carpentier, P. Gorden, R. G. Anderson, J. L. Goldstein, M. S. Brown, S. Cohen, and L. Orci. Co-localization of 125Iepidermal growth factor and ferritin-low density lipoprotein in coated pits: a quantitative electron microscopic study in normal and mutant human fibroblasts. Journal of Cell Biology, 95(1):73–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/95/1/73.

Carlin:1980:ICP

[CGCS80] R. K. Carlin, D. J. Grab, R. S. Cohen, and P. Siekevitz. Isolation and characterization of postsynaptic densities from various brain regions: enrichment of different types of postsynaptic densities. Journal of Cell Biology, 86(3):831–??, September 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/86/3/
831.

Connolly:1981:PFR

[CGG81] J. L. Connolly, S. A. Green, and L. A. Greene. Pit formation and rapid changes in surface morphology of sympathetic neurons in response to nerve growth factor. Journal of Cell Biology, 90(1):176–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/90/1/176.

Connolly:1984:CRC

[CGG84] J. L. Connolly, S. A. Green, and L. A. Greene. Comparison of rapid changes in surface morphology and coated pit formation of PC12 cells in response to nerve growth factor, epidermal growth factor, and dibutyryl cyclic AMP. Journal of Cell Biology, 98(2):457–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/2/457.

Chambard:1981:ICG

[CGM81] M. Chambard, J. Gabrion, and J. Mauchamp. Influence of collagen gel on the orientation of epithelial cell polarity: follicle formation from isolated thyroid cells and from preformed monolayers. Journal of Cell Biology, 91(1):157–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/91/1/
157.

Carlin:1981:FCP

[CGS81] R. K. Carlin, D. J. Grab, and P. Siekevitz. Function of a calmodulin in postsynaptic densities. III. Calmodulin-binding proteins of the postsynaptic density. Journal of Cell Biology, 89(3):449–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/89/3/449.

Cohen:1982:UCD

[CGW82] I. Cohen, J. M. Gerrard, and J. G. White. Ultrastructure of clots during isometric contraction. Journal of Cell Biology, 93 (3):775–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/93/3/775.

Ceccarelli:1980:CDR

[CH80a] B. Ceccarelli and W. P. Hurlbut. Ca^{2+}-dependent recycling of synaptic vesicles at the frog neuromuscular junction. Journal of Cell Biology, 87(1):297–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/87/1/297.

Chandler:1980:VLS

[CH80b] D. E. Chandler and J. Heuser. The vitelline layer of the sea urchin egg and its modification during fertilization. A freezefracture study using quick-freezing and deep-etching. Journal of Cell Biology, 84(3):618–??, March 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/84/3/618.

Chandler:1980:AMF

[CH80c] D. E. Chandler and J. E. Heuser. Arrest of membrane fusion events in mast cells by quick-freezing. Journal of Cell Biology, 86(2):666–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/86/2/666.

Cornell:1980:ACB

[CH80d] R. B. Cornell and A. F. Horwitz. Apparent coordination of the biosynthesis of lipids in cultured cells: its relationship to the regulation of the membrane sterol: phospholipid ratio and cell cycling. Journal of Cell Biology, 86(3):810–??, September 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/86/3/
810.

Cooper:1981:SDB

[CH81] J. A. Cooper and T. Hunter. Similarities and differences between the effects of epidermal growth factor and Rous sarcoma virus. Journal of Cell Biology, 91(3):878–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/91/3/878.

Cooper:1982:DPL

[CH82] J. A. Cooper and T. Hunter. Discrete primary locations of a tyrosine-protein kinase and of three proteins that contain phosphotyrosine in virally transformed chick fibroblasts. Journal of Cell Biology, 94(2):287–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/94/2/287.

Castellani:1983:CSS

[CH83a] L. Castellani and P. M. Hardwicke. Crystalline structure of sarcoplasmic reticulum from scallop. Journal of Cell Biology, 97 (2):557–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/2/557.

Cleveland:1983:AAC

[CH83b] D. W. Cleveland and J. C. Havercroft. Is apparent autoregulatory control of tubulin synthesis nontranscriptionally regulated? Journal of Cell Biology, 97(3):919–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/97/3/919.

Chrispeels:1982:RER

[CHCS82] M. J. Chrispeels, T. J. Higgins, S. Craig, and D. Spencer. Role of the endoplasmic reticulum in the synthesis of reserve proteins and the kinetics of their transport to protein bodies in developing pea cotyledons. Journal of Cell Biology, 93(1):5–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
93/1/5.

Chen:1981:MRT

[Che81] W. T. Chen. Mechanism of retraction of the trailing edge during fibroblast movement. Journal of Cell Biology, 90 (1):187–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/1/187.

Cook:1983:EPM

[CHH^{+}83] J. Cook, E. Hou, Y. Hou, A. Cairo, and D. Doyle. Establishment of plasma membrane domains in hepatocytes. I. Characterization and localization to the bile canaliculus of three antigens externally oriented in the plasma membrane. Journal of Cell Biology, 97(6):1823–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/6/1823.

Chiesi:1981:PRS

[CHI^{+}81] M. Chiesi, M. M. Ho, G. Inesi, A. V. Somlyo, and A. P. Somlyo. Primary role of sarcoplasmic reticulum in phasic contractile activation of cardiac myocytes with shunted myolemma. Journal of Cell Biology, 91(3):728–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/91/3/728.

Cerra:1984:EML

[CHRB84] R. F. Cerra, P. L. Haywood-Reid, and S. H. Barondes. Endogenous mammalian lectin localized extracellularly in lung elastic fibers. Journal of Cell Biology, 98(4):1580–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/4/
1580.

Couchman:1983:AGM

[CHRT83] J. R. Couchman, M. Höök, D. A. Rees, and R. Timpl. Adhesion, growth, and matrix production by fibroblasts on laminin substrates. Journal of Cell Biology, 96(1):177–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/1/
177.

Chrispeels:1982:ASP

[CHS82] M. J. Chrispeels, T. J. V. Higgins, and D. Spencer. Assembly of storage protein oligomers in the endoplasmic reticulum and processing of the polypeptides in the protein bodies of developing pea cotyledons. Journal of Cell Biology, 93(2):306–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/93/2/
306.

Carlson:1980:ASC

[CK80] S. S. Carlson and R. B. Kelly. An antiserum specific for cholinergic synaptic vesicles from electric organ. Journal of Cell Biology, 87(1):98–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/87/1/98.

Criscuolo:1982:STR

[CK82] B. A. Criscuolo and S. S. Krag. Selection of tunicamycinresistant Chinese hamster ovary cells with increased Nacetylglucosaminyltransferase activity. Journal of Cell Biology, 94(3):586–??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/94/3/586.

Caulfield:1980:AHN

[CKB^{+}80a] J. P. Caulfield, G. Korman, A. E. Butterworth, M. Hogan, and J. R. David. The adherence of human neutrophils and eosinophils to schistosomula: evidence for membrane fusion between cells and parasites. Journal of Cell Biology, 86(1):46–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
86/1/46.

Caulfield:1980:PCD

[CKB^{+}80b] J. P. Caulfield, G. Korman, A. E. Butterworth, M. Hogan, and J. R. David. Partial and complete detachment of neutrophils and eosinophils from schistosomula: evidence for the establishment of continuity between a fused and normal parasite membrane. Journal of Cell Biology, 86(1):64–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/86/1/64.

Cox:1981:CCE

[CKE81] G. N. Cox, M. Kusch, and R. S. Edgar. Cuticle of Caenorhabditis elegans: its isolation and partial characterization. Journal of Cell Biology, 90(1):7–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/90/1/7.

Colman:1982:SIM

[CKFS82] D. R. Colman, G. Kreibich, A. B. Frey, and D. D. Sabatini. Synthesis and incorporation of myelin polypeptides into CNS myelin. Journal of Cell Biology, 95(2):598–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/95/2/598.

Courtoy:1980:FLR

[CKHF80] P. J. Courtoy, Y. S. Kanwar, R. O. Hynes, and M. G. Farquhar. Fibronectin localization in the rat glomerulus. Journal of Cell Biology, 87(3):691–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/87/3/691.

Charp:1983:GCC

[CKJ83] P. A. Charp, R. J. Kinders, and T. C. Johnson. G2 cell cycle arrest induced by glycopeptides isolated from the bovine cerebral cortex. Journal of Cell Biology, 97(2):311–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/2/

Capco:1984:NMP

[CKP84] D. G. Capco, G. Krochmalnic, and S. Penman. A new method of preparing embeddment-free sections for transmission electron microscopy: applications to the cytoskeletal framework and other three-dimensional networks. Journal of Cell Biology, 98(5):1878–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/5/1878.

Caulfield:1982:HNE

[CKS82] J. P. Caulfield, G. Korman, and J. C. Samuelson. Human neutrophils endocytose multivalent ligands from the surface of schistosomula of Schistosoma mansoni before membrane fusion. Journal of Cell Biology, 94(2):370–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/94/2/370.

Chin:1980:UIL

[CL80a] N. W. Chin and K. W. Lanks. Use of immobilized lactoperoxidase to label L cell proteins involved in adhesion to polystyrene. Journal of Cell Biology, 85(2):402–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/85/2/402.

Craig:1980:CIB

[CL80b] S. W. Craig and C. L. Lancashire. Comparison of intestinal brush-border 95-Kdalton polypeptide and alpha-actinins. Journal of Cell Biology, 84(3):655–??, March 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/84/3/655.

Culpepper:1981:ITA

[CL81] J. A. Culpepper and A. Y. Liu. Induction of tyrosine aminotransferase in H-35 hepatoma cells by cAMP captured in phospholipid vesicles. Journal of Cell Biology, 88(1):89–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/88/1/
89.

Church:1982:MDM

[CL82] K. Church and H. P. Lin. Meiosis in Drosophila melanogaster. II. The prometaphase–I kinetochore microtubule bundle and kinetochore orientation in males. Journal of Cell Biology, 93 (2):365–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/93/2/365.

Clegg:1984:IWC

[Cle84] J. S. Clegg. Intracellular water and the cytomatrix: some methods of study and current views. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/1/167s.

Celis:1983:PKV

[CLFC83] J. E. Celis, P. M. Larsen, S. J. Fey, and A. Celis. Phosphorylation of keratin and vimentin polypeptides in normal and transformed mitotic human epithelial amnion cells: behavior of keratin and vimentin filaments during mitosis. Journal of Cell Biology, 97(5):1429–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/97/5/1429.

Caulfield:1980:SDH

[CLHA80] J. P. Caulfield, R. A. Lewis, A. Hein, and K. F. Austen. Secretion in dissociated human pulmonary mast cells. Evidence for solubilization of granule contents before discharge. Journal of Cell Biology, 85(2):299–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/85/2/299.

Christensen:1981:DSP

[CLJ^{+}81] M. E. Christensen, W. M. LeStourgeon, M. Jamrich, G. C. Howard, L. A. Serunian, L. M. Silver, and S. C. Elgin. Distribution studies on polytene chromosomes using antibodies directed against hnRNP. Journal of Cell Biology, 90(1):18–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
90/1/18.

Chinkers:1981:RRH

[CMC81] M. Chinkers, J. A. McKanna, and S. Cohen. Rapid rounding of human epidermoid carcinoma cells A-431 induced by epidermal growth factor. Journal of Cell Biology, 88(2): 422–??, February 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/88/2/422.

Chernoff:1983:ICL

[CMC83] E. A. G. Chernoff, G. A. Maresh, and L. A. Culp. Isolation and characterization of a large, neurite-associated glycoconjugate from neuroblastoma cells. Journal of Cell Biology, 96 (3):661–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/3/661.

Chou:1982:ISA

[CMF82] J. Y. Chou, T. Mano, and M. Feldman. Inhibition of synthesis of alpha-fetoprotein by glucocorticoids in cultured hepatoma cells. Journal of Cell Biology, 93(2):314–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/93/2/314.

Colman:1981:FSP

[CML^{+}81] A. Colman, J. Morser, C. Lane, J. Besley, C. Wylie, and G. Valle. Fate of secretory proteins trapped in oocytes of Xenopus laevis by disruption of the cytoskeleton or by imbalanced subunit synthesis. Journal of Cell Biology, 91(3): 770–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/91/3/770.

Cande:1981:PCM

[CMM81] W. Z. Cande, K. McDonald, and R. L. Meeusen. A permeabilized cell model for studying cell division: a comparison of anaphase chromosome movement and cleavage furrow constriction in lysed PtK1 cells. Journal of Cell Biology, 88(3):618–??, March 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
88/3/618.

Cohen:1980:ACM

[CN80] W. D. Cohen and I. Nemhauser. Association of centrioles with the marginal band of a molluscan erythrocyte. Journal of Cell Biology, 86(1):286–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/86/1/286.

Cornell:1980:CAM

[CNH80] R. B. Cornell, S. M. Nissley, and A. F. Horwitz. Cholesterol availability modulates myoblast fusion. Journal of Cell Biology, 86(3):820–??, September 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/86/3/820.

Chilton:1980:ICR

[CNSW80] B. Chilton, S. Nicosia, J. Sowinski, and D. Wolf. Isolation and characterization of rabbit endocervical cells. Journal of Cell Biology, 86(1):172–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/86/1/172.

Chen:1984:ETA

[COBC84] W. T. Chen, K. Olden, B. A. Bernard, and F. F. Chu. Expression of transformation-associated protease(s) that degrade fibronectin at cell contact sites. Journal of Cell Biology, 98 (4):1546–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/4/1546.

Connolly:1984:RCF

[Con84] J. A. Connolly. Role of the cytoskeleton in the formation, stabilization, and removal of acetylcholine receptor clusters in cultured muscle cells. Journal of Cell Biology, 99(1):148–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/1/148.

Crow:1983:MLC

[COS83] M. T. Crow, P. S. Olson, and F. E. Stockdale. Myosin lightchain expression during avian muscle development. Journal of Cell Biology, 96(3):736–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/3/736.

Crowther:1981:APC

[CP81] R. A. Crowther and B. M. Pearse. Assembly and packing of clathrin into coats. Journal of Cell Biology, 91(3):790–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
91/3/790.

Chow:1982:RCS

[CP82a] I. Chow and M. M. Poo. Redistribution of cell surface receptors induced by cell–cell contact. Journal of Cell Biology, 95(2): 510–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/95/2/510.

Chowrashi:1982:ZBD

[CP82b] P. K. Chowrashi and F. A. Pepe. The Z-band: 85,000dalton amorphin and alpha-actinin and their relation to structure. Journal of Cell Biology, 94(3):565–??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/94/3/565.

Capco:1983:MAC

[CP83] D. G. Capco and S. Penman. Mitotic architecture of the cell: the filament networks of the nucleus and cytoplasm. Journal of Cell Biology, 96(3):896–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/3/896.

Cavagnaro:1980:APP

[CPLC80] J. Cavagnaro, D. A. Pierce, J. C. Lucchesi, and C. B. Chae. Association of a protease with polytene chromosomes of Drosophila melanogaster. Journal of Cell Biology, 87(2): 415–??, November 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/87/2/415.

Chamberlain:1984:SSM

[CPS84] J. W. Chamberlain, J. W. Pollard, and C. P. Stanners. Selective synthesis of mitochondrial proteins by Chinese hamster ovary cells severely starved for various amino acids. Journal of Cell Biology, 98(4):1603–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/4/1603.

Courtoy:1984:SED

[CQB84] P. J. Courtoy, J. Quintart, and P. Baudhuin. Shift of equilibrium density induced by 3,3’-diaminobenzidine cytochemistry: a new procedure for the analysis and purification of peroxidase-containing organelles. Journal of Cell Biology, 98 (3):870–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/3/870.

Cohen:1982:SSS

[CRC^{+}82] C. Cohen, B. Reinhardt, L. Castellani, P. Norton, and M. Stirewalt. Schistosome surface spines are “crystals” of actin. Journal of Cell Biology, 95(3):987–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/95/3/987.

Creutz:1981:CUF

[Cre81] C. E. Creutz. cis-unsaturated fatty acids induce the fusion of chromaffin granules aggregated by synexin. Journal of Cell Biology, 91(1):247–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/91/1/247.

Couchman:1982:FDR

[CRGS82] J. R. Couchman, D. A. Rees, M. R. Green, and C. G. Smith. Fibronectin has a dual role in locomotion and anchorage of primary chick fibroblasts and can promote entry into the division cycle. Journal of Cell Biology, 93(2):402–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/93/2/402.

Carter:1981:SCA

[CRH81] W. G. Carter, H. Rauvala, and S. I. Hakomori. Studies on cell adhesion and recognition. II. The kinetics of cell adhesion and cell spreading on surfaces coated with carbohydrate-reactive proteins (glycosidases and lectins) and fibronectin. Journal of Cell Biology, 88(1):138–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/88/1/138.

Chen:1982:IMS

[CS82] W. T. Chen and S. J. Singer. Immunoelectron microscopic studies of the sites of cell-substratum and cell–cell contacts in cultured fibroblasts. Journal of Cell Biology, 95(1):205–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
95/1/205.

Crisona:1983:ICC

[CS83] N. J. Crisona and R. C. Strohman. Inhibition of contraction of cultured muscle fibers results in increased turnover of myofibrillar proteins but not of intermediate-filament proteins. Journal of Cell Biology, 96(3):684–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/3/684.

Chou:1981:SLD

[CSH81] J. Y. Chou and S. E. Schlegel-Haueter. Study of liver differentiation in vitro. Journal of Cell Biology, 89(2):216–??, May 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/89/2/
216.

Chaitin:1984:APC

[CSHP84] M. H. Chaitin, B. G. Schneider, M. O. Hall, and D. S. Papermaster. Actin in the photoreceptor connecting cilium: immunocytochemical localization to the site of outer segment disk formation. Journal of Cell Biology, 99(1):239–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/1/
239.

Catterall:1980:PSO

[CSK^{+}80] J. F. Catterall, J. P. Stein, P. Kristo, A. R. Means, and B. W. O’Malley. Primary sequence of ovomucoid messenger RNA as determined from cloned complementary DNA. Journal of Cell Biology, 87(2):480–??, November 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/87/2/480.

Cartaud:1981:CAT

[CSR^{+}81] J. Cartaud, A. Sobel, A. Rousselet, P. F. Devaux, and J. P. Changeux. Consequences of alkaline treatment for the ultrastructure of the acetylcholine-receptor-rich membranes from Torpedo marmorata electric organ. Journal of Cell Biology, 90 (2):418–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/2/418.

Cantini:1980:MTC

[CSS80] M. Cantini, S. Sartore, and S. Schiaffino. Myosin types in cultured muscle cells. Journal of Cell Biology, 85(3):903–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
85/3/903.

Czosnek:1980:SBN

[CSW80] H. Czosnek, D. Soifer, and H. M. Wisniewski. Studies on the biosynthesis of neurofilament proteins. Journal of Cell Biology, 85(3):726–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/85/3/726.

Chalfie:1982:SFD

[CT82] M. Chalfie and J. N. Thomson. Structural and functional diversity in the neuronal microtubules of Caenorhabditis elegans. Journal of Cell Biology, 93(1):15–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/93/1/15.

Coluccio:1983:UPC

[CT83] L. M. Coluccio and L. G. Tilney. Under physiological conditions actin disassembles slowly from the nonpreferred end of an actin filament. Journal of Cell Biology, 97(5):1629–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
97/5/1629.

Coluccio:1984:PEA

[CT84] L. M. Coluccio and L. G. Tilney. Phalloidin enhances actin assembly by preventing monomer dissociation. Journal of Cell Biology, 99(2):529–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/2/529.

Cande:1983:RCT

[CTKJ83] W. Z. Cande, P. J. Tooth, and J. Kendrick-Jones. Regulation of contraction and thick filament assembly-disassembly in glycerinated vertebrate smooth muscle cells. Journal of Cell Biology, 97(4):1062–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/4/1062.

Condeelis:1982:CPR

[CV82] J. Condeelis and M. Vahey. A calcium- and pH-regulated protein from Dictyostelium discoideum that cross-links actin filaments. Journal of Cell Biology, 94(2):466–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/94/2/466.

Clark:1980:ELE

[CVAK80] J. M. Clark, D. W. Vaughan, B. M. Aiken, and H. M. Kagan. Elastase-like enzymes in human neutrophils localized by ultrastructural cytochemistry. Journal of Cell Biology, 84(1): 102–??, January 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/84/1/102.

Condeelis:1984:PDA

[CVC^{+}84] J. Condeelis, M. Vahey, J. M. Carboni, J. DeMey, and S. Ogihara. Properties of the 120,000- and 95,000-dalton actinbinding proteins from Dictyostelium discoideum and their possible functions in assembling the cytoplasmic matrix. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/1/119s.

Chambard:1983:PTC

[CVGM83] M. Chambard, B. Verrier, J. Gabrion, and J. Mauchamp. Polarization of thyroid cells in culture: evidence for the basolateral localization of the iodide “pump” and of the thyroidstimulating hormone receptor-adenyl cyclase complex. Journal of Cell Biology, 96(4):1172–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/4/1172.

Carpentier:1981:SRI

[CVGO81] J. L. Carpentier, E. Van Obberghen, P. Gorden, and L. Orci. Surface redistribution of 125I-insulin in cultured human lymphocytes. Journal of Cell Biology, 91(1):17–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/91/1/17.

Chicheportiche:1984:CCO

[CVT84] Y. Chicheportiche, P. Vassalli, and A. M. Tartakoff. Characterization of cytoplasmically oriented Golgi proteins with a monoclonal antibody. Journal of Cell Biology, 99(6):2200–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/6/2200.

Cohen:1980:LAR

[CW80] M. W. Cohen and P. R. Weldon. Localization of acetylcholine receptors and synaptic ultrastructure at nerve-muscle contacts in culture: dependence on nerve type. Journal of Cell Biology, 86(2):388–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/86/2/388.

Cabral:1983:TRM

[CWBB83] F. Cabral, L. Wible, S. Brenner, and B. R. Brinkley. Taxolrequiring mutant of Chinese hamster ovary cells with impaired mitotic spindle assembly. Journal of Cell Biology, 97(1):30–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
97/1/30.

Christman:1980:HDD

[CWS^{+}80] J. K. Christman, N. Weich, B. Schoenbrun, N. Schneiderman, and G. Acs. Hypomethylation of DNA during differentiation of friend erythroleukemia cells. Journal of Cell Biology, 86 (2):366–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/86/2/366.

Dlugosz:1984:RBS

[DANH84] A. A. Dlugosz, P. B. Antin, V. T. Nachmias, and H. Holtzer. The relationship between stress fiber-like structures and nascent myofibrils in cultured cardiac myocytes. Journal of Cell Biology, 99(6):2268–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/6/2268.

Deamer:1980:EDM

[DB80] D. W. Deamer and D. E. Boatman. An enzymatically driven membrane reconstitution from solubilized components. Journal of Cell Biology, 84(2):461–??, February 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/84/2/461.

David:1981:TCR

[DB81a] G. David and M. Bernfield. Type I collagen reduces the degradation of basal lamina proteoglycan by mammary epithelial cells. Journal of Cell Biology, 91(1):281–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/91/1/281.

DeDuve:1981:SHT

[DB81b] C. De Duve and H. Beaufay. A short history of tissue fractionation. Journal of Cell Biology, 91(3):??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/91/3/
293s.

Deery:1983:CMA

[DB83] W. J. Deery and B. R. Brinkley. Cytoplasmic microtubule assembly-disassembly from endogenous tubulin in a Brij-lysed cell model. Journal of Cell Biology, 96(6):1631–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/6/
1631.

Dolby:1981:DRC

[DBBN81] T. W. Dolby, A. Belmont, T. W. Borun, and C. Nicolini. DNA replication, chromatin structure, and histone phosphorylation altered by theophylline in synchronized HeLa S3 cells. Journal of Cell Biology, 89(1):78–??, April 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/89/1/78.

DeBrabander:1981:IML

[DBG^{+}81] M. De Brabander, J. C. Bulinski, G. Geuens, J. De Mey, and G. G. Borisy. Immunoelectron microscopic localization of the 210,000-mol wt microtubule-associated protein in cultured cells of primates. Journal of Cell Biology, 91(2):438– ??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/91/2/438.

Davison:1980:GMR

[DBK80] P. M. Davison, K. Bensch, and M. A. Karasek. Growth and morphology of rabbit marginal vessel endothelium in cell culture. Journal of Cell Biology, 85(2):187–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/85/2/187.

Dyrberg:1982:SPB

[DBL82] T. Dyrberg, S. Baekkeskov, and A. Lernmark. Specific pancreatic beta-cell surface antigens recognized by a xenogenic antiserum. Journal of Cell Biology, 94(2):472–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/94/2/472.

DeCamilli:1983:SPNa

[DCG83] P. De Camilli, R. Cameron, and P. Greengard. Synapsin I (protein I), a nerve terminal-specific phosphoprotein. I. Its general distribution in synapses of the central and peripheral nervous system demonstrated by immunofluorescence in frozen and plastic sections. Journal of Cell Biology, 96(5):1337–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
96/5/1337.

Drubin:1984:SEM

[DCK84] D. G. Drubin, D. Caput, and M. W. Kirschner. Studies on the expression of the microtubule-associated protein, tau, during mouse brain development, with newly isolated complementary DNA probes. Journal of Cell Biology, 98(3):1090–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/3/
1090.

Dennis:1984:ALO

[DCS84] J. W. Dennis, J. P. Carver, and H. Schachter. Asparaginelinked oligosaccharides in murine tumor cells: comparison of a WGA-resistant (WGAr) nonmetastatic mutant and a related WGA-sensitive (WGAs) metastatic line. Journal of Cell Biology, 99(3):1034–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/3/1034.

deCouet:1984:MAA

[dCSB84] H. G. de Couet, S. Stowe, and A. D. Blest. Membraneassociated actin in the rhabdomeral microvilli of crayfish photoreceptors. Journal of Cell Biology, 98(3):834–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/3/
834.

Dentler:1980:MMIa

[Den80] W. L. Dentler. Microtubule-membrane interactions in cilia. I. Isolation and characterization of ciliary membranes from Tetrahymena pyriformis. Journal of Cell Biology, 84(2):364– ??, February 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/84/2/364.

Deutscher:1984:EAT

[Deu84] M. P. Deutscher. The eucaryotic aminoacyl-tRNA synthetase complex: suggestions for its structure and function. Journal of Cell Biology, 99(2):373–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/2/373.

Danto:1984:IAI

[DF84] S. I. Danto and D. A. Fischman. Immunocytochemical analysis of intermediate filaments in embryonic heart cells with monoclonal antibodies to desmin. Journal of Cell Biology, 98 (6):2179–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/6/2179.

Detmers:1983:EFT

[DGC83] P. A. Detmers, U. W. Goodenough, and J. Condeelis. Elongation of the fertilization tubule in Chlamydomonas: new observations on the core microfilaments and the effect of transient intracellular signals on their structural integrity. Journal of Cell Biology, 97(2):522–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/97/2/522.

Decker:1984:AMI

[DGD^{+}84] C. Decker, R. Greggs, K. Duggan, J. Stubbs, and A. Horwitz. Adhesive multiplicity in the interaction of embryonic fibroblasts and myoblasts with extracellular matrices. Journal of Cell Biology, 99(4):1398–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/4/1398.

Drenckhahn:1980:LMA

[DGS80] D. Drenckhahn and U. Gröschel-Stewart. Localization of myosin, actin, and tropomyosin in rat intestinal epithelium: immunohistochemical studies at the light and electron microscope levels. Journal of Cell Biology, 86(2):475–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/86/2/
475.

Dembo:1981:MPA

[DH81] M. Dembo and A. K. Harris. Motion of particles adhering to the leading lamella of crawling cells. Journal of Cell Biology, 91(2):528–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/91/2/528.

Das:1983:DRD

[DH83] O. Prem Das and E. J. Henderson. Developmental regulation of Dictyostelium discoideum plasma membrane proteins. Journal of Cell Biology, 97(5):1544–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/5/1544.

Dunn:1984:RME

[DH84] W. A. Dunn and A. L. Hubbard. Receptor-mediated endocytosis of epidermal growth factor by hepatocytes in the perfused rat liver: ligand and receptor dynamics. Journal of Cell Biology, 98(6):2148–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/6/2148.

DeCamilli:1983:SPNb

[DHHG83] P. De Camilli, S. M. Harris, W. B. Huttner, and P. Greengard. Synapsin I (Protein I), a nerve terminal-specific phosphoprotein. II. Its specific association with synaptic vesicles demonstrated by immunocytochemistry in agarose-embedded synaptosomes. Journal of Cell Biology, 96(5):1355–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/5/

Dutcher:1984:GDC

[DHL84] S. K. Dutcher, B. Huang, and D. J. Luck. Genetic dissection of the central pair microtubules of the flagella of Chlamydomonas reinhardtii. Journal of Cell Biology, 98(1):229–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/1/
229.

deHarven:1984:NAS

[dHLC84] E. de Harven, R. Leung, and H. Christensen. A novel approach for scanning electron microscopy of colloidal gold-labeled cell surfaces. Journal of Cell Biology, 99(1):53–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/99/1/53.

Dembitzer:1980:DDV

[DHS^{+}80] H. M. Dembitzer, F. Herz, A. Schermer, R. C. Wolley, and L. G. Koss. Desmosome development in an in vitro model. Journal of Cell Biology, 85(3):695–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/85/3/695.

Dvorak:1984:DBC

[DHS^{+}84] A. M. Dvorak, I. Hammel, E. S. Schulman, S. P. Peters, D. W. MacGlashan, R. P. Schleimer, H. H. Newball, K. Pyne, H. F. Dvorak, and L. M. Lichtenstein. Differences in the behavior of cytoplasmic granules and lipid bodies during human lung mast cell degranulation. Journal of Cell Biology, 99(5): 1678–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/5/1678.

Davison:1981:FPC

[DJ81] P. F. Davison and R. N. Jones. Filament proteins in central, cranial, and peripheral mammalian nerves. Journal of Cell Biology, 88(1):67–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/88/1/67.

Davis:1981:EML

[DK81a] R. Davis and G. B. Koelle. Electron microscope localization of acetylcholinesterase and butyrylcholinesterase in the superior cervical ganglion of the cat. II. Preganglionically denervated ganglion. Journal of Cell Biology, 88(3):581–??, March 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/88/3/581.

Davis:1981:DMD

[DK81b] R. L. Davis and J. A. Kiger. Dunce mutants of Drosophila melanogaster: mutants defective in the cyclic AMP phosphodiesterase enzyme system. Journal of Cell Biology, 90(1):101–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
90/1/101.

DAmore:1984:ECM

[DK84a] P. A. D’Amore and M. Klagsbrun. Endothelial cell mitogens derived from retina and hypothalamus: biochemical and biological similarities. Journal of Cell Biology, 99(4):1545–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/4/1545.

Darbre:1984:PSA

[DK84b] P. Darbre and R. J. King. Progression to steroid autonomy in S115 mouse mammary tumor cells: role of DNA methylation. Journal of Cell Biology, 99(4):1410–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/4/1410.

Damsky:1981:MCC

[DKDB81] C. H. Damsky, K. A. Knudsen, R. J. Dorio, and C. A. Buck. Manipulation of cell–cell and cell-substratum interactions in mouse mammary tumor epithelial cells using broad spectrum antisera. Journal of Cell Biology, 89(2):173–??, May 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/89/2/173.

deLaat:1983:QAM

[dLTvdSB83] S. W. de Laat, L. G. Tertoolen, P. T. van der Saag, and J. G. Bluemink. Quantitative analysis of modulations in numerical and lateral distribution of intramembrane particles during the cell cycle of neuroblastoma cells. Journal of Cell Biology, 96 (4):1047–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/4/1047.

Daukas:1983:RPP

[DLZ83] G. Daukas, D. A. Lauffenburger, and S. Zigmond. Reversible pinocytosis in polymorphonuclear leukocytes. Journal of Cell Biology, 96(6):1642–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/6/1642.

DesJardins:1983:PGM

[DM83] K. C. DesJardins and P. Morell. Phosphate groups modifying myelin basic proteins are metabolically labile; methyl groups are stable. Journal of Cell Biology, 97(2):438–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/97/2/438.

Deery:1984:CMA

[DMB84] W. J. Deery, A. R. Means, and B. R. Brinkley. Calmodulinmicrotubule association in cultured mammalian cells. Journal of Cell Biology, 98(3):904–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/3/904.

Dinauer:1980:CARa

[DMD80] M. C. Dinauer, S. A. MacKay, and P. N. Devreotes. Cyclic 3’,5’-AMP relay in Dictyostelium discoideum III. The relationship of cAMP synthesis and secretion during the cAMP signaling response. Journal of Cell Biology, 86(2):537–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/86/2/
537.

Daniel:1984:SCM

[DPK84] J. C. Daniel, B. U. Pauli, and K. E. Kuettner. Synthesis of cartilage matrix by mammalian chondrocytes in vitro. III. Effects of ascorbate. Journal of Cell Biology, 99(6):1960–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/6/
1960.

Mathieu:1981:DAC

[dPMMF^{+}81] D. di Padua Mathieu, C. V. Mura, L. L. Frado, C. L. Woodcock, and B. D. Stollar. Differing accessibility in chromatin of the antigenic sites of regions 1–58 and 63–125 of histone H2B. Journal of Cell Biology, 91(1):135–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/91/1/135.

Dentler:1980:MMIb

[DPS80] W. L. Dentler, M. M. Pratt, and R. E. Stephens. Microtubulemembrane interactions in cilia. II. Photochemical cross-linking of bridge structures and the identification of a membraneassociated dynein-like ATPase. Journal of Cell Biology, 84(2): 381–??, February 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/84/2/381.

Drake:1982:IPE

[DR82] D. K. Drake and S. D. Rosen. Identification and purification of an endogenous receptor for the lectin pallidin from Polysphondylium pallidum. Journal of Cell Biology, 93(2):383–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/93/2/
383.

Dunphy:1983:CAL

[DR83] W. G. Dunphy and J. E. Rothman. Compartmentation of asparagine-linked oligosaccharide processing in the Golgi apparatus. Journal of Cell Biology, 97(1):270–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/97/1/270.

Day:1984:IRM

[DRF84] D. A. Day, I. J. Ryrie, and N. Fuad. Investigations of the role of the main light-harvesting chlorophyll-protein complex in thylakoid membranes. Reconstitution of depleted membranes from intermittent-light-grown plants with the isolated complex. Journal of Cell Biology, 98(1):163–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/98/1/163.

Draper:1980:EDT

[DS80] R. K. Draper and M. I. Simon. The entry of diphtheria toxin into the mammalian cell cytoplasm: evidence for lysosomal involvement. Journal of Cell Biology, 87(3):849–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/87/3/
849.

Dinauer:1980:CARc

[DSD80a] M. Dinauer, T. L. Steck, and P. Devreotes. Cyclic 3’, 5’-AMP relay Dictyostelium discoideum. V. Adaptation of the cAMP signaling response during cAMP stimulation. Journal of Cell Biology, 86(2):554–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/86/2/554.

Dinauer:1980:CARb

[DSD80b] M. Dinauer, T. L. Steck, and P. Devreotes. Cyclic 3’, 5’-AMP relay in Dictyostelium discoideum. IV. Recovery of the cAMP signaling response after adaptation to cAMP. Journal of Cell Biology, 86(2):545–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/86/2/545.

daSilva:1984:LFM

[dSK84] P. Pinto da Silva and F. W. Kan. Label-fracture: a method for high resolution labeling of cell surfaces. Journal of Cell Biology, 99(3):1156–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/3/1156.

Downing:1984:HSG

[DSKS84] S. W. Downing, R. H. Spitzer, E. A. Koch, and W. L. Salo. The hagfish slime gland thread cell. I. A unique cellular system for the study of intermediate filaments and intermediate filament-microtubule interactions. Journal of Cell Biology, 98 (2):653–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/2/653.

Dennis:1984:LCP

[DSRF84] J. E. Dennis, T. Shimizu, F. C. Reinach, and D. A. Fischman. Localization of C-protein isoforms in chicken skeletal muscle: ultrastructural detection using monoclonal antibodies. Journal of Cell Biology, 98(4):1514–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/4/1514.

daSilva:1982:FFC

[dST82] P. P. da Silva and M. R. Torrisi. Freeze-fracture cytochemistry: partition of glycophorin in freeze-fractured human erythrocyte membranes. Journal of Cell Biology, 93(2):463–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/93/2/463.

daSilva:1981:FFC

[dSTK81] P. Pinto da Silva, M. R. Torrisi, and B. Kachar. Freezefracture cytochemistry: localization of wheat-germ agglutinin and concanavalin A binding sites on freeze-fractured pancreatic cells. Journal of Cell Biology, 91(2):361–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/91/2/
361.

DeRosier:1982:CTA

[DTBF82] D. J. DeRosier, L. G. Tilney, E. M. Bonder, and P. Frankl. A change in twist of actin provides the force for the extension of the acrosomal process in Limulus sperm: the false-discharge reaction. Journal of Cell Biology, 93(2):324–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/93/2/324.

Deutsch:1984:PHH

[DTP84] C. Deutsch, J. S. Taylor, and M. Price. pH homeostasis in human lymphocytes: modulation by ions and mitogen. Journal of Cell Biology, 98(3):885–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/3/885.

Deng:1981:NNA

[DTT81] J. S. Deng, Y. Takasaki, and E. M. Tan. Nonhistone nuclear antigens reactive with autoantibodies. Immunofluorescence studies on distribution in synchronized cells. Journal of Cell Biology, 91(3):654–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/91/3/654.

Dayer:1981:CSP

[DVB^{+}81] J. M. Dayer, J. D. Vassalli, J. L. Bobbitt, R. N. Hull, E. Reich, and S. M. Krane. Calcitonin stimulates plasminogen activator in porcine renal tubular cells: LLC–PK1. Journal of Cell Biology, 91(1):195–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/91/1/195.

Dessau:1981:EMF

[DVvdMvdM81] W. Dessau, B. M. Vertel, H. von der Mark, and K. von der Mark. Extracellular matrix formation by chondrocytes in monolayer culture. Journal of Cell Biology, 90(1):78–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/90/1/
78.

Diglio:1983:TRC

[DWM83] C. A. Diglio, D. E. Wolfe, and P. Meyers. Transformation of rat cerebral endothelial cells by Rous sarcoma virus. Journal of Cell Biology, 97(1):15–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/97/1/15.

Dickson:1981:AMA

[DWP81] R. B. Dickson, M. C. Willingham, and I. Pastan. alpha 2macroglobulin adsorbed to colloidal gold: a new probe in the study of receptor-mediated endocytosis. Journal of Cell Biology, 89(1):29–??, April 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/89/1/29.

Dennis:1984:IAL

[DWS84] J. W. Dennis, C. A. Waller, and V. Schirrmacher. Identification of asparagine-linked oligosaccharides involved in tumor cell adhesion to laminin and type IV collagen. Journal of Cell Biology, 99(4):1416–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/4/1416.

Dang:1983:AMT

[DYP83] C. V. Dang, D. C. Yang, and T. D. Pollard. Association of methionyl-tRNA synthetase with detergent-insoluble components of the rough endoplasmic reticulum. Journal of Cell Biology, 96(4):1138–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/4/1138.

Eaton:1983:PEM

[EB83] D. L. Eaton and J. B. Baker. Phorbol ester and mitogens stimulate human fibroblast secretions of plasmin-activatable plasminogen activator and protease nexin, an antiactivator/ antiplasmin. Journal of Cell Biology, 97(2):323–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/2/
323.

Einck:1984:FHA

[EB84] L. Einck and M. Bustin. Functional histone antibody fragments traverse the nuclear envelope. Journal of Cell Biology, 98(1):205–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/1/205.

Eichner:1984:CEK

[EBS84] R. Eichner, P. Bonitz, and T. T. Sun. Classification of epidermal keratins according to their immunoreactivity, isoelectric point, and mode of expression. Journal of Cell Biology, 98 (4):1388–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/4/1388.

Ecarot-Charrier:1983:OIM

[ECGvdRP83] B. Ecarot-Charrier, F. H. Glorieux, M. van der Rest, and G. Pereira. Osteoblasts isolated from mouse calvaria initiate matrix mineralization in culture. Journal of Cell Biology, 96 (3):639–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/3/639.

Erickson:1981:FMV

[ECM81] H. P. Erickson, N. Carrell, and J. McDonagh. Fibronectin molecule visualized in electron microscopy: a long, thin, flexible strand. Journal of Cell Biology, 91(3):673–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/91/3/
673.

East:1984:MAP

[ED84] I. J. East and J. Dean. Monoclonal antibodies as probes of the distribution of ZP-2, the major sulfated glycoprotein of the murine zona pellucida. Journal of Cell Biology, 98(3):795–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
98/3/795.

Ernst:1980:SDO

[EDK80] S. A. Ernst, W. C. Dodson, and K. J. Karnaky. Structural diversity of occluding junctions in the low-resistance chloridesecreting opercular epithelium of seawater-adapted killifish (Fundulus heteroclitus). Journal of Cell Biology, 87(2):488– ??, November 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/87/2/488.

Eckert:1982:AKP

[EDP82] B. S. Eckert, R. A. Daley, and L. M. Parysek. Assembly of keratin onto PtK1 cytoskeletons: evidence for an intermediate filament organizing center. Journal of Cell Biology, 92(2): 575–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/2/575.

Enenstein:1984:ICE

[EF84] J. Enenstein and L. T. Furcht. Isolation and characterization of epinectin, a novel adhesion protein for epithelial cells. Journal of Cell Biology, 99(2):464–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/2/464.

Egyhazi:1981:ASI

[EH81] E. Egyhazi and M. Holst. Adenosine selectively inhibits labeling of chromosomal RNA, especially hnRNA, probably by acting at or near the site of chain initiation. Journal of Cell Biology, 89(1):1–??, April 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/89/1/1.

Earnshaw:1984:KPM

[EHCR84] W. C. Earnshaw, N. Halligan, C. Cooke, and N. Rothfield. The kinetochore is part of the metaphase chromosome scaffold. Journal of Cell Biology, 98(1):352–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/1/352.

Evinger-Hodges:1982:IMD

[EHESF82] M. J. Evinger-Hodges, D. Z. Ewton, S. C. Seifert, and J. R. Florini. Inhibition of myoblast differentiation in vitro by a protein isolated from liver cell medium. Journal of Cell Biology, 93(2):395–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/93/2/395.

Eveloff:1980:SRM

[EHK80] J. Eveloff, W. Haase, and R. Kinne. Separation of renal medullary cells: isolation of cells from the thick ascending limb of Henle’s loop. Journal of Cell Biology, 87(3):672–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
87/3/672.

Emr:1980:MPL

[EHS80] S. D. Emr, M. N. Hall, and T. J. Silhavy. A mechanism of protein localization: the signal hypothesis and bacteria. Journal of Cell Biology, 86(3):701–??, September 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/86/3/701.

Euteneuer:1982:PSM

[EJM82] U. Euteneuer, W. T. Jackson, and J. R. McIntosh. Polarity of spindle microtubules in Haemanthus endosperm. Journal of Cell Biology, 94(3):644–??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/94/3/644.

Elhammer:1984:TEI

[EK84] A. Elhammer and S. Kornfeld. Two enzymes involved in the synthesis of O-linked oligosaccharides are localized on membranes of different densities in mouse lymphoma BW5147 cells. Journal of Cell Biology, 99(1):327–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/1/327.

Ekblom:1981:FBM

[Ekb81] P. Ekblom. Formation of basement membranes in the embryonic kidney: an immunohistological study. Journal of Cell Biology, 91(1):1–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/91/1/1.

Evercooren:1982:FPR

[EKS^{+}82] A. Baron-Van Evercooren, H. K. Kleinman, H. E. Seppä, B. Rentier, and M. Dubois-Dalcq. Fibronectin promotes rat Schwann cell growth and motility. Journal of Cell Biology, 93 (1):211–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/93/1/211.

Eckert:1980:ACP

[EKSZ80] B. S. Eckert, S. J. Koons, A. W. Schantz, and C. R. Zobel. Association of creatine phosphokinase with the cytoskeleton of cultured mammalian cells. Journal of Cell Biology, 86(1):1–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
86/1/1.

Eisen:1984:TSS

[EKWR84] A. Eisen, D. P. Kiehart, S. J. Wieland, and G. T. Reynolds. Temporal sequence and spatial distribution of early events of fertilization in single sea urchin eggs. Journal of Cell Biology, 99(5):1647–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/5/1647.

Earnshaw:1983:AMC

[EL83] W. C. Earnshaw and U. K. Laemmli. Architecture of metaphase chromosomes and chromosome scaffolds. Journal of Cell Biology, 96(1):84–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/1/84.

Evans:1984:ARE

[EL84] E. Evans and A. Leung. Adhesivity and rigidity of erythrocyte membrane in relation to wheat germ agglutinin binding. Journal of Cell Biology, 98(4):1201–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/4/1201.

Enns:1983:CMI

[ELS^{+}83] C. A. Enns, J. W. Larrick, H. Suomalainen, J. Schroder, and H. H. Sussman. Co-migration and internalization of transferrin and its receptor on K562 cells. Journal of Cell Biology, 97 (2):579–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/2/579.

Ekblom:1981:SCT

[ELST81] P. Ekblom, E. Lehtonen, L. Saxén, and R. Timpl. Shift in collagen type as an early response to induction of the metanephric mesenchyme. Journal of Cell Biology, 89(2):276–??, May 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/89/2/276.

Euteneuer:1980:PMP

[EM80] U. Euteneuer and J. R. McIntosh. Polarity of midbody and phragmoplast microtubules. Journal of Cell Biology, 87(2): 509–??, November 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/87/2/509.

Euteneuer:1981:SPK

[EM81] U. Euteneuer and J. R. McIntosh. Structural polarity of kinetochore microtubules in PtK1 cells. Journal of Cell Biology, 89(2):338–??, May 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/89/2/338.

Endo:1984:DED

[EM84] T. Endo and T. Masaki. Differential expression and distribution of chicken skeletal- and smooth-muscle-type alphaactinins during myogenesis in culture. Journal of Cell Biology, 99(6):2322–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/6/2322.

Ellisman:1983:MMT

[EMA83] M. H. Ellisman, J. A. Miller, and W. S. Agnew. Molecular morphology of the tetrodotoxin-binding sodium channel protein from Electrophorus electricus in solubilized and reconstituted preparations. Journal of Cell Biology, 97(6):1834–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
97/6/1834.

English:1983:VSK

[EMC83] L. H. English, I. G. Macara, and L. C. Cantley. Vanadium stimulates the (Na+,K+) pump in friend erythroleukemia cells and blocks erythropoiesis. Journal of Cell Biology, 97(4): 1299–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/4/1299.

Erickson:1984:EFM

[EN84] C. A. Erickson and R. Nuccitelli. Embryonic fibroblast motility and orientation can be influenced by physiological electric fields. Journal of Cell Biology, 98(1):296–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/98/1/296.

Ellisman:1980:MSA

[EP80] M. H. Ellisman and K. R. Porter. Microtrabecular structure of the axoplasmic matrix: visualization of cross-linking structures and their distribution. Journal of Cell Biology, 87(2): 464–??, November 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/87/2/464.

Erikson:1980:MEC

[EPE^{+}80] R. L. Erikson, A. F. Purchio, E. Erikson, M. S. Collett, and J. S. Brugge. Molecular events in cells transformed by Rous Sarcoma virus. Journal of Cell Biology, 87(2):319–??, November 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
87/2/319.

Ehmann:1984:GMM

[EPM84] U. K. Ehmann, W. D. Peterson, and D. S. Misfeldt. To grow mouse mammary epithelial cells in culture. Journal of Cell Biology, 98(3):1026–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/3/1026.

Egyhazi:1984:PSC

[EPO^{+}84] E. Egyházi, A. Pigon, A. Ossoinak, M. Holst, and U. Tayip. Phosphorylation of some chromosomal nonhistone proteins in active genes is blocked by the transcription inhibitor 5,6– dichloro–1–beta–D–ribofuranosylbenzimidazole (DRB). Journal of Cell Biology, 98(3):954–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/3/954.

Eppenberger:1981:MMP

[EPRS81] H. M. Eppenberger, J. C. Perriard, U. B. Rosenberg, and E. E. Strehler. The Mr 165,000 M-protein myomesin: a specific protein of cross-striated muscle cells. Journal of Cell Biology, 89 (2):185–??, May 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/89/2/185.

Eisen:1984:CTD

[ER84] A. Eisen and G. T. Reynolds. Calcium transients during early development in single starfish (Asterias forbesi) oocytes. Journal of Cell Biology, 99(5):1878–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/5/1878.

Euteneuer:1983:PKM

[ERB83] U. Euteneuer, H. Ris, and G. G. Borisy. Polarity of kinetochore microtubules in Chinese hamster ovary cells after recovery from a colcemid block. Journal of Cell Biology, 97 (1):202–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/1/202.

Earnshaw:1982:HSU

[ERH82] W. C. Earnshaw, O. P. Rekvig, and K. Hannestad. Histones synthesized for use in early development of Xenopus laevis are stored as a complex with antigenic properties similar to those of the octamer core of nucleosomes. Journal of Cell Biology, 92 (3):871–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/3/871.

Eisenbach:1983:BSE

[ERN83] L. Eisenbach, R. Ramanathan, and D. L. Nelson. Biochemical studies of the excitable membrane of paramecium tetraurelia. IX. Antibodies against ciliary membrane proteins. Journal of Cell Biology, 97(5):1412–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/5/1412.

Ernst:1981:ULK

[ES81a] S. A. Ernst and J. H. Schreiber. Ultrastructural localization of Na+,K+-ATPase in rat and rabbit kidney medulla. Journal of Cell Biology, 91(3):803–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/91/3/803.

Ernster:1981:MHR

[ES81b] L. Ernster and G. Schatz. Mitochondria: a historical review. Journal of Cell Biology, 91(3):??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/91/3/227s.

Emr:1982:MCS

[ES82] S. D. Emr and T. J. Silhavy. Molecular components of the signal sequence that function in the initiation of protein export. Journal of Cell Biology, 95(3):689–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/95/3/689.

Edidin:1982:LDH

[EW82] M. Edidin and T. Wei. Lateral diffusion of H-2 antigens on mouse fibroblasts. Journal of Cell Biology, 95(2):458–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
95/2/458.

Erdos:1983:FDI

[EW83] G. W. Erdos and D. Whitaker. Failure to detect immunocytochemically reactive endogenous lectin on the cell surface of Dictyostelium discoideum. Journal of Cell Biology, 97(4): 993–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/4/993.

Easter:1983:SIH

[EWB83] D. W. Easter, J. B. Wade, and J. L. Boyer. Structural integrity of hepatocyte tight junctions. Journal of Cell Biology, 96(3):745–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/3/745.

Edidin:1982:DRC

[EWG82] M. Edidin, T. Wei, and L. Gotlib. Diffusion rates of cell surface antigens of mouse-human heterokaryons. III. Regulation of lateral diffusion rates by calcium ions. Journal of Cell Biology, 95(2):453–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/95/2/453.

Edidin:1984:LDW

[EZ84] M. Edidin and M. Zuniga. Lateral diffusion of wild-type and mutant Ld antigens in L cells. Journal of Cell Biology, 99(6): 2333–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/6/2333.

Fowler:1982:PAP

[FA82] W. E. Fowler and U. Aebi. Polymorphism of actin paracrystals induced by polylysine. Journal of Cell Biology, 93(2):452–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
93/2/452.

Fowler:1983:CPA

[FA83] W. E. Fowler and U. Aebi. A consistent picture of the actin filament related to the orientation of the actin molecule. Journal of Cell Biology, 97(1):264–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/97/1/264.

Franzini-Armstrong:1981:SMC

[FAP81] C. Franzini-Armstrong and L. D. Peachey. Striated musclecontractile and control mechanisms. Journal of Cell Biology, 91(3):??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/91/3/166s.

Feramisco:1980:DFL

[FB80] Feramisco and S. H. Blose. Distribution of flourescently labeled α-actinin in living and fixed fibroblasts. Journal of Cell Biology, 86(2):608–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/86/2/608.

Fletcher:1982:DCL

[FB82] W. H. Fletcher and C. V. Byus. Direct cytochemical localization of catalytic subunits dissociated from cAMP-dependent protein kinase in Reuber H-35 hepatoma cells. I. Development and validation of fluorescinated inhibitor. Journal of Cell Biology, 93(3):719–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/93/3/719.

Fisher:1982:ICP

[FBB82] P. A. Fisher, M. Berrios, and G. Blobel. Isolation and characterization of a proteinaceous subnuclear fraction composed of nuclear matrix, peripheral lamina, and nuclear pore complexes from embryos of Drosophila melanogaster. Journal of Cell Biology, 92(3):674–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/92/3/674.

Fitzgerald:1983:ILM

[FBH83] P. G. Fitzgerald, D. Bok, and J. Horwitz. Immunocytochemical localization of the main intrinsic polypeptide (MIP) in ultrathin frozen sections of rat lens. Journal of Cell Biology, 97(5):1491–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/5/1491.

Finne:1982:EBL

[FBP82] J. Finne, M. M. Burger, and J. P. Prieels. Enzymatic basis for a lectin-resistant phenotype: increase in a fucosyltransferase in mouse melanoma cells. Journal of Cell Biology, 92 (2):277–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/2/277.

Fox:1984:AFC

[FBRP84] J. E. Fox, J. K. Boyles, C. C. Reynolds, and D. R. Phillips. Actin filament content and organization in unstimulated platelets. Journal of Cell Biology, 98(6):1985–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/6/
1985.

Fussle:1981:MMD

[FBS^{+}81] R. Füssle, S. Bhakdi, A. Sziegoleit, J. Tranum-Jensen, T. Kranz, and H. J. Wellensiek. On the mechanism of membrane damage by Staphylococcus aureus alpha-toxin. Journal of Cell Biology, 91(1):83–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/91/1/83.

Friedman:1984:RIH

[FBST84] M. J. Friedman, T. Blankenberg, G. Sensabaugh, and T. S. Tenforde. Recognition and invasion of human erythrocytes by malarial parasites: contribution of sialoglycoproteins to attachment and host specificity. Journal of Cell Biology, 98 (5):1672–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/5/1672.

Fechheimer:1982:PLM

[FC82] M. Fechheimer and J. J. Cebra. Phosphorylation of lymphocyte myosin catalyzed in vitro and in intact cells. Journal of Cell Biology, 93(2):261–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/93/2/261.

Fey:1984:ESR

[FCKP84] E. G. Fey, D. G. Capco, G. Krochmalnic, and S. Penman. Epithelial structure revealed by chemical dissection and unembedded electron microscopy. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/1/203s.

Fehlmann:1982:BME

[FCL^{+}82] M. Fehlmann, J. L. Carpentier, A. Le Cam, P. Thamm, D. Saunders, D. Brandenburg, L. Orci, and P. Freychet. Biochemical and morphological evidence that the insulin receptor is internalized with insulin in hepatocytes. Journal of Cell Biology, 93(1):82–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/93/1/82.

Furie:1984:CEC

[FCNS84] M. B. Furie, E. B. Cramer, B. L. Naprstek, and S. C. Silverstein. Cultured endothelial cell monolayers that restrict the transendothelial passage of macromolecules and electrical current. Journal of Cell Biology, 98(3):1033–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/3/
1033.

Feldherr:1983:EMP

[FCO83] C. M. Feldherr, R. J. Cohen, and J. A. Ogburn. Evidence for mediated protein uptake by amphibian oocyte nuclei. Journal of Cell Biology, 96(5):1486–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/5/1486.

Fifkova:1982:CAN

[FD82] E. Fifková and R. J. Delay. Cytoplasmic actin in neuronal processes as a possible mediator of synaptic plasticity. Journal of Cell Biology, 95(1):345–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/95/1/345.

Fielding:1982:RST

[FDKF82] P. E. Fielding, P. M. Davison, M. A. Karasek, and C. J. Fielding. Regulation of sterol transport in human microvascular endothelial cells. Journal of Cell Biology, 94(2):350–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/94/2/
350.

Fougere-Deschatrette:1984:SCC

[FDSWW84] C. Fougere-Deschatrette, R. T. Schimke, D. Weil, and M. C. Weiss. A study of chromosomal changes associated with amplified dihydrofolate reductase genes in rat hepatoma cells and their dedifferentiated variants. Journal of Cell Biology, 99(2): 497–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/2/497.

Fesus:1981:HBM

[FFEL81] L. Fésüs, A. Falus, A. Erdei, and K. Laki. Human beta 2microglobulin is a substrate of tissue transglutaminase: polymerization in solution and on the cell surface. Journal of Cell Biology, 89(3):706–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/89/3/706.

Fujikawa:1984:BMC

[FFGC84] L. S. Fujikawa, C. S. Foster, I. K. Gipson, and R. B. Colvin. Basement membrane components in healing rabbit corneal epithelial wounds: immunofluorescence and ultrastructural studies. Journal of Cell Biology, 98(1):128–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/98/1/128.

Fay:1983:DAA

[FFRF83] F. S. Fay, K. Fujiwara, D. D. Rees, and K. E. Fogarty. Distribution of alpha-actinin in single isolated smooth muscle cells. Journal of Cell Biology, 96(3):783–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/3/783.

Fujiki:1982:PPC

[FFS^{+}82] Y. Fujiki, S. Fowler, H. Shio, A. L. Hubbard, and P. B. Lazarow. Polypeptide and phospholipid composition of the membrane of rat liver peroxisomes: comparison with endoplasmic reticulum and mitochondrial membranes. Journal of Cell Biology, 93(1):103–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/93/1/103.

Findly:1980:ORG

[FG80a] R. C. Findly and J. G. Gall. Organization of ribosomal genes in Paramecium tetraurelia. Journal of Cell Biology, 84(3):547–??, March 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
84/3/547.

Fink:1980:ATP

[FG80b] D. J. Fink and H. Gainer. Axonal transport of proteins. A new view using in vivo covalent labeling. Journal of Cell Biology, 85(2):175–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/85/2/175.

Fallon:1981:FHH

[FG81] R. F. Fallon and D. A. Goodenough. Five-hour half-life of mouse liver gap-junction protein. Journal of Cell Biology, 90 (2):521–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/2/521.

Fesce:1980:FFS

[FGHC80] R. Fesce, F. Grohovaz, W. P. Hurlbut, and B. Ceccarelli. Freeze-fracture studies of frog neuromuscular junctions during intense release of neurotransmitter. III. A morphometric analysis of the number and diameter of intramembrane particles. Journal of Cell Biology, 85(2):337–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/85/2/337.

Fitch:1982:DBM

[FGS^{+}82] J. M. Fitch, E. Gibney, R. D. Sanderson, R. Mayne, and T. F. Linsenmayer. Domain and basement membrane specificity of a monoclonal antibody against chicken type IV collagen. Journal of Cell Biology, 95(2):641–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/95/2/641.

Ferrell:1984:PMM

[FH84] J. E. Ferrell and W. H. Huestis. Phosphoinositide metabolism and the morphology of human erythrocytes. Journal of Cell Biology, 98(6):1992–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/6/1992.

Fayet:1982:RPT

[FHDL82] G. Fayet, S. Hovsépian, J. G. Dickson, and S. Lissitzky. Reorganization of porcine thyroid cells into functional follicles in a chemically defined, serum- and thyrotropin-free medium.

Journal of Cell Biology, 93(2):479–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/93/2/479.

Fujiki:1982:IIM

[FHFL82] Y. Fujiki, A. L. Hubbard, S. Fowler, and P. B. Lazarow. Isolation of intracellular membranes by means of sodium carbonate treatment: application to endoplasmic reticulum. Journal of Cell Biology, 93(1):97–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/93/1/97.

Franke:1981:AMI

[FHG^{+}81] W. W. Franke, H. W. Heid, C. Grund, S. Winter, C. Freudenstein, E. Schmid, E. D. Jarasch, and T. W. Keenan. Antibodies to the major insoluble milk fat globule membrane-associated protein: specific location in apical regions of lactating epithelial cells. Journal of Cell Biology, 89(3):485–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/89/3/485.

Fisher:1982:MFF

[Fis82a] K. A. Fisher. Monolayer freeze-fracture autoradiography: quantitative analysis of the transmembrane distribution of radioiodinated concanavalin A. Journal of Cell Biology, 93 (1):155–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/93/1/155.

Fisher:1982:SAM

[Fis82b] K. A. Fisher. Spectroscopic assays for measuring quantities of erythrocyte membrane “halves”. Journal of Cell Biology, 92 (1):44–??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/1/44.

Fishman:1982:IDC

[Fis82c] P. H. Fishman. Internalization and degradation of cholera toxin by cultured cells: relationship to toxin action. Journal of Cell Biology, 93(3):860–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/93/3/860.

Freedman:1982:HIPa

[FJ82a] S. D. Freedman and J. D. Jamieson. Hormone-induced protein phosphorylation. I. Relationship between secretagogue action and endogenous protein phosphorylation in intact cells from the exocrine pancreas and parotid. Journal of Cell Biology, 95(3):903–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/95/3/903.

Freedman:1982:HIPb

[FJ82b] S. D. Freedman and J. D. Jamieson. Hormone-induced protein phosphorylation. II. Localization to the ribosomal fraction from rat exocrine pancreas and parotid of a 29,000-dalton protein phosphorylated in situ in response to secretagogues. Journal of Cell Biology, 95(3):909–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/95/3/909.

Freedman:1982:HIPc

[FJ82c] S. D. Freedman and J. D. Jamieson. Hormone-induced protein phosphorylation. III. regulation of the phosphorylation of the secretagogue-responsive 29,000-dalton protein by both Ca^{2+} and cAMP in vitro. Journal of Cell Biology, 95(3):918–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
95/3/918.

Fukui:1980:DNA

[FK80] Y. Fukui and H. Katsumaru. Dynamics of nuclear actin bundle induction by dimethyl sulfoxide and factors affecting its development. Journal of Cell Biology, 84(1):131–??, January 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/84/1/
131.

Feig:1983:MPM

[FKB83] L. A. Feig, M. Klagsbrun, and A. R. Bellvé. Mitogenic polypeptide of the mammalian seminiferous epithelium: biochemical characterization and partial purification. Journal of Cell Biology, 97(5):1435–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/5/1435.

Forbes:1983:SNR

[FKK83] D. J. Forbes, T. B. Kornberg, and M. W. Kirschner. Small nuclear RNA transcription and ribonucleoprotein assembly in early Xenopus development. Journal of Cell Biology, 97 (1):62–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/1/62.

Franke:1981:NSP

[FKS^{+}81] W. W. Franke, J. A. Kleinschmidt, H. Spring, G. Krohne, C. Grund, M. F. Trendelenburg, M. Stoehr, and U. Scheer. A nucleolar skeleton of protein filaments demonstrated in amplified nucleoli of Xenopus laevis. Journal of Cell Biology, 90 (2):289–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/2/289.

Feldherr:1984:MKP

[FKS84] C. M. Feldherr, E. Kallenbach, and N. Schultz. Movement of a karyophilic protein through the nuclear pores of oocytes. Journal of Cell Biology, 99(6):2216–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/6/2216.

Fleischer:1981:OGG

[Fle81] B. Fleischer. Orientation of glycoprotein galactosyltransferase and sialyltransferase enzymes in vesicles derived from rat liver Golgi apparatus. Journal of Cell Biology, 89(2):246–??, May 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/89/2/
246.

Fowler:1981:SPA

[FLH^{+}81] V. M. Fowler, E. J. Luna, W. R. Hargreaves, D. L. Taylor, and D. Branton. Spectrin promotes the association of F-actin with the cytoplasmic surface of the human erythrocyte membrane. Journal of Cell Biology, 88(2):388–??, February 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/88/2/388.

Fakan:1984:UDN

[FLM84] S. Fakan, G. Leser, and T. E. Martin. Ultrastructural distribution of nuclear ribonucleoproteins as visualized by immunocytochemistry on thin sections. Journal of Cell Biology, 98 (1):358–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/1/358.

Fry:1980:REC

[FM80] D. M. Fry and M. F. Morales. A reexamination of the effects of creatine on muscle protein synthesis in tissue culture. Journal of Cell Biology, 84(2):294–??, February 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/84/2/294.

Fitch:1983:DAB

[FML83] J. M. Fitch, R. Mayne, and T. F. Linsenmayer. Developmental acquisition of basement membrane heterogeneity: type IV collagen in the avian lens capsule. Journal of Cell Biology, 97(3): 940–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/3/940.

Falcone:1984:LHF

[FMS^{+}84] D. J. Falcone, N. Mated, H. Shio, C. R. Minick, and S. D. Fowler. Lipoprotein-heparin-fibronectin-denatured collagen complexes enhance cholesteryl ester accumulation in macrophages. Journal of Cell Biology, 99(4):1266–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/4/
1266.

Fallon:1980:LCS

[FN80] J. R. Fallon and V. T. Nachmias. Localization of cytoplasmic and skeletal myosins in developing muscle cells by doublelabel immunofluorescence. Journal of Cell Biology, 87(1): 237–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/87/1/237.

Ferro-Novick:1984:GRC

[FNHSS84] S. Ferro-Novick, W. Hansen, I. Schauer, and R. Schekman. Genes required for completion of import of proteins into the endoplasmic reticulum in yeast. Journal of Cell Biology, 98 (1):44–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/1/44.

Ferro-Novick:1984:YSM

[FNNFS84] S. Ferro-Novick, P. Novick, C. Field, and R. Schekman. Yeast secretory mutants that block the formation of active cell surface enzymes. Journal of Cell Biology, 98(1):35–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/1/
35.

Frantz:1981:IUC

[FNS81] C. N. Frantz, D. G. Nathan, and C. D. Scher. Intracellular univalent cations and the regulation of the BALB/c-3T3 cell cycle. Journal of Cell Biology, 88(1):51–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/88/1/51.

Fischer:1980:FRL

[FNSB80] H. D. Fischer, M. Natowicz, W. S. Sly, and R. K. Bretthauer. Fibroblast receptor for lysosomal enzymes mediates pinocytosis of multivalent phosphomannan fragment. Journal of Cell Biology, 84(1):77–??, January 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/84/1/77.

Feldherr:1980:MSN

[FO80] C. M. Feldherr and J. A. Ogburn. Mechanism for the selection of nuclear polypeptides in Xenopus oocytes. II. Twodimensional gel analysis. Journal of Cell Biology, 87(3):589– ??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/87/3/589.

Farquhar:1981:GAC

[FP81a] M. G. Farquhar and G. E. Palade. The Golgi apparatus (complex)-(1954–1981)-from artifact to center stage. Journal of Cell Biology, 91(3):??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/91/3/77s.

Findly:1981:RTG

[FP81b] R. C. Findly and T. Pederson. Regulated transcription of the genes for actin and heat-shock proteins in culture Drosophila cells. Journal of Cell Biology, 88(2):323–??, February 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/88/2/323.

Fulton:1981:DRS

[FPFP81] A. B. Fulton, J. Prives, S. R. Farmer, and S. Penman. Developmental reorganization of the skeletal framework and its surface lamina in fusing muscle cells. Journal of Cell Biology, 91(1):103–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/91/1/103.

Fletcher:1981:CPP

[FQBN81] D. J. Fletcher, J. P. Quigley, G. E. Bauer, and B. D. Noe. Characterization of proinsulin- and proglucagon-converting activities in isolated islet secretory granules. Journal of Cell Biology, 90(2):312–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/90/2/312.

Fries:1981:TAG

[FR81] E. Fries and J. E. Rothman. Transient activity of Golgi-like membranes as donors of vesicular stomatitis viral glycoprotein in vitro. Journal of Cell Biology, 90(3):697–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/90/3/
697.

Friend:1982:PMD

[Fri82a] D. S. Friend. Plasma-membrane diversity in a highly polarized cell. Journal of Cell Biology, 93(2):243–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/93/2/243.

Fristrom:1982:SJI

[Fri82b] D. K. Fristrom. Septate junctions in imaginal disks of Drosophila: a model for the redistribution of septa during cell rearrangement. Journal of Cell Biology, 94(1):77–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/94/1/
77.

Frixione:1983:FSA

[Fri83] E. Frixione. Firm structural associations between migratory pigment granules and microtubules in crayfish retinula cells. Journal of Cell Biology, 96(5):1258–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/5/1258.

Froehner:1984:PPP

[Fro84] S. C. Froehner. Peripheral proteins of postsynaptic membranes from Torpedo electric organ identified with monoclonal antibodies. Journal of Cell Biology, 99(1):88–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/99/1/88.

Field:1980:LSA

[FS80] C. Field and R. Schekman. Localized secretion of acid phosphatase reflects the pattern of cell surface growth in saccharomyces cerevisiae. Journal of Cell Biology, 86(1):123–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/86/1/
123.

Freter:1984:RMO

[FS84] R. R. Freter and R. M. Schultz. Regulation of murine oocyte meiosis: evidence for a gonadotropin-induced, cAMPdependent reduction in a maturation inhibitor. Journal of Cell Biology, 98(3):1119–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/3/1119.

Falkner:1981:TDM

[FSB81] F. G. Falkner, H. Saumweber, and H. Biessmann. Two Drosophila melanogaster proteins related to intermediate filament proteins of vertebrate cells. Journal of Cell Biology, 91 (1):175–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/91/1/175.

Florkiewicz:1983:ISM

[FSBR83] R. Z. Florkiewicz, A. Smith, J. E. Bergmann, and J. K. Rose. Isolation of stable mouse cell lines that express cell surface and secreted forms of the vesicular stomatitis virus glycoprotein. Journal of Cell Biology, 97(5):1381–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/5/
1381.

Franke:1980:ISF

[FSF^{+}80] W. W. Franke, E. Schmid, C. Freudenstein, B. Appelhans, M. Osborn, K. Weber, and T. W. Keenan. Intermediate-sized filaments of the prekeratin type in myoepithelial cells. Journal of Cell Biology, 84(3):633–??, March 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/84/3/633.

Ferguson:1984:SSJ

[FSFA84] D. G. Ferguson, H. W. Schwartz, and C. Franzini-Armstrong. Subunit structure of junctional feet in triads of skeletal muscle: a freeze-drying, rotary-shadowing study. Journal of Cell Biology, 99(5):1735–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/5/1735.

Franke:1981:NEA

[FSKJ81] W. W. Franke, U. Scheer, G. Krohne, and E. D. Jarasch. The nuclear envelope and the architecture of the nuclear periphery. Journal of Cell Biology, 91(3):??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/91/3/39s.

Franke:1980:PPR

[FSVW80] W. W. Franke, E. Schmid, J. Vandekerckhove, and K. Weber. Permanently proliferating rat vascular smooth muscle cell with maintained expression of smooth muscle characteristics, including actin of the vascular smooth muscle type. Journal of Cell Biology, 87(3):594–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/87/3/594.

Franke:1983:CCF

[FSW^{+}83] W. W. Franke, E. Schmid, J. Wellsteed, C. Grund, O. Gigi, and B. Geiger. Change of cytokeratin filament organization during the cell cycle: selective masking of an immunologic determinant in interphase PtK2 cells. Journal of Cell Biology, 97(4):1255–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/4/1255.

Fowler:1980:SPB

[FT80] V. Fowler and D. L. Taylor. Spectrin plus band 4.1 crosslink actin. Regulation by micromolar calcium. Journal of Cell Biology, 85(2):361–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/85/2/361.

Futrelle:1982:CBD

[FTM82] R. P. Futrelle, J. Traut, and W. G. McKee. Cell behavior in Dictyostelium discoideum: preaggregation response to localized cyclic AMP pulses. Journal of Cell Biology, 92(3):807–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
92/3/807.

Fukui:1980:FMC

[Fuk80] Y. Fukui. Formation of multinuclear cells induced by dimethyl sulfoxide: inhibition of cytokinesis and occurrence of novel nuclear division in Dictyostelium cells. Journal of Cell Biology, 86(1):181–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/86/1/181.

Fulton:1984:AAC

[Ful84] A. B. Fulton. Assembly associated with the cytomatrix. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/1/209s.

Fulton:1980:CDF

[FW80] C. Fulton and C. Walsh. Cell differentiation and flagellar elongation in Naegleria gruberi. Dependence on transcription and translation. Journal of Cell Biology, 85(2):346–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/85/2/346.

Franke:1981:ICD

[FWG^{+}81] W. W. Franke, S. Winter, C. Grund, E. Schmid, D. L. Schiller, and E. D. Jarasch. Isolation and characterization of desmosome-associated tonofilaments from rat intestinal brush border. Journal of Cell Biology, 90(1):116–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/90/1/116.

Fey:1984:ECF

[FWP84] E. G. Fey, K. M. Wan, and S. Penman. Epithelial cytoskeletal framework and nuclear matrix-intermediate filament scaffold: three-dimensional organization and protein composition. Journal of Cell Biology, 98(6):1973–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/6/1973.

Flanagan:1982:PMP

[FYC82] S. D. Flanagan, B. Yost, and G. Crawford. Putative 51,000-Mr protein marker for postsynaptic densities is virtually absent in cerebellum. Journal of Cell Biology, 94(3):743–??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/94/3/
743.

Goud:1984:ESI

[GA84] B. Goud and J. C. Antoine. Emergence of a surface immunoglobulin recycling process during B lymphocyte differentiation. Journal of Cell Biology, 98(4):1238–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/4/
1238.

Grover:1983:RLE

[GAA83] A. Grover, G. Andrews, and E. D. Adamson. Role of laminin in epithelium formation by F9 aggregates. Journal of Cell Biology, 97(1):137–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/1/137.

Gunther:1982:FAI

[GAAG82] S. Gunther, R. W. Alexander, W. J. Atkinson, and M. A. Gimbrone. Functional angiotensin II receptors in cultured vascular smooth muscle cells. Journal of Cell Biology, 92(2): 289–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/2/289.

Gall:1981:CSC

[Gal81] J. G. Gall. Chromosome structure and the C-value paradox. Journal of Cell Biology, 91(3):??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/91/3/3s.

Geiger:1984:MOC

[GAR^{+}84] B. Geiger, Z. Avnur, G. Rinnerthaler, H. Hinssen, and V. J. Small. Microfilament-organizing centers in areas of cell contact: cytoskeletal interactions during cell attachment and locomotion. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/1/83s.

Geiger:1982:RMM

[GAS82] B. Geiger, Z. Avnur, and J. Schlessinger. Restricted mobility of membrane constituents in cell-substrate focal contacts of chicken fibroblasts. Journal of Cell Biology, 93(2):495–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/93/2/
495.

Glass:1983:DEM

[GAS^{+}83a] R. H. Glass, J. Aggeler, A. Spindle, R. A. Pederson, and Z. Werb. Degradation of extracellular matrix by mouse trophoblast outgrowths: a model for implantation. Journal of Cell Biology, 96(4):1108–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/4/1108.

Gruenberg:1983:SEM

[GAS83b] J. Gruenberg, D. R. Allred, and I. W. Sherman. Scanning electron microscope-analysis of the protrusions (knobs) present on the surface of Plasmodium falciparum-infected erythrocytes. Journal of Cell Biology, 97(3):795–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/97/3/795.

Gundersen:1980:CTR

[GB80] R. W. Gundersen and J. N. Barrett. Characterization of the turning response of dorsal root neurites toward nerve growth factor. Journal of Cell Biology, 87(3):546–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/87/3/546.

Goldman:1981:VBC

[GB81] B. M. Goldman and G. Blobel. In vitro biosynthesis, core glycosylation and membrane integration of opsin. Journal of Cell Biology, 90(1):236–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/90/1/236.

Goldberg:1982:BSTb

[GB82] B. D. Goldberg and R. E. Burgeson. Binding of soluble type I collagen to fibroblasts: specificities for native collagen types, triple helical structure, telopeptides, propeptides, and cyanogen bromide-derived peptides. Journal of Cell Biology, 95(3):752–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/95/3/752.

Gasson:1983:NDG

[GB83] J. C. Gasson and S. Bourgeois. A new determinant of glucocorticoid sensitivity in lymphoid cell lines. Journal of Cell Biology, 96(2):409–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/2/409.

Griffiths:1982:VMP

[GBB^{+}82] G. Griffiths, R. Brands, B. Burke, D. Louvard, and G. Warren. Viral membrane proteins acquire galactose in trans Golgi cisternae during intracellular transport. Journal of Cell Biology, 95(3):781–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/95/3/781.

Greenberg:1984:CDK

[GBE84] M. E. Greenberg, R. Brackenbury, and G. M. Edelman. Changes in the distribution of the 34-kdalton tyrosine kinase substrate during differentiation and maturation of chicken tissues. Journal of Cell Biology, 98(2):473–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/98/2/473.

Gibson:1984:SLM

[GBF84] G. J. Gibson, B. W. Beaumont, and M. H. Flint. Synthesis of a low molecular weight collagen by chondrocytes from the presumptive calcification region of the embryonic chick sterna: the influence of culture with collagen gels. Journal of Cell Biology, 99(1):208–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/1/208.

Gingrich:1982:RSCa

[GBG82] J. C. Gingrich, L. K. Blaha, and A. N. Glazer. Rod substructure in cyanobacterial phycobilisomes: analysis of Synechocystis 6701 mutants low in phycoerythrin. Journal of Cell Biology, 92(2):261–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/92/2/261.

Gordon:1981:HBH

[GBL81] J. S. Gordon, J. Bruno, and J. J. Lucas. Heterogeneous binding of high mobility group chromosomal proteins to nuclei. Journal of Cell Biology, 88(2):373–??, February 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/88/2/373.

Gauthier:1983:MIN

[GBLH83] G. F. Gauthier, R. E. Burke, S. Lowey, and A. W. Hobbs. Myosin isozymes in normal and cross-reinnervated cat skeletal muscle fibers. Journal of Cell Biology, 97(3):756–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
97/3/756.

Gambino:1984:EMT

[GBM84] J. Gambino, L. G. Bergen, and N. R. Morris. Effects of mitotic and tubulin mutations on microtubule architecture in actively growing protoplasts of Aspergillus nidulans. Journal of Cell Biology, 99(3):830–??, September 1984. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/99/3/830.

Giddings:1980:VPC

[GBS80] T. H. Giddings, D. L. Brower, and L. A. Staehelin. Visualization of particle complexes in the plasma membrane of Micrasterias denticulata associated with the formation of cellulose fibrils in primary and secondary cell walls. Journal of Cell Biology, 84(2):327–??, February 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/84/2/327.

Gilmore:1982:PTAa

[GBW82] R. Gilmore, G. Blobel, and P. Walter. Protein translocation across the endoplasmic reticulum. I. Detection in the microsomal membrane of a receptor for the signal recognition particle. Journal of Cell Biology, 95(2):463–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/95/2/463.

Godelaine:1983:AMB

[GBWR83] D. Godelaine, H. Beaufay, M. Wibo, and A. M. Ravoet. Alteration of membrane barrier in stripped rough microsomes from rat liver on incubation with GTP: its relevance to the stimulation by this nucleotide of the dolichol pathway for protein glycosylation. Journal of Cell Biology, 97(2):340–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/2/
340.

Gusse:1980:EME

[GC80] M. Gusse and P. Chevaillier. Electron microscope evidence for the presence of globular structures in different sperm chromatins. Journal of Cell Biology, 87(1):280–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/87/1/280.

Gorelick:1983:CSP

[GCF^{+}83] F. S. Gorelick, J. A. Cohn, S. D. Freedman, N. G. Delahunt, J. M. Gershoni, and J. D. Jamieson. Calmodulin-stimulated protein kinase activity from rat pancreas. Journal of Cell Biology, 97(4):1294–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/4/1294.

Gould:1984:DTP

[GCH84] K. L. Gould, J. A. Cooper, and T. Hunter. The 46,000dalton tyrosine protein kinase substrate is widespread, whereas the 36,000-dalton substrate is only expressed at high levels in certain rodent tissues. Journal of Cell Biology, 98(2): 487–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/2/487.

Glickman:1983:GMC

[GCKAA83] J. Glickman, K. Croen, S. Kelly, and Q. Al-Awqati. Golgi membranes contain an electrogenic H+ pump in parallel to a chloride conductance. Journal of Cell Biology, 97(4): 1303–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/4/1303.

Gospodarowicz:1983:BBP

[GCL83] D. Gospodarowicz, J. Cheng, and M. Lirette. Bovine brain and pituitary fibroblast growth factors: comparison of their abilities to support the proliferation of human and bovine vascular endothelial cells. Journal of Cell Biology, 97(6):1677–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
97/6/1677.

Grab:1981:FCPa

[GCS81a] D. J. Grab, R. K. Carlin, and P. Siekevitz. Function of calmodulin in postsynaptic densities. I. Presence of a calmodulinactivatable cyclic nucleotide phosphodiesterase activity. Journal of Cell Biology, 89(3):433–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/89/3/433.

Grab:1981:FCPb

[GCS81b] D. J. Grab, R. K. Carlin, and P. Siekevitz. Function of calmodulin in postsynaptic densities. II. Presence of a calmodulin- activatable protein kinase activity. Journal of Cell Biology, 89 (3):440–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/89/3/440.

Goodenough:1982:ACF

[GDH82] U. W. Goodenough, P. A. Detmers, and C. Hwang. Activation for cell fusion in Chlamydomonas: analysis of wild-type gametes and nonfusing mutants. Journal of Cell Biology, 92(2): 378–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/2/378.

Gennaro:1983:NTC

[GDH^{+}83] R. Gennaro, B. Dewald, U. Horisberger, H. U. Gubler, and M. Baggiolini. A novel type of cytoplasmic granule in bovine neutrophils. Journal of Cell Biology, 96(6):1651–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/6/
1651.

Goodenough:1980:LMC

[GDL80] D. A. Goodenough, J. S. Dick, and J. E. Lyons. Lens metabolic cooperation: a study of mouse lens transport and permeability visualized with freeze-substitution autoradiography and electron microscopy. Journal of Cell Biology, 86(2):576–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/86/2/
576.

Glaser:1980:DVA

[GDM^{+}80] B. M. Glaser, P. A. D’Amore, R. G. Michels, A. Patz, and A. Fenselau. Demonstration of vasoproliferative activity from mammalian retina. Journal of Cell Biology, 84(2): 298–??, February 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/84/2/298.

Galli:1982:MCC

[GDM^{+}82] S. J. Galli, A. M. Dvorak, J. A. Marcum, T. Ishizaka, G. Nabel, H. Der Simonian, K. Pyne, J. M. Goldin, R. D. Rosenberg, H. Cantor, and H. F. Dvorak. Mast cell clones: a model for the analysis of cellular maturation. Journal of Cell Biology, 95 (2):435–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/95/2/435.

Gelman:1981:SCL

[GDMG81] B. B. Gelman, M. H. Davis, R. E. Morris, and E. Gruenstein. Structural changes in lysosomes from cultured human fibroblasts in Duchenne’s muscular dystrophy. Journal of Cell Biology, 88(2):329–??, February 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/88/2/329.

Gajdusek:1980:ECD

[GDRS80] C. Gajdusek, P. DiCorleto, R. Ross, and S. M. Schwartz. An endothelial cell-derived growth factor. Journal of Cell Biology, 85(2):467–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/85/2/467.

Griepp:1983:PPM

[GDRS83] E. B. Griepp, W. J. Dolan, E. S. Robbins, and D. D. Sabatini. Participation of plasma membrane proteins in the formation of tight junction by cultured epithelial cells. Journal of Cell Biology, 96(3):693–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/3/693.

Geiger:1981:IMS

[GDTS81] B. Geiger, A. H. Dutton, K. T. Tokuyasu, and S. J. Singer. Immunoelectron microscope studies of membrane-microfilament interactions: distributions of alpha-actinin, tropomyosin, and vinculin in intestinal epithelial brush border and chicken gizzard smooth muscle cells. Journal of Cell Biology, 91(3): 614–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/91/3/614.

Gurney:1980:IDL

[GE80] T. Gurney and G. L. Eliceiri. Intracellular distribution of low molecular weight RNA species in HeLa cells. Journal of Cell Biology, 87(2):398–??, November 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/87/2/398.

Grumet:1984:HBB

[GE84] M. Grumet and G. M. Edelman. Heterotypic binding between neuronal membrane vesicles and glial cells is mediated by a specific cell adhesion molecule. Journal of Cell Biology, 98 (5):1746–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/5/1746.

Gardner:1983:FED

[GF83] J. M. Gardner and D. M. Fambrough. Fibronectin expression during myogenesis. Journal of Cell Biology, 96(2):474– ??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/2/474.

Ghitescu:1984:SCD

[GF84] L. Ghitescu and A. Fixman. Surface charge distribution on the endothelial cell of liver sinusoids. Journal of Cell Biology, 99 (2):639–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/2/639.

Gonzalez-Fernandez:1984:ERB

[GFLG^{+}84] F. Gonzalez-Fernandez, R. A. Landers, P. A. Glazebrook, S. L. Fong, G. I. Liou, D. M. Lam, and C. D. Bridges. An extracellular retinol-binding glycoprotein in the eyes of mutant rats with retinal dystrophy: development, localization, and biosynthesis. Journal of Cell Biology, 99(6):2092–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/6/2092.

Gibbons:1980:CIQ

[GG80] B. H. Gibbons and I. R. Gibbons. Calcium-induced quiescence in reactivated sea urchin sperm. Journal of Cell Biology, 84 (1):13–??, January 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/84/1/13.

Gardiner:1983:MJX

[GG83] D. M. Gardiner and R. D. Grey. Membrane junctions in Xenopus eggs: their distribution suggests a role in calcium regulation. Journal of Cell Biology, 96(4):1159–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/4/

Gabel:1982:LEO

[GGK82] C. A. Gabel, D. E. Goldberg, and S. Kornfeld. Lysosomal enzyme oligosaccharide phosphorylation in mouse lymphoma cells: specificity and kinetics of binding to the mannose 6phosphate receptor in vivo. Journal of Cell Biology, 95(2): 536–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/95/2/536.

Goldberg:1983:SBM

[GGK83] D. E. Goldberg, C. A. Gabel, and S. Kornfeld. Studies of the biosynthesis of the mannose 6-phosphate receptor in receptorpositive and -deficient cell lines. Journal of Cell Biology, 97(6): 1700–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/6/1700.

Grunwald:1980:EDE

[GGL80] G. B. Grunwald, R. L. Geller, and J. Lilien. Enzymatic dissection of embryonic cell adhesive mechanisms. Journal of Cell Biology, 85(3):766–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/85/3/766.

Garofalo:1983:CAI

[GGS83] R. S. Garofalo, D. M. Gilligan, and B. H. Satir. Calmodulin antagonists inhibit secretion in Paramecium. Journal of Cell Biology, 96(4):1072–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/4/1072.

Gipson:1983:HFV

[GGSB83] I. K. Gipson, S. M. Grill, S. J. Spurr, and S. J. Brennan. Hemidesmosome formation in vitro. Journal of Cell Biology, 97(3):849–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/3/849.

Gonzalez:1981:EGF

[GGV81] A. González, J. Garrido, and J. D. Vial. Epidermal growth factor inhibits cytoskeleton-related changes in the surface of parietal cells. Journal of Cell Biology, 88(1):108–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/88/1/
108.

Glabe:1982:CSS

[GGVR82] C. G. Glabe, L. B. Grabel, V. D. Vacquier, and S. D. Rosen. Carbohydrate specificity of sea urchin sperm bindin: a cell surface lectin mediating sperm-egg adhesion. Journal of Cell Biology, 94(1):123–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/94/1/123.

Glenney:1983:SRM

[GGW83] J. R. Glenney, P. Glenney, and K. Weber. The spectrinrelated molecule, TW-260/240, cross-links the actin bundles of the microvillus rootlets in the brush borders of intestinal epithelial cells. Journal of Cell Biology, 96(5):1491–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/5/
1491.

Goodenough:1982:SOD

[GH82a] U. W. Goodenough and J. E. Heuser. Substructure of the outer dynein arm. Journal of Cell Biology, 95(3):798–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/95/3/
798.

Greenburg:1982:ESC

[GH82b] G. Greenburg and E. D. Hay. Epithelia suspended in collagen gels can lose polarity and express characteristics of migrating mesenchymal cells. Journal of Cell Biology, 95(1):333–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
95/1/333.

Geetha-Habib:1982:SMF

[GHB82] M. Geetha-Habib and G. B. Bouck. Synthesis and mobilization of flagellar glycoproteins during regeneration in Euglena. Journal of Cell Biology, 93(2):432–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/93/2/432.

Gospodarowicz:1981:FCP

[GHGT81] D. Gospodarowicz, K. Hirabayashi, L. Giguère, and J. P. Tauber. Factors controlling the proliferative rate, final cell density, and life span of bovine vascular smooth muscle cells in culture. Journal of Cell Biology, 89(3):568–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/89/3/568.

Grinnell:1982:FCS

[GHH82] F. Grinnell, J. R. Head, and J. Hoffpauir. Fibronectin and cell shape in vivo: studies on the endometrium during pregnancy. Journal of Cell Biology, 94(3):597–??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/94/3/597.

Goodloe-Holland:1984:MCD

[GHL84] C. M. Goodloe-Holland and E. J. Luna. A membrane cytoskeleton from Dictyostelium discoideum. III. Plasma membrane fragments bind predominantly to the sides of actin filaments. Journal of Cell Biology, 99(1):71–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/99/1/71.

Githens:1980:CDI

[GHWR80] S. Githens, D. R. Holmquist, J. F. Whelan, and J. R. Ruby. Characterization of ducts isolated from the pancreas of the rat. Journal of Cell Biology, 85(1):122–??, April 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/85/1/122.

Geisow:1981:TCL

[GHY81] M. J. Geisow, P. D’Arcy Hart, and M. R. Young. Temporal changes of lysosome and phagosome pH during phagolysosome formation in macrophages: studies by fluorescence spectroscopy. Journal of Cell Biology, 89(3):645–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/89/3/645.

Gibbons:1980:ISL

[Gib80] B. H. Gibbons. Intermittent swimming in live sea urchin sperm. Journal of Cell Biology, 84(1):1–??, January 1980.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/84/1/1.

Gibbons:1981:CFE

[Gib81] I. R. Gibbons. Cilia and flagella of eukaryotes. Journal of Cell Biology, 91(3):??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/91/3/107s.

Grab:1984:GAG

[GIKR84] D. J. Grab, S. Ito, U. A. Kara, and L. Rovis. Glycosyltransferase activities in Golgi complex and endoplasmic reticulum fractions isolated from African trypanosomes. Journal of Cell Biology, 99(2):569–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/2/569.

Gilkey:1983:RCP

[Gil83] J. C. Gilkey. Roles of calcium and pH in activation of eggs of the medaka fish, Oryzias latipes. Journal of Cell Biology, 97(3):669–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/3/669.

Gatmaitan:1983:RGD

[GJRO^{+}83] Z. Gatmaitan, D. M. Jefferson, N. Ruiz-Opazo, L. Biempica, I. M. Arias, G. Dudas, L. A. Leinwand, and L. M. Reid. Regulation of growth and differentiation of a rat hepatoma cell line by the synergistic interactions of hormones and collagenous substrata. Journal of Cell Biology, 97(4):1179–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/4/
1179.

Goldstein:1981:DPB

[GK81a] L. Goldstein and C. Ko. Distribution of proteins between nucleus and cytoplasm of amoeba proteus. Journal of Cell Biology, 88(3):516–??, March 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/88/3/516.

Goldstein:1981:SML

[GK81b] S. Goldstein and L. B. Korczack. Status of mitochondria in living human fibroblasts during growth and senescence in vitro: use of the laser dye rhodamine 123. Journal of Cell Biology, 91(2):392–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/91/2/392.

Gounon:1981:ICP

[GK81c] P. Gounon and E. Karsenti. Involvement of contractile proteins in the changes in consistency of oocyte nucleoplasm of the newt Pleurodeles waltlii. Journal of Cell Biology, 88(2): 410–??, February 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/88/2/410.

Green:1981:CPC

[GK81d] K. J. Green and D. L. Kirk. Cleavage patterns, cell lineages, and development of a cytoplasmic bridge system in Volvox embryos. Journal of Cell Biology, 91(3):743–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/91/3/743.

Green:1982:RCL

[GK82] K. J. Green and D. L. Kirk. A revision of the cell lineages recently reported for Volvox carteri embryos. Journal of Cell Biology, 94(3):741–??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/94/3/741.

Gabel:1984:TBG

[GK84] C. A. Gabel and S. Kornfeld. Targeting of beta-glucuronidase to lysosomes in mannose 6-phosphate receptor-deficient MOPC 315 cells. Journal of Cell Biology, 99(1):296–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/1/
296.

Garoff:1983:ESF

[GKKPB83] H. Garoff, C. Kondor-Koch, R. Pettersson, and B. Burke. Expression of Semliki Forest virus proteins from cloned complementary DNA. II. The membrane-spanning glycoprotein E2 is transported to the cell surface without its normal cytoplasmic domain. Journal of Cell Biology, 97(3):652–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/3/
652.

Grove:1984:NDS

[GKL^{+}84] B. K. Grove, V. Kurer, C. Lehner, T. C. Doetschman, J. C. Perriard, and H. M. Eppenberger. A new 185,000-dalton skeletal muscle protein detected by monoclonal antibodies. Journal of Cell Biology, 98(2):518–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/2/518.

Granzow:1981:RNG

[GKZ81] C. Granzow, M. Kopun, and H. P. Zimmermann. Role of nuclear glycogen synthase and cytoplasmic UDP glucose pyrophosphorylase in the biosynthesis of nuclear glycogen in HD33 Ehrlich–Lettré ascites tumor cells. Journal of Cell Biology, 89(3):475–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/89/3/475.

Gallin:1980:NCV

[GL80a] E. K. Gallin and D. R. Livengood. Nonlinear current-voltage relationships in cultured macrophages. Journal of Cell Biology, 85(1):160–??, April 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/85/1/160.

Griffin:1980:RRN

[GL80b] C. G. Griffin and P. C. Letourneau. Rapid retraction of neurites by sensory neurons in response to increased concentrations of nerve growth factor. Journal of Cell Biology, 86 (1):156–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/86/1/156.

Garner:1981:CAT

[GL81] J. A. Garner and R. J. Lasek. Clathrin is axonally transported as part of slow component b: the microfilament complex. Journal of Cell Biology, 88(1):172–??, January 1981.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/88/1/172.

Granger:1982:SIM

[GL82] D. L. Granger and A. L. Lehninger. Sites of inhibition of mitochondrial electron transport in macrophage-injured neoplastic cells. Journal of Cell Biology, 95(2):527–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/95/2/527.

Gomer:1983:HHF

[GL83a] R. H. Gomer and E. Lazarides. Highly homologous filamin polypeptides have different distributions in avian slow and fast muscle fibers. Journal of Cell Biology, 97(3):818–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
97/3/818.

Gomer:1983:SFP

[GL83b] R. H. Gomer and E. Lazarides. Switching of filamin polypeptides during myogenesis in vitro. Journal of Cell Biology, 96 (2):321–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/2/321.

Glacy:1983:PTC

[Gla83a] S. D. Glacy. Pattern and time course of rhodamine-actin incorporation in cardiac myocytes. Journal of Cell Biology, 96 (4):1164–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/4/1164.

Glacy:1983:SDR

[Gla83b] S. D. Glacy. Subcellular distribution of rhodamine-actin microinjected into living fibroblastic cells. Journal of Cell Biology, 97(4):1207–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/4/1207.

Gauthier:1982:DPM

[GLBH82] G. F. Gauthier, S. Lowey, P. A. Benfield, and A. W. Hobbs. Distribution and properties of myosin isozymes in developing avian and mammalian skeletal muscle fibers. Journal of Cell Biology, 92(2):471–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/92/2/471.

Gunning:1981:DSA

[GLBS81] P. W. Gunning, G. E. Landreth, M. A. Bothwell, and E. M. Shooter. Differential and synergistic actions of nerve growth factor and cyclic AMP in PC12 cells. Journal of Cell Biology, 89(2):240–??, May 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/89/2/240.

Gospodarowicz:1984:CAB

[GLMW84] D. Gospodarowicz, J. Lepine, S. Massoglia, and I. Wood. Comparison of the ability of basement membranes produced by corneal endothelial and mouse-derived Endodermal PF–HR–9 cells to support the proliferation and differentiation of bovine kidney tubule epithelial cells in vitro. Journal of Cell Biology, 99(3):947–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/3/947.

Guy:1980:LDF

[GLP80] J. Guy, R. Leclerc, and G. Pelletier. Localization of a 16,000dalton fragment of the common precursor of adrenocorticotropin and beta-lipotropin in the rat and human pituitary gland. Journal of Cell Biology, 86(3):825–??, September 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/86/3/825.

Greene:1983:RHM

[GLS83] L. A. Greene, R. K. Liem, and M. L. Shelanski. Regulation of a high molecular weight microtubule-associated protein in PC12 cells by nerve growth factor. Journal of Cell Biology, 96 (1):76–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/1/76.

Gerton:1984:GFC

[GM84] G. L. Gerton and C. F. Millette. Generation of flagella by cultured mouse spermatids. Journal of Cell Biology, 98(2): 619–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/2/619.

Gross:1982:PAC

[GMJR82] J. L. Gross, D. Moscatelli, E. A. Jaffe, and D. B. Rifkin. Plasminogen activator and collagenase production by cultured capillary endothelial cells. Journal of Cell Biology, 95(3): 974–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/95/3/974.

Gudewicz:1980:FMU

[GML^{+}80] P. W. Gudewicz, J. Molnar, M. Z. Lai, D. W. Beezhold, G. E. Siefring, R. B. Credo, and L. Lorand. Fibronectinmediated uptake of gelatin-coated latex particles by peritoneal macrophages. Journal of Cell Biology, 87(2):427–??, November 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/87/2/
427.

Goshima:1984:IIF

[GMO84] K. Goshima, A. Masuda, and K. Owaribe. Insulin-induced formation of ruffling membranes of KB cells and its correlation with enhancement of amino acid transport. Journal of Cell Biology, 98(3):801–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/3/801.

Green:1981:GDD

[GMRB81] R. F. Green, H. K. Meiss, and E. Rodriguez-Boulan. Glycosylation does not determine segregation of viral envelope proteins in the plasma membrane of epithelial cells. Journal of Cell Biology, 89(2):230–??, May 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/89/2/230.

Gotlieb:1981:DMO

[GMSK81] A. I. Gotlieb, L. M. May, L. Subrahmanyan, and V. I. Kalnins. Distribution of microtubule organizing centers in migrating sheets of endothelial cells. Journal of Cell Biology, 91(2): 589–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/91/2/589.

Gilmartin:1984:DRI

[GMVF84] M. E. Gilmartin, J. Mitchell, A. Vidrich, and I. M. Freedberg. Dual regulation of intermediate filament phosphorylation. Journal of Cell Biology, 98(3):1144–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/3/1144.

Gonnella:1981:PAM

[GN81] P. A. Gonnella and V. T. Nachmias. Platelet activation and microfilament bundling. Journal of Cell Biology, 89(1):146–??, April 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
89/1/146.

Gonnella:1984:MBF

[GN84] P. A. Gonnella and M. R. Neutra. Membrane-bound and fluidphase macromolecules enter separate prelysosomal compartments in absorptive cells of suckling rat ileum. Journal of Cell Biology, 99(3):909–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/99/3/909.

Gonzalez-Noriega:1980:CIL

[GNGTS80] A. Gonzalez-Noriega, J. H. Grubb, V. Talkad, and W. S. Sly. Chloroquine inhibits lysosomal enzyme pinocytosis and enhances lysosomal enzyme secretion by impairing receptor recycling. Journal of Cell Biology, 85(3):839–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/85/3/839.

Godfrey:1984:CTE

[GNW^{+}84] E. W. Godfrey, R. M. Nitkin, B. G. Wallace, L. L. Rubin, and U. J. McMahan. Components of Torpedo electric organ and muscle that cause aggregation of acetylcholine receptors on cultured muscle cells. Journal of Cell Biology, 99(2):615–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/2/615.

Gershoni:1980:CCI

[GO80] J. M. Gershoni and I. Ohad. Chloroplast-cytoplasmic interrelations involved in chloroplast development in Chlamydomonas reinhardi y-1: effect of selective depletion of chloroplast translates. Journal of Cell Biology, 87(1):124–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/87/1/124.

Grover:1983:ELF

[GOA83] A. Grover, R. G. Oshima, and E. D. Adamson. Epithelial layer formation in differentiating aggregates of F9 embryonal carcinoma cells. Journal of Cell Biology, 96(6):1690–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/6/
1690.

Gerace:1982:IMP

[GOKK82] L. Gerace, Y. Ottaviano, and C. Kondor-Koch. Identification of a major polypeptide of the nuclear pore complex. Journal of Cell Biology, 95(3):826–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/95/3/826.

Goldberg:1982:BSTa

[Gol82] B. D. Goldberg. Binding of soluble type I collagen to fibroblasts: effects of thermal activation of ligand, ligand concentration, pinocytosis, and cytoskeletal modifiers. Journal of Cell Biology, 95(3):747–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/95/3/747.

Goodenough:1983:MDE

[Goo83] U. W. Goodenough. Motile detergent-extracted cells of Tetrahymena and Chlamydomonas. Journal of Cell Biology, 96(6):1610–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/6/1610.

Gillies:1982:NMR

[GOSW82] R. J. Gillies, T. Ogino, R. G. Shulman, and D. C. Ward. 31p nuclear magnetic resonance evidence for the regulation of intracellular pH by Ehrlich ascites tumor cells. Journal of Cell Biology, 95(1):24–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/95/1/24.

Griffin:1981:IAT

[GPDM81] J. W. Griffin, D. L. Price, D. B. Drachman, and J. Morris. Incorporation of axonally transported glycoproteins into axolemma during nerve regeneration. Journal of Cell Biology, 88 (1):205–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/88/1/205.

Gieselmann:1983:BTC

[GPHV83] V. Gieselmann, R. Pohlmann, A. Hasilik, and K. Von Figura. Biosynthesis and transport of cathepsin D in cultured human fibroblasts. Journal of Cell Biology, 97(1):1–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/97/1/1.

Granelli-Piperno:1983:PAP

[GPR83] A. Granelli-Piperno and E. Reich. Plasminogen activators of the pituitary gland: enzyme characterization and hormonal modulation. Journal of Cell Biology, 97(4):1029–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/4/
1029.

Griffiths:1983:DGC

[GQW83] G. Griffiths, P. Quinn, and G. Warren. Dissection of the Golgi complex. I. Monensin inhibits the transport of viral membrane proteins from medial to trans Golgi cisternae in baby hamster kidney cells infected with Semliki Forest virus. Journal of Cell Biology, 96(3):835–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/3/835.

Garavito:1980:TDC

[GR80] R. M. Garavito and J. P. Rosenbusch. Three-dimensional crystals of an integral membrane protein: an initial X-ray analysis. Journal of Cell Biology, 86(1):327–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/86/1/327.

Gulley:1981:COP

[GR81] R. L. Gulley and T. S. Reese. Cytoskeletal organization at the postsynaptic complex. Journal of Cell Biology, 91(1):

298–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/91/1/298.

Geiger:1982:SRM

[GRB82] B. Geiger, D. Rosen, and G. Berke. Spatial relationships of microtubule-organizing centers and the contact area of cytotoxic T lymphocytes and target cells. Journal of Cell Biology, 95(1):137–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/95/1/137.

Gupta:1982:LCG

[GRDR82] P. Gupta, J. M. Rosen, P. D’Eustachio, and F. H. Ruddle. Localization of the casein gene family to a single mouse chromosome. Journal of Cell Biology, 93(1):199–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/93/1/199.

Grinnell:1980:FRC

[Gri80] F. Grinnell. Fibroblast receptor for cell-substratum adhesion: studies on the interaction of baby hamster kidney cells with latex beads coated by cold insoluble globulin (plasma fibronectin). Journal of Cell Biology, 86(1):104–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/86/1/104.

Granger:1982:SVC

[GRL82] B. L. Granger, E. A. Repasky, and E. Lazarides. Synemin and vimentin are components of intermediate filaments in avian erythrocytes. Journal of Cell Biology, 92(2):299–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/92/2/
299.

Gasc:1984:PRC

[GRR^{+}84] J. M. Gasc, J. M. Renoir, C. Radanyi, I. Joab, P. Tuohimaa, and E. E. Baulieu. Progesterone receptor in the chick oviduct: an immunohistochemical study with antibodies to distinct receptor components. Journal of Cell Biology, 99(4): 1193–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/4/1193.

Gulati:1983:CBM

[GRZ83] A. K. Gulati, A. H. Reddi, and A. A. Zalewski. Changes in the basement membrane zone components during skeletal muscle fiber degeneration and regeneration. Journal of Cell Biology, 97(4):957–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/4/957.

Giddings:1980:RBS

[GS80a] T. H. Giddings and L. A. Staehelin. Ribosome binding sites visualized on freeze-fractured membranes of the rough endoplasmic reticulum. Journal of Cell Biology, 85(1):147–??, April 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/85/1/
147.

Guilbert:1980:SIM

[GS80b] L. J. Guilbert and E. R. Stanley. Specific interaction of murine colony-stimulating factor with mononuclear phagocytic cells. Journal of Cell Biology, 85(1):153–??, April 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/85/1/153.

Gorbsky:1981:IIG

[GS81] G. Gorbsky and M. S. Steinberg. Isolation of the intercellular glycoproteins of desmosomes. Journal of Cell Biology, 90 (1):243–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/1/243.

Gilligan:1983:SIS

[GS83] D. M. Gilligan and B. H. Satir. Stimulation and inhibition of secretion in Paramecium: role of divalent cations. Journal of Cell Biology, 97(1):224–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/97/1/224.

Garofalo:1984:PSG

[GS84a] R. S. Garofalo and B. H. Satir. Paramecium secretory granule content: quantitative studies on in vitro expansion and its regulation by calcium and pH. Journal of Cell Biology, 99(6): 2193–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/6/2193.

Gilbert:1984:BTF

[GS84b] S. P. Gilbert and R. D. Sloboda. Bidirectional transport of fluorescently labeled vesicles introduced into extruded axoplasm of squid Loligo pealei. Journal of Cell Biology, 99(2):445–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/2/445.

Green:1984:GJC

[GS84c] C. R. Green and N. J. Severs. Gap junction connexon configuration in rapidly frozen myocardium and isolated intercalated disks. Journal of Cell Biology, 99(2):453–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/99/2/453.

Green:1984:SDB

[GS84d] R. Green and D. Shields. Somatostatin discriminates between the intracellular pathways of secretory and membrane proteins. Journal of Cell Biology, 99(1):97–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/1/97.

Gundersen:1984:SSP

[GS84e] G. G. Gundersen and B. M. Shapiro. Sperm surface proteins persist after fertilization. Journal of Cell Biology, 99(4): 1343–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/4/1343.

Gupta:1984:RDR

[GS84f] P. K. Gupta and M. A. Sirover. Regulation of DNA repair in serum-stimulated xeroderma pigmentosum cells. Journal of Cell Biology, 99(4):1275–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/4/1275.

Gobel:1981:FSA

[GSB81] U. Göbel, V. Speth, and W. Bredt. Filamentous structures in adherent Mycoplasma pneumoniae cells treated with nonionic detergents. Journal of Cell Biology, 91(2):537–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/91/2/
537.

Gibson:1982:EMM

[GSG82] G. J. Gibson, S. L. Schor, and M. E. Grant. Effects of matrix macromolecules on chondrocyte gene expression: synthesis of a low molecular weight collagen species by cells cultured within collagen gels. Journal of Cell Biology, 93(3):767–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/93/3/767.

Goodman:1980:EGF

[GSH80] R. Goodman, E. Slater, and H. R. Herschman. Epidermal growth factor induces tyrosine hydroxylase in a clonal pheochromocytoma cell line, PC–G2. Journal of Cell Biology, 84(3):495–??, March 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/84/3/495.

Gonatas:1984:EGV

[GSH^{+}84] N. K. Gonatas, A. Stieber, W. F. Hickey, S. H. Herbert, and J. O. Gonatas. Endosomes and Golgi vesicles in adsorptive and fluid phase endocytosis. Journal of Cell Biology, 99(4): 1379–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/4/1379.

Gotlieb:1983:MOC

[GSK83] A. I. Gotlieb, L. Subrahmanyan, and V. I. Kalnins. Microtubule-organizing centers and cell migration: effect of inhibition of migration and microtubule disruption in endothelial cells. Journal of Cell Biology, 96(5):1266–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/5/
1266.

Galvin:1984:ACA

[GSMHF84] N. J. Galvin, D. Stockhausen, B. L. Meyers-Hutchins, and W. A. Frazier. Association of the cyclic AMP chemotaxis receptor with the detergent-insoluble cytoskeleton of Dictyostelium discoideum. Journal of Cell Biology, 98(2):584– ??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/2/584.

Grabel:1983:TSC

[GSMR83] L. B. Grabel, M. S. Singer, G. R. Martin, and S. D. Rosen. Teratocarcinoma stem cell adhesion: the role of divalent cations and a cell surface lectin. Journal of Cell Biology, 96(6): 1532–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/6/1532.

Gonatas:1980:PIF

[GSOG80] J. Gonatas, A. Steiber, S. Olsnes, and N. K. Gonatas. Pathways involved in fluid phase and adsorptive endocytosis in neuroblastoma. Journal of Cell Biology, 87(3):579–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/87/3/
579.

Gal:1982:DIL

[GSPS82] D. Gal, E. R. Simpson, J. C. Porter, and J. M. Snyder. Defective internalization of low density lipoprotein in epidermoid cervical cancer cells. Journal of Cell Biology, 92(3):597–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
92/3/597.

Geuze:1982:ILR

[GSS^{+}82a] H. J. Geuze, J. W. Slot, G. J. Strous, H. F. Lodish, and A. L. Schwartz. Immunocytochemical localization of the receptor for asialoglycoprotein in rat liver cells. Journal of Cell Biology, 92 (3):865–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/3/865.

Gorza:1982:MTF

[GSS82b] L. Gorza, S. Sartore, and S. Schiaffino. Myosin types and fiber types in cardiac muscle. II. Atrial myocardium. Journal of Cell Biology, 95(3):838–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/95/3/838.

Geuze:1984:ULM

[GSS^{+}84] H. J. Geuze, J. W. Slot, G. J. Strous, A. Hasilik, and K. Von Figura. Ultrastructural localization of the mannose 6-phosphate receptor in rat liver. Journal of Cell Biology, 98 (6):2047–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/6/2047.

Geuze:1981:UCG

[GSvdLS81] H. J. Geuze, J. W. Slot, P. A. van der Ley, and R. C. Scheffer. Use of colloidal gold particles in double-labeling immunoelectron microscopy of ultrathin frozen tissue sections. Journal of Cell Biology, 89(3):653–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/89/3/653.

Goldberg:1984:PCC

[GT84] R. L. Goldberg and B. P. Toole. Pericellular coat of chick embryo chondrocytes: structural role of hyaluronate. Journal of Cell Biology, 99(6):2114–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/6/2114.

Gown:1982:MAI

[GV82] A. M. Gown and A. M. Vogel. Monoclonal antibodies to intermediate filament proteins of human cells: unique and crossreacting antibodies. Journal of Cell Biology, 95(2):414–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
95/2/414.

Green:1981:MFU

[GVK81] K. J. Green, G. I. Viamontes, and D. L. Kirk. Mechanism of formation, ultrastructure, and function of the cytoplasmic bridge system during morphogenesis in Volvox. Journal of Cell Biology, 91(3):756–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/91/3/756.

Gitelman:1980:PCC

[GW80] S. E. Gitelman and G. B. Witman. Purification of calmodulin from Chlamydomonas: calmodulin occurs in cell bodies and flagella. Journal of Cell Biology, 87(3):764–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/87/3/
764.

Garbi:1982:BLF

[GW82] C. Garbi and S. H. Wollman. Basal lamina formation on thyroid epithelia in separated follicles in suspension culture. Journal of Cell Biology, 94(2):489–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/94/2/489.

Gilmore:1982:PTAb

[GWB82] R. Gilmore, P. Walter, and G. Blobel. Protein translocation across the endoplasmic reticulum. II. Isolation and characterization of the signal recognition particle receptor. Journal of Cell Biology, 95(2):470–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/95/2/470.

Gingrich:1982:RSCb

[GWG82] J. C. Gingrich, R. C. Williams, and A. N. Glazer. Rod substructure in cyanobacterial phycobilisomes: phycoerythrin assembly in synechocystis 6701 phycobilisomes. Journal of Cell Biology, 95(1):170–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/95/1/170.

Gerhart:1984:CCD

[GWK84] J. Gerhart, M. Wu, and M. Kirschner. Cell cycle dynamics of an M-phase-specific cytoplasmic factor in Xenopus laevis oocytes and eggs. Journal of Cell Biology, 98(4):1247–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/4/
1247.

Gordon:1982:ILC

[GWP82] R. E. Gordon, K. B. Williams, and S. Puszkin. Immune localization of calmodulin in the ciliated cells of hamster tracheal epithelium. Journal of Cell Biology, 95(1):57–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
95/1/57.

Griffiths:1984:DNS

[GWQ^{+}84] G. Griffiths, G. Warren, P. Quinn, O. Mathieu-Costello, and H. Hoppeler. Density of newly synthesized plasma membrane proteins in intracellular membranes. I. Stereological studies. Journal of Cell Biology, 98(6):2133–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/6/2133.

Giffard:1984:CDB

[GWS84] R. G. Giffard, A. G. Weeds, and J. A. Spudich. Ca^{2+}- dependent binding of severin to actin: a one-to-one complex is formed. Journal of Cell Biology, 98(5):1796–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/5/
1796.

Ginsberg:1983:BFA

[GWWP83] M. H. Ginsberg, J. D. Wencel, J. G. White, and E. F. Plow. Binding of fibronectin to alpha-granule-deficient platelets. Journal of Cell Biology, 97(2):571–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/2/571.

Giannattasio:1980:MOP

[GZR^{+}80] G. Giannattasio, A. Zanini, P. Rosa, J. Meldolesi, R. K. Margolis, and R. U. Margolis. Molecular organization of prolactin granules. III. Intracellular transport of sulfated glycosaminoglycans and glycoproteins of the bovine prolactin granule matrix. Journal of Cell Biology, 86(1):273–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/86/1/273.

Howard:1980:CMF

[HA80]  R. J. Howard and J. R. Aist.  Cytoplasmic micro-

tubules and fungal morphogenesis: ultrastructural effects of methyl benzimidazole-2-ylcarbamate determined by freezesubstitution of hyphal tip cells. Journal of Cell Biology, 87 (1):55–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/87/1/55.

Hayden:1984:DSM

[HA84a] J. H. Hayden and R. D. Allen. Detection of single microtubules in living cells: particle transport can occur in both directions along the same microtubule. Journal of Cell Biology, 99(5): 1785–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/5/1785.

Herman:1984:TLV

[HA84b] B. Herman and D. F. Albertini. A time-lapse video image intensification analysis of cytoplasmic organelle movements during endosome translocation. Journal of Cell Biology, 98(2): 565–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/2/565.

Handley:1981:HBI

[HAEC81] D. A. Handley, C. M. Arbeeny, H. A. Eder, and S. Chien. Hepatic binding and internalization of low density lipoprotein-gold conjugates in rats treated with 17 alpha-ethinylestradiol. Journal of Cell Biology, 90(3):778–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/90/3/778.

Haugli:1982:DRP

[HAF82] F. Haugli, R. Andreassen, and S. Funderud. DNA replication in Physarum polycephalum: electron microscopic analysis of patterns of DNA replication in the presence of cycloheximide. Journal of Cell Biology, 95(1):323–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/95/1/323.

Hertzberg:1982:CAM

[HAFG82] E. L. Hertzberg, D. J. Anderson, M. Friedlander, and N. B. Gilula. Comparative analysis of the major polypeptides from liver gap junctions and lens fiber junctions. Journal of Cell Biology, 92(1):53–??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/92/1/53.

Herlands:1982:TRH

[HAP82] L. Herlands, V. G. Allfrey, and D. Poccia. Translational regulation of histone synthesis in the sea urchin strongylocentrotus purpuratus. Journal of Cell Biology, 94(1):219–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/94/1/219.

Hartwell:1980:MSC

[Har80a] L. H. Hartwell. Mutants of Saccharomyces cerevisiae unresponsive to cell division control by polypeptide mating hormone. Journal of Cell Biology, 85(3):811–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/85/3/811.

Hartzell:1980:DMA

[Har80b] H. C. Hartzell. Distribution of muscarinic acetylcholine receptors and presynaptic nerve terminals in amphibian heart. Journal of Cell Biology, 86(1):6–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/86/1/6.

Howard:1981:MAA

[HAS^{+}81] G. C. Howard, S. M. Abmayr, L. A. Shinefeld, V. L. Sato, and S. C. Elgin. Monoclonal antibodies against a specific nonhistone chromosomal protein of Drosophila associated with active genes. Journal of Cell Biology, 88(1):219–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/88/1/219.

Hatten:1981:CAP

[Hat81] M. E. Hatten. Cell assembly patterns of embryonic mouse cerebellar cells on carbohydrate-derivatized polylysine culture substrata. Journal of Cell Biology, 89(1):54–??, April 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/89/1/54.

Hay:1981:EM

[Hay81] E. D. Hay. Extracellular matrix. Journal of Cell Biology, 91 (3):??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/91/3/205s.

Hohman:1984:HSC

[HB84] T. C. Hohman and B. Bowers. Hydrolase secretion is a consequence of membrane recycling. Journal of Cell Biology, 98 (1):246–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/1/246.

Hoshina:1982:CLC

[HBB82] M. Hoshina, M. Boothby, and I. Boime. Cytological localization of chorionic gonadotropin alpha and placental lactogen mRNAs during development of the human placenta. Journal of Cell Biology, 93(1):190–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/93/1/190.

Heathcote:1984:BSMb

[HBO84] J. G. Heathcote, R. R. Bruns, and R. W. Orkin. Biosynthesis of sulphated macromolecules by rabbit lens epithelium. II. Relationship to basement membrane formation. Journal of Cell Biology, 99(3):861–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/3/861.

Heath:1981:AFI

[HBR81] J. Heath, S. Bell, and A. R. Rees. Appearance of functional insulin receptors during the differentiation of embryonal carcinoma cells. Journal of Cell Biology, 91(1):293–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/91/1/
293.

Hesketh:1984:SRH

[HBR^{+}84] T. R. Hesketh, M. A. Beaven, J. Rogers, B. Burke, and G. B. Warren. Stimulated release of histamine by a rat mast cell line is inhibited during mitosis. Journal of Cell Biology, 98 (6):2250–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/6/2250.

Huttner:1981:DMG

[HBS^{+}81] K. M. Huttner, J. A. Barbosa, G. A. Scangos, D. D. Pratcheva, and F. H. Ruddle. DNA-mediated gene transfer without carrier DNA. Journal of Cell Biology, 91(1):153–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/91/1/153.

Hays:1982:RAH

[HBS^{+}82] R. M. Hays, J. Bourguet, B. H. Satir, N. Franki, and J. Rapoport. Retention of antidiuretic hormone-induced particle aggregates by luminal membranes separated from toad bladder epithelial cells. Journal of Cell Biology, 92(1):237– ??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/1/237.

Hsieh:1983:BCS

[HC83] P. Hsieh and L. B. Chen. Behavior of cells seeded in isolated fibronectin matrices. Journal of Cell Biology, 96(5):1208–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
96/5/1208.

Havercroft:1984:PEB

[HC84] J. C. Havercroft and D. W. Cleveland. Programmed expression of beta-tubulin genes during development and differentiation of the chicken. Journal of Cell Biology, 99(6):1927–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/6/
1927.

Horwitz:1981:TIP

[HCH^{+}81] S. B. Horwitz, G. H. Chia, C. Harracksingh, S. Orlow, S. Pifko-Hirst, J. Schneck, L. Sorbara, M. Speaker, E. W. Wilk, and O. M. Rosen. Trifluoperazine inhibits phagocytosis in a macrophagelike cultured cell line. Journal of Cell Biology, 91(3):798–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/91/3/798.

Hedman:1983:CSP

[HCJV83] K. Hedman, J. Christner, I. Julkunen, and A. Vaheri. Chondroitin sulfate at the plasma membranes of cultured fibroblasts. Journal of Cell Biology, 97(4):1288–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/4/

Hooper:1980:ASV

[HCK80] J. E. Hooper, S. S. Carlson, and R. B. Kelly. Antibodies to synaptic vesicles purified from Narcine electric organ bind a subclass of mammalian nerve terminals. Journal of Cell Biology, 87(1):104–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/87/1/104.

Hollingshead:1981:CPS

[HCMK81] C. J. Hollingshead, D. L. D. Caspar, V. Melchior, and D. A. Kirschner. Compaction and particle segregation in myelin membrane arrays. Journal of Cell Biology, 89(3):631–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/89/3/
631.

Herman:1981:RBC

[HCP81] I. M. Herman, N. J. Crisona, and T. D. Pollard. Relation between cell activity and the distribution of cytoplasmic actin and myosin. Journal of Cell Biology, 90(1):84–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/90/1/84.

Holbrook:1982:AEK

[HDB82] K. A. Holbrook, B. A. Dale, and K. S. Brown. Abnormal Epidermal Keratinization in the repeated epilation mutant mouse. Journal of Cell Biology, 92(2):387–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/92/2/387.

Hoover:1983:RCC

[HDRK83] R. L. Hoover, E. A. Dawidowicz, J. M. Robinson, and M. J. Karnovsky. Role of cholesterol in the capping of surface immunoglobulin receptors on murine lymphocytes. Journal of Cell Biology, 97(1):73–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/1/73.

Hootman:1981:CMA

[HE81] S. R. Hootman and S. A. Ernst. Characterization of muscarinic acetylcholine receptors in the avian salt gland. Journal of Cell Biology, 91(3):781–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/91/3/781.

Heath:1980:BKD

[Hea80] I. B. Heath. Behavior of kinetochores during mitosis in the fungus Saprolegnia ferax. Journal of Cell Biology, 84(3):531–??, March 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
84/3/531.

Hayman:1982:CAR

[HEA^{+}82] E. G. Hayman, E. Engvall, E. A’Hearn, D. Barnes, M. Pierschbacher, and E. Ruoslahti. Cell attachment on replicas of SDS polyacrylamide gels reveals two adhesive plasma proteins. Journal of Cell Biology, 95(1):20–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/95/1/20.

Huff:1981:NGF

[HEG81] K. Huff, D. End, and G. Guroff. Nerve growth factor-induced alteration in the response of PC12 pheochromocytoma cells to epidermal growth factor. Journal of Cell Biology, 88(1): 189–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/88/1/189.

Herzlinger:1982:MEC

[HEO82] D. A. Herzlinger, T. G. Easton, and G. K. Ojakian. The MDCK epithelial cell line expresses a cell surface antigen of the kidney distal tubule. Journal of Cell Biology, 93(2):269–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
93/2/269.

Hepler:1980:MMA

[Hep80] P. K. Hepler. Membranes in the miotic apparatus of barley cells. Journal of Cell Biology, 86(2):490–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/86/2/490.

Herth:1980:CWC

[Her80] W. Herth. Calcofluor white and Congo red inhibit chitin microfibril assembly of Poterioochromonas: evidence for a gap between polymerization and microfibril formation. Journal of Cell Biology, 87(2):442–??, November 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/87/2/442.

Hayman:1981:CLC

[HER81] E. G. Hayman, E. Engvall, and E. Ruoslahti. Concomitant loss of cell surface fibronectin and laminin from transformed rat kidney cells. Journal of Cell Biology, 88(2):352–??, February 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/88/2/
352.

Herzog:1983:TTF

[Her83] V. Herzog. Transcytosis in thyroid follicle cells. Journal of Cell Biology, 97(3):607–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/97/3/607.

Heuser:1980:TDV

[Heu80] J. Heuser. Three-dimensional visualization of coated vesicle formation in fibroblasts. Journal of Cell Biology, 84(3):560–??, March 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
84/3/560.

Higashi-Fujime:1980:AMV

[HF80] S. Higashi-Fujime. Active movement in vitro of bundle of microfilaments isolated from Nitella cell. Journal of Cell Biology, 87(3):569–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/87/3/569.

Hyman:1983:DAR

[HF83] C. Hyman and S. C. Froehner. Degradation of acetylcholine receptors in muscle cells: effect of leupeptin on turnover rate, intracellular pool sizes, and receptor properties. Journal of Cell Biology, 96(5):1316–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/5/1316.

Haga:1982:MCT

[HFSK82] N. Haga, M. Forte, Y. Saimi, and C. Kung. Microinjection of cytoplasm as a test of complementation in Paramecium.

Journal of Cell Biology, 92(2):559–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/92/2/559.

Hoffstein:1982:DMA

[HFW82] S. T. Hoffstein, R. S. Friedman, and G. Weissmann. Degranulation, membrane addition, and shape change during chemotactic factor-induced aggregation of human neutrophils. Journal of Cell Biology, 95(1):234–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/95/1/234.

Hall:1980:BTC

[HG80a] W. M. Hall and P. Ganguly. Binding of thrombin to cultured human fibroblasts: evidence for receptor modulation. Journal of Cell Biology, 87(3):601–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/87/3/601.

Hall:1980:DET

[HG80b] W. M. Hall and P. Ganguly. Differential effect of thrombin on the growth of human fibroblasts. Journal of Cell Biology, 85(1):70–??, April 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/85/1/70.

Hoffman:1980:EDF

[HG80c] J. L. Hoffman and U. W. Goodenough. Experimental dissection of flagellar surface motility in chlamydomonas. Journal of Cell Biology, 86(2):656–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/86/2/656.

Hoffman:1984:CAC

[HGP84] P. N. Hoffman, J. W. Griffin, and D. L. Price. Control of axonal caliber by neurofilament transport. Journal of Cell Biology, 99 (2):705–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/2/705.

Hirokawa:1984:OMN

[HGW84] N. Hirokawa, M. A. Glicksman, and M. B. Willard. Organization of mammalian neurofilament polypeptides within the neuronal cytoskeleton. Journal of Cell Biology, 98(4):1523–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
98/4/1523.

Hirokawa:1981:QFD

[HH81a] N. Hirokawa and J. E. Heuser. Quick-freeze, deep-etch visualization of the cytoskeleton beneath surface differentiations of intestinal epithelial cells. Journal of Cell Biology, 91(2): 399–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/91/2/399.

Hirokawa:1981:SEB

[HH81b] N. Hirokawa and J. E. Heuser. Structural evidence that botulinum toxin blocks neuromuscular transmission by impairing the calcium influx that normally accompanies nerve depolarization. Journal of Cell Biology, 88(1):160–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/88/1/160.

Hampton:1983:ECO

[HH83a] R. Y. Hampton and R. W. Holz. Effects of changes in osmolality on the stability and function of cultured chromaffin cells and the possible role of osmotic forces in exocytosis. Journal of Cell Biology, 96(4):1082–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/4/1082.

Henry:1983:EUP

[HH83b] S. M. Henry and L. D. Hodge. Evidence for a unique profile of phosphatidylcholine synthesis in late mitotic cells. Journal of Cell Biology, 97(1):166–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/97/1/166.

Huang:1983:HPD

[HHD83] J. S. Huang, S. S. Huang, and T. F. Deuel. Human plateletderived growth factor: radioimmunoassay and discovery of a specific plasma-binding protein. Journal of Cell Biology, 97 (2):383–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/2/383.

Hoshina:1983:CDC

[HHPB83] M. Hoshina, R. Hussa, R. Pattillo, and I. Boime. Cytological distribution of chorionic gonadotropin subunit and placental lactogen messenger RNA in neoplasms derived from human placenta. Journal of Cell Biology, 97(4):1200–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/4/
1200.

Hiramoto:1981:QSPb

[HHS^{+}81] Y. Hiramoto, Y. Hamaguchi, Y. Shóji, T. E. Schroeder, S. Shimoda, and S. Nakamura. Quantitative studies on the polarization optical properties of living cells II. The role of microtubules in birefringence of the spindle of the sea urchin egg. Journal of Cell Biology, 89(1):121–??, April 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/89/1/121.

Harding:1983:RME

[HHS83] C. Harding, J. Heuser, and P. Stahl. Receptor-mediated endocytosis of transferrin and recycling of the transferrin receptor in rat reticulocytes. Journal of Cell Biology, 97(2):329–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
97/2/329.

Hunziker:1984:CUAa

[HHS^{+}84] E. B. Hunziker, W. Herrmann, R. K. Schenk, M. Mueller, and H. Moor. Cartilage ultrastructure after high pressure freezing, freeze substitution, and low temperature embedding. I. Chondrocyte ultrastructure–implications for the theories of mineralization and vascular invasion. Journal of Cell Biology, 98 (1):267–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/1/267.

Hiramoto:1981:QSPa

[HHSS81] Y. Hiramoto, Y. Hamaguchi, Y. Shôji, and S. Shimoda. Quantitative studies on the polarization optical properties of living cells. I. Microphotometric birefringence detection system. Journal of Cell Biology, 89(1):115–??, April 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/89/1/115.

Hiwatashi:1980:RME

[HHT80] K. Hiwatashi, N. Haga, and M. Takahashi. Restoration of membrane excitability in a behavioral mutant of Paramecium caudatum during conjugation and by microinjection of wildtype cytoplasm. Journal of Cell Biology, 84(2):476–??, February 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
84/2/476.

Heidemann:1984:SOA

[HHT^{+}84] S. R. Heidemann, M. A. Hamborg, S. J. Thomas, B. Song, S. Lindley, and D. Chu. Spatial organization of axonal microtubules. Journal of Cell Biology, 99(4):1289–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/4/
1289.

Hirokawa:1982:CLS

[Hir82] N. Hirokawa. Cross-linker system between neurofilaments, microtubules and membranous organelles in frog axons revealed by the quick-freeze, deep-etching method. Journal of Cell Biology, 94(1):129–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/94/1/129.

Harper:1980:ICC

[HJ80] P. A. Harper and R. L. Juliano. Isolation and characterization of Chinese hamster ovary cell variants defective in adhesion to fibronectin-coated collagen. Journal of Cell Biology, 87(3): 755–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/87/3/755.

Harper:1981:FIA

[HJ81] P. A. Harper and R. L. Juliano. Fibronectin-independent adhesion of fibroblasts to the extracellular matrix: mediation by a high molecular weight membrane glycoprotein. Journal of Cell Biology, 91(3):647–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/91/3/647.

Heuser:1980:FOR

[HK80a] J. E. Heuser and M. W. Kirschner. Filament organization revealed in platinum replicas of freeze-dried cytoskeletons. Journal of Cell Biology, 86(1):212–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/86/1/212.

Hsiung:1980:HGE

[HK80b] N. Hsiung and R. Kucherlapati. Histone gene expression and chromatin structure in mammalian cell hybrids. Journal of Cell Biology, 87(1):227–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/87/1/227.

Hirokawa:1983:MBB

[HKCM83] N. Hirokawa, T. C. Keller, R. Chasan, and M. S. Mooseker. Mechanism of brush border contractility studied by the quickfreeze, deep-etch method. Journal of Cell Biology, 96(5): 1325–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/5/1325.

Helenius:1980:ESF

[HKSF80] Ari Helenius, J. Kartenbeck, K. Simons, and E. Fries. On the entry of semliki forest virus into BHK-21 cells. Journal of Cell Biology, 84(2):404–??, February 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/84/2/404.

Hatten:1981:ACP

[HL81] M. E. Hatten and R. K. Liem. Astroglial cells provide a template for the positioning of developing cerebellar neurons in vitro. Journal of Cell Biology, 90(3):622–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/90/3/622.

Heidemann:1981:POA

[HLH81] S. R. Heidemann, J. M. Landers, and M. A. Hamborg. Polarity orientation of axonal microtubules. Journal of Cell Biology, 91(3):661–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/91/3/661.

Hatten:1984:TFC

[HLM84a] M. E. Hatten, R. K. Liem, and C. A. Mason. Two forms of cerebellar glial cells interact differently with neurons in vitro. Journal of Cell Biology, 98(1):193–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/1/193.

Hochman:1984:CAR

[HLM^{+}84b] J. Hochman, E. Levy, N. Mador, M. M. Gottesman, G. M. Shearer, and E. Okon. Cell adhesiveness is related to tumorigenicity in malignant lymphoid cells. Journal of Cell Biology, 99(4):1282–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/4/1282.

Hall:1981:CAP

[HLS81] Z. W. Hall, B. W. Lubit, and J. H. Schwartz. Cytoplasmic actin in postsynaptic structures at the neuromuscular junction. Journal of Cell Biology, 90(3):789–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/90/3/789.

Hutchison:1982:SHE

[HLSWH82] N. J. Hutchison, P. R. Langer-Safer, D. C. Ward, and B. A. Hamkalo. In situ hybridization at the electron microscope level: hybrid detection by autoradiography and colloidal gold. Journal of Cell Biology, 95(2):609–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/95/2/609.

Howe:1983:CKA

[HM83a] C. L. Howe and M. S. Mooseker. Characterization of the 110-kdalton actin-calmodulin-, and membrane-binding protein from microvilli of intestinal epithelial cells. Journal of Cell Biology, 97(4):974–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/4/974.

Hubbard:1983:IRHb

[HM83b] A. L. Hubbard and A. Ma. Isolation of rat hepatocyte plasma membranes. II. Identification of membrane-associated cytoskeletal proteins. Journal of Cell Biology, 96(1):230– ??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/1/230.

Horowitz:1984:SPG

[HM84a] S. B. Horowitz and D. S. Miller. Solvent properties of ground substance studied by cryomicrodissection and intracellular reference-phase techniques. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/1/172s.

Horwitz:1984:LPI

[HM84b] M. A. Horwitz and F. R. Maxfield. Legionella pneumophila inhibits acidification of its phagosome in human monocytes. Journal of Cell Biology, 99(6):1936–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/6/
1936.

Hosoya:1984:MWP

[HM84c] H. Hosoya and I. Mabuchi. A 45,000-mol-wt protein-actin complex from unfertilized sea urchin egg affects assembly properties of actin. Journal of Cell Biology, 99(3):994–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/3/
994.

Howard:1984:CPM

[HM84d] T. H. Howard and W. H. Meyer. Chemotactic peptide modulation of actin assembly and locomotion in neutrophils. Journal of Cell Biology, 98(4):1265–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/4/1265.

Huber:1984:ILM

[HM84e] G. Huber and A. Matus. Immunocytochemical localization of microtubule-associated protein 1 in rat cerebellum using monoclonal antibodies. Journal of Cell Biology, 98(2):777– ??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/2/777.

Hill:1982:NRS

[HMB^{+}82] R. J. Hill, M. R. Mott, E. J. Burnett, S. M. Abmayr, K. Lowenhaupt, and S. C. Elgin. Nucleosome repeat structure is present in native salivary chromosomes of Drosophila melanogaster. Journal of Cell Biology, 95(1):262–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/95/1/262.

Habener:1980:EEC

[HMDP80] J. F. Habener, R. Maunus, P. C. Dee, and J. T. Potts. Early events in the cellular formation of proparathyroid hormone. Journal of Cell Biology, 85(2):292–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/85/2/292.

Howe:1980:BBC

[HMG80] C. L. Howe, M. S. Mooseker, and T. A. Graves. Brushborder calmodulin. A major component of the isolated microvillus core. Journal of Cell Biology, 85(3):916–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/85/3/
916.

Hoober:1982:RAM

[HMKM82] J. K. Hoober, D. B. Marks, B. J. Keller, and M. M. Margulies. Regulation of accumulation of the major thylakoid polypeptides in Chlamydomonas reinhardtii y-1 at 25 degrees C and 38 degrees C. Journal of Cell Biology, 95(2):552–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/95/2/
552.

Hohman:1982:PLF

[HMM82] T. C. Hohman, P. L. McNeil, and L. Muscatine. Phagosomelysosome fusion inhibited by algal symbionts of Hydra viridis. Journal of Cell Biology, 94(1):56–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/94/1/56.

Hatzfeld:1982:ARE

[HMR82] J. Hatzfeld, R. Miskin, and E. Reich. Acetylcholine receptor: effects of proteolysis on receptor metabolism. Journal of Cell Biology, 92(1):176–??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/92/1/176.

Heathcote:1984:BSMa

[HO84a] J. G. Heathcote and R. W. Orkin. Biosynthesis of sulphated macromolecules by rabbit lens epithelium. I. Identification of the major macromolecules synthesized by lens epithelial cells in vitro. Journal of Cell Biology, 99(3):852–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/3/
852.

Herzlinger:1984:SDM

[HO84b] D. A. Herzlinger and G. K. Ojakian. Studies on the development and maintenance of epithelial cell surface polarity with monoclonal antibodies. Journal of Cell Biology, 98(5):1777–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
98/5/1777.

Hayman:1982:CHS

[HOMR82] E. G. Hayman, A. Oldberg, G. R. Martin, and E. Ruoslahti. Codistribution of heparan sulfate proteoglycan, laminin, and fibronectin in the extracellular matrix of normal rat kidney cells and their coordinate absence in transformed cells. Journal of Cell Biology, 94(1):28–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/94/1/28.

Hall:1984:ESS

[HOR84] D. J. Hall, J. J. O’Leary, and A. Rosenberg. Early synthesis of specific cytoplasm proteins is correlated with the rate of exit of lymphocytes from the resting state. Journal of Cell Biology, 99(5):1814–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/5/1814.

Hall:1981:RCR

[HOT81] P. F. Hall, S. Osawa, and C. L. Thomasson. A role for calmodulin in the regulation of steroidogenesis. Journal of Cell Biology, 90(2):402–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/90/2/402.

Harris:1980:DTC

[HOW80] P. Harris, M. Osborn, and K. Weber. Distribution of tubulincontaining structures in the egg of the sea urchin Strongylocentrotus purpuratus from fertilization through first cleavage. Journal of Cell Biology, 84(3):668–??, March 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/84/3/668.

Herman:1981:EML

[HP81] I. M. Herman and T. D. Pollard. Electron microscopic localization of cytoplasmic myosin with ferritin-labeled antibodies. Journal of Cell Biology, 88(2):346–??, February 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/88/2/346.

Howell:1982:HGF

[HP82a] K. E. Howell and G. E. Palade. Hepatic Golgi fractions resolved into membrane and content subfractions. Journal of Cell Biology, 92(3):822–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/92/3/822.

Howell:1982:HLP

[HP82b] K. E. Howell and G. E. Palade. Heterogeneity of lipoprotein particles in hepatic Golgi fractions. Journal of Cell Biology, 92 (3):833–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/3/833.

Hume:1983:ILM

[HPG83] D. A. Hume, V. H. Perry, and S. Gordon. Immunohistochemical localization of a macrophage-specific antigen in developing mouse retina: phagocytosis of dying neurons and differentiation of microglial cells to form a regular array in the plexiform layers. Journal of Cell Biology, 97(1):253–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/97/1/253.

Howard:1982:IPA

[HPM82] F. D. Howard, H. R. Petty, and H. M. McConnell. Identification of phagocytosis-associated surface proteins of macrophages by two-dimensional gel electrophoresis. Journal of Cell Biology, 92(2):283–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/92/2/283.

Halaban:1983:RTH

[HPM^{+}83] R. Halaban, S. H. Pomerantz, S. Marshall, D. T. Lambert, and A. B. Lerner. Regulation of tyrosinase in human melanocytes grown in culture. Journal of Cell Biology, 97(2):480–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
97/2/480.

Huang:1981:RSC

[HPRL81] B. Huang, G. Piperno, Z. Ramanis, and D. J. Luck. Radial spokes of Chlamydomonas flagella: genetic analysis of assembly and function. Journal of Cell Biology, 88(1):80–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/88/1/
80.

Hanks:1980:IDS

[HR80] S. K. Hanks and P. N. Rao. Initiation of DNA synthesis in the prematurely condensed chromosomes of G1 cells. Journal of Cell Biology, 87(1):285–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/87/1/285.

Haimo:1981:CFM

[HR81a] L. T. Haimo and J. L. Rosenbaum. Cilia, flagella, and microtubules. Journal of Cell Biology, 91(3):??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/91/3/
125s.

Heuser:1981:SCA

[HR81b] J. E. Heuser and T. S. Reese. Structural changes after transmitter release at the frog neuromuscular junction. Journal of Cell Biology, 88(3):564–??, March 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/88/3/564.

Hoessli:1980:LAA

[HRBJG80] D. Hoessli, E. Rungger-Brändle, B. M. Jockusch, and G. Gabbiani. Lymphocyte alpha-actinin. Relationship to cell membrane and co-capping with surface receptors. Journal of Cell Biology, 84(2):305–??, February 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/84/2/305.

Haslam:1980:PRN

[HS80a] S. Z. Haslam and G. Shyamala. Progesterone receptors in normal mammary gland: receptor modulations in relation to differentiation. Journal of Cell Biology, 86(3):730–??, September 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/86/3/
730.

Hsieh:1980:AIM

[HS80b] P. Hsieh and N. Sueoka. Antisera inhibiting mammalian cell spreading and possible cell surface antigens involved. Journal of Cell Biology, 86(3):866–??, September 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/86/3/866.

Howard:1981:CCC

[HS81] R. F. Howard and J. R. Sheppard. Cell cycle changes in the adenylate cyclase of C6 glioma cells. Journal of Cell Biology, 90(1):169–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/1/169.

Hertel:1983:RBA

[HS83] C. Hertel and M. Staehelin. Reappearance of beta-adrenergic receptors after isoproterenol treatment in intact C6-cells. Journal of Cell Biology, 97(5):1538–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/5/1538.

Hunziker:1984:CUAb

[HS84] E. B. Hunziker and R. K. Schenk. Cartilage ultrastructure after high pressure freezing, freeze substitution, and low temperature embedding. II. Intercellular matrix ultrastructurepreservation of proteoglycans in their native state. Journal of Cell Biology, 98(1):277–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/1/277.

Hammerschlag:1982:EAN

[HSB^{+}82] R. Hammerschlag, G. C. Stone, F. A. Bolen, J. D. Lindsey, and M. H. Ellisman. Evidence that all newly synthesized proteins destined for fast axonal transport pass through the Golgi apparatus. Journal of Cell Biology, 93(3):568–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/93/3/568.

Haeuptle:1983:ECS

[HSB^{+}83] M. Haeuptle, Y. L. M. Suard, E. Bogenmann, H. Reggio, L. Racine, and J. Kraehenbuhl. Effect of cell shape change on the function and differentiation of rabbit mammary cells in culture. Journal of Cell Biology, 96(5):1425–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/5/
1425.

Hsieh:1980:SFM

[HSC80] P. Hsieh, R. Segal, and L. B. Chen. Studies of fibronectin matrices in living cells with fluoresceinated gelatin. Journal of Cell Biology, 87(1):14–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/87/1/14.

Hollenbeck:1984:CDL

[HSC84] P. J. Hollenbeck, F. Suprynowicz, and W. Z. Cande. Cytoplasmic dynein-like ATPase cross-links microtubules in an ATPsensitive manner. Journal of Cell Biology, 99(4):1251–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/4/1251.

Hui:1980:ECL

[HSCS80] S. W. Hui, C. M. Stewart, M. P. Carpenter, and T. P. Stewart. Effects of cholesterol on lipid organization in human erythrocyte membrane. Journal of Cell Biology, 85(2):283–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/85/2/

Hafeman:1982:LMC

[HSFM82] D. G. Hafeman, L. M. Smith, D. T. Fearon, and H. M. Mc-Connell. Lipid monolayer-coated solid surfaces do not perturb the lateral motion and distribution of C3b receptors on neutrophils. Journal of Cell Biology, 94(1):224–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/94/1/224.

Huttner:1983:SPN

[HSGD83] W. B. Huttner, W. Schiebler, P. Greengard, and P. De Camilli. Synapsin I (protein I), a nerve terminal-specific phosphoprotein. III. Its association with synaptic vesicles studied in a highly purified synaptic vesicle preparation. Journal of Cell Biology, 96(5):1374–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/5/1374.

Hatfield:1984:GFA

[HSME84] J. S. Hatfield, R. P. Skoff, H. Maisel, and L. Eng. Glial fibrillary acidic protein is localized in the lens epithelium. Journal of Cell Biology, 98(5):1895–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/5/1895.

Hurkman:1981:SCM

[HSRL81] W. J. Hurkman, L. D. Smith, J. Richter, and B. A. Larkins. Subcellular compartmentalization of maize storage proteins in Xenopus oocytes injected with zein messenger RNAs. Journal of Cell Biology, 89(2):292–??, May 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/89/2/292.

Heltianu:1982:HRM

[HSS82] C. Heltianu, M. Simionescu, and N. Simionescu. Histamine receptors of the microvascular endothelium revealed in situ with a histamine-ferritin conjugate: characteristic high-affinity binding sites in venules. Journal of Cell Biology, 93(2):357–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
93/2/357.

Haga:1983:IIC

[HSTK83] N. Haga, Y. Saimi, M. Takahashi, and C. Kung. Intra- and interspecific complementation of membrane-inexcitable mutants of Paramecium. Journal of Cell Biology, 97(2):378–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/2/
378.

Haston:1982:LLA

[HSW82] W. S. Haston, J. M. Shields, and P. C. Wilkinson. Lymphocyte locomotion and attachment on two-dimensional surfaces and in three-dimensional matrices. Journal of Cell Biology, 92 (3):747–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/3/747.

Heiple:1980:IPS

[HT80] J. M. Heiple and D. L. Taylor. Intracellular pH in single motile cells. Journal of Cell Biology, 86(3):885–??, September 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/86/3/885.

Heiple:1982:PCP

[HT82a] J. M. Heiple and D. L. Taylor. pH changes in pinosomes and phagosomes in the ameba, Chaos carolinensis. Journal of Cell Biology, 94(1):143–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/94/1/143.

Hirokawa:1982:IBA

[HT82b] N. Hirokawa and L. G. Tilney. Interactions between actin filaments and between actin filaments and membranes in quickfrozen and deeply etched hair cells of the chick ear. Journal of Cell Biology, 95(1):249–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/95/1/249.

Hopkins:1983:IPT

[HT83] C. R. Hopkins and I. S. Trowbridge. Internalization and processing of transferrin and the transferrin receptor in human carcinoma A431 cells. Journal of Cell Biology, 97(2):508–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
97/2/508.

Hirokawa:1982:OAM

[HTFH82] N. Hirokawa, L. G. Tilney, K. Fujiwara, and J. E. Heuser. Organization of actin, myosin, and intermediate filaments in the brush border of intestinal epithelial cells. Journal of Cell Biology, 94(2):425–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/94/2/425.

Hickman:1984:PMI

[HTGB84] S. Hickman, J. L. Theodorakis, J. M. Greco, and P. H. Brown. Processing of MOPC 315 immunoglobulin A oligosaccharides: evidence for endoplasmic reticulum and trans Golgi alpha 1,2-mannosidase activity. Journal of Cell Biology, 98(2): 407–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/2/407.

Hogan:1982:SLT

[HTKC82] B. L. Hogan, A. Taylor, M. Kurkinen, and J. R. Couchman. Synthesis and localization of two sulphated glycoproteins associated with basement membranes and the extracellular matrix. Journal of Cell Biology, 95(1):197–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/95/1/197.

Hartwig:1980:ABP

[HTS80] J. H. Hartwig, J. Tyler, and T. P. Stossel. Actin-binding protein promotes the bipolar and perpendicular branching of actin filaments. Journal of Cell Biology, 87(3):841–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/87/3/
841.

Hoops:1983:ODH

[HW83] H. J. Hoops and G. B. Witman. Outer doublet heterogeneity reveals structural polarity related to beat direction in Chlamydomonas flagella. Journal of Cell Biology, 97(3): 902–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/3/902.

Harford:1983:MII

[HWAK83] J. Harford, A. W. Wolkoff, G. Ashwell, and R. D. Klausner. Monensin inhibits intracellular dissociation of asialoglycoproteins from their receptor. Journal of Cell Biology, 96 (6):1824–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/6/1824.

Haigler:1982:AVC

[HWBC82] C. H. Haigler, A. R. White, R. M. Brown, and K. M. Cooper. Alteration of in vivo cellulose ribbon assembly by carboxymethylcellulose and other cellulose derivatives. Journal of Cell Biology, 94(1):64–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/94/1/64.

Hoops:1984:FWR

[HWJW84] H. J. Hoops, R. L. Wright, J. W. Jarvik, and G. B. Witman. Flagellar waveform and rotational orientation in a Chlamydomonas mutant lacking normal striated fibers. Journal of Cell Biology, 98(3):818–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/3/818.

Haring:1984:AIB

[HWK^{+}84] H. U. Haring, M. F. White, C. R. Kahn, M. Kasuga, V. Lauris, R. Fleischmann, M. Murray, and J. Pawelek. Abnormality of insulin binding and receptor phosphorylation in an insulinresistant melanoma cell line. Journal of Cell Biology, 99(3): 900–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/3/900.

Hubbard:1983:IRHa

[HWM83] A. L. Hubbard, D. A. Wall, and A. Ma. Isolation of rat hepatocyte plasma membranes. I. Presence of the three major domains. Journal of Cell Biology, 96(1):217–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/96/1/217.

Hays:1982:TFK

[HWS82] T. S. Hays, D. Wise, and E. D. Salmon. Traction force on a kinetochore at metaphase acts as a linear function of kinetochore fiber length. Journal of Cell Biology, 93(2):374–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/93/2/
374.

Hynes:1982:FMM

[HY82] R. O. Hynes and K. M. Yamada. Fibronectins: multifunctional modular glycoproteins. Journal of Cell Biology, 95(2): 369–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/95/2/369.

Heidemann:1980:EMS

[HZM80] S. R. Heidemann, G. W. Zieve, and J. R. McIntosh. Evidence for microtubule subunit addition to the distal end of mitotic structures in vitro. Journal of Cell Biology, 87(1):152–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
87/1/152.

Ingraham:1981:FPL

[IBB^{+}81] L. M. Ingraham, C. P. Burns, L. A. Boxer, R. L. Baehner, and R. A. Haak. Fluidity properties and liquid composition of erythrocyte membranes in Chediak–Higashi syndrome. Journal of Cell Biology, 89(3):510–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/89/3/510.

Ip:1983:DDC

[IDF83] W. Ip, S. I. Danto, and D. A. Fischman. Detection of desmincontaining intermediate filaments in cultured muscle and nonmuscle cells by immunoelectron microscopy. Journal of Cell Biology, 96(2):401–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/2/401.

Ill:1984:APL

[IER84] C. R. Ill, E. Engvall, and E. Ruoslahti. Adhesion of platelets to laminin in the absence of activation. Journal of Cell Biology, 99(6):2140–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/6/2140.

Iwanij:1982:SCR

[IHJ82] V. Iwanij, B. E. Hull, and J. D. Jamieson. Structural characterization of a rat acinar cell tumor. Journal of Cell Biology, 95(3):727–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/95/3/727.

Iwanij:1982:BAS

[IJ82a] V. Iwanij and J. D. Jamieson. Biochemical analysis of secretory proteins synthesized by normal rat pancreas and by pancreatic acinar tumor cells. Journal of Cell Biology, 95(3): 734–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/95/3/734.

Iwanij:1982:CSP

[IJ82b] V. Iwanij and J. D. Jamieson. Comparison of secretory protein profiles in developing rat pancreatic rudiments and rat acinar tumor cells. Journal of Cell Biology, 95(3):742–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/95/3/
742.

Inoue:1983:URM

[ILL83] S. Inoué, C. P. Leblond, and G. W. Laurie. Ultrastructure of Reichert’s membrane, a multilayered basement membrane in the parietal wall of the rat yolk sac. Journal of Cell Biology, 97(5):1524–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/5/1524.

Ierardi:1983:SCI

[IMB83] L. A. Ierardi, S. B. Moss, and A. R. Bellvé. Synaptonemal complexes are integral components of the isolated mouse spermatocyte nuclear matrix. Journal of Cell Biology, 96(6):1717–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
96/6/1717.

Ishiguro:1982:ECD

[IMS82] K. Ishiguro, H. Murofushi, and H. Sakai. Evidence that cAMPdependent protein kinase and a protein factor are involved in reactivation of Triton X-100 models of sea urchin and starfish spermatozoa. Journal of Cell Biology, 92(3):777–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/92/3/
777.

Inoue:1981:CDM

[Ino81a] S. Inoué. Cell division and the mitotic spindle. Journal of Cell Biology, 91(3):??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/91/3/131s.

Inoue:1981:VIP

[Ino81b] S. Inoué. Video image processing greatly enhances contrast, quality, and speed in polarization-based microscopy. Journal of Cell Biology, 89(2):346–??, May 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/89/2/346.

Iozzo:1984:BHS

[Ioz84] R. V. Iozzo. Biosynthesis of heparan sulfate proteoglycan by human colon carcinoma cells and its localization at the cell surface. Journal of Cell Biology, 99(2):403–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/99/2/403.

Ibbotson:1984:ICO

[IRMM84] K. J. Ibbotson, G. D. Roodman, L. M. McManus, and G. R. Mundy. Identification and characterization of osteoclast-like cells and their progenitors in cultures of feline marrow mononuclear cells. Journal of Cell Biology, 99(2):471–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/99/2/471.

Isaacs:1981:ECD

[ISGS81] J. Isaacs, N. Savion, D. Gospodarowicz, and M. A. Shuman. Effect of cell density on thrombin binding to a specific site on bovine vascular endothelial cells. Journal of Cell Biology, 90(3):670–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/90/3/670.

Inoue:1982:ART

[IT82] S. Inoué and L. G. Tilney. Acrosomal reaction of thyone sperm. I. Changes in the sperm head visualized by high resolution video microscopy. Journal of Cell Biology, 93(3):812–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/93/3/
812.

Isom:1980:TIR

[ITT80] H. C. Isom, M. J. Tevethia, and J. M. Taylor. Transformation of isolated rat hepatocytes with simian virus 40. Journal of Cell Biology, 85(3):651–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/85/3/651.

Izant:1983:MAP

[IWM83] J. G. Izant, J. A. Weatherbee, and J. R. McIntosh. A microtubule-associated protein antigen unique to mitotic spindle microtubules in PtK1 cells. Journal of Cell Biology, 96(2): 424–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/2/424.

Iida:1984:DSH

[IY84a] H. Iida and I. Yahara. Durable synthesis of high molecular weight heat shock proteins in G0 cells of the yeast and other eucaryotes. Journal of Cell Biology, 99(1):199–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/99/1/199.

Iida:1984:HSR

[IY84b] H. Iida and I. Yahara. A heat shock-resistant mutant of Saccharomyces cerevisiae shows constitutive synthesis of two heat shock proteins and altered growth. Journal of Cell Biology, 99 (4):1441–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/4/1441.

Iida:1984:SEG

[IY84c] H. Iida and I. Yahara. Specific early–G1 blocks accompanied with stringent response in Saccharomyces cerevisiae lead to growth arrest in resting state similar to the G0 of higher eucaryotes. Journal of Cell Biology, 98(4):1185–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/4/
1185.

Ince:1983:EMP

[IYVV83] C. Ince, D. L. Ypey, R. Van Furth, and A. A. Verveen. Estimation of the membrane potential of cultured macrophages from the fast potential transient upon microelectrode entry. Journal of Cell Biology, 96(3):796–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/3/796.

Jaken:1981:CBS

[JB81a] S. Jaken and P. H. Black. Correlation between a specific molecular weight form of plasminogen activator and metabolic activity of 3T3 cells. Journal of Cell Biology, 90(3):721–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
90/3/721.

Jaken:1981:RPA

[JB81b] S. Jaken and P. H. Black. Regulation of plasminogen activator in 3T3 cells: effect of phorbol myristate acetate on subcellular distribution and molecular weight. Journal of Cell Biology, 90(3):727–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/90/3/727.

James:1981:PMD

[JB81c] T. W. James and R. Bohman. Proliferation of mitochondria during the cell cycle of the human cell line (HL-60). Journal of Cell Biology, 89(2):256–??, May 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/89/2/256.

Janero:1981:TMB

[JB81d] D. R. Janero and R. Barrnett. Thylakoid membrane biogenesis in Chlamydomonas reinhardtii 137+: cell cycle variations in the synthesis and assembly of polar glycerolipid. Journal of Cell Biology, 91(1):126–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/91/1/126.

Janero:1982:TMB

[JB82] D. R. Janero and R. Barrnett. Thylakoid membrane biogenesis in Chlamydomonas reinhardtii 137+. II. Cell-cycle variations in the synthesis and assembly of pigment. Journal of Cell Biology, 93(2):411–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/93/2/411.

Jorgensen:1984:EPC

[JC84] A. O. Jorgensen and K. P. Campbell. Evidence for the presence of calsequestrin in two structurally different regions of myocardial sarcoplasmic reticulum. Journal of Cell Biology, 98 (4):1597–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/4/1597.

Jackson:1982:MDD

[JD82] W. T. Jackson and B. G. Doyle. Membrane distribution in dividing endosperm cells of Haemanthus. Journal of Cell Biology, 94(3):637–??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/94/3/637.

Johnson:1980:RRO

[JDTW80] C. C. Johnson, W. E. Dawson, J. T. Turner, and J. H. Wyche. Regulation of rat ovarian cell growth and steroid secretion. Journal of Cell Biology, 86(2):483–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/86/2/483.

Johnson:1981:IPS

[JE81] C. H. Johnson and D. Epel. Intracellular pH of sea urchin eggs measured by the dimethyloxazolidinedione (DMO) method. Journal of Cell Biology, 89(2):284–??, May 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/89/2/284.

Jeffery:1982:MRC

[Jef82] W. R. Jeffery. Messenger RNA in the cytoskeletal framework: analysis by in situ hybridization. Journal of Cell Biology, 95 (1):1–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/95/1/1.

Jensen:1982:DSM

[Jen82] C. G. Jensen. Dynamics of spindle microtubule organization: kinetochore fiber microtubules of plant endosperm. Journal of Cell Biology, 92(2):540–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/92/2/540.

John:1982:NPM

[JFGC82] P. A. St John, S. C. Froehner, D. A. Goodenough, and J. B. Cohen. Nicotinic postsynaptic membranes from Torpedo: sidedness, permeability to macromolecules, and topography of major polypeptides. Journal of Cell Biology, 92(2): 333–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/2/333.

Jamieson:1984:TII

[JFS84] G. A. Jamieson, W. A. Frazier, and P. H. Schlesinger. Transient increase in intracellular pH during Dictyostelium differentiation. Journal of Cell Biology, 99(5):1883–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/5/
1883.

Jaffe:1982:SME

[JGSH82] L. A. Jaffe, M. Gould-Somero, and L. Z. Holland. Studies of the mechanism of the electrical polyspermy block using voltage clamp during cross-species fertilization. Journal of Cell Biology, 92(3):616–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/92/3/616.

Johansson:1984:SAR

[JH84] S. Johansson and M. Höök. Substrate adhesion of rat hepatocytes: on the mechanism of attachment to fibronectin. Journal of Cell Biology, 98(3):810–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/3/810.

Jung:1984:MCI

[JHCC84] G. Jung, R. M. Helm, C. A. Carraway, and K. L. Carraway. Mechanism of concanavalin A-induced anchorage of the major cell surface glycoproteins to the submembrane cytoskeleton in 13762 ascites mammary adenocarcinoma cells. Journal of Cell Biology, 98(1):179–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/1/179.

Johansson:1981:SAR

[JKHT81] S. Johansson, L. Kjellén, M. Höök, and R. Timpl. Substrate adhesion of rat hepatocytes: a comparison of laminin and fibronectin as attachment proteins. Journal of Cell Biology, 90 (1):260–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/1/260.

Janicke:1984:MHB

[JL84] M. A. Janicke and J. R. LaFountain. Malorientation in halfbivalents at anaphase: analysis of autosomal laggards in untreated, cold-treated, and cold-recovering crane fly spermatocytes. Journal of Cell Biology, 98(3):859–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/98/3/859.

Jesaitis:1984:RMF

[JNS^{+}84] A. J. Jesaitis, J. R. Naemura, L. A. Sklar, C. G. Cochrane, and R. G. Painter. Rapid modulation of N-formyl chemotactic peptide receptors on the surface of human granulocytes: formation of high-affinity ligand-receptor complexes in transient association with cytoskeleton. Journal of Cell Biology, 98(4):1378–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/4/1378.

Jacobson:1984:LDD

[JOA84] K. Jacobson, D. O’Dell, and J. T. August. Lateral diffusion of an 80,000-dalton glycoprotein in the plasma membrane of murine fibroblasts: relationships to cell structure and function. Journal of Cell Biology, 99(5):1624–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/5/

Jacobson:1984:RMC

[JOH^{+}84] K. Jacobson, D. O’Dell, B. Holifield, T. L. Murphy, and J. T. August. Redistribution of a major cell surface glycoprotein during cell movement. Journal of Cell Biology, 99(5):1613– ??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/5/1613.

Jones:1980:ILP

[JOM80] R. E. Jones, C. S. Okamura, and T. E. Martin. Immunofluorescent localization of the proteins of nuclear ribonucleoprotein complexes. Journal of Cell Biology, 86(1):235–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/86/1/235.

Johnson:1984:MFP

[JPS84] K. A. Johnson, M. E. Porter, and T. Shimizu. Mechanism of force production for microtubule-dependent movements. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/1/132s.

Jarvik:1980:OFM

[JR80] J. W. Jarvik and J. L. Rosenbaum. Oversized flagellar membrane protein in paralyzed mutants of Chlamydomonas reinhardrii. Journal of Cell Biology, 85(2):258–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/85/2/258.

Johnson:1980:MBSb

[JRB80] M. I. Johnson, C. D. Ross, and R. P. Bunge. Morphological and biochemical studies on the development of cholinergic properties in cultured sympathetic neurons. II. Dependence on postnatal age. Journal of Cell Biology, 84(3):692–??, March 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/84/3/
692.

Johnson:1980:MBSa

[JRM^{+}80] M. I. Johnson, C. D. Ross, M. Meyers, E. L. Spitznagel, and R. P. Bunge. Morphological and biochemical studies on the development of cholinergic properties in cultured sympathetic neurons. I. Correlative changes in choline acetyltransferase and synaptic vesicle cytochemistry. Journal of Cell Biology, 84 (3):680–??, March 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/84/3/680.

Joseph-Silverstein:1984:CSN

[JSC84] J. Joseph-Silverstein and W. D. Cohen. The cytoskeletal system of nucleated erythrocytes. III. Marginal band function in mature cells. Journal of Cell Biology, 98(6):2118–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/6/
2118.

Jorgensen:1983:ULC

[JSCM83] A. O. Jorgensen, A. C. Shen, K. P. Campbell, and D. H. MacLennan. Ultrastructural localization of calsequestrin in rat skeletal muscle by immunoferritin labeling of ultrathin frozen sections. Journal of Cell Biology, 97(5):1573–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/5/
1573.

Jorgensen:1982:LCM

[JSDM82] A. O. Jorgensen, A. C. Shen, P. Daly, and D. H. MacLennan. Localization of Ca^{2+} + Mg2+-ATPase of the sarcoplasmic reticulum in adult rat papillary muscle. Journal of Cell Biology, 93(3):883–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/93/3/883.

Jorgensen:1982:ULC

[JSMT82] A. O. Jorgensen, A. C. Shen, D. H. MacLennan, and K. T. Tokuyasu. Ultrastructural localization of the Ca^{2+} + Mg2+- dependent ATPase of sarcoplasmic reticulum in rat skeletal muscle by immunoferritin labeling of ultrathin frozen sections. Journal of Cell Biology, 92(2):409–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/92/2/409.

Johnson:1980:SVC

[JTM80] R. M. Johnson, G. Taylor, and D. B. Meyer. Shape and volume changes in erythrocyte ghosts and spectrin-actin networks. Journal of Cell Biology, 86(2):371–??, August 1980.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/86/2/371.

Jones-Villeneuve:1982:RAI

[JVMRK82] E. M. Jones-Villeneuve, M. W. McBurney, K. A. Rogers, and V. I. Kalnins. Retinoic acid induces embryonal carcinoma cells to differentiate into neurons and glial cells. Journal of Cell Biology, 94(2):253–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/94/2/253.

Johnson:1983:SMW

[JW83] K. A. Johnson and J. S. Wall. Structure and molecular weight of the dynein ATPase. Journal of Cell Biology, 96(3):669–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
96/3/669.

Johnson:1981:MRM

[JWBC81] L. V. Johnson, M. L. Walsh, B. J. Bockus, and L. B. Chen. Monitoring of relative mitochondrial membrane potential in living cells by fluorescence microscopy. Journal of Cell Biology, 88(3):526–??, March 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/88/3/526.

Johnson-Wint:1984:RCT

[JWG84] B. Johnson-Wint and J. Gross. Regulation of connective tissue collagenase production: stimulators from adult and fetal epidermal cells. Journal of Cell Biology, 98(1):90–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/1/
90.

Johnson:1981:IPM

[JZ81] M. H. Johnson and C. A. Ziomek. Induction of polarity in mouse 8-cell blastomeres: specificity, geometry, and stability. Journal of Cell Biology, 91(1):303–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/91/1/303.

Kilmartin:1984:SRT

[KA84] J. V. Kilmartin and A. E. Adams. Structural rearrangements of tubulin and actin during the cell cycle of the yeast Saccharomyces. Journal of Cell Biology, 98(3):922–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/98/3/922.

Kane:1980:IEC

[Kan80] R. E. Kane. Induction of either contractile or structural actin-based gels in sea urchin egg cytoplasmic extract. Journal of Cell Biology, 86(3):803–??, September 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/86/3/803.

Kane:1983:ISC

[Kan83] R. E. Kane. Interconversion of structural and contractile actin gels by insertion of myosin during assembly. Journal of Cell Biology, 97(6):1745–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/97/6/1745.

Kuriyama:1981:CCC

[KB81a] R. Kuriyama and G. G. Borisy. Centriole cycle in Chinese hamster ovary cells as determined by whole-mount electron microscopy. Journal of Cell Biology, 91(3):814–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/91/3/
814.

Kuriyama:1981:MNA

[KB81b] R. Kuriyama and G. G. Borisy. Microtubule-nucleating activity of centrosomes in Chinese hamster ovary cells is independent of the centriole cycle but coupled to the mitotic cycle. Journal of Cell Biology, 91(3):822–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/91/3/822.

Kurth:1981:LCP

[KB81c] P. D. Kurth and M. Bustin. Localization of chromosomal protein HMG-1 in polytene chromosomes of Chironomus thummi. Journal of Cell Biology, 89(1):70–??, April 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/89/1/70.

Kiino:1982:PSC

[KBD82] D. R. Kiino, D. E. Burger, and P. S. Dannies. Prolactin storage in a clonal strain of rat pituitary tumor cells is cell-cycle dependent. Journal of Cell Biology, 93(2):459–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/93/2/459.

Kramer:1984:BLF

[KBDK84] R. H. Kramer, K. G. Bensch, P. M. Davison, and M. A. Karasek. Basal lamina formation by cultured microvascular endothelial cells. Journal of Cell Biology, 99(2):692–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/2/
692.

Kielian:1980:PLF

[KC80a] M. C. Kielian and Z. A. Cohn. Phagosome-lysosome fusion. Characterization of intracellular membrane fusion in mouse macrophages. Journal of Cell Biology, 85(3):754–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/85/3/
754.

Kleve:1980:AAS

[KC80b] M. G. Kleve and W. H. Clark. Association of actin with sperm centrioles: isolation of centriolar complexes and immunofluorescent localization of actin. Journal of Cell Biology, 86(1):87–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/86/1/87.

Kielian:1982:IAPb

[KC82] M. C. Kielian and Z. A. Cohn. Intralysosomal accumulation of polyanions. II. Polyanion internalization and its influence on lysosomal pH and membrane fluidity. Journal of Cell Biology, 93(3):875–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/93/3/875.

Kayman:1983:RBA

[KC83] S. C. Kayman and M. Clarke. Relationship between axenic growth of Dictyostelium discoideum strains and their track morphology on substrates coated with gold particles. Journal of Cell Biology, 97(4):1001–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/4/1001.

Koonce:1984:LIC

[KCB84] M. P. Koonce, R. A. Cloney, and M. W. Berns. Laser irradiation of centrosomes in newt eosinophils: evidence of centriole role in motility. Journal of Cell Biology, 98(6):1999–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/6/
1999.

Kropf:1984:TIC

[KCGH84] D. L. Kropf, J. H. Caldwell, N. A. Gow, and F. M. Harold. Transcellular ion currents in the water mold Achlya. Amino acid proton symport as a mechanism of current entry. Journal of Cell Biology, 99(2):486–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/2/486.

Kaplan:1983:ISM

[KCM83] D. R. Kaplan, J. R. Colca, and M. L. McDaniel. Insulin as a surface marker on isolated cells from rat pancreatic islets. Journal of Cell Biology, 97(2):433–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/2/433.

Kinsey:1980:IPC

[KDL80] W. H. Kinsey, G. L. Decker, and W. J. Lennarz. Isolation and partial characterization of the plasma membrane of the sea urchin egg. Journal of Cell Biology, 87(1):248–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/87/1/
248.

Kistler:1984:MNP

[KDL84] J. Kistler, Y. Duncombe, and U. K. Laemmli. Mapping nucleolar proteins with monoclonal antibodies. Journal of Cell Biology, 99(6):1981–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/6/1981.

Keith:1983:MCC

[KDMS83] C. Keith, M. DiPaola, F. R. Maxfield, and M. L. Shelanski. Microinjection of Ca++-calmodulin causes a localized depolymerization of microtubules. Journal of Cell Biology, 97(6): 1918–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/6/1918.

Keith:1981:DVF

[KFS81] C. H. Keith, J. R. Feramisco, and M. Shelanski. Direct visualization of fluorescein-labeled microtubules in vitro and in microinjected fibroblasts. Journal of Cell Biology, 88(1): 234–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/88/1/234.

Kronenberg:1983:TCB

[KFV83] H. M. Kronenberg, B. J. Fennick, and T. J. Vasicek. Transport and cleavage of bacterial pre-β-lactamase by mammalian microsomes. Journal of Cell Biology, 96(4):1117–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/4/
1117.

Kensler:1980:IMM

[KG80] R. W. Kensler and D. A. Goodenough. Isolation of mouse myocardial gap junctions. Journal of Cell Biology, 86(3): 755–??, September 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/86/3/755.

Kaslow:1981:CTC

[KGAB81] H. R. Kaslow, V. E. Groppi, M. E. Abood, and H. R. Bourne. Cholera toxin can catalyze ADP-ribosylation of cytoskeletal proteins. Journal of Cell Biology, 91(2):410–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/91/2/
410.

Kirchhausen:1984:SDC

[KH84] T. Kirchhausen and S. C. Harrison. Structural domains of clathrin heavy chains. Journal of Cell Biology, 99(5):1725– ??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/5/1725.

Kanwar:1981:PCN

[KHF81] Y. S. Kanwar, V. C. Hascall, and M. G. Farquhar. Partial characterization of newly synthesized proteoglycans isolated from the glomerular basement membrane. Journal of Cell Biology, 90(2):527–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/90/2/527.

Kanwar:1984:EBD

[KHJG84] Y. S. Kanwar, V. C. Hascall, M. L. Jakubowski, and J. T. Gibbons. Effect of beta–D–xyloside on the glomerular proteoglycans. I. Biochemical studies. Journal of Cell Biology, 99 (2):715–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/2/715.

King:1982:AMS

[KHL82] S. M. King, J. S. Hyams, and A. Luba. Absence of microtubule sliding and an analysis of spindle formation and elongation in isolated mitotic spindles from the yeast Saccharomyces cerevisiae. Journal of Cell Biology, 94(2):341–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/94/2/341.

Khoury:1982:RBG

[Kho82] E. L. Khoury. Reexpression of blood group ABH antigens on the surface of human thyroid cells in culture. Journal of Cell Biology, 94(1):193–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/94/1/193.

Kaariainen:1980:MFI

[KHS^{+}80a] L. Kaariainen, K. Hashimoto, J. Saraste, I. Virtanen, and K. Penttinen. Monensin and FCCP inhibit the intracellular transport of alphavirus membrane glycoproteins. Journal of Cell Biology, 87(3):783–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/87/3/783.

Klymkowsky:1980:PES

[KHS80b] M. W. Klymkowsky, J. E. Heuser, and R. M. Stroud. Protease effects on the structure of acetylcholine receptor membranes from Torpedo californica. Journal of Cell Biology, 85 (3):823–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/85/3/823.

Kellems:1982:ADR

[KHS82] R. E. Kellems, M. E. Harper, and L. M. Smith. Amplified dihydrofolate reductase genes are located in chromosome regions containing DNA that replicates during the first half of S-phase. Journal of Cell Biology, 92(2):531–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/92/2/531.

Kosaka:1983:RBM

[KI83] T. Kosaka and K. Ikeda. Reversible blockage of membrane retrieval and endocytosis in the garland cell of the temperaturesensitive mutant of Drosophila melanogaster, shibirets1. Journal of Cell Biology, 97(2):499–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/2/499.

Kiehart:1981:SVS

[Kie81] D. P. Kiehart. Studies on the in vivo sensitivity of spindle microtubules to calcium ions and evidence for a vesicular calcium-sequestering system. Journal of Cell Biology, 88 (3):604–??, March 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/88/3/604.

Kirschner:1980:ITS

[Kir80] M. W. Kirschner. Implications of treadmilling for the stability and polarity of actin and tubulin polymers in vivo. Journal of Cell Biology, 86(1):330–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/86/1/330.

Kuchka:1982:AFS

[KJ82] M. R. Kuchka and J. W. Jarvik. Analysis of flagellar size control using a mutant of Chlamydomonas reinhardtii with a variable number of flagella. Journal of Cell Biology, 92(1): 170–??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/1/170.

Keller:1982:SPSb

[KJBR82] T. C. Keller, D. K. Jemiolo, W. H. Burgess, and L. I. Rebhun. Strongylocentrotus purpuratus spindle tubulin. II. Characteristics of its sensitivity to Ca++ and the effects of calmodulin isolated from bovine brain and S. purpuratus eggs. Journal of Cell Biology, 93(3):797–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/93/3/797.

Kaplan:1982:TSI

[KK82] J. Kaplan and E. A. Keogh. Temperature shifts induce the selective loss of alveolar-macrophage plasma membrane components. Journal of Cell Biology, 94(1):12–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/94/1/12.

Kondor-Koch:1983:ESF

[KKBG83] C. Kondor-Koch, B. Burke, and H. Garoff. Expression of Semliki Forest virus proteins from cloned complementary DNA. I. The fusion activity of the spike glycoprotein. Journal of Cell Biology, 97(3):644–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/97/3/644.

Kato:1984:ISL

[KKG84] M. Kato, K. Kato, and D. S. Goodman. Immunocytochemical studies on the localization of plasma and of cellular retinolbinding proteins and of transthyretin (prealbumin) in rat liver and kidney. Journal of Cell Biology, 98(5):1696–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/5/
1696.

Karnovsky:1982:CLD

[KKHK82] M. J. Karnovsky, A. M. Kleinfeld, R. L. Hoover, and R. D. Klausner. The concept of lipid domains in membranes. Journal of Cell Biology, 94(1):1–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/94/1/1.

Kielian:1984:MFM

[KKKH84] M. C. Kielian, S. Keränen, L. Kääriäinen, and A. Helenius. Membrane fusion mutants of Semliki Forest virus. Journal of Cell Biology, 98(1):139–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/1/139.

Kleinman:1981:RCM

[KKM81] H. K. Kleinman, R. J. Klebe, and G. R. Martin. Role of collagenous matrices in the adhesion and growth of cells. Journal of Cell Biology, 88(3):473–??, March 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/88/3/473.

Kruskal:1984:TRH

[KKM84] B. A. Kruskal, C. H. Keith, and F. R. Maxfield. Thyrotropinreleasing hormone-induced changes in intracellular [Ca^{2+}] measured by microspectrofluorometry on individual quin2loaded cells. Journal of Cell Biology, 99(3):1167–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/3/
1167.

Karsenti:1984:RCO

[KKMK84] E. Karsenti, S. Kobayashi, T. Mitchison, and M. Kirschner. Role of the centrosome in organizing the interphase microtubule array: properties of cytoplasts containing or lacking centrosomes. Journal of Cell Biology, 98(5):1763–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/5/
1763.

Kiehart:1984:DLM

[KKP84a] D. P. Kiehart, D. A. Kaiser, and T. D. Pollard. Direct localization of monoclonal antibody-binding sites on Acanthamoeba myosin–II and inhibition of filament formation by antibodies that bind to specific sites on the myosin–II tail. Journal of Cell Biology, 99(3):1015–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/3/1015.

Kiehart:1984:MAD

[KKP84b] D. P. Kiehart, D. A. Kaiser, and T. D. Pollard. Monoclonal antibodies demonstrate limited structural homology between myosin isozymes from Acanthamoeba. Journal of Cell Biology, 99(3):1002–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/3/1002.

Kinsey:1981:IGF

[KL81] W. H. Kinsey and W. J. Lennarz. Isolation of a glycopeptide fraction from the surface of the sea urchin egg that inhibits sperm-egg binding and fertilization. Journal of Cell Biology, 91(2):325–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/91/2/325.

Kensler:1982:EMO

[KL82] R. W. Kensler and R. J. Levine. An electron microscopic and optical diffraction analysis of the structure of Limulus telson muscle thick filaments. Journal of Cell Biology, 92(2): 443–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/2/443.

Klagsbrun:1980:BCS

[Kla80] M. Klagsbrun. Bovine colostrum supports the serum-free proliferation of epithelial cells but not of fibroblasts in long-term culture. Journal of Cell Biology, 84(3):808–??, March 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/84/3/808.

Konieczny:1983:AMP

[KLC83] S. F. Konieczny, J. B. Lawrence, and J. R. Coleman. Analysis of muscle protein expression in polyethylene glycol-induced chicken: rat myoblast heterokaryons. Journal of Cell Biology, 97(5):1348–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/5/1348.

Krauter:1982:SGM

[KLD^{+}82] K. Krauter, L. Leinwand, P. D’Eustachio, F. Ruddle, and J. E. Darnell. Structural genes of the mouse major urinary protein are on chromosome 4. Journal of Cell Biology, 94(2):414–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
94/2/414.

Kanwar:1980:IPG

[KLF80] Y. S. Kanwar, A. Linker, and M. G. Farquhar. Increased permeability of the glomerular basement membrane to ferritin after removal of glycosaminoglycans (heparan sulfate) by enzyme digestion. Journal of Cell Biology, 86(2):688–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/86/2/
688.

Keller:1982:CCD

[KM82] T. C. Keller and M. S. Mooseker. Ca++-calmodulin-dependent phosphorylation of myosin, and its role in brush border contraction in vitro. Journal of Cell Biology, 95(3): 943–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/95/3/943.

Klempner:1980:NPM

[KMCAS80] M. S. Klempner, R. B. Mikkelsen, D. H. Corfman, and J. André-Schwartz. Neutrophil plasma membranes. I. Highyield purification of human neutrophil plasma membrane vesicles by nitrogen cavitation and differential centrifugation. Journal of Cell Biology, 86(1):21–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/86/1/21.

Kachadorian:1981:ROF

[KMF81] W. A. Kachadorian, J. Muller, and A. Finkelstein. Role of osmotic forces in exocytosis: studies of ADH-induced fusion in toad urinary bladder. Journal of Cell Biology, 91(2): 584–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/91/2/584.

Kim:1984:EUL

[KMF84] K. H. Kim, D. Marchuk, and E. Fuchs. Expression of unusually large keratins during terminal differentiation: balance of type I and type II keratins is not disrupted. Journal of Cell Biology, 99(5):1872–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/5/1872.

Kiehart:1982:EMD

[KMI82] D. P. Kiehart, I. Mabuchi, and S. Inoué. Evidence that myosin does not contribute to force production in chromosome movement. Journal of Cell Biology, 94(1):165–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/94/1/165.

Klymkowsky:1983:MBI

[KML83] M. W. Klymkowsky, R. H. Miller, and E. B. Lane. Morphology, behavior, and interaction of cultured epithelial cells after the antibody-induced disruption of keratin filament organization. Journal of Cell Biology, 96(2):494–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/2/494.

Kuettner:1982:SCMb

[KMP^{+}82] K. E. Kuettner, V. A. Memoli, B. U. Pauli, N. C. Wrobel, E. J. Thonar, and J. C. Daniel. Synthesis of cartilage matrix by mammalian chondrocytes in vitro. II. Maintenance of collagen and proteoglycan phenotype. Journal of Cell Biology, 93(3):751–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/93/3/751.

Kopf:1982:ICS

[KMV82] G. S. Kopf, G. W. Moy, and V. D. Vacquier. Isolation and characterization of sea urchin egg cortical granules. Journal of Cell Biology, 95(3):924–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/95/3/924.

Kerjaschki:1984:MDG

[KNBSF84] D. Kerjaschki, L. Noronha-Blob, B. Sacktor, and M. G. Farquhar. Microdomains of distinctive glycoprotein composition in the kidney proximal tubule brush border. Journal of Cell Biology, 98(4):1505–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/4/1505.

Koehler:1980:CBP

[KNH80] J. K. Koehler, E. D. Nudelman, and S. Hakomori. A collagenbinding protein on the surface of ejaculated rabbit spermatozoa. Journal of Cell Biology, 86(2):529–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/86/2/529.

Karsenti:1984:IMI

[KNHK84] E. Karsenti, J. Newport, R. Hubble, and M. Kirschner. Interconversion of metaphase and interphase microtubule arrays, as studied by the injection of centrosomes and nuclei into Xenopus eggs. Journal of Cell Biology, 98(5):1730–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/5/
1730.

Karsenti:1984:RRC

[KNK84] E. Karsenti, J. Newport, and M. Kirschner. Respective roles of centrosomes and chromatin in the conversion of microtubule arrays from interphase to metaphase. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/1/47s.

Ko:1984:RAZ

[Ko84] C. P. Ko. Regeneration of the active zone at the frog neuromuscular junction. Journal of Cell Biology, 98(5):1685–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
98/5/1685.

Koppel:1982:SFD

[KOB82] D. E. Koppel, J. M. Oliver, and R. D. Berlin. Surface functions during mitosis. III. Quantitative analysis of ligand-receptor movement into the cleavage furrow: diffusion vs. flow. Journal of Cell Biology, 93(3):950–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/93/3/950.

Keski-Oja:1981:KFM

[KOLV81] J. Keski-Oja, V. P. Lehto, and I. Virtanen. Keratin filaments of mouse epithelial cells are rapidly affected by epidermal growth factor. Journal of Cell Biology, 90(2):537–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/90/2/537.

Kreibich:1982:RRR

[KORBS82] G. Kreibich, G. Ojakian, E. Rodriguez-Boulan, and D. D. Sabatini. Recovery of ribophorins and ribosomes in “inverted rough” vesicles derived from rat liver rough microsomes. Journal of Cell Biology, 93(1):111–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/93/1/111.

Knapp:1983:RKC

[KOS83] L. W. Knapp, W. M. O’Guin, and R. H. Sawyer. Rearrangement of the keratin cytoskeleton after combined treatment with microtubule and microfilament inhibitors. Journal of Cell Biology, 97(6):1788–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/97/6/1788.

Keski-Oja:1980:DAF

[KOST80] J. Keski-Oja, A. Sen, and G. J. Todaro. Direct association of fibronectin and actin molecules in vitro. Journal of Cell Biology, 85(3):527–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/85/3/527.

Kiehart:1984:IAA

[KP84] D. P. Kiehart and T. D. Pollard. Inhibition of Acanthamoeba actomyosin–II ATPase activity and mechanochemical function by specific monoclonal antibodies. Journal of Cell Biology, 99(3):1024–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/3/1024.

Kuettner:1982:SCMa

[KPG^{+}82] K. E. Kuettner, B. U. Pauli, G. Gall, V. A. Memoli, and R. K. Schenk. Synthesis of cartilage matrix by mammalian chondrocytes in vitro. I. Isolation, culture characteristics, and morphology. Journal of Cell Biology, 93(3):743–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/93/3/743.

Kellie:1983:CCT

[KPPC83] S. Kellie, B. Patel, E. J. Pierce, and D. R. Critchley. Capping of cholera toxin-ganglioside GM1 complexes on mouse lymphocytes is accompanied by co-capping of alpha-actinin. Journal of Cell Biology, 97(2):447–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/97/2/447.

Kelly:1983:TMC

[KPW^{+}83] W. G. Kelly, A. Passaniti, J. W. Woods, J. L. Daiss, and T. F. Roth. Tubulin as a molecular component of coated vesicles. Journal of Cell Biology, 97(4):1191–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/4/1191.

Kanwar:1982:CGB

[KR82a] Y. S. Kanwar and L. J. Rosenzweig. Clogging of the glomerular basement membrane. Journal of Cell Biology, 93(2):489–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
93/2/489.

Keller:1982:SPSa

[KR82b] T. C. Keller and L. I. Rebhun. Strongylocentrotus purpuratus spindle tubulin. I. Characteristics of its polymerization and depolymerization in vitro. Journal of Cell Biology, 93(3):788–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
93/3/788.

Keryer:1984:CDD

[KRB84a] G. Keryer, H. Ris, and G. G. Borisy. Centriole distribution during tripolar mitosis in Chinese hamster ovary cells. Journal of Cell Biology, 98(6):2222–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/6/2222.

Kravit:1984:RSP

[KRB84b] N. G. Kravit, C. S. Regula, and R. D. Berlin. A reevaluation of the structure of purified tubulin in solution: evidence for the prevalence of oligomers over dimers at room temperature. Journal of Cell Biology, 99(1):188–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/1/188.

Kuypers:1984:SCH

[KRB^{+}84c] F. A. Kuypers, B. Roelofsen, W. Berendsen, J. A. Op den Kamp, and L. L. van Deenen. Shape changes in human erythrocytes induced by replacement of the native phosphatidylcholine with species containing various fatty acids. Journal of Cell Biology, 99(6):2260–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/6/2260.

Kayman:1982:MMD

[KRC82] S. C. Kayman, M. Reichel, and M. Clarke. Motility mutants of Dictyostelium discoideum. Journal of Cell Biology, 93(3):705–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/93/3/705.

Kline:1983:ICM

[KRN83] D. Kline, K. R. Robinson, and R. Nuccitelli. Ion currents and membrane domains in the cleaving Xenopus egg. Journal of Cell Biology, 97(6):1753–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/6/1753.

Krawisz:1982:CPG

[KS82] B. R. Krawisz and R. E. Scott. Coupling of proadipocyte growth arrest and differentiation. I. Induction by heparinized medium containing human plasma. Journal of Cell Biology, 94 (2):394–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/94/2/394.

Kensler:1983:FSM

[KS83] R. W. Kensler and M. Stewart. Frog skeletal muscle thick filaments are three-stranded. Journal of Cell Biology, 96(6): 1797–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/6/1797.

Krystosek:1984:PNS

[KS84] A. Krystosek and N. W. Seeds. Peripheral neurons and Schwann cells secrete plasminogen activator. Journal of Cell Biology, 98(2):773–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/2/773.

Kielian:1982:IAPa

[KSC82] M. C. Kielian, R. M. Steinman, and Z. A. Cohn. Intralysosomal accumulation of polyanions. I. Fusion of pinocytic and phagocytic vacuoles with secondary lysosomes. Journal of Cell Biology, 93(3):866–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/93/3/866.

Kleinschmidt:1983:HMG

[KSD^{+}83] J. A. Kleinschmidt, U. Scheer, M. C. Dabauvalle, M. Bustin, and W. W. Franke. High mobility group proteins of amphibian oocytes: a large storage pool of a soluble high mobility group-1-like protein and involvement in transcriptional events. Journal of Cell Biology, 97(3):838–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/3/838.

Kerjaschki:1984:ICP

[KSF84] D. Kerjaschki, D. J. Sharkey, and M. G. Farquhar. Identification and characterization of podocalyxin–the major sialoprotein of the renal glomerular epithelial cell. Journal of Cell Biology, 98(4):1591–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/4/1591.

Kay:1981:VTC

[KSG81] B. K. Kay, O. Schmidt, and J. G. Gall. In vitro transcription of cloned 5S RNA genes of the newt Notophthalmus. Journal of Cell Biology, 90(2):323–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/90/2/323.

Koenig:1983:RCS

[KSI83] J. H. Koenig, K. Saito, and K. Ikeda. Reversible control of synaptic transmission in a single gene mutant of Drosophila melanogaster. Journal of Cell Biology, 96(6):1517–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/6/
1517.

Krohne:1982:ILM

[KSK^{+}82] G. Krohne, R. Stick, J. A. Kleinschmidt, R. Moll, W. W. Franke, and P. Hausen. Immunological localization of a major karyoskeletal protein in nucleoli of oocytes and somatic cells of Xenopus laevis. Journal of Cell Biology, 94(3):749–??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
94/3/749.

Kartenbeck:1984:AVF

[KSMF84] J. Kartenbeck, K. Schwechheimer, R. Moll, and W. W. Franke. Attachment of vimentin filaments to desmosomal plaques in human meningiomal cells and arachnoidal tissue. Journal of Cell Biology, 98(3):1072–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/3/1072.

Kane:1980:DIL

[KSR^{+}80] A. B. Kane, R. P. Stanton, E. G. Raymond, M. E. Dobson, M. E. Knafelc, and J. L. Farber. Dissociation of intracellular lysosomal rupture from the cell death caused by silica. Journal of Cell Biology, 87(3):643–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/87/3/643.

Kim:1983:AIS

[KSSP83] Y. J. Kim, J. Shuman, M. Sette, and A. Przybyla. Arsenate induces stress proteins in cultured rat myoblasts. Journal of Cell Biology, 96(2):393–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/2/393.

Keller:1984:TAB

[KSSR84] L. R. Keller, J. A. Schloss, C. D. Silflow, and J. L. Rosenbaum. Transcription of alpha- and beta-tubulin genes in vitro in isolated Chlamydomonas reinhardi nuclei. Journal of Cell Biology, 98(3):1138–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/3/1138.

Kuehl:1984:CHM

[KST84] L. Kuehl, B. Salmond, and L. Tran. Concentrations of highmobility-group proteins in the nucleus and cytoplasm of several rat tissues. Journal of Cell Biology, 99(2):648–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/2/
648.

Kuhlbrandt:1983:SMC

[KTW83] W. Kühlbrandt, T. Thaler, and E. Wehrli. The structure of membrane crystals of the light-harvesting chlorophyll a/b protein complex. Journal of Cell Biology, 96(5):1414–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/5/
1414.

Kubai:1982:MSC

[Kub82] D. F. Kubai. Meiosis in Sciara coprophila: structure of the spindle and chromosome behavior during the first meiotic division. Journal of Cell Biology, 93(3):655–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/93/3/655.

Kaneda:1984:EDT

[KUM^{+}84] Y. Kaneda, T. Uchida, E. Mekada, M. Nakanishi, and Y. Okada. Entry of diphtheria toxin into cells: possible existence of cellular factor(s) for entry of diphtheria toxin into cells was studied in somatic cell hybrids and hybrid toxins. Journal of Cell Biology, 98(2):466–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/2/466.

Kuo:1982:ADA

[Kuo82] M. T. Kuo. Analysis of DNA attached to the chromosome scaffold. Journal of Cell Biology, 93(2):278–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/93/2/278.

Klausner:1984:FRI

[KvRK^{+}84] R. D. Klausner, J. van Renswoude, C. Kempf, K. Rao, J. L. Bateman, and A. R. Robbins. Failure to release iron from transferrin in a Chinese hamster ovary cell mutant pleiotropically defective in endocytosis. Journal of Cell Biology, 98 (3):1098–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/3/1098.

Kubai:1981:NCS

[KW81a] D. F. Kubai and D. Wise. Nonrandom chromosome segregation in Neocurtilla (Gryllotalpa) hexadactyla: an ultrastructural study. Journal of Cell Biology, 88(2):281–??, February 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/88/2/
281.

Kuby:1981:IPO

[KW81b] J. M. Kuby and L. Wofsy. Intramembrane particles and the organization of lymphocyte membrane proteins. Journal of Cell Biology, 88(3):591–??, March 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/88/3/591.

Kamiya:1984:SLC

[KW84] R. Kamiya and G. B. Witman. Submicromolar levels of calcium control the balance of beating between the two flagella in demembranated models of Chlamydomonas. Journal of Cell Biology, 98(1):97–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/1/97.

Kilmartin:1982:RMA

[KWM82] J. V. Kilmartin, B. Wright, and C. Milstein. Rat monoclonal antitubulin antibodies derived by using a new nonsecreting rat cell line. Journal of Cell Biology, 93(3):576–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/93/3/576.

Kenter:1984:MMC

[KWSB84] A. L. Kenter, T. Warren, D. Shields, and B. K. Birshtein. Mouse myeloma cells that make short immunoglobulin heavy chains: pleiotropic effects on glycosylation and chain assembly. Journal of Cell Biology, 98(6):2215–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/6/2215.

Kabat-Zinn:1981:SUT

[KZS81] J. Kabat-Zinn and R. H. Singer. Sea urchin tube feet: unique structures that allow a cytological and molecular approach to the study of actin and its gene expression. Journal of Cell Biology, 89(1):109–??, April 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/89/1/109.

Lee:1980:CEC

[LA80a] H. C. Lee and N. Auersperg. Calcium in epithelial cell contraction. Journal of Cell Biology, 85(2):325–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/85/2/325.

Lovrien:1980:SWG

[LA80b] R. E. Lovrien and R. A. Anderson. Stoichiometry of wheat germ agglutinin as a morphology controlling agent and as a morphology controlling agent and as a morphology protective agent for the human erythrocyte. Journal of Cell Biology, 85 (3):534–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/85/3/534.

Lloyd:1983:MIR

[LA83] C. E. Lloyd and M. Ascoli. On the mechanisms involved in the regulation of the cell-surface receptors for human choriogonadotropin and mouse epidermal growth factor in cultured Leydig tumor cells. Journal of Cell Biology, 96(2): 521–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/2/521.

Loftus:1984:RFR

[LA84] J. C. Loftus and R. M. Albrecht. Redistribution of the fibrinogen receptor of human platelets after surface activation. Journal of Cell Biology, 99(3):822–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/3/822.

Levine:1984:ILI

[LAAS84] J. S. Levine, R. H. Allen, D. H. Alpers, and B. Seetharam. Immunocytochemical localization of the intrinsic factorcobalamin receptor in dog-ileum: distribution of intracellular receptor during cell maturation. Journal of Cell Biology, 98 (3):1111–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/3/1111.

Langford:1980:ASM

[Lan80] G. M. Langford. Arrangement of subunits in microtubules with 14 profilaments. Journal of Cell Biology, 87(2):521–??, November 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
87/2/521.

Lane:1982:MAP

[Lan82] E. B. Lane. Monoclonal antibodies provide specific intramolecular markers for the study of epithelial tonofilament organization. Journal of Cell Biology, 92(3):665–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/92/3/665.

Lawson:1983:ENP

[Law83] D. Lawson. Epinemin: a new protein associated with vimentin filaments in non-neural cells. Journal of Cell Biology, 97(6): 1891–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/6/1891.

Lawson:1984:DEC

[Law84] D. Lawson. Distribution of epinemin in colloidal gold-labelled, quick-frozen, deep-etched cytoskeletons. Journal of Cell Biology, 99(4):1451–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/4/1451.

Leech:1984:PFM

[LBA^{+}84] J. H. Leech, J. W. Barnwell, M. Aikawa, L. H. Miller, and R. J. Howard. Plasmodium falciparum malaria: association of knobs on the surface of infected erythrocytes with a histidine-rich protein and the erythrocyte skeleton. Journal of Cell Biology, 98(4):1256–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/4/1256.

Leitman:1984:GSC

[LBJ84] D. C. Leitman, S. C. Benson, and L. K. Johnson. Glucocorticoids stimulate collagen and noncollagen protein synthesis in cultured vascular smooth muscle cells. Journal of Cell Biology, 98(2):541–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/2/541.

Lewis:1982:GAA

[LBM82] J. A. Lewis, J. L. Biedler, and P. W. Melera. Gene amplification accompanies low level increases in the activity of dihydrofolate reductase in antifolate-resistant Chinese hamster lung cells containing abnormally banding chromosomes. Journal of Cell Biology, 94(2):418–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/94/2/418.

Lane:1980:DEE

[LC80] N. J. Lane and H. J. Chandler. Definitive Evidence for the existence of tight junctions in invertebrates. Journal of Cell Biology, 86(3):765–??, September 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/86/3/765.

Lanks:1981:ICC

[LC81] K. W. Lanks and N. W. Chin. Identification of cytoskeletal components involved in attachment of L929 cells and macrophages to polystyrene. Journal of Cell Biology, 89 (3):691–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/89/3/691.

Loftus:1984:PAC

[LCA84] J. C. Loftus, J. Choate, and R. M. Albrecht. Platelet activation and cytoskeletal reorganization: high voltage electron microscopic examination of intact and Triton-extracted whole mounts. Journal of Cell Biology, 98(6):2019–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/6/

Leys:1984:DRG

[LCK84] E. J. Leys, G. F. Crouse, and R. E. Kellems. Dihydrofolate reductase gene expression in cultured mouse cells is regulated by transcript stabilization in the nucleus. Journal of Cell Biology, 99(1):180–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/1/180.

Larson:1981:IUP

[LD81] D. E. Larson and A. D. Dingle. Isolation, ultrastructure, and protein composition of the flagellar rootlet of Naegleria gruberi. Journal of Cell Biology, 89(3):424–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/89/3/424.

Laliberte:1984:IDM

[LDF^{+}84] J. F. Laliberté, A. Dagenais, M. Filion, V. Bibor-Hardy, R. Simard, and A. Royal. Identification of distinct messenger RNAs for nuclear lamin C and a putative precursor of nuclear lamin A. Journal of Cell Biology, 98(3):980–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/3/
980.

Lin:1980:LCR

[LDM80] C. Lin, J. Dedman, and A. Means. Localization of calmodulin in rat cerebellum by immunoelectron microscopy. Journal of Cell Biology, 85(2):473–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/85/2/473.

Langanger:1984:ULA

[LdMM^{+}84] G. Langanger, J. de Mey, M. Moeremans, G. Daneels, M. de Brabander, and J. V. Small. Ultrastructural localization of alpha-actinin and filamin in cultured cells with the immunogold staining (IGS) method. Journal of Cell Biology, 99(4): 1324–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/4/1324.

Liesi:1983:LPE

[LDV83] P. Liesi, D. Dahl, and A. Vaheri. Laminin is produced by early rat astrocytes in primary culture. Journal of Cell Biology, 96 (3):920–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/3/920.

Lehrer:1981:DAF

[Leh81] S. S. Lehrer. Damage to actin filaments by glutaraldehyde: protection by tropomyosin. Journal of Cell Biology, 90(2): 459–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/2/459.

Letourneau:1983:DOA

[Let83] P. C. Letourneau. Differences in the organization of actin in the growth cones compared with the neurites of cultured neurons from chick embryos. Journal of Cell Biology, 97(4): 963–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/4/963.

Larson:1984:MCV

[LFAG84] D. M. Larson, K. Fujiwara, R. W. Alexander, and M. A. Gimbrone. Myosin in cultured vascular smooth muscle cells: immunofluorescence and immunochemical studies of alterations in antigenic expression. Journal of Cell Biology, 99(5):1582– ??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/5/1582.

Lambotte:1980:PSD

[LFC^{+}80] P. Lambotte, P. Falmagne, C. Capiau, J. Zanen, J. M. Ruysschaert, and J. Dirkx. Primary structure of diphtheria toxin fragment B: structural similarities with lipid-binding domains. Journal of Cell Biology, 87(3):837–??, December 1980. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/87/3/837.

Luduena:1981:ETV

[LFF^{+}81] R. F. Ludueña, A. Fellous, J. Francon, J. Nunez, and L. Mc-Manus. Effect of tau on the vinblastine-induced aggregation of tubulin. Journal of Cell Biology, 89(3):680–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/89/3/680.

Lander:1982:CFP

[LFGR82] A. D. Lander, D. K. Fujii, D. Gospodarowicz, and L. F. Reichardt. Characterization of a factor that promotes neurite outgrowth: evidence linking activity to a heparan sulfate proteoglycan. Journal of Cell Biology, 94(3):574–??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/94/3/
574.

Lindemann:1980:SEN

[LFR80] C. B. Lindemann, I. Fentie, and R. Rikmenspoel. A selective effect of Ni2+ on wave initiation in bull sperm flagella. Journal of Cell Biology, 87(2):420–??, November 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/87/2/420.

Luna:1981:MCD

[LFS^{+}81] E. J. Luna, V. M. Fowler, J. Swanson, D. Branton, and D. L. Taylor. A membrane cytoskeleton from Dictyostelium discoideum. I. Identification and partial characterization of an actin-binding activity. Journal of Cell Biology, 88(2): 396–??, February 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/88/2/396.

Linsenmayer:1983:MAA

[LFS^{+}83] T. F. Linsenmayer, J. M. Fitch, T. M. Schmid, Zak, N. B., E. Gibney, R. D. Sanderson, and R. Mayne. Monoclonal antibodies against chicken type V collagen: production, specificity, and use for immunocytochemical localization in embryonic cornea and other organs. Journal of Cell Biology, 96(1): 124–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/1/124.

Lasek:1984:ATC

[LGB84] R. J. Lasek, J. A. Garner, and S. T. Brady. Axonal transport of the cytoplasmic matrix. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/1/
212s.

Linsenmayer:1984:TSH

[LGF^{+}84] T. F. Linsenmayer, E. Gibney, J. M. Fitch, J. Gross, and R. Mayne. Thermal stability of the helical structure of type IV collagen within basement membranes in situ: determination with a conformation-dependent monoclonal antibody. Journal of Cell Biology, 99(4):1405–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/4/1405.

Luna:1984:MCD

[LGHI84] E. J. Luna, C. M. Goodloe-Holland, and H. M. Ingalls. A membrane cytoskeleton from Dictyostelium discoideum. II. Integral proteins mediate the binding of plasma membranes to F-actin affinity beads. Journal of Cell Biology, 99(1):58–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/99/1/58.

Levitt:1983:ICS

[LH83] D. Levitt and P. L. Ho. Induction of chondroitin sulfate proteoglycan synthesis and secretion in lymphocytes and monocytes. Journal of Cell Biology, 97(2):351–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/2/351.

Lim:1984:RDE

[LH84] R. W. Lim and S. D. Hauschka. A rapid decrease in epidermal growth factor-binding capacity accompanies the terminal differentiation of mouse myoblasts in vitro. Journal of Cell Biology, 98(2):739–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/2/739.

Liscum:1982:EBT

[LHHH82] L. Liscum, P. J. Hauptman, D. C. Hood, and E. Holtzman. Effect of barium and tetraethylammonium on membrane circulation in frog retinal photoreceptors. Journal of Cell Biology, 95 (1):296–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/95/1/296.

Lange:1982:RRS

[LHS82] Y. Lange, R. A. Hadesman, and T. L. Steck. Role of the reticulum in the stability and shape of the isolated human erythrocyte membrane. Journal of Cell Biology, 92(3):714–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
92/3/714.

Lizardi:1980:IGS

[Liz80] P. M. Lizardi. Isolation of giant silk fibroin polysomes and fibroin mRNP particles using a novel ribonuclease inhibitor, hydroxystilbamidine. Journal of Cell Biology, 87(1):292–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
87/1/292.

Lanks:1983:DIG

[LK83] K. W. Lanks and E. J. Kasambalides. Dexamethasone induces gelsolin synthesis and altered morphology in L929 cells. Journal of Cell Biology, 96(2):577–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/2/577.

Lodish:1984:GRL

[LK84] H. F. Lodish and N. Kong. Glucose removal from N-linked oligosaccharides is required for efficient maturation of certain secretory glycoproteins from the rough endoplasmic reticulum to the Golgi complex. Journal of Cell Biology, 98(5):1720–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
98/5/1720.

Longenecker:1984:GIV

[LKJ84] J. P. Longenecker, L. A. Kilty, and L. K. Johnson. Glucocorticoid inhibition of vascular smooth muscle cell proliferation: influence of homologous extracellular matrix and serum mitogens. Journal of Cell Biology, 98(2):534–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/98/2/534.

Levine:1983:SPC

[LKR^{+}83] R. J. Levine, R. W. Kensler, M. C. Reedy, W. Hofmann, and H. A. King. Structure and paramyosin content of tarantula thick filaments. Journal of Cell Biology, 97(1):186–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/1/

Linck:1981:RFB

[LL81] R. W. Linck and G. L. Langevin. Reassembly of flagellar B (alpha beta) tubulin into singlet microtubules: consequences for cytoplasmic microtubule structure and assembly. Journal of Cell Biology, 89(2):323–??, May 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/89/2/323.

Lapolla:1982:MRA

[LL82a] R. J. Lapolla and A. M. Lambowitz. Mitochondrial ribosome assembly in Neurospora. Structural analysis of mature and partially assembled ribosomal subunits by equilibrium centrifugation in CsCl gradients. Journal of Cell Biology, 95(1): 267–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/95/1/267.

Levin:1982:CBE

[LL82b] E. G. Levin and D. J. Loskutoff. Cultured bovine endothelial cells produce both urokinase and tissue-type plasminogen activators. Journal of Cell Biology, 94(3):631–??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/94/3/
631.

Lyles:1984:BDCb

[LLB84] J. M. Lyles, D. Linnemann, and E. Bock. Biosynthesis of the D2-cell adhesion molecule: post-translational modifications, intracellular transport, and developmental changes. Journal of Cell Biology, 99(6):2082–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/6/2082.

Laurie:1982:LTI

[LLM82] G. W. Laurie, C. P. Leblond, and G. R. Martin. Localization of type IV collagen, laminin, heparan sulfate proteoglycan, and fibronectin to the basal lamina of basement membranes. Journal of Cell Biology, 95(1):340–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/95/1/340.

Leterrier:1981:PPM

[LLS81] J. F. Leterrier, R. K. Liem, and M. L. Shelanski. Preferential phosphorylation of the 150,000 molecular weight component of neurofilaments by a cyclic AMP-dependent, microtubuleassociated protein kinase. Journal of Cell Biology, 90(3): 755–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/3/755.

Leterrier:1982:IBN

[LLS82] J. F. Leterrier, R. K. Liem, and M. L. Shelanski. Interactions between neurofilaments and microtubule-associated proteins: a possible mechanism for intraorganellar bridging. Journal of Cell Biology, 95(3):982–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/95/3/982.

Lohka:1984:RCC

[LM84] M. J. Lohka and Y. Masui. Roles of cytosol and cytoplasmic particles in nuclear envelope assembly and sperm pronuclear formation in cell-free preparations from amphibian eggs. Journal of Cell Biology, 98(4):1222–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/4/1222.

Leven:1983:RSA

[LMN83] R. M. Leven, W. H. Mullikin, and V. T. Nachmias. Role of sodium in ADP- and thrombin-induced megakaryocyte spreading. Journal of Cell Biology, 96(5):1234–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/5/
1234.

Lin:1984:TEA

[LMYM84] J. J. Lin, F. Matsumura, and S. Yamashiro-Matsumura. Tropomyosin-enriched and alpha-actinin-enriched microfilaments isolated from chicken embryo fibroblasts by monoclonal antibodies. Journal of Cell Biology, 98(1):116–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/1/

Lewis:1981:BSE

[LN81] R. M. Lewis and D. L. Nelson. Biochemical studies of the excitable membrane of Paramecium tetraurelia VI. Endogenous protein substrates for in vitro and in vivo phosphorylation in cilia and ciliary membranes. Journal of Cell Biology, 91(1): 167–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/91/1/167.

Leven:1982:CMC

[LN82] R. M. Leven and V. T. Nachmias. Cultured megakaryocytes: changes in the cytoskeleton after ADP-induced spreading. Journal of Cell Biology, 92(2):313–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/92/2/313.

Levine:1981:HIF

[LNA81] J. S. Levine, P. K. Nakane, and R. H. Allen. Human intrinsic factor secretion: immunocytochemical demonstration of membrane-associated vesicular transport in parietal cells. Journal of Cell Biology, 90(3):644–??, September 1981. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/90/3/644.

Lyles:1984:BDCa

[LNB84] J. M. Lyles, B. Norrild, and E. Bock. Biosynthesis of the D2 cell adhesion molecule: pulse-chase studies in cultured fetal rat neuronal cells. Journal of Cell Biology, 98(6):2077–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/6/
2077.

Liau:1981:IRC

[LOC81] G. Liau, D. E. Ong, and F. Chytil. Interaction of the retinol/ cellular retinol-binding protein complex with isolated nuclei and nuclear components. Journal of Cell Biology, 91(1): 63–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/91/1/63.

Linck:1981:ATS

[LOL81] R. W. Linck, G. E. Olson, and G. L. Langevin. Arrangement of tubulin subunits and microtubule-associated proteins in the central-pair microtubule apparatus of squid (Loligo pealei) sperm flagella. Journal of Cell Biology, 89(2):309–??, May 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/89/2/
309.

Langmore:1983:LAX

[LP83a] J. P. Langmore and Paulson. Low angle X-ray diffraction studies of chromatin structure in vivo and in isolated nuclei and metaphase chromosomes. Journal of Cell Biology, 96 (4):1120–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/4/1120.

Leslie:1983:IMS

[LP83b] B. A. Leslie and J. W. Putney. Ionic mechanisms in secretagogue-induced morphological changes in rat parotid gland. Journal of Cell Biology, 97(4):1119–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/4/
1119.

Lee:1984:MSP

[LPB84] E. Y. Lee, G. Parry, and M. J. Bissell. Modulation of secreted proteins of mouse mammary epithelial cells by the collagenous substrata. Journal of Cell Biology, 98(1):146–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/1/
146.

Leslie:1983:UMI

[LPH83] R. J. Leslie and J. D. Pickett-Heaps. Ultraviolet microbeam irradiations of mitotic diatoms: investigation of spindle elongation. Journal of Cell Biology, 96(2):548–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/96/2/548.

Lin:1982:MAR

[LQ82] J. J. Lin and S. A. Queally. A monoclonal antibody that recognizes Golgi-associated protein of cultured fibroblast cells. Journal of Cell Biology, 92(1):108–??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/92/1/108.

Landis:1981:AMS

[LR81] D. M. Landis and T. S. Reese. Astrocyte membrane structure: changes after circulatory arrest. Journal of Cell Biology, 88 (3):660–??, March 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/88/3/660.

Landis:1983:COC

[LR83a] D. M. Landis and T. S. Reese. Cytoplasmic organization in cerebellar dendritic spines. Journal of Cell Biology, 97(4): 1169–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/4/1169.

LHernault:1983:CAT

[LR83b] S. W. L’Hernault and J. L. Rosenbaum. Chlamydomonas alpha-tubulin is posttranslationally modified in the flagella during flagellar assembly. Journal of Cell Biology, 97(1): 258–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/1/258.

Letourneau:1984:INI

[LR84] P. C. Letourneau and A. H. Ressler. Inhibition of neurite initiation and growth by taxol. Journal of Cell Biology, 98 (4):1355–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/4/1355.

Lemmon:1982:BFG

[LRT^{+}82] S. K. Lemmon, M. C. Riley, K. A. Thomas, G. A. Hoover, T. Maciag, and R. A. Bradshaw. Bovine fibroblast growth factor: comparison of brain and pituitary preparations. Journal of Cell Biology, 95(1):162–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/95/1/162.

Louvard:1982:AGC

[LRW82] D. Louvard, H. Reggio, and G. Warren. Antibodies to the Golgi complex and the rough endoplasmic reticulum. Journal of Cell Biology, 92(1):92–??, January 1982. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/92/1/92.

Lubit:1980:AAD

[LS80] B. W. Lubit and J. H. Schwartz. An antiactin antibody that distinguishes between cytoplasmic and skeletal muscle actins. Journal of Cell Biology, 86(3):891–??, September 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/86/3/891.

Lee:1981:OIP

[LS81a] S. C. Lee and R. A. Steinhardt. Observations on intracellular pH during cleavage of eggs of Xenopus laevis. Journal of Cell Biology, 91(2):414–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/91/2/414.

Lingwood:1981:LSS

[LS81b] C. Lingwood and H. Schachter. Localization of sulfatoxygalactosylacylalkylglycerol at the surface of rat testicular germinal cells by immunocytochemical techniques: pH dependence of a nonimmunological reaction between immunoglobulin and germinal cells. Journal of Cell Biology, 89(3):621–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/89/3/621.

Larson:1982:IJT

[LS82] D. M. Larson and J. D. Sheridan. Intercellular junctions and transfer of small molecules in primary vascular endothelial cultures. Journal of Cell Biology, 92(1):183–??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/92/1/183.

Landis:1984:RVB

[LSBPG84] W. J. Landis, C. F. Sanzone, D. Brickley-Parsons, and M. J. Glimcher. Radioautographic visualization and biochemical identification of O-phosphoserine- and O-phosphothreonine-containing phosphoproteins in mineralizing embryonic chick bone. Journal of Cell Biology, 98(3):986–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/98/3/986.

Laterra:1983:CSH

[LSC83] J. Laterra, J. E. Silbert, and L. A. Culp. Cell surface heparan sulfate mediates some adhesive responses to glycosaminoglycan-binding matrices, including fibronectin. Journal of Cell Biology, 96(1):112–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/1/112.

Lemasters:1983:CSC

[LSJT83] J. J. Lemasters, C. J. Stemkowski, S. Ji, and R. G. Thurman. Cell surface changes and enzyme release during hypoxia and reoxygenation in the isolated, perfused rat liver. Journal of Cell Biology, 97(3):778–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/3/778.

LaVail:1983:LAT

[LSM83] J. H. LaVail, I. K. Sugino, and D. M. McDonald. Localization of axonally transported 125I-wheat germ agglutinin beneath the plasma membrane of chick retinal ganglion cells. Journal of Cell Biology, 96(2):373–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/2/373.

Leslie:1984:APF

[LSM^{+}84] R. J. Leslie, W. M. Saxton, T. J. Mitchison, B. Neighbors, E. D. Salmon, and J. R. McIntosh. Assembly properties of fluorescein-labeled tubulin in vitro before and after fluorescence bleaching. Journal of Cell Biology, 99(6):2146–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/6/2146.

Larsson:1984:DUT

[LSN^{+}84] L. I. Larsson, L. Skriver, L. S. Nielsen, J. Grøndahl-Hansen, P. Kristensen, and K. Danø. Distribution of urokinase-type plasminogen activator immunoreactivity in the mouse. Journal of Cell Biology, 98(3):894–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/3/894.

Lefort-Tran:1981:CEP

[LTAP^{+}81] M. Lefort-Tran, K. Aufderheide, M. Pouphile, M. Rossignol, and J. Beisson. Control of exocytotic processes: cytological and physiological studies of trichocyst mutants in Paramecium tetraurelia. Journal of Cell Biology, 88(2):301–??, February 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/88/2/
301.

Laug:1980:AIG

[LTBS80] W. E. Laug, Z. A. Tokes, W. F. Benedict, and N. Sorgente. Anchorage independent growth and plasminogen activator production by bovine endothelial cells. Journal of Cell Biology, 84 (2):281–??, February 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/84/2/281.

Landschulz:1984:LIC

[LTE84] W. Landschulz, I. Thesleff, and P. Ekblom. A lipophilic iron chelator can replace transferrin as a stimulator of cell proliferation and differentiation. Journal of Cell Biology, 98(2): 596–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/2/596.

Lin:1980:CIN

[LTGL80] D. C. Lin, K. D. Tobin, M. Grumet, and S. Lin. Cytochalasins inhibit nuclei-induced actin polymerization by blocking filament elongation. Journal of Cell Biology, 84(2):455–??, February 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
84/2/455.

Luck:1984:GBD

[Luc84] D. J. Luck. Genetic and biochemical dissection of the eucaryotic flagellum. Journal of Cell Biology, 98(3):789–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/3/
789.

Ledger:1980:ILP

[LUT80] P. W. Ledger, N. Uchida, and M. L. Tanzer. Immunocytochemical localization of procollagen and fibronectin in human fibroblasts: effects of the monovalent ionophore, monensin. Journal of Cell Biology, 87(3):663–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/87/3/663.

Levinger:1981:HSP

[LV81] L. Levinger and A. Varshavsky. Heat-shock proteins of Drosophila are associated with nuclease-resistant, high-salt-resistant nuclear structures. Journal of Cell Biology, 90(3): 793–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/3/793.

Lehto:1983:INC

[LV83] V. P. Lehto and I. Virtanen. Immunolocalization of a novel, cytoskeleton-associated polypeptide of Mr 230,000 daltons (p230). Journal of Cell Biology, 96(3):703–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/96/3/703.

Laurila:1982:EDV

[LVL^{+}82] P. Laurila, I. Virtanen, V. P. Lehto, T. Vartio, and S. Stenman. Expression and distribution of vimentin and keratin filaments in heterokaryons of human fibroblasts and amnion epithelial cells. Journal of Cell Biology, 94(2):308–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/94/2/
308.

Levine:1981:FAT

[LW81] J. Levine and M. Willard. Fodrin: axonally transported polypeptides associated with the internal periphery of many cells. Journal of Cell Biology, 90(3):631–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/90/3/631.

LaFond:1983:GIM

[LWW^{+}83] R. E. LaFond, H. Woodcock, C. L. Woodcock, E. R. Kundahl, and J. J. Lucas. Generation of an internal matrix in mature avian erythrocyte nuclei during reactivation in cytoplasts. Journal of Cell Biology, 96(6):1815–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/6/1815.

Lew:1984:MCF

[LWWP84] P. D. Lew, C. B. Wollheim, F. A. Waldvogel, and T. Pozzan. Modulation of cytosolic-free calcium transients by changes in intracellular calcium-buffering capacity: correlation with exocytosis and O2-production in human neutrophils. Journal of Cell Biology, 99(4):1212–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/4/1212.

Lundell:1981:TEA

[LYG81] D. J. Lundell, G. Yamanaka, and A. N. Glazer. A terminal energy acceptor of the phycobilisome: the 75,000-dalton polypeptide of Synechococcus 6301 phycobilisomes — a new biliprotein. Journal of Cell Biology, 91(1):315–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/91/1/
315.

Linder:1982:PCG

[LZR82] S. Linder, S. H. Zuckerman, and N. R. Ringertz. Pattern of chick gene activation in chick erythrocyte heterokaryons. Journal of Cell Biology, 95(3):885–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/95/3/885.

Mimura:1982:CLA

[MA82] N. Mimura and A. Asano. Characterization and localization of actinogelin, a Ca^{2+}-sensitive actin accessory protein, in nonmuscle cells. Journal of Cell Biology, 93(3):899–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/93/3/899.

McGookey:1983:MCC

[MA83a] D. J. McGookey and R. G. Anderson. Morphological characterization of the cholesteryl ester cycle in cultured mouse macrophage foam cells. Journal of Cell Biology, 97(4):1156– ??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/4/1156.

Miyata:1983:BRS

[MA83b] S. Miyata and T. Akazawa. Biosynthesis of rice seed alphaamylase: proteolytic processing and glycosylation of precursor polypeptides by microsomes. Journal of Cell Biology, 96 (3):802–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/3/802.

Mabuchi:1983:ADP

[Mab83] I. Mabuchi. An actin-depolymerizing protein (depactin) from starfish oocytes: properties and interaction with actin. Journal of Cell Biology, 97(5):1612–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/5/1612.

Marcantonio:1984:SPT

[MACK84] E. E. Marcantonio, A. Amar-Costesec, and G. Kreibich. Segregation of the polypeptide translocation apparatus to regions of the endoplasmic reticulum containing ribophorins and ribosomes. II. Rat liver microsomal subfractions contain equimolar amounts of ribophorins and ribosomes. Journal of Cell Biology, 99(6):2254–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/6/2254.

Madara:1983:IGP

[Mad83] J. L. Madara. Increases in guinea pig small intestinal transepithelial resistance induced by osmotic loads are accompanied by rapid alterations in absorptive-cell tight-junction structure. Journal of Cell Biology, 97(1):125–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/1/125.

Marzella:1982:IAV

[MAG82] L. Marzella, J. Ahlberg, and H. Glaumann. Isolation of autophagic vacuoles from rat liver: morphological and biochemical characterization. Journal of Cell Biology, 93(1):144–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
93/1/144.

Mannella:1982:SOM

[Man82] C. A. Mannella. Structure of the outer mitochondrial membrane: ordered arrays of porelike subunits in outer-membrane fractions from neurospora crassa mitochondria. Journal of Cell Biology, 94(3):680–??, September 1982. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/94/3/680.

Martz:1980:ILT

[Mar80] E. Martz. Immune T lymphocyte to tumor cell adhesion. Magnesium sufficient, calcium insufficient. Journal of Cell Biology, 84(3):584–??, March 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/84/3/584.

Masters:1984:IBG

[Mas84] C. Masters. Interactions between glycolytic enzymes and components of the cytomatrix. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/1/222s.

Maxfield:1982:WBI

[Max82] F. R. Maxfield. Weak bases and ionophores rapidly and reversibly raise the pH of endocytic vesicles in cultured mouse fibroblasts. Journal of Cell Biology, 95(2):676–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/95/2/
676.

Marty:1980:ACB

[MB80a] F. Marty and D. Branton. Analytical characterization of beetroot vacuole membrane. Journal of Cell Biology, 87(1): 72–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/87/1/72.

Mueller:1980:EIC

[MB80b] S. C. Mueller and R. M. Brown. Evidence for an intramembrane component associated with a cellulose microfibrilsynthesizing complex in higher plants. Journal of Cell Biology, 84(2):315–??, February 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/84/2/315.

McNeill:1981:CBA

[MB81] P. A. McNeill and M. W. Berns. Chromosome behavior after laser microirradiation of a single kinetochore in mitotic PtK2 cells. Journal of Cell Biology, 88(3):543–??, March 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/88/3/543.

Matsudaira:1982:OCF

[MB82a] P. T. Matsudaira and D. R. Burgess. Organization of the crossfilaments in intestinal microvilli. Journal of Cell Biology, 92 (3):657–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/3/657.

Matsudaira:1982:PRM

[MB82b] P. T. Matsudaira and D. R. Burgess. Partial reconstruction of the microvillus core bundle: characterization of villin as a Ca(++)-dependent, actin-bundling/depolymerizing protein. Journal of Cell Biology, 92(3):648–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/92/3/648.

Miao:1982:GPA

[MB82c] P. Miao and V. H. Black. Guinea pig adrenocortical cells: in vitro characterization of separated zonal cell types. Journal of Cell Biology, 94(2):241–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/94/2/241.

Majack:1984:HRG

[MB84a] R. A. Majack and P. Bornstein. Heparin and related glycosaminoglycans modulate the secretory phenotype of vascular smooth muscle cells. Journal of Cell Biology, 99(5): 1688–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/5/1688.

Mangeat:1984:AMI

[MB84b] P. Mangeat and K. Burridge. Actin-membrane interaction in fibroblasts: what proteins are involved in this association? Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/1/95s.

Mangeat:1984:INS

[MB84c] P. H. Mangeat and K. Burridge. Immunoprecipitation of nonerythrocyte spectrin within live cells following microinjection of specific antibodies: relation to cytoskeletal structures. Journal of Cell Biology, 98(4):1363–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/4/1363.

Mueller:1984:ICV

[MB84d] S. C. Mueller and D. Branton. Identification of coated vesicles in Saccharomyces cerevisiae. Journal of Cell Biology, 98(1): 341–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/1/341.

Mole-Bajer:1983:ATM

[MBB83] J. Molè-Bajer and A. S. Bajer. Action of taxol on mitosis: modification of microtubule arrangements and function of the mitotic spindle in Haemanthus endosperm. Journal of Cell Biology, 96(2):527–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/2/527.

Meeusen:1980:EME

[MBC80a] R. L. Meeusen, J. Bennett, and W. Z. Cande. Effect of microinjected N-ethylmaleimide-modified heavy meromyosin on cell division in amphibian eggs. Journal of Cell Biology, 86(3): 858–??, September 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/86/3/858.

Molano:1980:DCY

[MBC80b] J. Molano, B. Bowers, and E. Cabib. Distribution of chitin in the yeast cell wall. An ultrastructural and chemical study. Journal of Cell Biology, 85(2):199–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/85/2/199.

Milks:1983:EPT

[MBC83] L. C. Milks, M. J. Brontoli, and E. B. Cramer. Epithelial permeability and the transepithelial migration of human neutrophils. Journal of Cell Biology, 96(5):1241–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/5/
1241.

Mooseker:1984:BBC

[MBC^{+}84] M. S. Mooseker, E. M. Bonder, K. A. Conzelman, D. J. Fishkind, C. L. Howe, and T. C. Keller. Brush border cytoskeleton and integration of cellular functions. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/1/104s.

Miljanich:1982:PPP

[MBK82] G. P. Miljanich, A. R. Brasier, and R. B. Kelly. Partial purification of presynaptic plasma membrane by immunoadsorption. Journal of Cell Biology, 94(1):88–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/94/1/88.

Mollenhauer:1984:RAC

[MBLvdM84] J. Mollenhauer, J. A. Bee, M. A. Lizarbe, and K. von der Mark. Role of anchorin CII, a 31,000-mol-wt membrane protein, in the interaction of chondrocytes with type II collagen. Journal of Cell Biology, 98(4):1572–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/4/1572.

Matlin:1983:TTV

[MBP^{+}83] K. Matlin, D. F. Bainton, M. Pesonen, D. Louvard, N. Genty, and K. Simons. Transepithelial transport of a viral membrane glycoprotein implanted into the apical plasma membrane of Madin–Darby canine kidney cells. I. Morphological evidence. Journal of Cell Biology, 97(3):627–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/3/
627.

Mehrabian:1984:IRL

[MBR84] M. Mehrabian, K. J. Bame, and L. H. Rome. Interaction of rat liver lysosomal membranes with actin. Journal of Cell Biology, 99(2):680–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/2/680.

Marsh:1983:ISF

[MBWH83] M. Marsh, E. Bolzau, J. White, and A. Helenius. Interactions of Semliki Forest virus spike glycoprotein rosettes and vesicles with cultured cells. Journal of Cell Biology, 96(2): 455–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/2/455.

Mannino:1981:IAC

[MBZB81] R. J. Mannino, K. Ballmer, D. Zeltner, and M. M. Burger. An inhibitor of animal cell growth increases cell-to-cell adhesion. Journal of Cell Biology, 91(3):855–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/91/3/855.

Mori:1980:MAL

[MC80] H. Mori and A. K. Christensen. Morphometric analysis of Leydig cells in the normal rat testis. Journal of Cell Biology, 84 (2):340–??, February 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/84/2/340.

Munson:1982:MCS

[MCG82] R. Munson, K. L. Caldwell, and L. Glaser. Multiple controls for the synthesis of muscle-specific proteins in BC3H1 cells. Journal of Cell Biology, 92(2):350–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/92/2/350.

McIntosh:1981:P

[McI81] R. McIntosh. Preface I. Journal of Cell Biology, 91(3):??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
91/3/v.

Marchisio:1984:CSI

[MCN^{+}84] P. C. Marchisio, D. Cirillo, L. Naldini, M. V. Primavera, A. Teti, and A. Zambonin-Zallone. Cell-substratum interaction of cultured avian osteoclasts is mediated by specific adhesion structures. Journal of Cell Biology, 99(5):1696– ??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/5/1696.

Meldolesi:1980:LBN

[MCPB80] J. Meldolesi, G. Corte, G. Pietrini, and N. Borgese. Localization and biosynthesis of NADH-cytochrome b5 reductase, an integral membrane protein, in rat liver cells. II. Evidence that a single enzyme accounts for the activity in its various subcellular locations. Journal of Cell Biology, 85(3):516–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/85/3/
516.

Meyer:1980:ICM

[MD80a] D. I. Meyer and B. Dobberstein. Identification and characterization of a membrane component essential for the translocation of nascent proteins across the membrane of the endoplasmic reticulum. Journal of Cell Biology, 87(2):503–??, November 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
87/2/503.

Meyer:1980:MCE

[MD80b] D. I. Meyer and B. Dobberstein. A membrane component essential for vectorial translocation of nascent proteins across the endoplasmic reticulum: requirements for its extraction and reassociation with the membrane. Journal of Cell Biology, 87(2): 498–??, November 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/87/2/498.

Malawista:1982:CPS

[MD82] S. E. Malawista and A. De Boisfleury Chevance. The cytokineplast: purified, stable, and functional motile machinery from human blood polymorphonuclear leukocytes. Journal of Cell Biology, 95(3):960–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/95/3/960.

Murphy:1983:GCP

[MDA83] T. L. Murphy, G. Decker, and J. T. August. Glycoproteins of coated pits, cell junctions, and the entire cell surface revealed by monoclonal antibodies and immunoelectron microscopy. Journal of Cell Biology, 97(2):533–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/2/533.

Mosher:1982:SST

[MDJ82] D. F. Mosher, M. J. Doyle, and E. A. Jaffe. Synthesis and secretion of thrombospondin by cultured human endothelial cells. Journal of Cell Biology, 93(2):343–??, May 1982. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/93/2/343.

Maihle:1981:PII

[MDM^{+}81] N. J. Maihle, J. R. Dedman, A. R. Means, J. G. Chafouleas, and B. H. Satir. Presence and indirect immunofluorescent localization of calmodulin in Paramecium tetraurelia. Journal of Cell Biology, 89(3):695–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/89/3/695.

Mroczkowski:1980:UUM

[MDSH80] B. Mroczkowski, H. P. Dym, E. J. Siegel, and S. M. Heywood. Uptake and utilization of mRNA by myogenic cells in culture. Journal of Cell Biology, 87(1):65–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/87/1/65.

McCarthy:1980:PSA

[MdV80] K. D. McCarthy and J. de Vellis. Preparation of separate astroglial and oligodendroglial cell cultures from rat cerebral tissue. Journal of Cell Biology, 85(3):890–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/85/3/890.

Mains:1981:CES

[ME81] R. E. Mains and B. A. Eipper. Coordinate, equimolar secretion of smaller peptide products derived from pro-ACTH/ endorphin by mouse pituitary tumor cells. Journal of Cell Biology, 89(1):21–??, April 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/89/1/21.

Milam:1982:VNA

[ME82] L. Milam and H. P. Erickson. Visualization of a 21-nm axial periodicity in shadowed keratin filaments and neurofilaments. Journal of Cell Biology, 94(3):592–??, September 1982.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/94/3/592.

McIntosh:1984:THP

[ME84] J. R. McIntosh and U. Euteneuer. Tubulin hooks as probes for microtubule polarity: an analysis of the method and an evaluation of data on microtubule polarity in the mitotic spindle. Journal of Cell Biology, 98(2):525–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/2/525.

Mechler:1981:MBRb

[Mec81a] B. Mechler. Membrane-bound ribosomes of myeloma cells. V. Subcellular distribution of immunoglobulin mRNA molecules. Journal of Cell Biology, 88(1):37–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/88/1/37.

Mechler:1981:MBRc

[Mec81b] B. Mechler. Membrane-bound ribosomes of myeloma cells. VI. Initiation of immunoglobulin mRNA translation occurs on free ribosomes. Journal of Cell Biology, 88(1):42–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/88/1/42.

Matsuura:1984:FIM

[MEKT84] S. Matsuura, S. Eto, K. Kato, and Y. Tashiro. Ferritin immunoelectron microscopic localization of 5’-nucleotidase on rat liver cell surface. Journal of Cell Biology, 99(1):166–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/1/
166.

Manthorpe:1983:LPN

[MER^{+}83] M. Manthorpe, E. Engvall, E. Ruoslahti, F. M. Longo, G. E. Davis, and S. Varon. Laminin promotes neuritic regeneration from cultured peripheral and central neurons. Journal of Cell Biology, 97(6):1882–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/97/6/1882.

McBeath:1984:IFI

[MF84a] E. McBeath and K. Fujiwara. Improved fixation for immunofluorescence microscopy using light-activated 1,3,5triazido-2,4,6-trinitrobenzene (TTB). Journal of Cell Biology, 99(6):2061–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/6/2061.

McCarthy:1984:LFP

[MF84b] J. B. McCarthy and L. T. Furcht. Laminin and fibronectin promote the haptotactic migration of B16 mouse melanoma cells in vitro. Journal of Cell Biology, 98(4):1474–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/4/
1474.

Morejohn:1984:TIR

[MF84c] L. C. Morejohn and D. E. Fosket. Taxol-induced rose microtubule polymerization in vitro and its inhibition by colchicine. Journal of Cell Biology, 99(1):141–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/1/141.

Morimoto:1984:CSE

[MF84d] R. Morimoto and E. Fodor. Cell-specific expression of heat shock proteins in chicken reticulocytes and lymphocytes. Journal of Cell Biology, 99(4):1316–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/4/1316.

Murray:1984:HII

[MF84e] S. A. Murray and W. H. Fletcher. Hormone-induced intercellular signal transfer dissociates cyclic AMP-dependent protein kinase. Journal of Cell Biology, 98(5):1710–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/5/
1710.

McGookey:1983:FCC

[MFA83] D. J. McGookey, K. Fagerberg, and R. G. Anderson. Filipincholesterol complexes form in uncoated vesicle membrane derived from coated vesicles during receptor-mediated endocytosis of low density lipoprotein. Journal of Cell Biology, 96 (5):1273–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/5/1273.

MacLean-Fletcher:1980:VAG

[MFP80] S. D. MacLean-Fletcher and T. D. Pollard. Viscometric analysis of the gelation of Acanthamoeba extracts and purification of two gelation factors. Journal of Cell Biology, 85(2):414–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
85/2/414.

Merisko:1982:CVI

[MFP82] E. M. Merisko, M. G. Farquhar, and G. E. Palade. Coated vesicle isolation by immunoadsorption on Staphylococcus aureus cells. Journal of Cell Biology, 92(3):846–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/92/3/846.

Moll:1982:DKP

[MFVPK82] R. Moll, W. W. Franke, B. Volc-Platzer, and R. Krepler. Different keratin polypeptides in epidermis and other epithelia of human skin: a specific cytokeratin of molecular weight 46,000 in epithelia of the pilosebaceous tract and basal cell epitheliomas. Journal of Cell Biology, 95(1):285–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/95/1/285.

McDonald:1981:GRA

[MG81] R. A. McDonald and T. D. Gelehrter. Glucocorticoid regulation of amino acid transport in anucleate rat hepatoma (HTC) cells. Journal of Cell Biology, 88(3):536–??, March 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/88/3/
536.

McAbee:1983:FMB

[MG83a] D. D. McAbee and F. Grinnell. Fibronectin-mediated binding and phagocytosis of polystyrene latex beads by baby hamster kidney cells. Journal of Cell Biology, 97(5):1515–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/5/

Murti:1983:IFV

[MG83b] K. G. Murti and R. Goorha. Interaction of frog virus-3 with the cytoskeleton. I. Altered organization of microtubules, intermediate filaments, and microfilaments. Journal of Cell Biology, 96(5):1248–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/5/1248.

Moss:1983:EET

[MGG83] R. L. Moss, G. G. Giulian, and M. L. Greaser. Effects of EDTA treatment upon the protein subunit composition and mechanical properties of mammalian single skeletal muscle fibers. Journal of Cell Biology, 96(4):970–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/4/970.

Moore:1983:SPS

[MGK83] H. P. Moore, B. Gumbiner, and R. B. Kelly. A subclass of proteins and sulfated macromolecules secreted by AtT-20 (mouse pituitary tumor) cells is sorted with adrenocorticotropin into dense secretory granules. Journal of Cell Biology, 97(3):810– ??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/3/810.

Miller:1983:CKL

[MGLF83] D. K. Miller, E. Griffiths, J. Lenard, and R. A. Firestone. Cell killing by lysosomotropic detergents. Journal of Cell Biology, 97(6):1841–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/6/1841.

Mecham:1984:ACR

[MGMS84] R. P. Mecham, G. L. Griffin, J. G. Madaras, and R. M. Senior. Appearance of chemotactic responsiveness to elastin peptides by developing fetal bovine ligament fibroblasts parallels the onset of elastin production. Journal of Cell Biology, 98 (5):1813–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/5/1813.

McCue:1984:SBI

[MGSC84] P. A. McCue, M. L. Gubler, M. I. Sherman, and B. N. Cohen. Sodium butyrate induces histone hyperacetylation and differentiation of murine embryonal carcinoma cells. Journal of Cell Biology, 98(2):602–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/2/602.

Mooseker:1980:RMS

[MGW^{+}80] M. S. Mooseker, T. A. Graves, K. A. Wharton, N. Falco, and C. L. Howe. Regulation of microvillus structure: calciumdependent solation and cross-linking of actin filaments in the microvilli of intestinal epithelial cells. Journal of Cell Biology, 87(3):809–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/87/3/809.

Miller:1984:ESV

[MH84] T. M. Miller and J. E. Heuser. Endocytosis of synaptic vesicle membrane at the frog neuromuscular junction. Journal of Cell Biology, 98(2):685–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/2/685.

Menon:1984:SOI

[MHB84] A. K. Menon, D. Holowka, and B. Baird. Small oligomers of immunoglobulin E (IgE) cause large-scale clustering of IgE receptors on the surface of rat basophilic leukemia cells. Journal of Cell Biology, 98(2):577–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/2/577.

Melchior:1980:SLV

[MHC80] V. Melchior, C. J. Hollingshead, and M. E. Cahoon. Stacking in lipid vesicle-tubulin mixtures is an artifact of negative staining. Journal of Cell Biology, 86(3):881–??, September 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/86/3/
881.

Mesland:1980:FTA

[MHCG80] D. A. Mesland, J. L. Hoffman, E. Caligor, and U. W. Goodenough. Flagellar tip activation stimulated by membrane adhesions in Chlamydomonas gametes. Journal of Cell Biology, 84 (3):599–??, March 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/84/3/599.

Martinez-Hernandez:1982:ULT

[MHGM82] A. Martinez-Hernandez, S. Gay, and E. J. Miller. Ultrastructural localization of type V collagen in rat kidney. Journal of Cell Biology, 92(2):343–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/92/2/343.

McGuffee:1981:ASS

[MHLS81] L. J. McGuffee, L. Hurwitz, S. A. Little, and B. E. Skipper. A 45Ca autoradiographic and stereological study of freeze-dried smooth muscle of the guinea pig vas deferens. Journal of Cell Biology, 90(1):201–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/90/1/201.

Moroi:1981:DKC

[MHNT81] Y. Moroi, A. L. Hartman, P. K. Nakane, and E. M. Tan. Distribution of kinetochore (centromere) antigen in mammalian cell nuclei. Journal of Cell Biology, 90(1):254–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/90/1/254.

McGarry:1983:IDD

[MHRR83] T. McGarry, R. Hough, S. Rogers, and M. Rechsteiner. Intracellular distribution and degradation of immunoglobulin G and immunoglobulin G fragments injected into HeLa cells. Journal of Cell Biology, 96(2):338–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/2/338.

Maciag:1981:SPH

[MHSW81] T. Maciag, G. A. Hoover, M. B. Stemerman, and R. Weinstein. Serial propagation of human endothelial cells in vitro. Journal of Cell Biology, 91(2):420–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/91/2/420.

Murphy:1983:IOAa

[MHW83] D. B. Murphy, R. R. Hiebsch, and K. T. Wallis. Identity and Origin of the ATPase activity associated with neuronal microtubules. I. The ATPase activity is associated with membrane vesicles. Journal of Cell Biology, 96(5):1298–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/5/
1298.

Miller:1980:RGP

[Mil80] D. K. Miller. Reconstituted G protein-lipid vesicles from vesicular stomatitus virus and their inhition of VSV infection. Journal of Cell Biology, 84(2):421–??, February 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/84/2/421.

Miller:1981:NCV

[Mil81a] O. L. Miller. The nucleolus, chromosomes, and visualization of genetic activity. Journal of Cell Biology, 91(3):??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/91/3/
15s.

Milstone:1981:ICT

[Mil81b] L. M. Milstone. Isolation and characterization of two polypeptides that form intermediate filaments in bovine esophageal epithelium. Journal of Cell Biology, 88(2):317–??, February 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/88/2/
317.

Merlie:1984:DSS

[MIRS84] J. P. Merlie, K. E. Isenberg, S. D. Russell, and J. R. Sanes. Denervation supersensitivity in skeletal muscle: analysis with a cloned cDNA probe. Journal of Cell Biology, 99(1):332–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/1/332.

Meza:1980:OJC

[MIS^{+}80] I. Meza, G. Ibarra, M. Sabanero, A. Martinez-Palomo, and M. Cereijido. Occluding junctions and cytoskeletal components in a cultured transporting epithelium. Journal of Cell Biology, 87(3):746–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/87/3/746.

Miller:1983:TDC

[MJ83] K. R. Miller and J. S. Jacob. Two-dimensional crystals formed from photosynthetic reaction centers. Journal of Cell Biology, 97(4):1266–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/4/1266.

Manney:1983:TTS

[MJM83] T. R. Manney, P. Jackson, and J. Meade. Two temperaturesensitive mutants of Saccharomyces cerevisiae with altered expression of mating-type functions. Journal of Cell Biology, 96 (6):1592–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/6/1592.

Mastro:1984:DAC

[MK84a] A. M. Mastro and A. D. Keith. Diffusion in the aqueous compartment. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/1/
180s.

McDonald:1984:SBF

[MK84b] J. A. McDonald and D. G. Kelley. Specific binding of fibronectin–antifibronectin immune complexes to procollagen: a new pitfall in immunostaining. Journal of Cell Biology, 98 (3):1042–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/3/1042.

Melmed:1981:CCV

[MKB81] R. N. Melmed, P. J. Karanian, and R. D. Berlin. Control of cell volume in the J774 macrophage by microtubule disassembly and cyclic AMP. Journal of Cell Biology, 90(3): 761–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/3/761.

McDonald:1982:RFC

[MKB82] J. A. McDonald, D. G. Kelley, and T. J. Broekelmann. Role of fibronectin in collagen deposition: Fab’ to the gelatin-binding domain of fibronectin inhibits both fibronectin and collagen organization in fibroblast extracellular matrix. Journal of Cell Biology, 92(2):485–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/92/2/485.

Muller:1980:EAS

[MKD80] J. Muller, W. A. Kachadorian, and V. A. DiScala. Evidence that ADH-stimulated intramembrane particle aggregates are transferred from cytoplasmic to luminal membranes in toad bladder epithelial cells. Journal of Cell Biology, 85(1):83–??, April 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
85/1/83.

Mangelsdorf:1984:DDI

[MKD^{+}84] D. J. Mangelsdorf, H. P. Koeffler, C. A. Donaldson, J. W. Pike, and M. R. Haussler. 1,25-dihydroxyvitamin D3-induced differentiation in a human promyelocytic leukemia cell line (HL-60): receptor-mediated maturation to macrophage-like cells. Journal of Cell Biology, 98(2):391–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/2/391.

Merkel:1981:CCC

[MKG81] S. J. Merkel, E. S. Kaneshiro, and E. I. Gruenstein. Characterization of the cilia and ciliary membrane proteins of wild-type Paramecium tetraurelia and a pawn mutant. Journal of Cell Biology, 89(2):206–??, May 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/89/2/206.

Meda:1982:DCH

[MKK^{+}82] P. Meda, E. Kohen, C. Kohen, A. Rabinovitch, and L. Orci. Direct communication of homologous and heterologous endocrine islet cells in culture. Journal of Cell Biology, 92(1): 221–??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/1/221.

Maciag:1982:OBH

[MKW^{+}82] T. Maciag, J. Kadish, L. Wilkins, M. B. Stemerman, and R. Weinstein. Organizational behavior of human umbilical vein endothelial cells. Journal of Cell Biology, 94(3):511– ??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/94/3/511.

Miller:1980:IVS

[ML80] D. K. Miller and J. Lenard. Inhibition of vesicular stomatitis virus infection by spike glycoprotein. Evidence for an intracellular, G protein-requiring step. Journal of Cell Biology, 84(2): 430–??, February 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/84/2/430.

Morris:1982:SPA

[ML82] J. R. Morris and R. J. Lasek. Stable polymers of the axonal cytoskeleton: the axoplasmic ghost. Journal of Cell Biology, 92 (1):192–??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/1/192.

Moon:1983:CIV

[ML83] R. T. Moon and E. Lazarides. Canavanine inhibits vimentin assembly but not its synthesis in chicken embryo erythroid cells. Journal of Cell Biology, 97(4):1309–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/4/1309.

Marchi:1984:RCS

[ML84a] F. Marchi and C. P. Leblond. Radioautographic characterization of successive compartments along the rough endoplasmic reticulum-Golgi pathway of collagen precursors in foot pad fibroblasts of [3H]proline-injected rats. Journal of Cell Biology, 98(5):1705–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/5/1705.

Marsh:1984:GNF

[ML84b] L. Marsh and P. C. Letourneau. Growth of neurites without filopodial or lamellipodial activity in the presence of cytochalasin B. Journal of Cell Biology, 99(6):2041–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/6/
2041.

Moon:1984:BAE

[ML84c] R. T. Moon and E. Lazarides. Biogenesis of the avian erythroid membrane skeleton: receptor-mediated assembly and stabilization of ankyrin (goblin) and spectrin. Journal of Cell Biology, 98(5):1899–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/5/1899.

Morris:1984:MPE

[ML84d] J. R. Morris and R. J. Lasek. Monomer-polymer equilibria in the axon: direct measurement of tubulin and actin as polymer and monomer in axoplasm. Journal of Cell Biology, 98 (6):2064–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/6/2064.

Meyer:1982:CMI

[MLD82] D. I. Meyer, D. Louvard, and B. Dobberstein. Characterization of molecules involved in protein translocation using a specific antibody. Journal of Cell Biology, 92(2):579–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/92/2/
579.

McKeown-Longo:1984:BDP

[MLHM84] P. J. McKeown-Longo, R. Hanning, and D. F. Mosher. Binding and degradation of platelet thrombospondin by cultured fibroblasts. Journal of Cell Biology, 98(1):22–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/98/1/22.

McKeown-Longo:1983:BPF

[MLM83] P. J. McKeown-Longo and D. F. Mosher. Binding of plasma fibronectin to cell layers of human skin fibroblasts. Journal of Cell Biology, 97(2):466–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/97/2/466.

Mecham:1981:ESL

[MLMS81] R. P. Mecham, G. Lange, J. Madaras, and B. Starcher. Elastin synthesis by ligamentum nuchae fibroblasts: effects of culture conditions and extracellular matrix on elastin production. Journal of Cell Biology, 90(2):332–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/90/2/332.

Miake-Lye:1983:MPF

[MLNK83] R. Miake-Lye, J. Newport, and M. Kirschner. Maturationpromoting factor induces nuclear envelope breakdown in cycloheximide-arrested embryos of Xenopus laevis. Journal of Cell Biology, 97(1):81–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/97/1/81.

McCloskey:1984:LED

[MLP84] M. A. McCloskey, Z. Y. Liu, and M. M. Poo. Lateral electromigration and diffusion of Fc epsilon receptors on rat basophilic leukemia cells: effects of IgE binding. Journal of Cell Biology, 99(3):778–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/3/778.

McKeithan:1983:MFT

[MLSR83] T. W. McKeithan, P. A. Lefebvre, C. D. Silflow, and J. L. Rosenbaum. Multiple forms of tubulin in polytomella and chlamydomonas: evidence for a precursor of flagellar α-tubulin. Journal of Cell Biology, 96(4):1056–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/4/
1056.

Manuelidis:1982:HRM

[MLSW82] L. Manuelidis, P. R. Langer-Safer, and D. C. Ward. Highresolution mapping of satellite DNA using biotin-labeled DNA probes. Journal of Cell Biology, 95(2):619–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/95/2/619.

Mattaj:1983:NET

[MLZD83] I. W. Mattaj, S. Lienhard, R. Zeller, and E. M. DeRobertis. Nuclear exclusion of transcription factor IIIA and the 42s particle transfer RNA-binding protein in Xenopus oocytes: a possible mechanism for gene control? Journal of Cell Biology, 97(4):1261–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/4/1261.

Milstone:1981:DPF

[MM81a] L. M. Milstone and J. McGuire. Different polypeptides form the intermediate filaments in bovine hoof and esophageal epithelium and in aortic endothelium. Journal of Cell Biology, 88 (2):312–??, February 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/88/2/312.

Morrow:1981:SAS

[MM81b] J. S. Morrow and V. T. Marchesi. Self-assembly of spectrin oligomers in vitro: a basis for a dynamic cytoskeleton. Journal of Cell Biology, 88(2):463–??, February 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/88/2/463.

Mrsny:1981:PII

[MM81c] R. J. Mrsny and S. Meizel. Potassium ion influx and Na+,K+- ATPase activity are required for the hamster sperm acrosome reaction. Journal of Cell Biology, 91(1):77–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/91/1/77.

Mullins:1982:IIC

[MM82] J. M. Mullins and J. R. McIntosh. Isolation and initial characterization of the mammalian midbody. Journal of Cell Biology, 94(3):654–??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/94/3/654.

Montesano:1983:CMP

[MMAO83] R. Montesano, P. Mouron, M. Amherdt, and L. Orci. Collagen matrix promotes reorganization of pancreatic endocrine cell monolayers into islet-like organoids. Journal of Cell Biology, 97(3):935–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/3/935.

McConlogue:1983:GRE

[MMC83] L. C. McConlogue, L. J. Marton, and P. Coffino. Growth regulatory effects of cyclic AMP and polyamine depletion are dissociable in cultured mouse lymphoma cells. Journal of Cell Biology, 96(3):762–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/3/762.

Morel:1982:IPP

[MMIGK82] N. Morel, R. Manaranche, M. Israël, and T. Gulik-Krzywicki. Isolation of a presynaptic plasma membrane fraction from Torpedo cholinergic synaptosomes: evidence for a specific protein. Journal of Cell Biology, 93(2):349–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/93/2/349.

Mills:1981:IOP

[MMJ^{+}81] J. W. Mills, A. D. C. MacKnight, J. A. Jarrell, J. M. Dayer, and D. A. Ausiello. Interaction of ouabain with the Na(+) pump in intact epithelial cells. Journal of Cell Biology, 88 (3):637–??, March 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/88/3/637.

Margolis:1981:UAA

[MMKL81] T. P. Margolis, C. M. Marchand, H. B. Kistler, and L. H. LaVail. Uptake and anterograde axonal transport of wheat germ agglutinin from retina to optic tectum in the chick. Journal of Cell Biology, 89(1):152–??, April 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/89/1/152.

McNeil:1984:MIM

[MMLT84] P. L. McNeil, R. F. Murphy, F. Lanni, and D. L. Taylor. A method for incorporating macromolecules into adherent cells. Journal of Cell Biology, 98(4):1556–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/4/1556.

Matsuura:1981:DIC

[MMO^{+}81] S. Matsuura, R. Masuda, K. Omori, M. Negishi, and Y. Tashiro. Distribution and induction of cytochrome P-450 in rat liver nuclear envelope. Journal of Cell Biology, 91(1): 212–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/91/1/212.

Montesano:1984:LII

[MMR^{+}84] R. Montesano, A. Mossaz, J. E. Ryser, L. Orci, and P. Vassalli. Leukocyte interleukins induce cultured endothelial cells to produce a highly organized, glycosaminoglycan-rich pericellular matrix. Journal of Cell Biology, 99(5):1706–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/5/
1706.

Marnell:1984:CHO

[MMS^{+}84a] M. H. Marnell, L. S. Mathis, M. Stookey, S. P. Shia, D. K. Stone, and R. K. Draper. A Chinese hamster ovary cell mutant with a heat-sensitive, conditional-lethal defect in vacuolar function. Journal of Cell Biology, 99(6):1907–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/6/
1907.

Mecham:1984:EMS

[MMS84b] R. P. Mecham, J. G. Madaras, and R. M. Senior. Extracellular matrix-specific induction of elastogenic differentiation and maintenance of phenotypic stability in bovine ligament fibroblasts. Journal of Cell Biology, 98(5):1804–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/5/
1804.

Montesano:1983:SMP

[MMVO83] R. Montesano, A. Mossaz, P. Vassalli, and L. Orci. Specialization of the macrophage plasma membrane at sites of interaction with opsonized erythrocytes. Journal of Cell Biology, 96(5):1227–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/5/1227.

Matus:1984:SAB

[MNP^{+}84] A. Matus, M. Ng, G. Pehling, M. Ackermann, and K. Hauser. Surface antigens of brain synapses: identification of minor proteins using polyclonal antisera. Journal of Cell Biology, 98(1):

237–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/1/237.

Matsuura:1982:DAR

[MNST82] S. Matsuura, H. Nakada, T. Sawamura, and Y. Tashiro. Distribution of an asialoglycoprotein receptor on rat hepatocyte cell surface. Journal of Cell Biology, 95(3):864–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/95/3/
864.

Mevel-Ninio:1981:IAT

[MNW81] M. Mével-Ninio and M. C. Weiss. Immunofluorescence analysis of the time-course of extinction, reexpression, and activation of albumin production in rat hepatoma-mouse fibroblast heterokaryons and hybrids. Journal of Cell Biology, 90 (2):339–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/2/339.

Moos:1981:FMS

[Moo81] C. Moos. Fluorescence microscope study of the binding of added C protein to skeletal muscle myofibrils. Journal of Cell Biology, 90(1):25–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/90/1/25.

Madshus:1984:MEC

[MOS84] I. H. Madshus, S. Olsnes, and K. Sandvig. Mechanism of entry into the cytosol of poliovirus type 1: requirement for low pH. Journal of Cell Biology, 98(4):1194–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/4/1194.

Montesano:1983:VRO

[MOV83] R. Montesano, L. Orci, and P. Vassalli. In vitro rapid organization of endothelial cells into capillary-like networks is promoted by collagen matrices. Journal of Cell Biology, 97(5): 1648–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/5/1648.

MacBride:1980:PLS

[MP80] R. G. MacBride and R. J. Przybylski. Purified lectin from skeletal muscle inhibits myotube formation in vitro. Journal of Cell Biology, 85(3):617–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/85/3/617.

Mueckler:1981:SFR

[MP81] M. M. Mueckler and H. C. Pitot. Structure and function of rat liver polysome populations. I. Complexity, frequency distribution, and degree of uniqueness of free and membranebound polysomal polyadenylate-containing RNA populations. Journal of Cell Biology, 90(2):495–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/90/2/495.

Mueckler:1982:SFR

[MP82] M. M. Mueckler and H. C. Pitot. Structure and function of rat liver polysome populations. II. Characterization of polyadenylate-containing mRNA associated with subpopulations of membrane-bound particles. Journal of Cell Biology, 94 (2):297–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/94/2/297.

Maupin:1983:IPS

[MP83] P. Maupin and T. D. Pollard. Improved preservation and staining of HeLa cell actin filaments, clathrin-coated membranes, and other cytoplasmic structures by tannic acidglutaraldehyde-saponin fixation. Journal of Cell Biology, 96 (1):51–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/1/51.

McNiven:1984:CMS

[MP84a] M. A. McNiven and K. R. Porter. Chromatophores–models for studying cytomatrix translocations. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/1/152s.

Mellman:1984:IDM

[MP84b] I. Mellman and H. Plutner. Internalization and degradation of macrophage Fc receptors bound to polyvalent immune complexes. Journal of Cell Biology, 98(4):1170–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/4/
1170.

Myles:1984:LSA

[MP84c] D. G. Myles and P. Primakoff. Localized surface antigens of guinea pig sperm migrate to new regions prior to fertilization. Journal of Cell Biology, 99(5):1634–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/5/
1634.

Matus:1980:BPD

[MPAM80] A. Matus, G. Pehling, M. Ackermann, and J. Maeder. Brain postsynaptic densities: the relationship to glial and neuronal filaments. Journal of Cell Biology, 87(2):346–??, November 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/87/2/
346.

Murphy:1984:EPM

[MPC84] R. F. Murphy, S. Powers, and C. R. Cantor. Endosome pH measured in single cells by dual fluorescence flow cytometry: rapid acidification of insulin to pH 6. Journal of Cell Biology, 98(5):1757–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/5/1757.

McCarthy:1983:MHS

[MPF83a] J. B. McCarthy, S. L. Palm, and L. T. Furcht. Migration by haptotaxis of a Schwann cell tumor line to the basement membrane glycoprotein laminin. Journal of Cell Biology, 97(3): 772–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/3/772.

Mitchell:1983:PMI

[MPF83b] R. D. Mitchell, P. Palade, and S. Fleischer. Purification of morphologically intact triad structures from skeletal muscle.

Journal of Cell Biology, 96(4):1008–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/4/1008.

Manfredi:1982:TBC

[MPH82] J. J. Manfredi, J. Parness, and S. B. Horwitz. Taxol binds to cellular microtubules. Journal of Cell Biology, 94(3):688– ??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/94/3/688.

Maylie-Pfenninger:1980:DCS

[MPJ80] M. F. Maylie-Pfenninger and J. D. Jamieson. Development of cell surface saccharides on embryonic pancreatic cells. Journal of Cell Biology, 86(1):96–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/86/1/96.

Myles:1984:LSP

[MPK84] D. G. Myles, P. Primakoff, and D. E. Koppel. A localized surface protein of guinea pig sperm exhibits free diffusion in its domain. Journal of Cell Biology, 98(5):1905–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/5/
1905.

Martinez-Palomo:1980:EMO

[MPMBC80] A. Martinez-Palomo, I. Meza, G. Beaty, and M. Cereijido. Experimental modulation of occluding junctions in a cultured transporting epithelium. Journal of Cell Biology, 87(3):736– ??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/87/3/736.

Mellman:1983:IDM

[MPS^{+}83] I. S. Mellman, H. Plutner, R. M. Steinman, J. C. Unkeless, and Z. A. Cohn. Internalization and degradation of macrophage Fc receptors during receptor-mediated phagocytosis. Journal of Cell Biology, 96(3):887–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/3/887.

Mellman:1984:IRR

[MPU84] I. Mellman, H. Plutner, and P. Ukkonen. Internalization and rapid recycling of macrophage Fc receptors tagged with monovalent antireceptor antibody: possible role of a prelysosomal compartment. Journal of Cell Biology, 98(4):1163–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/4/
1163.

Mooseker:1982:NPA

[MPW82] M. S. Mooseker, T. D. Pollard, and K. A. Wharton. Nucleated polymerization of actin from the membrane-associated ends of microvillar filaments in the intestinal brush border. Journal of Cell Biology, 95(1):223–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/95/1/223.

Marin:1980:RDD

[MR80a] F. T. Marin and F. G. Rothman. Regulation of development in Dictyostelium discoideum. IV. Effects of ions on the rate of differentiation and cellular response to cyclic AMP. Journal of Cell Biology, 87(3):823–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/87/3/823.

McPherson:1980:IDS

[MR80b] M. A. McPherson and J. Ramachandran. Inhibition of DNA synthesis in adrenocortical cells by cytochalasin B. Journal of Cell Biology, 86(1):129–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/86/1/129.

McKeithan:1981:MFT

[MR81a] T. W. McKeithan and J. L. Rosenbaum. Multiple forms of tubulin in the cytoskeletal and flagellar microtubules of polytomella. Journal of Cell Biology, 91(2):352–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/91/2/
352.

Mechler:1981:MBRa

[MR81b] B. Mechler and T. H. Rabbitts. Membrane-bound ribosomes of myeloma cells. IV. mRNA complexity of free and membranebound polysomes. Journal of Cell Biology, 88(1):29–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
88/1/29.

Mottola:1982:CMM

[MR82] C. Mottola and D. Romeo. Calcium movement and membrane potential changes in the early phase of neutrophil activation by phorbol myristate acetate: a study with ion-selective electrodes. Journal of Cell Biology, 93(1):129–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/93/1/129.

Mumby:1984:ITE

[MRB84] S. M. Mumby, G. J. Raugi, and P. Bornstein. Interactions of thrombospondin with extracellular matrix proteins: selective binding to type V collagen. Journal of Cell Biology, 98(2): 646–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/2/646.

Madri:1980:ULF

[MRFF80] J. A. Madri, F. J. Roll, H. Furthmayr, and J. M. Foidart. Ultrastructural localization of fibronectin and laminin in the basement membranes of the murine kidney. Journal of Cell Biology, 86(2):682–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/86/2/682.

Matlin:1981:IEP

[MRHS81] K. S. Matlin, H. Reggio, A. Helenius, and K. Simons. Infectious entry pathway of influenza virus in a canine kidney cell line. Journal of Cell Biology, 91(3):601–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/91/3/601.

Mishkind:1982:ILW

[MRPK82] M. Mishkind, N. V. Raikhel, B. A. Palevitz, and K. Keegstra. Immunocytochemical localization of wheat germ agglutinin in wheat. Journal of Cell Biology, 92(3):753–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/92/3/753.

McDonnel:1980:ABL

[MS80a] A. McDonnel and L. A. Staehelin. Adhesion between liposomes mediated by the chlorophyll a/b light-harvesting complex isolated from chloroplast membranes. Journal of Cell Biology, 84 (1):40–??, January 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/84/1/40.

Mooseker:1980:BBA

[MS80b] M. S. Mooseker and R. E. Stephens. Brush-border alphaactinin? Comparison of two proteins of the microvillus core with alpha-actinin by two-dimensional peptide mapping. Journal of Cell Biology, 86(2):466–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/86/2/466.

Merion:1983:RIV

[MS83a] M. Merion and W. S. Sly. The role of intermediate vesicles in the adsorptive endocytosis and transport of ligand to lysosomes by human fibroblasts. Journal of Cell Biology, 96 (3):644–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/3/644.

Morschel:1983:RCF

[MS83b] E. Mörschel and L. A. Staehelin. Reconstitution of cytochrome f/b6 and CF0-CF1 ATP synthetase complexes into phospholipid and galactolipid liposomes. Journal of Cell Biology, 97 (2):301–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/2/301.

Marton:1984:DAM

[MS84a] L. S. Marton and K. Stefansson. Developmental alterations in molecular weights of proteins in the human central nervous system that react with antibodies against myelin-associated glycoprotein. Journal of Cell Biology, 99(5):1642–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/5/

Matlin:1984:SAP

[MS84b] K. S. Matlin and K. Simons. Sorting of an apical plasma membrane glycoprotein occurs before it reaches the cell surface in cultured epithelial cells. Journal of Cell Biology, 99(6): 2131–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/6/2131.

McMahan:1984:IBL

[MS84c] U. J. McMahan and C. R. Slater. The influence of basal lamina on the accumulation of acetylcholine receptors at synaptic sites in regenerating muscle. Journal of Cell Biology, 98(4):1453–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
98/4/1453.

Miller:1984:KIM

[MS84d] M. Miller and F. Solomon. Kinetics and intermediates of marginal band reformation: evidence for peripheral determinants of microtubule organization. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/1/70s.

Mohraz:1984:SKA

[MS84e] M. Mohraz and P. R. Smith. Structure of (Na+,K+)-ATPase as revealed by electron microscopy and image processing. Journal of Cell Biology, 98(5):1836–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/5/1836.

McGraw:1980:LCP

[MSB80] C. F. McGraw, A. V. Somlyo, and M. P. Blaustein. Localization of calcium in presynaptic nerve terminals. An ultrastructural and electron microprobe analysis. Journal of Cell Biology, 85(2):228–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/85/2/228.

Meier:1983:ABA

[MSB^{+}83] K. E. Meier, M. D. Snavely, S. L. Brown, J. H. Brown, and P. A. Insel. alpha 1- and beta 2-adrenergic receptor expression in the Madin–Darby canine kidney epithelial cell line. Journal of Cell Biology, 97(2):405–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/97/2/405.

Muller:1980:MPVa

[MSC80a] W. A. Muller, R. M. Steinman, and Z. A. Cohn. The membrane proteins of the vacuolar system I. Analysis of a novel method of intralysosomal iodination. Journal of Cell Biology, 86(1):292–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/86/1/292.

Muller:1980:MPVb

[MSC80b] W. A. Muller, R. M. Steinman, and Z. A. Cohn. The membrane proteins of the vacuolar system. II. Bidirectional flow between secondary lysosomes and plasma membrane. Journal of Cell Biology, 86(1):304–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/86/1/304.

Muller:1983:MPV

[MSC83] W. A. Muller, R. M. Steinman, and Z. A. Cohn. Membrane proteins of the vacuolar system. III. Further studies on the composition and recycling of endocytic vacuole membrane in cultured macrophages. Journal of Cell Biology, 96 (1):29–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/1/29.

Marnell:1982:MBT

[MSD82a] M. H. Marnell, M. Stookey, and R. K. Draper. Monensin blocks the transport of diphtheria toxin to the cell cytoplasm. Journal of Cell Biology, 93(1):57–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/93/1/57.

Murtaugh:1982:LCS

[MSD82b] M. P. Murtaugh, A. L. Steiner, and P. J. Davies. Localization of the catalytic subunit of cyclic AMP-dependent. Protein kinase in cultured cells using a specific antibody. Journal of Cell Biology, 95(1):64–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/95/1/64.

Mayrand:1981:SNR

[MSGP81] S. Mayrand, B. Setyono, J. R. Greenberg, and T. Pederson. Structure of nuclear ribonucleoprotein: identification of proteins in contact with poly(A)+ heterogeneous nuclear RNA in living HeLa cells. Journal of Cell Biology, 90(2):380–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
90/2/380.

Maruyama:1984:CFS

[MSK^{+}84] K. Maruyama, H. Sawada, S. Kimura, K. Ohashi, H. Higuchi, and Y. Umazume. Connectin filaments in stretched skinned fibers of frog skeletal muscle. Journal of Cell Biology, 99(4): 1391–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/4/1391.

Means:1984:PST

[MSP84] A. R. Means, G. R. Slaughter, and J. A. Putkey. Postreceptor signal transduction by cyclic adenosine monophosphate and the Ca^{2+}-calmodulin complex. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/1/226s.

Mitchell:1983:MIT

[MSPF83] R. D. Mitchell, A. Saito, P. Palade, and S. Fleischer. Morphology of isolated triads. Journal of Cell Biology, 96(4):1017–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
96/4/1017.

Meier:1984:SFP

[MSRB84] P. J. Meier, E. S. Sztul, A. Reuben, and J. L. Boyer. Structural and functional polarity of canalicular and basolateral plasma membrane vesicles isolated in high yield from rat liver. Journal of Cell Biology, 98(3):991–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/3/991.

Mellman:1980:SIP

[MSUC80] I. S. Mellman, R. M. Steinman, J. C. Unkeless, and Z. A. Cohn. Selective iodination and polypeptide composition of pinocytic vesicles. Journal of Cell Biology, 86(3):712–??, September 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/86/3/
712.

Malone:1982:ROP

[MTG^{+}82] J. D. Malone, S. L. Teitelbaum, G. L. Griffin, R. M. Senior, and A. J. Kahn. Recruitment of osteoclast precursors by purified bone matrix constituents. Journal of Cell Biology, 92 (1):227–??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/1/227.

Mazariegos:1984:ATJ

[MTH84] M. R. Mazariegos, L. W. Tice, and A. R. Hand. Alteration of tight junctional permeability in the rat parotid gland after isoproterenol stimulation. Journal of Cell Biology, 98(5):1865–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
98/5/1865.

Manaranche:1980:EVG

[MTI80] R. Manaranche, M. Thieffry, and M. Israel. Effect of the venom of Glycera convoluta on the spontaneous quantal release of transmitter. Journal of Cell Biology, 85(2):446–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/85/2/446.

Marshall:1984:EES

[MTKH84] L. M. Marshall, A. Thureson-Klein, and R. C. Hunt. Exclusion of erythrocyte-specific membrane proteins from clathrincoated pits during differentiation of human erythroleukemic cells. Journal of Cell Biology, 98(6):2055–??, June 1984. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/98/6/
2055.

Morel:1983:BGC

[MTM83] N. Morel, M. Thieffry, and R. Manaranche. Binding of a Glycera convoluta neurotoxin to cholinergic nerve terminal plasma membranes. Journal of Cell Biology, 97(6):1737–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/6/
1737.

McNeil:1983:API

[MTMT83] P. L. McNeil, L. Tanasugarn, J. B. Meigs, and D. L. Taylor. Acidification of phagosomes is initiated before lysosomal enzyme activity is detected. Journal of Cell Biology, 97(3): 692–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/3/692.

Matthew:1981:ISV

[MTR81] W. D. Matthew, L. Tsavaler, and L. F. Reichardt. Identification of a synaptic vesicle-specific membrane protein with a wide distribution in neuronal and neurosecretory tissue. Journal of Cell Biology, 91(1):257–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/91/1/257.

Mikuni-Takagaki:1980:CSA

[MTT80] Y. Mikuni-Takagaki and B. P. Toole. Cell-substratum attachment and cell surface hyaluronate of Rous sarcoma virustransformed chondrocytes. Journal of Cell Biology, 85(2): 481–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/85/2/481.

Mikkelsen:1982:MPP

[MTW82] R. B. Mikkelsen, K. Tanabe, and D. F. Wallach. Membrane potential of Plasmodium-infected erythrocytes. Journal of Cell Biology, 93(3):685–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/93/3/685.

Milligan:1982:VCR

[MU82] R. A. Milligan and P. N. Unwin. In vitro crystallization of ribosomes from chick embryos. Journal of Cell Biology, 95(2): 648–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/95/2/648.

Murphy:1980:IMA

[Mur80] D. B. Murphy. Identification of microtubule-associated proteins in the meiotic spindle of surf clam oocytes. Journal of Cell Biology, 84(2):235–??, February 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/84/2/235.

Murray:1984:DRM

[Mur84a] J. M. Murray. Disassembly and reconstitution of a membranemicrotubule complex. Journal of Cell Biology, 98(4):1481–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
98/4/1481.

Murray:1984:TDS

[Mur84b] J. M. Murray. Three-dimensional structure of a membranemicrotubule complex. Journal of Cell Biology, 98(1):283– ??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/1/283.

Mitchell:1980:IDA

[MW80] D. R. Mitchell and F. D. Warner. Interactions of dynein arms with b subfibers of Tetrahymena cilia: quantitation of the effects of magnesium and adenosine triphosphate. Journal of Cell Biology, 87(1):84–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/87/1/84.

Madri:1983:CEC

[MW83a] J. A. Madri and S. K. Williams. Capillary endothelial cell cultures: phenotypic modulation by matrix components. Journal of Cell Biology, 97(1):153–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/97/1/153.

Malinoff:1983:ICS

[MW83b] H. L. Malinoff and M. S. Wicha. Isolation of a cell surface receptor protein for laminin from murine fibrosarcoma cells. Journal of Cell Biology, 96(5):1475–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/5/1475.

Mirsky:1980:MSP

[MWA^{+}80] R. Mirsky, J. Winter, E. R. Abney, R. M. Pruss, J. Gavrilovic, and M. C. Raff. Myelin-specific proteins and glycolipids in rat Schwann cells and oligodendrocytes in culture. Journal of Cell Biology, 84(3):483–??, March 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/84/3/483.

Murphy:1983:IOAb

[MWH83] D. B. Murphy, K. T. Wallis, and R. R. Hiebsch. Identity and origin of the ATPase activity associated with neuronal microtubules. II. Identification of a 50,000-dalton polypeptide with ATPase activity similar to F-1 ATPase from mitochondria. Journal of Cell Biology, 96(5):1306–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/5/1306.

Mayne:1984:MAA

[MWI^{+}84] R. Mayne, H. Wiedemann, M. H. Irwin, R. D. Sanderson, J. M. Fitch, T. F. Linsenmayer, and K. Kühn. Monoclonal antibodies against chicken type IV and V collagens: electron microscopic mapping of the epitopes after rotary shadowing. Journal of Cell Biology, 98(5):1637–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/5/1637.

Madore:1984:PUS

[MWKP84] S. J. Madore, E. D. Wieben, G. R. Kunkel, and T. Pederson. Precursors of U4 small nuclear RNA. Journal of Cell Biology, 99(3):1140–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/3/1140.

Moon:1981:PPN

[MWM^{+}81] H. M. Moon, T. Wisniewski, P. Merz, J. De Martini, and H. M. Wisniewski. Partial purification of neurofilament subunits from bovine brains and studies on neurofilament assembly. Journal of Cell Biology, 89(3):560–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/89/3/560.

Madore:1984:ISU

[MWP84] S. J. Madore, E. D. Wieben, and T. Pederson. Intracellular site of U1 small nuclear RNA processing and ribonucleoprotein assembly. Journal of Cell Biology, 98(1):188–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/1/
188.

Meyer:1981:ICN

[MYR81] D. J. Meyer, S. B. Yancey, and J. P. Revel. Intercellular communication in normal and regenerating rat liver: a quantitative analysis. Journal of Cell Biology, 91(2):505–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/91/2/
505.

Nachmias:1980:CHP

[Nac80] V. T. Nachmias. Cytoskeleton of human platelets at rest and after spreading. Journal of Cell Biology, 86(3):795–??, September 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
86/3/795.

Nachmias:1981:HPM

[Nac81] V. T. Nachmias. Hybrids of Physarum myosin light chains and desensitized scallop myofibrils. Journal of Cell Biology, 90 (2):408–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/2/408.

Niederman:1983:TDS

[NAH83] R. Niederman, P. C. Amrein, and J. Hartwig. Threedimensional structure of actin filaments and of an actin gel made with actin-binding protein. Journal of Cell Biology, 96 (5):1400–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/5/1400.

Nelson:1982:PCI

[NAK82] G. A. Nelson, M. L. Andrews, and M. J. Karnovsky. Participation of calmodulin in immunoglobulin capping. Journal of Cell Biology, 95(3):771–??, December 1982. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/95/3/771.

Nelson:1983:CES

[NAK83] G. A. Nelson, M. L. Andrews, and M. J. Karnovsky. Control of erythrocyte shape by calmodulin. Journal of Cell Biology, 96 (3):730–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/3/730.

Nakajima:1981:AFS

[NB81] Y. Nakajima and P. C. Bridgman. Absence of filipin-sterol complexes from the membranes of active zones and acetylcholine receptor aggregates at frog neuromuscular junctions. Journal of Cell Biology, 88(2):453–??, February 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/88/2/453.

Nierzwicki-Bauer:1983:TDU

[NBBS83] S. A. Nierzwicki-Bauer, D. L. Balkwill, and S. E. Stevens. Three-dimensional ultrastructure of a unicellular cyanobacterium. Journal of Cell Biology, 97(3):713–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/97/3/713.

Nixon:1982:PMN

[NBM82] R. A. Nixon, B. A. Brown, and C. A. Marotta. Posttranslational modification of a neurofilament protein during axoplasmic transport: implications for regional specialization of CNS axons. Journal of Cell Biology, 94(1):150–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/94/1/150.

Neff:1981:DPM

[NBMD81] N. T. Neff, L. Bourret, P. Miao, and J. F. Dice. Degradation of proteins microinjected into IMR-90 human diploid fibroblasts. Journal of Cell Biology, 91(1):184–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/91/1/184.

Nothnagel:1981:FSM

[NBSW81] E. A. Nothnagel, L. S. Barak, J. W. Sanger, and W. W. Webb. Fluorescence studies on modes of cytochalasin B and phallotoxin action on cytoplasmic streaming in Chara. Journal of Cell Biology, 88(2):364–??, February 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/88/2/364.

Nicholson:1984:STS

[NCBP84] N. B. Nicholson, S. Chen, G. Blanck, and R. Pollack. SV40 transformation of Swiss 3T3 cells can cause a stable reduction in the calcium requirement for growth. Journal of Cell Biology, 99(6):2314–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/6/2314.

Nigg:1983:ILD

[NCH83] E. A. Nigg, J. A. Cooper, and T. Hunter. Immunofluorescent localization of a 39,000-dalton substrate of tyrosine protein kinases to the cytoplasmic surface of the plasma membrane. Journal of Cell Biology, 96(6):1601–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/6/1601.

Ngai:1984:DME

[NCL84] J. Ngai, Y. G. Capetanaki, and E. Lazarides. Differentiation of murine erythroleukemia cells results in the rapid repression of vimentin gene expression. Journal of Cell Biology, 99 (1):306–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/1/306.

Neubauer:1983:EMI

[NCM^{+}83] B. L. Neubauer, L. W. Chung, K. A. McCormick, O. Taguchi, T. C. Thompson, and G. R. Cunha. Epithelial-mesenchymal interactions in prostatic development. II. Biochemical observations of prostatic induction by urogenital sinus mesenchyme in epithelium of the adult rodent urinary bladder. Journal of Cell Biology, 96(6):1671–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/6/1671.

Nir:1984:ILO

[NCP84] I. Nir, D. Cohen, and D. S. Papermaster. Immunocytochemical localization of opsin in the cell membrane of developing rat retinal photoreceptors. Journal of Cell Biology, 98(5):1788–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
98/5/1788.

Nunnally:1980:FCC

[NDC80] M. H. Nunnally, J. M. D’Angelo, and S. W. Craig. Filamin concentration in cleavage furrow and midbody region: frequency of occurrence compared with that of alpha-actinin and myosin. Journal of Cell Biology, 87(1):219–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/87/1/219.

Norenberg:1980:GSL

[NDRL80] M. D. Norenberg, K. Dutt, and L. Reif-Lehrer. Glutamine synthetase localization in cortisol-induced chick embryo retinas. Journal of Cell Biology, 84(3):803–??, March 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/84/3/803.

Noe:1984:CPP

[NDS84] B. D. Noe, G. Debo, and J. Spiess. Comparison of prohormoneprocessing activities in islet microsomes and secretory granules: evidence for distinct converting enzymes for separate islet prosomatostatins. Journal of Cell Biology, 99(2):578–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/2/578.

Nath:1982:TTH

[NFG82] J. Nath, M. Flavin, and J. I. Gallin. Tubulin tyrosinolation in human polymorphonuclear leukocytes: studies in normal subjects and in patients with the Chediak–Higashi syndrome. Journal of Cell Biology, 95(2):519–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/95/2/519.

Nath:1981:STT

[NFS81] J. Nath, M. Flavin, and E. Schiffmann. Stimulation of tubulin tyrosinolation in rabbit leukocytes evoked by the chemoattractant formyl-methionyl-leucyl-phenylalanine. Journal of Cell Biology, 91(1):232–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/91/1/232.

Nelson:1983:ALS

[NGL83] W. J. Nelson, B. L. Granger, and E. Lazarides. Avian lens spectrin: subunit composition compared with erythrocyte and brain spectrin. Journal of Cell Biology, 97(4):1271–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/4/
1271.

Natowicz:1983:RRM

[NHF^{+}83] M. Natowicz, D. W. Hallett, C. Frier, M. Chi, P. H. Schlesinger, and J. U. Baenziger. Recognition and receptormediated uptake of phosphorylated high mannose-type oligosaccharides by cultured human fibroblasts. Journal of Cell Biology, 96(3):915–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/3/915.

Nelson:1984:LRG

[NHMT84] T. Nelson, M. H. Harpster, S. P. Mayfield, and W. C. Taylor. Light-regulated gene expression during maize leaf development. Journal of Cell Biology, 98(2):558–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/98/2/558.

Nicklas:1983:MFP

[Nic83] R. B. Nicklas. Measurements of the force produced by the mitotic spindle in anaphase. Journal of Cell Biology, 97(2): 542–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/2/542.

Nichols:1980:EMC

[NJR80] K. M. Nichols, A. Jacklet, and R. Rikmenspoel. Effects of Mg2+ and Ca^{2+} on photoinduced Euglena flagellar responses. Journal of Cell Biology, 84(2):355–??, February 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/84/2/355.

Nemhauser:1983:CMO

[NJSC83] I. Nemhauser, J. Joseph-Silverstein, and W. D. Cohen. Centriole as microtubule-organizing centers for marginal bands of molluscan erythrocytes. Journal of Cell Biology, 96(4):979–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
96/4/979.

Newman:1981:DCM

[NKH81] P. J. Newman, R. A. Kahn, and A. Hines. Detection and characterization of monoclonal antibodies to platelet membrane proteins. Journal of Cell Biology, 90(1):249–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/90/1/249.

Nicklas:1982:SMT

[NKH82] R. B. Nicklas, D. F. Kubai, and T. S. Hays. Spindle microtubules and their mechanical associations after micromanipulation in anaphase. Journal of Cell Biology, 95(1):91–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
95/1/91.

Nandini-Kishore:1980:EPS

[NKT80] S. G. Nandini-Kishore and G. A. Thompson. Evidence for a pronounced secretion of cyclic AMP by Tetrahymena. Journal of Cell Biology, 85(3):910–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/85/3/910.

Neff:1982:MAD

[NLD^{+}82] N. T. Neff, C. Lowrey, C. Decker, A. Tovar, C. Damsky, C. Buck, and A. F. Horwitz. A monoclonal antibody detaches embryonic skeletal muscle from extracellular matrices. Journal of Cell Biology, 95(2):654–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/95/2/654.

Nagy:1981:IKR

[NLN^{+}81] I. Z. Nagy, G. Lustyik, V. Z. Nagy, B. Zarándi, and C. Bertoni-Freddari. Intracellular Na+: K+ ratios in human cancer cells as revealed by energy dispersive X-ray microanalysis. Journal of Cell Biology, 90(3):769–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/90/3/769.

Novikoff:1983:ILL

[NLN^{+}83] P. M. Novikoff, N. F. La Russo, A. B. Novikoff, R. J. Stockert, A. Yam, and G. D. Le Sage. Immunocytochemical localization of lysosomal beta-galactosidase in rat liver. Journal of Cell Biology, 97(5):1559–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/97/5/1559.

Nagy:1982:NLA

[NLPH82] M. Nagy, J. Laporte, B. Penverne, and G. Hervé. Nuclear localization of aspartate transcabamoylase in Saccharomyces cerevisiae. Journal of Cell Biology, 92(3):790–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/92/3/790.

Noe:1984:ANS

[NM84] B. D. Noe and M. N. Moran. Association of newly synthesized islet prohormones with intracellular membranes. Journal of Cell Biology, 99(2):418–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/2/418.

Nishida:1983:FPE

[NNA^{+}83] T. Nishida, S. Nakagawa, T. Awata, Y. Ohashi, K. Watanabe, and R. Manabe. Fibronectin promotes epithelial migration of cultured rabbit cornea in situ. Journal of Cell Biology, 97(5): 1653–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/5/1653.

Novikoff:1980:ORC

[NNRR80] A. B. Novikoff, P. M. Novikoff, O. M. Rosen, and C. S. Rubin. Organelle relationships in cultured 3T3-L1 preadipocytes. Journal of Cell Biology, 87(1):180–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/87/1/180.

Newburger:1980:DOP

[NPG^{+}80] P. E. Newburger, J. S. Pagano, J. S. Greenberger, A. Karpas, and H. J. Cohen. Dissociation of opsonized particle phagocytosis and respiratory burst activity in an Epstein–Barr virusinfected myeloid cell line. Journal of Cell Biology, 85(3):

549–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/85/3/549.

Nagata:1984:BAM

[NRMB84] Y. Nagata, O. M. Rosen, M. H. Makman, and B. R. Bloom. Biochemical analysis of mutants of a macrophage cell line resistant to the growth-inhibitory activity of interferon. Journal of Cell Biology, 98(4):1342–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/4/1342.

Nelson:1982:CES

[NRW82] G. A. Nelson, T. M. Roberts, and S. Ward. Caenorhabditis elegans spermatozoan locomotion: amoeboid movement with almost no actin. Journal of Cell Biology, 92(1):121–??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/92/1/
121.

Nelson:1983:KKC

[NS83a] W. G. Nelson and T. T. Sun. The 50- and 58-kdalton keratin classes as molecular markers for stratified squamous epithelia: cell culture studies. Journal of Cell Biology, 97(1):244–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/1/
244.

Novick:1983:EMC

[NS83b] P. Novick and R. Schekman. Export of major cell surface proteins is blocked in yeast secretory mutants. Journal of Cell Biology, 96(2):541–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/2/541.

Nagata:1980:CCP

[NSI80] K. Nagata, J. Sagara, and Y. Ichikawa. Changes in contractile proteins during differentiation of myeloid leukemia cells. I. Polymerization of actin. Journal of Cell Biology, 85(2):273–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
85/2/273.

Nagata:1982:CCP

[NSI82] K. Nagata, J. Sagara, and Y. Ichikawa. Changes in contractile proteins during differentiation of myeloid leukemia cells. II. Purification and characterization of actin. Journal of Cell Biology, 93(2):470–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/93/2/470.

Nothnagel:1982:EEP

[NSW82] E. A. Nothnagel, J. W. Sanger, and W. W. Webb. Effects of exogenous proteins on cytoplasmic streaming in perfused Chara cells. Journal of Cell Biology, 93(3):735–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/93/3/735.

Nijweide:1982:BFC

[NvIvGdH^{+}82] P. J. Nijweide, A. S. van Iperen-van Gent, E. W. Kawilarang de Haas, A. van der Plas, and A. M. Wassenaar. Bone formation and calcification by isolated osteoblastlike cells. Journal of Cell Biology, 93(2):318–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/93/2/318.

Nitsch:1980:UIS

[NW80] L. Nitsch and S. H. Wollman. Ultrastructure of intermediate stages in polarity reversal of thyroid epithelium in follicles in suspension culture. Journal of Cell Biology, 86(3): 875–??, September 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/86/3/875.

Nothnagel:1982:HMV

[NW82] E. A. Nothnagel and W. W. Webb. Hydrodynamic models of viscous coupling between motile myosin and endoplasm in characean algae. Journal of Cell Biology, 94(2):444–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/94/2/
444.

Nijhof:1983:ICE

[NW83] W. Nijhof and P. K. Wierenga. Isolation and characterization of the erythroid progenitor cell: CFU–E. Journal of Cell Biology, 96(2):386–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/2/386.

Okuno:1981:EAA

[OAOB81] M. Okuno, D. J. Asai, K. Ogawa, and C. J. Brokaw. Effects of antibodies against dynein and tubulin on the stiffness of flagellar axonemes. Journal of Cell Biology, 91(3):689–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/91/3/
689.

OBrien:1980:TAS

[O’B80] J. C. O’Brien. Tyrosine aminotransferase sensitivity to bromodeoxyuridine during restricted intervals of S phase in hepatoma cells. Journal of Cell Biology, 87(3):629–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/87/3/
629.

OConnor:1981:ADC

[OB81] P. O’Connor and B. Burnside. Actin-dependent cell elongation in teleost retinal rods: requirement for actin filament assembly. Journal of Cell Biology, 89(3):517–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/89/3/517.

OConnor:1982:ECA

[OB82] P. O’Connor and B. Burnside. Elevation of cyclic AMP activates an actin-dependent contraction in teleost retinal rods. Journal of Cell Biology, 95(2):445–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/95/2/445.

Opperdoes:1984:PMA

[OBC^{+}84] F. R. Opperdoes, P. Baudhuin, I. Coppens, C. De Roe, S. W. Edwards, P. J. Weijers, and O. Misset. Purification, morphometric analysis, and characterization of the glycosomes (microbodies) of the protozoan hemoflagellate Trypanosoma brucei. Journal of Cell Biology, 98(4):1178–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/4/1178.

Osawa:1984:RAR

[OBH84] S. Osawa, G. Betz, and P. F. Hall. Role of actin in the responses of adrenal cells to ACTH and cyclic AMP: inhibition by DNase I. Journal of Cell Biology, 99(4):1335–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/4/
1335.

Olmsted:1984:CRM

[OCA^{+}84] J. B. Olmsted, J. V. Cox, C. F. Asnes, L. M. Parysek, and H. D. Lyon. Cellular regulation of microtubule organization. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/1/28s.

Okagaki:1981:MNC

[OCT81] T. Okagaki, B. A. Clark, and L. B. Twiggs. Measurement of number and cross-sectional area of basal cell pseudopodia: a new morphometric method. Journal of Cell Biology, 91(3): 629–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/91/3/629.

Ohmori:1982:EMH

[OFH^{+}82] K. Ohmori, L. J. Fretto, R. L. Harrison, M. E. Switzer, H. P. Erickson, and P. A. McKee. Electron microscopy of human factor VIII/Von Willebrand glycoprotein: effect of reducing reagents on structure and function. Journal of Cell Biology, 95 (2):632–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/95/2/632.

Ocklind:1983:CSL

[OFO83] C. Ocklind, U. Forsum, and B. Obrink. Cell surface localization and tissue distribution of a hepatocyte cell–cell adhesion glycoprotein (cell–CAM 105). Journal of Cell Biology, 96 (4):1168–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/4/1168.

Oi:1982:FPC

[OGS82] V. T. Oi, A. N. Glazer, and L. Stryer. Fluorescent phycobiliprotein conjugates for analyses of cells and molecules.

Journal of Cell Biology, 93(3):981–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/93/3/981.

Ozaki:1984:MRA

[OH84] K. Ozaki and S. Hatano. Mechanism of regulation of actin polymerization by Physarum profilin. Journal of Cell Biology, 98(6):1919–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/6/1919.

Orci:1984:NAA

[OHA^{+}84] L. Orci, P. Halban, M. Amherdt, M. Ravazzola, J. D. Vassalli, and A. Perrelet. Nonconverted, amino acid analog-modified proinsulin stays in a Golgi-derived clathrin-coated membrane compartment. Journal of Cell Biology, 99(6):2187–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/6/2187.

OMorchoe:1984:TDP

[OJJ^{+}84] C. C. O’Morchoe, W. R. Jones, H. M. Jarosz, P. J. O’Morchoe, and L. M. Fox. Temperature dependence of protein transport across lymphatic endothelium in vitro. Journal of Cell Biology, 98(2):629–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/2/629.

Omoto:1980:RTC

[OK80] C. K. Omoto and C. Kung. Rotation and twist of the centralpair microtubules in the cilia of Paramecium. Journal of Cell Biology, 87(1):33–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/87/1/33.

Olson:1982:ISA

[OK82] J. A. Olson and A. Kilejian. Involvement of spectrin and ATP in infection of resealed erythrocyte ghosts by the human malarial parasite, Plasmodium falciparum. Journal of Cell Biology, 95(3):757–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/95/3/757.

Ohad:1984:MPD

[OKA84] I. Ohad, D. J. Kyle, and C. J. Arntzen. Membrane protein damage and repair: removal and replacement of inactivated 32-kilodalton polypeptides in chloroplast membranes. Journal of Cell Biology, 99(2):481–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/2/481.

Owaribe:1981:DCC

[OKE81] K. Owaribe, R. Kodama, and G. Eguchi. Demonstration of contractility of circumferential actin bundles and its morphogenetic significance in pigmented epithelium in vitro and in vivo. Journal of Cell Biology, 90(2):507–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/90/2/507.

Okuno:1980:IRS

[Oku80] M. Okuno. Inhibition and relaxation of sea urchin sperm flagella by vanadate. Journal of Cell Biology, 85(3):712–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/85/3/
712.

Ostlund:1980:RMA

[OLH80] R. E. Ostlund, J. T. Leung, and S. V. Hajek. Regulation of microtubule assembly in cultured fibroblasts. Journal of Cell Biology, 85(2):386–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/85/2/386.

Olden:1981:IFE

[OLH^{+}81] K. Olden, J. Law, V. A. Hunter, R. Romain, and J. B. Parent. Inhibition of fusion of embryonic muscle cells in culture by tunicamycin is prevented by leupeptin. Journal of Cell Biology, 88(1):199–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/88/1/199.

OLoughlin:1981:ECE

[OLHH81] J. O’Loughlin, L. Lehr, A. Havaranis, and S. M. Heywood. Encapsulation of “core” eIF3, regulatory components of eIF3 and mRNA into liposomes, and their subsequent uptake into myogenic cells in culture. Journal of Cell Biology, 90(1):160–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
90/1/160.

Oliver:1982:EPL

[Oli82] C. Oliver. Endocytic pathways at the lateral and basal cell surfaces of exocrine acinar cells. Journal of Cell Biology, 95 (1):154–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/95/1/154.

Olmsted:1981:TPD

[Olm81] J. B. Olmsted. Tubulin pools in differentiating neuroblastoma cells. Journal of Cell Biology, 89(3):418–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/89/3/418.

Owaribe:1982:ICC

[OM82] K. Owaribe and H. Masuda. Isolation and characterization of circumferential microfilament bundles from retinal pigmented epithelial cells. Journal of Cell Biology, 95(1):310–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/95/1/
310.

Oakley:1983:MAN

[OM83] B. R. Oakley and R. Morris. A mutation in aspergillus nidulans that blocks the transition from interphase to prophase. Journal of Cell Biology, 96(4):1155–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/4/1155.

Oppenheimer-Marks:1982:IFR

[OMG82] N. Oppenheimer-Marks and F. Grinnell. Inhibition of fibronectin receptor function by antibodies against baby hamster kidney cell wheat germ agglutinin receptors. Journal of Cell Biology, 95(3):876–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/95/3/876.

Ohashi:1982:LZP

[OMM82] K. Ohashi, T. Mikawa, and K. Maruyama. Localization of Zprotein in isolated Z-disk sheets of chicken leg muscle. Journal of Cell Biology, 95(1):85–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/95/1/85.

Ogawa:1982:IDP

[ONO82] K. Ogawa, S. Negishi, and M. Obika. Immunological dissimilarity in protein component (dynein 1) between outer and inner arms within sea urchin sperm axonemes. Journal of Cell Biology, 92(3):706–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/92/3/706.

Owens:1982:PCR

[OO82] G. C. Owens and I. Ohad. Phosphorylation of chlamydomonas reinhardi chloroplast membrane proteins in vivo and in vitro. Journal of Cell Biology, 93(3):712–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/93/3/712.

Oliver:1981:PNP

[OOP^{+}81] J. M. Oliver, W. R. Osborne, J. R. Pfeiffer, F. M. Child, and R. D. Berlin. Purine nucleoside phosphorylase is associated with centrioles and basal bodies. Journal of Cell Biology, 91(3):837–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/91/3/837.

Olins:1980:VNT

[OOZF80] A. L. Olins, D. E. Olins, H. Zentgraf, and W. W. Franke. Visualization of nucleosomes in thin sections by stereo electron microscopy. Journal of Cell Biology, 87(3):833–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/87/3/
833.

Ohkuma:1981:CVM

[OP81] S. Ohkuma and B. Poole. Cytoplasmic vacuolation of mouse peritoneal macrophages and the uptake into lysosomes of weakly basic substances. Journal of Cell Biology, 90(3):656– ??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/3/656.

Ornberg:1981:BEC

[OR81] R. L. Ornberg and T. S. Reese. Beginning of exocytosis captured by rapid-freezing of Limulus amebocytes. Journal of Cell Biology, 90(1):40–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/90/1/40.

Orci:1984:ICP

[ORA^{+}84] L. Orci, M. Ravazzola, M. Amherdt, C. Yanaihara, N. Yanaihara, P. Halban, A. E. Renold, and A. Perrelet. Insulin, not C-peptide (proinsulin), is present in crinophagic bodies of the pancreatic B-cell. Journal of Cell Biology, 98(1):222–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
98/1/222.

Oates:1980:VFA

[OT80a] P. J. Oates and O. Touster. In vitro fusion of Acanthamoeba phagolysosomes. III. Evidence that cyclic nucleotides and vacuole subpopulations respectively control the rate and the extent of vacuole fusion in Acanthamoeba homogenates. Journal of Cell Biology, 85(3):804–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/85/3/804.

Orkin:1980:CEF

[OT80b] R. W. Orkin and B. P. Toole. Chick embryo fibroblasts produce two forms of hyaluronidase. Journal of Cell Biology, 85 (2):248–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/85/2/248.

Ogihara:1982:NDA

[OT82] S. Ogihara and Y. Tonomura. A novel 36,000-dalton actinbinding protein purified from microfilaments in Physarum plasmodia which aggregates actin filaments and blocks actinmyosin interaction. Journal of Cell Biology, 93(3):604–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/93/3/

Otto:1983:DVB

[Ott83] J. J. Otto. Detection of vinculin-binding proteins with an 125Ivinculin gel overlay technique. Journal of Cell Biology, 97(4): 1283–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/4/1283.

Oohira:1982:BUS

[OWMB82] A. Oohira, T. N. Wight, J. McPherson, and P. Bornstein. Biochemical and ultrastructural studies of proteoheparan sulfates synthesized by PYS-2, a basement membrane-producing cell line. Journal of Cell Biology, 92(2):357–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/92/2/357.

Omlin:1982:ILB

[OWPC82] F. X. Omlin, H. D. Webster, C. G. Palkovits, and S. R. Cohen. Immunocytochemical localization of basic protein in major dense line regions of central and peripheral myelin. Journal of Cell Biology, 95(1):242–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/95/1/242.

Ogou:1983:CDC

[OYNT83] S. I. Ogou, C. Yoshida-Noro, and M. Takeichi. Calciumdependent cell–cell adhesion molecules common to hepatocytes and teratocarcinoma stem cells. Journal of Cell Biology, 97(3): 944–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/3/944.

Pollack:1983:CGF

[PA83] L. Pollack and P. H. Atkinson. Correlation of glycosylation forms with position in amino acid sequence. Journal of Cell Biology, 97(2):293–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/2/293.

Paine:1983:PLD

[PADH83] P. L. Paine, C. F. Austerberry, L. J. Desjarlais, and S. B. Horowitz. Protein loss during nuclear isolation. Journal of Cell Biology, 97(4):1240–??, October 1983. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/97/4/1240.

Papasozomenos:1981:RAO

[PAGG81] S. C. Papasozomenos, L. Autilio-Gambetti, and P. Gambetti. Reorganization of axoplasmic organelles following beta, beta’- iminodipropionitrile administration. Journal of Cell Biology, 91(3):866–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/91/3/866.

Paine:1984:DNP

[Pai84] P. L. Paine. Diffusive and nondiffusive proteins in vivo. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/1/188s.

Parysek:1984:MOM

[PAO84] L. M. Parysek, C. F. Asnes, and J. B. Olmsted. MAP 4: occurrence in mouse tissues. Journal of Cell Biology, 99(4): 1309–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/4/1309.

Pesonen:1984:TGPa

[PAS84] M. Pesonen, W. Ansorge, and K. Simons. Transcytosis of the G protein of vesicular stomatitis virus after implantation into the apical plasma membrane of Madin–Darby canine kidney cells. I. Involvement of endosomes and lysosomes. Journal of Cell Biology, 99(3):796–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/3/796.

Pozzan:1982:AIC

[PATR82] T. Pozzan, P. Arslan, R. Y. Tsien, and T. J. Rink. Antiimmunoglobulin, cytoplasmic free calcium, and capping in B lymphocytes. Journal of Cell Biology, 94(2):335–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/94/2/

Porter:1981:IRB

[PB81] K. R. Porter and H. S. Bennett. Introduction: recollections on the beginnings of the Journal of Cell Biology. Journal of Cell Biology, 91(3):??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/91/3/vii.

Paiement:1983:LGS

[PB83a] J. Paiement and J. J. Bergeron. Localization of GTPstimulated core glycosylation to fused microsomes. Journal of Cell Biology, 96(6):1791–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/6/1791.

Pumplin:1983:LDA

[PB83b] D. W. Pumplin and R. J. Bloch. Lipid domains of acetylcholine receptor clusters detected with saponin and filipin. Journal of Cell Biology, 97(4):1043–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/97/4/1043.

Peterson:1984:NBL

[PB84a] A. C. Peterson and G. M. Bray. Normal basal laminas are realized on dystrophic Schwann cells in dystrophic in equilibrium shiverer chimera nerves. Journal of Cell Biology, 99(5): 1831–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/5/1831.

Porrello:1984:RRC

[PB84b] K. Porrello and B. Burnside. Regulation of reactivated contraction in teleost retinal cone models by calcium and cyclic adenosine monophosphate. Journal of Cell Biology, 98(6):2230–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
98/6/2230.

Patzak:1984:EER

[PBFC^{+}84] A. Patzak, G. Böck, R. Fischer-Colbrie, K. Schauenstein, W. Schmidt, G. Lingg, and H. Winkler. Exocytotic exposure and retrieval of membrane antigens of chromaffin granules: quantitative evaluation of immunofluorescence on the surface of chromaffin cells. Journal of Cell Biology, 98(5):1817–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
98/5/1817.

Paiement:1980:CMV

[PBG80] J. Paiement, H. Beaufay, and D. Godelaine. Coalescence of microsomal vesicles from rat liver: a phenomenon occurring in parallel with enhancement of the glycosylation activity during incubation of stripped rough microsomes with GTP. Journal of Cell Biology, 86(1):29–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/86/1/29.

Pardue:1983:PMA

[PBPD83] R. L. Pardue, R. C. Brady, G. W. Perry, and J. R. Dedman. Production of monoclonal antibodies against calmodulin by in vitro immunization of spleen cells. Journal of Cell Biology, 96 (4):1149–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/4/1149.

Pesonen:1984:TGPb

[PBS84] M. Pesonen, R. Bravo, and K. Simons. Transcytosis of the G protein of vesicular stomatitis virus after implantation into the apical membrane of Madin–Darby canine kidney cells. II. Involvement of the Golgi complex. Journal of Cell Biology, 99(3): 803–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/3/803.

Patsalos:1980:PMB

[PBW80] P. N. Patsalos, M. E. Bell, and R. C. Wiggins. Pattern of myelin breakdown during sciatic nerve Wallerian degeneration: reversal of the order of assembly. Journal of Cell Biology, 87 (1):1–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/87/1/1.

Phillips:1983:VEE

[PBWK83] J. H. Phillips, K. Burridge, S. P. Wilson, and N. Kirshner. Visualization of the exocytosis/endocytosis secretory cycle in cultured adrenal chromaffin cells. Journal of Cell Biology, 97(6):

1906–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/6/1906.

Peters:1980:TDG

[PC80] K. E. Peters and D. E. Commings. Two-dimensinal gel electrophoresis of rat liver nuclear washes, nuclear matrix, and hnRNA proteins. Journal of Cell Biology, 86(1):135–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/86/1/
135.

Porrello:1983:EMS

[PCB83] K. Porrello, W. Z. Cande, and B. Burnside. N-ethylmaleimide-modified subfragment-1 and heavy meromyosin inhibit reactivated contraction in motile models of retinal cones. Journal of Cell Biology, 96(2):449–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/2/449.

Penningroth:1982:MCR

[PCOK82] S. M. Penningroth, A. Cheung, K. Olehnik, and R. Koslosky. Mechanochemical coupling in the relaxation of rigor-wave sea urchin sperm flagella. Journal of Cell Biology, 92(3):733–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
92/3/733.

Pederson:1980:MRP

[PD80] T. Pederson and N. G. Davis. Messenger RNA processing and nuclear structure: isolation of nuclear ribonucleoprotein particles containing beta-globin messenger RNA precursors. Journal of Cell Biology, 87(1):47–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/87/1/47.

Pfeffer:1983:ITC

[PDK83] S. R. Pfeffer, D. G. Drubin, and R. B. Kelly. Identification of three coated vesicle components as alpha- and beta-tubulin linked to a phosphorylated 50,000-dalton polypeptide. Journal of Cell Biology, 97(1):40–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/97/1/40.

Pavelka:1983:ECG

[PE83] M. Pavelka and A. Ellinger. Effect of colchicine on the Golgi complex of rat pancreatic acinar cells. Journal of Cell Biology, 97(3):737–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/3/737.

Pederson:1983:NRP

[Ped83] T. Pederson. Nuclear RNA-protein interactions and messenger RNA processing. Journal of Cell Biology, 97(5):1321– ??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/5/1321.

Peng:1983:COP

[Pen83] H. B. Peng. Cytoskeletal organization of the presynaptic nerve terminal and the acetylcholine receptor cluster in cell cultures. Journal of Cell Biology, 97(2):489–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/2/489.

Peng:1984:PCC

[Pen84] H. B. Peng. Participation of calcium and calmodulin in the formation of acetylcholine receptor clusters. Journal of Cell Biology, 98(2):550–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/2/550.

Perkins:1981:IEE

[Per81a] M. Perkins. Inhibitory effects of erythrocyte membrane proteins on the in vitro invasion of the human malarial parasite (Plasmodium falciparum) into its host cell. Journal of Cell Biology, 90(3):563–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/90/3/563.

Perry:1981:RPC

[Per81b] R. P. Perry. RNA processing comes of age. Journal of Cell Biology, 91(3):??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/91/3/28s.

Patel:1981:SPS

[PF81] V. P. Patel and G. Fairbanks. Spectrin phosphorylation and shape change of human erythrocyte ghosts. Journal of Cell Biology, 88(2):430–??, February 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/88/2/430.

Palm:1983:PLF

[PF83a] S. L. Palm and L. T. Furcht. Production of laminin and fibronectin by Schwannoma cells: cell-protein interactions in vitro and protein localization in peripheral nerve in vivo. Journal of Cell Biology, 96(5):1218–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/5/1218.

Pumplin:1983:KAC

[PF83b] D. W. Pumplin and D. M. Fambrough. (Na^{+} + K^{+})-ATPase correlated with a major group of intramembrane particles in freeze-fracture replicas of cultured chick myotubes. Journal of Cell Biology, 97(4):1214–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/97/4/1214.

Pfefferkorn:1984:TSI

[Pfe84] L. C. Pfefferkorn. Transmembrane signaling: an ion-flux-independent model for signal transduction by complexed Fc receptors. Journal of Cell Biology, 99(6):2231–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/6/
2231.

Pearl:1984:SSL

[PFM^{+}84] M. Pearl, D. Fishkind, M. Mooseker, D. Keene, and T. Keller. Studies on the spectrin-like protein from the intestinal brush border, TW 260/240, and characterization of its interaction with the cytoskeleton and actin. Journal of Cell Biology, 98 (1):66–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/1/66.

Prives:1982:ICF

[PFP^{+}82] J. Prives, A. B. Fulton, S. Penman, M. P. Daniels, and C. N. Christian. Interaction of the cytoskeletal framework with acetylcholine receptor on th surface of embryonic muscle cells in culture. Journal of Cell Biology, 92(1):231–??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/92/1/
231.

Painter:1982:CII

[PG82] R. G. Painter and M. Ginsberg. Concanavalin A induces interactions between surface glycoproteins and the platelet cytoskeleton. Journal of Cell Biology, 92(2):565–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/92/2/
565.

Paul:1983:VSM

[PG83a] D. L. Paul and D. A. Goodenough. In vitro synthesis and membrane insertion of bovine MP26, an integral protein from lens fiber plasma membrane. Journal of Cell Biology, 96(3):633–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
96/3/633.

Paul:1983:PCL

[PG83b] D. L. Paul and D. A. Goodenough. Preparation, characterization, and localization of antisera against bovine MP26, an integral protein from lens fiber plasma membrane. Journal of Cell Biology, 96(3):625–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/3/625.

Pinder:1983:SDS

[PG83c] J. C. Pinder and W. B. Gratzer. Structural and dynamic states of actin in the erythrocyte. Journal of Cell Biology, 96 (3):768–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/3/768.

Pemrick:1984:QAS

[PG84] S. M. Pemrick and R. C. Grebenau. Qualitative analysis of skeletal myosin as substrate of Ca^{2+}-activated neutral protease: comparison of filamentous and soluble, native, and L2deficient myosin. Journal of Cell Biology, 99(6):2297–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/6/2297.

Pozzan:1984:CDI

[PGD^{+}84] T. Pozzan, G. Gatti, N. Dozio, L. M. Vicentini, and J. Meldolesi. Ca^{2+}-dependent and -independent release of neurotransmitters from PC12 cells: a role for protein kinase C activation? Journal of Cell Biology, 99(2):628–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/99/2/628.

Parness:1981:TBP

[PH81] J. Parness and S. B. Horwitz. Taxol binds to polymerized tubulin in vitro. Journal of Cell Biology, 91(2):479–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/91/2/
479.

Petty:1981:DMS

[PHM81] H. R. Petty, D. G. Hafeman, and H. M. McConnell. Disappearance of macrophage surface folds after antibody-dependent phagocytosis. Journal of Cell Biology, 89(2):223–??, May 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/89/2/223.

Piperno:1981:RSC

[PHRL81] G. Piperno, B. Huang, Z. Ramanis, and D. J. Luck. Radial spokes of Chlamydomonas flagella: polypeptide composition and phosphorylation of stalk components. Journal of Cell Biology, 88(1):73–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/88/1/73.

Pickett-Heaps:1984:CMS

[PHST84] J. Pickett-Heaps, T. Spurck, and D. Tippit. Chromosome motion and the spindle matrix. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/1/137s.

Pierce:1982:QGG

[Pie82] M. Pierce. Quantification of ganglioside GM1 synthetase activity on intact chick neural retinal cells. Journal of Cell Biology, 93(1):76–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/93/1/76.

Piperno:1984:MAD

[Pip84] G. Piperno. Monoclonal antibodies to dynein subunits reveal the existence of cytoplasmic antigens in sea urchin egg. Journal of Cell Biology, 98(5):1842–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/5/1842.

Pfenninger:1983:MBS

[PJ83] K. H. Pfenninger and M. P. Johnson. Membrane biogenesis in the sprouting neuron. I. Selective transfer of newly synthesized phospholipid into the growing neurite. Journal of Cell Biology, 97(4):1038–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/4/1038.

Phillips:1980:IMP

[PJE80] D. Phillips, L. Jennings, and H. Edwards. Identification of membrane proteins mediating the interaction of human platelets. Journal of Cell Biology, 86(1):77–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/86/1/77.

Pfeiffer:1981:DTL

[PJM^{+}81] S. E. Pfeiffer, H. Jakob, K. Mikoshiba, P. Dubois, J. L. Guenet, J. F. Nicolas, J. Gaillard, G. Chevance, and F. Jacob. Differentiation of a teratocarcinoma line: preferential development of cholinergic neurons. Journal of Cell Biology, 88(1):57–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
88/1/57.

Pfeffer:1981:IMC

[PK81] S. R. Pfeffer and R. B. Kelly. Identification of minor components of coated vesicles by use of permeation chromatography. Journal of Cell Biology, 91(2):385–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/91/2/385.

Pepper:1984:SMA

[PKB84] D. A. Pepper, H. Y. Kim, and M. W. Berns. Studies of a microtubule-associated protein using a monoclonal antibody elicited against mammalian mitotic spindles. Journal of Cell Biology, 99(2):503–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/2/503.

Peltz:1981:DMC

[PKS81] G. Peltz, E. R. Kuczmarski, and J. A. Spudich. Dictyostelium myosin: characterization of chymotryptic fragments and localization of the heavy-chain phosphorylation site. Journal of Cell Biology, 89(1):104–??, April 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/89/1/104.

Paulson:1983:LAX

[PL83a] J. R. Paulson and J. P. Langmore. Low angle X-ray diffraction studies of HeLa metaphase chromosomes: effects of histone phosphorylation and chromosome isolation procedure. Journal of Cell Biology, 96(4):1132–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/4/1132.

Peterson:1983:IDS

[PL83b] S. W. Peterson and V. Lerch. Inhibition of DNA synthesis in SV3T3 cultures by isolated 3T3 plasma membranes. Journal of Cell Biology, 97(1):276–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/97/1/276.

Price:1983:EIF

[PL83c] M. G. Price and E. Lazarides. Expression of intermediate filament-associated proteins paranemin and synemin in chicken development. Journal of Cell Biology, 97(6):1860–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
97/6/1860.

Pesacreta:1984:PMC

[PL84] T. C. Pesacreta and W. J. Lucas. Plasma membrane coat and a coated vesicle-associated reticulum of membranes: their structure and possible interrelationship in Chara corallina. Journal of Cell Biology, 98(4):1537–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/4/1537.

Pagano:1981:MIL

[PLMS81] R. E. Pagano, K. J. Longmuir, O. C. Martin, and D. K. Struck. Metabolism and intracellular localization of a fluorescently labeled intermediate in lipid biosynthesis within cultured fibroblasts. Journal of Cell Biology, 91(3):872–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/91/3/872.

Pantazis:1981:ILD

[PLP81] P. Pantazis, S. A. Lazarou, and N. M. Papadopoulos. Isoenzymes of lactate dehydrogenase in human leukemic cells in culture treated with inducers of differentiation. Journal of Cell Biology, 90(2):396–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/90/2/396.

Pollard:1981:DMA

[PM81] T. D. Pollard and M. S. Mooseker. Direct measurement of actin polymerization rate constants by electron microscopy of actin filaments nucleated by isolated microvillus cores. Journal of Cell Biology, 88(3):654–??, March 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/88/3/654.

Pullman:1983:RNC

[PM83] J. M. Pullman and T. E. Martin. Reconstitution of nucleoprotein complexes with mammalian heterogeneous nuclear ribonucleoprotein (hnRNP) core proteins. Journal of Cell Biology, 97(1):99–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/1/99.

Pfenninger:1981:LLSa

[PMP81a] K. H. Pfenninger and M. F. Maylié-Pfenninger. Lectin labeling of sprouting neurons. I. Regional distribution of surface glycoconjugates. Journal of Cell Biology, 89(3):536–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/89/3/

Pfenninger:1981:LLSb

[PMP81b] K. H. Pfenninger and M. F. Maylié-Pfenninger. Lectin labeling of sprouting neurons. II. Relative movement and appearance of glycoconjugates during plasmalemmal expansion. Journal of Cell Biology, 89(3):547–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/89/3/547.

Prochniewicz-Nakayama:1983:SCF

[PNYO83] E. Prochniewicz-Nakayama, T. Yanagida, and F. Oosawa. Studies on conformation of F-actin in muscle fibers in the relaxed state, rigor, and during contraction using fluorescent phalloidin. Journal of Cell Biology, 97(6):1663–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/6/
1663.

Poole:1981:EWB

[PO81] B. Poole and S. Ohkuma. Effect of weak bases on the intralysosomal pH in mouse peritoneal macrophages. Journal of Cell Biology, 90(3):665–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/90/3/665.

Pollard:1981:CCP

[Pol81] T. D. Pollard. Cytoplasmic contractile proteins. Journal of Cell Biology, 91(3):??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/91/3/156s.

Pollard:1982:SPA

[Pol82] T. D. Pollard. Structure and polymerization of Acanthamoeba myosin–II filaments. Journal of Cell Biology, 95(3):816–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
95/3/816.

Pollard:1984:PAA

[Pol84a] T. D. Pollard. Polymerization of ADP-actin. Journal of Cell Biology, 99(3):769–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/99/3/769.

Pollard:1984:PHM

[Pol84b] T. D. Pollard. Purification of a high molecular weight actin filament gelation protein from Acanthamoeba that shares antigenic determinants with vertebrate spectrins. Journal of Cell Biology, 99(6):1970–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/6/1970.

Porter:1984:CSH

[Por84] K. R. Porter. The cytomatrix: a short history of its study. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/1/3s.

Pratt:1980:DLM

[POS80] M. M. Pratt, T. Otter, and E. D. Salmon. Dynein-like Mg^{2+}- ATPase in mitotic spindles isolated from sea urchin embryos (Strongylocentrotus droebachiensis). Journal of Cell Biology, 86(3):738–??, September 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/86/3/738.

Phillips:1981:ECB

[POYF81] M. J. Phillips, M. Oda, I. M. Yousef, and K. Funatsu. Effects of cytochalasin B on membrane-associated microfilaments in a cell-free system. Journal of Cell Biology, 91(2):524–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
91/2/524.

Peracchia:1980:GJDa

[PP80a] C. Peracchia and L. L. Peracchia. Gap junction dynamics: reversible effects of divalent cations. Journal of Cell Biology, 87(3):708–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/87/3/708.

Peracchia:1980:GJDb

[PP80b] C. Peracchia and L. L. Peracchia. Gap junction dynamics: reversible effects of hydrogen ions. Journal of Cell Biology, 87(3): 719–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/87/3/719.

Pease:1981:EMU

[PP81] D. C. Pease and K. R. Porter. Electron microscopy and ultramicrotomy. Journal of Cell Biology, 91(3):??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/91/3/
287s.

Peng:1984:ECS

[PP84] H. B. Peng and K. A. Phelan. Early cytoplasmic specialization at the presumptive acetylcholine receptor cluster: a meshwork of thin filaments. Journal of Cell Biology, 99(1):344–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/1/
344.

Poole:1982:MEA

[PPR^{+}82a] A. R. Poole, I. Pidoux, A. Reiner, L. Cöster, and J. R. Hassell. Mammalian eyes and associated tissues contain molecules that are immunologically related to cartilage proteoglycan and link protein. Journal of Cell Biology, 93(3):910–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/93/3/910.

Poole:1982:RPE

[PPR82b] A. R. Poole, I. Pidoux, and L. Rosenberg. Role of proteoglycans in endochondral ossification: immunofluorescent localization of link protein and proteoglycan monomer in bovine fetal epiphyseal growth plate. Journal of Cell Biology, 92(2): 249–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/2/249.

Poole:1984:AEP

[PPR^{+}84] A. R. Poole, I. Pidoux, A. Reiner, H. Choi, and L. C. Rosenberg. Association of an extracellular protein (chondrocalcin) with the calcification of cartilage in endochondral bone formation. Journal of Cell Biology, 98(1):54–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/98/1/54.

Poole:1982:IMS

[PPRR82] A. R. Poole, I. Pidoux, A. Reiner, and L. Rosenberg. An immunoelectron microscope study of the organization of proteoglycan monomer, link protein, and collagen in the matrix of articular cartilage. Journal of Cell Biology, 93(3):921–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/93/3/
921.

Peters:1983:FSP

[PPSP83] K. R. Peters, G. E. Palade, B. G. Schneider, and D. S. Papermaster. Fine structure of a periciliary ridge complex of frog retinal rod cells revealed by ultrahigh resolution scanning electron microscopy. Journal of Cell Biology, 96(1):265–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
96/1/265.

Phillips:1983:CDC

[PR83] M. A. Phillips and R. H. Rice. Convergent differentiation in cultured rat cells from nonkeratinized epithelia: keratinocyte character and intrinsic differences. Journal of Cell Biology, 97(3):686–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/3/686.

Parsegian:1984:WNI

[PR84] V. A. Parsegian and D. C. Rau. Water near intracellular surfaces. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/1/196s.

Paiement:1982:MFG

[PRK^{+}82] J. Paiement, R. A. Rachubinski, N. M. Ng Ying Kin, R. A. Sikstrom, and J. J. Bergeron. Membrane fusion and glycosylation in the rat hepatic Golgi apparatus. Journal of Cell Biology, 92(1):147–??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/92/1/147.

Pike:1981:ANM

[PS81] M. C. Pike and R. Snyderman. Alterations of new methylated phospholipid synthesis in the plasma membranes of macrophages exposed to chemoattractants. Journal of Cell Biology, 91(1):221–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/91/1/221.

Pardee:1982:MKI

[PS82] J. D. Pardee and J. A. Spudich. Mechanism of K+-induced actin assembly. Journal of Cell Biology, 93(3):648–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/93/3/
648.

Paavola:1983:ULS

[PS83a] L. G. Paavola and J. F. Strauss. Uptake of lipoproteins by in situ perfused rat ovaries: identification of binding sites for high density lipoproteins. Journal of Cell Biology, 97(3): 593–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/3/593.

Pesonen:1983:TTV

[PS83b] M. Pesonen and K. Simons. Transepithelial transport of a viral membrane glycoprotein implanted into the apical plasma membrane of Madin–Darby canine kidney cells. II. Immunological quantitation. Journal of Cell Biology, 97(3):638–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/3/
638.

Pilch:1983:CVP

[PSBF83] P. F. Pilch, M. A. Shia, R. J. Benson, and R. E. Fine. Coated vesicles participate in the receptor-mediated endocytosis of insulin. Journal of Cell Biology, 96(1):133–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/96/1/133.

Pardo:1983:VCE

[PSC83] J. V. Pardo, J. D. Siliciano, and S. W. Craig. Vinculin is a component of an extensive network of myofibril-sarcolemma attachment regions in cardiac muscle fibers. Journal of Cell Biology, 97(4):1081–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/4/1081.

Pollard:1984:IAF

[PSM84] T. D. Pollard, S. C. Selden, and P. Maupin. Interaction of actin filaments with microtubules. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/1/33s.

Patzer:1982:RCC

[PSR82] E. J. Patzer, D. M. Schlossman, and J. E. Rothman. Release of clathrin from coated vesicles dependent upon a nucleoside triphosphate and a cytosol fraction. Journal of Cell Biology, 93(1):230–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/93/1/230.

Pardee:1982:AFU

[PSSS82] J. D. Pardee, P. A. Simpson, L. Stryer, and J. A. Spudich. Actin filaments undergo limited subunit exchange in physiological salt conditions. Journal of Cell Biology, 94(2):316–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
94/2/316.

Pacifici:1983:ICM

[PST^{+}83] M. Pacifici, R. Soltesz, G. Thal, D. J. Shanley, D. Boettiger, and H. Holtzer. Immunological characterization of the major chick cartilage proteoglycan and its intracellular localization in cultured chondroblasts: a comparison with Type II procollagen. Journal of Cell Biology, 97(6):1724–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/6/
1724.

Pitelka:1983:MTI

[PT83] D. R. Pitelka and B. N. Taggart. Mechanical tension induces lateral movement of intramembrane components of the tight junction: studies on mouse mammary cells in culture. Journal of Cell Biology, 96(3):606–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/3/606.

Pitelka:1983:EEC

[PTH83] D. R. Pitelka, B. N. Taggart, and S. T. Hamamoto. Effects of extracellular calcium depletion on membrane topography and occluding junctions of mammary epithelial cells in culture. Journal of Cell Biology, 96(3):613–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/3/613.

Podell:1984:WGA

[PV84] S. B. Podell and V. D. Vacquier. Wheat germ agglutinin blocks the acrosome reaction in Strongylocentrotus purpuratus sperm by binding a 210,000-mol-wt membrane protein. Journal of Cell Biology, 99(5):1598–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/5/1598.

Petersen:1983:SSA

[PvD83] O. W. Petersen and B. van Deurs. Serial-section analysis of coated pits and vesicles involved in adsorptive pinocytosis in cultured fibroblasts. Journal of Cell Biology, 96(1): 277–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/1/277.

Posner:1982:ECU

[PVPB82] B. I. Posner, A. K. Verma, B. A. Patel, and J. J. Bergeron. Effect of colchicine on the uptake of prolactin and insulin into Golgi fractions of rat liver. Journal of Cell Biology, 93 (3):560–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/93/3/560.

Painter:1981:ECB

[PWM81] R. G. Painter, J. Whisenand, and A. T. McIntosh. Effects of cytochalasin B on actin and myosin association with particle binding sites in mouse macrophages: implications with regard to the mechanism of action of the cytochalasins. Journal of Cell Biology, 91(2):373–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/91/2/373.

Parysek:1984:MMA

[PWO84] L. M. Parysek, J. J. Wolosewick, and J. B. Olmsted. MAP 4: a microtubule-associated protein specific for a subset of tissue microtubules. Journal of Cell Biology, 99(6):2287–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/6/2287.

Phaire-Washington:1980:PMAb

[PWSW80] L. Phaire-Washington, S. C. Silverstein, and E. Wang. Phorbol myristate acetate stimulates microtubule and 10-nm filament extension and lysosome redistribution in mouse macrophages. Journal of Cell Biology, 86(2):641–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/86/2/641.

Pfeffer:1980:IEM

[PWT80] L. M. Pfeffer, E. Wang, and I. Tamm. Interferon effects on microfilament organization cellular fibronectin distribution, and cell motility in human fibroblasts. Journal of Cell Biology, 85(1):9–??, April 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/85/1/9.

Plattner:1982:CSV

[PWT82] H. Plattner, C. Westphal, and R. Tiggemann. Cytoskeletonsecretory vesicle interactions during the docking of secretory vesicles at the cell membrane in Paramecium tetraurelia cells. Journal of Cell Biology, 92(2):368–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/92/2/368.

Phaire-Washington:1980:PMAa

[PWWS80] L. Phaire-Washington, E. Wang, and S. C. Silverstein. Phorbol myristate acetate stimulates pinocytosis and membrane spreading in mouse peritoneal macrophages. Journal of Cell Biology, 86(2):634–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/86/2/634.

Papasozomenos:1982:RPF

[PYC^{+}82] S. C. Papasozomenos, M. Yoon, R. Crane, L. Autilio-Gambetti, and P. Gambetti. Redistribution of proteins of fast axonal transport following administration of beta,beta’- iminodipropionitrile: a quantitative autoradiographic study. Journal of Cell Biology, 95(2):672–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/95/2/672.

Quintart:1984:RME

[QCB84] J. Quintart, P. J. Courtoy, and P. Baudhuin. Receptormediated endocytosis in rat liver: purification and enzymic characterization of low density organelles involved in uptake of galactose-exposing proteins. Journal of Cell Biology, 98 (3):877–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/3/877.

Quinn:1983:DGC

[QGW83] P. Quinn, G. Griffiths, and G. Warren. Dissection of the Golgi complex. II. Density separation of specific Golgi functions in virally infected cells treated with monensin. Journal of Cell Biology, 96(3):851–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/3/851.

Quinn:1984:DNS

[QGW84] P. Quinn, G. Griffiths, and G. Warren. Density of newly synthesized plasma membrane proteins in intracellular membranes II. Biochemical studies. Journal of Cell Biology, 98(6):2142–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
98/6/2142.

Reaven:1981:EVH

[RA81] E. Reaven and S. Azhar. Effect of various hepatic membrane fractions on microtubule assembly-with special emphasis on the role of membrane phospholipids. Journal of Cell Biology, 89(2):300–??, May 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/89/2/300.

Ravoet:1981:QAS

[RACGB81] A. Ravoet, A. Amar-Costesec, D. Godelaine, and H. Beaufay. Quantitative assay and subcellular distribution of enzymes acting on dolichyl phosphate in rat liver. Journal of Cell Biology, 91(3):679–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/91/3/679.

Raff:1984:GMS

[Raf84] E. C. Raff. Genetics of microtubule systems. Journal of Cell Biology, 99(1):1–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/1/1.

Raghavan:1981:DPC

[Rag81] V. Raghavan. Distribution of poly(A)-containing RNA during normal pollen development and during induced pollen embryogenesis in Hyoscyamus niger. Journal of Cell Biology, 89 (3):593–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/89/3/593.

Rao:1981:CPH

[Rao81] M. V. Rao. Cell population heterogeneity in the inducibility of DNA synthesis in human diploid fibroblasts. Journal of Cell Biology, 89(2):194–??, May 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/89/2/194.

Rauvala:1984:NON

[Rau84] H. Rauvala. Neurite outgrowth of neuroblastoma cells: dependence on adhesion surface–cell surface interactions. Journal of Cell Biology, 98(3):1010–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/3/1010.

Rogalski:1980:CLF

[RB80] A. A. Rogalski and G. B. Bouck. Characterization and localization of a flagellar-specific membrane glycoprotein in Euglena. Journal of Cell Biology, 86(2):424–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/86/2/424.

Rogalski:1982:FSA

[RB82a] A. A. Rogalski and G. B. Bouck. Flagellar surface antigens in Euglena: immunological evidence for an external glycoprotein pool and its transfer to the regenerating flagellum. Journal of Cell Biology, 93(3):758–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/93/3/758.

Roth:1982:ILG

[RB82b] J. Roth and E. G. Berger. Immunocytochemical localization of galactosyltransferase in HeLa cells: codistribution with thiamine pyrophosphatase in trans-Golgi cisternae. Journal of Cell Biology, 93(1):223–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/93/1/223.

Roberson:1983:XLL

[RB83] M. M. Roberson and S. H. Barondes. Xenopus laevis lectin is localized at several sites in Xenopus oocytes, eggs, and embryos. Journal of Cell Biology, 97(6):1875–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/6/
1875.

Reggio:1984:AAL

[RBH^{+}84] H. Reggio, D. Bainton, E. Harms, E. Coudrier, and D. Louvard. Antibodies against lysosomal membranes reveal a 100,000-mol-wt protein that cross-reacts with purified H+,K+ ATPase from gastric mucosa. Journal of Cell Biology, 99(4): 1511–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/4/1511.

Roy:1982:SAL

[RBMM82] H. Roy, M. Bloom, P. Milos, and M. Monroe. Studies on the assembly of large subunits of ribulose bisphosphate carboxylase in isolated pea chloroplasts. Journal of Cell Biology, 94(1):20–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/94/1/20.

Roth:1983:RDA

[RBO83] J. Roth, D. Brown, and L. Orci. Regional distribution of N– acetyl–D–galactosamine residues in the glycocalyx of glomerular podocytes. Journal of Cell Biology, 96(5):1189–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/5/

Rothman:1980:TMG

[RBPF80] J. E. Rothman, H. Bursztyn-Pettegrew, and R. E. Fine. Transport of the membrane glycoprotein of vesicular stomatitis virus to the cell surface in two stages by clathrin-coated vesicles. Journal of Cell Biology, 86(1):162–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/86/1/162.

Rodriguez-Boulan:1983:AEV

[RBPS83] E. Rodriguez-Boulan, K. T. Paskiet, and D. D. Sabatini. Assembly of enveloped viruses in Madin–Darby canine kidney cells: polarized budding from single attached cells and from clusters of cells in suspension. Journal of Cell Biology, 96 (3):866–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/3/866.

Rodriguez-Boulan:1984:ITI

[RBPSB84] E. Rodriguez-Boulan, K. T. Paskiet, P. J. Salas, and E. Bard. Intracellular transport of influenza virus hemagglutinin to the apical surface of Madin–Darby canine kidney cells. Journal of Cell Biology, 98(1):308–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/1/308.

Rogalski:1984:EMA

[RBS84] A. A. Rogalski, J. E. Bergmann, and S. J. Singer. Effect of microtubule assembly status on the intracellular processing and surface expression of an integral protein of the plasma membrane. Journal of Cell Biology, 99(3):1101–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/3/
1101.

Rubenstein:1983:CSM

[RC83] J. L. Rubenstein and T. G. Chappell. Construction of a synthetic messenger RNA encoding a membrane protein. Journal of Cell Biology, 96(5):1464–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/5/1464.

Rauvala:1981:SCAa

[RCH81] H. Rauvala, W. G. Carter, and S. I. Hakomori. Studies on cell adhesion and recognition. I. Extent and specificity of cell adhesion triggered by carbohydrate-reactive proteins (glycosidases and lectins) and by fibronectin. Journal of Cell Biology, 88 (1):127–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/88/1/127.

Roberts:1982:DCB

[RCJS82] P. J. Roberts, A. R. Cross, O. T. G. Jones, and A. W. Segal. Development of cytochrome b and an active oxidase system in association with maturation of a human promyelocytic (HL-60) cell line. Journal of Cell Biology, 95(3):720–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/95/3/
720.

Radke:1983:MAD

[RCM^{+}83] K. Radke, V. C. Carter, P. Moss, P. Dehazya, M. Schliwa, and G. S. Martin. Membrane association of a 36,000-dalton substrate for tyrosine phosphorylation in chicken embryo fibroblasts transformed by avian sarcoma viruses. Journal of Cell Biology, 97(5):1601–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/97/5/1601.

Rovasio:1983:NCC

[RDY^{+}83] R. A. Rovasio, A. Delouvee, K. M. Yamada, R. Timpl, and J. P. Thiery. Neural crest cell migration: requirements for exogenous fibronectin and high cell density. Journal of Cell Biology, 96(2):462–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/2/462.

Rutishauser:1980:EFO

[RE80] U. Rutishauser and G. M. Edelman. Effects of fasciculation on the outgrowth of neurites from spinal ganglia in culture. Journal of Cell Biology, 87(2):370–??, November 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/87/2/370.

Rebois:1982:EGR

[Reb82] R. V. Rebois. Establishment of gonadotropin-responsive murine Leydig tumor cell line. Journal of Cell Biology, 94 (1):70–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/94/1/70.

Rivin:1980:CCP

[RF80a] C. J. Rivin and W. L. Fangman. Cell cycle phase expansion in nitrogen-limited cultures of Saccharomyces cerevisiae. Journal of Cell Biology, 85(1):96–??, April 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/85/1/96.

Rivin:1980:RFR

[RF80b] C. J. Rivin and W. L. Fangman. Replication fork rate and origin activation during the S phase of Saccharomyces cerevisiae. Journal of Cell Biology, 85(1):108–??, April 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/85/1/108.

Rothman:1981:TNS

[RF81] J. E. Rothman and E. Fries. Transport of newly synthesized vesicular stomatitis viral glycoprotein to purified Golgi membranes. Journal of Cell Biology, 89(1):162–??, April 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/89/1/162.

Rebhun:1984:ITW

[RF84] L. I. Rebhun and G. W. Fisher. Initiation of two-way cortical traffic after fertilization in sea urchin eggs. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/1/129s.

Rubenstein:1981:PCV

[RFLR81] J. L. Rubenstein, R. E. Fine, B. D. Luskey, and J. E. Rothman. Purification of coated vesicles by agarose gel electrophoresis. Journal of Cell Biology, 89(2):357–??, May 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/89/2/357.

Rachubinski:1984:ACO

[RFML84] R. A. Rachubinski, Y. Fujiki, R. M. Mortensen, and P. B. Lazarow. Acyl-coa oxidase and hydratase-dehydrogenase, two enzymes of the peroxisomal beta-oxidation system, are synthesized on free polysomes of clofibrate-treated rat liver. Journal of Cell Biology, 99(6):2241–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/6/2241.

Rubino:1984:LAM

[RFUC84] S. Rubino, M. Fighetti, E. Unger, and P. Cappuccinelli. Location of actin, myosin, and microtubular structures during directed locomotion of Dictyostelium amebae. Journal of Cell Biology, 98(2):382–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/2/382.

Ratner:1984:PCS

[RGB84] N. Ratner, L. Glaser, and R. P. Bunge. PC12 cells as a source of neurite-derived cell surface mitogen, which stimulates Schwann cell division. Journal of Cell Biology, 98(3): 1150–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/3/1150.

Rutishauser:1983:NCA

[RGE83] U. Rutishauser, M. Grumet, and G. M. Edelman. Neural cell adhesion molecule mediates initial interactions between spinal cord neurons and muscle cells in culture. Journal of Cell Biology, 97(1):145–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/1/145.

Rojkind:1980:CTB

[RGM^{+}80] M. Rojkind, Z. Gatmaitan, S. Mackensen, M. A. Giambrone, P. Ponce, and L. M. Reid. Connective tissue biomatrix: its isolation and utilization for long-term cultures of normal rat hepatocytes. Journal of Cell Biology, 87(1):255–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/87/1/

Raviola:1980:SRF

[RGR80] E. Raviola, D. A. Goodenough, and G. Raviola. Structure of rapidly frozen gap junctions. Journal of Cell Biology, 87(1): 273–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/87/1/273.

Rauvala:1981:SCAb

[RH81] H. Rauvala and S. Hakomori. Studies on cell adhesion and recognition. III. The occurrence of α-mannosidase at the fibroblast cell surface, and its possible role in cell recognition. Journal of Cell Biology, 88(1):149–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/88/1/149.

Roof:1982:SRPa

[RH82] D. J. Roof and J. E. Heuser. Surfaces of rod photoreceptor disk membranes: integral membrane components. Journal of Cell Biology, 95(2):487–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/95/2/487.

Roman:1983:DSM

[RH83] L. M. Roman and A. L. Hubbard. A domain-specific marker for the hepatocyte plasma membrane: localization of leucine aminopeptidase to the bile canalicular domain. Journal of Cell Biology, 96(6):1548–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/6/1548.

Roman:1984:DSMa

[RH84a] L. M. Roman and A. L. Hubbard. A domain-specific marker for the hepatocyte plasma membrane. II. Ultrastructural localization of leucine aminopeptidase to the bile canalicular domain of isolated rat liver plasma membranes. Journal of Cell Biology, 98(4):1488–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/4/1488.

Roman:1984:DSMb

[RH84b] L. M. Roman and A. L. Hubbard. A domain-specific marker for the hepatocyte plasma membrane. III. Isolation of bile canalicular membrane by immunoadsorption. Journal of Cell Biology, 98(4):1497–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/4/1497.

Ring:1982:MCM

[RHK82] D. Ring, R. Hubble, and M. Kirschner. Mitosis in a cell with multiple centrioles. Journal of Cell Biology, 94(3):549– ??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/94/3/549.

Richman:1980:LCC

[Ric80] D. P. Richman. Lymphocyte cell-cycle analysis by flow cytometry. Evidence for a specific postmitotic phase before return to G0. Journal of Cell Biology, 85(2):459–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/85/2/459.

Rindler:1984:VGD

[RIP^{+}84] M. J. Rindler, I. E. Ivanov, H. Plesken, E. Rodriguez-Boulan, and D. D. Sabatini. Viral glycoproteins destined for apical or basolateral plasma membrane domains traverse the same Golgi apparatus during their intracellular transport in doubly infected Madin–Darby canine kidney cells. Journal of Cell Biology, 98(4):1304–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/4/1304.

Rossomando:1981:IND

[RJVS81] E. F. Rossomando, E. G. Jahngen, B. Varnum, and D. R. Soll. Inhibition of a nutrient-dependent pinocytosis in Dictyostelium discoideum by the amino acid analogue hadacidin. Journal of Cell Biology, 91(1):227–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/91/1/227.

Robinson:1980:SFM

[RK80] J. M. Robinson and M. J. Karnovsky. Specializations in filopodial membranes at points of attachment to the substrate. Journal of Cell Biology, 87(3):562–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/87/3/562.

Rubinstein:1981:DMF

[RK81] N. A. Rubinstein and A. M. Kelly. Development of muscle fiber specialization in the rat hindlimb. Journal of Cell Biology, 90 (1):128–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/1/128.

Rodewald:1984:RMT

[RK84] R. Rodewald and J. P. Kraehenbuhl. Receptor-mediated transport of IgG. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/1/
159s.

Reeves:1980:AGF

[RKF80] W. H. Reeves, Y. S. Kanwar, and M. G. Farquhar. Assembly of the glomerular filtration surface. Differentiation of anionic sites in glomerular capillaries of newborn rat kidney. Journal of Cell Biology, 85(3):735–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/85/3/735.

Roof:1982:SRPb

[RKH82] D. J. Roof, J. I. Korenbrot, and J. E. Heuser. Surfaces of rod photoreceptor disk membranes: light-activated enzymes. Journal of Cell Biology, 95(2):501–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/95/2/501.

Robinson:1982:GAP

[RKK82] J. M. Robinson, M. L. Karnovsky, and M. J. Karnovsky. Glycogen accumulation in polymorphonuclear leukocytes, and other intracellular alterations that occur during inflammation. Journal of Cell Biology, 95(3):933–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/95/3/933.

Rapoport:1981:SVI

[RKM^{+}81] J. Rapoport, W. A. Kachadorian, J. Muller, N. Franki, and R. M. Hays. Stabilization of vasopressin-induced membrane events by bifunctional imidoesters. Journal of Cell Biology, 89(2):261–??, May 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/89/2/261.

Rodman:1984:PEC

[RKMF84] J. S. Rodman, D. Kerjaschki, E. Merisko, and M. G. Farquhar. Presence of an extensive clathrin coat on the apical plasmalemma of the rat kidney proximal tubule cell. Journal of Cell Biology, 98(5):1630–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/5/1630.

Rosenfeld:1982:BLH

[RKP^{+}82] MG Rosenfeld, G. Kreibich, D. Popov, K. Kato, and DD Sabatini. Biosynthesis of lysosomal hydrolases: their synthesis in bound polysomes and the role of co- and post-translational processing in determining their subcellular distribution. Journal of Cell Biology, 93(1):135–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/93/1/135.

Rubin:1984:CPC

[RLP84] R. W. Rubin, A. K. Lyubkin, and B. C. Pressman. Comparison of the protein content of three different bovine secretory granule membrane types: a search for exocytosis-specific shared proteins. Journal of Cell Biology, 99(1):356–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/1/
356.

Russo:1981:PLM

[RLT81] R. G. Russo, L. A. Liotta, and U. Thorgeirsson. Polymorphonuclear leukocyte migration through human amnion membrane. Journal of Cell Biology, 91(2):459–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/91/2/459.

Raugi:1982:TSS

[RMABB82] G. J. Raugi, S. M. Mumby, D. Abbott-Brown, and P. Bornstein. Thrombospondin: synthesis and secretion by cells in culture. Journal of Cell Biology, 95(1):351–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/95/1/351.

Roll:1980:CCT

[RMAF80] F. J. Roll, J. A. Madri, J. Albert, and H. Furthmayr. Codistribution of collagen types IV and AB2 in basement membranes and mesangium of the kidney. an immunoferritin study of ultrathin frozen sections. Journal of Cell Biology, 85(3):597–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
85/3/597.

Reinach:1983:CCP

[RMF83] F. C. Reinach, T. Masaki, and D. A. Fischman. Characterization of the C-protein from posterior latissimus dorsi muscle of the adult chicken: heterogeneity within a single sarcomere. Journal of Cell Biology, 96(1):297–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/1/297.

Rosenfeld:1984:BPR

[RMH^{+}84] M. G. Rosenfeld, E. E. Marcantonio, J. Hakimi, V. M. Ort, P. H. Atkinson, D. Sabatini, and G. Kreibich. Biosynthesis and processing of ribophorins in the endoplasmic reticulum. Journal of Cell Biology, 99(3):1076–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/3/1076.

Rosenzweig:1983:ILC

[RMJ83] S. A. Rosenzweig, L. J. Miller, and J. D. Jamieson. Identification and localization of cholecystokinin-binding sites on rat pancreatic plasma membranes and acinar cells: a biochemical and autoradiographic study. Journal of Cell Biology, 96 (5):1288–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/5/1288.

Rosenzweig:1984:ACB

[RMJ84] S. A. Rosenzweig, L. D. Madison, and J. D. Jamieson. Analysis of cholecystokinin-binding proteins using endo–beta–N– acetylglucosaminidase F. Journal of Cell Biology, 99(3):1110– ??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/3/1110.

Reinach:1982:ICP

[RMS^{+}82] F. C. Reinach, T. Masaki, S. Shafiq, T. Obinata, and D. A. Fischman. Isoforms of C-protein in adult chicken skeletal muscle: detection with monoclonal antibodies. Journal of Cell Biology, 95(1):78–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/95/1/78.

Ranscht:1984:NSG

[RMT84] B. Ranscht, D. J. Moss, and C. Thomas. A neuronal surface glycoprotein associated with the cytoskeleton. Journal of Cell Biology, 99(5):1803–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/5/1803.

Rothman:1984:ITG

[RMU84] J. E. Rothman, R. L. Miller, and L. J. Urbani. Intercompartmental transport in the Golgi complex is a dissociative process: facile transfer of membrane protein between two Golgi populations. Journal of Cell Biology, 99(1):260–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/99/1/260.

Rauh:1981:CMC

[RN81] J. J. Rauh and D. L. Nelson. Calmodulin is a major component of extruded trichocysts from Paramecium tetraurelia. Journal of Cell Biology, 91(3):860–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/91/3/860.

Rieder:1983:IMF

[RN83] C. L. Rieder and R. Nowogrodzki. Intranuclear membranes and the formation of the first meiotic spindle in Xenos peckii (Acroschismus wheeleri) oocytes. Journal of Cell Biology, 97 (4):1144–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/4/1144.

Ryo:1983:NSP

[RNH^{+}83] R. Ryo, A. Nakeff, S. S. Huang, M. Ginsberg, and T. F. Deuel. New synthesis of a platelet-specific protein: platelet factor 4 synthesis in a megakaryocyte-enriched rabbit bone marrow culture system. Journal of Cell Biology, 96(2):515–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/2/
515.

Robertson:1981:MS

[Rob81] J. D. Robertson. Membrane structure. Journal of Cell Biology, 91(3):??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/91/3/189s.

Robbins:1984:SMC

[ROB^{+}84] A. R. Robbins, C. Oliver, J. L. Bateman, S. S. Krag, C. J. Galloway, and I. Mellman. A single mutation in Chinese hamster ovary cells impairs both Golgi and endosomal functions. Journal of Cell Biology, 99(4):1296–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/4/1296.

Rodewald:1980:DIG

[Rod80] R. Rodewald. Distribution of immunoglobulin G receptors in the small intestine of the young rat. Journal of Cell Biology, 85(1):18–??, April 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/85/1/18.

Roth:1984:CLT

[Rot84] J. Roth. Cytochemical localization of terminal N–acetyl–D– galactosamine residues in cellular compartments of intestinal goblet cells: implications for the topology of O-glycosylation. Journal of Cell Biology, 98(2):399–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/2/399.

Ronne:1983:LDR

[ROW^{+}83] H. Ronne, C. Ocklind, K. Wiman, L. Rask, B. Obrink, and P. A. Peterson. Ligand-dependent regulation of intracellular protein transport: effect of vitamin a on the secretion of the retinol-binding protein. Journal of Cell Biology, 96(3):907–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
96/3/907.

Regula:1981:MAD

[RPB81] C. S. Regula, Pfeiffer, and R. D. Berlin. Microtubule assembly and disassembly at alkaline pH. Journal of Cell Biology, 89 (1):45–??, April 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/89/1/45.

Rodman:1981:TBC

[RPHA81] T. C. Rodman, F. H. Pruslin, H. P. Hoffmann, and V. G. Allfrey. Turnover of basic chromosomal proteins in fertilized eggs: a cytoimmunochemical study of events in vivo. Journal of Cell Biology, 90(2):351–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/90/2/351.

Robbins:1983:MCH

[RPM83] A. R. Robbins, S. S. Peng, and J. L. Marshall. Mutant Chinese hamster ovary cells pleiotropically defective in receptormediated endocytosis. Journal of Cell Biology, 96(4):1064–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
96/4/1064.

Reddy:1981:IIP

[RQHR81] M. K. Reddy, S. A. Qreshi, P. F. Hollenberg, and J. K. Reddy. Immunochemical identity of peroxisomal enoyl–CoA hydratase with the peroxisome-proliferation-associated 80,000 mol wt polypeptide in rat liver. Journal of Cell Biology, 89(3):406–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
89/3/406.

Reaven:1980:EMP

[RR80a] E. P. Reaven and G. M. Reaven. Evidence that microtubules play a permissive role in hepatocyte very low density lipoprotein secretion. Journal of Cell Biology, 84(1):28–??, January 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/84/1/
28.

Rodriguez:1980:PRS

[RR80b] N. Rodríguez and F. L. Renaud. On the possible role of serotonin in the regulation of regeneration of cilia. Journal of Cell Biology, 85(2):242–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/85/2/242.

Roos:1983:CFI

[RRD83] D. S. Roos, J. M. Robinson, and R. L. Davidson. Cell fusion and intramembrane particle distribution in polyethylene glycol-resistant cells. Journal of Cell Biology, 97(3):909– ??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/3/909.

Reddy:1984:IOH

[RRQ^{+}84] J. K. Reddy, M. S. Rao, S. A. Qureshi, M. K. Reddy, D. G. Scarpelli, and N. D. Lalwani. Induction and origin of hepatocytes in rat pancreas. Journal of Cell Biology, 98(6):2082–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
98/6/2082.

Rao:1982:RCI

[RRRS82] M. S. Rao, M. K. Reddy, J. K. Reddy, and D. G. Scarpelli. Response of chemically induced hepatocytelike cells in hamster pancreas to methyl clofenapate, a peroxisome proliferator. Journal of Cell Biology, 95(1):50–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/95/1/50.

Rubenstein:1982:SIR

[RRS82] P. Rubenstein, T. Ruppert, and A. Sandra. Selective isoactin release from cultured embryonic skeletal muscle cells. Journal of Cell Biology, 92(1):164–??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/92/1/164.

Rao:1981:DRC

[RS81] P. N. Rao and M. L. Smith. Differential response of cycling and noncycling cells to inducers of DNA synthesis and mitosis. Journal of Cell Biology, 88(3):649–??, March 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/88/3/649.

Rosenberg:1982:EAB

[RS82] S. Rosenberg and A. Stracher. Effect of actin-binding protein on the sedimentation properties of actin. Journal of Cell Biology, 94(1):51–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/94/1/51.

Rambo:1983:EAE

[RS83] C. O. Rambo and C. M. Szego. Estrogen action at endometrial membranes: alterations in luminal surface detectable within seconds. Journal of Cell Biology, 97(3):679–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/3/
679.

Rogalski:1984:AEG

[RS84] A. A. Rogalski and S. J. Singer. Associations of elements of the Golgi apparatus with microtubules. Journal of Cell Biology, 99(3):1092–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/3/1092.

Repasky:1984:SID

[RSB84] E. A. Repasky, D. E. Symer, and R. B. Bankert. Spectrin immunofluorescence distinguishes a population of naturally capped lymphocytes in situ. Journal of Cell Biology, 99 (1):350–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/1/350.

Rattner:1982:UOY

[RSDH82] J. B. Rattner, C. Saunders, J. R. Davie, and B. A. Hamkalo. Ultrastructural organization of yeast chromatin. Journal of Cell Biology, 93(1):217–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/93/1/217.

Rosenberg:1981:ICA

[RSL81] S. Rosenberg, A. Stracher, and R. C. Lucas. Isolation and characterization of actin and actin-binding protein from human platelets. Journal of Cell Biology, 91(1):201–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/91/1/
201.

Ramanathan:1983:ACM

[RSP^{+}83a] R. Ramanathan, Y. Saimi, J. B. Peterson, D. L. Nelson, and C. Kung. Antibodies to the ciliary membrane of Paramecium tetraurelia alter membrane excitability. Journal of Cell Biology, 97(5):1421–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/5/1421.

Rao:1983:IDS

[RSP83b] P. N. Rao and K. L. Satya-Prakash. Inducers of DNA synthesis: levels higher in transformed cells than in normal cells. Journal of Cell Biology, 96(2):571–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/2/571.

Rink:1982:CPF

[RTP82] T. J. Rink, R. Y. Tsien, and T. Pozzan. Cytoplasmic pH and free Mg2+ in lymphocytes. Journal of Cell Biology, 95(1): 189–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/95/1/189.

Rothman:1984:TPB

[RUB84] J. E. Rothman, L. J. Urbani, and R. Brands. Transport of protein between cytoplasmic membranes of fused cells: correspondence to processes reconstituted in a cell-free system. Journal of Cell Biology, 99(1):248–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/1/248.

Rudzinska:1980:IMM

[Rud80] M. A. Rudzinska. Internalization of macromolecules from the medium in Suctoria. Journal of Cell Biology, 84(1): 172–??, January 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/84/1/172.

Rifas:1982:SED

[RUM^{+}82] L. Rifas, J. Uitto, V. A. Memoli, K. E. Kuettner, R. W. Henry, and W. A. Peck. Selective emergence of differentiated chondrocytes during serum-free culture of cells derived from fetal rat calvaria. Journal of Cell Biology, 92(2):493–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/92/2/
493.

Robertson:1980:ASI

[RV80] J. D. Robertson and J. Vergara. Analysis of the structure of intramembrane particles of the mammalian urinary bladder. Journal of Cell Biology, 86(2):514–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/86/2/514.

Rachubinski:1980:SRL

[RVB80] R. A. Rachubinski, D. P. Verma, and J. J. Bergeron. Synthesis of rat liver microsomal cytochrome b5 by free ribosomes. Journal of Cell Biology, 84(3):705–??, March 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/84/3/705.

Robinson:1984:SPU

[RVH84] J. Robinson, M. Viti, and M. Höök. Structure and properties of an under-sulfated heparan sulfate proteoglycan synthesized by a rat hepatoma cell line. Journal of Cell Biology, 98(3):946–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
98/3/946.

Roos:1983:FAE

[RVM83] D. Roos, A. A. Voetman, and L. J. Meerhof. Functional activity of enucleated human polymorphonuclear leukocytes. Journal of Cell Biology, 97(2):368–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/2/368.

Remillard:1982:STU

[RW82a] S. P. Remillard and G. B. Witman. Synthesis, transport, and utilization of specific flagellar proteins during flagellar regeneration in Chlamydomonas. Journal of Cell Biology, 93 (3):615–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/93/3/615.

Roberts:1982:CFP

[RW82b] T. M. Roberts and S. Ward. Centripetal flow of pseudopodial surface components could propel the amoeboid movement of Caenorhabditis elegans spermatozoa. Journal of Cell Biology, 92(1):132–??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/92/1/132.

Roberts:1982:MFD

[RW82c] T. M. Roberts and S. Ward. Membrane flow during nematode spermiogenesis. Journal of Cell Biology, 92(1):113–??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/92/1/
113.

Resing:1984:ITI

[RWD84] K. A. Resing, K. A. Walsh, and B. A. Dale. Identification of two intermediates during processing of profilaggrin to filaggrin in neonatal mouse epidermis. Journal of Cell Biology, 99(4): 1372–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/4/1372.

Roberts:1983:ASF

[RZSW83] D. M. Roberts, R. E. Zielinski, M. Schleicher, and D. M. Watterson. Analysis of suborganellar fractions from spinach and pea chloroplasts for calmodulin-binding proteins. Journal of Cell Biology, 97(5):1644–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/5/1644.

Smit:1982:CSP

[SA82] J. Smit and N. Agabian. Cell surface patterning and morphogenesis: biogenesis of a periodic surface array during Caulobacter development. Journal of Cell Biology, 95(1):41–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/95/1/
41.

Staehelin:1983:RCM

[SA83a] L. A. Staehelin and C. J. Arntzen. Regulation of chloroplast membrane function: protein phosphorylation changes the spatial organization of membrane components. Journal of Cell Biology, 97(5):1327–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/97/5/1327.

Stya:1983:MDE

[SA83b] M. Stya and D. Axelrod. Mobility and detergent extractability of acetylcholine receptors on cultured rat myotubes: a correlation. Journal of Cell Biology, 97(1):48–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/1/48.

Sanders:1981:IHH

[San81] M. M. Sanders. Identification of histone H2b as a heatshock protein in Drosophila. Journal of Cell Biology, 91(2): 579–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/91/2/579.

Sanes:1982:LFC

[San82] J. R. Sanes. Laminin, fibronectin, and collagen in synaptic and extrasynaptic portions of muscle fiber basement membrane. Journal of Cell Biology, 93(2):442–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/93/2/442.

Sloat:1981:RCG

[SAP81] B. F. Sloat, A. Adams, and J. R. Pringle. Roles of the CDC24 gene product in cellular morphogenesis during the Saccharomyces cerevisiae cell cycle. Journal of Cell Biology, 89 (3):395–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/89/3/395.

Satir:1984:CMO

[Sat84] P. Satir. Cytoplasmic matrix: old and new questions. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/1/235s.

Schor:1983:LMT

[SAW83] S. L. Schor, T. D. Allen, and B. Winn. Lymphocyte migration into three-dimensional collagen matrices: a quantitative study. Journal of Cell Biology, 96(4):1089–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/4/1089.

Salmon:1980:FIC

[SB80a] E. D. Salmon and D. A. Begg. Functional implications of coldstable microtubules in kinetochore fibers of insect spermatocytes during anaphase. Journal of Cell Biology, 85(3):853–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
85/3/853.

Salzer:1980:SSCa

[SB80b] J. L. Salzer and R. P. Bunge. Studies of Schwann cell proliferation. I. An analysis in tissue culture of proliferation during development, Wallerian degeneration, and direct injury. Journal of Cell Biology, 84(3):739–??, March 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/84/3/739.

Springer:1980:CAM

[SB80c] W. R. Springer and S. H. Barondes. Cell adhesion molecules: detection with univalent second antibody. Journal of Cell Biology, 87(3):703–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/87/3/703.

Shadle:1982:AHP

[SB82a] P. J. Shadle and S. H. Barondes. Adhesion of human platelets to immobilized trimeric collagen. Journal of Cell Biology, 95 (1):361–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/95/1/361.

Sohal:1982:INT

[SB82b] G. S. Sohal and W. R. Boydston. Influence of neostigmine treatment on embryonic development of acetylcholine receptors and neuromuscular junctions. Journal of Cell Biology, 94(3):540–??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/94/3/540.

Soranno:1982:CDS

[SB82c] T. Soranno and E. Bell. Cytostructural dynamics of spreading and translocating cells. Journal of Cell Biology, 95(1): 127–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/95/1/127.

Skutelsky:1983:CTR

[SB83a] E. Skutelsky and E. A. Bayer. Cell-type-related segregation of surface galactosyl-containing components at an early developmental stage in hemopoietic bone marrow cells in the rabbit. Journal of Cell Biology, 96(1):184–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/1/184.

Sluder:1983:CMC

[SB83b] G. Sluder and D. A. Begg. Control mechanisms of the cell cycle: role of the spatial arrangement of spindle components in the timing of mitotic events. Journal of Cell Biology, 97(3): 877–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/3/877.

Shadle:1984:PCAa

[SB84] P. J. Shadle and S. H. Barondes. Platelet-collagen adhesion: evidence for participation of antigenically distinct entities. Journal of Cell Biology, 99(6):2048–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/6/
2048.

Snabes:1981:DAB

[SBB81] M. C. Snabes, A. E. Boyd, and J. Bryan. Detection of actin-binding proteins in human platelets by 125I-actin overlay of polyacrylamide gels. Journal of Cell Biology, 90(3): 809–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/3/809.

Steer:1984:DMC

[SBBS84] C. J. Steer, M. Bisher, R. Blumenthal, and A. C. Steven. Detection of membrane cholesterol by filipin in isolated rat liver coated vesicles is dependent upon removal of the clathrin coat. Journal of Cell Biology, 99(1):315–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/1/315.

Sampedro:1981:PDA

[SBC81] M. N. Sampedro, C. M. Bussineau, and C. W. Cotman. Postsynaptic density antigens: preparation and characterization of an antiserum against postsynaptic densities. Journal of Cell Biology, 90(3):675–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/90/3/675.

Snell:1982:LRI

[SBC82] W. J. Snell, M. Buchanan, and A. Clausell. Lidocaine reversibly inhibits fertilization in Chlamydomonas: a possible role for calcium in sexual signalling. Journal of Cell Biology, 94(3):607–??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/94/3/607.

Schwartz:1982:CCE

[SBCS^{+}82] E. Schwartz, R. S. Bienkowski, B. Coltoff-Schiller, S. Goldfischer, and O. O. Blumenfeld. Changes in the components of extracellular matrix and in growth properties of cultured aortic smooth muscle cells upon ascorbate feeding. Journal of Cell Biology, 92(2):462–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/92/2/462.

Schwartz:1984:RAR

[SBF84] A. L. Schwartz, A. Bolognesi, and S. E. Fridovich. Recycling of the asialoglycoprotein receptor and the effect of lysosomotropic amines in hepatoma cells. Journal of Cell Biology, 98(2): 732–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/2/732.

Salzer:1980:SSCc

[SBG80] J. L. Salzer, R. P. Bunge, and L. Glaser. Studies of Schwann cell proliferation. III. Evidence for the surface localization of the neurite mitogen. Journal of Cell Biology, 84(3):767–??, March 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
84/3/767.

Schmidt:1981:BPT

[SBG^{+}81] G. W. Schmidt, S. G. Bartlett, A. R. Grossman, A. R. Cashmore, and N. H. Chua. Biosynthetic pathways of two polypeptide subunits of the light-harvesting chlorophyll a/b protein complex. Journal of Cell Biology, 91(2):468–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/91/2/
468.

Schedl:1984:CCR

[SBGD84] T. Schedl, T. G. Burland, K. Gull, and W. F. Dove. Cell cycle regulation of tubulin RNA level, tubulin protein synthesis, and assembly of microtubules in Physarum. Journal of Cell Biology, 99(1):155–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/1/155.

Small:1984:CPMb

[SBGP84] R. K. Small, M. Blank, R. Ghez, and K. H. Pfenninger. Components of the plasma membrane of growing axons. II. Diffusion of membrane protein complexes. Journal of Cell Biology, 98(4):1434–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/4/1434.

Singer:1984:BFY

[SBJ84] R. A. Singer, D. P. Bedard, and G. C. Johnston. Bud formation by the yeast Saccharomyces cerevisiae is directly dependent on “start”. Journal of Cell Biology, 98(2):678–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/2/
678.

Sandoval:1984:RMO

[SBK^{+}84] I. V. Sandoval, J. S. Bonifacino, R. D. Klausner, M. Henkart, and J. Wehland. Role of microtubules in the organization and localization of the Golgi apparatus. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/1/113s.

Stallcup:1984:GIAb

[SBM84] K. C. Stallcup, S. J. Burakoff, and M. F. Mescher. Growthinhibitory activity of lymphoid cell plasma membranes. II. Partial characterization of the inhibitor. Journal of Cell Biology, 99(4):1227–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/4/1227.

Saubermann:1981:ASEb

[SBP81] A. J. Saubermann, R. Beeuwkes, and P. D. Peters. Application of scanning electron microscopy to X-ray analysis of frozen-hydrated sections. II. Analysis of standard solutions and artificial electrolyte gradients. Journal of Cell Biology, 88(2): 268–??, February 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/88/2/268.

Spiegel:1980:FDS

[SBS80] E. Spiegel, M. Burger, and M. Spiegel. Fibronectin in the developing sea urchin embryo. Journal of Cell Biology, 87(1): 309–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/87/1/309.

Spiegel:1984:AEM

[SBSL84] J. E. Spiegel, D. S. Beardsley, F. S. Southwick, and S. E. Lux. An analogue of the erythroid membrane skeletal protein 4.1 in nonerythroid cells. Journal of Cell Biology, 99(3): 886–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/3/886.

Salisbury:1984:SFR

[SBSM84] J. L. Salisbury, A. Baron, B. Surek, and M. Melkonian. Striated flagellar roots: isolation and partial characterization of a calcium-modulated contractile organelle. Journal of Cell Biology, 99(3):962–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/3/962.

Steck:1982:GCC

[SBVW82] P. A. Steck, J. Blenis, P. G. Voss, and J. L. Wang. Growth control in cultured 3T3 fibroblasts II. Molecular properties of a fraction enriched in growth inhibitory activity. Journal of Cell Biology, 92(2):523–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/92/2/523.

Sizemore:1981:AAN

[SC81] S. R. Sizemore and R. D. Cole. Asynchronous appearance of newly synthesized histone H1 subfractions in HeLa chromatin. Journal of Cell Biology, 90(2):415–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/90/2/415.

Samuelson:1982:LCL

[SC82a] J. C. Samuelson and J. P. Caulfield. Loss of covalently labeled glycoproteins and glycolipids from the surface of newly transformed schistosomula of Schistosoma mansoni. Journal of Cell Biology, 94(2):363–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/94/2/363.

Schreiner:1982:LIB

[SC82b] G. F. Schreiner and R. S. Cotran. Localization of an Ia-bearing glomerular cell in the mesangium. Journal of Cell Biology, 94 (2):483–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/94/2/483.

Stoscheck:1984:REG

[SC84a] C. M. Stoscheck and G. Carpenter. Down regulation of epidermal growth factor receptors: direct demonstration of receptor degradation in human fibroblasts. Journal of Cell Biology, 98 (3):1048–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/3/1048.

Straus:1984:GCV

[SC84b] D. S. Straus and D. L. Coppock. Growth control variant cell line having increased serum requirement and decreased response to platelet-derived growth factor: reversion by 5azacytidine. Journal of Cell Biology, 99(5):1838–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/5/

Sullivan:1984:SHD

[SC84c] K. F. Sullivan and D. W. Cleveland. Sequence of a highly divergent beta tubulin gene reveals regional heterogeneity in the beta tubulin polypeptide. Journal of Cell Biology, 99(5): 1754–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/5/1754.

Samuelson:1982:SSM

[SCD82] J. C. Samuelson, J. P. Caulfield, and J. R. David. Schistosomula of Schistosoma mansoni clear concanavalin A from their surface by sloughing. Journal of Cell Biology, 94(2):355–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
94/2/355.

Scheer:1981:INC

[Sch81a] U. Scheer. Identification of a novel class of tandemly repeated genes transcribed on lampbrush chromosomes of Pleurodeles waltlii. Journal of Cell Biology, 88(3):599–??, March 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/88/3/599.

Schroeder:1981:MMS

[Sch81b] T. E. Schroeder. Microfilament-mediated surface change in starfish oocytes in response to 1-methyladenine: implications for identifying the pathway and receptor sites for maturationinducing hormones. Journal of Cell Biology, 90(2):362–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
90/2/362.

Schwartz:1981:FEV

[Sch81c] E. A. Schwartz. First events in vision: the generation of responses in vertebrate rods. Journal of Cell Biology, 90(2): 271–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/2/271.

Schliwa:1982:ACD

[Sch82] M. Schliwa. Action of cytochalasin D on cytoskeletal networks. Journal of Cell Biology, 92(1):79–??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/92/1/79.

Schmidt:1983:CMP

[Sch83a] S. Y. Schmidt. Cytidine metabolism in photoreceptor cells of the rat. Journal of Cell Biology, 97(3):824–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/3/
824.

Schmidt:1983:LCD

[Sch83b] S. Y. Schmidt. Light- and cytidine-dependent phosphatidylinositol synthesis in photoreceptor cells of the rat. Journal of Cell Biology, 97(3):832–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/3/832.

Schulman:1984:PMA

[Sch84] H. Schulman. Phosphorylation of microtubule-associated proteins by a Ca^{2+}/calmodulin-dependent protein kinase. Journal of Cell Biology, 99(1):11–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/1/11.

Snell:1983:FAC

[SCM83] W. J. Snell, A. Clausell, and W. S. Moore. Flagellar adhesion in chlamydomonas induces synthesis of two high molecular weight cell surface proteins. Journal of Cell Biology, 96 (3):589–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/3/589.

Sugahara:1984:ECF

[SCM84] K. Sugahara, J. H. Caldwell, and R. J. Mason. Electrical currents flow out of domes formed by cultured epithelial cells. Journal of Cell Biology, 99(4):1541–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/4/1541.

Skosey:1981:EOT

[SCN^{+}81] J. L. Skosey, D. C. Chow, S. Nusinow, J. May, V. Gestautas, and Y. Niwa. Effect of oxygen tension on human peripheral blood leukocytes: lysosomal enzyme release and metabolic responses during phagocytosis. Journal of Cell Biology, 88(2): 358–??, February 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/88/2/358.

Singh:1983:PMR

[SCP^{+}83] J. P. Singh, M. A. Chaikin, W. J. Pledger, C. D. Scher, and C. D. Stiles. Persistence of the mitogenic response to plateletderived growth factor (competence) does not reflect a longterm interaction between the growth factor and the target cell. Journal of Cell Biology, 96(5):1497–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/5/1497.

Salisbury:1980:RCV

[SCS80] J. L. Salisbury, J. S. Condeelis, and P. Satir. Role of coated vesicles, microfilaments, and calmodulin in receptor-mediated endocytosis by cultured B lymphoblastoid cells. Journal of Cell Biology, 87(1):132–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/87/1/132.

Singh:1982:PAP

[SCS82] J. P. Singh, M. A. Chaikin, and C. D. Stiles. Phylogenetic analysis of platelet-derived growth factor by radio-receptor assay. Journal of Cell Biology, 95(2):667–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/95/2/667.

Sheetz:1984:ADM

[SCS84] M. P. Sheetz, R. Chasan, and J. A. Spudich. ATP-dependent movement of myosin in vitro: characterization of a quantitative assay. Journal of Cell Biology, 99(5):1867–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/5/
1867.

Summerhayes:1981:EKV

[SCSC81] I. C. Summerhayes, Y. S. Cheng, T. T. Sun, and L. B. Chen. Expression of keratin and vimentin intermediate filaments in rabbit bladder epithelial cells at different stages of benzo[a]pyrene-induced neoplastic progression. Journal of Cell Biology, 90(1):63–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/90/1/63.

Sloboda:1980:SCC

[SD80] R. D. Sloboda and K. Dickersin. Structure and composition of the cytoskeleton of nucleated erythrocytes I. The presence of microtubule-associated protein 2 in the marginal band. Journal of Cell Biology, 87(1):170–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/87/1/170.

Suprenant:1982:ABE

[SD82] K. A. Suprenant and W. L. Dentler. Association between endocrine pancreatic secretory granules and in-vitro-assembled microtubules is dependent upon microtubule-associated proteins. Journal of Cell Biology, 93(1):164–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/93/1/164.

Staufenbiel:1984:PNM

[SD84] M. Staufenbiel and W. Deppert. Preparation of nuclear matrices from cultured cells: subfractionation of nuclei in situ. Journal of Cell Biology, 98(5):1886–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/5/1886.

Segal:1980:KFC

[SDC80] A. W. Segal, J. Dorling, and S. Coade. Kinetics of fusion of the cytoplasmic granules with phagocytic vacuoles in human polymorphonuclear leukocytes. Biochemical and morphological studies. Journal of Cell Biology, 85(1):42–??, April 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/85/1/42.

Simmen:1984:EST

[SDG^{+}84] R. C. Simmen, B. S. Dunbar, V. Guerriero, J. G. Chafouleas, J. H. Clark, and A. R. Means. Estrogen stimulates the transient association of calmodulin and myosin light chain kinase with the chicken liver nuclear matrix. Journal of Cell Biology, 99(2):588–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/2/588.

Smith:1984:EMA

[SDGYJ84] Z. D. Smith, F. D’Eugenio-Gumkowski, K. Yanagisawa, and J. D. Jamieson. Endogenous and monoclonal antibodies to the rat pancreatic acinar cell Golgi complex. Journal of Cell Biology, 98(6):2035–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/6/2035.

Spitzer:1984:HSG

[SDK^{+}84] R. H. Spitzer, S. W. Downing, E. A. Koch, W. L. Salo, and L. J. Saidel. Hagfish slime gland thread cells. II. Isolation and characterization of intermediate filament components associated with the thread. Journal of Cell Biology, 98(2): 670–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/2/670.

Sattilaro:1981:MAP

[SDL81] R. F. Sattilaro, W. L. Dentler, and E. L. LeCluyse. Microtubule-associated proteins (MAPs) and the organization of actin filaments in vitro. Journal of Cell Biology, 90 (2):467–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/2/467.

Scharff:1982:CCC

[SDL82] D. J. Scharff, A. M. Delegeane, and A. S. Lee. Characterization of a cell cycle mutant derived from hamster fibroblast: reversion analysis. Journal of Cell Biology, 92(3):629–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/92/3/
629.

Stallcup:1984:GIAa

[SDM84] K. C. Stallcup, A. Dawson, and M. F. Mescher. Growthinhibitory activity of lymphoid cell plasma membranes. I. Inhibition of lymphocyte and lymphoid tumor cell growth. Journal of Cell Biology, 99(4):1221–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/4/1221.

Scheider:1981:FPM

[SDT81] Y. Scheider, C. Duve, and A. Trouet. Fate of plasma membrane during endocytosis. III. Evidence for incomplete breakdown of immunoglobulins in lysosomes of cultured fibroblasts. Journal of Cell Biology, 88(2):380–??, February 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/88/2/380.

Sullivan:1984:ADC

[SDZ84] S. J. Sullivan, G. Daukas, and S. H. Zigmond. Asymmetric distribution of the chemotactic peptide receptor on polymorphonuclear leukocytes. Journal of Cell Biology, 99(4): 1461–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/4/1461.

Sealock:1982:CSS

[Sea82] R. Sealock. Cytoplasmic surface structure in postsynaptic membranes from electric tissue visualized by tannic-acid-mediated negative contrasting. Journal of Cell Biology, 92(2): 514–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/2/514.

Steiner:1984:CAO

[SEE84] E. K. Steiner, J. C. Eissenberg, and S. C. Elgin. A cytological approach to the ordering of events in gene activation using the Sgs-4 locus of Drosophila melanogaster. Journal of Cell Biology, 99(1):233–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/1/233.

Steere:1980:PCF

[SEM80] R. L. Steere, E. F. Erbe, and J. M. Moseley. Prefracture and cold-fracture images of yeast plasma membranes. Journal of Cell Biology, 86(1):113–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/86/1/113.

Saubermann:1981:ASEa

[SEPB81] A. J. Saubermann, P. Echlin, P. D. Peters, and R. Beeuwkes. Application of scanning electron microscopy to X-ray analysis of frozen-hydrated sections. I. Specimen handling techniques. Journal of Cell Biology, 88(2):257–??, February 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/88/2/257.

Salpeter:1981:HRA

[SF81] M. M. Salpeter and M. G. Farquhar. High resolution analysis of the secretory pathway in mammotrophs of the rat anterior pituitary. Journal of Cell Biology, 91(1):240–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/91/1/240.

Schubart:1984:ICR

[SF84a] U. K. Schubart and K. L. Fields. Identification of a calcium-regulated insulinoma cell phosphoprotein as an islet cell keratin. Journal of Cell Biology, 98(3):1001–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/3/
1001.

Smith:1984:IDR

[SF84b] D. E. Smith and P. A. Fisher. Identification, developmental regulation, and response to heat shock of two antigenically related forms of a major nuclear envelope protein in Drosophila embryos: application of an improved method for affinity purification of antibodies using polypeptides immobilized on nitrocellulose blots. Journal of Cell Biology, 99(1):20–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/1/
20.

Sonderegger:1984:FAP

[SFB^{+}84] P. Sonderegger, M. C. Fishman, M. Bokoum, H. C. Bauer, E. A. Neale, and P. G. Nelson. A few axonal proteins distinguish ventral spinal cord neurons from dorsal root ganglion neurons. Journal of Cell Biology, 98(1):364–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/98/1/364.

Schnitzer:1981:IDV

[SFS81] J. Schnitzer, W. W. Franke, and M. Schachner. Immunocytochemical demonstration of vimentin in astrocytes and ependymal cells of developing and adult mouse nervous system. Journal of Cell Biology, 90(2):435–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/90/2/435.

Schiff:1984:RMB

[SFU84] J. M. Schiff, M. M. Fisher, and B. J. Underdown. Receptormediated biliary transport of immunoglobulin A and asialoglycoprotein: sorting and missorting of ligands revealed by two radiolabeling methods. Journal of Cell Biology, 98(1): 79–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/1/79.

Schloss:1980:MTC

[SG80] J. A. Schloss and R. D. Goldman. Microfilaments and tropomyosin of cultured mammalian cells: isolation and characterization. Journal of Cell Biology, 87(3):633–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/87/3/
633.

Slot:1981:SPC

[SG81] J. W. Slot and H. J. Geuze. Sizing of protein A-colloidal gold probes for immunoelectron microscopy. Journal of Cell Biology, 90(2):533–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/90/2/533.

Schuetze:1982:DTB

[SG82a] S. M. Schuetze and D. A. Goodenough. Dye transfer between cells of the embryonic chick lens becomes less sensitive to CO2 treatment with development. Journal of Cell Biology, 92(3): 694–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/3/694.

Stahl:1982:EMF

[SG82b] P. Stahl and S. Gordon. Expression of a mannosyl-fucosyl receptor for endocytosis on cultured primary macrophages and their hybrids. Journal of Cell Biology, 93(1):49–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/93/1/49.

Seglen:1984:AAC

[SG84a] P. O. Seglen and P. B. Gordon. Amino acid control of autophagic sequestration and protein degradation in isolated rat hepatocytes. Journal of Cell Biology, 99(2):435–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/2/
435.

Stevenson:1984:ZOJ

[SG84b] B. R. Stevenson and D. A. Goodenough. Zonulae occludentes in junctional complex-enriched fractions from mouse liver: preliminary morphological and biochemical characterization. Journal of Cell Biology, 98(4):1209–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/4/1209.

Schonfeld:1980:DAC

[SGA80] G. Schonfeld, N. Grimme, and D. Alpers. Detection of apolipoprotein C in human and rat enterocytes. Journal of Cell Biology, 86(2):562–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/86/2/562.

Sartore:1981:MTF

[SGB^{+}81] S. Sartore, L. Gorza, S. Pierobon Bormioli, L. Dalla Libera, and S. Schiaffino. Myosin types and fiber types in cardiac muscle. I. Ventricular myocardium. Journal of Cell Biology, 88 (1):226–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/88/1/226.

Senior:1983:CAP

[SGH^{+}83] R. M. Senior, G. L. Griffin, J. S. Huang, D. A. Walz, and T. F. Deuel. Chemotactic activity of platelet alpha granule proteins for fibroblasts. Journal of Cell Biology, 96(2):382–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/2/
382.

Salditt-Georgieff:1980:APN

[SGHS^{+}80] M. Salditt-Georgieff, M. Harpold, S. Sawicki, J. Nevins, and J. E. Darnell. Addition of poly(A) to nuclear RNA occurs soon after RNA synthesis. Journal of Cell Biology, 86(3): 844–??, September 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/86/3/844.

Satter:1982:EAF

[SGKC82] R. L. Satter, R. C. Garber, L. Khairallah, and Y. S. Cheng. Elemental analysis of freeze-dried thin sections of Samanea motor organs: barriers to ion diffusion through the apoplast. Journal of Cell Biology, 95(3):893–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/95/3/893.

Spooncer:1983:RHL

[SGKD83] E. Spooncer, J. T. Gallagher, F. Krizsa, and T. M. Dexter. Regulation of haemopoiesis in long-term bone marrow cultures. IV. Glycosaminoglycan synthesis and the stimulation of haemopoiesis by beta–D–xylosides. Journal of Cell Biology, 96(2):510–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/2/510.

Senior:1984:VGV

[SGM^{+}84] R. M. Senior, G. L. Griffin, R. P. Mecham, D. S. Wrenn, K. U. Prasad, and D. W. Urry. Val-Gly-val-ala-Pro-gly, a repeating peptide in elastin, is chemotactic for fibroblasts and monocytes. Journal of Cell Biology, 99(3):870–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/99/3/870.

Shadle:1984:PCAb

[SGPB84] P. J. Shadle, M. H. Ginsberg, E. F. Plow, and S. H. Barondes. Platelet-collagen adhesion: inhibition by a monoclonal antibody that binds glycoprotein IIb. Journal of Cell Biology, 99(6):2056–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/6/2056.

Seppa:1982:PDG

[SGS^{+}82] H. Seppä, G. Grotendorst, S. Seppä, E. Schiffmann, and G. R. Martin. Platelet-derived growth factor in chemotactic for fibroblasts. Journal of Cell Biology, 92(2):584–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/92/2/
584.

Somlyo:1981:CRI

[SGSS^{+}81] A. V. Somlyo, H. G. Gonzalez-Serratos, H. Shuman, G. Mc-Clellan, and A. P. Somlyo. Calcium release and ionic changes in the sarcoplasmic reticulum of tetanized muscle: an electronprobe study. Journal of Cell Biology, 90(3):577–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/90/3/
577.

Schmidt:1982:SIM

[SGW^{+}82] R. A. Schmidt, J. A. Glomset, T. N. Wight, A. J. Habenicht, and R. Ross. A study of the influence of mevalonic acid and its metabolites on the morphology of Swiss 3T3 cells. Journal of Cell Biology, 95(1):144–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/95/1/144.

Sheterline:1981:TLB

[SH81a] P. Sheterline and C. R. Hopkins. Transmembrane linkage between surface glycoproteins and components of the cytoplasm in neutrophil leukocytes. Journal of Cell Biology, 90(3): 743–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/3/743.

Sugrue:1981:RBE

[SH81b] S. P. Sugrue and E. D. Hay. Response of basal epithelial cell surface and Cytoskeleton to solubilized extracellular matrix molecules. Journal of Cell Biology, 91(1):45–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/91/1/45.

Shur:1982:RMS

[SH82a] B. D. Shur and N. G. Hall. A role for mouse sperm surface galactosyltransferase in sperm binding to the egg zona pellucida. Journal of Cell Biology, 95(2):574–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/95/2/574.

Shur:1982:SSG

[SH82b] B. D. Shur and N. G. Hall. Sperm surface galactosyltransferase activities during in vitro capacitation. Journal of Cell Biology, 95(2):567–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/95/2/567.

Smith:1982:EWG

[SH82c] L. Smith and R. M. Hochmuth. Effect of wheat germ agglutinin on the viscoelastic properties of erythrocyte membrane. Journal of Cell Biology, 94(1):7–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/94/1/7.

Salpeter:1983:DTR

[SH83] M. M. Salpeter and R. Harris. Distribution and turnover rate of acetylcholine receptors throughout the junction folds at a vertebrate neuromuscular junction. Journal of Cell Biology, 96(6):1781–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/6/1781.

Springer:1980:ECS

[SHB80] W. R. Springer, P. L. Haywood, and S. H. Barondes. Endogenous cell surface lectin in Dictyostelium: quantitation, elution by sugar, and elicitation by divalent immunoglobulin. Journal of Cell Biology, 87(3):682–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/87/3/682.

Stendahl:1980:DAB

[SHBS80] O. I. Stendahl, J. H. Hartwig, E. A. Brotschi, and T. P. Stossel. Distribution of actin-binding protein and myosin in macrophages during spreading and phagocytosis. Journal of Cell Biology, 84(2):215–??, February 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/84/2/215.

Suard:1983:CPM

[SHFK83] Y. M. L. Suard, M. Haeuptle, E. Farinon, and J. Kraehenbuhl. Cell proliferation and milk protein gene expression in rabbit mammary cell cultures. Journal of Cell Biology, 96(5):1435–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
96/5/1435.

Schneider:1982:RBD

[SHH82] H. Schneider, M. Höchli, and C. R. Hackenbrock. Relationship between the density distribution of intramembrane particles and electron transfer in the mitochondrial inner membrane as revealed by cholesterol incorporation. Journal of Cell Biology, 94(2):387–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/94/2/387.

Salik:1981:EAS

[SHHP81] J. Salik, L. Herlands, H. P. Hoffmann, and D. Poccia. Electrophoretic analysis of the stored histone pool in unfertilized sea urchin eggs: quantification and identification by antibody binding. Journal of Cell Biology, 90(2):385–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/90/2/385.

Scheer:1984:DID

[SHHR84] U. Scheer, B. Hügle, R. Hazan, and K. M. Rose. Drug-induced dispersal of transcribed rRNA genes and transcriptional products: immunolocalization and silver staining of different nucleolar components in rat cells treated with 5,6–dichloro–beta– D–ribofuranosylbenzimidazole. Journal of Cell Biology, 99(2): 672–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/2/672.

Silve:1982:PHR

[SHJA82] C. M. Silve, G. T. Hradek, A. L. Jones, and C. D. Arnaud. Parathyroid hormone receptor in intact embryonic chicken bone: characterization and cellular localization. Journal of Cell Biology, 94(2):379–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/94/2/379.

Spearman:1984:SLS

[SHO^{+}84] T. N. Spearman, K. P. Hurley, R. Olivas, R. G. Ulrich, and F. R. Butcher. Subcellular location of stimulus-affected endogenous phosphoproteins in the rat parotid gland. Journal of Cell Biology, 99(4):1354–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/4/1354.

Sztul:1983:ITT

[SHP83] E. S. Sztul, K. E. Howell, and G. E. Palade. Intracellular and transcellular transport of secretory component and albumin in rat hepatocytes. Journal of Cell Biology, 97(5): 1582–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/5/1582.

Sasse:1984:SPM

[SHPH84] J. Sasse, A. Horwitz, M. Pacifici, and H. Holtzer. Separation of precursor myogenic and chondrogenic cells in early limb bud mesenchyme by a monoclonal antibody. Journal of Cell Biology, 99(5):1856–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/5/1856.

Segal:1984:MSC

[SHRL84] R. A. Segal, B. Huang, Z. Ramanis, and D. J. Luck. Mutant strains of Chlamydomonas reinhardtii that move backwards only. Journal of Cell Biology, 98(6):2026–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/6/
2026.

Steven:1983:EKF

[SHT^{+}83] A. C. Steven, J. F. Hainfeld, B. L. Trus, J. S. Wall, and P. M. Steinert. Epidermal keratin filaments assembled in vitro have masses-per-unit-length that scale according to average subunit mass: structural basis for homologous packing of subunits in intermediate filaments. Journal of Cell Biology, 97(6): 1939–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/6/1939.

Sargent:1984:STN

[SHT^{+}84] P. B. Sargent, B. E. Hedges, L. Tsavaler, L. Clemmons, S. Tzartos, and J. M. Lindstrom. Structure and transmembrane nature of the acetylcholine receptor in amphibian skeletal muscle as revealed by cross-reacting monoclonal antibodies.

Journal of Cell Biology, 98(2):609–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/2/609.

Schuler:1982:VCB

[SHW^{+}82a] G. Schuler, H. Hintner, K. Wolff, P. Fritsch, and G. Stingl. In vitro complement binding on cytoplasmic structures in normal human skin: immunoelectronmicroscopic studies. Journal of Cell Biology, 95(2):543–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/95/2/543.

Scott:1982:CPG

[SHW^{+}82b] R. E. Scott, B. J. Hoerl, J. J. Wille, D. L. Florine, B. R. Krawisz, and K. Yun. Coupling of proadipocyte growth arrest and differentiation. II. A cell cycle model for the physiological control of cell proliferation. Journal of Cell Biology, 94 (2):400–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/94/2/400.

Steven:1983:MDC

[SHWS83] A. C. Steven, J. F. Hainfeld, J. S. Wall, and C. J. Steer. Mass distributions of coated vesicles isolated from liver and brain: analysis by scanning transmission electron microscopy. Journal of Cell Biology, 97(6):1714–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/6/1714.

Schmidt:1983:BPM

[SHZN83] B. Schmidt, B. Hennig, R. Zimmermann, and W. Neupert. Biosynthetic pathway of mitochondrial ATPase subunit 9 in Neurospora crassa. Journal of Cell Biology, 96(1):248–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
96/1/248.

Siekevitz:1981:PI

[Sie81] P. Siekevitz. Preface II. Journal of Cell Biology, 91(3):??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
91/3/vi.

Sutoh:1984:ICH

[SIM^{+}84] K. Sutoh, M. Iwane, F. Matsuzaki, M. Kikuchi, and A. Ikai. Isolation and characterization of a high molecular weight actinbinding protein from Physarum polycephalum plasmodia. Journal of Cell Biology, 98(5):1611–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/5/1611.

Sinensky:1980:AAP

[Sin80] M. Sinensky. Adaptive alteration in phospholipid composition of plasma membranes from a somatic cell mutant defective in the regulation of cholesterol biosynthesis. Journal of Cell Biology, 85(1):166–??, April 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/85/1/166.

Singer:1982:AFV

[Sin82] I. I. Singer. Association of fibronectin and vinculin with focal contacts and stress fibers in stationary hamster fibroblasts. Journal of Cell Biology, 92(2):398–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/92/2/398.

Smith:1980:SSC

[SJ80] R. M. Smith and L. Jarett. Surface structure changes of rat adipocytes during lypolysis stimulated by various lypolysis stimulated by various lypolytic agents. Journal of Cell Biology, 84(1):57–??, January 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/84/1/57.

Scheele:1980:MCS

[SJC80] G. Scheele, R. Jacoby, and T. Carne. Mechanism of compartmentation of secretory proteins: transport of exocrine pancreatic proteins across the microsomal membrane. Journal of Cell Biology, 87(3):611–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/87/3/611.

Steinert:1984:IF

[SJG84] P. M. Steinert, J. C. Jones, and R. D. Goldman. Intermediate filaments. Journal of Cell Biology, 99(1):??, July 1984.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/99/1/22s.

Siegel:1981:ALP

[SJM81] C. O. Siegel, A. E. Jordan, and K. R. Miller. Addition of lipid to the photosynthetic membrane: effects on membrane structure and energy transfer. Journal of Cell Biology, 91(1): 113–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/91/1/113.

Shen:1984:UUF

[SJS84] B. W. Shen, R. Josephs, and T. L. Steck. Ultrastructure of unit fragments of the skeleton of the human erythrocyte membrane. Journal of Cell Biology, 99(3):810–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/99/3/810.

Steimer:1981:SFG

[SK81] K. S. Steimer and M. Klagsbrun. Serum-free growth of normal and transformed fibroblasts in milk: differential requirements for fibronectin. Journal of Cell Biology, 88(2):294–??, February 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/88/2/
294.

Stamatoglou:1983:CBC

[SK83] S. C. Stamatoglou and J. M. Keller. Correlation between cell substrate attachment in vitro and cell surface heparan sulfate affinity for fibronectin and collagen. Journal of Cell Biology, 96(6):1820–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/6/1820.

Sudo:1983:VDC

[SKA^{+}83] H. Sudo, H. A. Kodama, Y. Amagai, S. Yamamoto, and S. Kasai. In vitro differentiation and calcification in a new clonal osteogenic cell line derived from newborn mouse calvaria. Journal of Cell Biology, 96(1):191–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/1/191.

Sauk:1984:RCF

[SKCJW84] J. J. Sauk, M. Krumweide, D. Cocking-Johnson, and J. G. White. Reconstitution of cytokeratin filaments in vitro: further evidence for the role of nonhelical peptides in filament assembly. Journal of Cell Biology, 99(5):1590–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/5/
1590.

Sariola:1984:COF

[SKE84] H. Sariola, P. Kuusela, and P. Ekblom. Cellular origin of fibronectin in interspecies hybrid kidneys. Journal of Cell Biology, 99(6):2099–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/6/2099.

Singer:1984:VCD

[SKKC84] I. I. Singer, D. W. Kawka, D. M. Kazazis, and R. A. Clark. In vivo co-distribution of fibronectin and actin fibers in granulation tissue: immunofluorescence and electron microscope studies of the fibronexus at the myofibroblast surface. Journal of Cell Biology, 98(6):2091–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/6/2091.

Sabatini:1982:MIP

[SKMA82] D. D. Sabatini, G. Kreibich, T. Morimoto, and M. Adesnik. Mechanisms for the incorporation of proteins in membranes and organelles. Journal of Cell Biology, 92(1):1–??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/92/1/
1.

Sato:1983:MOC

[SKN83] C. Sato, R. Kuriyama, and K. Nishizawa. Microtubuleorganizing centers abnormal in number, structure, and nucleating activity in x-irradiated mammalian cells. Journal of Cell Biology, 96(3):776–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/3/776.

Schmid:1982:PIM

[SKS82] H. P. Schmid, K. Köhler, and B. Setyono. Possible involvement of messenger RNA-associated proteins in protein synthesis. Journal of Cell Biology, 93(3):893–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/93/3/893.

Scherson:1984:DIF

[SKS^{+}84a] T. Scherson, T. E. Kreis, J. Schlessinger, U. Z. Littauer, G. G. Borisy, and B. Geiger. Dynamic interactions of fluorescently labeled microtubule-associated proteins in living cells. Journal of Cell Biology, 99(2):425–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/2/425.

Schick:1984:HSE

[SKS84b] J. Schick, H. Kern, and G. Scheele. Hormonal stimulation in the exocrine pancreas results in coordinate and anticoordinate regulation of protein synthesis. Journal of Cell Biology, 99(5): 1569–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/5/1569.

Spiegel:1984:DVR

[SKWF84] S. Spiegel, S. Kassis, M. Wilchek, and P. H. Fishman. Direct visualization of redistribution and capping of fluorescent gangliosides on lymphocytes. Journal of Cell Biology, 99(5): 1575–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/5/1575.

Stirling:1980:AFR

[SL80] C. E. Stirling and A. Lee. [3h]ouabain autoradiography of frog retina. Journal of Cell Biology, 85(2):313–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/85/2/313.

Schubert:1982:PEA

[SL82a] D. Schubert and M. LaCorbiere. Properties of extracellular adhesion-mediating particles in myoblast clone and its adhesion-deficient variant. Journal of Cell Biology, 94(1): 108–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/94/1/108.

Strauch:1982:CSE

[SL82b] A. R. Strauch and J. R. LaFountain. Centrifugation shearing exposes filamentous networks in cortical regions of crane-fly spermatocytes. Journal of Cell Biology, 93(3):670–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/93/3/
670.

Swan:1980:SFU

[SLIF80] M. A. Swan, R. W. Linck, S. Ito, and D. W. Fawcett. Structure and function of the undulating membrane in spermatozoan propulsion in the toad Bufo marinus. Journal of Cell Biology, 85(3):866–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/85/3/866.

Skriver:1984:ILU

[SLK^{+}84] L. Skriver, L. I. Larsson, V. Kielberg, L. S. Nielsen, P. B. Andresen, P. Kristensen, and K. Danø. Immunocytochemical localization of urokinase-type plasminogen activator in Lewis lung carcinoma. Journal of Cell Biology, 99(2):753–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/2/
753.

Schubert:1983:RAN

[SLKB83] D. Schubert, M. LaCorbiere, F. G. Klier, and C. Birdwell. A role for adherons in neural retina cell adhesion. Journal of Cell Biology, 96(4):990–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/4/990.

Strauch:1980:BAA

[SLL80] A. R. Strauch, E. J. Luna, and J. R. LaFountain. Biochemical analysis of actin in crane-fly gonial cells: evidence for actin in spermatocytes and spermatids–but not sperm. Journal of Cell Biology, 86(1):315–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/86/1/315.

Schinz:1982:FFS

[SLLP82] R. H. Schinz, M. V. Lo, D. C. Larrivee, and W. L. Pak. Freeze-fracture study of the Drosophila photoreceptor membrane: mutations affecting membrane particle density. Journal of Cell Biology, 93(3):961–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/93/3/961.

Scheer:1983:MRR

[SLR^{+}83] U. Scheer, G. Lanfranchi, K. M. Rose, W. W. Franke, and N. R. Ringertz. Migration of rat RNA polymerase I into chick erythrocyte nuclei undergoing reactivation in chick-rat heterokaryons. Journal of Cell Biology, 97(5):1641–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/5/
1641.

Suchard:1982:SCS

[SLRP82] S. J. Suchard, F. A. Lattanzio, R. W. Rubin, and B. C. Pressman. Stimulation of catecholamine secretion from cultured chromaffin cells by an ionophore-mediated rise in intracellular sodium. Journal of Cell Biology, 94(3):531–??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/94/3/
531.

Sommer:1981:RBG

[SLS81] I. Sommer, C. Lagenaur, and M. Schachner. Recognition of Bergmann glial and ependymal cells in the mouse nervous system by monoclonal antibody. Journal of Cell Biology, 90(2): 448–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/2/448.

Simionescu:1983:RMS

[SLS83] N. Simionescu, F. Lupu, and M. Simionescu. Rings of membrane sterols surround the openings of vesicles and fenestrae, in capillary endothelium. Journal of Cell Biology, 97(5): 1592–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/5/1592.

Salmon:1984:SMD

[SLS^{+}84] E. D. Salmon, R. J. Leslie, W. M. Saxton, M. L. Karow, and J. R. McIntosh. Spindle microtubule dynamics in sea urchin embryos: analysis using a fluorescein-labeled tubulin and measurements of fluorescence redistribution after laser photobleaching. Journal of Cell Biology, 99(6):2165–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/6/2165.

Saul:1984:DMD

[SLSK84] A. Saul, G. Lamont, W. H. Sawyer, and C. Kidson. Decreased membrane deformability in Melanesian ovalocytes from Papua New Guinea. Journal of Cell Biology, 98(4):1348–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/4/
1348.

Snell:1980:ADT

[SM80] W. J. Snell and W. S. Moore. Aggregation-dependent turnover of flagellar adhesion molecules in chlamydomonas gametes. Journal of Cell Biology, 84(1):203–??, January 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/84/1/203.

Solomon:1981:CSM

[SM81] F. Solomon and M. Magendantz. Cytochalasin separates microtubule disassembly from loss of asymmetric morphology. Journal of Cell Biology, 89(1):157–??, April 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/89/1/157.

Sherline:1982:EGFb

[SM82a] P. Sherline and R. Mascardo. Epidermal growth factor-induced centrosomal separation: mechanism and relationship to mitogenesis. Journal of Cell Biology, 95(1):316–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/95/1/316.

Sherline:1982:EGFa

[SM82b] P. Sherline and R. N. Mascardo. Epidermal growth factor induces rapid centrosomal separation in HeLa and 3T3 cells.

Journal of Cell Biology, 93(2):507–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/93/2/507.

Sinicropi:1983:CAC

[SM83a] D. V. Sinicropi and D. L. McIlwain. Changes in the amounts of cytoskeletal proteins within the perikarya and axons of regenerating frog motoneurons. Journal of Cell Biology, 96(1): 240–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/1/240.

Soll:1983:FRF

[SM83b] D. R. Soll and L. H. Mitchell. Filament ring formation in the dimorphic yeast Candida albicans. Journal of Cell Biology, 96 (2):486–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/2/486.

Sugino:1983:FIT

[SM83c] H. Sugino and F. Matsumura. Fragmin induces tension reduction of actomyosin threads in the presence of micromolar levels of Ca^{2+}. Journal of Cell Biology, 96(1):199–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/96/1/199.

Small:1981:OAL

[Sma81] J. V. Small. Organization of actin in the leading edge of cultured cells: influence of osmium tetroxide and dehydration on the ultrastructure of actin meshworks. Journal of Cell Biology, 91(3):695–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/91/3/695.

Sabban:1981:EMP

[SMAS81] E. Sabban, V. Marchesi, M. Adesnik, and D. D. Sabatini. Erythrocyte membrane protein band 3: its biosynthesis and incorporation into membranes. Journal of Cell Biology, 91(3): 637–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/91/3/637.

Studer:1981:SBM

[SMG81] D. Studer, H. Moor, and H. Gross. Single bacteriorhodopsin molecules revealed on both surfaces of freeze-dried and heavy metal-decorated purple membranes. Journal of Cell Biology, 90(1):153–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/1/153.

Schmidt:1984:CRP

[SMGB84] R. J. Schmidt, A. M. Myers, N. W. Gillham, and J. E. Boynton. Chloroplast ribosomal proteins of Chlamydomonas synthesized in the cytoplasm are made as precursors. Journal of Cell Biology, 98(6):2011–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/6/2011.

Salmon:1984:RRT

[SMH84] E. D. Salmon, M. McKeel, and T. Hays. Rapid rate of tubulin dissociation from microtubules in the mitotic spindle in vivo measured by blocking polymerization with colchicine. Journal of Cell Biology, 99(3):1066–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/3/1066.

Skriver:1980:UKT

[SMJ80] E. Skriver, A. B. Maunsbach, and P. L. Jørgensen. Ultrastructure of Na,K-transport vesicles reconstituted with purified renal Na,K–ATPase. Journal of Cell Biology, 86(3): 746–??, September 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/86/3/746.

Sage:1983:FCL

[SMJB83] H. Sage, R. Mecham, C. Johnson, and P. Bornstein. Fetal calf ligament fibroblasts in culture secrete a low molecular weight collagen with a unique resistance to proteolytic degradation. Journal of Cell Biology, 97(6):1933–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/6/

Steinman:1983:ERP

[SMMC83] R. M. Steinman, I. S. Mellman, W. A. Muller, and Z. A. Cohn. Endocytosis and the recycling of plasma membrane. Journal of Cell Biology, 96(1):1–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/1/1.

Sanger:1984:AMS

[SMS84a] J. W. Sanger, B. Mittal, and J. M. Sanger. Analysis of myofibrillar structure and assembly using fluorescently labeled contractile proteins. Journal of Cell Biology, 98(3):825–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/3/
825.

Sanger:1984:IFL

[SMS84b] J. W. Sanger, B. Mittal, and J. M. Sanger. Interaction of fluorescently-labeled contractile proteins with the cytoskeleton in cell models. Journal of Cell Biology, 99(3):918–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/3/918.

Schonbrunn:1980:EGF

[SMT80] A. Schonbrunn, Westendorf J. Krasnoff M., and A. Tashjian. Epidermal growth factor and thyrotropin-releasing hormone act similarly on a clonal pituitary cell strain. Modulation of hormone production and inhibition of cell proliferation. Journal of Cell Biology, 85(3):786–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/85/3/786.

Shepherd:1980:CCO

[SMTP80] J. H. Shepherd, E. R. Mulvihill, P. S. Thomas, and R. D. Palmiter. Commitment of chick oviduct tubular gland cells to produce ovalbumin mRNA during hormonal withdrawal and restimulation. Journal of Cell Biology, 87(1):142–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/87/1/

Stenn:1983:MMD

[SMTT83] K. S. Stenn, J. A. Madri, T. Tinghitella, and V. P. Terranova. Multiple mechanisms of dissociated epidermal cell spreading. Journal of Cell Biology, 96(1):63–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/1/63.

Specian:1980:MRM

[SN80] R. D. Specian and M. R. Neutra. Mechanism of rapid mucus secretion in goblet cells stimulated by acetylcholine. Journal of Cell Biology, 85(3):626–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/85/3/626.

Spray:1984:SBA

[SNdC^{+}84] D. C. Spray, J. Nerbonne, A. Campos de Carvalho, A. L. Harris, and M. V. Bennett. Substituted benzyl acetates: a new class of compounds that reduce gap junctional conductance by cytoplasmic acidification. Journal of Cell Biology, 99(1):174–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/1/174.

Schliwa:1984:TPI

[SNPE84] M. Schliwa, T. Nakamura, K. R. Porter, and U. Euteneuer. A tumor promoter induces rapid and coordinated reorganization of actin and vinculin in cultured cells. Journal of Cell Biology, 99(3):1045–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/3/1045.

Sung:1983:YMI

[SNS83] S. J. Sung, R. S. Nelson, and S. C. Silverstein. Yeast Mannans inhibit binding and phagocytosis of zymosan by mouse peritoneal macrophages. Journal of Cell Biology, 96 (1):160–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/1/160.

Sandvig:1980:DTE

[SO80] K. Sandvig and S. Olsnes. Diphtheria toxin entry into cells is facilitated by low pH. Journal of Cell Biology, 87(3):

828–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/87/3/828.

Stearns:1982:FVM

[SO82] M. E. Stearns and R. L. Ochs. A functional in vitro model for studies of intracellular motility in digitonin-permeabilized erythrophores. Journal of Cell Biology, 94(3):727–??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/94/3/
727.

Sobel:1984:MRS

[Sob84] J. S. Sobel. Myosin rings and spreading in mouse blastomeres. Journal of Cell Biology, 99(3):1145–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/3/
1145.

Solomon:1981:SCM

[Sol81] F. Solomon. Specification of cell morphology by endogenous determinants. Journal of Cell Biology, 90(3):547–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
90/3/547.

Sowers:1984:CEF

[Sow84] A. E. Sowers. Characterization of electric field-induced fusion in erythrocyte ghost membranes. Journal of Cell Biology, 99(6):1989–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/6/1989.

Sarris:1982:IDE

[SP82a] A. H. Sarris and G. E. Palade. Immunofluorescent detection of erythrocyte sialoglycoprotein antigens on murine erythroid cells. Journal of Cell Biology, 93(3):591–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/93/3/591.

Sarris:1982:IPC

[SP82b] A. H. Sarris and G. E. Palade. Isolation and partial characterization of the sialoglycoprotein fraction of murine erythrocyte ghosts. Journal of Cell Biology, 93(3):583–??, June 1982.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/93/3/583.

Severinsson:1984:BMI

[SP84a] L. Severinsson and P. A. Peterson. Beta 2-microglobulin induces intracellular transport of human class I transplantation antigen heavy chains in Xenopus laevis oocytes. Journal of Cell Biology, 99(1):226–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/1/226.

Sleight:1984:TFP

[SP84b] R. G. Sleight and R. E. Pagano. Transport of a fluorescent phosphatidylcholine analog from the plasma membrane to the Golgi apparatus. Journal of Cell Biology, 99(2):742–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/2/742.

Small:1984:CPMa

[SP84c] R. K. Small and K. H. Pfenninger. Components of the plasma membrane of growing axons. I. Size and distribution of intramembrane particles. Journal of Cell Biology, 98(4):1422–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
98/4/1422.

Smith:1984:SDD

[SPBB84] H. C. Smith, E. Puvion, L. A. Buchholtz, and R. Berezney. Spatial distribution of DNA loop attachment and replicational sites in the nuclear matrix. Journal of Cell Biology, 99(5): 1794–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/5/1794.

Snyderman:1984:CRM

[SPEL84] R. Snyderman, M. C. Pike, S. Edge, and B. Lane. A chemoattractant receptor on macrophages exists in two affinity states regulated by guanine nucleotides. Journal of Cell Biology, 98 (2):444–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/2/444.

Strehler:1980:BUA

[SPHE80] E. E. Strehler, G. Pelloni, C. W. Heizmann, and H. M. Eppenberger. Biochemical and ultrastructural aspects of Mr 165,000 M-protein in cross-striated chicken muscle. Journal of Cell Biology, 86(3):775–??, September 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/86/3/775.

Storrie:1984:EBP

[SPS^{+}84] B. Storrie, R. R. Pool, M. Sachdeva, K. M. Maurey, and C. Oliver. Evidence for both prelysosomal and lysosomal intermediates in endocytic pathways. Journal of Cell Biology, 98 (1):108–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/1/108.

Shriver:1981:OPS

[SR81a] K. Shriver and L. Rohrschneider. Organization of pp60src and selected cytoskeletal proteins within adhesion plaques and junctions of Rous sarcoma virus-transformed rat cells. Journal of Cell Biology, 89(3):525–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/89/3/525.

Sieber:1981:QAI

[SR81b] F. Sieber and S. Roseman. Quantitative analysis of intercellular adhesive specificity in freshly explanted and cultured cells. Journal of Cell Biology, 90(1):55–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/90/1/55.

Schindler:1982:CMM

[SR82a] M. Schindler and J. P. Rosenbusch. Chemical modification of matrix porin from Escherichia coli: probing the pore topology of a transmembrane protein. Journal of Cell Biology, 92 (3):742–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/3/742.

Schnapp:1982:CSR

[SR82b] B. J. Schnapp and T. S. Reese. Cytoplasmic structure in rapidfrozen axons. Journal of Cell Biology, 94(3):667–??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/94/3/
667.

Stoolman:1983:PRC

[SR83a] L. M. Stoolman and S. D. Rosen. Possible role for cell-surface carbohydrate-binding molecules in lymphocyte recirculation. Journal of Cell Biology, 96(3):722–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/3/722.

Stump:1983:XNC

[SR83b] R. F. Stump and K. R. Robinson. Xenopus neural crest cell migration in an applied electrical field. Journal of Cell Biology, 97(4):1226–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/4/1226.

Sampath:1984:IGE

[SR84a] T. K. Sampath and A. H. Reddi. Importance of geometry of the extracellular matrix in endochondral bone differentiation. Journal of Cell Biology, 98(6):2192–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/6/2192.

Suprenant:1984:PCO

[SR84b] K. A. Suprenant and L. I. Rebhun. Purification and characterization of oocyte cytoplasmic tubulin and meiotic spindle tubulin of the surf clam Spisula solidissima. Journal of Cell Biology, 98(1):253–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/1/253.

Seeley:1984:DIN

[SRCG84] P. J. Seeley, A. Rukenstein, J. L. Connolly, and L. A. Greene. Differential inhibition of nerve growth factor and epidermal growth factor effects on the PC12 pheochromocytoma line. Journal of Cell Biology, 98(2):417–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/2/417.

Schmidt:1983:SSC

[SRGB83] R. J. Schmidt, C. B. Richardson, N. W. Gillham, and J. E. Boynton. Sites of synthesis of chloroplast ribosomal proteins in Chlamydomonas. Journal of Cell Biology, 96(5):1451–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/5/
1451.

Salmon:1980:CLM

[SS80a] E. D. Salmon and R. R. Segall. Calcium-labile mitotic spindles isolated from sea urchin eggs (Lytechinus variegatus). Journal of Cell Biology, 86(2):355–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/86/2/355.

Sanger:1980:BPA

[SS80b] J. M. Sanger and J. W. Sanger. Banding and polarity of actin filaments in interphase and cleaving cells. Journal of Cell Biology, 86(2):568–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/86/2/568.

Spudich:1980:RCC

[SS80c] J. L. Spudich and R. Sager. Regulation of the Chlamydomonas cell cycle by light and dark. Journal of Cell Biology, 85 (1):136–??, April 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/85/1/136.

Spudich:1981:PPH

[SS81] E. N. Spudich and J. L. Spudich. Photosensitive phosphoproteins in Halobacteria: regulatory coupling of transmembrane proton flux and protein dephosphorylation. Journal of Cell Biology, 91(3):895–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/91/3/895.

Shine:1984:IMB

[SS84a] H. D. Shine and R. L. Sidman. Immunoreactive myelin basic proteins are not detected when shiverer mutant Schwann cells and fibroblasts are co-cultured with normal neurons. Journal of Cell Biology, 98(4):1291–??, April 1984. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/4/1291.

Stafstrom:1984:ALD

[SS84b] J. P. Stafstrom and L. A. Staehelin. Are annulate lamellae in the Drosophila embryo the result of overproduction of nuclear pore components? Journal of Cell Biology, 98(2): 699–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/2/699.

Swan:1984:RMB

[SS84c] J. A. Swan and F. Solomon. Reformation of the marginal band of avian erythrocytes in vitro using calf-brain tubulin: peripheral determinants of microtubule form. Journal of Cell Biology, 99(6):2108–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/6/2108.

Stommel:1980:MSC

[SSA80] E. W. Stommel, R. E. Stephens, and D. L. Alkon. Motile statocyst cilia transmit rather than directly transduce mechanical stimuli. Journal of Cell Biology, 87(3):652–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/87/3/652.

Schatten:1982:TIN

[SSBB82] G. Schatten, H. Schatten, T. H. Bestor, and R. Balczon. Taxol inhibits the nuclear movements during fertilization and induces asters in unfertilized sea urchin eggs. Journal of Cell Biology, 94(2):455–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/94/2/455.

Schlossman:1984:ERC

[SSBR84] D. M. Schlossman, S. L. Schmid, W. A. Braell, and J. E. Rothman. An enzyme that removes clathrin coats: purification of an uncoating ATPase. Journal of Cell Biology, 99(2):723–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/2/723.

Stanley:1980:DIC

[SSC80] P. Stanley, T. Sudo, and J. P. Carver. Differential involvement of cell surface sialic acid residues in wheat germ agglutinin binding to parental and wheat germ agglutinin-resistant Chinese hamster ovary cells. Journal of Cell Biology, 85(1):60–??, April 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
85/1/60.

Saito:1984:PMS

[SSCF84] A. Saito, S. Seiler, A. Chu, and S. Fleischer. Preparation and morphology of sarcoplasmic reticulum terminal cisternae from rabbit skeletal muscle. Journal of Cell Biology, 99(3): 875–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/3/875.

Steinberg:1980:ETD

[SSCP80] B. M. Steinberg, K. Smith, M. Colozzo, and R. Pollack. Establishment and transformation diminish the ability of fibroblasts to contract a native collagen gel. Journal of Cell Biology, 87 (1):304–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/87/1/304.

Schreiber:1984:IBM

[SSE84] A. B. Schreiber, J. Schlessinger, and M. Edidin. Interaction between major histocompatibility complex antigens and epidermal growth factor receptors on human cells. Journal of Cell Biology, 98(2):725–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/2/725.

Shimizu:1981:CCA

[SSF81] N. Shimizu, Y. Shimizu, and B. B. Fuller. Cell-cycle analysis of insulin binding and internalization on mouse melanoma cells. Journal of Cell Biology, 88(1):241–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/88/1/241.

Spiegel:1984:IFG

[SSF84] S. Spiegel, J. Schlessinger, and P. H. Fishman. Incorporation of fluorescent gangliosides into human fibroblasts: mobility, fate, and interaction with fibronectin. Journal of Cell Biology, 99 (2):699–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/2/699.

Santos-Sacchi:1980:IAR

[SSG80] J. Santos-Sacchi and M. Gordon. Induction of the acrosome reaction in guinea pig spermatozoa by cGMP analogues. Journal of Cell Biology, 85(3):798–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/85/3/798.

Schmid:1983:TTS

[SSG^{+}83a] E. Schmid, D. L. Schiller, C. Grund, J. Stadler, and W. W. Franke. Tissue type-specific expression of intermediate filament proteins in a cultured epithelial cell line from bovine mammary gland. Journal of Cell Biology, 96(1):37–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
96/1/37.

Schneider:1983:UAA

[SSG^{+}83b] W. J. Schneider, C. J. Slaughter, J. L. Goldstein, R. G. Anderson, J. D. Capra, and M. S. Brown. Use of antipeptide antibodies to demonstrate external orientation of the NH2-terminus of the low density lipoprotein receptor in the plasma membrane of fibroblasts. Journal of Cell Biology, 97(5):1635–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/5/
1635.

Salpeter:1982:BEC

[SSHP82] M. M. Salpeter, S. Spanton, K. Holley, and T. R. Podleski. Brain extract causes acetylcholine receptor redistribution which mimics some early events at developing neuromuscular junctions. Journal of Cell Biology, 93(2):417–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/93/2/417.

Sanger:1983:DSF

[SSJ83] J. W. Sanger, J. M. Sanger, and B. M. Jockusch. Differences in the stress fibers between fibroblasts and epithelial cells. Journal of Cell Biology, 96(4):961–??, April 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/4/961.

Sasaki-Sherrington:1984:ICA

[SSJS84] S. E. Sasaki-Sherrington, J. R. Jacobs, and J. K. Stevens. Intracellular control of axial shape in non-uniform neurites: a serial electron microscopic analysis of organelles and microtubules in AI and AII retinal amacrine neurites. Journal of Cell Biology, 98(4):1279–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/4/1279.

Salmon:1984:DCF

[SSL^{+}84a] E. D. Salmon, W. M. Saxton, R. J. Leslie, M. L. Karow, and J. R. McIntosh. Diffusion coefficient of fluorescein-labeled tubulin in the cytoplasm of embryonic cells of a sea urchin: video image analysis of fluorescence redistribution after photobleaching. Journal of Cell Biology, 99(6):2157–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/6/
2157.

Saxton:1984:TDC

[SSL^{+}84b] W. M. Saxton, D. L. Stemple, R. J. Leslie, E. D. Salmon, M. Zavortink, and J. R. McIntosh. Tubulin dynamics in cultured mammalian cells. Journal of Cell Biology, 99(6): 2175–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/6/2175.

Stenberg:1984:RAG

[SSLB84] P. E. Stenberg, M. A. Shuman, S. P. Levine, and D. F. Bainton. Redistribution of alpha-granules and their contents in thrombin-stimulated platelets. Journal of Cell Biology, 98(2): 748–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/2/748.

Stommel:1982:SLS

[SSMH82] E. W. Stommel, R. E. Stephens, H. R. Masure, and J. F. Head. Specific localization of scallop gill epithelial calmodulin in cilia. Journal of Cell Biology, 92(3):622–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/92/3/622.

Sandvig:1984:EDT

[SSO84] K. Sandvig, A. Sundan, and S. Olsnes. Evidence that diphtheria toxin and modeccin enter the cytosol from different vesicular compartments. Journal of Cell Biology, 98(3):963–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
98/3/963.

Simionescu:1981:DMLa

[SSP81] N. Simionescu, M. Simionescu, and G. E. Palade. Differentiated microdomains on the luminal surface of the capillary endothelium. I. Preferential distribution of anionic sites. Journal of Cell Biology, 90(3):605–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/90/3/605.

Simionescu:1982:DML

[SSP82a] M. Simionescu, N. Simionescu, and G. E. Palade. Differentiated microdomains on the luminal surface of capillary endothelium: distribution of lectin receptors. Journal of Cell Biology, 94(2):406–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/94/2/406.

Simionescu:1982:PDA

[SSP82b] M. Simionescu, N. Simionescu, and G. E. Palade. Preferential distribution of anionic sites on the basement membrane and the abluminal aspect of the endothelium in fenestrated capillaries. Journal of Cell Biology, 95(2):425–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/95/2/425.

Simpson:1984:MAP

[SSP84a] P. A. Simpson, J. A. Spudich, and P. Parham. Monoclonal antibodies prepared against Dictyostelium actin: characterization and interactions with actin. Journal of Cell Biology, 99(1):287–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/1/287.

Strichartz:1984:CPM

[SSP84b] G. R. Strichartz, R. K. Small, and K. H. Pfenninger. Components of the plasma membrane of growing axons. III. Saxitoxin binding to sodium channels. Journal of Cell Biology, 98 (4):1444–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/4/1444.

Subjeck:1983:ABM

[SSSJ83] J. R. Subjeck, T. Shyy, J. Shen, and R. J. Johnson. Association between the mammalian 110,000-dalton heat-shock protein and nucleoli. Journal of Cell Biology, 97(5):1389–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/5/
1389.

Simionescu:1981:DMLb

[SSSP81] M. Simionescu, N. Simionescu, J. E. Silbert, and G. E. Palade. Differentiated microdomains on the luminal surface of the capillary endothelium. II. Partial characterization of their anionic sites. Journal of Cell Biology, 90(3):614–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/90/3/614.

Saber:1983:CCA

[SSZ83] M. A. Saber, D. A. Shafritz, and M. A. Zern. Changes in collagen and albumin mRNA in liver tissue of mice infected with Schistosoma mansoni as determined by in situ hybridization. Journal of Cell Biology, 97(4):986–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/4/986.

Schwab:1983:MUR

[ST83] M. E. Schwab and H. Thoenen. Mechanism of uptake and retrograde axonal transport of noradrenaline in sympathetic neurons in culture: reserpine-resistant large dense-core vesicles as transport vehicles. Journal of Cell Biology, 96(6):1538–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
96/6/1538.

Stadler:1981:ASE

[Sta81] J. Stadler. Absence of sugars in electrophoretically purified cytochrome b5 demonstrated by combined gas chromatographymass spectrometry. Journal of Cell Biology, 89(3):615–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
89/3/615.

Stearns:1982:HVE

[Ste82a] M. E. Stearns. High voltage electron microscopy studies of axoplasmic transport in neurons: a possible regulatory role for divalent cations. Journal of Cell Biology, 92(3):765–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/92/3/
765.

Stephens:1982:EHM

[Ste82b] R. E. Stephens. Equimolar heterodimers in microtubules. Journal of Cell Biology, 94(2):263–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/94/2/263.

Steinberg:1983:SPM

[Ste83a] R. A. Steinberg. Sites of phosphorylation and mutation in regulatory subunit of cyclic AMP-dependent protein kinase from S49 mouse lymphoma cells: mapping to structural domains. Journal of Cell Biology, 97(4):1072–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/4/1072.

Stephens:1983:RCM

[Ste83b] R. E. Stephens. Reconstitution of ciliary membranes containing tubulin. Journal of Cell Biology, 96(1):68–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/1/
68.

Stearns:1984:CC

[Ste84a] M. E. Stearns. Cytomatrix in chromatophores. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/1/144s.

Steinberg:1984:CAD

[Ste84b] R. A. Steinberg. Cyclic AMP-dependent phosphorylation of the precursor to beta subunit of mitochondrial F1-ATPase: a physiological mistake? Journal of Cell Biology, 98(6):2174–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
98/6/2174.

Stossel:1984:CAS

[Sto84] T. P. Stossel. Contribution of actin to the structure of the cytoplasmic matrix. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/1/
15s.

Stoolman:1984:PRM

[STR84] L. M. Stoolman, T. S. Tenforde, and S. D. Rosen. Phosphomannosyl receptors may participate in the adhesive interaction between lymphocytes and high endothelial venules. Journal of Cell Biology, 99(4):1535–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/4/1535.

Seyedin:1983:VIC

[STRP83] S. M. Seyedin, A. Y. Thompson, D. M. Rosen, and K. A. Piez. In vitro induction of cartilage-specific macromolecules by a bone extract. Journal of Cell Biology, 97(6):1950–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
97/6/1950.

Surprenant:1982:CBE

[Sur82] A. Surprenant. Correlation between electrical activity and ACTH/beta-endorphin secretion in mouse pituitary tumor cells. Journal of Cell Biology, 95(2):559–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/95/2/559.

Schliwa:1981:SIC

[SvB81] M. Schliwa and J. van Blerkom. Structural interaction of cytoskeletal components. Journal of Cell Biology, 90(1):222–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
90/1/222.

Strous:1983:TTG

[SVW^{+}83] G. J. Strous, P. Van Kerkhof, R. Willemsen, H. J. Geuze, and E. G. Berger. Transport and topology of galactosyltransferase in endomembranes of HeLa cells. Journal of Cell Biology, 97(3):723–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/3/723.

Skene:1981:CAT

[SW81a] J. H. Skene and M. Willard. Changes in axonally transported proteins during axon regeneration in toad retinal ganglion cells. Journal of Cell Biology, 89(1):86–??, April 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/89/1/86.

Skene:1981:ATP

[SW81b] J. H. P. Skene and M. Willard. Axonally transported proteins associated with axon growth in rabbit central and peripheral nervous systems. Journal of Cell Biology, 89(1):96–??, April 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/89/1/
96.

Schlegel:1984:TRH

[SW84] W. Schlegel and C. B. Wollheim. Thyrotropin-releasing hormone increases cytosolic free Ca^{2+} in clonal pituitary cells (GH3 cells): direct evidence for the mobilization of cellular calcium. Journal of Cell Biology, 99(1):83–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/1/83.

Sato:1983:RPL

[SWA83] M. Sato, T. Z. Wong, and R. D. Allen. Rheological properties of living cytoplasm: endoplasm of Physarum plasmodium. Journal of Cell Biology, 97(4):1089–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/4/1089.

Stidwill:1984:BBC

[SWB84] R. P. Stidwill, T. Wysolmerski, and D. R. Burgess. The brush border cytoskeleton is not static: in vivo turnover of proteins. Journal of Cell Biology, 98(2):641–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/2/641.

Sealock:1984:ULM

[SWF84] R. Sealock, B. E. Wray, and S. C. Froehner. Ultrastructural localization of the Mr 43,000 protein and the acetylcholine receptor in Torpedo postsynaptic membranes using monoclonal antibodies. Journal of Cell Biology, 98(6):2239–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
98/6/2239.

Salzer:1980:SSCb

[SWGB80] J. L. Salzer, A. K. Williams, L. Glaser, and R. P. Bunge. Studies of Schwann cell proliferation. II. Characterization of the stimulation and specificity of the response to a neurite membrane fraction. Journal of Cell Biology, 84(3):753–??, March 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/84/3/
753.

Schliwa:1981:LOA

[SWP81] M. Schliwa, K. Weber, and K. R. Porter. Localization and organization of actin in melanophores. Journal of Cell Biology, 89(2):267–??, May 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/89/2/267.

Strous:1983:VSV

[SWvK^{+}83] G. J. Strous, R. Willemsen, P. van Kerkhof, J. W. Slot, H. J. Geuze, and H. F. Lodish. Vesicular stomatitis virus glycoprotein, albumin, and transferrin are transported to the cell surface via the same Golgi vesicles. Journal of Cell Biology, 97(6): 1815–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/6/1815.

Stanchfield:1980:PIV

[SY80] J. E. Stanchfield and J. D. Yager. Primary induction of vitellogenin synthesis in monolayer cultures of amphibian hepatocytes. Journal of Cell Biology, 84(2):468–??, February 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/84/2/468.

Stanley:1983:SEP

[SY83] J. R. Stanley and S. H. Yuspa. Specific epidermal protein markers are modulated during calcium-induced terminal differentiation. Journal of Cell Biology, 96(6):1809–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/6/
1809.

Sullivan:1980:CPR

[SZ80] S. J. Sullivan and S. H. Zigmond. Chemotactic peptide receptor modulation in polymorphonuclear leukocytes. Journal of Cell Biology, 85(3):703–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/85/3/703.

Siekevitz:1981:RPS

[SZ81] P. Siekevitz and P. C. Zamecnik. Ribosomes and protein synthesis. Journal of Cell Biology, 91(3):??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/91/3/53s.

Travis:1981:SMF

[TA81] J. L. Travis and R. D. Allen. Studies on the motility of the foraminifera. I. Ultrastructure of the reticulopodial network of Allogromia laticollaris (Arnold). Journal of Cell Biology, 90(1):211–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/1/211.

Tyler:1980:SCS

[TAB80] J. M. Tyler, J. M. Anderson, and D. Branton. Structural comparison of several actin-binding macromolecules. Journal of Cell Biology, 85(2):489–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/85/2/489.

Thompson:1984:HAM

[TAC84] W. C. Thompson, D. J. Asai, and D. H. Carney. Heterogeneity among microtubules of the cytoplasmic microtubule complex detected by a monoclonal antibody to alpha tubulin. Journal of Cell Biology, 98(3):1017–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/3/1017.

Tessler:1980:AGD

[TAGG80] A. Tessler, A. Autilio-Gambetti, and P. Gambetti. Axonal growth during regeneration: a quantitative autoradiographic study. Journal of Cell Biology, 87(1):197–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/87/1/197.

Tamm:1982:FEB

[Tam82] S. L. Tamm. Flagellated ectosymbiotic bacteria propel a eucaryotic cell. Journal of Cell Biology, 94(3):697–??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/94/3/
697.

Tilney:1983:ATS

[TBCM83] L. G. Tilney, E. M. Bonder, L. M. Coluccio, and M. S. Mooseker. Actin from Thyone sperm assembles on only one end of an actin filament: a behavior regulated by profilin. Journal of Cell Biology, 97(1):112–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/1/112.

Tilney:1981:AFE

[TBD81] L. G. Tilney, E. M. Bonder, and D. J. DeRosier. Actin filaments elongate from their membrane-associated ends. Journal of Cell Biology, 90(2):485–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/90/2/485.

Taylor:1980:CBAb

[TBR80] D. L. Taylor, J. R. Blinks, and G. Reynolds. Contractile basis of ameboid movement. VII. Aequorin luminescence during ameboid movement, endocytosis, and capping. Journal of Cell Biology, 86(2):599–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/86/2/599.

Tsanev:1982:UFR

[TD82] R. G. Tsanev and L. P. Djondjurov. Ultrastructure of free ribonucleoprotein complexes in spread mammalian nuclei. Journal of Cell Biology, 94(3):662–??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/94/3/662.

Theibert:1983:CAR

[TD83] A. Theibert and P. N. Devreotes. Cyclic 3’, 5’-AMP relay in Dictyostelium discoideum: adaptation is independent of activation of adenylate cyclase. Journal of Cell Biology, 97 (1):173–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/1/173.

Tanaka:1980:MIB

[TDM80a] Y. Tanaka, P. De Camilli, and J. Meldolesi. Membrane interactions between secretion granules and plasmalemma in three exocrine glands. Journal of Cell Biology, 84(2):438–??, February 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/84/2/
438.

Tilney:1980:OAF

[TDM80b] L. G. Tilney, D. J. Derosier, and M. J. Mulroy. The organization of actin filaments in the stereocilia of cochlear hair cells. Journal of Cell Biology, 86(1):244–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/86/1/244.

Tadvalkar:1983:VRA

[TdS83a] G. Tadvalkar and P. Pinto da Silva. In vitro, rapid assembly of gap junctions is induced by cytoskeleton disruptors. Journal of Cell Biology, 96(5):1279–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/5/1279.

Tokuyasu:1983:IMSb

[TDS83b] K. T. Tokuyasu, A. H. Dutton, and S. J. Singer. Immunoelectron microscopic studies of desmin (skeletin) localization and intermediate filament organization in chicken cardiac muscle. Journal of Cell Biology, 96(6):1736–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/6/1736.

Tokuyasu:1983:IMSa

[TDS83c] K. T. Tokuyasu, A. H. Dutton, and S. J. Singer. Immunoelectron microscopic studies of desmin (skeletin) localization and intermediate filament organization in chicken skeletal muscle. Journal of Cell Biology, 96(6):1727–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/6/1727.

Torrisi:1984:CIM

[TdS84] M. R. Torrisi and P. Pinto da Silva. Compartmentalization of intracellular membrane glycocomponents is revealed by fracture-label. Journal of Cell Biology, 98(1):29–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/1/
29.

Tolbert:1981:MPG

[TE81] N. E. Tolbert and E. Essner. Microbodies: peroxisomes and glyoxysomes. Journal of Cell Biology, 91(3):??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/91/3/
271s.

Tilney:1983:AFSb

[TEDS83] L. G. Tilney, E. H. Egelman, D. J. DeRosier, and J. C. Saunder. Actin filaments, stereocilia, and hair cells of the bird cochlea. II. Packing of actin filaments in the stereocilia and in the cuticular plate and what happens to the organization when the stereocilia are bent. Journal of Cell Biology, 96(3):822–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
96/3/822.

Takasaki:1984:VDP

[TEL84] S. Takasaki, F. Emling, and L. Leive. Variants deficient in phagocytosis of latex beads isolated from the murine macrophagelike cell line J774. Journal of Cell Biology, 98 (6):2198–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/6/2198.

Tennekoon:1981:CSP

[TFAP81] G. I. Tennekoon, J. Frangia, S. Aitchison, and D. L. Price. Cerebroside sulfotransferase: preparation of antibody and localization of antigen in kidney. Journal of Cell Biology, 91(2): 332–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/91/2/332.

Tarone:1984:CSI

[TFGC84] G. Tarone, R. Ferracini, G. Galetto, and P. Comoglio. A cell surface integral membrane glycoprotein of 85,000 mol wt (gp85) associated with Triton X-100-insoluble cell skeleton. Journal of Cell Biology, 99(2):512–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/2/512.

Traeger:1983:TFU

[TG83] L. Traeger and M. A. Goldstein. Thin filaments are not of uniform length in rat skeletal muscle. Journal of Cell Biology, 96 (1):100–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/1/100.

Tarone:1982:CSM

[TGPC82] G. Tarone, G. Galetto, M. Prat, and P. M. Comoglio. Cell surface molecules and fibronectin-mediated cell adhesion: effect of proteolytic digestion of membrane proteins. Journal of Cell Biology, 94(1):179–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/94/1/179.

Telzer:1981:DSM

[TH81] B. R. Telzer and L. T. Haimo. Decoration of spindle microtubules with Dynein: evidence for uniform polarity. Journal of Cell Biology, 89(2):373–??, May 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/89/2/373.

Tomasek:1984:ARM

[TH84] J. J. Tomasek and E. D. Hay. Analysis of the role of microfilaments and microtubules in acquisition of bipolarity and elongation of fibroblasts in hydrated collagen gels. Journal of Cell Biology, 99(2):536–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/2/536.

Tseng:1984:ESK

[THT^{+}84] S. C. Tseng, D. Hatchell, N. Tierney, A. J. Huang, and T. T. Sun. Expression of specific keratin markers by rabbit corneal, conjunctival, and esophageal epithelia during vitamin A deficiency. Journal of Cell Biology, 99(6):2279–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/6/
2279.

Tsukita:1980:MMO

[TI80] S. Tsukita and H. Ishikawa. The movement of membranous organelles in axons. Electron microscopic identification of anterogradely and retrogradely transported organelles. Journal of Cell Biology, 84(3):513–??, March 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/84/3/513.

Taniguchi:1982:SRM

[TI82a] M. Taniguchi and H. Ishikawa. In situ reconstitution of myosin filaments within the myosin-extracted myofibril in cultured skeletal muscle cells. Journal of Cell Biology, 92(2): 324–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/2/324.

Tilney:1982:ART

[TI82b] L. G. Tilney and S. Inoué. Acrosomal reaction of Thyone sperm. II. The kinetics and possible mechanism of acrosomal process elongation. Journal of Cell Biology, 93(3):820–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/93/3/

Tsukita:1981:INF

[TIK81] S. Tsukita, H. Ishikawa, and M. Kurokawa. Isolation of 10-nm filaments from astrocytes in the mouse optic nerve. Journal of Cell Biology, 88(1):245–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/88/1/245.

Trapp:1981:ILP

[TIS^{+}81] B. D. Trapp, Y. Itoyama, N. H. Sternberger, R. H. Quarles, and H. Webster. Immunocytochemical localization of P0 protein in Golgi complex membranes and myelin of developing rat Schwann cells. Journal of Cell Biology, 90(1):1–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/90/1/1.

Tilney:1980:AMF

[TJ80] L. G. Tilney and L. A. Jaffe. Actin, microvilli, and the fertilization cone of sea urchin eggs. Journal of Cell Biology, 87(3): 771–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/87/3/771.

Tranum-Jensen:1983:FFA

[TJB83] J. Tranum-Jensen and S. Bhakdi. Freeze-fracture analysis of the membrane lesion of human complement. Journal of Cell Biology, 97(3):618–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/97/3/618.

Travis:1983:SMF

[TKA83] J. L. Travis, J. F. Kenealy, and R. D. Allen. Studies on the motility of the foraminifera. II. The dynamic microtubular cytoskeleton of the reticulopodial network of Allogromia laticollaris. Journal of Cell Biology, 97(6):1668–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/6/
1668.

Tycko:1983:RAE

[TKM83] B. Tycko, C. H. Keith, and F. R. Maxfield. Rapid acidification of endocytic vesicles containing asialoglycoprotein in cells of a human hepatoma line. Journal of Cell Biology, 97(6):

1762–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/6/1762.

Tougard:1983:RER

[TLPTV83] C. Tougard, D. Louvard, R. Picart, and A. Tixier-Vidal. The rough endoplasmic reticulum and the Golgi apparatus visualized using specific antibodies in normal and tumoral prolactin cells in culture. Journal of Cell Biology, 96(5):1197–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/5/
1197.

Turner:1984:TIB

[TM84] P. F. Turner and R. L. Margolis. Taxol-induced bundling of brain-derived microtubules. Journal of Cell Biology, 99(3): 940–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/3/940.

Tooney:1983:SSE

[TMA^{+}83] N. M. Tooney, M. W. Mosesson, D. L. Amrani, J. F. Hainfeld, and J. S. Wall. Solution and surface effects on plasma fibronectin structure. Journal of Cell Biology, 97(6):1686– ??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/6/1686.

Tanasugarn:1984:MDI

[TMRT84] L. Tanasugarn, P. McNeil, G. T. Reynolds, and D. L. Taylor. Microspectrofluorometry by digital image processing: measurement of cytoplasmic pH. Journal of Cell Biology, 98(2): 717–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/2/717.

Tokuyasu:1984:DVD

[TMS84] K. T. Tokuyasu, P. A. Maher, and S. J. Singer. Distributions of vimentin and desmin in developing chick myotubes in vivo. I. Immunofluorescence study. Journal of Cell Biology, 98 (6):1961–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/6/1961.

Tanabe:1982:CTP

[TMW82] K. Tanabe, R. B. Mikkelsen, and D. F. Wallach. Calcium transport of Plasmodium chabaudi-infected erythrocytes. Journal of Cell Biology, 93(3):680–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/93/3/680.

Talian:1983:RPP

[TOG83] J. C. Talian, J. B. Olmsted, and R. D. Goldman. A rapid procedure for preparing fluorescein-labeled specific antibodies from whole antiserum: its use in analyzing cytoskeletal architecture. Journal of Cell Biology, 97(4):1277–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/4/
1277.

Tokuyasu:1983:VLO

[Tok83] K. T. Tokuyasu. Visualization of longitudinally-oriented intermediate filaments in frozen sections of chicken cardiac muscle by a new staining method. Journal of Cell Biology, 97 (2):562–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/2/562.

Takei:1984:RLS

[TOK^{+}84] F. Takei, F. Oyama, K. Kimura, A. Hyodo, S. Mizuno, and K. Shimura. Reduced level of secretion and absence of subunit combination for the fibroin synthesized by a mutant silkworm, Nd(2). Journal of Cell Biology, 99(6):2005–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/6/2005.

Takemura:1984:QIL

[TOT^{+}84] S. Takemura, K. Omori, K. Tanaka, K. Omori, S. Matsuura, and Y. Tashiro. Quantitative immunoferritin localization of [Na+,K+]ATPase on canine hepatocyte cell surface. Journal of Cell Biology, 99(4):1502–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/4/1502.

Tseng:1982:MAA

[TP82] P. C. Tseng and T. D. Pollard. Mechanism of action of Acanthamoeba profilin: demonstration of actin species specificity and regulation by micromolar concentrations of MgCl2. Journal of Cell Biology, 94(1):213–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/94/1/213.

Tippit:1980:CDT

[TPHL80] D. H. Tippit, J. D. Pickett-Heaps, and R. Leslie. Cell division in two large pennate diatoms Hantzschia and Nitzschia III. A new proposal for kinetochore function during prometaphase. Journal of Cell Biology, 86(2):402–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/86/2/402.

Tippit:1980:OSM

[TPPH80] D. H. Tippit, L. Pillus, and J. Pickett-Heaps. Organization of spindle microtubules in Ochromonas danica. Journal of Cell Biology, 87(3):531–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/87/3/531.

Tsien:1982:CHI

[TPR82] R. Y. Tsien, T. Pozzan, and T. J. Rink. Calcium homeostasis in intact lymphocytes: cytoplasmic free calcium monitored with a new, intracellularly trapped fluorescent indicator. Journal of Cell Biology, 94(2):325–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/94/2/325.

Trapp:1982:PMA

[TQ82] B. D. Trapp and R. H. Quarles. Presence of the myelinassociated glycoprotein correlates with alterations in the periodicity of peripheral myelin. Journal of Cell Biology, 92 (3):877–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/3/877.

Trapp:1984:MAG

[TQG84] B. D. Trapp, R. H. Quarles, and J. W. Griffin. Myelinassociated glycoprotein and myelinating Schwann cell-axon interaction in chronic B,B’-iminodipropionitrile neuropathy. Journal of Cell Biology, 98(4):1272–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/4/1272.

Trapp:1984:ISQ

[TQS84] B. D. Trapp, R. H. Quarles, and K. Suzuki. Immunocytochemical studies of quaking mice support a role for the myelinassociated glycoprotein in forming and maintaining the periaxonal space and periaxonal cytoplasmic collar of myelinating Schwann cells. Journal of Cell Biology, 99(2):594–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/2/
594.

Tseng:1984:PIF

[TRC^{+}84] P. C. Tseng, M. S. Runge, J. A. Cooper, R. C. Williams, and T. D. Pollard. Physical, immunochemical, and functional properties of Acanthamoeba profilin. Journal of Cell Biology, 98 (1):214–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/1/214.

Trinkaus-Randall:1984:RCC

[TRG84] V. Trinkaus-Randall and I. K. Gipson. Role of calcium and calmodulin in hemidesmosome formation in vitro. Journal of Cell Biology, 98(4):1565–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/4/1565.

Thorens:1982:ILV

[TRN^{+}82] B. Thorens, J. Roth, A. W. Norman, A. Perrelet, and L. Orci. Immunocytochemical localization of the vitamin D-dependent calcium binding protein in chick duodenum. Journal of Cell Biology, 94(1):115–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/94/1/115.

Taylor:1981:DAA

[TRSS81] D. L. Taylor, J. Reidler, J. A. Spudich, and L. Stryer. Detection of actin assembly by fluorescence energy transfer. Journal of Cell Biology, 89(2):362–??, May 1981. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/89/2/362.

Toyota:1981:DTC

[TS81] N. Toyota and Y. Shimada. Differentiation of troponin in cardiac and skeletal muscles in chicken embryos as studied by immunofluorescence microscopy. Journal of Cell Biology, 91(2): 497–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/91/2/497.

Thompson:1983:PGS

[TS83a] H. A. Thompson and B. S. Spooner. Proteoglycan and glycosaminoglycan synthesis in embryonic mouse salivary glands: effects of beta–D–xyloside, an inhibitor of branching morphogenesis. Journal of Cell Biology, 96(5):1443–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/5/
1443.

Tilney:1983:AFSa

[TS83b] L. G. Tilney and J. C. Saunders. Actin filaments, stereocilia, and hair cells of the bird cochlea. I. Length, number, width, and distribution of stereocilia of each hair cell are related to the position of the hair cell on the cochlea. Journal of Cell Biology, 96(3):807–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/3/807.

Tseng:1983:MCS

[TSGS83] S. C. Tseng, N. Savion, D. Gospodarowicz, and R. Stern. Modulation of collagen synthesis by a growth factor and by the extracellular matrix: comparison of cellular response to two different stimuli. Journal of Cell Biology, 97(3):803–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
97/3/803.

Tietze:1982:MSE

[TSS82] C. Tietze, P. Schlesinger, and P. Stahl. Mannose-specific endocytosis receptor of alveolar macrophages: demonstration of two functionally distinct intracellular pools of receptor and their roles in receptor recycling. Journal of Cell Biology, 92 (2):417–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/2/417.

Tamm:1981:CRR

[TT81] S. L. Tamm and S. Tamm. Ciliary reversal without rotation of axonemal structures in ctenophore comb plates. Journal of Cell Biology, 89(3):495–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/89/3/495.

Tamm:1983:DSS

[TT83] S. L. Tamm and S. Tamm. Distribution of sterol-specific complexes in a continually shearing region of a plasma membrane and at procaryotic-eucaryotic cell junctions. Journal of Cell Biology, 97(4):1098–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/97/4/1098.

Tamm:1984:APD

[TT84a] S. L. Tamm and S. Tamm. Alternate patterns of doublet microtubule sliding in ATP-disintegrated macrocilia of the ctenophore Beroë. Journal of Cell Biology, 99(4):1364–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/4/1364.

Tilney:1984:OHA

[TT84b] L. G. Tilney and M. S. Tilney. Observations on how actin filaments become organized in cells. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/1/76s.

Tamir:1982:SSP

[TTGA82] H. Tamir, T. C. Theoharides, M. D. Gershon, and P. W. Askenase. Serotonin storage pools in basophil leukemia and mast cells: characterization of two types of serotonin binding protein and radioautographic analysis of the intracellular distribution of [3H]serotonin. Journal of Cell Biology, 93(3):638–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/93/3/

Tsukita:1980:CNU

[TTI80] S. Tsukita, S. Tsukita, and H. Ishikawa. Cytoskeletal network underlying the human erythrocyte membrane. Thin-section electron microscopy. Journal of Cell Biology, 85(3):567–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
85/3/567.

Tsukita:1981:EMS

[TTI^{+}81] S. Tsukita, S. Tsukita, H. Ishikawa, S. Sato, and M. Nakao. Electron microscope study of reassociation of spectrin and actin with the human erythrocyte membrane. Journal of Cell Biology, 90(1):70–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/90/1/70.

Tsukita:1983:BSC

[TTI^{+}83] S. Tsukita, S. Tsukita, H. Ishikawa, M. Kurokawa, K. Morimoto, K. Sobue, and S. Kakiuchi. Binding sites of calmodulin and actin on the brain spectrin, calspectin. Journal of Cell Biology, 97(2):574–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/2/574.

Tsukita:1984:BPG

[TTI84] S. Tsukita, S. Tsukita, and H. Ishikawa. Bidirectional polymerization of G-actin on the human erythrocyte membrane. Journal of Cell Biology, 98(3):1102–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/3/1102.

Tsukita:1983:ADS

[TTUI83] S. Tsukita, S. Tsukita, J. Usukura, and H. Ishikawa. ATPdependent structural changes of the outer dynein arm in Tetrahymena cilia: a freeze-etch replica study. Journal of Cell Biology, 96(5):1480–??, May 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/5/1480.

Tucker:1984:SOM

[Tuc84] J. B. Tucker. Spatial organization of microtubule-organizing centers and microtubules. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/1/55s.

Tartakoff:1983:LBS

[TV83] A. M. Tartakoff and P. Vassalli. Lectin-binding sites as markers of Golgi subcompartments: proximal-to-distal maturation of oligosaccharides. Journal of Cell Biology, 97(4):1243–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
97/4/1243.

Tanaka:1983:AST

[TWG83] K. Tanaka, L. Waxman, and A. L. Goldberg. ATP serves two distinct roles in protein degradation in reticulocytes, one requiring and one independent of ubiquitin. Journal of Cell Biology, 96(6):1580–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/6/1580.

Taylor:1980:CBAa

[TWH80] D. L. Taylor, Y. L. Wang, and J. M. Heiple. Contractile basis of ameboid movement. VII. The distribution of fluorescently labeled actin in living amebas. Journal of Cell Biology, 86 (2):590–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/86/2/590.

Tank:1982:EMD

[TWW82] D. W. Tank, E. S. Wu, and W. W. Webb. Enhanced molecular diffusibility in muscle membrane blebs: release of lateral constraints. Journal of Cell Biology, 92(1):207–??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/92/1/207.

Tiwari:1984:CIC

[TWWG84] S. C. Tiwari, S. M. Wick, R. E. Williamson, and B. E. Gunning. Cytoskeleton and integration of cellular function in cells of higher plants. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/99/1/63s.

Tennekoon:1983:TCS

[TZW83] G. Tennekoon, M. Zaruba, and J. Wolinsky. Topography of cerebroside sulfotransferase in Golgi-enriched vesicles from rat brain. Journal of Cell Biology, 97(4):1107–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/4/
1107.

Uddin:1984:RVD

[UAL84] M. Uddin, G. G. Altmann, and C. P. Leblond. Radioautographic visualization of differences in the pattern of [3H]uridine and [3H]orotic acid incorporation into the RNA of migrating columnar cells in the rat small intestine. Journal of Cell Biology, 98(5):1619–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/5/1619.

Unwin:1983:CMC

[UE83] P. N. Unwin and P. D. Ennis. Calcium-mediated changes in gap junction structure: evidence from the low angle X-ray pattern. Journal of Cell Biology, 97(5):1459–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/5/
1459.

Unwin:1982:LPA

[UM82] P. N. Unwin and R. A. Milligan. A large particle associated with the perimeter of the nuclear pore complex. Journal of Cell Biology, 93(1):63–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/93/1/63.

Ungar:1984:UCE

[UPLH84] F. Ungar, I. Piscopo, J. Letizia, and E. Holtzman. Uptake of calcium by the endoplasmic reticulum of the frog photoreceptor. Journal of Cell Biology, 98(5):1645–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/5/
1645.

Vierling:1983:RSP

[VA83] E. Vierling and R. S. Alberte. Regulation of synthesis of the photosystem I reaction center. Journal of Cell Biology, 97(6):

1806–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/6/1806.

Vahey:1983:MWP

[Vah83] M. Vahey. A 72,000-mol-wt protein from tomato inhibits rabbit acto–S–1 ATPase activity. Journal of Cell Biology, 96 (6):1761–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/6/1761.

Vallee:1982:TDP

[Val82] R. B. Vallee. A taxol-dependent procedure for the isolation of microtubules and microtubule-associated proteins (MAPs). Journal of Cell Biology, 92(2):435–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/92/2/435.

VanHaastert:1983:SAD

[Van83] P. J. Van Haastert. Sensory adaptation of Dictyostelium discoideum cells to chemotactic signals. Journal of Cell Biology, 96(6):1559–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/6/1559.

Vaccaro:1981:SFA

[VB81] C. A. Vaccaro and J. S. Brody. Structural features of alveolar wall basement membrane in the adult rat lung. Journal of Cell Biology, 91(2):427–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/91/2/427.

Virtanen:1984:DCD

[VBPL84] I. Virtanen, R. A. Badley, R. Paasivuo, and V. P. Lehto. Distinct cytoskeletal domains revealed in sperm cells. Journal of Cell Biology, 99(3):1083–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/3/1083.

Vallee:1984:MAP

[VBT84] R. B. Vallee, G. S. Bloom, and W. E. Theurkauf. Microtubuleassociated proteins: subunits of the cytomatrix. Journal of Cell Biology, 99(1):??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/1/38s.

VanZant:1983:ECR

[VC83] G. Van Zant and B. D. Chen. Erythropoietin causes down regulation of colony-stimulating factor (CSF-1) receptors on peritoneal exudate macrophages of the mouse. Journal of Cell Biology, 97(6):1945–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/97/6/1945.

Vitale:1984:TAB

[VC84] A. Vitale and M. J. Chrispeels. Transient N-acetylglucosamine in the biosynthesis of phytohemagglutinin: attachment in the Golgi apparatus and removal in protein bodies. Journal of Cell Biology, 99(1):133–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/1/133.

vanDeurs:1980:MSN

[vDB80] B. van Deurs and O. Behnke. Membrane structure of nonactivated and activated human blood platelets as revealed by freeze-fracture: evidence for particle redistribution during platelet contraction. Journal of Cell Biology, 87(1):209–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
87/1/209.

Vallee:1981:PKB

[VDT81] R. B. Vallee, M. J. DiBartolomeis, and W. E. Theurkauf. A protein kinase bound to the projection portion of MAP 2 (microtubule-associated protein 2). Journal of Cell Biology, 90(3):568–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/90/3/568.

vandeWater:1981:PGL

[vdWSCH81] L. van de Water, S. Schroeder, E. B. Crenshaw, and R. O. Hynes. Phagocytosis of gelatin-latex particles by a murine macrophage line is dependent on fibronectin and heparin. Journal of Cell Biology, 90(1):32–??, July 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/90/1/32.

Virtanen:1980:SCS

[VEL80] I. Virtanen, P. Ekblom, and P. Laurila. Subcellular compartmentalization of saccharide moieties in cultured normal and malignant cells. Journal of Cell Biology, 85(2):429–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/85/2/
429.

vanEekelen:1981:HAN

[vEvV81] C. A. van Eekelen and W. J. van Venrooij. hnRNA and its attachment to a nuclear protein matrix. Journal of Cell Biology, 88(3):554–??, March 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/88/3/554.

Valkirs:1982:TRC

[VF82] G. E. Valkirs and G. Feher. Topography of reaction center subunits in the membrane of the photosynthetic bacterium, rhodopseudomonas sphaeroides. Journal of Cell Biology, 95 (1):179–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/95/1/179.

Volloch:1982:TDM

[VH82] V. Volloch and D. Housman. Terminal differentiation of murine erythroleukemia cells: physical stabilization of end- stage cells. Journal of Cell Biology, 93(2):390–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/93/2/390.

Villiger:1981:HAM

[VKE^{+}81] B. Villiger, D. G. Kelley, W. Engleman, C. Kuhn, and J. A. McDonald. Human alveolar macrophage fibronectin: synthesis, secretion, and ultrastructural localization during gelatincoated latex particle binding. Journal of Cell Biology, 90(3): 711–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/3/711.

Vandewalle:1982:CCP

[VKHH82] A. Vandewalle, B. Köpfer-Hobelsberger, and H. G. Heidrich. Cortical cell populations from rabbit kidney isolated by free-flow electrophoresis: characterization by measurement of hormone-sensitive adenylate cyclase. Journal of Cell Biology, 92(2):505–??, February 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/92/2/505.

Valtorta:1984:SLA

[VMMC84] F. Valtorta, L. Madeddu, J. Meldolesi, and B. Ceccarelli. Specific localization of the alpha-latrotoxin receptor in the nerve terminal plasma membrane. Journal of Cell Biology, 99(1):124–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/1/124.

VanNest:1980:PSS

[VMRR80] G. A. Van Nest, R. J. MacDonald, R. K. Raman, and W. J. Rutter. Proteins synthesized and secreted during rat pancreatic development. Journal of Cell Biology, 86(3):784– ??, September 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/86/3/784.

vanMeer:1983:EMI

[vMS83] G. van Meer and K. Simons. An efficient method for introducing defined lipids into the plasma membrane of mammalian cells. Journal of Cell Biology, 97(5):1365–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/5/
1365.

VanDeurs:1982:PMF

[VN82] B. Van Deurs and K. Nilausen. Pinocytosis in mouse Lfibroblasts: ultrastructural evidence for a direct membrane shuttle between the plasma membrane and the lysosomal compartment. Journal of Cell Biology, 94(2):279–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/94/2/279.

Vale:1984:PPM

[VPMF84] R. D. Vale, S. W. Peterson, N. V. Matiuck, and C. F. Fox. Purified plasma membranes inhibit polypeptide growth factorinduced DNA synthesis in subconfluent 3T3 cells. Journal of Cell Biology, 98(3):1129–??, March 1984. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/3/1129.

Vogel:1980:RSC

[VRR80] A. Vogel, R. Ross, and E. Raines. Role of serum components in density-dependent inhibition of growth of cells in culture. Platelet-derived growth factor is the major serum determinant of saturation density. Journal of Cell Biology, 85(2):377–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
85/2/377.

Vale:1982:ABP

[VS82] R. D. Vale and E. M. Shooter. Alteration of binding properties and cytoskeletal attachment of nerve growth factor receptors in PC12 cells by wheat germ agglutinin. Journal of Cell Biology, 94(3):710–??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/94/3/710.

Varnum:1984:ECS

[VS84] B. Varnum and D. R. Soll. Effects of cAMP on single cell motility in Dictyostelium. Journal of Cell Biology, 99(3): 1151–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/3/1151.

Velazquez:1983:MDH

[VSBL83] J. M. Velazquez, S. Sonoda, G. Bugaisky, and S. Lindquist. Is the major Drosophila heat shock protein present in cells that have not been heat shocked? Journal of Cell Biology, 96(1): 286–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/1/286.

Vicker:1984:CCD

[VSD84] M. G. Vicker, W. Schill, and K. Drescher. Chemoattraction and chemotaxis in Dictyostelium discoideum: myxamoeba cannot read spatial gradients of cyclic adenosine monophosphate. Journal of Cell Biology, 98(6):2204–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/6/

Via:1980:ESB

[VSLS80] D. P. Via, S. Sramek, G. Larriba, and S. Steiner. Effects of sodium butyrate on the membrane glycoconjugates of murine sarcoma virus-transformed rat cells. Journal of Cell Biology, 84(2):225–??, February 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/84/2/225.

Vasile:1983:VBE

[VSS83] E. Vasile, M. Simionescu, and N. Simionescu. Visualization of the binding, endocytosis, and transcytosis of low-density lipoprotein in the arterial endothelium in situ. Journal of Cell Biology, 96(6):1677–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/6/1677.

Vogel:1980:MPD

[VTSS80] G. Vogel, L. Thilo, H. Schwarz, and R. Steinhart. Mechanism of phagocytosis in Dictyostelium discoideum: phagocytosis is mediated by different recognition sites as disclosed by mutants with altered phagocytotic properties. Journal of Cell Biology, 86(2):456–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/86/2/456.

VanHaaster:1983:EAD

[VV83] P. J. M. Van Haaster and P. R. Van Der Heijden. Excitation, adaption, and deadaptation of the cAMP-mediated cGMP Response in Dictyostelium discoideum. Journal of Cell Biology, 96(2):347–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/2/347.

VanHaastert:1984:TKC

[VV84] P. J. Van Haastert and M. M. Van Lookeren Campagne. Transient kinetics of a cGMP-dependent cGMP-specific phosphodiesterase from Dictyostelium discoideum. Journal of Cell Biology, 98(2):709–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/2/709.

Voyta:1984:IIE

[VVBZ84] J. C. Voyta, D. P. Via, C. E. Butterfield, and B. R. Zetter. Identification and isolation of endothelial cells based on their increased uptake of acetylated-low density lipoprotein. Journal of Cell Biology, 99(6):2034–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/6/2034.

VanDeurs:1981:VTC

[VVM81] B. Van Deurs, F. Von Bülow, and M. Møller. Vesicular transport of cationized ferritin by the epithelium of the rat choroid plexus. Journal of Cell Biology, 89(1):131–??, April 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/89/1/131.

VanHaastert:1982:NKC

[VVP82] P. J. M. Van Haastert, H. Van Walsum, and F. J. Pasveer. Nonequilibrium kinetics of a cyclic GMP-binding protein in Dictyostelium discoideum. Journal of Cell Biology, 94(2): 271–??, August 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/94/2/271.

Vorobjev:1982:CCC

[VY82] I. A. Vorobjev and Chentsov YuS. Centrioles in the cell cycle. I. Epithelial cells. Journal of Cell Biology, 93(3):938–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/93/3/
938.

Wright:1983:RRM

[WA83] W. E. Wright and J. Aronoff. Regulation of rat myosin lightchain synthesis in heterokaryons between 5-bromodeoxyuridine-blocked rat myoblasts and differentiated chick myocytes. Journal of Cell Biology, 96(6):1571–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/6/1571.

Wenkert:1984:TAM

[WA84] D. Wenkert and C. D. Allis. Timing of the appearance of macronuclear-specific histone variant hv1 and gene expression in developing new macronuclei of Tetrahymena thermophila.

Journal of Cell Biology, 98(6):2107–??, June 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/6/2107.

Walz:1982:CSSa

[Wal82a] B. Walz. Ca^{2+}-sequestering smooth endoplasmic reticulum in an invertebrate photoreceptor. I. Intracellular topography as revealed by OsFeCN staining and in situ Ca accumulation. Journal of Cell Biology, 93(3):839–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/93/3/839.

Walz:1982:CSSb

[Wal82b] B. Walz. Ca^{2+}-sequestering smooth endoplasmic reticulum in an invertebrate photoreceptor. II. Its properties as revealed by microphotometric measurements. Journal of Cell Biology, 93 (3):849–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/93/3/849.

Walsh:1984:SAC

[Wal84] C. Walsh. Synthesis and assembly of the cytoskeleton of Naegleria gruberi flagellates. Journal of Cell Biology, 98(2): 449–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/2/449.

Ward:1981:SMW

[WAN81] S. Ward, Y. Argon, and G. A. Nelson. Sperm morphogenesis in wild-type and fertilization-defective mutants of Caenorhabditis elegans. Journal of Cell Biology, 91(1):26–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/91/1/
26.

Wang:1984:RAF

[Wan84] Y. L. Wang. Reorganization of actin filament bundles in living fibroblasts. Journal of Cell Biology, 99(4):1478–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/4/

Watt:1984:SMT

[Wat84] F. M. Watt. Selective migration of terminally differentiating cells from the basal layer of cultured human epidermis. Journal of Cell Biology, 98(1):16–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/98/1/16.

Walter:1981:TPAb

[WB81a] P. Walter and G. Blobel. Translocation of proteins across the endoplasmic reticulum. II. Signal recognition protein (SRP) mediates the selective binding to microsomal membranes of in-vitro-assembled polysomes synthesizing secretory protein. Journal of Cell Biology, 91(2):551–??, November 1981. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/91/2/551.

Walter:1981:TPAc

[WB81b] P. Walter and G. Blobel. Translocation of proteins across the endoplasmic reticulum III. Signal recognition protein (SRP) causes signal sequence-dependent and site-specific arrest of chain elongation that is released by microsomal membranes. Journal of Cell Biology, 91(2):557–??, November 1981. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/91/2/557.

Weidner:1982:MSIa

[WB82] E. Weidner and W. Byrd. The microsporidian spore invasion tube. II. Role of calcium in the activation of invasion tube discharge. Journal of Cell Biology, 93(3):970–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/93/3/970.

Walter:1983:SDS

[WB83a] P. Walter and G. Blobel. Subcellular distribution of signal recognition particle and 7SL–RNA determined with polypeptide-specific antibodies and complementary DNA probe. Journal of Cell Biology, 97(6):1693–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/6/

Wurster:1983:SSA

[WB83b] B. Wurster and U. Butz. A study on sensing and adaptation in Dictyostelium discoideum: guanosine 3’,5’-phosphate accumulation and light-scattering responses. Journal of Cell Biology, 96(6):1566–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/6/1566.

Walter:1984:ISF

[WB84] M. F. Walter and H. Biessmann. Intermediate-sized filaments in Drosophila tissue culture cells. Journal of Cell Biology, 99 (4):1468–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/4/1468.

Witt:1983:TSI

[WBG83] D. P. Witt, D. J. Brown, and J. A. Gordon. Transformationsensitive isoactin in passaged chick embryo fibroblasts transformed by Rous sarcoma virus. Journal of Cell Biology, 96 (6):1766–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/6/1766.

Walter:1980:PER

[WBPO80] R. J. Walter, R. D. Berlin, J. R. Pfeiffer, and J. M. Oliver. Polarization of endocytosis and receptor topography on cultured macrophages. Journal of Cell Biology, 86(1):199–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/86/1/199.

Wasmuth:1980:LCC

[WC80] J. J. Wasmuth and L. Y. Chu. Linkage in cultured Chinese hamster cells of two genes, emtB and leuS, involved in protein synthesis and isolation of cell lines with mutations in three linked genes. Journal of Cell Biology, 87(3):697–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/87/3/
697.

Werb:1983:OAS

[WC83] Z. Werb and J. R. Chin. Onset of apoprotein E secretion during differentiation of mouse bone marrow-derived mononuclear phagocytes. Journal of Cell Biology, 97(4):1113–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/4/
1113.

Weisenberg:1984:AIG

[WC84] R. C. Weisenberg and C. Cianci. ATP-induced gelation– contraction of microtubules assembled in vitro. Journal of Cell Biology, 99(4):1527–??, October 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/4/1527.

Wright:1983:ABB

[WCJ83] R. L. Wright, B. Chojnacki, and J. W. Jarvik. Abnormal basalbody number, location, and orientation in a striated fiberdefective mutant of Chlamydomonas reinhardtii. Journal of Cell Biology, 96(6):1697–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/6/1697.

Wang:1983:IFA

[WCY^{+}83] E. Wang, J. G. Cairncross, W. K. Yung, E. A. Garber, and R. K. Liem. An intermediate filament-associated protein, p50, recognized by monoclonal antibodies. Journal of Cell Biology, 97(5):1507–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/5/1507.

Whiteley:1984:TPP

[WCZG84] B. Whiteley, D. Cassel, Y. X. Zhuang, and L. Glaser. Tumor promoter phorbol 12-myristate 13-acetate inhibits mitogenstimulated Na+/H+ exchange in human epidermoid carcinoma A431 cells. Journal of Cell Biology, 99(3):1162– ??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/3/1162.

Walcott:1980:LTR

[WD80] B. Walcott and M. M. Dewey. Length-tension relation in Limulus striated muscle. Journal of Cell Biology, 87(1):204–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
87/1/204.

Wick:1983:IMT

[WD83] S. M. Wick and J. Duniec. Immunofluorescence microscopy of tubulin and microtubule arrays in plant cells. I. Preprophase band development and concomitant appearance of nuclear envelope-associated tubulin. Journal of Cell Biology, 97(1):235–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/1/235.

Wollman:1984:CBC

[WD84] F. A. Wollman and P. Delepelaire. Correlation between changes in light energy distribution and changes in thylakoid membrane polypeptide phosphorylation in Chlamydomonas reinhardtii. Journal of Cell Biology, 98(1):1–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/98/1/1.

Wallimann:1983:NSP

[WDE83] T. Wallimann, T. C. Doetschman, and H. M. Eppenberger. Novel staining pattern of skeletal muscle M-lines upon incubation with antibodies against MM-creatine kinase. Journal of Cell Biology, 96(6):1772–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/96/6/1772.

Wiley:1980:RDP

[WE80] C. A. Wiley and M. H. Ellisman. Rows of dimeric-particles within the axolemma and juxtaposed particles within glia, incorporated into a new model for the paranodal glial-axonal junction at the node of Ranvier. Journal of Cell Biology, 84 (2):261–??, February 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/84/2/261.

Weigel:1980:RHB

[Wei80] P. H. Weigel. Rat hepatocytes bind to synthetic galactoside surfaces via a patch of asialoglycoprotein receptors. Journal of Cell Biology, 87(3):855–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/87/3/855.

Weidner:1982:MSIb

[Wei82] E. Weidner. The microsporidian spore invasion tube. III. Tube extrusion and assembly. Journal of Cell Biology, 93(3):976–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
93/3/976.

Weinberg:1983:OMB

[Wei83] R. A. Weinberg. Oncogenes and the molecular biology of cancer. Journal of Cell Biology, 97(6):1661–??, December 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/6/
1661.

Weiss:1984:MAA

[WES84] R. A. Weiss, R. Eichner, and T. T. Sun. Monoclonal antibody analysis of keratin expression in epidermal diseases: a 48- and 56-kdalton keratin as molecular markers for hyperproliferative keratinocytes. Journal of Cell Biology, 98(4):1397–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/98/4/
1397.

Warren:1983:NSG

[WFGB83] G. Warren, C. Featherstone, G. Griffiths, and B. Burke. Newly synthesized G protein of vesicular stomatitis virus is not transported to the cell surface during mitosis. Journal of Cell Biology, 97(5):1623–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/5/1623.

Woodcock:1984:HOS

[WFR84] C. L. Woodcock, L. L. Frado, and J. B. Rattner. The higherorder structure of chromatin: evidence for a helical ribbon arrangement. Journal of Cell Biology, 99(1):42–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/99/1/42.

Watt:1981:ISC

[WG81] F. M. Watt and H. Green. Involucrin synthesis is correlated with cell size in human epidermal cultures. Journal of Cell Biology, 90(3):738–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/90/3/738.

Witt:1982:MIF

[WG82a] D. P. Witt and J. A. Gordon. Most iodinatable fibroblast surface proteins accompany the cytoplast membrane during cytochalasin B-mediated enucleation of chick embryo fibroblasts. Journal of Cell Biology, 94(3):557–??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/94/3/557.

Wojtyk:1982:CSC

[WG82b] R. I. Wojtyk and S. Goldstein. Clonal selection in cultured human fibroblasts: role of protein synthetic errors. Journal of Cell Biology, 95(3):704–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/95/3/704.

White:1983:FIE

[WGF83] G. E. White, M. A. Gimbrone, and K. Fujiwara. Factors influencing the expression of stress fibers in vascular endothelial cells in situ. Journal of Cell Biology, 97(2):416–??, August 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/2/
416.

Williams:1980:CPP

[WGG80] R. Williams, J. Gingrich, and A. Glazer. Cyanobacterial Phycobilisomes. Particles from synechocystis 6701 and two pigment mutants. Journal of Cell Biology, 85(3):558–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/85/3/
558.

Warshawsky:1980:DVD

[WGRB80] H. Warshawsky, D. Goltzman, M. F. Rouleau, and J. J. Bergeron. Direct in vivo demonstration by radioautography of specific binding sites for calcitonin in skeletal and renal tissues of the rat. Journal of Cell Biology, 85(3):682–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/85/3/682.

Wick:1980:LCC

[WH80] S. M. Wick and P. K. Hepler. Localization of Ca++-containing antimonate precipitates during mitosis. Journal of Cell Biology, 86(2):500–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/86/2/500.

Wall:1981:GSR

[WH81] D. A. Wall and A. L. Hubbard. Galactose-specific recognition system of mammalian liver: receptor distribution on the hepatocyte cell surface. Journal of Cell Biology, 90(3): 687–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/3/687.

Wood:1982:DPG

[WH82] J. S. Wood and L. H. Hartwell. A dependent pathway of gene functions leading to chromosome segregation in Saccharomyces cerevisiae. Journal of Cell Biology, 94(3):718–??, September 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/94/3/
718.

Wight:1983:PPA

[WH83] T. N. Wight and V. C. Hascall. Proteoglycans in primate arteries. III. Characterization of the proteoglycans synthesized by arterial smooth muscle cells in culture. Journal of Cell Biology, 96(1):167–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/1/167.

Whitson:1984:FBB

[WHD^{+}84] S. W. Whitson, W. Harrison, M. K. Dunlap, D. E. Bowers, L. W. Fisher, P. G. Robey, and J. D. Termine. Fetal bovine bone cells synthesize bone-specific matrix proteins. Journal of Cell Biology, 99(2):607–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/2/607.

Wolniak:1980:DMC

[WHJ80] S. M. Wolniak, P. K. Hepler, and W. T. Jackson. Detection of the membrane-calcium distribution during mitosis in Haemanthus endosperm with chlorotetracycline. Journal of Cell Biology, 87(1):23–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/87/1/23.

Wolniak:1983:ICM

[WHJ83] S. M. Wolniak, P. K. Hepler, and W. T. Jackson. Ionic changes in the mitotic apparatus at the metaphase/anaphase transition. Journal of Cell Biology, 96(3):598–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/96/3/598.

Woods:1984:RHS

[WHK^{+}84] A. Woods, M. Höök, L. Kjellén, C. G. Smith, and D. A. Rees. Relationship of heparan sulfate proteoglycans to the cytoskeleton and extracellular matrix of cultured fibroblasts. Journal of Cell Biology, 99(5):1743–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/5/1743.

Warrick:1980:DMC

[WHSK80] H. Warrick, N. Hsiung, T. B. Shows, and R. Kucherlapati. DNA-mediated cotransfer of unlinked mammalian cell markers into mouse L cells. Journal of Cell Biology, 86(1):341–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/86/1/
341.

Walter:1981:TPAa

[WIB81] P. Walter, I. Ibrahimi, and G. Blobel. Translocation of proteins across the endoplasmic reticulum. I. Signal recognition protein (SRP) binds to in-vitro-assembled polysomes synthesizing secretory protein. Journal of Cell Biology, 91(2):545–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
91/2/545.

Wieben:1981:RSL

[Wie81] E. D. Wieben. Regulation of the synthesis of lactate dehydrogenase–X during spermatogenesis in the mouse. Journal of Cell Biology, 88(3):492–??, March 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/88/3/492.

Williams:1984:SID

[Wil84] G. T. Williams. Specific inhibition of the differentiation of Trypanosoma cruzi. Journal of Cell Biology, 99(1):79–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/1/
79.

Wiche:1983:OIP

[WKA^{+}83] G. Wiche, R. Krepler, U. Artlieb, R. Pytela, and H. Denk. Occurrence and immunolocalization of plectin in tissues. Journal of Cell Biology, 97(3):887–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/97/3/887.

Wolkoff:1984:ISA

[WKAH84] A. W. Wolkoff, R. D. Klausner, G. Ashwell, and J. Harford. Intracellular segregation of asialoglycoproteins and their receptor: a prelysosomal event subsequent to dissociation of the ligand-receptor complex. Journal of Cell Biology, 98(2): 375–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/2/375.

White:1980:FSF

[WKH80] J. White, J. Kartenbeck, and A. Helenius. Fusion of Semliki forest virus with the plasma membrane can be induced by low pH. Journal of Cell Biology, 87(1):264–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/87/1/264.

Wilcox:1982:IPR

[WKW82] D. K. Wilcox, R. P. Kitson, and C. C. Widnell. Inhibition of pinocytosis in rat embryo fibroblasts treated with monensin. Journal of Cell Biology, 92(3):859–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/92/3/859.

Woodrum:1980:SBM

[WL80] D. T. Woodrum and R. W. Linck. Structural basis of motility in the microtubular axostyle: implications for cytoplasmic microtubule structure and function. Journal of Cell Biology, 87(2):404–??, November 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/87/2/404.

Wright:1984:CBR

[WLCS84] S. D. Wright, M. R. Licht, L. S. Craigmyle, and S. C. Silverstein. Communication between receptors for different ligands on a single cell: ligation of fibronectin receptors induces a reversible alteration in the function of complement receptors on cultured human monocytes. Journal of Cell Biology, 99(1):336–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/1/336.

Wilkinson:1984:CAH

[WLFD84] P. C. Wilkinson, J. M. Lackie, J. V. Forrester, and G. A. Dunn. Chemokinetic accumulation of human neutrophils on immune complex-coated substrata: analysis at a boundary. Journal of Cell Biology, 99(5):1761–??, November 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/99/5/1761.

Warner:1981:PDM

[WM81] F. D. Warner and D. R. Mitchell. Polarity of dyneinmicrotubule interactions in vitro: cross-bridging between parallel and antiparallel microtubules. Journal of Cell Biology, 89(1):35–??, April 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/89/1/35.

Wagner:1984:BWP

[WM84] D. D. Wagner and V. J. Marder. Biosynthesis of von Willebrand protein by human endothelial cells: processing steps and their intracellular localization. Journal of Cell Biology, 99(6): 2123–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/6/2123.

Woodcock-Mitchell:1982:IKP

[WMENS82] J. Woodcock-Mitchell, R. Eichner, W. G. Nelson, and T. T. Sun. Immunolocalization of keratin polypeptides in human epidermis using monoclonal antibodies. Journal of Cell Biology, 95(2):580–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/95/2/580.

Watt:1984:CIR

[WMG84] F. M. Watt, D. L. Mattey, and D. R. Garrod. Calcium-induced reorganization of desmosomal components in cultured human keratinocytes. Journal of Cell Biology, 99(6):2211–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/6/2211.

White:1981:CFS

[WMH81] J. White, K. Matlin, and A. Helenius. Cell fusion by Semliki Forest, influenza, and vesicular stomatitis viruses. Journal of Cell Biology, 89(3):674–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/89/3/674.

Wieben:1983:USN

[WMP83] E. D. Wieben, S. J. Madore, and T. Pederson. U1 small nuclear ribonucleoprotein studied by in vitro assembly. Journal of Cell Biology, 96(6):1751–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/6/1751.

Wang:1984:ISP

[WMP^{+}84a] E. Wang, J. Michl, L. M. Pfeffer, S. C. Silverstein, and I. Tamm. Interferon suppresses pinocytosis but stimulates phagocytosis in mouse peritoneal macrophages: related changes in cytoskeletal organization. Journal of Cell Biology, 98(4):1328–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/4/1328.

Wilson:1984:CBA

[WMP^{+}84b] L. Wilson, H. P. Miller, T. A. Pfeffer, K. F. Sullivan, and H. W. Detrich. Colchicine-binding activity distinguishes sea urchin egg and outer doublet tubulins. Journal of Cell Biology, 99(1):37–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/1/37.

Warn:1984:DFA

[WMW84] R. M. Warn, R. Magrath, and S. Webb. Distribution of Factin during cleavage of the Drosophila syncytial blastoderm. Journal of Cell Biology, 98(1):156–??, January 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/1/156.

Webb:1981:DMI

[WN81] D. J. Webb and R. Nuccitelli. Direct measurement of intracellular pH changes in Xenopus eggs at fertilization and cleavage. Journal of Cell Biology, 91(2):562–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/91/2/562.

Wu:1982:DCH

[WNB82] R. S. Wu, D. Nishioka, and W. M. Bonner. Differential conservation of histone 2A variants between mammals and sea urchins. Journal of Cell Biology, 93(2):426–??, May 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/93/2/426.

Wollman:1980:OPI

[WOBR80] F. A. Wollman, J. Olive, P. Bennoun, and M. Recouvreur. Organization of the photosystem II centers and their associated antennae in the thylakoid membranes: a comparative ultrastructural, biochemical, and biophysical study of Chlamydomonas wild type and mutants lacking in photosystem II reaction centers. Journal of Cell Biology, 87(3):728–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
87/3/728.

Wolfe:1980:PSS

[Wol80a] J. Wolfe. A possible skeletal substructure of the macronucleus of Tetrahymena. Journal of Cell Biology, 84(1):160–??, January 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
84/1/160.

Wolosewick:1980:APG

[Wol80b] J. J. Wolosewick. The application of polyethylene glycol (PEG) to electron microscopy. Journal of Cell Biology, 86 (2):675–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/86/2/675.

Warburton:1981:CMC

[WOM^{+}81] M. J. Warburton, E. J. Ormerod, P. Monaghan, S. Ferns, and P. S. Rudland. Characterization of a myoepithelial cell line derived from a neonatal rat mammary gland. Journal of Cell Biology, 91(3):827–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/91/3/827.

Wagner:1982:IWP

[WOM82] D. D. Wagner, J. B. Olmsted, and V. J. Marder. Immunolocalization of von Willebrand protein in Weibel–Palade bodies of human endothelial cells. Journal of Cell Biology, 95(1): 355–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/95/1/355.

Woodcock:1980:NAI

[Woo80] C. L. Woodcock. Nucleus-associated intermediate filaments from chicken erythrocytes. Journal of Cell Biology, 85(3): 881–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/85/3/881.

Wachsberger:1980:IBV

[WP80] P. R. Wachsberger and F. A. Pepe. Interaction between vertebrate skeletal and uterine muscle myosins and light meromyosins. Journal of Cell Biology, 85(1):33–??, April 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/85/1/33.

Westermark:1982:HIC

[WP82a] B. Westermark and K. R. Portor. Hormonally induced changes in the cytoskeleton of human thyroid cells in culture. Journal of Cell Biology, 94(1):42–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/94/1/42.

Willingham:1982:TEG

[WP82b] M. C. Willingham and I. H. Pastan. Transit of epidermal growth factor through coated pits of the Golgi system. Journal of Cell Biology, 94(1):207–??, July 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/94/1/207.

Wong:1983:MRN

[WPKB83] C. H. Wong, C. E. Pankhurst, A. Kondorosi, and W. J. Broughton. Morphology of root nodules and nodule-like structures formed by Rhizobium and Agrobacterium strains containing a Rhizobium meliloti megaplasmid. Journal of Cell Biology, 97(3):787–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/97/3/787.

Weiss:1981:AFD

[WR81] R. E. Weiss and A. H. Reddi. Appearance of fibronectin during the differentiation of cartilage, bone, and bone marrow. Journal of Cell Biology, 88(3):630–??, March 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/88/3/630.

Wang:1983:ICA

[WR83] Z. W. Wang and E. Rozengurt. Interplay of cyclic AMP and microtubules in modulating the initiation of DNA synthesis in 3T3 cells. Journal of Cell Biology, 96(6):1743–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/6/
1743.

Weinberg:1981:TSO

[WRH81] C. B. Weinberg, C. G. Reiness, and Z. W. Hall. Topographical segregation of old and new acetylcholine receptors at developing ectopic endplates in adult rat muscle. Journal of Cell Biology, 88(1):215–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/88/1/215.

Wright:1981:SRM

[Wri81] W. E. Wright. Synthesis of rat myosin light chains in heterokaryons formed between undifferentiated rat myoblasts and chick skeletal myocytes. Journal of Cell Biology, 91(1): 11–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/91/1/11.

Wright:1984:CDH

[Wri84a] W. E. Wright. Control of differentiation in heterokaryons and hybrids involving differentiation-defective myoblast variants. Journal of Cell Biology, 98(2):436–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/2/436.

Wright:1984:IMG

[Wri84b] W. E. Wright. Induction of muscle genes in neural cells. Journal of Cell Biology, 98(2):427–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/2/427.

Wu:1981:CSM

[WRK81] L. Wu, M. Rechsteiner, and L. Kuehl. Comparative studies on microinjected high-mobility-group chromosomal proteins, HMG1 and HMG2. Journal of Cell Biology, 91(2): 488–??, November 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/91/2/488.

Wang:1983:NTL

[WRM83] K. Wang and R. Ramirez-Mitchell. A network of transverse and longitudinal intermediate filaments is associated with sarcomeres of adult vertebrate skeletal muscle. Journal of Cell Biology, 96(2):562–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/96/2/562.

Wasserman:1981:CTR

[WS81a] W. J. Wasserman and L. D. Smith. Calmodulin triggers the resumption of meiosis in amphibian oocytes. Journal of Cell Biology, 89(3):389–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/89/3/389.

Willard:1981:ADN

[WS81b] M. Willard and C. Simon. Antibody decoration of neurofilaments. Journal of Cell Biology, 89(2):198–??, May 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/89/2/198.

Wadsworth:1983:MFT

[WS83] P. Wadsworth and R. D. Sloboda. Microinjection of fluorescent tubulin into dividing sea urchin cells. Journal of Cell Biology, 97(4):1249–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/4/1249.

Wang:1984:MWP

[WS84] L. L. Wang and J. A. Spudich. A 45,000-mol-wt protein from unfertilized sea urchin eggs severs actin filaments in a calciumdependent manner and increases the steady-state concentration of nonfilamentous actin. Journal of Cell Biology, 99(3): 844–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/99/3/844.

Willard:1980:AAT

[WSB^{+}80] M. Willard, C. Simon, C. Baitinger, J. Levine, and P. Skene. Association of an axonally transported polypeptide (H) with 100-A filaments. Use of immunoaffinity electron microscope grids. Journal of Cell Biology, 85(3):587–??, June 1980. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/85/3/587.

Wills:1984:MRS

[WSH84] J. W. Wills, R. V. Srinivas, and E. Hunter. Mutations of the Rous sarcoma virus env gene that affect the transport and subcellular location of the glycoprotein products. Journal of Cell Biology, 99(6):2011–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/6/2011.

Wojcieszyn:1983:SMP

[WSLSJ83] J. W. Wojcieszyn, R. A. Schlegel, K. Lumley-Sapanski, and K. A. Jacobson. Studies on the mechanism of polyethylene glycol-mediated cell fusion using fluorescent membrane and cytoplasmic probes. Journal of Cell Biology, 96(1):151–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
96/1/151.

Wolfe:1980:CCR

[WSM80] R. A. Wolfe, G. H. Sato, and D. B. McClure. Continuous culture of rat C6 glioma in serum-free medium. Journal of Cell Biology, 87(2):434–??, November 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/87/2/434.

Weatherbee:1982:MAP

[WSM^{+}82] J. A. Weatherbee, P. Sherline, R. N. Mascardo, J. G. Izant, R. B. Luftig, and R. R. Weihing. Microtubule-associated proteins of HeLa cells: heat stability of the 200,000 mol wt HeLa MAPs and detection of the presence of MAP-2 in HeLa cell extracts and cycled microtubules. Journal of Cell Biology, 92 (1):155–??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/92/1/155.

Wick:1981:IMO

[WSO^{+}81] S. M. Wick, R. W. Seagull, M. Osborn, K. Weber, and B. E. Gunning. Immunofluorescence microscopy of organized microtubule arrays in structurally stabilized meristematic plant cells. Journal of Cell Biology, 89(3):685–??, June 1981. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/89/3/685.

White:1983:ITD

[WTK83] E. White, E. M. Tolbert, and E. R. Katz. Identification of tubulin in Dictyostelium discoideum: characterization of some unique properties. Journal of Cell Biology, 97(4):1011–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
97/4/1011.

Wibo:1981:ASM

[WTSAC^{+}81] M. Wibo, D. Thinès-Sempoux, A. Amar-Costesec, H. Beaufay, and D. Godelaine. Analytical study of microsomes and isolated subcellular membranes from rat liver VIII. Subfractionation of preparations enriched with plasma membranes, outer mitochondrial membranes, or Golgi complex membranes. Journal of Cell Biology, 89(3):456–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/89/3/456.

Wolf:1984:DRM

[WV84] D. E. Wolf and J. K. Voglmayr. Diffusion and regionalization in membranes of maturing ram spermatozoa. Journal of Cell Biology, 98(5):1678–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/5/1678.

Wehland:1983:RMAb

[WW83] J. Wehland and M. C. Willingham. A rat monoclonal antibody reacting specifically with the tyrosylated form of alphatubulin. II. Effects on cell movement, organization of microtubules, and intermediate filaments, and arrangement of Golgi elements. Journal of Cell Biology, 97(5):1476–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/5/
1476.

Wan:1984:MAL

[WWCD84] Y. J. Wan, T. C. Wu, A. E. Chung, and I. Damjanov. Monoclonal antibodies to laminin reveal the heterogeneity of basement membranes in the developing and adult mouse tissues. Journal of Cell Biology, 98(3):971–??, March 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/3/971.

Wilcox:1983:RIP

[WWDYW83] D. K. Wilcox, P. A. Whitaker-Dowling, J. S. Youngner, and C. C. Widnell. Rapid inhibition of pinocytosis in baby hamster kidney (BHK-21) cells following infection with vesicular stomatitis virus. Journal of Cell Biology, 97(5):1444–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/97/5/
1444.

Wiser:1983:MAP

[WWES83] M. F. Wiser, P. A. Wood, J. W. Eaton, and J. R. Sheppard. Membrane-associated phosphoproteins in Plasmodium berghei-infected murine erythrocytes. Journal of Cell Biology, 97(1):196–??, July 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/97/1/196.

Wang:1983:SCD

[WWKW83] R. J. Wang, W. Wissinger, E. J. King, and G. Wang. Studies on cell division in mammalian cells. VII. A temperaturesensitive cell line abnormal in centriole separation and chromosome movement. Journal of Cell Biology, 96(1):301–??, January 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
96/1/301.

Wallace:1984:CPI

[WWPL84] P. J. Wallace, R. P. Wersto, C. H. Packman, and M. A. Lichtman. Chemotactic peptide-induced changes in neutrophil actin conformation. Journal of Cell Biology, 99(3):1060–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
99/3/1060.

Wehland:1983:RMAa

[WWS83] J. Wehland, M. C. Willingham, and I. V. Sandoval. A rat monoclonal antibody reacting specifically with the tyrosylated form of alpha-tubulin. I. Biochemical characterization, effects on microtubule polymerization in vitro, and microtubule polymerization and organization in vivo. Journal of Cell Biology, 97(5):1467–??, November 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/5/1467.

Wood:1980:ILC

[WWWC80] J. G. Wood, R. W. Wallace, J. N. Whitaker, and W. Y. Cheung. Immunocytochemical localization of calmodulin and a heat-labile calmodulin-binding protein (CaM–BP80) in basal ganglia of mouse brain. Journal of Cell Biology, 84(1): 66–??, January 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/84/1/66.

Woda:1981:MSP

[WYF81] B. A. Woda, J. Yguerabide, and J. D. Feldman. Mobility of surface proteins on normal rat macrophages and on a “macrophagelike” rat tumor. Journal of Cell Biology, 90(3): 705–??, September 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/90/3/705.

Wang:1984:UGL

[WYK^{+}84] E. Wang, H. L. Yin, J. G. Krueger, L. A. Caliguiri, and I. Tamm. Unphosphorylated gelsolin is localized in regions of cell-substratum contact or attachment in Rous sarcoma virus-transformed rat cells. Journal of Cell Biology, 98(2): 761–??, February 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/98/2/761.

Warner:1980:PMS

[WZ80] F. D. Warner and N. C. Zanetti. Properties of microtubule sliding disintegration in isolated Tetrahymena cilia. Journal of Cell Biology, 86(2):436–??, August 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/86/2/436.

Wolf:1983:RLD

[WZ83] D. E. Wolf and C. A. Ziomek. Regionalization and lateral diffusion of membrane proteins in unfertilized and fertilized mouse eggs. Journal of Cell Biology, 96(6):1786–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/6/
1786.

Yin:1981:IGC

[YAF81] H. L. Yin, J. H. Albrecht, and A. Fattoum. Identification of gelsolin, a Ca^{2+}-dependent regulatory protein of actin gel-sol transformation, and its intracellular distribution in a variety of cells and tissues. Journal of Cell Biology, 91(3):901–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
91/3/901.

Yao:1982:ESD

[Yao82] M. C. Yao. Elimination of specific DNA sequences from the somatic nucleus of the ciliate Tetrahymena. Journal of Cell Biology, 92(3):783–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/92/3/783.

Yao:1981:TRC

[YBG81] M. C. Yao, E. Blackburn, and J. Gall. Tandemly repeated C– C–C–C–A–A hexanucleotide of Tetrahymena rDNA is present elsewhere in the genome and may be related to the alteration of the somatic genome. Journal of Cell Biology, 90(2):515–??, August 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
90/2/515.

Yoshino:1984:DPR

[YDL^{+}84] J. E. Yoshino, M. P. Dinneen, B. L. Lewis, J. H. Meador-Woodruff, and G. H. Devries. Differential proliferative responses of cultured Schwann cells to axolemma- and myelinenriched fractions. I. Biochemical studies. Journal of Cell Biology, 99(6):2309–??, December 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/6/2309.

Yen:1981:ANG

[YF81] S. H. Yen and K. L. Fields. Antibodies to neurofilament, glial filament, and fibroblast intermediate filament proteins bind to different cell types of the nervous system. Journal of Cell Biology, 88(1):115–??, January 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/88/1/115.

Yumura:1983:FPI

[YF83] S. Yumura and Y. Fukui. Filopodelike projections induced with dimethyl sulfoxide and their relevance to cellular polarity in Dictyostelium. Journal of Cell Biology, 96(3):857–??, March 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/96/3/
857.

Yamashiro:1983:AEV

[YFM83] D. J. Yamashiro, S. R. Fluss, and F. R. Maxfield. Acidification of endocytic vesicles by an ATP-dependent proton pump. Journal of Cell Biology, 97(3):929–??, September 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/97/3/929.

Yahara:1982:CBE

[YHS^{+}82] I. Yahara, F. Harada, S. Sekita, K. Yoshihira, and S. Natori. Correlation between effects of 24 different cytochalasins on cellular structures and cellular events and those on actin in vitro. Journal of Cell Biology, 92(1):69–??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/92/1/69.

Yamada:1984:DNA

[YK84] K. M. Yamada and D. W. Kennedy. Dualistic nature of adhesive protein function: fibronectin and its biologically active peptide fragments can autoinhibit fibronectin function. Journal of Cell Biology, 99(1):29–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/1/29.

Yang:1980:AFP

[YKJF80] N. S. Yang, W. Kirkland, T. Jorgensen, and P. Furmanski. Absence of fibronectin and presence of plasminogen activator in both normal and malignant human mammary epithelial cells in culture. Journal of Cell Biology, 84(1):120–??, January 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/84/1/120.

Yumura:1984:LAM

[YMF84] S. Yumura, H. Mori, and Y. Fukui. Localization of actin and myosin for the study of ameboid movement in Dictyostelium using improved immunofluorescence. Journal of Cell Biology, 99(3):894–??, September 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/99/3/894.

Yamamoto:1982:MID

[YPR^{+}82] K. Yamamoto, J. D. Pardee, J. Reidler, L. Stryer, and J. A. Spudich. Mechanism of interaction of Dictyostelium severin with actin filaments. Journal of Cell Biology, 95(3):711–??, December 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
95/3/711.

Yamaguchi:1983:EAI

[YRS83] M. Yamaguchi, R. M. Robson, and M. H. Stromer. Evidence for actin involvement in cardiac Z-lines and Z-line analogues.

Journal of Cell Biology, 96(2):435–??, February 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/96/2/435.

Yoshioka:1984:CMC

[YS84] T. Yoshioka and A. P. Somlyo. Calcium and magnesium contents and volume of the terminal cisternae in caffeine-treated skeletal muscle. Journal of Cell Biology, 99(2):558–??, August 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/99/2/
558.

Yamamoto:1981:DPI

[YSJ81] M. Yamamoto, H. W. Steinbusch, and T. M. Jessell. Differentiated properties of identified serotonin neurons in dissociated cultures of embryonic rat brain stem. Journal of Cell Biology, 91(1):142–??, October 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/91/1/142.

Yarden:1982:NAE

[YSS82] Y. Yarden, A. B. Schreiber, and J. Schlessinger. A nonmitogenic analogue of epidermal growth factor induces early responses mediated by epidermal growth factor. Journal of Cell Biology, 92(3):687–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/92/3/687.

Yamin:1982:ARR

[YT82] M. A. Yamin and S. L. Tamm. ATP reactivation of the rotary axostyle in termite flagellates: effects of dynein ATPase inhibitors. Journal of Cell Biology, 95(2):589–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/95/2/589.

Zieske:1982:MCS

[ZB82] J. D. Zieske and I. A. Bernstein. Modification of cell surface glycoprotein: addition of fucosyl residues during epidermal differentiation. Journal of Cell Biology, 95(2):626–??, October 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/95/2/

Zavala:1983:LPU

[ZB83] M. E. Zavala and D. L. Brandon. Localization of a phytohormone using immunocytochemistry. Journal of Cell Biology, 97(4):1235–??, October 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/97/4/1235.

Zirkin:1980:IAP

[ZCH80] B. R. Zirkin, T. S. K. Chang, and J. Heaps. Involement of an acrosinlike proteinase in the sulfhydryl-induced degradation of rabbit sperm nuclear protamine. Journal of Cell Biology, 85 (1):116–??, April 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/85/1/116.

Zampighi:1980:GJS

[ZCR80] G. Zampighi, J. M. Corless, and J. D. Robertson. On gap junction structure. Journal of Cell Biology, 86(1):190–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/86/1/
190.

Zucker-Franklin:1981:EHP

[ZF81]  D. Zucker-Franklin.  Endocytosis by human platelets:

metabolic and freeze-fracture studies. Journal of Cell Biology, 91(3):706–??, December 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/91/3/706.

Zentgraf:1984:DSO

[ZF84] H. Zentgraf and W. W. Franke. Differences of supranucleosomal organization in different kinds of chromatin: cell typespecific globular subunits containing different numbers of nucleosomes. Journal of Cell Biology, 99(1):272–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/99/1/272.

Zucker-Franklin:1984:TAM

[ZFP84] D. Zucker-Franklin and S. Petursson. Thrombocytopoiesis– analysis by membrane tracer and freeze-fracture studies on fresh human and cultured mouse megakaryocytes. Journal of Cell Biology, 99(2):390–??, August 1984. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/99/2/390.

Zackroff:1984:ICK

[ZGJ^{+}84] R. V. Zackroff, A. E. Goldman, J. C. Jones, P. M. Steinert, and R. D. Goldman. Isolation and characterization of keratinlike proteins from cultured cells with fibroblastic morphology. Journal of Cell Biology, 98(4):1231–??, April 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/98/4/1231.

Zanini:1980:MOP

[ZGN^{+}80] A. Zanini, G. Giannattasio, G. Nussdorfer, R. K. Margolis, R. U. Margolis, and J. Meldolesi. Molecular organization of prolactin granules. II. Characterization of glycosaminoglycans and glycoproteins of the bovine prolactin matrix. Journal of Cell Biology, 86(1):260–??, July 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/86/1/260.

Zeitlin:1982:CSD

[ZH82] P. L. Zeitlin and A. L. Hubbard. Cell surface distribution and intracellular fate of asialoglycoproteins: a morphological and biochemical study of isolated rat hepatocytes and monolayer cultures. Journal of Cell Biology, 92(3):634–??, March 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/92/3/634.

Zieve:1980:IPC

[ZHM80] G. W. Zieve, S. R. Heidemann, and J. R. McIntosh. Isolation and partial characterization of a cage of filaments that surrounds the mammalian mitotic spindle. Journal of Cell Biology, 87(1):160–??, October 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/87/1/160.

Zannis-Hadjopoulos:1980:IDC

[ZHTH80] M. Zannis-Hadjopoulos, M. W. Taylor, and R. Hand. Inhibition of DNA chain elongation in a purine-auxotrophic mutant of Chinese hamster. Journal of Cell Biology, 85(3):777–??, June 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
85/3/777.

Zigmond:1981:CCP

[Zig81] S. H. Zigmond. Consequences of chemotactic peptide receptor modulation for leukocyte orientation. Journal of Cell Biology, 88(3):644–??, March 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/88/3/644.

Zaremba:1983:APC

[ZK83] S. Zaremba and J. H. Keen. Assembly polypeptides from coated vesicles mediate reassembly of unique clathrin coats. Journal of Cell Biology, 97(5):1339–??, November 1983. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/97/5/
1339.

Zampighi:1984:SOK

[ZKF84] G. Zampighi, J. Kyte, and W. Freytag. Structural organization of (Na^{+} + K^{+})-ATPase in purified membranes. Journal of Cell Biology, 98(5):1851–??, May 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/98/5/1851.

Zigmond:1981:CPE

[ZLK81] S. H. Zigmond, H. I. Levitsky, and B. J. Kreel. Cell polarity: an examination of its behavioral expression and its consequences for polymorphonuclear leukocyte chemotaxis. Journal of Cell Biology, 89(3):585–??, June 1981. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/89/3/585.

Zielinski:1980:STM

[ZP80] R. E. Zielinski and C. A. Price. Synthesis of thylakoid membrane proteins by chloroplasts isolated from spinach. Cytochrome b559 and P700-chlorophyll a-protein. Journal of Cell Biology, 85(2):435–??, May 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/85/2/435.

Zampighi:1980:CAB

[ZRW80] G. Zampighi, J. A. Reynolds, and R. M. Watt. Characterization of apolipoprotein B from human serum low density lipoprotein in n-dodecyl octaethyleneglycol monoether: an electron microscope study. Journal of Cell Biology, 87(3): 555–??, December 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/87/3/555.

Zaner:1982:SPV

[ZS82] K. S. Zaner and T. P. Stossel. Some perspectives on the viscosity of actin filaments. Journal of Cell Biology, 93(3):987–??, June 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
93/3/987.

Zanetti:1984:ICL

[ZS84] N. C. Zanetti and M. Solursh. Induction of chondrogenesis in limb mesenchymal cultures by disruption of the actin cytoskeleton. Journal of Cell Biology, 99(1):115–??, July 1984. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/99/1/115.

Ziomek:1980:RMP

[ZSE80] C. A. Ziomek, S. Schulman, and M. Edidin. Redistribution of membrane proteins in isolated mouse intestinal epithelial cells. Journal of Cell Biology, 86(3):849–??, September 1980. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/86/3/849.

Zigmond:1982:KAC

[ZSL82] S. H. Zigmond, S. J. Sullivan, and D. A. Lauffenburger. Kinetic analysis of chemotactic peptide receptor modulation. Journal of Cell Biology, 92(1):34–??, January 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/92/1/34.

Zimmermann:1983:FMC

[ZSM^{+}83] A. Zimmermann, J. C. Schaer, D. E. Muller, J. Schneider, N. M. Miodonski-Maculewicz, and R. Schindler. Formation of mast cell granules in cell cycle mutants of an undifferentiated mastocytoma line: evidence for two different states of reversible proliferative quiescence. Journal of Cell Biology, 96 (6):1756–??, June 1983. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/96/6/1756.

Zampighi:1982:SOI

[ZSR^{+}82] G. Zampighi, S. A. Simon, J. D. Robertson, T. J. McIntosh, and M. J. Costello. On the structural organization of isolated bovine lens fiber junctions. Journal of Cell Biology, 93 (1):175–??, April 1982. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/93/1/175.